A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2020–2024)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 January 2025 Version 1.44

Title word cross-reference

[AHPS22]. (−1, 1) [RR22b]. (F.IES) [Rab21]. (k, h) [CL24c]. (λ, ϵ) [HTOC20]. (m, n) [LLZ24b]. (n − k + 1) [Ozk21]. (x + y) ∗ (x − y) [Jea20]. + [WMF20]. 0 < p < 1 [LL24f]. 1 [CA23, DC20b, DFK20, LLGC24b]. 1 < r < 2 [RVS^{+}22]. 2 [BYE^{+}23, BM24a, CJ23, CSS24, DL21, DK24, DT24, FFS24, GPHHA23, HW21a, HL24a, LYZ24, Mia24, SESH23, WTL21, XLLA21, YZ24d, ZE23]. 2 × 2 [CMVV20]. 3 [CUN24, CMS20b, DOT23, DFK20, FGPP23, HLLW21, LL24c, Mia24, Qiu20, RDG23, SGD20, WDY20]. 4 [OOO22]. 6 [BEIR23]. 9 [FHM23]. ^{∗} [USTG23]. _{2}φ_{2} [KJGN20]. _{3}φ_{2} [KJGN20]. _{N} [BS23, PMVB22]. _{q} [dS24]. _{σ} [QWY24]. A [KZ23b]. α [HÇYK24, HSC21, SUW23]. αl_{1} − βl_{2} [DH24b]. AXB = C [WLZ22]. B [HQS20, Uça23]. C(Ω, E) [Pie20]. C^{∗}X^{∗}AXC + C^{∗}X^{∗}B + B^{∗}XC + D = 0 [YZLL22]. C^{0} [BK21b, GW23]. C^{1} [CCM22, Min20]. C^{2} [BEL23a, FKV21, JLLL23, YKY24]. C^{n,γ}(R_{+}) [KKA20]. D [HbX24]. δ [Cha24, LNP19]. E [HQS20]. ℓ^{p} [BR20b]. ℓ^{q} [BR20b]. ϵ [HTOC20]. F [SA24, ZZZL21]. G [Kuz22, YZ24c]. G^{1} [CCM22, WSYPD24]. G^{2} [AH23, KPS24b, Zag23]. G_{b} [PH24]. H [CZ23a, DL20c, HH21a, HVM22, LZLL20, LM21b]. H(div) [ZZW^{+}24a]. H^{−s} [YTA22]. H^{1} [MMF24, QWY24, RlLZZ21, XFW23, ZHW20]. H_{+} [Wu22]. H_{−} [DSCZ24]. H_{0}^{1} [KWYN23]. H_{1} [HS20]. H_{∞} [DSCZ24]. H ∈ (0, 12) [Liu23]. hp [CLHL20]. I(0, 1; t) [BMS22]. k [Asa24, DNS21, DJ24, GZ24, JB21, JZY22, JS22b, MKS21, New24, WTWZ20, Zhu22, dF20]. khp [GEG23]. L [KRT23]. L^{1} [ZW21]. L^{2} [KK22, LLGC24a, LV23]. L^{p} [CWW23b, CWW23a, GGHY24, LWW23a]. L^{p,q}(R × R^{d}) [JL21]. l_{0} [SCY23]. L_{0/1} [LHS^{+}24]. l_{1} [JGZ^{+}24, LZZ20, LL24d, LNN20, LNN20]. L_{2} [HG22, Xu20, ZY20, WXZ21, LNN20]. L_{2}(^{1}) [HJL20]. l_{2} − l_{1} [WYW^{+}24]. l_{∞} [JGW^{+}23]. l_{p} [GXC24, LL24f, ZZLY23]. M [ABA19, Ans21, CMZ22, FHM23, Guo24, LDC20, LWXL21, ZZZL21, Gol24, Ozk21]. A [KZ23b]. R [BHJB21]. R^{2} [FNH20]. δ [LN22]. H [TXZ24]. H(div) [KD23]. H(div) [WWH21b]. A, q [HBLZ21]. I [FS24b]. M [HBLZ21, Miy22]. SO(n) [HHL20]. m × n [LWG24]. μ [AHPS22, LXZ23]. N [NMB22, Asa24, CSLQ22, DNS21, DJ24, GZ24, KDOT24, MA21a, Ozk21, YCX20]. ωl_{q} [GXC24]. P [EFJK22, BA23, CDCVV22, CLM21, DAM23, LL24f, TR22, Uga24, WLZY21]. P_{1} [YZZ20a]. ϕ [CCZ23]. Ψ [JBSN22, SuR23]. q [DOP23, DU19, GJ19, MPR24, Zho24]. Q_{1} [HW20]. QR [JW24]. R [JNCY23, Mim20]. {R, s + 1, k} [CLST20]. {R, s + 1, k, ∗} [CLST20]. R_{II} [SS23b]. R_{n} [SZ20]. Rn [HHL20]. S [Kes23, LZ24b, PMVB22, CLM20, Zhu22]. t [HSRA24, JD22, MGL20, RU22]. τ [ZYZ22]. θ [LY23b, YPLM22]. U_{S}(λ) [NZ21]. u_{xx}u_{yy} = 1 [ZB22]. φ [AFJGJJ22]. W [KHAW21, PP21, PS24c, dS24]. w′′ = f(t, w, w′) [SSRK23]. x [dS24]. X^{p} = A + M^{T} (X♯B)M [LKM20]. Z [CZ23a, SN21b]. |x| [BCGAR22].

-adaptive [GEG23]. -alkane [CSLQ22]. -alkane/nitrogen [CSLQ22]. -bases [AHPS22]. -Bernstein [MPR24]. -Bézier [DOP23]. -body [NMB22]. -Buzano [KZ23b]. -Camassa [LXZ23]. -compartment [OOO22]. -conforming [HH21a, MMF24, ZZW^{+}24a]. -constrained [LL24d]. -D [CSS24]. -decomposition [CMZ22]. -derivative [ABA19]. -dimensional [Gol24, KDOT24, MA21a]. -divergence [FS24b]. -divergences [AFJGJJ22]. -eigenvalues [CZ23a, SN21b]. -elliptic [TXZ24, KD23]. -equation [HVM22]. -equilibria [HTOC20]. -Eulerian [DU19]. -field [WTL21]. -finite [HW20]. -fractional [SuR23]. -framework [YZ24c]. -function [Kuz22]. -functions [CCZ23]. -gap [HbX24]. -Hilfer [JBSN22]. -Laplace [GJ19, BA23]. -Laplacian [CDCVV22, DAM23]. -matrices [Wu22]. -matrix [DL20c, EFJK22, FHM23, LWXL21, ZZZL21]. -means [JZY22]. -methods [LY23b, KHAW21]. -metric [PH24]. -multigrid [TR22]. -norm [HS20, LL24f, RlLZZ21, ZW21]. -norms [LL24f]. -numerical [KZ23b]. -orthogonal [Xu20]. -out-of- [Asa24, DNS21, DJ24, GZ24, Ozk21, Zhu22]. -point [FHM23]. -polarized [CLM20, CLM21]. -policy [SA24]. -populations [JD22]. -potent [CLST20]. -primal [LWW23a]. -product [RU22]. -projection [KWYN23]. -ratio [Zho24]. -record [JB21, MKS21, WTWZ20]. -records [JS22b, New24]. -regularized [WXZ21]. -Riemann [CL24c]. -robust [HSC21]. -robustness [HTOC20]. -shock [Cha24, LNP19, LN22]. -space [AHPS22]. -splines [KRT23]. -split [BEIR23]. -stability [HTOC20, KK22, QWY24, ZY20]. -stable [SUW23]. -strictly [LZ24b]. -tensor [Guo24, LDC20, LM21b, Miy22]. -tensors [HBLZ21, LZLL20]. -th [Uga24, WLZY21]. -Toeplitz [dF20]. -type [Kes23]. -uncontrollability [Mim20]. -value [MGL20]. -Wright [Ans21].

1 [Ano20a, Ano21a, Ano21b, Ano21c,

Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, QWY24, Tak20]. 1-median [TBA22]. 15 [Ano20b, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, APW^{+}21]. 16 [AHK^{+}23]. 19 [OOO22, SMR23, YBAE23].

2 [HCCSCGAM24, ÖHS23]. 2020 [Ano20a, Ano20b]. 2021 [Ano21l, Ano21a, Ano21g, Ano21b, Ano21h, Ano21-31, Ano21c, Ano21i, Ano21-33, Ano21-32, Ano21d, Ano21j, Ano21e, Ano21-34, Ano21f, Ano21k, Ano21-35]. 2022 [Ano22m, Ano22a, Ano22g, Ano22b, Ano22h, Ano22-32, Ano22c, Ano22i, Ano22-34, Ano22-33, Ano22d, Ano22j, Ano22k, Ano22e, Ano22-35, Ano22f, Ano22-36]. 2023 [Ano23m, Ano23a, Ano23g, Ano23b, Ano23h, Ano23-32, Ano23c, Ano23i, Ano23-34, Ano23-33, Ano23d, Ano23j, Ano23k, Ano23e, Ano23-35, Ano23f, Ano23l, Ano23-36]. 2024 [Ano24k, Ano24a, Ano24f, Ano24b, Ano24g, Ano24-29, Ano24h, Ano24-31, Ano24-30, Ano24c, Ano24i, Ano24d, Ano24-32, Ano24e, Ano24j, Ano24-33]. 2D [BHRW22]. 2D/3D [BHRW22]. 2nd [DSCF24].

319 [MHY21]. 322 [YHC20]. 343 [GMSR21]. 354 [GAA21]. 358 [BKMO20]. 366 [FAVM20]. 369 [LLG^{+}22]. 371 [AA21c]. 372 [AA20a]. 373 [BPR21]. 3D [BHRW22, GÇSA23]. 3rd [BSZ21].

407 [AGMZ22a]. 420 [KZ24]. 421 [MR25]. 445 [ZZW^{+}24b].

512 [CKT21].

63 [DS21].

8T [PP22]. 8T-LE [PP22].

a-posteriori [BPR20a, BPR21]. Abate [HMT23]. ABC [AGNGG^{+}23, PA24a]. ABC/GBRT [AGNGG^{+}23]. ABC/MLP [AGNGG^{+}23]. Abel [CST19, Dju24, HAD24, WLZ23]. ability [LFZ24]. ablation [AGH20]. absolute [HL22c, MER20, WCLW20, ZWL21, ZCSS24]. absorbing [Kim22, PJA24]. absorption [AL20]. academia [MRG^{+}21]. Accelerated [WLC23, BSNL19, DV21a, ER20a, ER21, GHTC21, GH20, HH24c, JH24, JWSG22, LM24, LZ21, LZZZ21, LZG24, MSW23, NP23, PN21a, Ros23, WLCH23, WY24b, XHG22, XHLD21, ZZZL21]. Accelerating [MT20b, Saâ20, ZL23b]. Acceleration [AS21a, CV24, DHLT24, GRVZ23, LZH21, PMVB22, YTA22, YYZB24, ZGZ23, ZVS21]. acceptability [FPGSR21]. according [VZ21]. accounts [CSZP20]. accretion [BBSG23]. AccSumK [LGG^{+}22]. Accuracy [KKO23, CZY22, DV21b, FHM23, Jea20, JYY21, Jia20, KJO23, LL20b, MH21, MGD22, MZY23, SBKR24, XLKL23, Yan22, YKY24, ZHW20]. Accurate [DOP21, LWXL21, Nov24, SDLZ23, ADDG22, All24, BBTZ20, BP22a, Cal23, DDNZ21, FdOPS21, Kop20, KAA22, LLZY22, LJ23a, MPR24, MO20, SMDVA22, SDC20, WWW^{+}24, XWXZ22, YVXX24, YC23, YTD23, ZHWD20, ZCY^{+}20, ZOW^{+}21]. acoustic [EFTW21, LMV24, LRQV24, RCN21, RL24, vtW24, WDY20, XLY22, ZAH^{+}24]. across [WHC^{+}20]. action [CCZ23]. activation [LWX24]. active [LLL24b]. activities [LPLP21]. actuarial [Pou21]. actuation [JKJ^{+}23]. adaptation [NMST21, WW24b]. adapted [TDD23]. Adaptive [ASML23, ANR20, BK20, BK21b, GJ21, JMA21, Maj23a, PM20a, PSWZ21, WLHJ20, WYH22, XWL24, Ais23, AAM20, ACPV24, BJ22, CGP22, CGS20, CD23, CDD^{+}24, CLX21, CZ23a, CHY20, CJL^{+}24b, CJLL21, DV21a, ER20b, EL22, FS24a, FDDM24, GS21a, GEG23, HYW20, HSM20, HPW22, HWLD23, HCCW20, KS22b, KAD22b, KZM21, LZD21, LY22a, LLLL20, MM20a, MM20b, NNL^{+}24, OBB23, RS23, RS22b, SGY22, SWW23, SK23a, SSRK23, TK24a, TH24b, WHY^{+}21, XH20a, XXH^{+}22, YCG^{+}23, YTT23, ZLZL23, ZSHD24, ZS20b]. Adaptive-multilevel [PSWZ21]. Adaptivity [BFVY23, AESW22, KLM22, PR24, TGS21]. added [PSZ22]. adding [KLS21]. addition [vdBS21]. Additional [LWT24, WX22]. Additive [YS24b, AUAA20, CEKN20, DCP22, DJ22b, MLY22, SLZD20, VS21, YLXY23, ZGJ21, ZVS21, ZGZW23]. addressing [KFK^{+}23]. Adesi [BO22]. ADI [GZ23, HL24a]. adjoint [HH24b, LS22a, MM21]. adjoint-based [HH24b]. Adjusted [EWCQ23, WCLW20]. Adjustment [CMdMRFG22, JYWZ21, WRZW21]. administration [KS22b]. administrative [LLW21]. admissible [BCKSV24]. ADMM [GHG24, HLLZ24, JHW21, JW23, MWLY24, QXZF23, SZZ21, WSLW23, WYW^{+}24, ZSCH21]. ADMM-based [HLLZ24, SZZ21, WSLW23]. ADMM-type [GHG24]. adopting [HC20a]. Adroit [YSUY24]. adsorption [GIV20]. Advanced [ZDF^{+}20, BBBU22, CMW24, MHW23]. advances [BBGI24, FP21, FMRS24, HMRZ23]. Advancing [KA24]. advection [AAM20, CDL22, CS23, LG23, MD20, PM20b, PLS22, RFGGM20, TH24a, WZZ21a, ZWY22, ZEW23]. AEFEM [XZL20]. aerodynamic [VS23]. aerosol [CLL20b]. aerothermodynamics [TFR^{+}20]. aesthetic [CFJVG21]. affecting [Gök23]. Affine [AQ20, AM22a, CJL^{+}24b, JWS23, XCY24]. affine-jump-diffusion [XCY24]. Age [OKE24, EP20, OEH23, WFZ22]. age-space [WFZ22]. ageing [BCLS24]. agglomerated [FFT20]. agglomeration [DBSS^{+}21]. aggregate [SHZ20]. aggregated [JHW21]. aggregation [BZ20, KK23, dHSR22]. agreement [LCZ19]. agricultural [HKU22]. AI [DDT23, TDD23]. AIDS [Sal21a, WFZ22]. aircraft [MBG21b]. airflow [SH21]. airfoil [HF21]. Airy [GST20, SKE20]. Airy-type [GST20, SKE20]. Ait [EM21a]. Ait-Sahalia-type [EM21a]. al [MGG^{+}24]. Albany [CWT23, LPH23]. ALE [GLYS21]. aleatory [HH23]. algal [GNGGL^{+}19]. algebra [CCM22, MO20]. Algebraic [APW^{+}21, RVX20, AS21a, AAMR24, AHK^{+}23, BSV20, BLV20, CT20, DC23, DL20c, HBGS21, HM21a, HM21b, HCH20, LWXL21, PNS21, SW20, SB24b, WSYPD24, Wen21b, WY21b, YQZS22, YCXW24, ZFwC20, ZKT20, ZTS21]. algebraic-trigonometric [WY21b]. Algorithm [Li24a, RV20, ABBB23, Ais20, AD21, ASA^{+}21, AF23, AORS22, BTC21, BA20, BHRW22, BYE20, BSMM20, BBBK24, BFF22, BM21, CD23, CZ23a, CKY23, CL24b, CLPZ21, CUUdB24, DKM23a, DMZ23, Dad20, DMP23, DS22a, DS22b, DOP23, DXW21, DBR24, DG21, DRD^{+}23, DHLT24, DV21a, DDKD19, EFJK22, FS20, FS24b, GR22, GXL22, GGO21, GWHZ23, GDZN23, GW24, HW21a, HYY21, HW22, HWCJ21, HZC^{+}23, HBLZ21, HLLZ24, HO22, HB22, HL24b, HWH^{+}21, Hu22, HLR22, HWLD23, Hur22, Ibr20, ISD23, JLY^{+}21, JMLX23, JQ22, JW23, JYYL23, JH24, JWS23, JWSG22, KMZ21, KLZ^{+}24, KLMP21, KJO23, KDLY22, KAM^{+}23, KL23b, La 22, LGG^{+}22, LZD21, LWLW23, LWG24, LZLL20, LLX^{+}22b, LLL22a, LCGW24, LH23, LZG24, LN21, MM20b, MVB20, MGL20, MJ20, MSS20, Mim20, MP21, NMF^{+}22, PS24a, PHDD22, PC24, PSWZ21, PDBdS24, PN21a, PSS24, QGW21, QHPL21]. algorithm [QXZF23, RDHA20, RYH24, Ros23, San21, Sha23, SVK20, SZZ21, SCY23, SWW23, SDW24a, SL24a, SKZ23, SD24, TBA22, TLY23, TLD^{+}23, TDAT23, TY23, TH24b, VPHE23, WEEA20, WXZ21, WX21, WDS22, WT22, WSLW23, WPX23, WY24b, WDPY24, WYW^{+}24, WAH22, WLDZ24, WHY^{+}21, Wen21a, WLC23, WZL^{+}23, Wör22, WZC22b, WSL^{+}23, XHLD21, XLZ20a, XH20b, YKAA21, Yam21a, YLK24, YL21b, YYBL24, hYbJzJ21, YYT24, YYZB24, YWL24, ZC21, ZSH20, Zha20a, ZZZL21, Zha21a, ZSCH21, ZCSL21, ZL22, ZLZL23, ZZLY23, ZSHD24, ZJBY23, ZCCK24, ZLL^{+}20b, ZWX24, dOdO23, FPCV^{+}22]. Algorithms [BSZ21, TSMZ20, AVLMRMGG23, BKMA24, BV22, BBSG23, BP22a, BM24a, CGCMRP24, CWY21, Cho24b, DBSS^{+}21, DHF23, DH24a, DXM^{+}20, DDT23, DS21, ERB^{+}22, FGM21, HSM20, HTD21, HC21, JGW^{+}23, JGZ^{+}24, JHYZ24, JLLL23, JNCY23, Kan20, KEKT21, KRT23, LWD20, LLLD20, MHW23, MPM23, MDLY23, MBKK24, NL23b, PK24, PRW^{+}22, PL20, PKMM24, RTT24, RH23, Rou24, SH23, SNMK21, SDL22, SW21a, SZZM23, TDD23, TV22, TDL^{+}22, TN24, UPPZ24, WLBG22, WLCH23, WLZY21, WSX21, Xin21, XA21, YL22, YZ24a, YJJW23, ZuIKA22, ZCY^{+}20, ZWZ23b, dSFSS24]. alignment [ZM21b]. alkane/nitrogen [CSLQ22]. all-speed [HF23, JZZH22]. Allee [NEYZ22]. Allen [CHY20, Fab22, GR22, HZZ24, KW23, LCW20, NS24a, TCVC21, WCLY23, WZY23, WZD^{+}24, YX20, YZ24d, ZCY^{+}20, ZCYP20]. alleviation [WQ20]. allocation [JO24a, LY24a, Tor21]. alloys [HYY21]. Almost [APS22, AK22b, ZH24b]. almost-maximal [ZH24b]. along [FM22]. Alternate [JNCY23]. alternated [SDL22]. Alternating [KCL20, MV22, SRS21, Axe20, BBBK24, BM24a, CDD20, DVRV22, GBM21]. alternating-direction [DVRV22]. alternative [BCM20, OGR22, RO20, RLGGAV21, SCY23, WLZY21, YZLL22]. American [HSYZ24, Lee21, LH21, MVB20, MNH21, WSC23, XW22]. American-style [LH21]. AMF [GPHHA23]. AMF-W [GPHHA23]. AMFR [GPHAPR21]. AMFR-W-methods [GPHAPR21]. AMIB [YZL24]. amount [ZYZ23]. amphiphilic [DP21a]. amplitude [MHHV20]. analogs [OK23]. analyses [BS24a, DL23b, LL24e, YBM23]. Analysis [ADKH20, BFQ21a, BFLQ24, BLN21, BJMR20, CJLL21, CLM20, CLM21, CGM21, DS21, EKCF22, GS20b, GS20a, HC21, HLYZ24, JBSN22, JGLH20, JHZY20, Lee23, LZ23b, MCD24, MDVR22, NMB22, PRW^{+}22, PKE20, RVX20, SMM22, SKVA22, SMR23, Uça23, WK20b, WTL21, XCWY20, YT19, ZC22, Zha20c, ZZM20, ZEW23, ZMFL20, AAA21, AA23, AZL21, ABHK20, Bac19, BYWA22, BZ20, BFQ22, BGV20, Bha23, BB20a, BDPR23, BLRZ23, Bro23, CZHQ23, CE24, CLH22, CWLM22, CKY23, CJL^{+}24a, CLS24, CRRB21, CDJ20, CECCCR23, CDMM22, CHsL22, DKSQ22, DS20, DRR22, DG22a, DJ22b, DG22b, DC23, DWWZ22, DAZ23, DWZ20, EY23, Fab22, FM21c, FS20, FHB20, FAVM19, FAVM20, FPS22, FP24, FGP24, FMNT24, GXL22, Gao22, GANR22, GMSR18, GMSR21, GGVRB22, GXC24, GST20, GS21b, GT23, GLLZ21, GWWY21, HW21b, HH21a, HJM24, HM21a]. analysis [HZC^{+}23, HYW20, HCCSCGAM24, HH21c, HZZ24, HZL24, HLF19, HS20, HY21b, HL21c, HLLQ22, HSYZ24, HTTO20, JS22a, JT20, JYWZ21, JWZ21, JMLX23, JQ22, JZM^{+}23, JLL21, JCLC20, JVAD22, KA24, KZ23a, KZ24, KKO23, KUA20, KMUR24, KK24a, KP22, KV19, Kra22, LNSZ22, LYK23, Let24, LMZ22, LPY20, LWC20, LLS22, LZL^{+}23, LS23, LJ23b, LP24, LX24, LL24c, LCH23, LLLL20, LMS21, LWW^{+}22b, LYL23, LDQF23, LY24b, Liu24, LCZL24, LDM23, LY23b, LS22b, MYC22, MD22, MDG22, MH22, MS20a, MKN20, MSD21, MRG^{+}21, Men24, MH23, MGV23, MVFAMF21, NMMH23, NPP^{+}21, NNL^{+}24, OKW23, OPK23, PJM21, PJA24, PS22a, PW23, Pro23, PN21b, PNS21, Pul22, Rab21, RCVA24, RR22c, Rou22, RG23, RR23b, SL22, SA24, SDV21, SRMD22, SS24a, SC21, SLK24, SW24b, SZ24, SKVA23, SL24a, SLA^{+}24, SD21, TV22, TGJY22]. analysis [TTK23, UÖA23, USTG23, Ust21, WK20a, WCPW20, WC21, WW22, WLH23, WWW^{+}24, WKN22, WLZY21, XW21, XD23, XDX24, XH20b, YSN24, YX20, YTMK19, YYY21, YZ21a, YPLM22, YY24a, YHF22, YY21, YLS20, YLS23, ZZW^{+}24a, ZZG20, ZLG20, ZGJ21, ZCSL21, ZDL22, ZC24, ZG24, Zha24, ZC20b, ZY24, ZCT21, ZGG23, ZTS21]. Analytic [JCY20, Mik23, Mil22, BCA20, FU19, GJL23, GHJM24, JCLC20, NS20, SI23, Tam20, WTP^{+}21, XLZ20b]. analytic-signal [WTP^{+}21]. Analytical [BT23, CL24c, MMJ21, MHM23, Wan21b, BEG^{+}22, FFT20, FV24, KAS23, MNH21, OOO22, RKS^{+}22, Ruj23, ZUZ20]. Analytically [LLH24]. analyze [HZ24a]. Analyzing [HW24a, Kan24, GVT21, TFK23]. and/or [FHM23, GO20, WSLJ23]. Anderson [CV24, DHLT24, GRVZ23, YTA22, YYZB24]. angiogenesis [FJ24]. angular [WW23]. Anisotropic [RFGGM20, BM24a, CGP22, FSC24, HET24, HW20, HKL21, Kop20, LLYM21, LXY24, Mia24, MZY23, RB23b, SY24, SZ24, TL20, TL22, Ter20, WSY22, WLS24, ZCY^{+}20, ZW22]. annealing [ZMOW23]. annuities [SPS21]. annuity [JK21a, Wan24]. annular [MJK20, WKC20]. annuli [TL22]. anomalous [LLFT20, MD20]. anomaly [LGLC20]. antennas [PKK^{+}22]. anti [CNRR20, PPSM24, ZZMK23]. anti-community [CNRR20]. anti-Gaussian [PPSM24]. anti-triangular [ZZMK23]. antibody [TKM21]. any [Grz21, LL24f]. aortic [KCJ23]. aperture [JMA21, ZMYL23]. appealing [Els21b]. Appell [ÖÇ23]. Appl [AGMZ22a, AA20a, AA21b, AA21c, BPR21, BKMO20, FAVM20, GMSR21, GAA21, KZ24, LLG^{+}22, MHY21, MR25, YHC20]. Application [CX24, CNQRR22, DLL21, HJS21, HMA24, Kim20, MJK20, PV22, San21, SSS24, TZBK21, VLR23, VVB21, WZ23b, ADKH20, AS20a, ASON23, AD21, AAKA24, Alh24, AAA21, BKMA24, BCLS24, BGR20, BSH24, BGH22, BEL22, BBBU22, BAIB24, CGSVBB^{+}24, CKY23, CK20, CST19, Cso20, DG20, DJ22a, DV21a, EL23, ERB^{+}22, FFT20, FGM21, GM23, GBTS22, GAA19, GAA21, GX24, HZ24a, HL22a, HOGO23, Hon22, HLR22, HLF19, HLYZ24, Hun22, IAH24, JD22, JHW21, JLY^{+}21, JS22b, KKA20, KS22b, KD20b, KTK21, LZD21, LWG24, LHL19, LL24f, MHW23, MKS21, PH24, PC24, RS22b, Ret24, RPQ24, SZ20, SK24b, SCY23, SLK24, SS20b, SAZW24, SB24a, SCZ20, TSG22, TGLa21, WWS^{+}21, WT22, WSC23, WY24b, WYW^{+}24, WZL^{+}23, WMK22, WZC22b, XGYY23, YKAA21, YHW22, YLL20, YZ24a, YL21b, YJJW23, YYT24, ZXZ24]. Application [ACT22]. applications [AKM^{+}24, ABFSC24, AM21, AM22a, AB22, AAA^{+}23, Asa24, BTBR21, BSV20, BKM22, CZH21, DMP23, DAMA22, DO21, DH24a, DDT23, FMRS24, FB24, FH24, FS21, GJ19, GHG24, GPP^{+}20, GDSCO21, GPST23, HLW20, HLLZ24, HIM24, HL24b, HHL20, JBSN22, JLZ23, KWI24, KPW21, KWYN23, KAM^{+}23, LNSZ22, LWLW23, LTF^{+}20, LYL^{+}24, LZ24b, LL24e, MSW23, MYZ23, MAA^{+}22, MS24b, OVGI20, PR24, QXZF23, RTT24, RS22a, SK24a, SY23, SI23, SRBK23, SJZ20, SKZ23, TFM24, Vab22, WLL21, WSLW23, WDPY24, WJVG24, WLDZ24, WSL^{+}23, XJK19, XL23, XTL23, Yam23, Yam21b, YL22, YHF22, hYbJzJ21, YHQW21, ZTY^{+}20, ZZ24, ZL23a, ZTY^{+}24, eBI23, GPHAPR21]. Applied [Sug20, ZZW^{+}24b, ADDG22, BHPV23, CCJ20, LLL22a, MM22a, MR23, MR25, OBdSD24, ORS23, PP22, RU22, TFR^{+}20, VNV23]. Applying [GeOJD24, HQL21]. Approach [ACD22, CLM21, AHB20, AMZ24b, AMZ24a, AID24a, AJSGL22, ACT22, ÁGKM21, BCMP23, BD23, BJB22, BYE20, BCM20, BBTW24, BVBO23, BK22, BPD20, BCCW22, BGIR24, CDL24, CFS21, CW22, CLS20, CG22, CPBG23, CLM22b, CSS24, CMJN22, CMRS21, DR23a, Dax20, DO22, DU19, FMRT20, FC23, FV24, GJN20, GHL21, HL21a, HM21b, HSGLE19, HA20b, HKU22, Ixa24, JC24, KA24, KH24, KPW21, Kog21, KTK21, LYM22, Li20b, LH21, Ngo22a, PT20, PT21, PGG20, PCWY20, RLGGAV21, Ruj23, SM20, SUW23, SMKV21, SK23a, SAL21b, SYG19, WCW20, WL24b, WHC^{+}20, XH20a, YSBT20, YZLL22, ZZW^{+}24b, ZZW^{+}24c, ZSH20, ZZ21, ZaZWZ23, Zhu21, CLM20]. approaches [CLS21, GLHL24, HBPV21, RZA^{+}24, VS23]. Approaching [MGV23]. approximants [CS24]. Approximate [FPCV^{+}22, HK21, Kan20, RO20, Vab23, AHGG20, BWS21, DRR22, FGM21, KHAW21, LMM22, LZWL23, Mam23, MORO20, OOO22, RVS^{+}22, RKS^{+}22, TZTC20, Vab21, Vab24, VOJS24, YJN20, ZYZ22, ZHW20]. approximated [PT24]. Approximating [Kaz21, FMRT20, GHG24, GW24, LMZ22, Sha23, XZ22]. Approximation [BGH22, BA23, DFK20, FPS21, Jia20, KN20, MPPR23, PS21, SAL21b, ZuIH20, AMS23, AGS24, AIDAJ23, AET23, AV21, ACG20, BJ22, BZH24, BC23, BP21, BCGAR22, Bos24, BHJB21, BGIR24, BM24b, CGP22, CMFMSQ22, CLHL20, CLH22, CDS21, CJLL21, CRRB21, DIAJ^{+}22, DS22b, DMNO24, DPW24, FAUU20, FLZ21, FW22, FRT23, FS24b, GW23, HJL20, HCCC21, HCCZ24, HSW21, HG22, HIM24, HKL21, HH24c, JJR21, JS22a, JWZ21, JHYZ24, JQ22, JNCY23, KWI24, KEKT21, KAMW21, KNM24, KMN20, KZM21, Lem24, LP23a, LWYZ21, LG23, LL24c, LFZ24, LHL19, Liu23, LH23, LRQV24, LS22b, LWZ21, LXYL24, MWHC20, MYC22, MJ20, MDS20, MA19, MP21, OSK20, OT21, OCS24, PHV24, PSRVA24, PSS24, RFGGM20, Saj23, SDV21, Seg23, SS20b, SC23b, SLA^{+}24, Ust21, WZ20a, WLL21, WZZ21a, WC24b, WSYPD24, XLLA21, YZ20a]. approximation [YQZS22, YZ21a, YLZ^{+}22, ZuIKA22, ZZW^{+}24a, ZRQ22, ZDY20, ZY20, ZEW23, ZH24c]. approximation-based [LH23]. Approximations [RR22b, SZK20, AA24c, BS23, BCVV20, CWY21, Dax20, Dax24, DJ22b, DN20a, DN20b, EGG22, EMR21, HN22b, LLY21, LV23, MVY19, MP24, Mas21, Mil22, MJ24, SI23, Sou22, WLJ24, Wol24, Wu24, ZG22]. April [Ano21l, Ano22m, Ano23m, Ano24k]. arbitrage [DKL20]. arbitrage-free [DKL20]. Arbitrarily [BIZ20, FHB20]. arbitrarily-oriented [FHB20]. arbitrary [DKM^{+}23b, Han24, LS21, LLZ24a, McA21, SMDVA22, XLKL23]. Arc [Zag23, KPS24b]. arch [SKS22]. Archimedean [BABR20, CFL21, FBJ20]. architecture [Gra20]. architectures [ERB^{+}22]. arcs [VZ21, Zag23]. arcsine [LLGZ20]. arcsinh [CECCCR23]. Arctic [FMTB21]. ARDL [WHC^{+}20]. area [HSRA24]. areas [HKU22]. argument [AN20, DdO22]. arguments [HAD24, SSL24, YW23, ZSL22, ZSL23]. arising [BFQ22, BMV23, BKMO19, BKMO20, DYM21, FAVM19, FAVM20, GST20, HR22, HBMA22, LJCL20, LCD21, LWT24, LS22b, MKS21, SMM22, TKT^{+}24, VAMO24, YT19]. arithmetic [IY20]. armed [DR23b]. Arnold [CM22]. Arnoldi [APR24, EMR21, FW22, HWH^{+}21, MT20b, WHY^{+}21].

ARPIST [LJ23a]. array [PKK^{+}22]. arrays [ZTY^{+}24]. Arrival [GMSR18, GMSR21, GHF24]. Arrow [GRVZ23]. art [GDG^{+}21, BK20]. arteriosclerosis [LJCL20]. artery [WAAM20]. article [Sug20]. artificial [DXW21, GS20b, LN23, SUW23, SSV23, ŚK23b, Tia24, VSSV20, ZX23a]. Asian [MW20, MST22, RG22, Rou22, Wan21b, WHC^{+}20]. ASOR [GH20]. aspects [BDR23, SC23b]. assemblies [CTGV20]. assembly [BPDC20, LLK23, PJM21]. Assessment [Kap21, LNP19, BD23, Cho24b, DLL21, MBG21a, MLS22]. assessments [BT23, MS20b]. asset [Boe20, BitH21, DKL20, Fan22a, JO24a, KKKL22, YSBT20, ZCJL20]. asset-liability [ZCJL20]. asset-or-nothing [YSBT20]. assignment [YL22]. assimilation [DS21, GANR22, SLP20]. assistance [MHY17, MHY21]. assisted [CW21a, GGVRB22, IY20]. associated [AKP23, DL20c, HL22c, KS22a, Ver20]. assumptions [LN21]. asymmetric [ATO22, LL24e, MTP21, NY24a, ZZWS22, ZQL24]. asymmetrical [GMK20, NN24]. asymptotes [BPD20, PDBdS24]. Asymptotic [ABFSC24, BSH24, CYC21, HOGO23, MPV23, Pro23, WXS21, AAP24, GST20, HCH^{+}23, HH21c, KK20b, Let24, LLY21, LL21, LPP23, LPP24, LH24, OH23, SKE20, SBM21, VVB21, ZIZ21, ZFW22]. asymptotical [XXL21, ZG20a]. Asynchronous [GBM21, GBM23, ANR20]. Atangana [LYL^{+}24, WK20b]. atmospheric [OBB23]. attraction [GGT24]. Attractivity [AL16, FL23]. atypical [GNGGL^{+}19]. augmented [CE24, CGM21, DV21a, JHYZ24, LZ20a, Mia20, PL24b, YJN20, ZIZ21]. August [Ano21a, Ano21g, Ano22a, Ano22g, Ano23a, Ano23g, Ano24a, Ano24f]. auto [NGNB23]. auto-tuning [NGNB23]. autocorrelated [LSZ22]. autoencoder [ZJ24]. Automated [BBTW24, TKT^{+}24, VPHE23]. automated-vehicle [TKT^{+}24]. Automatic [CWT23, GBG23, HS24a, SL21]. automobile [NK22]. autonomous [CT20, CFdD23, Fal21, LLW22]. AutoPas [NGNB23]. autoregressive [CLW^{+}22, CECCCR23, SUW23, SLK24, XGYY23, ZW23, ZGZW23]. Auxiliary [QH23, XDP24, CKA21, DXM^{+}20, EFTW21, GLHL24, GK21b, LC23, MHM23, MIdMR24, PS24a, YSUY24]. availability [Kan24]. average [ACK21, GDZN23, MW22, WCL22, XGYY23, YY21]. Averaged [RS22a, DDRS21, DDRS24, DG22b, EK24b, RS24c]. averaging [Ais23, DRB22, LRP24, XY24, YSFS22]. aversion [CL24b]. AVF [HL21b]. AVX [CKT21]. AVX-512 [CKT21]. axisymmetric [Kog21, LRQV24].

B [AAA20, BEL22, BPDC20, BC21, BC22a, BL21, DU19, GS21d, LNP22, MA19, MA21a, PJM21, RG22, WZNT24, ZM20]. B-spline [BL21, DU19, GS21d, LNP22, MA19, MA21a, RG22, WZNT24, ZM20]. B-spline-like [BEL22]. B-splines [BPDC20, BC21, BC22a, PJM21]. back [HHM22a, JLL21, CQBL20]. backbone [WTP^{+}21]. background [ZLZ^{+}24]. backtracking [TLY23]. backward [BPRZ21, CG20, CLYS24, FZQ22, ISD23, LD21, LF21, NZA24, SI23, SXX21, YZ21d]. Bakhvalov [NMMH23]. Bakhvalov-type [NMMH23]. balance [Hri22]. Balanced [RB23a, SV23, RAN20, YC20b]. Balancing [LMS21, KK20a, LLK23, Pat23]. Baleanu [LYL^{+}24, WK20b]. balking [BS24a]. ball [MPPR23, SZZM23]. BAMPHI [CCZ23]. Banach [GG24, GHTC21, HWLD23, LY23b, NB23, WLJ24]. Band [NT20, CLS24, LNP22, ZZNL24]. bandit [DR23b]. bank [ÜY20, WK20b, YC20a]. bank-specific [ÜY20]. banks [XCWY20, ZK20]. Barca [GNGGL^{+}19].

Barnes [Kuz22]. Barone [BO22]. barrier [DLS20, JCY20, KKO23, KYP21, LH21]. barriers [WWC23]. barycenter [KDLY22]. barycentric [ABP24, BE21, IpAmK20, JK21b, LHH21, Ma24b, Tia22, eBI23]. Barzilai [ZL22]. Based [PM20a, ASADB23, AAP24, AVRLRMGG23, AHB20, AHB21, AFJGJJ22, APR24, AUAA20, Ant20, AID24a, AID24b, AJSGL22, AB23, AK22a, AAk20, AKP23, ARK20, ÁGKM21, BJB20, BV22, BHRW22, BLZL22, BYE20, BSMM20, BR22, BBTW24, BBBK24, BCA20, Bha23, BJF24, BYE^{+}23, BGIR24, BHPV23, BM24a, CZ20, CTGV20, CD23, CDLR24, CUN24, CSZP20, CCCMH24, CG22, CWLM22, CLYS24, CL24b, CJL^{+}24b, CJLL21, CLPZ21, CECCCR23, CES23, DMP23, DTMJ24, DR23a, DDF23, DO22, DP21c, DHLT24, DWZ20, DS21, DLZ22, Dum24, DN20a, DN20b, DRR^{+}21, EJHR24, EL22, EKÇ20, EMR21, FM21c, Fan22c, FS24a, FJMP24, FPGSR21, FMNT24, FHH21, GS20a, GXL22, GLHL24, GMW24, GNGGL^{+}19, GG24, GBM23, GEG23, GWHZ23, GHF24, GW24, HJ23, HLW22, HZC^{+}23, HV23, HA21b, HTJC24, HSW21, HR23, HTD21, HQS20, HLLW21, HC23, HTZ21]. based [HH24b, IS23, IpAmK20, ISD23, JB21, JLY^{+}21, JH24, JS22b, JJ24, KZ23b, KAD22b, KTK21, KDLY22, KL23b, KS21, KRV24, LGG^{+}22, LRV23, LZW20, LRT20, LWC20, LZD21, LMSZ21, LWZF21, LZH21, LDY22, LW22, LWL22, LXLL23, LW23a, LWLW23, LJ23b, Li24a, LZLL20, LLX^{+}22b, LWWN22, LV23, LCGW24, LST^{+}24, LH23, LLGZ20, LN23, LS24, MHW23, MWLY24, MÖ21, MRI22, MIdMR24, MI20, MDS20, MMJM22, MFH24, MDE24, MJK20, NKBJ20, New24, NY24a, NRGJ21, NP23, Oga20, OA20, OCS24, OMRdQ24, PS24a, PVV24, PWL23, PKPD21, PSRVA24, PT24, QGW21, QDHW21, RKA20, Ray23, Ri22, Ri24, RFB22, RG22, RPQ24, SuR23, SES20, SCHS24, SMDVA22, Sha23, SZZ21, SCY23, SDW24a, SKVA23, SZL20, SYTD24, SHZ20, SPC21, SZZM23, SL24b, TBA22, TFM24, TLY23, TSMZ20, TLD^{+}23, TWQ^{+}23, VdG20, WYTL19, WDTW20, Wan21b, WWH21b, WDS22]. based [WT22, WSLW23, WPX23, WC24b, WL24a, WLB^{+}21, WLDZ24, WD20, WZC22b, WLZ22, WSL^{+}23, WZZ24, XLKL23, XY22, XHLD21, XFW24, XZZ^{+}20, XLZ20a, XDP24, XH20a, XL20, XA21, YSN24, YN21, YLL20, YTA22, YLXY23, YZL24, YSFdRdP20, hYbJzJ21, YJJW23, YCD^{+}20, YL21c, YYZB24, YWW24, uDuIZ22, ZvS23, ZYZ22, ZZW^{+}24b, ZZW^{+}24c, ZWZ^{+}23a, ZG20b, Zha20b, Zha20a, ZWH^{+}20, Zha21a, ZLZL23, ZGZ23, ZZLY23, ZY24, ZWL21, ZZ23, ZCSS24, ZMZ^{+}24, Zhu20b, Zhu22, ZLL^{+}20b, HLLZ24]. bases [AHPS22, BEL22, GANU21, MMPP20, MDE24, PS22b, Rab21, TR22, WZNT24]. basic [BFRV21, Ust23]. basis [AID24a, BEIR23, BWS20, BWS21, Bra23, FGPP23, GJ21, KH21, LMLB22, LP23b, Men24, MBG21b, MMV20, NWD20, NNL^{+}24, PL24b, RG22, Seg23, Son24, WCK20, WKC20, Wör22, XTL23, YSBT20, ZH24c]. basis/cardinal [KH21]. bat [SVK20]. batch [LLZ24b]. Bauer [KCK24]. Baxter [Lu24]. Baxter-like [Lu24]. Bayes [ANN21]. Bayesian [ASTK24, BJ22, Bha23, CDL24, CDLR24, CYY23, SJZ20, WZZ24]. Bazykin [NEYZ22]. BB [RYH24]. BDDC [PSWZ21]. BDF [GT23, WA21]. BDF2 [HZZ24, KW23, LWA^{+}22, LL24c, LCZL24, WMH22]. BDF3 [LWA^{+}22]. beam [HZC^{+}23, OKW23]. Beavers [CHLY21]. bed [DL20b]. bee [DXW21]. behavior [BHJB21, OEH23, WXS21, YHF22]. behavioral [CL24b, MER20]. behaviour [Gök23, NS24b, OKW23, RTV20, SKE20]. Bellman [BK21b]. BEM [ADDG22, ACG20, WZ24a]. Bénard [ABHK20]. Benchmark [EGG22]. Bending [NNL^{+}24, JO24b, LLT^{+}21, MJK20]. benefit [JLL21, WRZW21]. benefits [GJM^{+}23, ZYY20, ZCZ23, ZSZ24]. Benney [HR22]. Benney-Lin [HR22]. Benz [JW24]. Berezovskaya [NEYZ22]. Berge [CFB24b, CFB24a]. Bernoulli [DIAJ^{+}22]. Bernstein [BCLS24, BEIR23, Bic23, BWS20, BWS21, GSZ22, HVM22, LL22, MPR24, SMM22, SMRN24, Ust21]. Bertrand [Yu22]. Bessel [KZ24, CYC21, Den24, KZW22, KWZX23, KZ23a, LCD21, LC24b, WK20a, ZuIKA22, ZSMG20]. best [IHK20]. beta [BKMO21, MKS21]. Bethe [BP22a]. better [HC20a]. between [AHPS22, AJ21, AAA20, GVT21, Juh21, Kan24, LWLD23, MAA^{+}22, MLS22, NT23, NPP^{+}21, PCWY20, SLC22, UDI20, VdG20, WZJQ21, XL20, ZH23, ZZ21]. Beyond [ZZWS20, CHsL22, HMT23, LP23a]. Bézier [AM23, DOP23, HF24, Juh21, LZW^{+}22, LHL19, OMRdQ24, SMRN24, WZ20b]. BFGS [AKM^{+}24, LZD21, LN21, Zho20]. BFGS-like [AKM^{+}24]. Bi [XZZ^{+}20, YT24]. Bi-directional [XZZ^{+}20]. bi-Hamiltonian [YT24]. bias [CM24a, WN23, ZMZ^{+}24]. bias-corrected [ZMZ^{+}24]. biased [YL20, YL21c]. Bicriteria [GJGDPR23]. Bidiagonal [DKM^{+}23b, DOP21]. bidiagonalization [ACR24]. bidomain [CK22]. Bifurcation [LZL^{+}23, CW21a, LHES21, MRG^{+}21, ML23, SDV21, YBM23]. Bifurcations [SS23c, YAE23]. big [CYY23, DLL21]. Biharmonic [YZZ20a, HY21b, JO24b, LWW23b, LWWZ24, MM22b, PO21, TZTC20, YMC22, YZ24b]. bilateral [KS23b]. bilevel [GM23]. bilinear [DXM^{+}20, GDZN23, GLLZ21, mHLZ20, Liu20, SS20a]. bimatrix [WYW^{+}24]. Bimodal [GDSCO21]. Binary [XLKL23, HYY21, LLLY21, MK24]. Bingham [ZHMD22]. binomial [OK23, QG20, SLK24, ZW23]. bioconvective [CGM21]. bioelectromagnetics [KA21]. biological [KK24a, RKS^{+}22]. biology [BTBR21]. biomarkers [BVBO23, TVB21]. biomedical [Yam23]. biomedicine [LTF^{+}20]. bioresource [Ret24]. biorthogonal [BCP20, LSC21, LS22b]. Biot [ZLC24]. bipartizing [CNRR20]. bisection [LL24f]. bit [GGPP21]. Bitcoin [CLS20]. Bivariate [LHH21, MJ20, BSNL19, BG21, BWS21, BKMO21, CS24, HBMA22, MA21b, PL24a, TE20]. Black [KDO20, RG20a, ACD22, AF20, CLS24, Kaz23, OT21, RHM22, RG20b]. Black-box [KDO20]. blades [RMG21]. blending [Kra22]. Blind [LLLY21, DRR^{+}21]. Bloch [XLLA21]. Block [MJR20, WZC^{+}22a, ZS22, Zho24, ASB22, ALKH21, BEJR22, Cao20, DDKD19, EFJK22, GS20b, HLLZ24, Hon22, Hua21, gHtSjT24, JLZ23, JLY^{+}21, JZY22, KM20, LZF20, LHL^{+}22, LZ21, LWT24, MW22, NL23b, ORS23, QDHW21, RR22a, RR23a, SZZ21, SLXF24, SSRK23, SZL20, SFI23, WL22, WSLW23, XY24, YZ20a, YQZS22, YCXW24, YHC20, YLS23, ZSCH21, ZG22, ZH24b]. block-based [QDHW21]. block-centered [LHL^{+}22, SLXF24]. block-Hessenberg [DDKD19]. block-Lanczos [JLZ23]. block-pulse [KM20]. Blockchain [XLZ20a, DXW21, Fan22a, GS20a]. blocks [BG23, CMVV20, KCS21]. Blomqvist [BKMO21]. blood [KCJ23, WAAM20]. blossoming [BEL23a]. blow [Fro24, MT20a]. blow-up [Fro24, MT20a]. blurring [JH24]. Board [Ano19, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, Ano24y, Ano24z, Ano24-27]. Board [Ano24-28]. bodies [LR22, OGR22]. body [CM22, HHL20, KP20, Khl21, LSE24, NT23, NMB22, Slo21]. bolt [Grz21]. Boltzmann [HBPV21, ZY24]. bond [AHB20, LST^{+}24, XCY24]. bond-based [LST^{+}24]. bonds [MN20, ÖS20]. Boolean [ABM^{+}22]. bootstrap [WHC^{+}20]. Bootstraps [Bon24]. Borges [PPSM24]. Borwein [ZL22]. boson [JS21, MD22]. boson-fermion [JS21]. both [BB21a, NY24b]. bottom [KD20a]. Boubaker [DR23a]. bound [DV21a, GDA22, HL24b, JPS24, JWS23, KWN20, LLY21, RKA20, RK23, SWW23, SDW24a, ZSHD24]. bound-preserving [RK23]. boundaries [BL23, DZLC20, LHG^{+}24, RR21]. Boundary [MJR20, VLR23, AGH20, BT23, BSMM20, CHXL20, CLX21, CTN22, CJM20, CDD20, CHH20, CKA21, CM24b, CD20b, CRRB21, DRR22, DL21, DDF24, DT24, EOR21, Ela23, FM21c, Fan22b, FP24, GCV20, GHJM24, GS21b, GO20, GS21c, GO24, HB23, IHK20, JO24b, JG21a, JLW23, JM21, Kim20, Kim22, KAA22, KS21, LMZ22, LZW^{+}22, LLS22, LLG24, LLLL20, MW20, MVB20, MY21, MH22, MS22, MP24, MBZ23, Mik23, NL23a, ON22, PT21, PJA24, PP24, PFG24, QXZG22, RR22a, SS24a, SMDVA22, SK23a, Sou22, SGD20, SCZ20, TKVA21, UL23, WLHJ20, WCL22, WSC23, vtW24, WZ20b, WDY20, YZ20a, YQZS22, YDC20, YCH^{+}21, Yüc21, ZvS23, ZSMG20, ZF24, ZY24, ZX23a, ZOK21, ZS22, ZHW20, ZWKS24].

Boundary-safe [VLR23]. boundary-value [JM21]. bounded [ATO22, FYL^{+}24, HTOC20, Hun22, JGW^{+}23, Juk20, LG23, OMRdQ24, TFK23, XZP22, YS24a, YYT24, ZY20]. Bounding [BE21]. Bounds [KZ23b, TG21, APR21, BBSV21, BG23, BKMO21, CX24, CW22, ERB^{+}22, Fro24, HTZ21, IM24, LMP20, LZ19, Liu20, LV23, MORO20, Rum20, RL23, SV23, tSjTsJ24, Wu22, YT20]. Boussinesq [DG22a, DHH24, GG24, LGY24, PKE20]. box [KDO20, LWG24]. box-constrained [LWG24]. BP [San21]. BPX [OF22]. branch [HL24b, JWS23, RKA20, SWW23, SDW24a, ZSHD24]. branch-and-bound [RKA20, SWW23, SDW24a, ZSHD24]. branch-reduction-bound [HL24b, JWS23]. branching [SWW23, SDW24a, ZSHD24]. Bratu [KLH21]. Bravais [LLL^{+}22b]. breaking [CW21a]. Bregman [JMLX23]. Bregman-type [JMLX23]. Brinkman [CRRB21, CE24, CGO24, DZZ24, WLH23]. Brownian [AHT21, AS21b, CN24, KKJ22, LHHW23, Liu23, NP22, SPMB21, TTLB22, WD20, YYY21, ZHZ24]. Broyden [LN21, YWL24, ZZ20b]. Broyden-like [ZZ20b]. Broyden-type [YWL24]. BSDE [SYG19]. BSDEs [HJ23]. buckling [MJK20, PH22]. Building [Wan20b]. built [ŚK23b]. bundle [WPX23]. Bürer [ACT22]. Bures [HKM24]. Burgers’ [GJL23, DGL24, JA22, SC21, SCK22]. Burr [QG20]. Burr-XII [QG20]. burst [DH22]. butterfly [CLU23]. butterfly-type [CLU23]. Buzano [KZ23b]. BVP [LG23, Vab21]. BVPs [BM24a, CKA21, LP24, MP24]. bypass [Kes23].

C [LPH23]. Cahn [BS20, CJL^{+}24a, BDS20, CHY20, CG22, DWWZ22, Fab22, GR22, GLHL24, GWWY21, HZZ24, JGLH20, KW23, LLZY22, LLX^{+}22a, LLG24, LCW20, LY21, MTIR22, MBZ23, MZY23, NS24a, TCVC21, WY20, WWJ22, WCLY23, WZY23, WWW^{+}24, WZD^{+}24, YX20, YY24a, YYC24, YOCY22, YZ24d, ZZG20, ZCY^{+}20, ZCYP20, ZJGZ20, ZOW^{+}21]. calculating [Xia22]. calculation [Oga20, PLZ21]. Calculations [SYTD24, HOOT20]. calculus [ABA19, ARF23, FU19, LYL^{+}24, MMM22, MMM23, TTK23, VORD20]. Calibration [GJNM24, MIdMR24, MNH21, MRI22, MIdMR23]. call [OT21, RG20a]. Camassa [LXZ23]. Can [VORD20]. cancer [KS22b, PC24, UÖA23]. Canonical [MMS^{+}23, YHQW21]. cantilever [HZC^{+}23]. Cantor [GN24]. cap [HH21b]. capacity [BS24a, FSC24, GM23, HNV23, NV20]. capillary [PH21]. capital [LY24a, LCZ19, XL20]. capital-exchange [LCZ19]. capped [ZL24]. capture [HLR22, YO20]. capturing [CUN24]. Caputo [AZD21, CHXL20, CBJ20, DO22, DHKV20, Evi23, FB24, GS21c, HZMD22, MY21, RKS^{+}22, SESH23, TZ20, Uça23, WOV22, WX22]. Caputo-Fabrizio [DO22]. CARA [HCH^{+}23, SL23]. Carathéodory [DPW24]. carbon [SZWH21]. carbonate [WCC^{+}19]. cardiac [CK22, CK24a]. cardinal [HR23, KH21]. Care [Yam23]. career [Tia24]. Carleman [ALNK24]. Carlo [VTBM21, DMZ23, HJ23, Har21, MMS^{+}23, XW22, XFW24, XHW24, YYBL24]. Cartesian [XLY22]. cartoon [WW24a]. cartoon-texture [WW24a]. Cascadic [XH20a]. case [BYE^{+}23, CC21b, FDDM24, GJNR21, GNGGL^{+}19, GS24, HZ23, HCCSCGAM24, MPPR23, MJ20, Ruj23, SB24a, Yam23, ZGZW23]. cases [Ngo22a]. cash [YSBT20]. cash-or-nothing [YSBT20]. catastrophic [ÖS20]. cattle [BYWA22]. Cauchy [CJM20, DdO22, GWHZ23, HLZ20, IuR23, KZW22, KWZX23, NNCH21, OCS24, OHK23, PL24a, Vab23, VCNR20, Wan20a, WYZZ22, YSN24, ZM21b]. Cauchy-Euler [IuR23]. Cauchy-Stokes [OHK23]. Cauchy-type [PL24a]. causal [LSY24, Ri24]. causality [PCWY20]. causes [WTL20]. cavity [WCL22]. Cayley [WGM21]. CCBM [GWHZ23]. CCBM-based [GWHZ23]. CCC [Bos24]. CCC-spline [Bos24]. CCT [KPD23]. CD [DAMA22]. CD4 [WMF20]. CDG [YZ24b]. CDS [MCW24]. celestial [OGR22]. Cell [TT20, Ter20, AGNGG^{+}23, BZ20, HW24a, Mia24]. Cell-centered [TT20, Ter20, Mia24]. cells [WMF20, YO20]. cellular [PV22]. CEM [CHP21]. CEM-GMsFEM [CHP21]. Censored [PM20a, VBRT24, Cha24, LMSZ21, LWLD23, Pou21, QG20, YCD^{+}20, ZG20b, Zhu20b]. censoring [DNK23, ER20a, ER21, GC19, WTL20, YWW24, Zhu23, Zhu24]. center [HTJC24]. centered [LHL^{+}22, Mia24, SLXF24, TT20, Ter20, WSY22, YLS23]. centers [BL20]. central [HZ23]. centres [CKA21]. certain [Fro24, JPS24, LZZ24, UPPZ24, WX21, YC23]. certificates [BSdCCGV23, HAFA22]. Certified [KLMP21, WSYPD24]. cervical [PC24]. CESTAC [NAS22]. CEV [Ara20, AV21, CKZ21, FM21b, Lee21]. CFD [MBG21a]. CG [AF23, BKMA24, LZD21, LH20, SL24a, WZZ21b, WY24b, WDPY24]. CGP [JH24]. CGPM [hYbJzJ21]. CGPM-based [hYbJzJ21]. Ch [USTG23]. Chain [SS23b, VTBM21, CQBL20, FSR24, GS20b, GXL22, Har21, MYC22, San21, SZL20, WQ20]. chains [DDKD19, MR20]. chance [TBA22]. Chandrasekhar [HVM22]. Change [SLK24, AS20a, NPP^{+}21, PHV24, SD21, YL20, YL21c, ZLL^{+}20b]. Change-point [SLK24, YL21c, ZLL^{+}20b]. change-points [YL20]. changed [KKJ22]. changes [Men24]. changing [YCH24]. channel [CWLM22, SH21, WAAM20]. chaos [Bha23, EPL21, GMCJGM24, LZW20, MLYZ21, SS23c, SPC21, SH24b]. chaotic [PZY^{+}24, SAN^{+}24]. characteristic [BJF24, CHLY21, CLL20b, LHL^{+}22, MTZ23, PFG24, SEAS22]. characteristic-finite [CLL20b]. characteristics [PKK^{+}22, ZAH^{+}24]. Characterization [AN24, JS22b, GBG23]. Characterizations [ZCT21, SS20b, WJ23]. charge [AGNGG^{+}23]. charged [BIZ20, LJL24, Wan21a]. charged-particle [Wan21a]. Charlier [LZ22, LSZ23]. chart [AK20, ACK21]. charts [ASON23, MÖ21, RZA^{+}24, SCHS24]. Chebyshev [AAA^{+}23, BJB20, BVDB20, BCGAR22, BCKSV24, DMNO24, DDRS21, ESA23, ER20b, GLCL20, HR23, Ibr20, JWSG22, LZ21, MT20b, Moi24]. Chebyshevian [BCR22]. check [YLL20]. chemical [CKT21]. chemistry [BMS22]. chemo [GGT24, JZW22]. chemo-attraction [GGT24]. chemo-repulsive [JZW22]. chemotaxis [BGG^{+}21]. Chen [HZZ21]. China [LLG^{+}22, LGMC20, PCWY20, SZWH21]. Chip [Yam23]. chirp [CJLL21]. choice [BPR20a, BPR21, FPLZ21, MYLL24, ZZ22a]. Cholera [ÖHS23]. Cholesky [WCK^{+}22]. Choosing [NK24]. Chorin [RVX20]. chosen [Six20]. Christoffel [BD24]. Ciarlet [PKE20]. CIEL [USTG23]. CIR [LET23, Wu24, YLWW21]. circle [BCA20, BM24b, DMST21, RR22b, Ver20, Vie21]. circuit [AM21, DDT23]. circuits [AAKA24]. circulant [Hon22]. circular [KMZ21, SC24, VZ21, Zag23]. circulation [AMS23]. citie [BD23]. city [BD23]. civilization [SZWH21]. Civita [FS21]. claim [NK22, ZYZ23]. claims [SHZ20]. clamped [BL21]. class [ASADB23, AHGG20, AC22, AH22, ADEEY21, AM22b, BKMA24, BLZL22, BGH22, BL20, BR20a, BIJ21, Bro23, ÇK24b, CYC21, DRR22, DDM20, DJR21, DL20c, DAM23, EP20, GZ23, GR22, GMK20, GS20c, GS22, HBGS21, HBLZ21, Hua21, HTOC20, HTZ21, JZ21, LDY22, LVW24, LC24b, LXYL24, Ma24b, MARH^{+}23, MYZ23, MVFAMF21, PH24, PT21, Pie20, RTT24, RWW22, SZ20, SYG19, THS24, WL22, WZ24a, WLB^{+}21, XWZ24, XLZ20b, XA21, YZ20a, YCCZ24, ZZZL21, Zha24, ZCT21, Zhu20a, dSFSS24]. classes [CT20, PP22, WK20a, WN23]. Classical [KDOT24, dS24, Cal23, FPCV^{+}22, KS22a, MMM23, OH23, ZD21]. classification [DJ22a, JD22, KBK^{+}22, LYYZ24, LHS^{+}24, SBC^{+}24, YC20a]. classifier [LHS^{+}24]. classifying [WXZ21]. clauses [LLW21]. Clement [AHK24]. Clenshaw [KEKT21]. climate [NPP^{+}21]. climb [NY24b]. closed [Cho24a, GL24, LC24b, OHCC22, SH21]. closed-form [LC24b]. closed-loop [GL24]. closely [XLY22]. clothoid [FPS21]. clouds [BLV22]. cluster [MMS^{+}23, MMV20, SŁK^{+}22]. Clustering [NN24, DR23b, FLMR20, FSR24, JZY22, LWG24, TFM24, VPHE23, WY24b, ZLL^{+}20b]. clusters [ARMP20, RL23]. clusterwise [BTC21]. CMMSE [BA22]. CMRH [GHC^{+}20, GS22]. CN [LL20b]. CN-WSGD [LL20b]. CNN [HL22b]. CNTs [FFT20]. co [ELTD23, HKU22]. co-registration [HKU22]. co-simulation [ELTD23]. coagulation [GSK20, GPK23]. coagulation-fragmentation [GSK20]. Coarse [ZMOW23, Axe20, RFF23, WCK20]. coarse-fine [Axe20]. Coarse-grid [ZMOW23]. coastal [FMTB21]. cobweb [CBJ20]. code [CTGV20, HS24a]. coded [JMA21]. Codes [DH22, Bra21]. Codimension [DP21a, YAE23]. coefficient [CLW^{+}22, CLM22a, DZ22, Fan22b, HYW20, HA20b, HTIA24, MDY20, SC23a, SVJG24, SCZ20, WX22, XA21, YZW21, bYW23, YH22, YZ21d, ZWY22, Zha20c, ZEW23, ZMZ^{+}24]. coefficients [AGH20, Abi20, Alh24, BGN^{+}20, BK21b, CMW24, CLL20a, CCETZ21, CG20, DFFM21, FM21a, FLT20, GZ23, HL24a, KRT23, KS21, Li24a, LV20, MkCZ20, McA21, PM20b, QS23, TM24, WYC24, XZL20, YWKM19, YLFT21, YZ21a, YLLA23]. coercive [DMTW22, PS24a]. Cohen [AB23, HLLQ22]. coherence [ACT22]. coherent [BCLS24, EP20, Jas22, Jas24]. cohort [XCY24]. coincident [CPBG23]. coinfection [BYWA22]. cointegration [Fra21]. cold [DNS21]. Cole [WZZ21b]. coli [BYWA22]. collaborative [SKZ23]. collision [CKT21]. Collocation [MH23, SYD21, ASADB23, AZD21, AID24a, AID24b, AS21b, BJB20, BJB22, BR20a, BDF21, CKA21, CM24b, DP21b, DO21, FM21a, Gu21, HM21a, HM21b, JWZ21, LBBB23, LGY24, MH21, MH22, MA21a, NAA20, Pat23, QXZG22, RCM20, SS24a, SZA24, WLZ23, YWZ24, ZM20, ZFM22, ZZ23, ZH24c]. Colloidal [RB23b, DRB22, YLK24]. colony [DXW21]. color [MHW23, USTG23, WY24a, WW24a]. colored [LCGW24]. coloring [LM24]. coloured [YZ23]. Combination [DOR21]. combinations [CCZ23, DOP23]. Combinatorial [AAMR24, JGW^{+}23, JGZ^{+}24]. combined [CDL22, OWV24]. Combining [CMdMRFG22, MHW23]. combustion [FDDM24, LH24, MDSRFO24]. Comment [For24, FL23]. Comments [YHC20, dF20, NRV21]. commercial [XCWY20]. commodity [GXL22, LWC20]. Common [JZ24, Kum22, BM21, BWS20, EK24b, HWLD23, RTT24, TH24b]. community [CNRR20, YBAE23]. Comonotonic [DKL20]. Comp [GMSR21, MHY21]. Compact [Cap20, CKL20, FHM23, Rab21, Abi20, Dax24, HZ23, HL24a, JZZH22, KWN20, KRV24, LL20a, LLS22, PM20b, RG20a, SLXF24]. compact-polynomial [KRV24]. compactly [MH22]. Comparative [EY23, CK22, CK24a]. compare [HM22]. Comparing [AVLMRMGG23].

Comparison

[BEG^{+}22, Bra21, BPR20a, BPR21, LZ20c, TFR^{+}20, VS23, DBSS^{+}21, GJ21, KHAW21, MAA^{+}22, PH20b, PLS22, RVX20, VdG20]. comparisons [BABR20, ZYZ23]. compartment [OOO22]. compartments [BYE^{+}23]. Compensated [ERB^{+}22]. competent [Ray23]. competing [BSNL19, BCMP23, BC24, Bra21, CCCMH24, DNK23, QG20, WTL20]. competition [AMZ24b, AMZ24a, AAk20, Gök23, MWW21, MBKK24, ML23]. competitive [BB20a, FPLZ21, YCX20]. complement [LZ24b]. complementarity [BFVY23, Fan22c, HWCJ21, HV23, HC23, gHtSjT24, KMZ21, KJO23, LDY22, LWL22, LW23a, MN21, SD24, WL24a, Wu22, YSFdRdP20, YYZB24, ZLV20]. Complementary [ACFM21, MKS21, XH20a]. complements [SSB20]. Complete [LD21, WYTL19, ZTY^{+}24, BC22b, CN24, JZ24, Kim20]. completely [PL20, Zhu20a]. completion [ABBB23, KCL20, MHW23, SWN23, WLC23, XWXX22]. Complex [IS23, NEYZ22, WMF20, AHGG20, AUAA20, AM22b, BCKSV24, CTGV20, CCETZ21, DL20c, FSR24, GS22, HSB24, Hua21, IMT23, KSPC22, Kuz22, LZL^{+}23, LW23a, NZ21, Nov24, PGG20, PT24, RDH21, Rum20, TG21, WWWY23, Wan24, XWZ24, ZZWS20, ZZwS21, ZDW24, ZDF^{+}20]. complexes [AHGG20, WWYZ21]. Complexity [PL20, CKY23, GBA21, RSS24]. compliance [PW23, WC21]. component [DG21, HYY21, HMWY23, IIS^{+}21, LLX^{+}22a, VSP23, WZZT20]. component-by-component [DG21]. components [AKN24, BABR20, DJ24, FBJ20, HLR22, Jas22, Jas24, KO22, MMMM20, MIdMR23, Tor21]. Composite [BN21, EGG22, FFRS23, LY23b, LH24, SIS^{+}22, WZL^{+}23, YZ24a, ZC21, ZZWS22]. composition [Vab24]. Compositions [CCETZ21]. compound [ASON23, CX24, GX24, LLL22a, TE20, VCNR20]. Compressed [BSV20]. compressibility [LN23]. compressible [AJR22, Hur22, LHL^{+}22, PFG24, SRAG23, WLH23, YCH24, YLS20, ZW21, ZGG23]. compression [BCP20, LM24]. compressive [JMA21, LWLW23]. comprising [LYL23]. Comput [AGMZ22a, AA20a, AA21b, AA21c, BPR21, BKMO20, FAVM20, GAA21, KZ24, LLG^{+}22, MR25, YHC20]. Computation [APR21, BJMR20, GHM21, IMMS21, KKKL22, Rod21, SL21, Uga24, AHPS22, Alj22, BV22, BEL23a, BWS20, BFRV21, CST20, CPS20, DS21, DDKD19, FV24, JdlCD20, JCLC20, KWZX23, Miy21, Miy22, NV20, RL23, Sad23, SI23, TSMZ20, VS23, VOJS24, VNA22, WWWY23, XLKL23, YL21b, YT19, YTT23, dlCYS22]. Computational [BGR20, GIV20, Sug20, TE24, Vab24, VAAR22, YHF22, YO20, ZZW^{+}24b, ANM24, AID24a, CMJN22, DFH21, DS22a, FS21, KLH21, PH20b, SMDVA22, TVAS23, YCCZ24, Zhu21]. computationally [HO22]. Computations [BSZ21, CUN24, DOP21, IMT20, IMT23, MPR24, RFF23]. compute [KA21, MGL20, PV22, SZL22]. computed [ACFM21, JMA21]. Computer [CW21a, HP22, IMMS21, IY20, PH20b, AS23, DH20, MHY17, MHY21, RLGGAV21]. Computer-assisted [CW21a, IY20]. Computing [AV21, BS23, BD20, BLV20, CZ24, GZ24, Guo24, Kuz22, MR20, Mia20, SN21b, SSB20, VS21, Yan22, BCR22, BPD20, CZ23a, CKL20, COT21, DIP^{+}22, DS22a, GN24, HWH^{+}21, Ibr20, JWSG22, LMAV24, LL24f, MDW23, MSS20, MT20b, PDBdS24, Rou22, RMM22, TZWG21, WHY^{+}21, WTP^{+}21, ZW24]. comrade [Mar19]. Comtet [Ade22]. concave [TM24]. concavity [BCLS24]. concentrated [GH23]. concentration [BEM^{+}24].

Concept

[Hri22, AJR22, CDMM22, EHV22, ZZ24]. concerning [KUA20]. concomitant [Ozk21]. concomitants [HBMA22]. concrete [TZBK21]. condensers [NV20]. Condition [WM20, ALNK24, BBAO23, CHLY21, CLX21, CDD20, CNYC23, DL21, DHF23, Ela23, Kim20, KAA22, LD21, MDY20, SL22, Sun20, WLHJ20, WX22, YJZ23, ZOK21]. Conditional [ZJ24, BBBK24, HG23, LD22, Wu24]. conditionally [MDE24]. Conditioning [HKL21]. conditions [AK21, BK21a, BGN^{+}20, CLL21, CHH20, CD20b, CRRB21, EOR21, FP24, GS21b, GO24, HB23, JO24b, Kar24, Kim22, LMM22, LLG24, MMMM20, MMM22, MMM23, MP24, MBZ23, MFA21, NS20, PJA24, PP24, PFG24, RR21, SS24a, SH24a, SLZL23, UL23, WZ20b, YSN24, ZR24b, ZGG23]. conducting [TSA23]. conduction [BM24a, EGM23, SDC20, SDLZ23, YCH^{+}21, YLS23, ZHWD20, ZC21]. conductive [WCL22]. conductivity [LW23b, NM24, ZPW^{+}24]. cone [HWCJ21, KMZ21]. cones [HL22c, YJZ23]. Conference [BSZ21]. confidence [WTWZ20]. Confluent [OVGI20, ÖÇ23]. conformable [CW24, DSCZ24, DXWC22, YT24]. Conformal [HNV23, Cho24a, CCM22, OSK20, ONT21]. conforming [DZZ24, HH21a, MVY19, MMF24, MDKG20, RS23, SC23b, WGC22, ZZW^{+}24a, ZR23]. congruences [AMRP22]. conical [ÇÖÖ23]. conjugacy [FP21]. Conjugate [IBW^{+}20, ACK23, CKY23, DAMA22, JYYL23, KAM^{+}23, LDC20, LH20, SWN23, WAH22, WSL^{+}23, ZJBY23, ZYL20, ZYZ24, SvG22]. conjugations [ARF23]. connected [DMST21]. connections [MM22a]. consecutive [Kon22, KT24]. Consensus [KLZ^{+}24, AKMO21]. Conservation [MGA24, AAP24, CCL23, CW24, DD23, Kri21, KS22c, MN22, RHM22, RK23, YKBH23, YF22, ZCY^{+}20]. conservative [CD22, HLA22, Lee20, LDM23, PH21, SRMD22, VV23, VP24, WW23, WSHW22, WZD^{+}24]. conserved [CZY22, LL20a, LLT^{+}21, WCLY23, WZY23]. conserves [MDSRFO24]. conserving [CQZ22]. considerations [ORS23]. considering [LLK23, NMF^{+}22]. Consistency [LRZ23, ZY24, BB20a]. Consistent [XJK19, BA20, BS20, HLA22, KH22, LS23, MD20, OGR22, PZL23, SL21, SW21a, YCX20, ZGL20, ZWKS24]. consisting [Jas22, Jas24, SCY23]. constant [BV22, BE21, ER21, KWYN23, KRT23, SCK22, Wan21a, ZJBY23]. constant-stress [ER21]. constants [BCKSV24]. Constrained [BBBK24, ANN21, Bha21, CLMO23, CLHL20, DMNO24, DV21a, GGO21, HA21a, HLLZ24, JHW21, JLY^{+}21, JMLX23, JG21a, JH24, Kar24, KCL20, LMM22, LWLW23, LWG24, LL24d, MER20, RYH24, SZZ21, SC23b, SL24a, SH24a, WSLW23, WPX23, WA21, WLZY21, WZC22b, WSL^{+}23, XLZ20b, hYbJzJ21, YTT23, YW23, ZG22, dSFSS24, dOdO23]. Constraint [CCL20, CN24, ELTD23, LMM22, LLLD20, LWZF21, LLZ24a, TLY23, WWS^{+}21, WC21, WAH22, ZGJ21]. constraints [BFVY23, Bre22, CSZP20, CLH22, FGPP23, GT23, JDBT22, LXQ21, LZWL23, LTZ24, MM22a, MYLL24, MG22, ZL22, ZYL20]. construct [OGR22, RK23]. constructed [DG21]. Constructing [JYWZ21, Lu24, EWCQ23]. Construction [CGSVBB^{+}24, CA23, FLMR20, GJ21, HF24, IpAmK20, JW24, KPS24b, LZW^{+}22, PJA24, WZ20b, BEL22, FNH20, Kra22, LZW20, Maj23b, MHY17, MHY21, Ri22, Ri24, RB24, RR23b, eBI23]. Constructive [KMN20]. Consumer [WZ23c]. consumption [GGT24, JO24a, MYLL24, MO20]. contact [CHLR22, DV21a, Grz21, HJO20, JLM24, LR22, LX24, NT23, PW23, RH23, Slo21, WC21, XL23]. contagion [KH22]. containing [ZuIH20]. contaminant [Kap21]. contaminated [BR20b, NN24]. contamination [ADKH20]. context [MSBSM23]. contingent [AZ22, SH24a]. Continuation [JS21, KS24, AK21, Mil22, WT22, WLC23]. continuity [LLX^{+}22b, LP23b]. Continuous [EL23, CBJ20, CCM22, DS21, FO22, JCY20, Juk20, KWI24, KPD23, MYC22, OK23, SSL24, WZLT23, YZ21a, YH22, ZFwC20, ZSL22, ZSL23]. continuous-installment [JCY20]. continuous-time [MYC22, WZLT23, ZFwC20]. Continuum [BZ20, KA24, PH20a, PCMH20, PCM21, SV24, TK24b, ZZWS20, ZC22]. contour [HLLW21, IMT20]. contoured [RHA20]. contraction [CK24a, GHM21, Luc23, PV22]. contraction-like [Luc23]. contractions [JZ24]. contractive [DL20a, DH24a, YTA22]. contractor [RKA20]. contractor-based [RKA20]. contracts [BD22]. contrast [DBR24, GCV20, KK20b, WYC24, ZSLG21, ZC20b]. contribution [LLW21, ZLL20a]. control [AGH20, ASON23, AK20, ACK21, AA23, AK22a, AA24c, AS23, BBTW24, BB20b, Bre22, CFS21, CCJ20, CLHL20, CLH22, CNQRR22, DMP23, EN20, EY23, FGM21, FLT20, FMAV23, FPR^{+}24, FK23, GYZ22, HÇYK24, He23, HS24b, JYWZ21, Juh21, Kar24, KAD22b, KAM^{+}23, LS22a, LMYZ21, Li23, LX24, LLZ24a, LWT24, LCH23, LW24, LF21, LTZ24, LWX24, MP23, MVB20, Mam23, MÖ21, MS22, MMMM20, Moi22, NMST21, PA24a, QGW21, RZA^{+}24, Ros23, RK23, SCHS24, SS20a, SVK20, SGY22, SS23c, SZA24, TK24a, TTK23, TY23, WZZ21a, WT22, WL24b, WWZ24, WZC22b, YKAA21, YSFdRdP20, YT19, YW23, Yüc21, Zha20a, ZG22, LS24]. control-dependent [YW23]. control-LSD [LS24]. control-systems [Moi22]. Controllability [JVAD22, KMD21, RVS^{+}22, YFW20]. controlled [FGP24, Hun21, Hun22]. controller [DTM20, Hu22, TDD23]. Controlling [FMAV23]. controls [EL23, SZK20]. Convection [KL23a, ABHK20, CWW23a, CUN24, CSS24, DO21, FPS22, FP24, FJ21, GLYS21, GS21b, GO24, JKW23, Kap21, KSVA22, LYZ24, NMMH23, OMRdQ24, SMM22, TY23, WZ21, WLM24, YCCZ24, ZC24, ZXC22]. convection-dominated [CUN24, DO21, OMRdQ24]. convective [CE24, CGO24, ZR24a]. Convergence [AJR22, AET23, BCT22a, BCV24, BLRZ23, CGS20, CG20, CJL^{+}24a, CQZ22, CHP21, CHS20, FP24, GGHY24, HM21a, HYW20, HQS20, HZZ24, HTTO20, JMLX23, JL24a, JK21b, gK21a, LMZ22, LLLL20, MW20, MYC22, MS20a, MKN20, MG22, OKW23, SGY22, SSL24, Six20, WWW^{+}24, WLZY21, XZL20, XXH^{+}22, XDX24, YH22, YY24a, YDC20, ZGJ21, ZSL22, ZVS21, Ais23, AAB^{+}20, AF23, AZL21, AAA^{+}23, BA22, BGG^{+}21, BC22b, Bic23, Bre23, CLL21, CGW21, CG23, DRR22, DDF23, DC24, DG22b, DWWZ22, EL22, EHVM22, EHV22, EHVMM24, Fan22c, FC23, FC22, GRVZ23, GPHHA23, Han24, HL21c, HN22b, gKZ21, gK23, KWYN23, KPS21, LGC23, LY20, LHHW21, LWW^{+}22b, LN21, MGG^{+}24, MM22b, MH23, Min20, MAA^{+}22, MS24b, PI24, RG20b, Saâ20, SRBK23, SL24a, Sou22, TV22, WCPW20, WLBG22, Wen21a, XD23, YTMK19, YYY21, YZZ22, YZ21d, ZXK23, ZSL23, ZG24]. convergence [ZYZ24, Zho22, ZTY^{+}24]. Convergent [EO24, Ais20, BM21, BCVV20, DHQZ24, HSM20, KS21, LPP23, LC24b, MDSRFO24, SLW21, SMDVA22, YZ21a, YTD23, ZZZL21, dOdO23]. converging [TDL^{+}22]. conversion [WFZ22]. converter [LLL24b]. Convex [FYL^{+}24, LLY21, AORS22, DOP23, DL20a, DH24a, FLT20, FHC22, GRS20, GGO21, GHTC21, HLLZ24, HW20, JZ24, KJO23, La 22, LYK23, LT24a, LL23, LZWL23, MS20c, PO21, RYH24, SZZ21, SCY23, SL20, WDS22, WAH22, WQ22, hYbJzJ21, ZL22, ZSHD24, ZYL20, FNH20]. conveying [FFT20]. convolution [AL16, ESS21, EEE21, FL23, Li20b, Yak21]. convolution-type [EEE21]. convolutions [CX24]. Conway [MÖ21]. cooling [CWLM22, KPD23]. cooperate [TDD23]. Cooperation [Ret24, ML23]. cooperative [ARK20, RKA20]. coordinate [Bal24, KDOT24]. coordinate-free [Bal24]. coordinates [Sur21, eBI23]. copolymer [LLZY22]. copolymers [WZY23, ZCYP20]. copositive [MN21]. copositiveness [WCWZ20]. copula [BABR20, FBJ20, SPKS^{+}24, SHZ20, ZZ21, Zhu22]. copulas [BKMO19, BKMO20, BKMO21, KO22, TFM24]. Cordes [BK21b]. cordial [SYD21]. core [Gra20, ZW20, ZCT21, FLMP24]. core-EP [ZCT21, FLMP24]. cores [Gra20]. corner [JYY21, RR23b]. corner-cutting [JYY21]. coronary [WAAM20]. corporate [HC20a]. corrected [Mag20, ZMZ^{+}24]. correction [CC21a, DYF24, EKÇ20, LZ20a, LWA^{+}22, RJ22, Sha21, WSC23, WZ24a, YX20, YY24a, ZWZ^{+}23a]. Corrector [GM24, ABP24, HVR22a, ZJGZ20]. Correctors [GJNR21]. correlated [KMT20, MCY23, RHA20, SPS21, YCD^{+}20, ZZM20]. correlation [ARMP20, Bon24, LWC20]. Correspondence [AJ21]. Corrigendum [AGMZ22a, AA20a, AA21b, BPR21, FAVM20, GMSR21, LLG^{+}22, ZZW^{+}24b]. corrosive [GJDL20]. cosine [DIP^{+}22]. cost [Bre22, JLL21, LMYZ21, SA24, WWS^{+}21, YTT23]. cost-benefit [JLL21]. cost-constrained [YTT23]. costly [YT19]. costs [JO24a, TBA22, WZ23c, YL21a, dFG21]. Cotes [Gol24]. Coulomb [CFS21, CHLR22, NT23]. count [MÖ21, Vie21]. countably [BM21]. counterpart [PT20]. counterparty [VLR23]. counts [FPGSR21, SLK24]. Coupled [CK22, CK24a, CLM20, CLM21, LP24, PHV24, ANR20, AGS^{+}23, AZL21, ACPV24, CC24, CRW20, CR23, CLM22b, CSS24, ESS21, FSC24, FJ24, HR23, Hu24, IS23, LLZC21, LLZY22, LWA^{+}22, LWT24, LCGW24, LM21a, LR24a, LR24b, MBKK24, Ngo22b, RSRW24, Sun20, TSG22, WZ23b, WW23, WSHW22, YOCY22, ZR23, ZLC24, dPdMQL23]. coupled-wave [CLM22b]. Coupling [FdOPS21, NA21, Sha23, CPBG23, CD22, ELTD23, YO20, YLLA23, ZR24a, ZG24]. coupon [MN20]. Cournot [Yu22]. CoV [BYE^{+}23, HCCSCGAM24, ÖHS23]. covariance [AKN24, NRGJ21, OR21, PT20, Psa24, Zha20c, ZY22]. covariate [MFH24]. covariate-based [MFH24]. covariates [KMT20]. coverage [Sal21a]. covered [TZBK21]. Covid [Ngo22a, OOO22, SMR23, YBAE23]. COVID-19 [OOO22, SMR23, YBAE23]. Covid-2019 [Ngo22a]. Cowell [YCXW24]. Cox [KLM22]. CP [WLCH23]. crack [JO24b, Kov22]. cracks [TSA23]. Cramér [LLL22a]. Crank [ADZ20, BSdCCGV23, BDF21, FC22, GH22, JF22, LW24, Ngo22b, NA21, WY20, YYBL24]. CRBC [Kim22]. Credit [ABO24, CZ20, DLL21, KH22, KH24, PLZ21, VLR23]. crime [AMT23]. criteria [ANR20, AB23, SLZD20, ZD21, ZBH24, dHSR22]. criterion [DTM21, Juk20, TBA22, TBA24]. Critical [BCM20, MSBSM23, SBKR24, WZZT20]. cross [BR20b, Dax20, DFK20, FHM23, KZM21, MJ20]. cross-interface [FHM23]. cross-product [Dax20]. cross-sections [DFK20]. cross-validation [BR20b]. crossed [NMF^{+}22]. crosswind [KLS21].

Crouzeix [Nud24, ZOK21]. crude [ZZ21]. cryptocurrencies [PGG20]. crystal [HYY21, HLLW21, LML21, LLYM21, LYL23, QH23, ZLS23, ZaZWZ23, WC24a]. crystalline [BP22a]. crystals [LLL^{+}22b, YLK24]. CT [HJS21, RV20]. CT-scanning [RV20]. cubature [BSV20, CDD^{+}24, FMRS24, For24, MJK20, TFR22]. Cubed [BBC22, Bel23b, BC23]. Cubic [CTN22, MA19, SLZL23, BEL23a, BK24, CZ23a, CVPM22, GS21d, MM20a, RG22, SKVA23, WZH^{+}21]. cubic-quintic-septic-nonic [BK24]. cumulants [LZ22]. cumulative [TA22, YO20, ZG20b]. CUR [ACG20]. cure [YCD^{+}20]. curl [CWW23b, HZ24b, TXZ24, WWZ23, YZZ20b]. currency [Fan22a, YWTL19]. Current [DHJ^{+}24]. currents [dL23a]. curse [FMFM20]. Curtis [KEKT21]. Curvature [CFJVG21, ZJBY23]. Curvatures [ZB22]. curve [BPD20, GVT21, Juh21, PDBdS24, Zho22]. curved [LWZ24, Mu21, YLK24, ZOK21]. curves [AH23, AHPS22, BAIB24, BL21, BLV22, CFJVG21, FPS21, GÇSA23, HY21a, HF24, HM22, KS24, KLMP21, LZW^{+}22, MSK21, OMRdQ24, RB24, RFB22, SS23a, WSYPD24, WY21b]. curvilinear [LMLB22, RR21]. customer [FPLZ21]. customers [BS24a, CJL^{+}22]. cut [AK22a, SS24b]. CutFE [ZDW24]. cutting [JYY21, ZLL^{+}20b]. cutting position [dL23a]. cutting tolerance [dL23a]. cycle [BCH20, PA24a]. cycles [LP20]. cyclicity [BL20]. cylindrical [BEH^{+}23]. cytotoxic [TKM21].

D [BM24a, CUN24, CMS20b, CA23, CJ23, CSS24, DL21, DOT23, DK24, DC20b, DT24, DFK20, FFS24, FGPP23, GPHHA23, HW21a, HL24a, HLLW21, LL24c, LLGC24b, LYZ24, Mia24, Qiu20, RDG23, SESH23, SGD20, VV23, WDY20, XLLA21, YZ24d, ZE23]. D-linear [GPHHA23]. Daehee [JKKK20]. DAEs [GT23, SL21]. dagger [FLMP24]. Dai [LH20]. daily [DKM23a]. dairy [BYWA22]. dam [GCV20]. damage [HJO20, Khl21, Let24]. damped [BCH20, BFF22, CNQRR22, PKE20, WC24b]. damping [CV24, JVAD22, JSRdS23]. Darbo [SuR20]. Darboux [SH23]. Darcy [ACS22, ACPV24, BMS21, CJCDLL22, CHLY21, CC24, CRW20, CR23, CWLM22, CO19, CRRB21, GLHL24, HCCC21, JF22, JQ22, LZWC21, LLZC21, LHL^{+}22, LWA^{+}22, LR24a, LR24b, Ngo22b, PHV24, QHPL21, SV24, TGJY22, WWH21b, WLH23, XH20b]. Darcy-transport [LR24a]. Data [AGNGG^{+}23, FS24a, ZZWS22, dLNO21, ABBB23, AS20a, AFJGJJ22, AUAA20, ASTK24, BC24, BVBO23, BL21, BEH^{+}23, CZ20, CDD^{+}24, CZM23, CYY23, DDUF22, Din23, DS21, DLL21, DNK23, EK24a, EKCF22, FGL23, FS20, FGM21, GMW24, GANR22, GO20, GS24, GHL21, HL23, HZ24a, HSB24, HJC24, HSW21, HG22, HLR22, JHZD20, KH21, KNOR20, LWC20, LMSZ21, LYB22, LWLD23, LCGW24, MÖ21, MCY23, MJ24, NK22, NN24, OBO23, ÖHS23, QX24, QG20, Seg21, SLP20, SACQ21, SW20, TVB21, TZTC20, TBT^{+}20, Uça23, VCNR20, VTB21, VPHE23, WK20b, WLJ24, XL20, YLL20, YCD^{+}20, ZG20b, ZS20a, ZMYL23, Zha20c, ZZM20, ZFH20, ZMZ^{+}24, Zhu20b]. Data-based [FS24a]. Data-driven [AGNGG^{+}23, ZZWS22, dLNO21]. data-mining [LWC20]. datasets [AMRP22, FGPP23, SKZ23, WXZ21]. DB [AZZ23]. DBH [LPLP21]. DBH-VNS [LPLP21]. dc [LLL24b, BTC21, HTD21, LXQ21, PZL23]. DCA [HTD21]. DCA-based [HTD21]. DCCA [LWZF21]. DCT [PL20]. de-blurring [JH24]. DEA [AJSGL22]. death [Gök23, ZYY20]. deautoconvolution [FS24b]. debt [ZCJL20]. decay [CJP20, MQ20]. decaying [HS24a, OSK20]. December [Ano20a, Ano20b, Ano21b, Ano21h, Ano22b, Ano22h, Ano23b, Ano23h, Ano24b, Ano24g]. Decision [PCT23, AVRLRMGG23, ASTK24, CGCMRP24, YC20a, dHSR22]. decline [WMF20]. decoder [NY24a]. decomposition [Ais23, Ars22, Axe20, BPR20b, CMZ22, CD20b, DOP21, DLC21, FYL^{+}24, FS20, FLMP24, GANR22, Gra20, HLLL23, JMHZ24, KK23, LYB22, LXLL23, MkCZ20, MQZ^{+}24, OA20, PK24, RKS^{+}22, SS23a, SLXF24, Shi21b, SPC21, UL23, Vab24, WLCH23, WJVG24, WW24a, WLDZ24, XA21, XTL23, YLXY23, YS24a, YC20b]. Decompositions [DDRS24, BAD^{+}22, DKM^{+}23b]. deconvolution [DRR^{+}21, LLLY21]. Decoupled [CHLY21, DYM21, JF22, MBKK24, WCLY23, ZPW^{+}24, DHH24, EKÇ20, GLHL24, Hu24, LLZY22, XYAH23, YY24a, YOCY22, ZHY20, ZSH20, ZaZWZ23]. Decoupling [UDI20, ZIZ21]. decreasing [ONT21]. dedicated [BSZ21]. Deep [FPR^{+}24, KPS24a, MR25, NY24a, XFW24, CLPZ21, Fan22a, GS20b, HHM22b, HLLL23, HQL21, KA24, Kap21, LXLL23, SSM23, XLZ20a, ZLZ^{+}24, MR23]. Default [YHSK22, ABO24, PLZ21, WRZW21, ZGL20]. defect [DYF24, EN20, Sha21]. defect-correction [Sha21]. defective [FPGSR21]. deferred [CC21a, EKÇ20, WSC23, YX20, ZWZ^{+}23a]. deficient [Bic21, BC22b, CGMM20]. defined [BWS21, LLW21, Liu20, Pet20, ZLL20a]. defined-contribution [LLW21]. definite [KH21, Liu20, LNP22, MDE24, WL22]. definiteness [LZLL20, LM21b]. Definitions [HL22a]. Deflated [SvG22]. deflating [SK24b]. Deflation [VNV23, CDW20]. deformation [MBG21b, TT20]. deformations [CFO22]. deforming [RCN21]. defuzzification [LS24]. degeneracy [RR22d]. degenerate [JKKK20, KSVA22, YVXX24, YLD19]. degenerations [MMM23]. deGennes [DWWZ22]. degradation [EKCF22, WWS^{+}21, ZS20a]. degree [AHPS22, BC21, BC22a, BL20, BWS20, FPS21, FKV21, LZ24b]. degrees [RSS24]. DEK [BD24]. DEK-type [BD24]. delaminated [KP20]. Delay [HCH20, HH21c, AAHK24, AA22, AA24b, AA24a, ACC23, Ask21, BJB22, BR22, CLL21, Cso20, DS20, DFFM21, DPW24, Ela23, EM21a, FZ21, GM22, GGHY24, HMA24, HAD24, Hu22, JMX24, KCKK24, KS21, LXYL24, Mam23, Ois20, WSY20, YZ20a, YQZS22, YCXW24, YFW20, YZ24c, YZ21d, ZVT^{+}23, ZQL24]. Delay-dependent [HCH20]. delay-differential [YQZS22]. delay-differential-algebraic [YCXW24]. delay-integro-differential [BR22]. delayed [AORS22, AB23, LDQF23, RCVA24, WMF20, WZZT20, YFW20, YW23]. delays [AUAA20, AID24b, CCJ20, HH20, SLZD20, WSY20]. delivery [FPS22, FGP24]. delta [ZSVL20, ACD22]. demand [GXL22, LLK23]. demosaicking [GJM^{+}23]. Dempster [HH23]. denoiser [HWG21]. denoising [FGL23, GJM^{+}23, HL22b, Shi21a, WSL^{+}23, XTL23, YLZ^{+}22, dLNO21]. dense [MORO20]. densities [Bra21, LL24d, LZ22, LSZ23]. Density [CMP21, HW22, SY23, Bra21, BCVV20, CMS20a, DHJ^{+}24, HCC24, HH24a, JCLC20, Kap21, LL24c, LN23, TFM24, WCLY23, WZNT24, WN23, YO20, YL21c]. density-suppressed [HH24a]. dependence [FBJ20, GMSR18, GMSR21, Kan24, LSY24, MCW24, SHZ20, Tor21, ZZ21]. dependency [KV20]. dependent [BSNL19, BLBH21, BZHB23, BABR20, BB21b, BCP20, CLL20a, CL24a, CTGV20, CFO22, DDKD19, FC23, FC22, GLYS21, GH21, HH21a, HJM24, HCH20, HH21c, JZ22, JF22, JLZ22, JLW23, KCKK24, LS23, LDM23, LRQV24, LR24a, MKN20, MSK21, MFH24, QS23, RZ20, RMM22, SA24, SN23, Sha23, SYG19, TN24, TL21, WL21, WW22, WWC23, WSZ20, WX22, WFZ22, XW21, bYW23, YL20, YW23, YY24b, ZZK21, ZGZ23, ZMYL23, ZYZ23, ZTY^{+}24, Zhu22]. depending [BMS21]. deposits [ÜY20]. depth [SZL22]. derivates [GPHAPR21]. Derivation [BK21a]. Derivative [AF20, ABA19, AZD21, BCT22b, Bic21, BKM22, BHJB21, CHXL20, Evi23, Fan22b, GW23, GS21c, HZZ21, HHL20, IAH24, JBSN22, LLLL20, MY21, MKN20, MDS20, MFA21, MP21, NRV21, RDH21, SESH23, SYTD24, TZ20, Uça23, Vab23, WK20b, WL21, WX22]. derivative-free [BKM22, IAH24, MP21]. derivatives [AH22, FB24, HY21a, IMMS21, JK21b, LL22, MX20, RR22a, Rod21, Ruj23, SNMK21, SH24a, WLL21, YLWW21, YT24]. derived [CJLL21]. descent [LDC20, MG23, RJ22, RPQ24, SWN23, WY24b]. described [BK24]. describing [UÖA23]. descriptor [VCNR20]. Design [HLLY24, MD22, MDG22, AM22a, BSV20, DSCZ24, ER20a, ER21, Hu22, JMA21, JZM^{+}23, KLM21, KKO23, KCK24, OPK23, RLGGAV21, WWS^{+}21, ZGG23]. Designing [RZA^{+}24, Els21b]. designs [ÇK24b, EWCQ23, WWWY23]. desorption [GIV20]. destriping [YZM^{+}20]. destruction [ÁGKM21]. destructive [BJP24, BC24]. Detachable [JMHZ24]. detailed [SESH23]. detect [GÇSA23]. detecting [CNRR20, HBLZ21]. Detection [Tam20, AM22a, BVBO23, DSCZ24, GRS20, LGLC20, ŚK23b, WCL22, ZLL^{+}20b]. determinant [OF20, Rum20]. Determination [SIS^{+}22, dL23a, CNYC23, SL21]. determine [ASX21]. determined [BA20, BM22a, BS24b]. Determining [NW23, WX22, PKPD21]. deterministic [DRD^{+}23, HRT20, SACQ21]. developable [BRLS23]. Developing [KD22]. Development [KRV24, MBG21a, YZZ22, SZWH21, WHC^{+}20, Yam23]. deviation [MER20, WCLW20]. device [BCMP23, BJP24, BC24, HLLY24, LLGC24b, YLS23]. devices [DRD^{+}23]. DEWMA [HZS^{+}24]. DFPM [YJJW23]. DFPM-based [YJJW23]. DG [FJ21, ZZW^{+}24b, EGG22, HH21a, HHD23, OBB23, WZZ21b, YY24b, ZZW^{+}24c]. Diabetes [AVLMRMGG23]. diagnosis [SL21]. diagnostic [AVRLRMGG23]. diagonal [GJGDPR23, HF24, HH24c, KCS21, LZW^{+}22, LHG^{+}24, ZYZ22]. diagonal-plus-Toeplitz [ZYZ22]. diagonalization [LNP22]. diagonally [GS20c, LZ24b]. diagonals [AM23]. DiagPara [LHG^{+}24]. diagram [RLGGAV21]. diagrams [KPD23, Yam21b]. Diblock [LLZY22, ZCYP20]. dielectric [BFQ22]. diffeomorphic [HW21a]. Difference [Abi20, BGG^{+}21, HH24a, WL21, AA21a, AAHK24, AA22, AA24b, BK22, CZHQ23, CLYS24, CC20, Dar20, DL21, DOT23, DVRV22, DCAA22, FHM23, GL24, GS21c, HL24a, HE24, HWRM24, Hor20, HL21b, HSYZ24, JT20, JLZ22, KSPC22, KCRVA22, LMZ22, LHL^{+}22, LLS22, LLLL20, LZ20c, MDSRFO24, NMMH23, Pat23, RG20a, SM22a, SRMD22, Son24, Sou22, UGG^{+}20, WW23, WY21a, Wol24, XWXZ22, XYAH23, YQZS21, YKBH23, YSBT20, YMC22, YLS23, ZVT^{+}23, ZZwS21, ZY24]. difference/Galerkin [ZVT^{+}23]. difference/local [WY21a]. differences [BL23, CDJ20, CPP23, LGMC20, LLG^{+}22]. different [ABA19, AAA^{+}23, AGGAV^{+}21, BB21a, ÇK24b, DDF24, FY22, HZS^{+}24, Jas24, MLS22, OPK23, TTLB22, VS23]. differentiability [RPQ24]. differentiable [CTN22, MS20c, MVFAMF21]. Differential [APW^{+}21, ADZ20, ASML23, AEST21, AM21, AAKA24, AS21a, ALK20, AN24, AZD21, ASA^{+}21, AA24a, AS21b, ACC23, AHK^{+}23, Ars22, Ask21, AA20a, AA20b, AA21b, AA21c, AI22b, BJB20, BTBR21, BJB22, BR22, BL20, BR20a, BuR24, BCVV20, CLL21, Cap20, CCETZ21, CF20, CHH20, CW21b, CW24, CKL23, CST19, Dar20, DRR22, DTM20, DTM21, DOR21, DO22, DdO22, DFFM21, DPW24, DHKV20, DKMT24, DAM23, DCAA22, ELTD23, EL23, Ela23, ER20b, FM21a, FZ21, FM22, FJ24, FLZ21, GGRCGG^{+}23, GM22, GJN20, GLCL20, GLHJ21, HJ20, HBGS21, HM21a, HM21b, HMA24, HWRM24, HCH20, HWD^{+}24, IpAmK20, JWG21, JL24a, JLM24, JCLC20, JVAD22, KCKK24, KKA20, KRT23, KK20c, KK21, KPS21, KMD21, Kum22, KS21, KRV24, LY20, Li20a, LZL21, LHHW23, LVW24, LLW22, Liu20, LG22, LY23b]. differential [LWZ21, LXYL24, MW20, MMJ21, MMJM22, MA19, MJK20, NNJ24, OKW23, ÖKG23, Pat23, PL24a, PNS21, RO20, RS24a, RS24b, RB19, RB23a, RGA20, RCM20, RG20a, RWW22, SuR23, SES20, SESH23, SM22a, SuR20, SSL24, SSL22, SW20, SYD21, SZA24, SB24b, SKVA22, SDW24b, TZBK21, TSG22, TM24, TS20, WEEA20, WXS21, WKC20, WSY20, WH24, bWLgZbX24, XFW24, Yam21b, YQZS22, YCXW24, YWKM19, YYY21, YZ21a, YH22, YTD23, YSFdRdP20, YZM23, YFW20, YZ24c, YZ21d, ZvS23, ZM20, ZG20a, ZLG20, ZGJ21, Zha21c, ZSL22, ZB22, ZSL23, ZCZ24, ZS22, ZHZ24, ZGL20, ZVS21, ZTS21]. differential-algebraic [SB24b]. differentiated [XZP22]. differentiation [ALK20, RDH21, SL21, Tia22]. differentiator [WLB^{+}21]. diffuse [AET23]. diffusion [AHB20, AHB21, AH22, AAHK24, AJ21, AGS24, AAM20, AVL22, BZ20, BB21a, BB21b, BGG^{+}21, BMV23, Boe20, BitH21, BGN^{+}20, CWW23a, CL24a, CX24, CG20, CJ23, CSS24, CPP23, CS23, DWWZ20, DZ22, DO21, DFFM21, EGG23, FSC24, FJ24, FLT20, FZQ22, FP24, FJ21, GJNR21, GZ23, GLYS21, GGVRB22, GM24, GS21b, GO20, GO24, HL21a, Hai23, HL24a, HW20, HS20, HKL21, HSC21, JWZ21, JHG^{+}20, JL24b, JKW23, KK22, Kar20, KLS21, KK20b, KSVA22, LLFT20, LFL20, LY22a, LG23, LLGC24a, LL20b, Liu22, Liu24, LXY24, LC24a, LT24b, LYZ24, LXF21, LWZ21, MW20, MGD22, MD20, MDVR22, MDG22, MSD21, Mia24, MBKK24, ML23, NRV21, NMMH23, NSL22, PAAI24, PLZ21, PM20b, PLS22, PSS24, Pul22, QX24, RS24a, RFGGM20, RG23, Rou24, RZ20, SM23, SMM22, SMDVA22, SPS21, Sha23]. diffusion [SDLZ23, SY24, SXX21, Shi21a, SC23a, SKVA23, SJZ20, Sun20, SS24b, TH24a, TM24, TY23, TBT^{+}20, TZ20, UV24, VSP23, VAMO24, Wan20b, WMF20, WZ21, WZZ21a, WOV22, WSY22, WLM24, WZZT20, WY21a, WX22, WLW23, WLS24, XA21, XCY24, YZW21, bYW23, YWKM19, YPL20, YLFT21, YLWW21, YH22, YLLA23, YCCZ24, YZ21d, ZVT^{+}23, ZYZ22, ZWY22, ZLJT22, ZGZ23, ZC24, ZF24, ZEW23, ZXC22, ZCZ23, ZW22, Zhu21]. diffusion-wave [AAHK24, EGG23, SM23, SDLZ23, SC23a, WX22, ZWY22, ZLJT22, ZGZ23]. diffusions [LP23a]. diffusive [BM22b, LSY23, WY20]. diffusive-viscous [LSY23]. diffusivity [TY23]. Digital [KA21, DP21c, Fan22a, MGL20]. dimension [Bac19, BT21, CLS20, FGL23, WZLX20]. dimensional [ADZ20, ALK20, BV22, BS20, BRLS23, BDF21, Bra23, CMW24, CS24, CL22, CH20, DWWZ20, DR23a, DDUF22, DDF24, DW20, DW23, DDKD19, EFTW21, ETKF23, FLT20, FDDM24, GZ23, GG24, Gol24, HZ23, HJC24, Hes20, HC20b, HC21, JHZD20, JJ24, KLH21, KM22, Kaz21, KDOT24, KMUR24, KWN20, KV19, LBBB23, LR22, LSE24, LZW20, LL20a, LWYZ21, LZZ24, LHH21, LMS21, LY24b, LXZ23, LLL^{+}22b, MD22, MS20a, MDS20, MH23, MCY23, MHHV20, MQ20, MA21a, Ngo24, OWV24, PJA24, PSWZ21, PSRVA24, PS22b, QX24, RDHA20, RAN20, RB19, Ray23, SM20, SH23, SNMK21, SZK20, SDW24a, SŁK^{+}22, Son24, SL24b, WEEA20, WYH22, Wen21b, Xin21, uDuIZ22, ZZwS21, ZZ22a, ZLJT22, ZC24, ZR23]. dimensionality [FMFM20]. dimensions [BB23, CL24c, HZMD22, HC24, KWI24, MH22, NM24, SZL22, WZ20a, YMC22]. diminishing [FL24, Kri21]. Dirac [FY22, FC22, KSZ21, LWYZ21, SH23, VV23]. Direct [CCJ20, OF22, HM21b, JLW23, LLX^{+}22a, SGD20, ZC24]. directed [KV20]. direction [DVRV22, LH20, PH22, ZL22]. directional [GS21d, LZ24a, XZZ^{+}20]. directions [SRS21]. directly [KS21, LJL24, SSRK23]. Dirichlet [AKN24, CHH20, EFTW21, LG23, MSD21, RR21, Yüc21]. Disc [Yam23]. discontinuities [GO20]. discontinuity [DP23]. Discontinuous [HLLQ22, LMV24, LZWC21, LTZ24, DP23, Bac19, BTBR21, BT21, BGN^{+}20, BLRZ23, CHZ21, CGS20, CRW20, CR23, Che21, CCL20, CM20, CA23, DZZ24, DW20, DW23, DKM^{+}20, FJ21, GEG23, LRZ23, LW23b, LSY23, LY21, LMLB22, LXY24, LC24a, LXZ23, MF21, MTIR22, Moo22, PK24, PW23, RS24b, SDW24b, TXZ24, TXS21, TR22, TGJY22, TCGY24, WLHJ20, WGC22, WSY22, WTX24, WSY20, WY21a, XA21, YX20, YKBH23, YZZ20b, YZM23, ZZW^{+}24a, ZC24, ZaZWZ23, ZXC22, ZLC24]. discount [DMP23]. discounted [SHZ20]. discrepancy [GGPP21, HH21b]. Discrete [CCL23, LS22a, Min20, SLW^{+}22, ZWJ22, AFLP22, AAP24, AGGAV^{+}21, Bel23b, BB20a, BLV22, Bro23, BPR20a, BPR21, BDR23, CZY22, CM22, DKSQ22, DTMJ24, DO22, DG20, DYZ24, EEE21, ETKF23, FPLZ21, FS24b, HLZ20, HCC24, HG22, HOGO23, Hor20, HX22, HPWM20, HLY22, Jia20, JZM^{+}23, KA24, KMUR24, KK24a, KMN20, KSG^{+}22, LZW20, LWD20, LPY20, LS23, LJ23b, Liu24, LNP19, MMMM20, MPV23, MZY23, MB24, NEYZ22, NY20, ORS23, Ret24, RWW22, Ruj23, SN23, SS23c, SKVA23, TGJY22, TTK23, TZTC20, TBT^{+}20, WSY22, WZY23, WLJ24, WZC22b, WLS24, YN21, YBM23, YT19, YTT23, YAE23, YY21, YW23, ZZG20, ZW20, Zha24, ZY20]. Discrete-time [SLW^{+}22, AGGAV^{+}21, FS24b, HPWM20, LPY20, LJ23b, NEYZ22, Ret24, SS23c, TTK23, WZC22b, YY21]. discretisation [FDDM24]. discretisations [BB21a]. discretization [AGST23, AST^{+}23, BS20, BLN21, CHXL20, LLX^{+}22a, LVW24, LLZ24a, LCH23, LW24, LF21, LV20, MDD21, NS24c, NB23, PJA24, DFMK20, Ust23, YZ24d, ZHY20]. discretizations [Abi20, DC23, DT24, FY22, FK23, GT23, HLLQ22, TR22, WWTL24, YVXX24]. discretize [HMWY23]. discretized [CPBG23, DKMT24, TH24a]. Discussions [WQ22]. disease [ÖHS23, PC24, Uça23, WZZT20]. disk [EOT22, For24, KSG^{+}22, SK24b, TFR22]. dislocation [NY24b]. dislocations [EO24]. dismantlable [Luc23]. disparity [NPP^{+}21]. Dispersion [SLA^{+}24, GS20c, RZA^{+}24]. dispersive [EOR21, HLLW21, WZZ21b]. displacement [BEG^{+}22, BEM^{+}24, LHL^{+}22, WZ20a, YCH24, YLS20]. displacements [PV22]. disposal [Kap21]. dissipation [Cho24b, GS20c, SC21, SCK22]. dissipative [AUAA20, Bha21, MD22, YVXX24]. dissipativity [HZMD22, MDSRFO24]. dissolution [GJDL20, GM24]. distance [DH22, Mim20, VZ21]. distances [MLS22]. Distributed [KZM21, LCGW24, ADZ20, CJCDLL22, EGG23, FJMP24, GPST23, GHF24, He23, HS24b, HH20, LLFT20, LFL20, LMSZ21, YZ20a, YLFT21, YWW24, ZLJT22, ZHW20]. distributed-order [ADZ20, EGG23, LLFT20, YLFT21, ZLJT22]. Distribution [MIdMR23, PM20a, ASON23, AAB24, Ask21, BSNL19, BJP24, Bay22, CM24a, CG23, CDS21, CMJN22, DDKD19, ER20a, ER21, EK24a, GDSCO21, GN24, HM22, JB21, JS22b, LLY21, LY24a, MKS21, MIdMR24, Mil22, MPGE24, NP22, NK23, NMB22, NN22, OVGI20, PA21, Pou21, Pro23, QG20, Ruj23, SCHS24, SMKV21, SVJG24, WTWZ20, WDTW20, YCD^{+}20, ZG20b, ZCF^{+}22, ZZWS22, Zhu20b]. distribution-free [HM22, ZZWS22]. distributional [ARMP20]. Distributionally [WW24b]. Distributions [Kon22, AB22, BN21, BSH24, CX24, CECCCR23, DJ22a, DNK23, HM22, HBMA22, LWLD23, LZ20c, NN24, OK23, SZ20, SC24, ZYB21, ZY22]. disturbance [KLK21]. Ditzian [GSZ22]. div [CWW23b, GANR22, GRVZ23, HH21a, Sha24, YZZ20b]. Divergence [CMP21, KK24b, BCMP23, CCL23, CC20, FS24b, GW23, LWW22a, ZTZZ20]. Divergence-free [KK24b]. divergences [AFJGJJ22, HQT20]. dividend [LMYZ21, LCZ19, LLL22a]. Dividends [LCZ19, SZ20]. divisor [BWS20]. DJI [AD21]. DNA [BDPR23]. DNID [DG20]. Do [ÜY20]. Does [HC20a]. Domain [Ars22, UL23, AK22a, Axe20, BWS21, DMST21, DVRV22, DLC21, GHJM24, Gra20, HLLY24, HLYZ24, LXLL23, Li24a, MHHV20, PK24, SLXF24, WW24b, XA21, YWZ24, ZLJT22, ZZ23]. domains [ATO22, BB21a, BA23, CM20, CHP21, DDF24, EOT22, FLT20, GLYS21, GH21, HNV23, HSB24, HQS20, LP24, MDKG20, NYL23, PO21, SM20, SVWC22, TCVC21, WKC20, ZDW24, ZY20, ZX23a]. dominant [CZ24, LZ24b]. dominated [CUN24, CDL22, DO21, KL23a, OMRdQ24]. domination [AN20]. Dormand [NNJ24]. dot [ERB^{+}22]. double [AK20, BS24a, CBF24, GJGDPR23, HS23, HC23, MNH21, TSMZ20, TDAT23, YZ24d]. double-ended [BS24a]. double-layered [GJGDPR23]. double-relaxation [HC23]. double-sampling [CBF24]. doubling [Wen21b]. doubly [PT21, WWC23, XBC22]. Douglas [CW21b]. Dow [AD21]. down [LLL24b]. Downside [Mit20, WZLT23]. Drazin [ZZMK23]. drift [LLGC24a, ZHZ24]. drift-diffusion [LLGC24a]. drive [HC20a, ÜY20]. driven [AMS23, AGNGG^{+}23, BJ22, EPL21, HWG21, JLM24, Kap21, Kov22, LHHW23, MFH24, SPMB21, ZGJ21, ZZWS22, ZHZ24, dLNO21]. droplet [Kog21]. droplets [BBSG23]. Drude [HLYZ24]. drug [FdOPS21, FPS22, FGP24]. DSMC [TFR^{+}20]. DtN [ZXZ24]. Dual [KT24, BJF24, BLRZ23, CWW23b, CWW23a, HCH^{+}23, KA24, KLH21, LRZ23, LWW22a, LWW23a, LQS23, NMST21, SMM22, SCY23, Wan20a, WZ21, WJ23, WSHW22, WZL^{+}23, YCH^{+}21, ZC21, Zha21a, ZC22]. dual-asymptotic [HCH^{+}23]. dual-continuum [KA24]. dual-phase-lag [LQS23]. dual-phase-lagging [ZC21]. dual-porosity [WSHW22]. dual-porosity-Navier [BJF24]. dual-wind [BLRZ23, LRZ23]. duality [LXQ21, PT24]. duct [TSA23]. Duffing [Ois20]. Duffy [XZZ^{+}20]. duopoly [AAk20, WZ23c, Yu22]. durable [YP21]. during [BYE^{+}23, VSSV20, YBAE23]. dwell [YY21]. DY [DAMA22]. Dyadic [LNP22]. Dynamic [AKN24, AAk20, HZC^{+}23, JO24a, LWC20, WZLT23, XZP22, Ais23, BGV20, BB20a, CKL20, CHLR22, CLL20b, DFH21, HCH^{+}23, HP22, JLM24, Lee23, LLS21, LX24, LLG24, MCD24, Men24, MBZ23, MDLY23, OEH23, PCWY20, Ret24, RV20, SN21a, TZBK21, YP21, Yu22, ZAH^{+}24, ZZ21].

Dynamic-model-based [LWC20].

Dynamical

[Gök23, KMUR24, KK24a, LY24b, NAS22, ADMV22, ABM^{+}22, AUAA20, ESS21, EPL21, KV20, PN21b, SPMB21, Zha24]. dynamically [MD20]. Dynamics [AA21a, SAN^{+}24, TKM21, AA23, AVL22, AMT23, BBMK22, BIZ20, COT21, CNQRR22, EO24, Evi23, GeOJD24, Kap21, KKKL22, Kog21, LZL^{+}23, LJL24, LWX24, MARH^{+}23, MRG^{+}21, NEYZ22, OBO23, OBB23, TK24a, WMF20, Wan21a, WT22, WCLY23, WCK^{+}22, WFZ22, YZ24d, ZZZ^{+}22, dPdMQL23, NGNB23]. dynamo [CMS20b].

e-MoM [FMRT20]. E3SM [NPP^{+}21]. early [DLL21, HQL21]. earnings [GS20a]. ecological [SZWH21]. econometrics [XL20]. economic [LYL^{+}24, WHC^{+}20, XL20, YS24b]. Economical [KFK^{+}24]. economies [WHC^{+}20]. ecosystem [WQ20]. Eddy [GQR23, MS20b, dL23a]. Edge [Li24a, HYW20, RBR23, SKS22, WZH^{+}21]. edges [Gua20, Mu21]. Editorial [BPR22, KHYL24, Sol24, VAAR22, ZDF^{+}20, Ano19, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano21y, Ano21z, Ano21-27, Ano21-28, Ano21-29, Ano21-30, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano23-29, Ano23-30, Ano23-31, Ano24l, Ano24m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, Ano24u, Ano24v]. Editorial [Ano24w, Ano24x, Ano24y, Ano24z, Ano24-27, Ano24-28]. EEG [LRP24]. Effect [HB23, MCL22, BVBO23, CCGT22, Kap21, LLK23, MP23, ML23, MLS22, NEYZ22, PA24a, PZY^{+}24, PCWY20, Sal21a, VS21, YBM23]. Effective [CMFMSQ22, DDKD19, LYK23, CNYC23, PT21, PDBdS24, SES20, SIS^{+}22, SD24, WZZ20, ZWZ^{+}23a]. Effects [BB21a, LRP24, FRT23, GH21, GH23, KMD21]. Efficiency [Cho24b, RR21, LGMC20, LLG^{+}22, PH20b, WXZ21]. Efficient [AA24c, BPDC20, BP22a, BR20a, BFRV21, BEH^{+}23, GZ23, GHC^{+}20, HR21, Hua21, JdlCD20, KWZX23, KMT20, LLT^{+}21, LML21, LLZY22, LM21a, LHE24, MHW23, Maj23b, MBG21b, PJM21, Rou24, XLL23, XHW24, YOCY22, ZCY^{+}20, ZCYP20, ZG22, ZLS23, ZFM22, ZCZ23, ZH24c, ASA^{+}21, AID24b, BA20, BR22, BBMK22, BFF22, CMS20b, CUUdB24, DWWZ20, DO22, ELTD23, EHVM22, GR22, GJDL20, GW24, HYY21, HSB24, HVR22a, HVR22b, HG23, HAD24, HL24b, HLLW21, KEKT21, KXW23, Lee21, LLYM21, LWA^{+}22, LNP22, LLX^{+}22b, LC23, LH20, MSBSM23, MP21, OOO22, PS24b, QXZG22, QXZF23, RZA^{+}24, RMG21, RG22, RG23, SKE20, SWW23, SW24a, SSRK23, SL24a, Son24, TL20, TLY23, TR22, TH24b, Wör22, WSL^{+}23, Yam21a, YC23, YYT24, ZSY24, ZSHD24]. Efficiently [YLJW24]. efflux [GPK23]. Ehrhart [McA21]. eigen [CZ23b, RL23]. eigen- [RL23]. eigen-representation [CZ23b]. eigenfunction [KBK^{+}22]. eigenfunctions [AN24, Ans21, BV22, LV23]. eigenpair [CZ24]. eigenpairs [FW22, MDW23, Mia20, RL23]. eigenproblems [GCV19, GS21a, IMT20]. eigensolver [HLLW21, IS23]. Eigenvalue [AM22b, BG23, Ais20, BT23, BP22a, BA23, CDCVV22, DC20b, HOOT20, IMT20, IMT23, JLZ23, LRV23, Liu20, LLL^{+}22b, MWBM22, MM22b, MMF24, MW21, Ned22, PT20, RK22, SK24b, TL21, XH20a, XBC21, XXH^{+}22, XBC22, XWL24, YC21, ZDL22, ZJY22, ZHW20]. eigenvalues [Alj22, BV22, CJP20, CZ23a, Dom21, DV21b, FSR24, RMM22, SN21b, WY24b, Yan22, ZYZ24, ZZWX23]. eigenvector [IMT23, TL21]. eigenvector-dependent [TL21]. eighth [KFK^{+}23, QXZG22]. eighth-order [QXZG22]. Einstein [EJHR24]. Ekeland [ZZ24]. Elastic [SW21b, AZ20, BRLS23, CFO22, DT24, FFS24, FPS22, KP20, Khl21, LSE24, LLT^{+}21, LD22, NT23, SGD20, SD21, TT20, WYC24, YY24b, ZSLG21]. elastic-net [LD22]. elastically [PSZ22]. elasticity [CPBG23, DTMJ24, Dom21, DV21a, GYZ22, GJW23, HL21c, JSRdS23, LQS23, MF21, WMZZ23, WZ23a, ZWZ23c]. elastodynamics [ADDG22, CJM20]. electric [AM21, FdOPS21]. electrical [ACFM21, Ma24a, SIS^{+}22, ZPW^{+}24, LW23b]. electroanalytical [BMS22]. electrochemical [KNOR20]. electrochemistry [BMV23]. electrohydrodynamic [ZPW^{+}24]. electromagnetic [ABA20, CLM22b, HLLL23, MDS20, VNA22, ZCSL21]. electromagnetics [ACEIJH22]. electromechanics [CK24a]. electron [DYM21]. electronic [HOOT20]. Electroosmotic [AAA20]. electrophysiology [CK22]. electroporoelasticity [HM23]. Element [AVC22, AGS^{+}23, ANM24, KVHC23, MWBM22, SVC20, SVWC22, TVSC20, VSSV20, ADKH20, AMS23, ATM21, ATHW21, ACS22, AMZ24b, AMZ24a, ACPV24, BDS20, CC24, CGO24, CUN24, CHY20, CLHL20, CL22, CO19, CJL^{+}24b, CHLR22, CK23, CKL23, CRRB21, CGM21, CFO22, DCY24, DZZ24, DBSS^{+}21, DG22a, DC20b, DHH24, DHLT24, DLZ22, DYF24, DN20b, ER20b, FFS24, FM21c, FSC24, FS24a, GLYS21, GW23, GGVRB22, GG24, GDA22, GPP^{+}20, Grz21, Gua20, GGT24, GLLZ21, GH22, GSWH22, HYW20, HDW20, mHLZ20, HCCC21, HLLY24, HCCZ24, HCC24, HW20, HQS20, HC20b, HC21, HS20, HKL21, HY21b, HSC21, HLYZ24, HZ24b, HWAT22, JS22a, JZ22, JGLH20, JL24b, JZ21, KP22, KNM24, KK24b, KD20a, KD20b, Ku24, KD22, LSC21, LS21, LYM22, Lee23, LLFT20, LZWC21, LWYZ21, LWW22a, LY22a, LWW23b, LWW23a, LWWY24, LL24c, LLGC24a, LLG24, LLZ24a]. element [LLGC24b, LZ19, LCH23, LCW20, LHMT20, LW21, LWWN22, LY23a, LV23, LL24e, LCZL24, LDM23, LRQV24, LR24a, LR24b, MF21, MGD22, MTZ23, MCD24, Ma24a, MSD21, MCL22, MM22b, MMF24, MDKG20, Mu21, NL23a, NSL22, Nud24, OBdSD24, PL21, PH20a, PKE20, PH21, QZW23, RS23, RSRW24, RK23, SEAS22, SS20a, Sha24, SGY22, SKS22, SSV23, SLLP24, SGD20, TXZ24, THS24, VdG20, WZZ20, Wan20a, WGS20, WTL21, WZ21, WWYZ21, WWH21a, WWH21b, WYH22, WGC22, WYZZ22, WCLY23, WMZZ23, WWZ23, WWW^{+}24, WWTL24, WYZ24, WLJ24, WC24a, WLM24, WWZ24, WSHW22, vtW24, WDY20, WLS24, XL23, XW21, XFW23, XH20a, YDC20, YZ20b, YZ21c, YZ21b, YCH24, YCH^{+}21, YLS20, ZZ20a, ZGJ21, ZG22, ZLCT22, ZR24a, Zha24, ZWZ23c, ZR23, ZMFL20, ZOK21, ZW22, ZZL20, ZR24b]. element-free [MCD24]. element-upwind [YLS20]. elements [AK22a, CHZ21, CH20, CDJ20, DKMT24, Grz21, Kop20, LWWZ24, OHCC22, Par20, RBR23, RR21, SMRN24, WZH^{+}21, YZ24b]. elimination [ALKH21]. ELLAM [CDL22, WZ20a]. ellipses [CDTFGV22]. ellipsoid [ABA20, AZ20]. ellipsoidal [Cho24a, FV24]. ellipsoids [CDTFGV22]. elliptic [AN24, AKP23, Bog21, CGS20, CUN24, CDD20, CM20, CA23, CV20, CSS24, DDF24, DC20b, DLC21, DZLC20, DLZ22, FHM23, GDA22, GLCL20, GS21a, GO24, GT23, GSWH22, HDW20, mHLZ20, HHM22b, He23, HH24a, HWAT22, IHK20, JT20, Kar20, KP22, KD23, KRV24, LZ20a, LWYZ21, LWWY24, LWZ24, LCH23, LWWN22, LZ23b, LTZ24, LV20, MkCZ20, MYZ23, MHHV20, NMMH23, PL21, PSRVA24, Pou24, SSL22, Tan20, TXZ24, Wan20a, WWH21a, WGC22, WWZ24, WKC20, WZZ24, YL21b, ZTZZ20, ZG22, ZDL22, ZDW24, ZXZ24]. elliptical [ANN21, LY24a, Vab21, ZYB21, ZY22]. elliptically [RHA20]. Elsässer [EKÇ20]. Elzaki [RKS^{+}22]. EMAC [HJM24]. Embedded [FCS22, CRW20, NY20, YN21]. embedded-hybridized [CRW20]. Embedding [SBC^{+}24, SW24b]. Emden [RR23a, SSVS20]. emergency [SVK20]. emerging [Hri22, HC20a]. emphasis [RL23]. Empirical [TFM24, KA24, MB24, ZMZ^{+}24]. Employing [SZL22]. emulators [MMV20]. encapsulating [Gök23]. encoder [NY24a]. end [LPP24, LPLP21, SLZL23]. ended [BS24a]. endpoint [Mag20, YSN24]. energetic [ADDG22, MV22]. Energy [SYS24, THS24, VdG20, ZJGZ20, AZL21, BSdCCGV23, BCCW22, BIZ20, CZY22, CG22, CQZ22, CCL20, DWWZ22, DHQZ24, EWT24, Fab22, HF24, HL21b, Hu24, JWG21, JZW22, LL20a, LLT^{+}21, LLYM21, LZW^{+}22, LHG^{+}24, LJL24, LY21, LY22b, LF24, MDMS22, MDSRFO24, MSD21, MZY23, MO20, PHV24, PZ21, QQSZ24, SL20, SD21, WY20, WCK20, Wan21a, WWJ22, WYJ23, WC24a, XLKL23, YQZS21, YLJW24, YLK24, ZZG20, ZHY20, ZGJ21, ZOW^{+}21, ZLS23, ZPW^{+}24, ZWKS24]. Energy-based [VdG20]. energy-conserved [LL20a]. energy-conserving [CQZ22]. energy-mass-preserving [LF24]. energy-minimizing [LZW^{+}22, LHG^{+}24, WCK20]. energy-preserving [BIZ20, HL21b, Hu24, JWG21, LJL24, MDMS22, Wan21a, YQZS21]. Energy-stable [THS24, ZJGZ20, Fab22, YLJW24, YLK24]. enforcement [RR21]. Engineering [VAAR22, CLS20, ZDF^{+}20]. enhance [FdOPS21, SBC^{+}24]. enhanced [FPS22, YW23]. enhancement [BS23, JMHZ24, RCVA24]. Enhancing [SBKR24, RFF23]. enriched [CA23, HLMY24, JZ24]. enrichment [CA23, Men24, Nud24]. enrichments [GLZ24]. ensemble [GJW23, JQ22, NRGJ21, SMR23, ŚK23b, YYBL24]. enterprise [CQBL20]. entire [Kuz22, LP20]. Entropy [YKBH23, YVXX24, BGIR24, TA22]. entropy-based [BGIR24]. Entropy-stable [YKBH23]. Enumerating [ADMV22]. enumerative [HH21b]. environment [BBBU22, GJNM24, Gök23, HTZ21, IIS^{+}21, PBR21]. environmental [XCWY20, YLL20, YO20, YTT23]. environments [LWX24]. EP [FLMP24, ZCT21]. epidemic [AA23, CGW21, Cso20, MP23, Sal21a, YPLM22]. epidemiological [BYE^{+}23, YBAE23]. epidemiology [Pat23, SB24a]. epistemic [HH23]. equal [BGdlR24, ZMFL20]. equal-order [ZMFL20]. equality [DV21a, MM22a]. Equation [DVRG21, ZCY^{+}20, ADZ20, Abi20, ATHW21, AL20, ACDV22, ALK20, AAHK24, ACD22, AA22, AA24b, AA24a, AAM20, ATO22, Bac19, BT21, BBAO23, BJB22, BK24, BB21b, BDPR23, Bra23, BDS20, Bro23, CHH20, CHY20, CG22, CLS24, CMS20b, CW24, DD23, DL21, DVRV22, DMNO24, DL23b, DW20, DW23, DZLC20, DWZ20, DV21b, EOT22, ESA23, Fab22, Fan22b, FSC24, FY22, Gao22, GRVZ23, GBTS22, GSK20, GPK23, GJL23, Gua20, GWWY21, Guo24, HMWY23, HLW22, HSGLE19, HVM22, HR22, HWRM24, HC20b, HC21, HZZ21, HZZ24, Hu24, HS20, HTIA24, IYA^{+}23, JS22a, JA22, JWZ21, JG21a, JLZ22, KA24, KK23, KF24, Kim20, KMN20, KXW23, KAD22b, KW23, KLK21, KK20c, KK21, KMD21, KD22, Lee20, LKM20, LYK23, LLFT20, LWD20, LG23, LLG24, LL20b, LH21, LYYZ24, LCW20, LY21, LY22b, Liu22, LY23a]. equation [Liu23, LTZ24, LH20, LXF21, Lu24, LT21, MGD22, MD22, MDSRFO24, MM20a, MK24, MSD21, MGA24, MTIR22, MHHV20, MA19, NS24a, NZA24, NAA20, Ois20, PKE20, QH23, QSW24, QX24, RDHA20, RHM22, RS24a, RS24b, RG20a, RR22c, Rou24, RZ20, RB23b, SM23, SM22a, SMM22, SS20a, SC21, SCK22, Sha23, SW24a, SY24, SXX21, SC23a, SZ24, SSL22, SSM23, SHV^{+}22, SKVA22, TKT^{+}24, TFR^{+}20, TY23, TZTC20, TBT^{+}20, TZ20, Vab23, VV23, WYZZ22, WC24b, WL24b, WC24a, WKN22, WY21a, WX22, WLW23, Wol24, WLZ22, XJK19, XFW24, XFW23, Xin21, YX20, YZW21, YQZS21, bYW23, YLD19, YPL20, YLFT21, YL21b, YT24, YZZ20a, YZ24b, YYBL24, YT19, YLLA23, YZLL22, YZ24d, ZBS22, ZC21, ZZG20, ZJGZ20, ZSMG20, ZKT20, ZGJ21, ZOW^{+}21, ZB22, ZWY22, ZLJT22, ZLS23, ZGZ23, ZF24, ZX23a, ZWL21, ZHZ24, BBGI24]. Equation [PP24, WZZ20]. Equations [APW^{+}21, ASADB23, AC21, ABP24, ASML23, ALNK24, AAP24, AA21a, AKM^{+}24, AEST21, AM21, ATM21, AAKA24, AS21a, AH22, All24, AGS24, ADEEY21, AL16, AZD21, ASA^{+}21, AZL21, AF20, AS21b, ACC23, AHK^{+}23, Ars22, Ask21, AA20a, AA20b, AA21b, AA21c, AI22b, BJB20, BTBR21, BV22, BA20, BS20, BVDB20, BLN21, BR22, BR20a, BuR24, Bog21, BGN^{+}20, BK21b, BGdlR24, BCVV20, BM22b, CMW24, CD20a, CK22, CK24a, CLL21, CZHQ23, CWW23a, Cap20, CM22, CCETZ21, CF20, CS24, CAN23, CLHL20, CW21b, CMZ22, CLYS24, CQZ22, CZ23b, CK23, CH20, CHSZ24, CKL23, CVPM22, CST19, Cso20, CLL20b, DWWZ20, DZZ24, DS20, Dar20, DTM20, DTM21, DGL24, DR23a, DDM20, DK24, DOR21, DO22, DS22b, DdO22, DO21, DLC21, DFFM21, DPW24, DHKV20, DG22b, DT24, DKMT24, DL20c]. equations [DYM21, DHLT24, DAM23, DN20a, DN20b, DCAA22, DKM^{+}20, Ela23, ER20b, EGG23, EEE21, FM21a, FFS24, FZ20, FZ21, FL23, FM22, FLT20, FLZ21, FO22, FMRS24, FP24, FK23, FJ21, Fro24, GJNR21, GGRCGG^{+}23, GZ23, GM22, GJN20, GGHY24, GLHL24, GANR22, GG24, GLCL20, GEG23, GQR23, Gol24, GGO21, GS22, GYZ22, GLHJ21, GW24, HJ20, HBGS21, HH21a, HJM24, HM21a, HM21b, HRT20, HMA24, HYW20, HLZ20, HL24a, HVMS22, HR23, HSMT20, HAD24, HL21b, HCH20, HM23, HWD^{+}24, HLMY24, HKL21, HY21b, HL22c, HKM24, IAH24, IpAmK20, JT20, JO24b, JZ22, JHG^{+}20, JWG21, JG21b, JZZH22, JQ22, JH24, JL24a, JL24b, JHZY20, JCLC20, JSRdS23, KK22, KN20, KAS23, KCKK24, KM20, KM22, KKA20, Kaz21, KAMW21, KD20a, KSZ21, KPS21, Kum22, KS21, KRV24, LBBB23, LSC21, LY20, LZF20, Li20a]. equations [Li20b, LFL20, LZL21, LHHW21, LWYZ21, LWW22a, LLS22, LHW22, LS23, LWW23b, LWLW23, LHHW23, LWWZ24, LVW24, LL24c, LLZ24a, Li24a, Li24b, LLW22, LCH23, LSY23, LDC20, LWXL21, LHH21, LLX^{+}22b, LG22, LZWL23, LZ23b, Liu24, LF24, LT24b, LY23b, LZG24, LZ20c, LXZ23, LXYL24, MW20, MH21, MTZ23, Ma24b, MD20, MDG22, MPM23, MY21, MMJ21, MMM20, MH22, MS20a, MKN20, MS22, MV22, MPV23, MH23, Mia24, MA21a, MDKG20, MMPM22, MVFAMF21, NKBJ20, Naj20, NNJ24, NSL22, NM24, NAS22, NB23, OWV24, ÖKG23, PH24, PO21, PL21, PJA24, PCM21, PRW^{+}22, PAAI24, PM20b, PMVB22, Pat23, PL24a, Pie20, PLS22, PNS21, Pul22, QQSZ24, Qiu20, QWY24, RO20, RYH24, RB19, Ray23, RRS24, RB23a, RGA20, RCM20, RWW22, RR23a, RR23b, SuR23, SES20, SESH23, SM20, SDV21, SH23, SuR20, SS24a, Sha21, Sha24]. equations [SBKR24, Shi21a, SSL24, SLXF24, SKVA23, SW20, SM22b, SYD21, SL24a, SZA24, SB24b, Sun20, SLW^{+}22, SDW24b, SL24b, SS24b, SSVS20, TH24a, TM24, TXS21, THS24, TCGY24, Ust21, Vab22, VNV23, WMF20, WEEA20, WYD20, WGS20, WXS21, WZZ21b, WZ21, WZZ21a, WYH22, WMH22, WOV22, WSY22, WZ23b, WW23, WLZ23, WCMW23, WWW^{+}24, WYZ24, WLJ24, WWZ24, WAH22, WSY20, WH24, Wen21b, WZH^{+}21, WZC22b, WSL^{+}23, WLS24, XHLD21, XY24, XZL20, XW21, XA21, XLLA21, YSN24, Yam20, Yam21a, Yam21b, YQZS22, YCXW24, YVXX24, YTMK19, YWKM19, YYY21, YZ21a, YH22, YSFS22, YTD23, YWZ24, YZM23, hYbJzJ21, YJJW23, YFW20, YHQW21, YCCZ24, YY24b, YZ24c, YZ21d, uDuIZ22, ZVT^{+}23, ZE23, ZvS23, ZM20, ZYZ22, ZZW^{+}24b, ZZW^{+}24c, ZTZZ20, ZHY20, ZFwC20, ZG20a, ZLG20, Zha21c, ZZwS21, ZSL22, ZZ22a, ZLCT22, ZLZL23, ZWZ23b, ZSL23, ZC24, Zha24, ZY20, ZFM22, ZFW22]. equations [ZH24a, ZYL20, ZEW23, ZCZ24, ZXC22, ZMFL20, ZZ20b, Zho20, ZOK21, ZS22, ZCSS24, Zhu21, ZVS21, ZTS21, dSFSS24, dOdO23]. equidistribution [KSVA22]. equilibrated [CHZ21]. equilibria [CFB24a, CFB24b, GL24, GHM21, HTOC20]. Equilibrium [hKhWjH21, KP20, LSE24, ZLL20a, ACK23, CSLQ22, DHF23, EL23, FHH21, GPK23, HSM20, Hun22, IM24, LYL^{+}24, OGR22, SH24a, WZ23c, YCX20, ZCJL20]. equipment [LPLP21, dL23a]. equipped [FM21b, HBLZ21]. equity [BD22, SPS21, Wan24, ZYY20]. equity-indexed [SPS21]. equity-linked [BD22, ZYY20]. equivalence [CT20, PT24, UPPZ24, YHQW21, ZXZ24]. equivalences [AQ20, AM22a, AHPS22, BLV20, GÇSA23, LLY21]. Equivalent [BAD^{+}22]. equivariant [ZZNL24]. ERBS [KD20b]. Ergodic [FLZ21, HS23]. ERKN [YZ23]. Erlang [GHL21]. Erlangization [DLS20]. erosion [FMTB21]. Errata [MHY21]. Erratum [BKMO20, KZ24, MR25]. Error [APR24, BGH22, Bro23, CMS20a, CLH22, CRRB21, Den24, ERB^{+}22, FS20, GANR22, GT23, GLLZ21, HSYZ24, JPS24, KR21, LRV23, LS23, Liu24, LDM23, LXZ23, QH23, RR22d, SZ23, YSN24, ZZG20, ZC24, ANR20, AESW22, APR21, Bac19, CHZ21, CZHQ23, CE24, CC21a, CHY20, CLM22a, DCY24, DJ22b, DHH24, DC23, DLZ22, DN20b, Fab22, FK23, GJNR21, GGVRB22, GDA22, GS20c, GS21b, GWWY21, HH21a, HG22, HSRA24, HY21b, HTZ21, HH24b, JHZY20, KP22, KMN20, KWYN23, Ku24, KCRVA22, LMP20, LY22a, LL24c, LLGC24a, LZ19, LCH23, LWWN22, LY23a, LV23, LCGW24, LCZL24, MVY19, MSD21, MS22, Mas21, MR23, MR25, MORO20, MLY22, NMST21, NL23a, Ngo24, OBdSD24, Pou24, Rab21, RS23, RlLZZ21, RL23, SM22a, SI23, SV23, SS24a, TXZ24, TXS21, TGJY22, WK20a, WZ20a, WW23, WZ24a, WLJ24, WZD^{+}24, Wu22]. error [YX20, YZ21a, YSFS22, YLXY23, YYC24, Yüc21, ZZW^{+}24a, Zha24, ZXK24]. errors [AS20b, DH22, DC23, LRP24, LMS21, RHA20, VS21, YJN20, ZZM20]. escape [NMB22]. Escherichia [BYWA22]. ESR [Yam23]. essentially [CZ24]. estimate [CMS20a, LY23a, Ngo24, RR22d, SM22a, TXS21, WZ20a]. estimates [Ade22, ANR20, APR24, BGH22, CDD20, DCY24, Den24, DJ22b, DHH24, GJNR21, GM24, HG22, KMN20, KWN20, KWYN23, KMT20, KR21, Ku24, KCRVA22, LRV23, LJCL20, LY22a, LLGC24a, LXZ23, MS22, MR23, MR25, MCY23, OBdSD24, Pou24, QH23, RlLZZ21, SZ23, TGJY22, WW23, WLJ24, WZD^{+}24, WN23, YX20, YSFS22, YYC24, ZC24, ZXK24]. Estimating [EK24a, JD22, Let24, GDZN23, Mim20, Zha21b]. Estimation [LRP24, LWLD23, PM20a, SACQ21, WZLX20, AS20a, AESW22, ALK20, BK22, CM24a, CSLQ22, CC21a, CHY20, CLM22a, CEKN20, DXM^{+}20, Din23, HSGLE19, HG23, HSRA24, HX22, HH24b, JKJ^{+}23, KTK21, Liu22, MVY19, MRI22, MIdMR23, MIdMR24, MS24a, MLY22, NL23a, OVGI20, RS23, RHA20, SL22, SLP20, SBM21, SVJG24, TFM24, WHlZH23, WZNT24, WM20, XGYY23, XD23, XDX24, XHW24, YL20, YL21c, Yüc21, ZZWS22, Zho22, Zhu22]. estimations [CD23, DN20b, PA24b, WD20]. estimator [ANN21, GK21b, JHZY20, LWWN22, NRGJ21, TXZ24]. estimators [AK22b, BK22, CHZ21, ÇK24b, DLZ22, GS24, HS23, SCHS24, WMH22, YSUY24, ZSY24, CMP21]. Eta [MI20]. Eta-based [MI20]. Euler [AIDAJ23, BLN21, BR22, BGN^{+}20, CM22, CG20, DHKV20, ÉM21b, HBGS21, HH21c, HN22b, IuR23, KAMW21, LD21, MWW21, OBO23, RS22b, TM24, TCGY24, YWKM19, YH22, YZ24c, YZ21d, ZLG20, ZSL23]. Eulerian [AFLP22, AAP24, BBSG23, DU19, HBPV21, LS21]. European [ACD22, Boe20, GJNM24, JCY20, Kaz23, LLH24, MN20, RG20a, Wan21b, YL21a]. eutrophication [FMAV23]. EV [ZMZ^{+}24]. evaluate [CC21b, CMRS21, GL20]. evaluating [BBBU22, CJL^{+}24b, MN20]. Evaluation [HZS^{+}24, BC22a, DOP23, GST20, JLL21, KZ23a, KZ24, KH22, PP21]. evaluations [BK20]. evasion [YSFdRdP20]. even [CZ23a, LM21b]. even-order [CZ23a, LM21b]. event [BVBO23, BYE^{+}23, MLS22, TVB21, VBRT24]. events [ZKP21]. evidence [WHC^{+}20, HC20a]. evolution [DAM23, HLZ20, JS21, JLM24, PGG20]. evolutionary [BJF24, CHLY21, CL24b, Hon22, LX24, Ngo22b, PK23, WQ20, ZCF^{+}22]. evolving [Slo21]. EWMA [SCHS24]. Exact [Alh24, FPLZ21, Wu24, ZB22, ZZMK23, BKMO21, BM22a, BS24b, HA20a, LYYZ24, Ois20, PJA24, RHM22]. exactly [BA20]. examples [Hri22]. excess [ZZK21]. excess-of-loss [ZZK21]. exchange [Hai23, LCZ19]. exchangeability [BKMO19, BKMO20]. excitation [Ben24, CK24a]. execution [CMFMSQ22]. exhibiting [SKVA23]. Existence [CHLR22, CFB24a, CFB24b, DPW24, GSK20, HM23, gHtSjT24, IHK20, MK24, Pie20, Pou24, RDHA20, SuR20, BCH20, CW21a, Juk20, Rab21, Sun20, bWLgZbX24]. expanded [Ars22, YCG^{+}23, ZXK24]. expansion [Bha23, CYC21, DIP^{+}22, GMCJGM24, KK20b, LH24, MGG^{+}24, MLYZ21, SPC21, WAAM20, XJK19]. expansions [GL20, LMP20, OH23, PA21, SH24b, WLZ23, ZXK23]. expectations [HG23]. expected [YLH23]. expensive [BKM22, SNMK21]. experiment [BJP24]. Experiments [ARK20, DS22a, Els21b, EGG22, FLMR20]. Experiments-based [ARK20]. expert [LS24]. experts [AVRLRMGG23, YLZ^{+}22]. explainable [HJC24]. explanatory [ZW23]. Explicit [Ade22, Ant20, IKL24, JWG21, LY20, Liu20, MRC23, MRC24, AGST23, AST^{+}23, CLL21, CGW21, CG23, DVRV22, FZ21, FCS22, FRR24, HH24a, HB22, KK20c, KK21, LL24b, Li20a, LY24b, LY23b, MPM23, Mim20, NS24a, Ngo22a, NS24c, SLW21, SSL24, SLLP24, XW21, XA21, ZZ22b, ZCF^{+}22]. Explicit-implicit [IKL24, XA21]. Exploration [CZ20, CQBL20, LSZ23, QJLQ20]. exponent [AHT21, NYL23]. Exponential [Ixa24, KCKK24, LZL21, MQ20, RWW22, Wan21a, WC24b, ASON23, AAB24, AIDAJ23, BABR20, Bha21, BCCW22, BK20, BM22b, CCZ23, CGSVBB^{+}24, CJP20, Cho24b, DIAJ^{+}22, ER21, FZ21, GKV21, HW22, HH21c, HS23, JL24a, JdlCD20, JS22b, KR21, LWLD23, LD21, LM21a, LHE24, MD20, Mas21, MT20a, SW24b, SVJG24, TSMZ20, WTWZ20, YWTL19, ZYY20, dFG21]. exponential-dissipation [Cho24b]. exponentiality [JS22b]. exponentially [BMS21, GHM21, HS24a, HWRM24, LMSZ21, NMMH23, OSK20]. exponentials [GAA19, GAA21]. exponentiated [DNK23, PM20a]. exposure [MSK21]. expression [PH22, VTB21]. Expressions [ZYB21, LC24b, WX21, ZZMK23]. Extended [BK21a, AORS22, BLZL22, BABR20, BEJR22, CC24, CC20, HTJC24, JJR21, Kan20, La 22, LY24a, LW21, LCGW24, MS24b, ON22, SMR23, WXZ21, WW22, WZ23b, WSLJ23, YCXW24, ZSCH21, ZTY^{+}24]. extending [WDPY24]. Extension [ÁGKM21, KSPC22, ME20, AHK24, CFB24a, CFB24b, DMN24, ER21, KZ23b, MDMS22, Ri22, Ri24, VLR23, WSL^{+}23, XLZ20b]. Extensions [Axe20, GDSCO21, gKZ21, LK21, RFF23]. external [KYP21, Sur21]. externalities [KMTZ21]. externally [Grz21]. Extraction [JO24b, SK24b, KD20b]. extragradient [DHF23, LH23, TV22, TDL^{+}22]. Extrapolation [BRZ22, GH22, Saâ20, TDAT23]. Extremal [BM24b, Guo24]. extreme [BSH24, New24, YCD^{+}20, ZYZ24]. extremum [MMM22]. extropy [JS22b, SN21a].

Fabrizio [DO22, TZ20]. face [PVV24]. face-based [PVV24]. faces [Gua20, JO24b]. factor [ABO24, JHG^{+}20, LLH24]. factorisation [ACT22, BWS21]. factorization [HW24a, JW24, KHAW21, KCK24, LLLD20, Ri22, Ri24, SB24b, TWQ^{+}23, WYTL19, WCK^{+}22]. factors [BD22, Els21b, JO24b, LGMC20, LLG^{+}22, MER20, NMF^{+}22, ÜY20]. failed [Jas22, Jas24]. Failure [KT24, CCCMH24, EKCF22, GMSR18, GMSR21, LWLD23, WTL20, ZS20a]. failures [LPLP21]. falling [HR22]. families [HW22, HP22, ZL23a]. family [BEL22, BCT22a, BGN^{+}20, BM21, BCP20, DAZ23, DNK23, EHVMM23, GBTS22, HBMA22, JK21b, KKA20, LG22, Moi24, PI24, SBKR24, WY21b, YSUY24, YKY24, YSFdRdP20, YJJW23, YS24b, ZW22]. far [HSB24]. farm [BYWA22, Kan24]. Fast [All24, Cho24a, DDNZ21, DHF23, JHG^{+}20, KS22a, KRT23, Miy22, RL23, SDL22, SQT24, TK24a, TLD^{+}23, VNA22, ZSLG21, ZX23a, ZM21b, BHRW22, CHH20, DS22a, DEH^{+}24, DRR^{+}21, FNH20, HW21a, Ibr20, JWZ21, LGG^{+}22, LLLD20, LCW20, LHL21, LH23, QWY24, QDHW21, SGD20, TH24a, TDL^{+}22, Tia22, WZL^{+}23, YL22, ZLZL23, ZVS21]. Faster [LYB22, SZZ21, TSL^{+}21]. fatigue [Let24]. fault [DSCZ24, ŚK23b]. fBms [DJ22b]. FCC [Maj23a, Maj23b]. FD [MDD21]. FDM [BM24a, FP24]. FDM-based [BM24a]. FDM/FEM [FP24]. FDTD [BHRW22, DVRG21, LL20a]. FE [PSZ22, SH21]. FE-holomorphic [PSZ22]. fear [Gök23]. fears [YBAE23]. feasibility [BP21, BHB21, HWLD23]. feasible [ZL22]. feature [HCC24, LWZF21, MGV23, NPP^{+}21, WM20]. features [LXLL23, NBLZ24, SBC^{+}24]. February [Ano21-31, Ano22-32, Ano23-32, Ano24-29]. feedback [HH20, Hu22, Mim20]. fees [LLW21]. Feller [DMP23]. FEM [BJF24, CE24, DWZ20, FP24, GS21a, GS21b, GYZ22, LMS21, RR22d, SZ24, SC23b, TSA23, YY24a, YYC24]. FEM’s [CA23, AK22a, CHLY21]. Fenchel [LXQ21].

Fermi [SS24a, ZW24]. fermion [JS21]. FETD [LYL23]. few [ZZWX23]. Feynman [Yam21b]. FFT [YZL24]. FFT-AMIB [YZL24]. Fibonacci [AEST21]. fictitious [AK22a]. fidelity [MLYZ21]. Field [ABHK20, HLA22, LLZY22, SZK20, YLZ^{+}22, AZ20, BGH22, CCS24, CZY22, CGO24, EGM23, FS21, GJDL20, HYY21, HSB24, KLK21, LYM22, LLT^{+}21, LML21, LLYM21, NY24b, QH23, RS22b, SvG22, SBC^{+}24, WTL21, Wan21a, WCLY23, WZY23, XLKL23, YLJW24, YLK24, YOCY22, ZCYP20, ZLS23, ZJBY23, WC24a]. Field-of-values [ABHK20]. fields [AAA^{+}23, Bal24, FdOPS21, ZM21a, BAPPF23]. figures [OGR22]. file [MGV23]. filling [FGPP23, HM22]. film [GDG^{+}21, HR22, Let24, WCPW20]. filter [KCK24, NRGJ21, Six20, SMR23, ZK20]. Filtered [ORT22, AFKM22, BDPR23, MW21]. filtering [FNH20, KTK21, LSZ22, SvG22, SQT24, CD20b]. filters [BCP20, HCCW20]. filtration [IKL23, IKL24]. Final [TBT^{+}20]. finance [BTBR21, QJLQ20, RPQ24, San21, XCWY20]. financial [BYE20, CC21b, CN24, DLL21, GS20b, GS20a, HQL21, JLL21, QGW21, SZWH21, SZL20, Son24, WQ20, WLL21, WHC^{+}20, XLZ20a, XGYY23, YHF22, Zha20a]. financing [CQBL20]. find [KLH21, WYW^{+}24]. Finding [Yam20, BM21, HWLD23, JZM^{+}23, LZ20a, NZ21, PS24a, PH20b, RSS24, SD24, TH24b, Yam21a, ZJBY23]. findings [TA22]. fine [Axe20, BCDF22]. fingers [BEG^{+}22, BEM^{+}24]. Finite [Abi20, AUAA20, AVC22, AGS^{+}23, ANM24, BGG^{+}21, CL22, CFB24b, DHLT24, GPP^{+}20, GQR23, Grz21, GGT24, HH24a, JS22a, KVHC23, KD20b, LYM22, LP20, LRQV24, MF21, MSD21, SS20a, SKS22, SVC20, SVWC22, TVSC20, VSSV20, Wol24, ATM21, ATHW21, ACS22, AJR22, AET23, AMZ24b, AMZ24a, ACPV24, BL23, Bra23, BDS20, CC24, CGO24, CUN24, CHY20, CO19, CDS21, CHLR22, CK23, CH20, CKL23, CRRB21, CGM21, CDJ20, CPP23, CC20, CFO22, CLL20b, DCY24, DZZ24, Dar20, DL21, DOT23, DVRV22, DG22a, DC20b, DHH24, DKMT24, DLZ22, DYF24, DN20b, EL23, FSC24, FS24a, FHM23, GLYS21, GW23, GGVRB22, GBTS22, GDA22, GS21c, Gua20, GLLZ21, GH22, GSWH22, GKV21, HET24, HYW20, HDW20, mHLZ20, HCCC21, HL24a, HLLY24, HCCZ24, HCC24, HE24, HW20, Hor20, HL21b, HQS20, HC20b, HC21]. finite [HS20, HKL21, HSC21, HLYZ24, HSYZ24, HZ24b, HWAT22, JZ22, JGLH20, JL24b, JZ21, KSPC22, KP22, KNM24, KK20a, Kop20, KK24b, KD20a, Ku24, KD22, KCRVA22, LSC21, LS21, LR22, Lee23, LL24a, LMZ22, LLFT20, LJCL20, LZWC21, LWYZ21, LHL^{+}22, LWW22a, LY22a, LWW23b, LWW23a, LWWY24, LL24c, LLGC24a, LLG24, LLGC24b, LZ19, LLLL20, LCW20, LHMT20, LZ20b, LW21, LSZ22, LWWN22, LY23a, LV23, LL24e, LDM23, LR24a, LR24b, MGD22, MTZ23, Ma24a, MDSRFO24, MM22b, Mia24, Mu21, NL23a, NYL23, NSL22, Nud24, ORT22, OBdSD24, PH20a, PKE20, PH21, RS23, RSRW24, RG20a, RR21, RK23, SEAS22, SA24, SM22a, SRMD22, Sha24, SGY22, SMRN24, Son24, SSV23, SLLP24, TK24a, TXZ24, TT20, Ter20, THS24, UGG^{+}20, VdG20, WZZ20, Wan20a, WGS20, WTL21, WZ21, WWYZ21, WWH21b, WYH22, WGC22, WSY22, WYZZ22, WCLY23, WLH23, WMZZ23, WW23, WWZ23, WWW^{+}24]. finite [WWTL24, WYZ24, WLJ24, WC24a, WLM24, WWZ24, WY21a, WSHW22, WLS24, XWXZ22, XW21, XZ22, XFW23, XH20a, XLL23, YQZS21, YSBT20, YLFT21, YMC22, YDC20, YZ20b, YZ21c, YZ21b, YZ24b, YLLA23, YLS20, YLS23, ZZ20a, ZGJ21, ZZwS21, ZG22, ZZ22a, ZR24a, Zha24, ZIZ21, ZWZ23c, ZR23, ZY24, ZMFL20, ZW22, ZZL20, ZR24b, ZGZW23]. finite-difference [CC20, DVRV22, JLZ22, MDSRFO24, SRMD22]. finite-difference-based [ZY24]. finite-part [XLL23]. finite-step [LSZ22]. Finite-time [AUAA20, GBTS22, NYL23, XZ22]. finite-variance [ZGZW23]. finite-volume [TT20, Ter20]. FinTech [CC21b]. First [CLW^{+}22, CGW21, LLX^{+}22a, MMMM20, Ade22, CDCVV22, CS24, CNQRR22, DRR22, DDRS21, LWW23a, LWLD23, QZW23, Vab23, ZZ22b]. First- [LLX^{+}22a, MMMM20]. first-failure [LWLD23]. First-order [CLW^{+}22, CNQRR22, LWW23a, Vab23]. Fisher [CG23, JA22, RR22c]. fit [MI20, ZG20b]. Fitted [GKV21, FHY^{+}21, HWRM24, LF24, OWV24]. Fitting [FGPP23, GHL21, Din23, HTD21, Ixa24, MJ24, Pou21, Seg21, ZFH20]. FitzHugh [RS22b]. five [YHQW21]. Fixed [ABM^{+}22, WSX21, ANR20, BGR20, BP21, BM21, BHJB21, DK24, EK24b, HLR22, HWLD23, JZ24, LMYZ21, MK24, New24, NB23, PH24, Saâ20, SuR20, SRBK23, SWN23, TSG22, TLD^{+}23, TH24b, Wen21a, YTA22, ZVT^{+}23]. Fixed-point [WSX21, HLR22, NB23, YTA22]. fixed-stress [ANR20]. flexible [CJL^{+}22, DDT23]. flight [BBSG23, RDG23]. Floater [TV24, CHS20]. Floater-Hormann [TV24]. floating [ERB^{+}22]. floating-point [ERB^{+}22]. Flory [CJL^{+}24a, DWWZ22]. Flow [EPL21, FRT23, ADKH20, ACS22, AJR22, AK21, AAA20, CS21, CZM23, DMST21, DYF24, HET24, HMWY23, Hur22, KCJ23, Kes23, LJCL20, LHMT20, LP23b, LN23, MHM23, NYL23, NA21, PS22a, PHV24, PH21, RCN21, RL24, RV20, SMKV21, SVC20, SVWC22, SV24, TSA23, TKT^{+}24, TK24b, WLHJ20, WY20, WWH21b, WCLY23, WLH23, WAAM20, XWXZ22, YN21, YOCY22, ZZWS20, ZOW^{+}21, Zha21b, ZGG23]. flow-coupled [YOCY22]. Flow-driven [EPL21]. flows [AN24, BM20, CMS20a, CHP21, CGM21, DL20b, FDDM24, FHH21, GRS20, HCC24, HBPV21, HF23, HLA22, LZWC21, Moo22, MBG21a, NY20, PCMH20, PFG24, SRAG23, VMP23, WZZ24, YZL24, YOCY22, ZW24, ZaZWZ23, ZHW20]. fluctuating [Gök23]. Fluid [ZAC^{+}23, ADKH20, AESW22, AAA20, BHPV23, CWLM22, CD22, DYF24, FFT20, GH22, IKL23, NA21, RCN21, Svá23, VS23, VMP23, WCLY23, WLH23, ZSH20]. fluid-fluid [ZSH20]. fluid-saturated [RCN21]. Fluid-Structure [ZAC^{+}23, AESW22, BHPV23, VS23]. fluids [Kes23, XLKL23, ZW21, ZHMD22]. fluorescence [LCB20]. flutter [PZY^{+}24]. flux [AET23, CD22, HCCC21, HCCZ24, KKO23, RK23, WV21, YF22]. flux-conservative [CD22]. FMQHI [Hun21]. FMQHI-fuzzy [Hun21]. focal [LRP24]. focus [WGM21]. Fokker [Wol24]. Folded [LTF^{+}20, BN21, BSH24]. folds [SH21]. following [AKMO21]. food [GJGDPR23]. footrule [BKMO21, BM22a, BS24b]. force [DTMJ24, Grz21, VS23, VNA22]. force-based [DTMJ24]. forced [BCH20, CVPM22, Liu23]. forces [KV19, PV22, YZL24]. Forchheimer [CE24, CGO24, CRRB21, HCCC21, LHL^{+}22, SV24, TGJY22]. forecast [GS20a, GNGGL^{+}19]. Forecasting [SUW23, ABO24, HCCW20, Tak20]. foreground [ZLZ^{+}24]. foreground-background [ZLZ^{+}24]. Foreword [BSZ21, NOR21, ÖYT23]. form [ALKH21, CMVV20, GW23, LWW22a, LC24b, MMF24, PRW^{+}22, RR23b, WMZZ23, YHQW21, ZTZZ20]. formal [CDMM22]. formalism [RHM22]. format [HSMT20, LZZZ21]. formation [AKMO21, HRT20]. formed [DOP23]. forms [gK23, Liu20, LZLL20, LK21, NZ21]. formula [AV21, BD24, BO22, HJL20, HH21b, Lem24, MNH21, MCMR23, ONT21, SD21, TSMZ20, YL21b, dC20a]. formulae [AAMR24, CDTFGV22]. formulas [Alh24, BD20, HJS21, MMJM22, RS22a]. Formulation [ACEIJH22, DMTW22, ELTD23, HJM24, KZM21, LRV23, RR23a, WWTL24]. formulations [CHLR22, JS22a, WHZ24, YSFdRdP20]. Fort [DL23b]. forward [BPRZ21, CLYS24, ISD23, LRP24, LF21, Ma24a, SI23, Tam20]. forward-backward [BPRZ21, LF21]. forward-reflected-backward [ISD23]. found [For24, TFR22]. foundation [NT23, TZBK21]. Four [YZ24b, AHK24, IY20, KR21, LZ24a, YKY24]. four-directional [LZ24a]. Four-order [YZ24b]. four-point [YKY24]. Fourier [CJLL21, DDF23, HJL20, KK22, KEKT21, LXLL23, MWLY24, MMJ21, MCMR23, Oga20, RG21, SS20b, VdG20, Wan24, WSX21, ZSMG20]. Fourier-type [RG21]. fourteen [LLL^{+}22b]. Fourth [MJR20, MMPM22, PM20b, SLXF24, BK24, BFQ24, DL21, DLZ22, HL21b, HTIA24, LL20a, LMLB22, LC24a, MGG^{+}24, QZW23, RR22a, TXS21, THS24, VV23, XWXZ22, XLY22, YKY24]. Fourth-order [SLXF24, BFQ24, DL21, HL21b, HTIA24, LL20a, LMLB22, LC24a, QZW23, TXS21, THS24, VV23, XLY22, YKY24]. Fowler [RR23a, SSVS20]. fractal [BV23, CTN22, CL24c, PA24a, Uça23, WK20b]. fractal-fractional [Uça23]. fraction [MDLY23, SS23a]. Fractional [AHT21, HHL20, SDC20, ASADB23, ABA19, AEST21, AM21, ARF23, AAKA24, ACDV22, ALK20, AAA21, AA23, AH22, AAHK24, AKMO21, AGS24, ADEEY21, AZD21, ASA^{+}21, Ara20, AB23, AS21b, AMT23, AI22b, BJB20, BHRW22, BBAO23, BJB22, BB21b, BR22, BuR24, Bic23, BYE^{+}23, BCVV20, CHXL20, CBJ20, CLH22, CLS24, CQZ22, CL24c, CW24, CV20, CST19, CN24, DWWZ20, DRR22, DTM20, DFH21, DK24, DOR21, DO22, DdO22, DHKV20, DKMT24, DSCZ24, DP24, DXWC22, ESA23, EGG23, EY23, Evi23, FM21a, FLT20, FZQ22, FU19, FB24, FK23, GZ23, GM22, GS21c, Gu21, HL21a, HZC^{+}23, HLZ20, HZMD22, HR22, HR23, HOGO23, HL21b, HL24b, HZZ24, HS20, HPWM20, HSC21, HTIA24, Hur22, IpAmK20, IuR23, JBSN22, JWZ21, JHG^{+}20, JVAD22, KCKK24, KGV23, KKA20, Kaz23, KH24, KKJ20, KKJ22, KPS21]. fractional [KMD21, Kum22, LHES21, LLZC21, LG23, LHHW23, LL23, LL20b, LLLL20, LCW20, LHH21, Liu23, LYL^{+}24, Liu24, LT24b, LZ20c, MGD22, MDG22, MDMS22, MY21, MMJ21, Mam23, MSD21, MP24, MDS20, MA19, ME20, MS20c, NRV21, NZA24, OOO22, OWV24, ÖKG23, ÖHS23, PA24a, PAAI24, PM20b, QX24, RDHA20, RO20, RVS^{+}22, RHM22, Ray23, RGA20, RG23, Rou24, RZ20, SuR23, SES20, SESH23, SM23, SBC20, SM22a, SMM22, SuR20, SRMD22, SPMB21, Sha23, SDLZ23, SW24a, SDW24a, SXX21, SC23a, SSS24, SKVA23, SZ23, SJZ20, SYTD24, Sou22, SHV^{+}22, SS24b, Sur21, TK24a, TKM21, TH24a, TSL^{+}21, TBT^{+}20, TZ20, TVAS23, Uça23, Ust21, Vab21, Vab22, VORD20, WK20b, WYD20, WXS21, WL21, WZZ21a, WOV22, WYJ23, WLB^{+}21, WY21a, WX22, WLW23, WD20, bWLgZbX24, XYAH23, XXL21, XLLA21, YZW21, bYW23, YPL20, YYY21]. fractional [YMC22, YT24, YFW20, YLLA23, YZ24d, ZVT^{+}23, ZE23, ZYZ22, ZWZ^{+}23a, Zha21c, ZZwS21, ZWY22, ZGZ23, ZF24, ZEW23, ZHZ24, ZS20b]. fractional-order [AA23, AB23, AMT23, BYE^{+}23, DSCZ24, ESA23, LL23, SXX21, TK24a, TSL^{+}21, XYAH23]. fracture [CL22, CO19, DYZ24, Kov22, KSG^{+}22, LYM22, YN21, ZR23]. fractured [ACS22, AVC22, ASTK24, FHB20, IKL23, LZWC21, SVC20, SV24, TVAS23, WCC^{+}19, WLH23, AGST23]. fractured-porous [IKL23]. fractures [NY20]. fragmentation [GSK20]. frailty [BVBO23]. frame [LRT20]. Frames [IN22, JW24]. framework [CBF24, CK23, CMRS21, CPS20, FP21, GLYS21, HWG21, HMT23, LPY20, MN22, PP24, RSRW24, SRS21, SCY23, SS20c, SPC21, YCX20, YTD23, YO20, YZ24c, ZAC^{+}23]. Frankel [DL23b]. Fréchet [Pie20, TTK23]. Fredholm [BR20a, CAN23, CHH20, DDM20, DOR21, DMNO24, DCAA22, Gol24, HVMS22, HWRM24, KM20, Kaz21, Kum22, MS20a, MKN20, MH23, Naj20, NAA20, NB23, Pat23, RS24a, RS24b, RGA20, SES20, SESH23, SM20, SL24b]. Free [KD22, ADKH20, AGH20, ATHW21, ALKH21, Bal24, BCT22b, BKM22, CCZ23, DP21b, DKL20, DHZ24, GCV20, GH21, GYZ22, HHM22b, HBPV21, HM22, IAH24, JO24b, JG21b, JG21a, KSPC22, KK24b, LLGC24a, MGD22, MCD24, MDSRFO24, MVB20, MWBM22, Mia24, MP21, NNJ24, TK24b, WSC23, YZ23, YZ20b, YZ21c, YZ21b, ZUZ20, ZZWS22, ZYZ24, ZZZ^{+}22]. free-boundary [JG21a]. free-stabilization [MWBM22]. free-tree [YZ23]. freezing [SSV23, VSSV20]. frequency [DKM23a, FHY^{+}21, Lee23, WW22, YZ23, ZuIKA22]. friction [CFS21, CHLR22, DL20b, NT23, WC21]. frictional [LX24, PW23, WLHJ20]. frog [SS24b]. front [AVL22]. frontier [TTK23]. fronts [SEAS22]. frozen [MAA^{+}22, Zha21b]. FSC [EPL21]. fuel [CTGV20]. Full [MDE24, WWTL24, KMN20, KCS21, TT20, YHW22, ZZG20]. full-discrete [ZZG20]. Full-rank [MDE24]. full-tensor [TT20]. full-waveform [YHW22]. Fully [CZY22, Hu24, LLZY22, MZY23, WZY23, ZHY20, CK22, CK24a, CHLY21, CGM21, CNQRR22, DHH24, FGN^{+}23, GLHL24, GG24, HCC24, LS23, LG22, Liu24, LR24a, SKVA23, SZ23, TGJY22, WLJ24, Zha24, ZaZWZ23]. fully-decoupled [ZaZWZ23]. Fully-discrete [CZY22, MZY23, TGJY22]. fully-mixed [CGM21]. function [AC22, BBSV21, Bay22, BKM22, Che22, EMR21, GHG24, GPST23, GN24, GJ21, IpAmK20, JJR21, JCLC20, Juk20, JC24, KUA20, Kuz22, LCD21, MIdMR23, MIdMR24, MBG21b, MMV20, NNL^{+}24, OT21, PS24a, PSZ22, PP21, Seg23, SK24b, Son24, Vab22, VTB21, WKC20, XZ22, YSBT20, YLXY23, ZG20b, Zha21b, Zhu20a, Zhu22, dS24]. function-differential [WKC20]. functional [AL16, Ask21, BLN21, Bra21, CVPM22, FL23, HLLY24, LY23b, MM22a, PS21, Pie20, PT24, SLZD20, SS23c, WZLX20, WXS21]. Functionalized [ZOW^{+}21]. functionally [HHL20]. functionals [Bre22, FMRT20]. functions [ABM^{+}22, Ans21, Ant20, AID24a, BAB20, BV23, BYE20, BD20, Bra23, CCZ23, CCGT22, CA23, DdO22, Den24, FO22, FGPP23, HL22a, HS24a, HR23, HbX24, HTZ21, Ixa24, JPS24, JdlCD20, KZW22, KWI24, KH21, LCD21, LP20, LP23b, LC24b, LWX24, MS20a, MHM23, MR20, MI20, MHY17, MHY21, MSM24, MA21b, MS20c, NS24b, NWD20, NS20, OSK20, ONT21, OH23, PS24a, PA21, PA24b, PL24b, PS22b, RO20, RRS24, RDH21, RG22, SI23, SLP20, Six20, SAL21b, Tak20, Tom24, Ust23, WCK20, WYG23, WLB^{+}21, WQ22, YT20, YFW20, ZuIH20, ZH23, ZXK23, ZXK24, ZH24c, Zhu20a, dHSR22, vdMOB22]. functions/wavelets [HR23]. Fundamental [CW24, DZLC20, EOT22, HIM24]. Funk [Bel23b]. Further [LSZ23]. fused [HJC24]. fusion [WY24a]. future [New24]. futures [YWTL19]. Fuzzy [Bay22, HA20a, PCT23, AJSGL22, AI22b, BYE20, HA20b, HA21b, HJC24, HTJC24, HTTO20, HTOC20, Hun21, HTZ21, HK23, LYL^{+}24, LS24, MS24a, NAA20, OHCC22, SLC22, TBA22, TBA24, Tak20, Ust23, bWLgZbX24, YY21, ZBS22, Zha20b, dHSR22].

GA [SA24]. galaxy [CKT21]. Galerkin [ADKH20, ADZ20, ATM21, ATHW21, ADEEY21, BSdCCGV23, Bac19, BTBR21, BT21, BLRZ23, CGS20, CZY22, CWW23b, CWW23a, CF20, CRW20, CR23, CAN23, CHH20, Che21, CCL20, CM20, DCY24, DZZ24, DG22a, DC23, DW20, DW23, DYF24, DKM^{+}20, FJ21, FC22, GDA22, GEG23, GeOJD24, GT23, Gua20, GSWH22, HLMY24, HLLQ22, HWAT22, JZ22, KKJ20, KD20a, KD22, KD23, LMV24, LRZ23, LLFT20, LWW23b, LW23b, LWWY24, LLGC24a, LLGC24b, Li24b, LWZ24, LSY23, LCW20, LHMT20, LY21, LMLB22, LWWN22, LZ23b, LTZ24, LXY24, LC24a, LYZ24, LXZ23, MF21, MGD22, MCD24, MTIR22, MZY23, MMPM22, Moo22, Mu21, Ned22, PL21, PP24, PW23, PS24c, QS23, QSW24, SEAS22, SSS24, SDW24b, TXZ24, TXS21, TR22, TGJY22, TCGY24, VdG20, WLHJ20, WZZ20, Wan20a, WGS20, WZ21, WWYZ21, WZZ21a, WWH21b, WZC^{+}22a, WGC22, WYZZ22, WZY23, WMZZ23, WWZ23, WZ23a]. Galerkin [WWTL24, WTX24, WYZ24, WWZ24, WSY20, WY21a, XFW23, YSN24, YX20, YKBH23, YTMK19, YMC22, YZZ20b, YZZ20a, YZ20b, YZ21c, YZ21b, YZM23, ZVT^{+}23, ZZW^{+}24a, ZTZZ20, ZZ20a, ZZ22b, ZR24a, ZC24, ZWZ23c, ZR23, ZEW23, ZaZWZ23, ZXC22, ZWJ22, ZLC24, ZR24b]. Galerkin-characteristic [SEAS22]. Galerkin/finite [JLZ22, YMC22]. game [AAk20, EL23, GL24, LLS21, OHK23, WQ20, WZJQ21, WYW^{+}24, Yu22, ZQL24, ZGL20]. games [CCS24, CLYS24, CFL21, CFB24a, CFB24b, HTOC20, HK23, LWZ21, Ret24, YL22, YSFdRdP20, YP21, SZK20]. gamma [AK22b, BKMO21, BM22a, Kuz22, PA21, Zhu20a]. gap [HbX24, HTZ21]. gaps [Alj22]. GARCH [AKN24, KTK21, ZZWS22]. GARCH-in-Mean [KTK21]. gas [CKL20, COT21, MHM23, ZZ21]. gases [ZW24]. Gauge [CMS20a, FC22]. Gauss [APR21, BEJR22, CC20, DOR21, DO22, DDRS21, DDRS24, JJR21, LL24e, MG22, MS24b, RS22a, RS24c, Tom24, ZJY22]. Gauss-type [BEJR22, Tom24]. Gaussian [AS20b, BJ22, BPDC20, BAPPF23, CDS21, CES23, Den24, GJ21, JPS24, LMAV24, LZZ24, Liu24, MSM24, MMV20, OH23, PCT23, PPSM24, RG21, TKVA21, ZRQ22, ZJ24]. GBM [EL22]. GBRT [AGNGG^{+}23]. GCD [FGM21]. GDM [CDL22]. GDP [HZ24a]. Gegenbauer [ER20b, JPS24, ÖKG23, RRS24]. General [CW22, AC21, ASML23, AM21, AHGG20, BB21b, BCCW22, BIJ21, CT20, DNK23, FHC22, HE24, HW20, JWG21, JZW22, JC24, KKKL22, LLS21, LWL22, LD21, LW21, LTZ24, LS24, MFA21, Nud24, OEH23, QWY24, SHB24, SRS21, WWH21a, XL23, Xin21, YCX20, ZYY20, ZSHD24, ZH24a, ZL23a, eBI23]. generalised [EM21a, RG20a]. Generalising [ARF23]. Generalization [MM21, CHS20, TV24, ZD21]. Generalized [ACS22, AMZ24b, AMZ24a, BPR20b, CHZ21, CLS21, CECCCR23, CDMM22, DTM20, DN20a, FGM21, HWCJ21, HCCC21, HH24a, HbX24, Hun21, HK23, KS23b, LNSZ22, LD22, LWW23b, LZ24a, LWWY24, LN22, MF21, Moo22, NB23, SVWC22, SSV23, WCC^{+}19, YT24, YP21, YYZB24, ZZ24, Ais20, APR21, ABM^{+}22, AH22, AAHK24, BLBH21, BZHB23, BBSV21, BK24, BEJ20, BMS22, CMVV20, Cao20, CM22, CCCMH24, CCL20, CDS21, CA23, CDJ20, CFB24a, CFB24b, DL21, DS22b, DOP21, DC20b, DHH24, DDRS21, DMTW22, EFJK22, ER20a, GCV19, GWHZ23, GANU21, HMA24, HLLZ24, HCCZ24, HSMT20, HOOT20, HAD24, HL24b, HZL24, gHtSjT24, HTOC20, HTZ21, IMT20, IMT23, JA22, JDBT22, JWS23, JZ24, Kes23, KUA20, KTK21, KPS21, LZF20, LLY21, LWLD23, LY24b, LCGW24, LNP19, MDW23, MD22, MLYZ21, MSM24, MS20c, NV20, NS24b, OH23]. generalized [PH21, QSW24, RRS24, RG20b, RR22c, SuR20, SC21, SCK22, SK24b, SZZ21, SA23, SZA24, SLLP24, TBA22, TSG22, UGG^{+}20, WTWZ20, WTL20, WLZ23, WDPY24, XZZ^{+}20, XYAH23, YCXW24, Yan22, YC23, hYbJzJ21, YCD^{+}20, ZvS23, ZUZ20, ZLG20, ZGJ21, ZDL22, ZRQ22, ZJBY23, ZWL21, ZCCK24, ZCSS24, ZZL20, Zhu20b, Zhu23, Zhu24, ZKP21, ZY22, dSFSS24, AVC22, AGS^{+}23, ANM24, BGG^{+}21, KVHC23, LTL^{+}24, SYS24, SVC20, TVSC20]. generated [ABA19, JL21, PP22, Son24]. Generating [Ros23, Ant20, IpAmK20]. generative [MWLY24]. generator [DDKD19]. generators [GGPP21, HMS22, Har21, LLGZ20]. generic [KLMP21]. genetic [BYE20, LHE24, San21]. genus [Cho24a]. genus-0 [Cho24a]. Geomechanical [VTBM21]. Geometric [JA22, WGM21, WY21b, dS24, AI22a, CCM22, DBSS^{+}21, DJ24, DFK20, EK24a, MST22, Men24, NP22, PA21, PA24b, TTLB22, VZ21, WWS^{+}21, ZZ21, ZJBY23]. geometrical [vdBS21]. geometries [AK22a, ACEIJH22]. Geometry [ZS20a, Kes23, PMVB22]. geostatistical [CGSVBB^{+}24]. Geostrophic [GQR23]. geothermal [EGM23]. GEQUIP [SYS24]. Gerber [XZ22]. GFDM [UGG^{+}20]. gH [RPQ24]. gH-differentiability [RPQ24]. Ghana [NK22, ÜY20]. Gibbs [BCA20, BCGAR22, MDSRFO24]. Gibbs-Wilbraham [BCA20]. Gibson [BFQ21a, BFQ22, BFLQ24]. Gini [BKMO21, BM22a]. Ginzburg [ZZwS21]. given [HF24, LZW^{+}22, LHG^{+}24]. GKB [GWHZ23]. Global [AF23, GS22, OEH23, Qiu20, SCK22, WFZ22, AJR22, BAB20, Bha23, BB20a, CLL21, DH20, ELTD23, EHVM22, EHV22, GBM21, JJR21, Juk20, KS24, NKBJ20, PS24a, SNMK21, YY24b]. globalization [BFF22]. globalize [MPM23]. Globally [WSYPD24, ALNK24, HL24b, YZ21a, ZSHD24]. glottal [SH21]. gluing [BBAO23]. GMRES [HHM22a, JWSG22, NT20, ORS23, SHZ23, YJN20]. GMsFEM [CHP21, WCFH21]. GMsHDG [Moo22]. Goal [DC23, MVY19, Yüc21, LY22a, YLH23]. Goal-oriented [DC23, MVY19, Yüc21, LY22a]. Golub [EJHR24, RU22]. Gompertz [Pro23]. goodness [ZG20b]. Gordon [Cal23, MDMS22, Bac19, HR23, YQZS21]. Gordon-Schrödinger [HR23]. governed [CLHL20, DLC21, GYZ22, JHZY20, LLZ24a, LT21, MS22, SZA24, TY23, WZZ21a, WWZ24, WZC22b]. governing [ACD22]. government [WQ20]. governs [JA22]. GPGPUs [SŁK^{+}22]. GPM [KDLY22]. GPM-based [KDLY22]. GPU [ARK20, BSMM20, LZH21]. GPU-acceleration [LZH21]. GPU-based [BSMM20]. GPUs [MO20]. GQVIP [HTZ21]. GQVIP-generalized [HTZ21]. grad [GANR22, GRVZ23, Sha24]. grad-div [GANR22, GRVZ23, Sha24]. graded [CM24b, DKMT24, HHL20, NMMH23, SZ23]. Gradient [DXM^{+}20, RPQ24, SKZ23, ACK23, AORS22, BFQ24, Ben24, BBBK24, CKY23, CLS21, DAMA22, FH24, FPR^{+}24, GDZN23, HKM24, JYYL23, Kar20, KAM^{+}23, LMM22, LD22, LWLW23, LLL24a, LWG24, LDC20, LZWL23, LHL21, LH20, NP23, RJ22, Ros23, RWW22, SL20, SWN23, Tak24, TY23, WY20, WAH22, WZC22b, WSL^{+}23, XDP24, XD23, XDX24, YL22, ZOW^{+}21, ZW24, ZJBY23, ZYL20, ZYZ24, ZHW20, ZM21b, IBW^{+}20, SvG22]. Gradient-based [RPQ24, WZC22b, WSL^{+}23]. gradient-type [LZWL23]. gradients [DSCF24, GW24, MPPR23]. Gram [LZ22, LSZ23]. Granger [PCWY20]. granular [LYL^{+}24, TK24b]. Graph [dLNO21, AAMR24, KL23a, LM24, SW24b]. graphical [BGR20]. graphs [ABFSC24, ADMV22, CES23, KV20, Luc23, RLGGAV21]. grating [CLM22b]. gratings [CLM20, CLM21]. Gray [HRT20, ZWZ^{+}23a]. greatest [BWS20]. Greedy [ZL24, MW22, WLZ22]. Greek [ACD22]. Greeks [CST20, CPS20, MYC22]. green [LGMC20, LLG^{+}22, MKN20, RRS24]. Greenwood [SAZW24]. grey [DXWC22]. grid [ABFSC24, BL23, CLX21, DDF23, HW21a, HC20b, HC21, KSVA22, LZH21, LLLL20, OWV24, RFF23, RFGGM20, XLY22, YZL24, ZMOW23, ZXC22]. grid-based [DDF23]. gridded [MJ24]. grids [DN20a, KDOT24, LMLB22, NMST21, PFG24, SK23a, WOV22, WLH23, vtW24, YCG^{+}23]. grooving [CL24a]. Gross [Bro23, EGG22, SRMD22, WW23]. Grossberg [AB23]. ground [EGG22, SSV23, VSSV20, ZW24]. groundwater [ADKH20, AK21, LGLC20]. Group [RHM22, WCLW20, HA21a, LZ24a, LYYZ24]. grouping [FBJ20]. groups [WGM21, WA21]. growing [BB21a, CG20]. growth [CL24b, LD21, MDD21, RS22b, SCZ20, VORD20, WHC^{+}20, XL20, ZSVL20]. GSAV [Hu24]. GSURE [Six20]. GSVD [ACR24]. guarantee [JK21a, ZWH^{+}20]. Guaranteed [NL23a, LV23, ZCZ23, ZSZ24]. gyrocenter [BIZ20]. gyroscopic [IM24].

H7 [BYWA22]. Haar

[ASA^{+}21, HMA24, SSVS20]. Hadamard [DP24, WOV22, WQ22, ZJBY23]. Hahn [DO22]. half [BR20a, ER20a, KS21]. half-line [BR20a, KS21]. half-normal [ER20a]. halving [MP24]. Hamilton [BK21b]. Hamiltonian [Ant20, Bha21, BCCW22, BB20b, ELS21a, HB22, JWG21, SYS24, WW22, YT24, YC21]. Hamiltonians [CLMO23]. Hammerstein [CAN23, MKN20, RRS24]. hand [MDY20]. handling [HJC24, KFK^{+}24, SBKR24]. Hankel [BZH24, MWHC20]. hard [SZZM23, SD24]. hardening [KPD23]. hardware [DDT23, TDD23]. harmonic [CLHL20, FFS24, GN24, HYW20, MMM20, WYH22, WTP^{+}21, XWXZ22, XZL20, YJN20]. harmonic-measure [GN24]. harmonics [BC23, CTGV20, FV24]. Hartley [ZSY24]. Hartley-Ross [ZSY24]. harvesting [MBKK24]. Hausdorff [SS20b, VZ21]. having [BA22, Ozk21]. Hawkes [PLZ21, JJ24, KH24]. hazard [NS24b, ZG20b]. HDG [CMZ22, Fab22, HS24b, SM22b]. Heart [Dax24]. heat [ACDV22, AAA21, AA24c, BS23, BM24a, CLL20a, DZ22, EGM23, FAVM19, FAVM20, HTIA24, KS23a, KMN20, LT21, SBC20, SDC20, SDLZ23, SSV23, Tam20, TKT^{+}24, VSSV20, WCL22, YYBL24, YCH^{+}21, YLS23, ZHWD20, ZC21, ZX23a]. heating [ZHWD20]. heavenly [YT24]. Heavy [NK22, SZZM23]. Heavy-ball-based [SZZM23]. hedging [CN24, KKJ22, YWTL19]. Hele [JGLH20, WWJ22]. helium [CKT21]. Helmholtz [CW20, DL21, DWZ20, DV21b, EOT22, GLLZ21, HLW22, KF24, Kim20, KZM21, PSWZ21, ZSMG20]. help [VORD20]. hemivariational [HJO20, HW21b, HbX24, Hun21, JLM24, LX24, MYZ23, WWH21a, XL23]. Hemker [HO22]. Hénon [ATO22]. hepatitis [Uça23]. Hermite [AH23, BL21, CHS20, DGL24, DIP^{+}22, DDNZ21, DP24, FAUU20, JK21b, KDOT24, MJ20, WY21b, WQ22, YSBT20, Zag23]. Hermite-Hadamard [DP24]. Hermitian [GHC^{+}20, IMT20, IMT23, JLZ23, Mia20, RL23, WYD20, ZYZ24]. Hessenberg [DDKD19, GS22, HSMT20, NKBJ20]. Hessian [AB22, HH24c]. Heston [GPHAPR21, HCH^{+}23, hKhWjH21, MNH21, YLWW21, ZZK21]. heterogeneous [AKN24, AVC22, BABR20, BM20, CS21, CHP21, FBJ20, GJNR21, KVHC23, Li24a, LP23b, LV20, MF21, Mia24, OKE24, SY24, SP22, SD21, SCZ20, TT20, Ter20, TVAS23, VMP23, VSP23, VVB21, VP24, WCFH21, WLS24, ZCSL21, ZYZ23]. heteroscedasticity [ZZQ21]. Heuristic [ZZ22a, FPLZ21]. hidden [PZL23]. Hierarchical [EK24b, GDZN23, PH20a, LMLB22, PJM21, QDHW21, SJZ20, Wen21a, XDP24]. Higgs [MD22]. High [BCT22a, Bos24, CH20, CV20, CC20, CLL20b, DL23b, DW20, ER20b, FZ21, FDDM24, GPHHA23, JZZH22, KD23, LWWZ24, LY22b, LXY24, MTZ23, NMST21, ON22, OH23, WLL21, XWXZ22, ZFW22, dA23, BT21, BBMK22, BIZ20, CCS24, Cap20, CKA21, CKL20, COT21, CD22, DKM23a, DS20, DBR24, DW23, DWZ20, FGL23, FHM23, GCV20, GZ23, GR22, HJC24, Hu24, JYWZ21, JWG21, KWI24, KK20b, KA21, LGG^{+}22, LLS22, LW23b, Li23, LZ22, LGMC20, LST^{+}24, LL24e, LPLP21, LN23, LLG^{+}22, MH21, MCY23, NS24a, Pat23, PI24, RR22c, SM23, SZL22, ST22, TCGY24, VAMO24, WEEA20, WW22, WSC23, WYC24, WKN22, Yan22, YTD23, YF22, ZuIKA22, ZC20b]. high-contrast [GCV20, KK20b]. high-dimensional [GZ23, HJC24, MCY23]. high-end [LPLP21]. High-order [CC20, CLL20b, DL23b, DW20, ER20b, FDDM24, KD23, MTZ23, OH23, dA23, BT21, BIZ20, CCS24, Cap20, CD22, DW23, GR22, Hu24, JWG21, Li23, LZ22, LST^{+}24, LL24e, LN23, NS24a, Pat23, SM23, TCGY24, WEEA20, WSC23, YTD23]. high-performance [COT21]. high-precision [LGG^{+}22]. high-quality [JYWZ21]. high-tech [LGMC20, LLG^{+}22]. high-temperature [COT21]. Higher [DGL24, DVRG21, DC20b, EGG22, GM22, HZL24, LVW24, SMRN24, SH24a, SSVS20, BCT22b, FK23, HM21a, IMMS21, Kar24, LY24b, LP23b, MJK20, RG20a, SC23b, YVXX24, ZBS22, ZM20]. higher-dimensional [LY24b]. higher-index [HM21a]. Higher-order [DVRG21, SH24a, BCT22b, FK23, Kar24, SC23b]. highest [HC24]. Highly [ADDG22, HYY21, LLYM21, ZOW^{+}21, DP21a, Gao22, GJ23, GGHY24, KZ23a, KZ24, KEKT21, KAD22b, KMT20, MF21, MI20, RG21, XLL23, uDuIZ22, ZuIH20, ZFM22, ZFW22, ZH24a]. Hilbert [BM21, DHF23, HZZ21, HbX24, KWN20, Li20a, Li20b, ORT22, TDL^{+}22, TLD^{+}23, TDAT23, TH24b, WZ23b, ZCCK24]. Hilfer [JBSN22]. Hilliard [CJL^{+}24a, GLHL24, JGLH20, WWJ22, YY24a, YYC24, BS20, BDS20, CG22, DWWZ22, GWWY21, LLZY22, LLX^{+}22a, LLG24, LY21, MTIR22, MBZ23, MZY23, WY20, WWW^{+}24, YOCY22, ZZG20, ZJGZ20, ZOW^{+}21]. Hilliard-Magnetohydrodynamics [WWW^{+}24]. Histopolating [BV23]. history [CFO22]. history-dependent [CFO22]. HIV [Sal21a, TKM21, WFZ22]. HIV/AIDS [Sal21a, WFZ22]. HJB [JG21a]. hodograph [BL21, FPS21, SS23a]. Hohenberg [CVPM22, Lee20, LYK23, LY22b, LC23, WC24a]. Hölder [HN22b, KKA20, YWKM19, YH22]. Hölder-norms [HN22b]. Holling [SS23c]. Holm [LXZ23]. holomorphic [PSZ22, RDH21]. Homogeneous [ŚK23b, ARMP20, CTN22, DCAA22, FH24, HKM24, LZLL20, Wan20b, Wol24, ZR24b, dA23]. Homogenization [MMM20, RH23, RL24, BM20, SSM23, VVB21]. homogenized [FFRS23, GJNR21, RCN21]. Homotopy [XHG22, CZ24, Liu22, NAS22, WT22]. Homotopy-Perturbation-Kaczmarz [XHG22]. Hopf [HVR22a, HVR22b, ML23]. Hopfield [TSL^{+}21]. horizon [EL23, LL24a, YP21]. Hormann [TV24, CHS20]. Hough [RFB22]. hp [GS21a, WSY20, LCH23]. hp-FEM [GS21a]. hp-version [WSY20]. hr [MM20a]. hr-adaptive [MM20a]. HS [KAM^{+}23]. HS-LS [KAM^{+}23]. Huber [Zha21b]. Hueso [MGG^{+}24]. Huff [FPLZ21]. Huggins [CJL^{+}24a, DWWZ22]. Hull [FNH20]. human [ASX21, XL20]. Hurst [AHT21, Liu23, SPMB21]. Hurwicz [GRVZ23]. Hurwitz [ZD21, ZH23]. Hybrid [GLCL20, Kim22, MJR20, PM20a, YZ24a, AKM^{+}24, AZZ23, AD21, AAA21, CKZ21, CLYS24, CDJ20, DMZ23, DVRV22, DAMA22, DNK23, GNGGL^{+}19, GC19, HZ23, HR23, HF23, JT20, JLZ22, JM21, KM20, Ku24, KAM^{+}23, LPLP21, MS20a, MTIR22, OBdSD24, PVV24, PHDD22, RR22a, SSRK23, WTL20, YZ20a, YZM^{+}20, YWZ24, hYbJzJ21, ZSL23, ZIZ21, ZR23, ZQL24, Zhu20b, Zhu23, Zhu24]. hybrid-dimensional [ZR23]. hybrid-mixed [PVV24]. hybrid-variable [HZ23]. hybridizable [CR23, CM20, MF21, Moo22]. Hybridization [CLM22b]. hybridized [CRW20, ZWZ23c]. hydraulic [Kov22]. hydraulically [FHB20]. hydraulically-fractured [FHB20]. Hydrodynamic [CKT21, KLM21, LCD21, ZHY20]. Hydrodynamics [HB23]. Hydrogen [CKT21]. Hydrogen-helium [CKT21]. hydrostatic [DL20b]. hydrothermal [WZC22b]. Hyperbolic [AI22a, AFLP22, AAP24, ATM21, BDPR23, CCL23, Cap20, CDS21, DIP^{+}22, DKM^{+}20, FGN^{+}23, LP23a, LLY21, MDVR22, MY21, PKS23, SK23a, Shi21b, UGG^{+}20, YKBH23, YF22, ZZ20a, ZZ22b]. hyperbolic-transport [AFLP22]. hyperbolicity [DKM^{+}20, KS22c]. hyperbolicity-preserving [DKM^{+}20, KS22c]. hypercube [SM20]. hypercubes [CZH21]. hyperexchange [Shi21b]. hyperfunction [Oga20]. hypergeometric [OVGI20]. Hyperpower [Saj23]. hyperrectangle [DP21c]. hypersingular [LZZ24, XLL23]. hyperspectral [LT24a, WHlZH23]. hypersurfaces [FP21]. hypothesis [ARMP20, MSBSM23, Zha21b]. hysteresis [BGV20, PS21].

I/IV [PL20]. Ice [GDG^{+}21, BBSG23, GBG23, RMG21, RDG23, USTG23, CWT23]. ice-shedding [RDG23]. icing [MBG21b]. ICMM [Kan20]. idea [BD23]. ideal [YCG^{+}23, ZGG23]. identically [FJMP24]. Identification [BB21b, ZGZW23, ARK20, DLC21, DXM^{+}20, GIV20, GDZN23, LCGW24, MDLY23, MLY22, OHK23, RKA20, WCWZ20, YLD19]. Identify [SC23a, YZW21]. Identifying [bYW23, LZLL20, LM21b, YPL20, ZGZ23]. identities [Pet20, Psa24]. identity [CHZ21, LL20a]. IDPCNN [HL22b]. IEFG [ADKH20]. IFE [HL21c]. IFRS [HC20a]. II [ABFSC24, AMZ24b, BC22b, CK24a, ER21, JKW23, LMSZ21, MQ20, PM20a, QG20, RSS24, SS23c, YZ21b, YWW24, ZG20b]. III [JO24b, BFQ21b, YZ21c]. ill [BPR20a, BPR21, BDR23, DH24b, HLY22, Jia20, LHL21, LZG24, MV22, NKBJ20, ORS23, Six20, TSL^{+}21, Wan20a, ZW20, ZZ22a]. ill-posed [BPR20a, BPR21, BDR23, DH24b, HLY22, Jia20, LHL21, LZG24, MV22, NKBJ20, ORS23, Six20, TSL^{+}21, Wan20a, ZW20, ZZ22a]. illiquid [HL21a]. illuminates [NPP^{+}21]. illumination [JMHZ24]. ILU [SFI23]. Image [Pul22, Shi21a, AKP23, BKMA24, BCP20, CKY23, CK20, HW21a, HWG21, HL24b, JMHZ24, JW23, JYYL23, JH24, Kan20, KAM^{+}23, LZ24a, LWG24, LNN20, MHW23, RYH24, RU22, SCY23, SBC^{+}24, WY24b, WY24a, WW24a, WZL^{+}23, WSL^{+}23, YLZ^{+}22, hYbJzJ21]. images [BR20b, DKSQ22, HJS21, HKU22, LLLY21, SvG22, WHlZH23, YZM^{+}20]. imaginary [JS21]. imaging [ACFM21, Ben24, BPRZ21, DHJ^{+}24, JHZD20, YHQW21, GMW24]. IMEX [OBB23, PLS22]. IMEX-DG [OBB23]. immersed [BAIB24, GDA22, GLLZ21, JZ21, LZ19, RS23, ZR24b]. immiscible [AJR22, CS21]. immune [TKM21]. Impact [AMRP22, JLL21, KCJ23, ÖS20, Sur21]. impatient [CJL^{+}22]. Impedance [LW23b, PH22, ACFM21, DHJ^{+}24, Ma24a]. IMPES [CS21]. implement [DDT23]. Implementation [AC21, PH22, BD23, CMFMSQ22, EL22, ERB^{+}22, NRGJ21, NAA20, RCVA24, Tia22, WZZ20]. implemented [Bün20, DRD^{+}23, TDD23]. Implicit [FRR24, SJZ20, AS21a, CK22, CK24a, DVRV22, ELS21a, Fan22c, GS20c, HBGS21, HH24a, Hor20, Hu24, HKL21, Hur22, IKL24, JHG^{+}20, JZZH22, LS22a, LDY22, LG22, LC23, LY23b, MN21, MS20b, MBZ23, MPV23, MDD21, PDBdS24, PFG24, RB23a, SRS21, SLW21, XA21, YX20, YVXX24, ZCF^{+}22]. Implicit-explicit [FRR24, LY23b]. implicitly [JLZ23]. Implied [NW23, GJNM24, LL24d]. Importance [KLS21, BBTW24]. improve [VORD20]. Improved [CX24, DOT23, GRVZ23, GLZ24, GS24, HV23, Men24, SLZD20, BZHB23, Bay22, BJF24, BO22, CG22, GK21b, GWWY21, JYY21, JMLX23, LZW20, QGW21, TKVA21]. improvement [EHVMM23]. Improvements [BV22, Pho22]. improving [JC24]. impulse [BR20b, DMP23, EL23, LL23, LNN20, YT19]. impulses [Zha21c]. impulsive [JVAD22, KCKK24, KMD21, ZZLY23]. imputation [VBRT24, KCL20]. in-flight [BBSG23, RDG23]. INAR [PBR21]. inbreeding [MRG^{+}21]. incentive [WQ20]. incidence [OEH23]. including [DC23, GDG^{+}21, Kan20, Men24]. Inclusion [KWYN23, JLM24, KP20, Khl21, Ois20, PC24, PKMM24, SGD20]. inclusions [KK20b, LR22, LSE24, VMP23, Wen21a]. incomplete [HLR22, MX20, WCK^{+}22, WSX21]. incompressible [CMS20a, CS21, DHLT24, FZ20, HH21a, HCC24, HLA22, HHD23, LJCL20, LZF20, LL24c, Li24b, NMST21, Ngo24, Qiu20, RVX20, Sha24, SW21a, SH21, TCGY24, WLHJ20, WCLY23, YZL24, ZC20a, ZY20, ZWKS24]. inconsistent [BLZL22]. Incorporating [LLL24a]. increase [CFO22, WCK^{+}22]. increasing [AB22, HW24b]. increasing-Hessian [AB22]. Incremental [BTC21, FS20, JHW21]. indefinite [CXS20, CUUdB24, DL20a, JW23, JSRdS23]. independence [CMJN22, LL22]. independent [FJMP24, ZSLG21]. Index [AD21, CLM22a, HBGS21, HM21a, HSB24, HBLZ21, JB21, KL23b, MCW24, SL21, SPMB21, SB24b, Wan24, YCXW24, YL21b, ZZM20, ZZQ21]. index-1 [HBGS21, YCXW24]. indexed [SPS21]. Indifference [CNS21, GKV21, YL21a]. indirect [CCJ20, HL23, WDY20]. individual [CY20]. induced [ABM^{+}22, BHPV23, HF21, IIS^{+}21, NY24a, SCY23, UV24, VNA22, ZHWD20, ZS20a]. induction [KPD23, WYH22]. inductionless [LDM23, Zha24]. industry [CC21b, LGMC20, LLG^{+}22]. inequalities [AA24a, CL24c, DP24, HbX24, Hun21, ISD23, MYZ23, MS20c, QZW23, SDL22, TV22, TDAT23, TH24b, WWH21a, WQ22, bWLgZbX24, XL23, ZCCK24]. Inequality [YL22, BP21, BLRZ23, EK24b, HJO20, HW21b, HTZ21, IHK20, JDBT22, JLM24, JHZY20, KZ23b, Kov22, LXQ21, LX24, LH23, TLD^{+}23]. Inertial [DL20a, WLCH23, DHF23, DH24a, HWLD23, IAH24, ISD23, JH24, PKMM24, SDL22, SL24a, TV22, TDL^{+}22, TLD^{+}23, TDAT23, WSLW23, WZL^{+}23, YJJW23, ZLZL23]. inertial-relaxed [YJJW23]. Inexact [BKM22, GCV19, BFVY23, GGO21, JW23, KJO23, LLL24a, LZG24, MVY19, WPX23, dSFSS24, dOdO23]. infected [BYE^{+}23]. infection [TKM21, WFZ22]. infectious [Ngo22a]. Inference [BSNL19, BZHB23, DJ22a, DNK23, LTL^{+}24, WTWZ20, WTL20, Zhu23, BLBH21, BC24, CMdMRFG22, DJ24, HK21, NN22, QG20, SN23, WDTW20, WZZ24, YLL20, ZMZ^{+}24, Zhu20b, Zhu24]. Inferences [JB21]. infiltration [NSL24]. infinite [AM22b, Ask21, BM21, HS24a, KWN20, KS21, LXQ21, LP24, OSK20, ONT21, OF20, SuR20, ZLJT22]. infinite-dimensional [KWN20]. infinitesimal [BK21a, FRR24]. Infinity [IMMS21, BHJB21]. inflated [ZTY^{+}20]. inflation [CCHW24]. Inflection [PS24a]. inflow [GO24]. inflow/outflow [GO24]. influence [CC21b, CES23, Wu24, XCWY20]. influencing [LGMC20, LLG^{+}22]. Information [HBMA22, dS24, DXWC22, DR23b, GBA21, LLL24a, MIdMR24, MP21, RCVA24, San21, XWXX22, ZQL24, dHSR22]. Informative [Cho24b]. Informed [DSCF24, ZG24]. Ingersoll [KLM22]. Inheritance [ZFwC20]. Inhomogeneous [HL23, KFK^{+}23, LR22, LSE24, MRC23, MRC24, SC23a, TTLB22, WYC24]. InitDAE [SL21]. initial [ALNK24, Bic23, CLL20a, DRR22, DCP22, DC24, GO20, GS21c, GJ21, KKJ20, MHHV20, OWV24, QX24, SV23, SK23a, SLP20, SKVA23, TGS21, TZ20, YPL20]. initial-boundary [GS21c, SK23a]. injection [BEK^{+}23, Kap21]. injectivity [YJZ23]. Inner [ALKH21, KL23b, NP23]. inner-outer [KL23b]. innovation [Din23, GDZN23, LGMC20, LLG^{+}22]. innovations [AKN24]. innovative [DLL21, XL20]. input [LCGW24, MDLY23, XDP24]. input-output [MDLY23]. input-output-error [LCGW24]. inputs [DYZ24]. insider [RCVA24]. inspired [BDPR23, ZWX24]. instability [Mas21, ZW21]. instable [FFT20]. installment [JCY20]. instances [CPRV22]. instantaneous [KCKK24, KMD21, Zha21c]. insulation [TSA23]. insurance [BD22, LLS21, LSY24, ZYZ23]. insurers [SYG19, WZJQ21, YCX20]. integer [CLW^{+}22, PS24b, Yam20]. integer-valued [CLW^{+}22]. Integral [Hri22, PP24, AC21, ABP24, ABBB23, All24, AL16, Ans21, AV21, BBAO23, BVDB20, BLN21, BBBU22, BMS22, CJP20, CHXL20, CLX21, CF20, CS24, CAN23, CJM20, CLHL20, CST19, DS20, DR23a, DDM20, DK24, DMNO24, DT24, ER20b, FL23, FO22, FMRS24, Fro24, Gao22, Gol24, HVMS22, HAD24, HLLW21, KN20, KM20, KM22, Kaz21, KAMW21, KXW23, KAA22, Kum22, LBBB23, Li20a, Li20b, LHHW21, LHW22, LG23, LH21, LHH21, MH21, Ma24b, MH22, MS20a, MK24, MKN20, MP24, MH23, MA21a, MS20c, MVFAMF21, Naj20, NP22, NAA20, NAS22, Pie20, PP21, RDHA20, RRS24, SM20, SS24a, SD21, SL24b, Ust21, WLZ23, WX22, WQ22, YSN24, YTMK19, YWZ24, uDuIZ22, ZE23, ZBS22, ZUZ20, Zha21c, ZG22, ZZ22a, ZLCT22, ZFM22, ZFW22, ZH24a]. Integral-balance [Hri22]. integrals [ADDG22, CL24c, CJL^{+}24b, DP24, GJ23, GST20, HS24a, HJL20, IMT20, KZW22, KEKT21, LCD21, LZZ24, LPP23, LC24b, Maj23a, ONT21, RG21, XZZ^{+}20, XLL23, ZuIH20]. integrands [XLL23]. integrated [ER20b, MJK20]. integrating [SSRK23]. integration [CCETZ21, CGL22, DP21c, EPL21, GJ23, HA21a, HKL21, JHG^{+}20, LJ23a, LM21a, Mas21, MCL22, NMST21, ORT22, DFMK20, SRS21, TGS21, WGM21]. integrations [GLZ24]. integrator [Cso20, EL22, WW22, WC24b]. integrators [ACC23, BM22b, CCZ23, Cal23, CCETZ21, CFdD23, JdlCD20, LZL21, WA21, dA23]. integro [ADZ20, AEST21, ASA^{+}21, AA24a, BJB20, BR22, BR20a, CF20, CHH20, Dar20, DRR22, DOR21, DO22, DCAA22, GJN20, HWRM24, JVAD22, KPS21, KS21, MW20, MA19, ÖKG23, Pat23, PL24a, RO20, RS24a, RS24b, RB19, RGA20, RCM20, SES20, SESH23, SM22a, SYD21, SZA24, SKVA22, TSG22, WEEA20, WH24, YYY21, YTD23, YZM23, ZM20, ZLG20, ZCZ24, ZS22, ZTS21]. integro-differential [AEST21, ASA^{+}21, AA24a, BJB20, BR20a, CF20, CHH20, Dar20, DRR22, DOR21, DO22, DCAA22, GJN20, JVAD22, KPS21, MW20, MA19, ÖKG23, PL24a, RO20, RS24a, RS24b, RB19, RGA20, RCM20, SES20, SESH23, SM22a, SYD21, SZA24, SKVA22, TSG22, WEEA20, WH24, YYY21, YTD23, YZM23, ZM20, ZLG20, ZCZ24, ZS22, ZTS21]. integro-differential-delay [KS21]. integro-differential-difference [Pat23]. integro-partial [ADZ20]. integrodifferential [RVS^{+}22, Vab23]. intelligence [Tia24]. intelligent [HCCW20, Zha20b]. intensities [KH24]. intensity [JO24b, SACQ21, Wan21b, WYH22]. intensity-based [Wan21b]. inter [GMSR18, GMSR21, GHF24]. inter-arrival [GHF24]. inter-failure [GMSR18, GMSR21]. interacting [LSE24, RL24]. Interaction [ZAC^{+}23, AESW22, AGMZ22a, AGMZ22b, BHPV23, KSZ21, LML21, Svá23, VS23, YCX20, ZSH20]. interactions [AGS^{+}23, AST^{+}23]. interarrival [LN22, PA21]. Interconnections [UDI20]. interdisciplinarity [MRG^{+}21]. interest [CMRS21, EM21a, WZLT23, ZLL20a]. interface [BGN^{+}20, CGS20, CHLY21, CA23, DC20b, FHM23, GJNR21, GDA22, GLZ24, GLLZ21, HLW22, HDW20, HHM22b, Hes20, HZL24, HL21c, JT20, JZ21, KD23, LS21, LWZ24, LPH23, LZ19, LT21, MTZ23, QS23, RS23, RSRW24, UL23, WGC22, WZ23a, ZDL22, ZIZ21, ZZL20, ZR24b]. interfaces [AET23, BDF21, CPBG23, JL24b, YZL24]. interference [CDTFGV22]. interior [BK21b, EFTW21, HHD23, KKO23, OF20, TXZ24]. Intermediate [BEM^{+}24]. internal [EFTW21, JS22a, ZM21a]. Internality [DDRS21]. International [BSZ21, PCWY20, SS20c]. internet [CZM23, DLL21, QGW21]. Internetg [QJLQ20]. interpolant [BE21]. interpolants [Bos24, CHS20, DFK20, JK21b, KPS24b, LNSZ22, TV24]. interpolating [ADKH20, BEL23a]. Interpolation [BL21, VOJS24, AH23, BEL22, BEIR23, BCA20, BCGAR22, BLV22, BB23, CD23, CDD^{+}24, CK20, CLU23, DDNZ21, DMN24, HHL20, IpAmK20, Ixa24, JLLL23, Kad22a, KA24, KH21, KDOT24, KR21, KRT23, LHH21, Ma24b, MMPP20, MJ20, MN22, MSS20, Mia24, MB24, PL24b, SLZL23, TFM24, VZ21, WYG23, WY21b, Zag23, ZL23a, dC20a, SRAG23]. interpolation-free [Mia24]. Interpolatory [ZK20, AI22a, vdBS21]. interpretation [JVAD22, MMV20, ZYZ24, vdBS21, MSBSM23]. interprovincial [XL20]. intersection [CFB24b, GVT21]. Interval [KNOR20, RKA20, BHJB21, Bra23, IY20, LG23, MP24, MER20, OSK20, RPQ24, Rum20, SBM21, TBA22, TGLa21, UPPZ24, Yam20, YS24a, YCD^{+}20]. interval-censored [YCD^{+}20]. interval-valued [RPQ24, UPPZ24]. intervals [Ngo22b]. intervention [YTT23]. INTLAB [Bün20]. intra [Gök23]. intra-specific [Gök23]. Introducing [Sug20]. introduction [ZDF^{+}20]. intrusion [AET23]. intrusive [AFKM22, KS22c, SPC21]. intuitionistic [TBA24]. Invariant [SYS24, BCCW22, CG22, JL21, MGV23, Sad23, WC24a, ZZG20]. invariants [SYS24]. invasion [KS22b]. Inverse [Bal24, CDS21, EGM23, LZZ20, OR21, Ais20, ASX21, AAMR24, ABA20, AZ20, BJ22, BGH22, BM24a, FMFM20, FHC22, GWHZ23, GHTC21, HMT23, HTIA24, JGW^{+}23, JGZ^{+}24, JLW23, KWI24, KS23a, KPW21, LRP24, LCB20, LLS22, LLL24a, MG23, NNCH21, OKW23, OCS24, RK22, RB24, SN21b, Six20, SQT24, SHV^{+}22, TBA22, WSLJ23, WZZ24, XHG22, XBC21, XBC22, YHW22, YC21, ZYZ22, ZG20a, ZRQ22, ZZMK23, ZMYL23, ZZNL24, ZHMD22, ZJY22, ZL23a]. inverses [DS22a, KS23b, MSS^{+}21, ZCT21]. inversion [LFL20, MA21b, PHDD22, RZ20, SJZ20, Ust23, WLW23, YHW22]. inverted [DNK23, LWLD23, PM20a, WTWZ20]. investigate [CS23]. investigation [PCWY20]. investigations [AAk20, HZMD22, SA24, XZP22]. Investment [Zha20b, CCHW24, Dad20, DXW21, HCH^{+}23, JLL21, MX20, PZL23, WZJQ21, WRZW21, XLZ20a, YCX20, ZLL20a, ZZK21, ZQL24, ZGL20]. involving [Den24, IuR23, LCD21, LC24b, OWV24, Pat23, PP21, SESH23, XLLA21, ZZZL21, dHSR22]. ion [ZAH^{+}24]. ion-acoustic [ZAH^{+}24]. IP [Kan20]. IPCDGM [CW20]. IPDPGM [CW20]. IRNN [PN21a]. IRNN-Iteratively [PN21a]. irreducible [LZLL20, Miy22]. irregularities [SSL22]. isentropic [JZZH22]. isochordal [RB24]. isochordal-viewed [RB24]. Isogeometric [FM21c, Kra22, SD21, WDY20, YCH^{+}21, ZZ23, BRLS23, CWLM22, HZL24, JYWZ21, PJM21, SGD20, ZDL22]. isolated [CDW20]. isolating [KLMP21, WX21]. Isomeric [BP22b]. isometries [AQ20]. isoptics [RB24]. isothermal [HMWY23, IKL24]. isotropic [LLL^{+}22b, YY24b]. issue [BSZ21, DFH21, FMRS24, TE24]. issues [DD23]. Itô [BVDB20, LG22, MA21a]. Iterated [QX24, BPR20b, GW24, RSS24, ZF24]. iteration [Axe20, Dax24, Fan22c, GRVZ23, HV23, Hua21, HC23, LDY22, LW22, LWL22, LZ21, LF21, LWW^{+}22b, LLL22a, MDW23, MVB20, SRBK23, Sha23, TZWG21, WYD20, WSC23, WL24a, XWZ24, XA21, YYT24, ZLV20, ZCSS24]. iterations [GBM21, GCV19, JC24, LC20, Mia20, ZL23a]. Iterative [ACK23, HVM22, HL22b, MDLY23, OA20, Ust23, ZL23b, ANR20, ALKH21, BKMA24, BVDB20, BBMK22, BCT22a, Bic23, BM24a, BDR23, BHB21, CCGT22, Cho24b, DMP23, DC23, DYM21, DHLT24, EHVM22, EHV22, Gol24, GWHZ23, HSB24, HVR22a, HVR22b, HWLD23, KL23b, LHL19, LZLL20, LZ20b, LM21b, NKBJ20, Ned22, PI24, RDHA20, RTT24, Saj23, SBKR24, TH24a, Wen21a, WLW23, ZCY^{+}20, ZGZ23, CD20b, KLZ^{+}24]. iteratively [Kan20, MG22, PN21a]. Ito [KAS23]. IUSAC [KLZ^{+}24]. IV [PL20]. IVP [Vab21]. IVPs [MRC23, MRC24].

J [AGMZ22a, AA20a, AA21b, AA21c, BPR21, BKMO20, FAVM20, GMSR21, GAA21, KZ24, LLG^{+}22, MHY21, MR25, YHC20]. Jacobi [BK21b, Bre23, HR22, KKJ20, Nov24, XBC21, XBC22, YTMK19]. Jacobi-Newton [Bre23]. Jacobian [KSPC22, NNJ24]. Jacobian-free [KSPC22, NNJ24]. Jaeger [BMS22]. jamming [BKMA24]. January [Ano21c, Ano21i, Ano22c, Ano22i, Ano23c, Ano23i, Ano24h]. Japan [SMR23]. jets [YCG^{+}23]. job [SL23]. Joint [GJM^{+}23, TVB21, ACR24, MLS22, SLP20, ZYB21]. Jones [AD21]. Jordan [Wör22]. Joseph [CHLY21]. Journal [ZZW^{+}24b, Sug20]. July [Ano21-32, Ano22-33, Ano23-33, Ano24-30]. jump [AHB21, AUAA20, Boe20, BitH21, CG20, HL21a, Hai23, MW20, ÖS20, PLZ21, PSS24, SPS21, XCY24, YLWW21, ZCZ23, ZSZ24, ZR24b]. jump-diffusion [AHB21, Boe20, BitH21, CG20, HL21a, Hai23, MW20, PLZ21, PSS24, SPS21]. jumps [DLS20, HSYZ24, KAMW21, LL24a]. June [Ano21-33, Ano22-34, Ano23-34, Ano24-31]. justification [CJCDLL22].

Kaczmarz [BLZL22, GHTC21, JZY22, gK21a, gKZ21, gK23, MW22, MS24b, NK24, WLBG22, WCMW23, WX20, WLZ22, XHG22, XY24, ZWZ23b, ZL24, ZH24b]. Kaczmarz-like [WCMW23]. Kaczmarz-Tanabe [gK21a]. Kaczmarz-type [gKZ21, ZWZ23b]. Kahan [RU22, EJHR24]. Kalman [KTK21, LSZ22, NRGJ21, SMR23]. Kannan [BP21]. Kansa [CKA21]. Kantorovich [BCV24, LN21]. Kantorovich-type [LN21]. Kaup [KAS23]. Kawahara [MGA24]. Kawarada [ZS20b]. Kawasaki [ZCYP20]. Kawaski [LLZY22]. KdV [XJK19]. Kelvin [DYF24]. Kernel [CES23, AAHK24, AAA20, CZHQ23, CD23, CDL24, CDLR24, GNGGL^{+}19, HVMS22, HCCW20, KXW23, LHS^{+}24, MKN20, MDE24, XTL23, YWZ24, uDuIZ22, ZLCT22, ZFM22]. Kernel-based [CES23, CDLR24, MDE24]. kernels [ARF23, Ans21, MO20, PL24a, PZ21, RRS24, YTMK19, ZBS22, ZFW22, ZS22]. kind [Ade22, AID24a, AID24b, BJB20, BBSV21, CS24, DRR22, DDRS21, JKKK20, KN20, KXW23, MH21, Ma24b, Miy22, NAS22, QZW23]. Kinetic [RB23b, DRB22, IIS^{+}21, TFR^{+}20, ZY20]. Kirchhoff [BRLS23, MCD24]. KKT [SH24a]. Klein [Cal23, HR23, MDMS22, YQZS21]. knot [BEL22]. known [CGSVBB^{+}24, SVJG24, YSUY24]. Kolmogorov [CW21a, CMW24, XFW24, YAE23]. König [LP20]. Kontorovich [Yak21]. Kopel [LZL^{+}23]. Korteweg [DD23, LF24, SC21, SCK22]. KP [MGA24, XJK19]. Krasnoselski [Wen21a].

Krasnosel’skiĭ [BM21]. kriging [GMCJGM24]. Krylov [BEJ20, BK20, ERB^{+}22, HJ20, HN22a, IS23, Jia20, KSPC22, Mia20, MW21, RMM22, TGS21, ZFwC20]. Kumaraswamy [WDTW20]. Kupershmidt [KAS23]. Kuramoto [AL20, FC23]. Kurchatov [MAA^{+}22]. kurtoses [MTP21]. Kutta [ASML23, KFK^{+}23, KFK^{+}24, Moi24, MRC24, ST22, ÁGKM21, Bha21, BP22b, ESS21, FZ21, FCS22, GS20c, HR21, HCH20, HW24b, HKL21, LF21, MM21, MF24, MRC23, NS24a, NNJ24, RAN20, SL20, SW23, WH24, YLK24]. Kuznetsov [BGdlR24].

L1 [WYJ23, YLLA23, ZVT^{+}23]. L1-type [YLLA23]. L2 [QWY24]. L2-1 [QWY24]. Lab [Yam23]. LabVIEW [AKP23]. laden [HBPV21]. lag [Fra21, LQS23]. lagging [ZC21]. Lagrange [BSdCCGV23, BCA20, BCGAR22, CDJ20, LXQ21, OBO23, WHZ24, dC20a]. Lagrangian [AFLP22, AAP24, BBSG23, CFdD23, DG22b, DV21a, HBPV21, JHYZ24, LS21, LMLB22, WLZY21, dA23]. Lagrangian-type [JHYZ24]. Laguerre [DGL24, MSM24, SZA24, ZXK23, ZXK24]. Lambert [dS24]. Lambert-Tsallis [dS24]. Lamé [Dom21]. laminar [HF21]. Lanczos [ACR24, BEJR22, CST19, DRR^{+}21, JJR21, JLZ23]. Lanczos-based [DRR^{+}21]. Land [CWT23]. Landau [Wol24, ZZwS21]. LANDSAT [DKM23a]. Landweber [Sha23]. Lane [RR23a]. Langevin [AM21]. Language [AST^{+}23, AGS^{+}23, AMZ24b, AMZ24a, MGV23]. languages [AGMZ22a, AGMZ22b]. Laplace [MR25, BA23, DZLC20, EGG23, Fan22b, GJ19, HMT23, LV23, LPP23, Mil22, MR23, MA21b, MCMR23, NK23, WZ24b]. Laplacian [AAMR24, BB21a, CDCVV22, DAM23, MWBM22, MMF24, SS24b, XLLA21]. large [ABFSC24, AS21a, BTC21, BBMK22, BD20, CW20, CNRR20, DKSQ22, Dax20, Dax24, DYM21, DV21a, Els21b, Fal21, Gra20, GW24, HOOT20, HL24b, HLY22, JLZ23, JZY22, LZL21, LZD21, LWLW23, Li24a, LLX^{+}22b, MS20b, MW22, NNJ24, Qiu20, RSS24, RYH24, WXZ21, WLCH23, WY24b, WAH22, Wen21b, WSL^{+}23, hYbJzJ21, YJJW23, Zha21a, ZOW^{+}21, ZLZL23, ZWZ23b, ZH24b, dOdO23, GQR23]. large-scale [BTC21, DYM21, DV21a, HOOT20, HL24b, HLY22, JLZ23, LZL21, LZD21, LWLW23, LLX^{+}22b, RYH24, WXZ21, WLCH23, WY24b, WAH22, Wen21b, WSL^{+}23, hYbJzJ21, YJJW23, Zha21a, ZLZL23, dOdO23]. large-size [DKSQ22]. largest [ZYZ23]. laser [AGH20, Ben24, KPD23, ZHWD20]. lasso [HJC24, AGNGG^{+}23]. last [BYE^{+}23]. latent [HK21]. lateral [CJM20]. Latest [DFH21]. Lattice [Ri22, Ri24, AHB20, AHB21, BZ20, CMRS21, DG21, HBPV21, ZY24, WGM21]. lattice-based [AHB20, AHB21]. lattices [LLL^{+}22b]. Law [Cho24b, LTL^{+}24, BFQ24, CJCDLL22, CGSVBB^{+}24, LLGZ20]. laws [AAP24, CCL23, CW24, GAA19, GAA21, Kri21, KS22c, MGA24, MN22, RHM22, RK23, YKBH23, YF22]. Lax [LLS22]. layer [AA22, AA24b, DL21, GHJM24, HLYZ24, Hur22, SMDVA22, TZBK21]. layered [ABA20, GJGDPR23]. layers [GJNR21, MJK20]. layout [RLGGAV21]. LCP [EFJK22, WYW^{+}24]. LCPs [tSjTsJ24]. LDG [YZZ22]. LDU [WLDZ24]. leader [AKMO21]. leader-following [AKMO21]. leading [BG23]. leap [SS24b]. leap-frog [SS24b]. LeapFrog [JF22]. Learning [CLMO23, PS22b, YCCZ24, ZG24, AVLMRMGG23, AGST23, AST^{+}23, AGNGG^{+}23, Bha23, CLS20, CC21b, CLPZ21, Fan22a, FRT23, FPR^{+}24, GS20b, HQL21, KA24, KDO20, KJO23, LLK23, LGLC20, NY24a, NPP^{+}21, OBO23, PWL23, DP23, SSM23, SKZ23, TN24, VTBM21, WHC^{+}20, XY22, XFW24, XLZ20a, ZZZ^{+}22]. learning-based [Bha23, PWL23]. Least [BC23, CCL23, CKL23, Din23, Hes20, JM21, MLY22, SAL21b, AKP23, BGIR24, CGMM20, CLM22a, CUUdB24, DMNO24, HM21a, HSW21, HG22, HH24c, KL23b, LCGW24, MH22, MI20, MDKG20, RR22b, Saj23, SL22, WCLW20, WYZZ22, WD20, YLXY23, YZZ20b, ZZWS22, ZJY22, ZFH20, ZH24c, ZWX24]. Least-Squares [SAL21b, CCL23, CKL23, JM21, AKP23, HM21a, HSW21, HG22, MH22, MI20, MDKG20, YZZ20b, ZFH20]. Lebedev [Yak21]. Lebesgue [BE21, BCKSV24, JL21]. Ledoit [NRGJ21]. Lee [HZZ21]. Leffler [DdO22, GHF24, LHES21, YFW20]. Legendre [ASADB23, DOR21, DO22, HZC^{+}23, HCH^{+}23, KM20, NM24, RO20, SES20, ZCZ24]. Legendre-Gauss [DO22]. Legendre-tau [NM24]. length [BCM20, SACQ21, Zag23]. lengths [KPS24b]. Leptokurtic [GMK20]. Leray [HÇYK24]. Leray- [HÇYK24]. level [BSMM20, BCP20, CL24b, DKM23a, DDKD19, Els21b, EWCQ23, GR22, Gra20, LZH21, Ngo22a, Ngo22b, Sha21, Sha24, WCK20, WHlZH23, ZZwS21]. level-dependent [BCP20, DDKD19]. level-set [LZH21]. Levenberg [FHC22, LZZZ21, dOdO23]. Levi [FS21]. Levitin [Hun21]. Lévy [AZ22, AZZ23, Dar20, XZ22, ZYY20, ZRQ22, ZX23b]. liability [ZCJL20]. Liao [LH20]. Lie [HA21a, IMMS21, Lem24, RHM22, WGM21, WA21]. life [AZ22, BD22, ER20a, ER21]. life-contingent [AZ22]. lifetime [BJP24]. lifetimes [BC24, DJ24]. lifting [GVT21]. light [CLM20, CLM21, FGP24, JMHZ24]. light-controlled [FGP24]. like [AKM^{+}24, AAA^{+}23, BEL22, CAJ24, DGL24, GH20, HZ23, HL22c, JG21a, Lu24, Luc23, WSLJ23, WCMW23, ZZ20b]. Likelihood [DJ24, CD23, CMJN22, HK21, WD20, XGYY23, ZMZ^{+}24]. limit [BCH20, Cal23, KSZ21]. limited [AST^{+}23, HL23, ZMYL23]. limited-aperture [ZMYL23]. limiters [YF22]. Limiting [BHJB21, LYM22, MJ20, NS24b]. Lin [HR22]. line [BR20a, KS21, LWYZ21, LLK23, MGA24, RRS24, SW24a, Zho20]. Linear [ACG20, JZW22, ZVS21, AC21, AS20a, AH23, ATM21, AAKA24, AS21a, ARMP20, AA22, AA24b, ACC23, ALKH21, BTC21, BLZL22, BR20a, BDPR23, BIJ21, BPR20a, BPR21, BDR23, CCZ23, Cao20, CZHQ23, CXS20, CHH20, CG22, CPBG23, Cho24b, CYY23, CJLL21, CUUdB24, CNQRR22, DDUF22, DK24, DHZ24, Din23, Dom21, DHQZ24, DWZ20, DFK20, ELTD23, EOR21, ER20b, EBVB21, Fal21, FHY^{+}21, FM22, FO22, FRT23, GJDL20, GPHHA23, GEU24, GHC^{+}20, GYZ22, GBA21, GLLZ21, HJ20, mHLZ20, HSRA24, HCH20, HH20, Hu22, HKL21, Hua21, HC23, IpAmK20, Jia20, JZY22, JCLC20, JJ24, gK21a, gK23, KFK^{+}23, KJO23, LBBB23, Let24, LLYM21, LWL22, LW23a, LT24a, LZ21, LWT24, LYL23, LY23a, LP23b, LL22, MF21, MH21, MMJ21, Mar19, MV22, MDY20, MPV23, MW22, MMJM22, MORO20, Min20, MA19, Miy22, ME20, MRC23]. linear [MRC24, MFA21, MO20, ORS23, Pat23, PBR21, PN21b, PNS21, QQSZ24, RHM22, RKA20, RS22b, SRS21, SN23, SLZD20, SWW23, SDW24a, SL24b, TGLa21, TV22, UGG^{+}20, VLR23, WYTL19, WY20, WCW20, WCFH21, WDS22, WMH22, WWJ22, WMZZ23, WZ24a, WL24a, WLB^{+}21, Wu22, XHLD21, XFW24, XWZ24, XYAH23, Yam20, YC23, YLJW24, YYZB24, ZSLG21, ZYZ22, ZHY20, ZGJ21, ZOW^{+}21, ZL22, ZZ22a, ZSHD24, ZZM20, ZWZ23c, ZH24b, ZS22, ZMZ^{+}24, ZM21b, ZKP21]. Linear-time [ACG20, Mar19]. linear/bilinear [GLLZ21]. linearization [Alh24, DC23, DN20b]. Linearized [HLLZ24, NNJ24, ABHK20, DTM21, FC23, LCZL24, SMKV21]. Linearly [ELS21a, LF24, ANN21, CG20, HLLZ24, Hu24, JHW21, JMLX23, MBKK24, RK22, SZZ21, WSLW23, YLK24]. Linearly-fitted [LF24]. linesearch [RPQ24]. link [CZM23, MQZ^{+}24]. link-load [MQZ^{+}24]. linked [BD22, ZYY20]. Liouville [BBAO23, CM24b, CL24c, DP24, GS21c, Gu21, IuR23, LLLL20, Mag20, Zha21c, ZS20b]. Liouville-Caputo [GS21c]. lipid [CZY22, LLT^{+}21]. Lipschitz [BAB20, CLL21, HQS20, LLX^{+}22b, MHHV20, SNMK21, YZ21a]. Lipschitzian [DHF23, ZZZL21]. liquid [CSLQ22, GDG^{+}21, ZaZWZ23]. liquidity [CCHW24, CN24]. Liu [AK22b, HZZ21]. Liu-type [AK22b]. Lizorkin [RR22d]. Ljusternik [YL21b]. LM [GGO21, KMZ21]. LMI [EBVB21]. load [MQZ^{+}24, WFZ22]. load-dependent [WFZ22]. loaded [BBAO23, Grz21]. loads [GH23]. Lobatto [APR21, BCGAR22, DO22, LL24e, RS24c]. Lobatto-type [RS24c]. LOC [Yam23]. Local [AAA^{+}23, ACEIJH22, BB20a, BKMO21, CLM22a, mHLZ20, HSW21, LWA^{+}22, LSY23, LC24a, MS22, MAA^{+}22, SDW24b, YSFS22, ABA19, ABFSC24, AAB^{+}20, AMRP22, Bac19, BGG^{+}21, CZHQ23, Che21, CJLL21, EO24, Ela23, GRS20, GBM21, HCH^{+}23, KZM21, LSC21, LXZ23, MYC22, NL23a, NRV21, SNMK21, SV24, TGJY22, WTX24, WY21a, YX20, ZC22]. local-stochastic [HCH^{+}23]. localization [HLW20, LRP24]. Localized [Ku24, NA21]. Locally [LZ23a, NNJ24, HAD24, Kim20, MHHV20, dOdO23]. locating [Che22, CNYC23]. location [Ben24, FPLZ21, LR22, LMSZ21, SVJG24, TBA22, TBA24]. location-scale [LMSZ21]. locking [DHZ24, GYZ22]. locking-free [DHZ24, GYZ22]. LOD [Yam23, ZDW24]. log [ARMP20, BCLS24, CM24a, LLY21, OF20]. log-determinant [OF20]. log-generalized [LLY21]. log-logistic [CM24a]. logarithm [TSMZ20]. Logarithmic [Ans21, LGC23, MDG22, YQZS21]. logistic [CM24a, JHW21, WXZ21]. loglinear [CMP21]. lognormal [BJP24, Mil22]. Lommel [BBBU22]. Long [QWY24, WW22, YZ24d, EPL21, FM21b, FY22, LML21, MP23, Mas21, Ngo22b, WZC22b]. long-memory [FM21b]. long-range [LML21]. Long-term [YZ24d, MP23, WZC22b]. Long-time [WW22, EPL21, FY22, Mas21]. longevity [XCY24]. longitudinal [TVB21, VNA22, Zha20c, ZMZ^{+}24]. lookback [AZ22, WZ24b]. loop [GL24, ZCJL20]. loops [NY24b]. Lorentz [FC22, LL20a, VNA22]. Lorentzian [ZZLY23]. Lorentzian- [ZZLY23]. Lorenz [DS21]. loss [CL24b, GL20, LHS^{+}24, MSK21, NY24a, YLL20, YLXY23, ZZK21, Zho22, vdMOB22]. loss-based [YLL20]. lot [FPGSR21]. Lotka [MWW21, YBM23]. Love [BRLS23]. Low [ACT22, GHG24, KF24, Lee23, WZ21, XWXX22, ZZW^{+}24a, AA24c, Bro23, CMFMSQ22, CWY21, CZM23, Dax20, Dax24, DRD^{+}23, DYM21, FDDM24, GGPP21, HR21, HLR22, JMHZ24, KCK24, MHW23, MJ24, PN21a, PS22b, QXZF23, SvG22, WDPY24, WN23, YZM^{+}20, YLZ^{+}22, ZWH^{+}20, ZMFL20]. low-dimensional [PS22b]. low-field [SvG22]. low-light [JMHZ24]. low-Mach [FDDM24]. low-order [ZMFL20]. low-power [DRD^{+}23]. Low-rank [GHG24, KF24, XWXX22, CMFMSQ22, Dax20, Dax24, DYM21, HLR22, MHW23, MJ24, PN21a, QXZF23, WDPY24, YZM^{+}20]. low-storage [HR21]. low-tubal-rank [ZWH^{+}20]. lower [HWCJ21, JS22b, Juk20, KWN20, LZF20, WLCH23]. lower-order [HWCJ21]. lower-rank [WLCH23]. lowest [LHMT20, MWBM22, MM22b, WZ20a]. lowest-order [LHMT20, MWBM22, WZ20a]. LP [Hun22, Yam21a]. LQ [GL24]. LS [KAM^{+}23, SL24b]. LS-SVM [SL24b]. LSD [LS24]. LSIAC [PR24]. LSIAC-MRA [PR24]. LSNN [CCL23]. LSPIA [RJ22]. LTI [SV23]. Luc [BPR22]. Lucas [DOR21]. Lundberg [LLL22a]. Lyapunov [MHY21, CVPM22, DS22b, LWD20, MHY17, NYL23, SLZD20]. lymphocyte [TKM21].

MA [OR21]. MAC [LS23, YZL24]. MacArthur [HE24]. MacCormack [Ngo22b, Ngo24]. MacCormack/Crank [Ngo22b]. Mach [FDDM24]. Machine [LGLC20, NPP^{+}21, RDH21, AVLMRMGG23, AGNGG^{+}23, CLS20, CC21b, KJO23, LPLP21, VTBM21, WHC^{+}20, YYT24]. machining [MGV23]. macro [ZVS21]. macroscale [TVAS23]. macroscopic [ANM24]. magma [KV19]. magnet [KKO23]. Magnetic [JKJ^{+}23, Wan21a, WYH22, GMW24]. magneto [ZHY20]. magneto-hydrodynamic [ZHY20]. Magnetohydrodynamic [KBK^{+}22]. Magnetohydrodynamics [WWW^{+}24, AZL21, HHD23]. magnitude [LN22]. Magnus [ACC23, Cso20]. Magnus-type [Cso20]. Mainardi [SBC20]. maintenance [Asa24, LPLP21, MFH24, Ret24]. major [WHC^{+}20]. Majorization [CW22, TWQ^{+}23]. Majorized [KCL20]. making [ASTK24, dHSR22]. management [QGW21, Ret24, YTT23, ZCJL20]. Manakov [BBTZ20]. Manifold [FGL23, SA23, WDPY24, ZS20a, ZJBY23]. manifolds [BFF22, SAL21b]. Mann [BM21, Wen21a]. Mann-type [BM21]. many [BEL22, CJL^{+}22]. many-to-many [CJL^{+}22]. map [OSK20, ONT21, SAN^{+}24, TTK23, USTG23].

Mapped [RRS24, MMPP20]. mapping [JMHZ24, LZ23a, XWXZ22, ZXZ24]. mappings [HTOC20, MLS22]. maps [CGP22]. March [Ano21d, Ano21j, Ano22d, Ano22j, Ano23d, Ano23j, Ano24c]. marching [QDHW21, YX20, YLJW24]. Marcum [BBSV21]. margin [LHS^{+}24]. marginal [XWXX22]. marine [BBBU22]. market [CN24, Hai23, HC20a, LLS21, LYL^{+}24, MSBSM23, MX20, PGG20, PCWY20, SZWH21, SZL20, XGYY23]. marketing [YC20a]. markets [BO22, DKL20, HL21a, WZ23c, Zha20a]. Markov [BAPPF23, DDKD19, FSR24, Har21, MYC22, MR20, PZL23, VTBM21, YLWW21]. Markov-modulated [YLWW21]. Markovian [GMSR21, AUAA20, AGGAVGG23, BBTW24, GMSR18, KK20c, KK21]. Markowitz [MSBSM23]. Marotto [LZW20]. Marquardt [FHC22, LZZZ21, dOdO23]. Marquardt-type [dOdO23]. Marsden [Bic21, FL24]. Marshall [BSNL19]. Maruyama [BGN^{+}20, DHKV20, HN22b, KAMW21, KW23, MWW21, RS22b, TM24, YWKM19, YH22, YZ24c, ZLG20]. masks [Han24]. mass [AAA21, BPDC20, DZ22, KV19, Lee20, LF24, MPR24, SSV23, WZY23, ZCY^{+}20]. mass-conserved [WZY23]. masses [PSZ22]. Matched [LLZ24b, CJL^{+}22, DL21, HLYZ24]. Matching [CMdMRFG22, CJL^{+}22, LLZ24b, WLH23]. material [BS23, CDCVV22, PKPD21]. materials [ANM24, DYM21, HHL20, LYL23, LH24, SD21, VVB21, ZC21, ZCSL21]. Math [AGMZ22a, AA20a, AA21b, AA21c, BPR21, BKMO20, FAVM20, GMSR21, GAA21, KZ24, LLG^{+}22, MHY21, MR25, YHC20]. Mathematical [BGV20, FFT20, FAVM19, FAVM20, HLF19, JKJ^{+}23, MRG^{+}21, Pet20, VAAR22, AAA21, BMS22, BYE^{+}23, DDT23, OOO22, SS20c, YBAE23, dPdMQL23]. Mathematics [ZZW^{+}24b, Sug20]. MATLAB [Bün20, HS24a, LMAV24]. MATLAB/INTLAB [Bün20]. matrices [AEST21, APS22, AM22b, ACG20, BAPPF23, BD20, Bon24, CLST20, CUUdB24, Dax20, Dax24, DOP21, DKM^{+}23b, FLMP24, FSR24, LM24, LNP22, Lu24, MPR24, Mia20, Miy21, Rod21, Rum20, RMM22, Shi21b, SS20b, WJVG24, WLDZ24, Wu22, XBC21, XBC22, Yan22, YC23, YHC20, YHQW21, YC21, ZZMK23, ZYZ24, ZCT21]. Matrix [BC21, CCZ23, MSS20, AAMR24, AHK24, AIDAJ23, ALKH21, BPDC20, BEJ20, BEJR22, BAD^{+}22, BK20, CMFMSQ22, CW21b, CZM23, DWWZ20, DR23a, DIAJ^{+}22, DIP^{+}22, DS22b, DdO22, DEH^{+}24, DL20c, DYM21, EFJK22, EMR21, Fan22c, FM22, FLMR20, FGM21, FHM23, FMRT20, GHG24, GHF24, GS22, HJ20, Han24, HMRZ23, HV23, HLLL23, HW24a, HC23, JJR21, JLZ23, JZM^{+}23, JC24, KHAW21, KCS21, KCL20, LKM20, LLLD20, LDY22, LW22, LWL22, LW23a, LJ23b, LWG24, LWXL21, LZ24b, Lu24, MPM23, Mas21, MR20, MDY20, NRGJ21, NL23b, OA20, OR21, PH24, PT20, PJM21, PHDD22, QXZF23, Ri22, Ri24, RL23, Sad23, SI23, SH23, SK24b, SB24b, TSMZ20, TWQ^{+}23, Uga24, WYTL19, WL24a, WLDZ24, WLC23, WLZ22, XWXX22, XLZ20b, YLZ^{+}22, YYZB24, YZLL22, ZE23, ZYZ22, ZD21, ZH23, ZBH24, ZZZL21, ZWL21, ZCSS24, dF20]. matrix-based [ZYZ22]. Matrix-free [CCZ23]. matrix-valued [BEJR22, ZH23]. matrix-vector [DEH^{+}24]. maturity [ZCZ23, ZSZ24]. Matusita [LWLD23]. Max [Dum24, BG21, BCV24]. max-min [BG21]. Max-product [Dum24, BCV24]. maxima [BSH24]. Maximal [MS24b, ZH24b, ZTY^{+}24]. maximization [AAk20, CES23]. Maximum [SZ20, TCGY24, XGYY23, YF22, CD23, CFB24a, CFB24b, HZZ24, Kop20, NS24a, SY24, WSY22, WWC23, WD20, WLS24, ZWZ23b, ZWJ22]. Maximum-principle-preserving [TCGY24, NS24a]. maxterm [ABM^{+}22]. Maxwell [MÖ21, BHRW22, CMZ22, CH20, DMTW22, FC22, HYW20, LLL^{+}22b, MMM20, NM24, WZZ21b, WYZ24, WZH^{+}21, XZL20]. May [Ano21e, Ano22k, Ano22e, Ano23k, Ano23e, Ano24i, Ano24d]. MBDyn [ZAC^{+}23]. MBE [WYJ23]. McKean [GGHY24, RS22b]. McKean-Vlasov [GGHY24]. MCMC [SZL22]. Mean [BCVV20, KTK21, LHHW21, LWZF21, LHHW23, LET23, TBA24, AHB20, AHB21, AK20, CCS24, ÇK24b, GK21b, HH21c, hKhWjH21, LS24, MMV20, NN24, PA21, PA24b, Psa24, RS22b, RL24, SYG19, TBA22, WZJQ21, WZLT23, Wu24, YSUY24, ZCJL20, Zha20c, ZSZ24, LY24a, MSBSM23, SZK20]. mean-covariance [Zha20c]. mean-downside [WZLT23]. mean-field [RS22b]. Mean-MF [LWZF21]. mean-mixture [NN24]. mean-reversion [Wu24]. Mean-reverting [LET23, AHB20, AHB21]. Mean-square [LHHW21]. Mean-variance [TBA24, SYG19, WZJQ21, ZCJL20, LY24a, MSBSM23]. mean-variance-utility [hKhWjH21]. means [CST20, FMAV23, JZY22, MR20]. measurable [IYA^{+}23]. measure [GN24, LZZ20, LWLD23, Pie20, RDHA20, TSG22, ZY22]. measured [Seg23]. measurement [CZM23, HSRA24, LSZ22, Mit20, VS21, WCL22, ZM21a]. measurements [DDF23, MQZ^{+}24, NYL23]. measures [DDRS21, FLZ21, GS24, HQT20, HBMA22, KKA20, KDLY22, MP23, MRI22, MR20, PZ21, Ver20]. measuring [Sur21]. mechanical [CLMO23, FMTB21, HA21a, WA21]. mechanics [WJVG24, XL23, dA23]. mechanism [CJL^{+}22, SZWH21, WQ20, YCX20]. mechanisms [ZZWS20]. media [ACS22, AJR22, AVC22, AGST23, ASTK24, CS21, CRRB21, DRB22, FHB20, GCV20, GIV20, HET24, HW21b, IKL23, KLM21, KVHC23, LZWC21, LP23b, Moo22, NY20, PH20a, RH23, RB23b, SEAS22, Sal21a, SIS^{+}22, SVC20, SV24, TT20, TVSC20, TVAS23, VSP23, WZZ21b, WCFH21, WLH23, YN21, ZZWS20]. median [TBA22, TBA24]. medium [GM24, Kov22, MHM23, WCL22]. medius [HY21b]. Mellin [AI22b]. mellitus [AVLMRMGG23]. melt [LLZY22]. Memoriam [BPR22]. Memory [Sal21a, WL21, AAHK24, DAM23, FM21b, GEG23, MN20, QDHW21, Vab23, WC21, ZL23b]. Memory-dependent [WL21]. memoryless [LZD21]. memristive [MP23]. memristor [AB23]. memristor-based [AB23]. mental [CSZP20]. Mercedes [JW24]. Mercedes-Benz [JW24]. merging [LL21]. meromorphic [Che22]. Merton [Boe20, BitH21]. mesh [AESW22, Axe20, CM24b, FLT20, HHM22b, HW20, Kog21, KS22b, MW20, MBG21b, NMST21, NMMH23, OBB23, PR24, RS23, SMDVA22, TK24b, WWWY23, WLM24, YZ21c, YZ21b, YZZ22]. mesh-free [HHM22b, TK24b]. meshes [AMS23, BCKSV24, CLU23, DKMT24, Gua20, HDW20, HO22, HKL21, HC24, Hur22, LLZ24a, LHMT20, LZ23b, LCZL24, PR24, QWY24, QDHW21, SRAG23, SZ23, TCGY24, WTL21, WWH21b, WYZ24, YZ20b, YZ24b, YCH24, ZWJ22, ZW22]. Meshfree [NSL24, NS24c, SM20, Son24]. meshing [KSG^{+}22]. Meshless [uDuIZ22, AID24b, DOT23, FGN^{+}23, Li24b, NA21, Sha23]. metabolic [AS20a]. metamaterial [HLYZ24]. metamaterials [HLF19]. Method [EOT22, HIM24, MWW21, MJR20, VTBM21, ADKH20, ADZ20, ALNK24, AAP24, ACK23, AKM^{+}24, ATM21, ATHW21, AFKM22, AAB^{+}20, AZD21, AA22, AA24b, AMZ24b, AMZ24a, AZL21, AK21, AID24a, AAM20, ACPV24, AS21b, Ars22, AA20a, AA20b, AA21b, AA21c, Axe20, AI22b, BSdCCGV23, BJB20, Bac19, BTBR21, BT21, BVDB20, BCR22, BR22, BBBU22, BEJ20, BEJR22, BDF21, BJF24, BuR24, Bic23, Bog21, BGN^{+}20, BLRZ23, Bre22, BDS20, Bre23, BR20b, BDR23, CGP22, CCL23, CGS20, CWW23b, CWW23a, CC24, CTGV20, CGO24, CDLR24, CLX21, CF20, CRW20, CR23, CAN23, CC21a, CG20, CHH20, CHY20, CKA21, CW21b, CZM23, CLS24, CDW20, Che21, CO19, CCL20, CJL^{+}24b, CM20, CHLR22, CK23, CJLL21, CKL23, CD20b, CJ23, CGM21, CNRR20, CDJ20, CD22, CS23, CZ24, DMP23, DZZ24]. method [DS20, DDM20, DVRV22, DG22a, DMN24, DDF24, DC20b, DL20a, DLC21, DHZ24, DHH24, DW23, DZLC20, DYF24, DCAA22, EOR21, EJHR24, ER20b, EM21a, EEE21, EFTW21, ETKF23, EKÇ20, EPL21, EHVMM24, Fab22, FM21a, FFT20, FFS24, FLMR20, FLT20, FW22, FZQ22, FO22, FGN^{+}23, FPR^{+}24, FC22, FHC22, GLHL24, GG24, GBM21, GBM23, GDA22, GEG23, GeOJD24, GÇSA23, GS21c, GK21b, GHC^{+}20, GHTC21, Gu21, Gua20, GJW23, GH20, GH22, GSWH22, GKV21, HB23, HH21a, HMWY23, HLW22, HZ23, HS24a, HSB24, HYW20, HDW20, mHLZ20, HCCC21, HHM22b, He23, HL24a, HCCZ24, HBPV21, HS24b, HCCSCGAM24, HTJC24, Hes20, HR22, HR23, HSMT20, HE24, HWRM24, HOOT20, HL21b, Hri22, HQS20, HC20b, HH21c, HW24b, HWD^{+}24, HS20, HQL21, HSC21, HLA22, HL22c, HLY22, HC23, HLYZ24, HWAT22, HH24c, IAH24, IuR23, JJR21]. method [JZ22, JGLH20, JF22, JLZ23, JHG^{+}20, JG21b, JLZ22, JLW23, JL24a, KSPC22, gK21a, KLH21, KS23a, KM20, KK23, Kaz23, KLM22, KH24, KP22, KPS24a, KKJ20, KWYN23, KLS21, KCK24, KXW23, KW23, KAA22, KD22, KSVA22, KL23b, KS21, KRV24, LMM22, LBBB23, LSC21, LS21, Lee20, Lee21, LL24b, LRZ23, LLFT20, LZ20a, LCB20, LZWC21, LHHW21, LLZC21, LZH21, LHL^{+}22, LD22, LDY22, LWW22a, LWL22, LHW22, LXLL23, LW23b, LHHW23, LWW23a, LVW24, LJL24, LLGC24a, LLG24, LLZ24a, LLGC24b, Li24b, LZZZ21, LLW22, LGY24, LCW20, LHMT20, LDC20, LZ20b, LHH21, LW21, LMLB22, Liu22, LWWN22, LZWL23, LZ23b, LY23a, LDQF23, LL24f, Liu24, LC24a, LYZ24, LCZL24, LHL21, LPP23, LPP24, LH20, LXF21, LN21, LT21, LN23, LR24b, MGG^{+}24, MkCZ20, MF21, MH21, MGD22, MCD24, Ma24a, MM20a, Mag20, MMS^{+}23, MSW23, MS20a, Mam23, MLYZ21, MHM23, MV22, MTIR22]. method [MM22b, MMF24, MH23, MT20b, MW21, MW22, MR23, MR25, MA21a, MG22, MG23, ME20, MDKG20, Moo22, MJK20, Naj20, NS24a, NNJ24, Ned22, NY24a, Ngo22b, Ngo24, NK24, NSL24, NS24c, NY20, NM24, NAA20, NAS22, NNL^{+}24, NT20, NZ21, NP23, OKW23, OF20, OGR22, PL21, PSZ22, PVV24, PH20a, PHDD22, PWL23, Pho22, PH21, QS23, QSW24, QX24, QXZG22, QWY24, QDHW21, Rab21, RR22a, RSS24, RKS^{+}22, RS24a, RMG21, RB19, RS23, RR22b, RCM20, RG20a, RG20b, RG22, RR23b, RMM22, SuR23, SM23, SEAS22, SRS21, SMDVA22, SS20a, Sha21, Sha24, Sha23, SGY22, SW24a, SN21b, SXX21, SSL24, SLXF24, SSS24, SL20, SYS24, SFMdD23, SSRK23, SM22b, SWN23, SZA24, SSV23, SV24, SW23, SP22, SHV^{+}22, SGD20, SFI23, Tan20, TH24a, TM24, TXZ24, TXS21, TT20, Ter20]. method [TLD^{+}23, THS24, TFR^{+}20, TK24b, TKVA21, UGG^{+}20, Ust21, VV23, VTB21, VBRT24, VdG20, WZZ20, Wan20a, WCK20, WGS20, WZZ21b, WWH21a, WWH21b, WYH22, WZC^{+}22a, WGC22, WMH22, WWJ22, WYZZ22, WZ23b, WSLJ23, WSC23, WLH23, WMZZ23, WPX23, WCMW23, WZ23a, WTX24, WYC24, WHZ24, WYZ24, Wan24, WL24a, WLM24, WKC20, WSY20, WY21a, WSHW22, WLW23, WZD^{+}24, vtW24, WDY20, WLS24, XHG22, XJK19, XL23, XFW24, XWZ24, XLY22, XYAH23, XFW23, XXL21, XH20a, XXH^{+}22, XSL22, XWL24, YSN24, YHW22, YN21, YWKM19, YLFT21, YH22, YS24a, YWZ24, YMC22, YSFdRdP20, YDC20, YZZ20b, YZZ20a, YZ20b, YZ21c, YZ21b, YZZ22, YZM23, YCH24, YCH^{+}21, YLLA23, YLS23, YZ24c, YZ21d, ZE23, ZvS23, ZBS22, ZM20, ZTZZ20, ZWZ^{+}23a, ZSMG20, ZLG20, ZZ20a, ZGJ21, Zha21c, ZZwS21, ZSL22, ZZ22b, ZLCT22, ZLJT22, ZGZ23, ZMYL23, ZSL23, ZDW24, ZC24, ZF24, ZC20b, ZIZ21]. method [ZJY22, ZR23, ZY24, ZYL20, ZLV20, ZX23a, ZaZWZ23, ZYZ24, ZCZ24, ZXC22, ZH24b, ZZ20b, Zho20, ZWL21, ZWJ22, ZZ23, ZX23b, ZLC24, ZCSS24, ZHZ24, ZZL20, ZWKS24, ZR24b, ZVS21, dFG21, AVC22, AGS^{+}23, ANM24, BGG^{+}21, KVHC23, MWBM22, SVWC22, TVSC20, VSSV20, SVC20]. method-based [TLD^{+}23, WPX23]. methodologies [PH20b]. Methodology [PKPD21]. Methods [DVRG21, VAAR22, AC21, ASML23, ASB22, ACS22, Ant20, AAA^{+}23, ALKH21, BLZL22, BK21a, BLN21, BCT22a, BFVY23, BR20a, Bha21, BPRZ21, BIJ21, BLyS20, BK21b, BIZ20, BHB21, BP22b, CK22, CK24a, CLL21, CMS20a, CL22, CMZ22, CM24b, CPP23, CLL20b, CST20, CPS20, DP21b, DCY24, DOT23, DMST21, DDF23, DFH21, DAMA22, DO21, DCP22, DC24, Din23, DYZ24, DW20, DYM21, DLZ22, DRR^{+}21, ESS21, EN20, ERB^{+}22, EHV22, EHVMM23, Fal21, FZ21, FHY^{+}21, Fan22c, FCS22, FHM23, FH24, FRR24, FJ21, GLYS21, GGVRB22, GJL23, GHJM24, GS20c, GPHAPR21, GPHHA23, GJ21, GS22, GLLZ21, HJ20, HM21a, HHM22a, HN22a, HV23, HR21, HCH20, Hu24, HLMY24, Hua21, HY21b, HL21c, HHD23, HZ24b, JS21, JT20, JWG21, JG21a, JZY22, JL24b, JKW23, JZ21, gKZ21, KZW22, gK23, KN20, KKA20, KHAW21, KD20a]. methods [KFK^{+}23, KSZ21, Ku24, KS22c, LS22a, Lee23, LGC23, LMV24, LMZ22, LJCL20, LLX^{+}22a, LW22, LY22a, LWW23b, LZ23a, LW23a, LJ23b, LP24, LWWY24, LWZ24, LZ21, LZ19, LSY23, LP20, LG22, LC23, LTZ24, LXY24, LT24b, LY23b, LXZ23, LHE24, LXYL24, MW20, MTZ23, MH22, MM21, Min20, Moi24, MF24, MMPM22, MRC23, MRC24, MFA21, MBG21b, MB24, MS24b, ON22, OBdSD24, ORS23, PKE20, PP24, PV22, PH20b, PW23, PI24, PS24c, QZW23, RAN20, RKA20, RVX20, RU22, RB23a, RR23a, SLW21, SvG22, SW24b, Six20, SYD21, SLLP24, SDW24b, Svá23, TZBK21, TE24, TZWG21, TGJY22, TCGY24, TGS21, TTLB22, Vab22, WLHJ20, WK20a, WYD20, Wan21a, WZ21, WLZ23, WWZ23, WC24a, WWZ24, WKN22, WH24, WX20, WLZ22, XY24, XW21, XDP24, XLL23, XD23, YZ20a, YX20, YQZS21, YQZS22, YKBH23, YCXW24, YTMK19, YPL20, YPLM22, YZ23]. methods [YS24b, ZSY24, ZFwC20, ZG20a, ZKT20, ZR24a, ZL24, Zha24, ZFM22, ZFW22, ZH24a, ZS22, ZL23b, ZH24c, ZDF^{+}20, ZM21b]. methods-convergence [MS24b]. metric [BGR20, BPRZ21, HKM24, JZ24, KL23a, PH24]. metrics [GHM21, OPK23]. MF [LWZF21]. MFEM [WZ20a]. MFS [CNYC23]. MHD [DHLT24, EKÇ20, HQS20, LL24c, LDM23, Qiu20, TSA23, ZC20a, Zha24]. MIASS [EFTW21]. micro [PA24a, PZY^{+}24, ZVS21]. micro-macro [ZVS21]. micro-plastic [PA24a]. microenvironment [SCZ20]. micromorphic [SMRN24]. micropores [RH23]. microtemperatures [BFLQ24, LQS23]. middle [AAB24, GN24]. middle-thirds [GN24]. Midpoint [LT24b, GH22, ZvS23]. midpoint-based [ZvS23]. mild [bWLgZbX24]. mildly [Bre23, Six20]. Milstein [HWD^{+}24, JL24a, KK20c, KK21, LY20, LHW22, LLW22, MP21, PSS24]. Milstein-type [HWD^{+}24, KK20c, KK21]. Mimetic [BL23, CC20, Mia24]. MIMO [FGM21]. min [BG21]. miniature [JKJ^{+}23]. minima [Juk20, PS24a]. Minimal [Fal21, AM22a, Bic21, CUUdB24, HF24, LZ23b]. minimax [LZ20b, SDW24a, YL21b]. Minimization [CDCVV22, ACK23, BKM22, BR20b, CES23, GXC24, LWX24, PN21a, QXZF23, SL24a, TWQ^{+}23, XLZ20b]. minimized [GS20c]. minimizers [DP21a, FPCV^{+}22]. minimizing [CCL20, GHG24, JHZD20, LZW^{+}22, LHG^{+}24, WCK20, YLH23, ZGJ21, Zho24]. Minimum [EWT24, PZ21, XWZ24, JGW^{+}23, JGZ^{+}24, MMM23, ZCZ23, ZSZ24, CMP21]. Minimum-energy [PZ21]. mining [CZ20, LWC20]. Minkowski [BL21, MLS22]. minorants [BAB20]. minterm [ABM^{+}22]. Minty [Hun21]. misaligned [BL23]. miscible [BEG^{+}22, BEM^{+}24, LHL^{+}22, WZ20a, YCH24]. MISSILES [ELTD23]. missing [GS24]. Mississippi [Tia24]. mitigation [KS22c]. Mittag [DdO22, GHF24, LHES21, YFW20]. Mittag-Leffler [DdO22, GHF24, LHES21, YFW20]. Mixed [HH23, JL24b, LLK23, DFMK20, WCFH21, YL22, YLS20, AKN24, AVRLRMGG23, Ara20, Ars22, BBAO23, BDS20, CE24, CGO24, Cha24, CLW^{+}22, CRRB21, CGM21, CN24, DMN24, DAZ23, DLZ22, DMTW22, DN20b, GZ24, GGVRB22, GG24, GPHAPR21, GS21b, HZ24a, HWCJ21, HCCC21, HCCZ24, HSRA24, HZZ21, HZ24b, Hun21, HTZ21, Hur22, HWAT22, ISD23, JL21, KM20, KKJ22, LS21, LJCL20, LCZL24, LNP19, LN22, LR24b, Ma24a, MM22b, MMF24, MO20, Ngo22b, NMF^{+}22, PVV24, PKE20, RGA20, TXZ24, VNV23, WMZZ23, WHZ24, WZZT20, XFW23, YCH24, ZR24a, dSFSS24, SVWC22]. mixed- [HSRA24]. mixed-flow [Hur22]. mixed-fractional [Ara20]. Mixed-mode [DFMK20]. Mixed-model [LLK23]. mixed-precision [MO20]. mixed-thinning [CLW^{+}22]. mixing [DH20]. mixture [AKN24, BJ22, CDS21, NN24, NK23, Ruj23, HW22]. mixtures [AB22, BFQ21b, CSLQ22, CYY23, DP21a, GHL21, JD22, JS21, NN24, NS24b]. MLP [AGNGG^{+}23]. MMOC [CDL22]. mobility [SW21a]. Möbius [CT20]. mock [DMNO24, Ibr20]. mock-Chebyshev [DMNO24, Ibr20]. modal [HZ24a, TR22, XY22]. mode [Ais23, DLL21, JA22, DFMK20, TK24a, ZLS23, KT24]. Model [AFJGJJ22, FJMP24, HLLL23, LMZ24, MZY23, MB24, AMS23, AZZ23, AAA21, AA23, ACS22, AET23, ANM24, Ara20, AMT23, AK22b, ABO24, AKP23, BSNL19, BRLS23, BG21, BK24, BJF24, BYE^{+}23, Boe20, BitH21, Bün20, CGSVBB^{+}24, CHLY21, CZY22, CLW^{+}22, CCCMH24, CGW21, CL22, CLM22a, CG23, CLS24, CL24b, CEKN20, CLPZ21, CRRB21, CGM21, CPP23, CFO22, CN24, DH20, DTMJ24, DLS20, DHZ24, DXM^{+}20, DHH24, DYF24, DXWC22, EGM23, EM21a, EY23, FM21b, FFRS23, FPLZ21, FPS22, FMTB21, GJDL20, GJNM24, GMW24, GM24, GPP^{+}20, GQR23, GeOJD24, GPHAPR21, GDZN23, GJW23, GGT24, GBA21, GDG^{+}21, GKV21, HÇYK24, HW21a, HYY21, HMWY23, HCCC21, HCH^{+}23, HSGLE19, HA20b, HA21b, HJC24, HTJC24, HE24, HSRA24, HLF19, HQL21, HLLQ22, HLYZ24, Hur22, IKL24, JF22, JMX24, JZW22, JZM^{+}23, JJ24, KLM21, hKhWjH21]. model [Kaz23, KLM22, KMUR24, KYP21, KKKL22, KKJ22, KO22, LHES21, Lee21, LL24a, LL20a, LWC20, LLT^{+}21, LML21, LLYM21, LWZF21, LLZY22, LZL^{+}23, LLK23, LJ23b, LSY24, LT24a, LX24, LL23, LLH24, LNN20, LLL22a, LC23, LET23, LNP19, MP23, Ma24a, MDSRFO24, MST22, MWW21, MS20b, MN20, MNH21, Men24, MBZ23, MER20, MDD21, MBKK24, NEYZ22, Ngo22b, NA21, NY24b, OOO22, OEH23, ÖHS23, PA24a, Pat23, PWL23, PBR21, Pro23, QHPL21, RS22b, RTV20, RG23, Sal21a, SV23, SDC20, SDLZ23, SLK24, SS23c, SMRN24, SZ23, SS20c, SW21b, SYG19, SLW^{+}22, SMR23, SL24b, ŚK23b, TKM21, Tam20, TGJY22, TFK23, TKVA21, TCVC21, TVAS23, UÖA23, UV24, VNV23, VCNR20, VMP23, VAMO24, WCC^{+}19, WK20b, WMF20, WZLX20, WTL20, WCPW20, WDTW20, WWS^{+}21, WZJQ21, WCLY23, WZY23, WYJ23, WY24a, WW24a, WSHW22, WZD^{+}24, WSX21, WFZ22]. model [Wu24, WZZ24, XLKL23, XZP22, XGYY23, XDP24, XL20, XCY24, XH20b, YLWW21, YPLM22, YY24a, YLJW24, YYC24, YOCY22, YHSK22, YCD^{+}20, YAE23, YBAE23, YWTL19, Yu22, ZHWD20, ZWZ^{+}23a, ZCYP20, ZS20a, Zha20b, ZTY^{+}20, ZZK21, Zha21b, ZC22, ZCF^{+}22, ZR23, Zhu22, Zhu23, ZPW^{+}24, Zhu24, ZGZW23, ZKP21, dPdMQL23]. model-based [GMW24, WZZ24, XDP24]. Modeling [GMSR18, GMSR21, PHV24, WLH23, YTT23, ZZ21, AGH20, Ais23, ARMP20, AGMZ22a, AGMZ22b, AMZ24b, AMZ24a, AK21, ABO24, BDPR23, CGCMRP24, EO24, FFT20, Grz21, JKJ^{+}23, KCJ23, Kes23, NK22, NRV21, ÖS20, PLZ21, PH21, RH23, SPC21, TFM24, TVB21, VVB21, WMF20, WL21, WSX21, YLK24, YY21, ZZWS20, ZLZ^{+}24]. Modelling [RCN21, BMS22, KKO23, MRG^{+}21, TK24b, WMK22]. models [AFLP22, AS20a, AKN24, AZ22, AKMO21, ANN21, BN21, BEG^{+}22, BZ20, BGV20, BKMO19, BKMO20, CMP21, Cha24, CBJ20, CECCCR23, Cso20, CES23, DZ22, DJR21, DYZ24, DAZ23, EBVB21, GM23, GMSR18, GMSR21, GMK20, GX24, GHF24, GBA21, HZMD22, HK21, HQS20, KS22b, KMT20, LPY20, LLGC24b, LTF^{+}20, LYL^{+}24, LN22, LWZ21, LWX24, MYC22, MDVR22, MDMS22, MDS20, MLY22, NPP^{+}21, NMF^{+}22, PHV24, PKS23, RKS^{+}22, RKA20, RHA20, SUW23, SN23, SW21a, SJZ20, SYTD24, TKT^{+}24, VORD20, Wan21b, WZL^{+}23, Wen21b, XZ22, YLL20, YLXY23, YHF22, YCCZ24, ZYY20, ZJ24, Zha20c, ZZM20, ZZQ21, ZZWS22, ZCZ23, ZSZ24, ZMZ^{+}24, HW22]. modification [BD24, LLW22, MSM24, PKMM24]. Modified [CW21b, FZ20, LDY22, MORO20, NWD20, SAZW24, Wen21a, AF23, ÁGKM21, BVBO23, BJF24, BGdlR24, CHLY21, CHXL20, CKY23, CL24b, DDRS21, EOT22, GSWH22, GW24, HW24b, HL22c, HWAT22, JGLH20, JL24a, KMZ21, LLFT20, LL20a, LML21, LWLW23, MG23, QH23, RR22b, SS24a, TM24, WYD20, WMZZ23, WYZ24, WAH22, WLZY21, XBC21, YH22, YWL24, ZFH20, ZZ20b, Zho20, ZCSS24]. MODIS [DKM23a]. modulated [CHsL22, YLWW21]. Modulating [WLB^{+}21]. Modulus [LW23a, Fan22c, GSZ22, HV23, HC23, LDY22, LWL22, WL24a, YYZB24]. Modulus-based [LW23a, Fan22c, HV23, HC23, LDY22, LWL22, WL24a, YYZB24]. Molecular [NGNB23, Bra21, LCB20, WCK^{+}22]. MoM [FMRT20]. moment [AFKM22, IS23, JO24b, ZTY^{+}24]. moment-based [IS23]. Moments [DG20, MKS21, SHZ20, BCVV20, DKSQ22, IMT23, KPS24a, Pou21, SLC22, Uga24, Wu24, Xia22, ZYB21]. momentum [TY23, WW23, SRAG23]. money [LWC20]. monic [ZBH24]. monitoring [AK20, HZS^{+}24, HQL21, LMSZ21, MÖ21, MLS22, SCHS24, YWW24]. Monk [HLLQ22]. monolithic [CK22, CK24a, RSRW24]. Monotone [HL22a, NY20, AKM^{+}24, ASB22, Bre23, FPS21, HET24, JH24, LLX^{+}22b, NSL22, RYH24, SDL22, SL24a, TV22, TDL^{+}22, WAH22, Wen21a, YN21, YJJW23, ZLZL23, ZJBY23, ZCCK24]. monotonic [ZZZL21, Zhu20a]. Monotonicity [BBSV21, DHF23, JDBT22]. Monte [VTBM21, DMZ23, HJ23, Har21, MMS^{+}23, XW22, XFW24, XHW24, YYBL24]. Monteiro [ACT22]. Moore [AAMR24, BFQ21a, BFQ22, BFLQ24]. Morley [AMS23, LWWZ24]. Morley-type [AMS23]. morphing [BAIB24]. Morse [YL21b]. mortality [ÖS20, Wan24, XCY24]. most [JNCY23]. motility [HH24a, UV24]. motion [AHT21, AS21b, CM22, CN24, HLR22, IM24, KV19, KAM^{+}23, LHHW23, Liu23, LLLY21, NP22, NY24b, SPMB21, TTLB22, USTG23, YKAA21, ZHZ24]. motions [CCM22, FKV21, VOJS24, YYY21]. motor [KKO23]. movement [BZ20, HB23]. Moving [SAL21b, ACK21, GLYS21, GDZN23, JA22, LS21, LT21, MW20, MTZ23, Mik23, SEAS22, SMDVA22, SK23a, Slo21, XGYY23, ZFH20]. moving/evolving [Slo21]. MPDATA [AF20]. MPHT [NWD20]. MPHT-splines [NWD20]. MR [SvG22]. MRA [PR24]. MRI [HL22b]. MS [TS20]. MS-Stability [TS20]. MSV [AA23]. MTH [NBLZ24]. MTRSVD [HLY22]. Multi [CSZP20, DH24a, DR23b, JHZD20, RDG23, YLD19, dHSR22, ASML23, ADEEY21, AAB^{+}20, AET23, BLBH21, BZHB23, BC21, BC22a, BSMM20, BCCW22, CMW24, CAN23, CL22, CLPZ21, CJ23, DJR21, Din23, DAZ23, DDKD19, FM21a, FHY^{+}21, GR22, GMW24, GANU21, Gra20, GHC^{+}20, GDZN23, HW21a, HJ23, HMWY23, HLLY24, HLLZ24, HSC21, JDBT22, JLZ23, JJ24, Kon22, KZM21, LL24b, LLX^{+}22a, LST^{+}24, LNP19, MD22, MDMS22, MLYZ21, Miy22, MER20, PH20a, PCMH20, PCM21, Rab21, RAN20, RGA20, SWN23, Son24, SV24, TKVA21, VSP23, WCW20, WSLW23, XLLA21, XCY24, YZ23, YWZ24, YBAE23, YWTL19, NGNB23]. Multi-armed [DR23b]. multi-block [HLLZ24, WSLW23]. multi-class [DJR21]. multi-cohort [XCY24]. multi-component [HMWY23, LLX^{+}22a, VSP23]. multi-continuum [PH20a, PCMH20, PCM21, SV24]. multi-core [Gra20]. Multi-criteria [dHSR22]. multi-degree [BC21, BC22a]. Multi-dimensional [JHZD20, CMW24, CL22, DDKD19, JJ24, MD22, RAN20, Son24]. multi-domain [YWZ24]. multi-fidelity [MLYZ21]. multi-fractional [MDMS22]. multi-frequency [YZ23]. multi-functional [HLLY24]. multi-Galerkin [CAN23]. multi-innovation [Din23, GDZN23]. multi-level [BSMM20, GR22, Gra20]. multi-linear [Miy22, WCW20]. multi-order [ADEEY21, FM21a]. Multi-parameters [YLD19]. multi-patch [GMW24]. Multi-period [CSZP20, MER20, YWTL19]. Multi-physics [RDG23, LL24b]. multi-point [AET23]. multi-shifted [GHC^{+}20]. Multi-Site [NGNB23]. multi-species [TKVA21]. multi-spline [GANU21]. multi-splitting [CJ23]. multi-stage [CLPZ21]. multi-state [BLBH21, BZHB23, DAZ23, Kon22, LNP19]. Multi-step [DH24a, ASML23, AAB^{+}20, FHY^{+}21, HJ23]. multi-symplectic [BCCW22, JLZ23]. multi-term [HSC21, RGA20, XLLA21]. multi-time [JDBT22]. multi-time-step [LST^{+}24]. multi-trace [KZM21]. multi-wavelet [Rab21]. Multichannel [CK20, LL21]. multicollinearity [MÖ21]. multicomponent [CJLL21, WDTW20, Zhu22, Zhu23, Zhu24]. multicontinua [SVC20]. multicontinuum [AGST23, AST^{+}23, BM20, TVSC20, WCC^{+}19, ZC22]. multicontinuum/multiscale [AGST23, AST^{+}23]. multicriteria [Ret24]. Multidimensional [DZ22, KGV23, CCGT22, EBVB21, ÉM21b, MIdMR24, ZCZ24]. multifactor [MST22, Wu24]. multifractal [LWZF21]. Multigoal [AESW22]. Multigoal-oriented [AESW22]. Multigrid [BLyS20, DKMT24, DBSS^{+}21, He23, KDO20, Liu22, PHDD22, TR22, XXH^{+}22, XWL24]. multigrid-homotopy [Liu22]. multigrid-Thomas [PHDD22]. multihead [GJGDPR23]. Multilevel [CGL22, FL24, HSGLE19, DLC21, HJ23, PSWZ21, XHW24]. multilinear [CWY21, LZZZ21]. multimesh [GS21a]. multimodal [VCNR20]. Multinode [DDF24]. multinomial [AFJGJJ22, CMP21]. Multiobjective [LMM22, CFB24a, CFB24b, HTOC20, HK23, PT24, TTK23, UPPZ24]. multiphase [HLA22, PKS23, SW21a]. multiphysics [LPH23]. Multiple [ACEIJH22, Kon22, BD22, BCT22b, BVBO23, BHB21, CCCMH24, CPRV22, ER20a, ER21, EWCQ23, GZ24, GGHY24, GC19, HZMD22, HA20b, HA21b, HCCW20, Jas22, JVAD22, LMAV24, LLL24a, MDY20, MHHV20, SBKR24, SZZ21, SYS24, WSLJ23, XLY22, XDP24, XXH^{+}22, YL21b, ZZWS20, ZL23b, ZLL^{+}20b, dPdMQL23]. multiple-block [SZZ21]. multiple-continuum [ZZWS20]. multiple-delay [GGHY24]. multiple-dimensional [MHHV20]. multiple-input [XDP24]. multiple-output [XDP24]. multiple-sets [BHB21]. multiplication [BAD^{+}22, NL23b, OA20]. Multiplicative [LRT20, CLM22a, CEKN20, DC24, DP24, JWS23, KW23, LVW24, MLY22, SWW23, WDS22, YLXY23, ZSHD24]. multiplier [WHZ24]. multiply [DMST21, JL21]. multipoint [HCCC21, HCCZ24, MWHC20, ZL23a]. multiquadric [GJ21]. Multirate [HBGS21, BK21a, FRR24, RSRW24]. Multiscale [AVC22, AGS^{+}23, ANM24, CM20, GJW23, KVHC23, Li24a, MF21, PCMH20, PCM21, RFF23, SVC20, SVWC22, TVSC20, TCVC21, VMP23, VSP23, ZCSL21, ACS22, AGST23, AST^{+}23, AMZ24b, AMZ24a, CW20, CHH20, CCL20, CK23, CLPZ21, DW20, DW23, GCV20, HCCC21, HCCZ24, LH24, MkCZ20, Moo22, NSL24, NS24c, PVV24, PH20a, PCWY20, PH21, SJZ20, SP22, SLLP24, Svá23, VTBM21, WCC^{+}19, WCK20, WYC24, YDC20, ZSLG21, ZC21, ZZWS20, ZGJ21, ZDW24, SSV23, VSSV20]. multispin [MMS^{+}23]. multisplitting [GBM23]. multisplitting-based [GBM23]. Multistep [ESS21, MDS20, WH24, Fal21, Min20]. Multivalue [DP21b]. multivalued [Pie20]. multivariable [Dum24, XD23]. Multivariate [Kad22a, Seg21, ZTY^{+}20, ZY22, AKN24, AB22, BSH24, BCT22a, GHL21, HM22, JD22, KDOT24, LLY21, LCGW24, PS22b, SHB24, TVB21, UDI20, VPHE23, WYG23, Xia22, ZKP21]. multiwavelet [AJ21, KCK24]. multiwavelets [Ri24]. Müntz [SES20]. Mutations [HCCSCGAM24].

n [AHPS22]. nabla [HOGO23]. Naghdi [KNM24]. Nagumo [RS22b]. Nakazawa [WZY23]. nano [DYM21]. nanofluid [FFT20]. nanofluids [AAA21]. nanoplates [MJK20]. nanoscale [SDC20, SDLZ23, ZHWD20]. NASCA18 [BJMR20]. Nash [CFB24b, CFB24a, EL23, GL24, OHK23]. Nash-equilibrium [EL23]. natural [AC22, AGMZ22a, AGMZ22b, ZZ21]. Navier [CJL^{+}24a, GLHL24, YY24a, YYC24, BS20, BJF24, CHLY21, CR23, CZ23b, DG22b, DN20a, DN20b, FZ20, GANR22, GRVZ23, HH21a, HJM24, JF22, JZZH22, LZF20, LLZY22, Li24b, NMST21, Ngo24, RR23b, Sha21, Sha24, XW21, XH20b, ZY20, ZR24b]. navigation [ŚK23b]. NCP [ZZZL21]. NCPs [AC22]. NDSS [XWZ24]. near [AZ20, BGH22, LPP24, RSS24]. near-field [AZ20, BGH22]. near-optimal [RSS24]. nearly [ELTD23, Men24, XZZ^{+}20]. necessarily [MVFAMF21]. Necessary [MMM22, MMM23, Sun20, Juk20, MMMM20]. Nédélec [SMRN24]. negative [OK23, ZZWX23]. negatively [ZTY^{+}24]. neighborhood [Slo21]. nematic [ZaZWZ23]. Nernst [DHQZ24, JZ22, LWW^{+}22b, QQSZ24, YSFS22]. nested [HG23, XHW24]. Nesterov [MSW23, Ros23, ZGZ23]. net [LD22, MGL20, SW21b]. nets [DP21c]. Network [KBK^{+}22, MQZ^{+}24, AS23, BHPV23, CQBL20, CCL23, CZM23, CO19, CNRR20, GS20b, GXL22, HHM22b, JLL21, JJ24, Kad22a, KWI24, KPS24a, KMTZ21, LFZ24, Pou24, QJLQ20, San21, SYTD24, TSL^{+}21, WCW20, WYG23, WL24b, XLZ20a, XCWY20, YL22, Zha20b, ZMYL23]. Networks [DSCF24, vdMOB22, AC22, AUAA20, AB23, BD22, BBTW24, CSLQ22, DJR21, GBTS22, GEU24, HLLL23, HOGO23, HB22, HPWM20, Hun22, HH24b, KA24, KSG^{+}22, LDQF23, LL21, MMV20, PGG20, PK23, DP23, SUW23, SMKV21, SDW24b, ŚK23b, TBA22, TBA24, ZG24, ZZNL24, ZZZ^{+}22]. Neumann [EOT22, FP24, NZ21]. Neural [BHPV23, DSCF24, KWI24, KBK^{+}22, WCW20, WYG23, vdMOB22, AC22, AB23, BD22, CQBL20, CCL23, CSLQ22, GS20b, GBTS22, GEU24, HHM22b, HLLL23, HOGO23, HB22, HPWM20, HH24b, JLL21, JJ24, Kad22a, KA24, KPS24a, KLK21, LFZ24, LDQF23, PK23, Pou24, QJLQ20, San21, DP23, SUW23, SYTD24, SDW24b, ŚK23b, TSL^{+}21, WL24b, Zha20b, ZMYL23, ZG24, ZZNL24, ZZZ^{+}22]. neural-network [KPS24a]. Neural-network-based [BHPV23]. neutral [AA24a, Ask21, BJB22, DFFM21, GGHY24, KCKK24, KMD21, LDQF23, LY23b, WXS21, ZS22]. Newell [IYA^{+}23]. Newmark [SFMdD23, TZBK21]. Newton [AA20a, AA21b, AA21c, ALNK24, AAB^{+}20, AA20b, BFVY23, BFF22, Bre23, DN20b, Gol24, HH24c, JG21a, KSPC22, LZ20a, LW22, LZG24, MSW23, MG22, Pho22, RSS24, SKZ23, SD24, UPPZ24, WXZ21, YKAA21, Zha21a, ZJY22, ZZ20b, Zho20, ZWL21, ZCSS24, dC20a]. Newton-based [LW22, ZWL21, ZCSS24]. Newton-correction [LZ20a]. Newton-type [BFF22, DN20b, LZG24, WXZ21, Zha21a]. Newtonian [Kes23]. Next [Tia24]. nexus [WHC^{+}20]. Nicolson [JF22, ADZ20, BSdCCGV23, BDF21, FC22, GH22, LW24, Ngo22b, NA21, WY20, YYBL24]. Nicolson/Galerkin [ADZ20]. Nikol’skij [RR22d]. Nikol’skij-Lizorkin [RR22d]. NIMOC [ZGG23]. nine [EWCQ23]. nine-level [EWCQ23]. Nipah [Evi23]. nitrogen [CSLQ22]. Nitsche [CHLR22]. NLMC [SV24, ZC20b]. NLS [YKAA21]. NN [LWX24]. no [ARMP20, SYG19]. no-short [SYG19]. Noda [GCV19, MDW23]. nodal [BCGAR22, LXY24, WV21]. nodes [BL23, BE21, DZLC20, Ibr20, LL21, vdBS21].

Noise [IIS^{+}21, BR20b, CEKN20, CNQRR22, DCP22, DC24, GDZN23, KW23, LRT20, LVW24, LL23, LNN20, LSZ22, Liu24, SvG22, SLW^{+}22, TKVA21, WHlZH23, ZZLY23, ZVS21, ZGZW23]. Noise-induced [IIS^{+}21]. noise-tolerant [SLW^{+}22]. noises [LCGW24, SLW^{+}22, ZGJ21]. noisy [Din23, HSW21, HG22, LWX24, MP21]. Non [ABA19, BJP24, BKMO19, BKMO20, CV24, CFL21, ÇÖÖ23, CLU23, GRS20, HKM24, LL23, MDKG20, PK24, SV24, SPC21, AGGAVGG23, AAA20, BGG^{+}21, BB20b, Bre22, BDR23, BM22b, CJCDLL22, Cal23, CLL21, CZHQ23, CMdMRFG22, CTN22, CPBG23, CFdD23, DK24, DHF23, EO24, Ela23, Fal21, FRT23, FHH21, GW23, GO24, GHC^{+}20, HVMS22, IKL23, IKL24, JYY21, JJ24, KLM21, KCKK24, KM22, Kov22, KMD21, Let24, Li20a, LWW22a, LLW22, LHL21, LP23b, MH21, MMM22, MPV23, MW21, MG23, MBKK24, NRV21, NT20, PO21, PBR21, RHM22, RS23, SCHS24, SCK22, UGG^{+}20, VLR23, WLH23, WLW23, WZC22b, XH20b, YCG^{+}23, YZ21a, YLXY23, YLLA23, ZTZZ20, ZCY^{+}20, ZKT20, ZZZL21, Zha21c, ZC22, ZGL20, Zhu21, ZR24b, ZGG23]. Non-Archimedean [CFL21]. non-autonomous [CFdD23, Fal21, LLW22]. non-coincident [CPBG23]. Non-conforming [MDKG20, RS23]. non-constant [SCK22]. Non-convex [GRS20, LL23, PO21]. Non-destructive [BJP24]. non-diffusive [BM22b]. non-divergence [GW23, LWW22a, ZTZZ20]. non-equilibrium [FHH21]. Non-exchangeability [BKMO19, BKMO20]. non-global [CLL21]. non-globally [YZ21a]. non-Hermitian [GHC^{+}20]. Non-homogeneous [HKM24, CTN22, ZR24b]. non-ideal [YCG^{+}23, ZGG23]. non-instantaneous [KCKK24, KMD21, Zha21c]. Non-intrusive [SPC21]. non-isothermal [IKL24]. non-iterative [ZCY^{+}20]. non-linear [DK24, FRT23, JJ24, Let24, LP23b, MPV23, PBR21, RHM22, UGG^{+}20, VLR23]. non-linearly [MBKK24]. non-Lipschitzian [DHF23, ZZZL21]. Non-local [ABA19, SV24, BGG^{+}21, CZHQ23, EO24, Ela23, GRS20, NRV21, ZC22]. non-matching [WLH23]. non-normal [Li20a, SCHS24]. Non-oscillatory [CLU23]. Non-overlapping [PK24]. Non-parabolic [ÇÖÖ23]. non-parametric [YLXY23]. non-penetrating [Kov22]. non-periodically [CJCDLL22]. non-probability [CMdMRFG22]. non-relativistic [Cal23]. non-separable [HVMS22]. non-singular [AAA20]. non-smooth [BB20b, Bre22, GO24, KM22, LHL21, MH21, MMM22, MG23, MDKG20, WZC22b, Zhu21]. Non-stationary [CV24, BDR23, WLW23, XH20b]. non-symmetric [MW21, NT20, ZKT20]. non-technical [FRT23]. non-uniform [JYY21, YLLA23]. non-uniqueness [IKL23]. non-zero-sum [ZGL20]. nonautonomous [ÁGKM21]. nonclassical [PMVB22]. noncompact [Juk20, YTMK19]. noncompactness [KKA20, RDHA20, TSG22]. Nonconforming [QZW23, CC24, GYZ22, HY21b, JZ21, LW21, MVY19, MCL22, SGY22, SZ24, vtW24]. nonconstant [BSNL19]. noncontractive [YTA22]. nonconvex [AF23, ISD23, JHW21, JLY^{+}21, JMLX23, LRT20, LZ24a, LNN20, LWX24, PN21a, WSLW23, WPX23, WY24b, WLZY21, YZM^{+}20, ZSCH21, ZZWX23]. nonequicontinuity [Pie20]. nonexpansive [BM21]. nonholonomic [SFMdD23, dA23]. Nonhomogeneous [JMX24, MMM20, Qiu20]. nonic [BK24]. Nonlinear [HOGO23, WLBG22, ZAH^{+}24, ZSVL20, ASADB23, ALNK24, ASB22, AAKA24, AAHK24, AJ21, ACD22, AZD21, AID24b, AVL22, BJB20, BA20, BVDB20, BK24, BLN21, BA22, BBMK22, BS23, BFVY23, Bog21, Bre23, CCL23, CD20a, CLX21, CAN23, CQZ22, CS23, DS20, DRR22, DR23a, DZ22, DL23b, DH24b, DC23, DYM21, DS21, DSCZ24, FSC24, FdOPS21, FP24, Fro24, FHC22, GGHY24, GeOJD24, Gol24, GGO21, GHTC21, Gu21, GW24, HM21a, HVMS22, Hor20, HL21b, HC20b, HC21, HZZ21, Hu24, HH24c, IAH24, IYA^{+}23, JHG^{+}20, JG21b, JZZH22, JH24, JL24b, Kaz21, KLS21, KSZ21, Kum22, KRV24, LHES21, LBBB23, LKM20, LYM22, LY20, LCB20, LPY20, LW22, LY22a, LWLW23, LDC20, LLX^{+}22b, Liu22, LZWL23, LYL23, LY23a, LY24b, LT24b, LHS^{+}24, LHL21, LM21a, MDVR22, MDG22, MDSRFO24, MM20a, MM22a, MS20a, MK24, MS22, MH23]. nonlinear [MHHV20, Moo22, Naj20, NS24c, NM24, NB23, OWV24, OEH23, PH24, PT21, PSZ22, RDHA20, RO20, RYH24, Ray23, RTV20, RR22c, Ruj23, SKE20, SBKR24, SY24, Shi21a, SSL24, SL24a, SZA24, SKVA22, SLW^{+}22, TK24a, TXS21, TKT^{+}24, THS24, TL21, TZTC20, TBT^{+}20, VAMO24, VP24, WEEA20, WYD20, WZ23b, WTX24, WAH22, WSY20, WZC22b, WSL^{+}23, WLS24, XHG22, XY24, XH20a, XA21, Yam21a, YZ20a, YQZS22, YHW22, YCXW24, YVXX24, YLFT21, YH22, YSFS22, YTD23, YWZ24, YZM23, YJJW23, YZ21d, ZVT^{+}23, ZZwS21, ZL22, ZSL22, ZB22, ZLZL23, ZWZ23b, ZSL23, ZR24a, ZC24, ZL24, ZIZ21, ZLV20, ZZ20b, Zho20, BBGI24]. nonlinearities [Kar20]. Nonlinearity [AK21, BK24, MT20a]. Nonlocal [MY21, BB21b, MSD21, MJK20, NS24a, PJA24, Shi21a, UL23, VAMO24, WMF20, WZD^{+}24, YLZ^{+}22, ZCY^{+}20]. nonlocal-piezoelasticity [MJK20]. nonmonotone [GSWH22, LWG24, YKAA21]. nonnegative [CZ24, Guo24, HW24a, KL23b, LLLD20, Miy22, TWQ^{+}23]. nonnormal [TS20]. Nonoverlapping [YS24b]. Nonparametric [CEKN20, HZ24a, WZNT24, XY22, HTJC24, RR21]. nonparaxial [CD20a]. nonregular [EWCQ23]. nonrelativistic [KSZ21]. nonsensitive [ZSY24]. nonseparable [JMLX23]. nonsingular [HBLZ21, WCW20, WCMW23, ZBS22]. Nonsmooth [WLJ24, FP24, Kar20, KJO23, LZ24a, PN21a, SH24a, WPX23, dOdO23]. Nonstandard [CPP23, HE24, MMJM22]. nonstationary [CD20b, DHH24, Vab24, ZF24]. nonsymmetric [CXS20, DL20c, SZK20, Wen21b]. Nonuniform [DVRG21, CK20, PR24, QWY24, WOV22, WZNT24]. nonuniformly [CVPM22]. nonuniqueness [ZM21a]. nonvariational [LZ20a]. nonzero [SV23]. norm [HS20, JGW^{+}23, JGZ^{+}24, JHZD20, Kop20, LZZ20, LL24f, RlLZZ21, YTA22, Zha21b, ZZLY23, ZW21, Zho22, PN21a]. normal [AH23, AB22, CDS21, ER20a, GDSCO21, HK21, HZS^{+}24, Li20a, LTF^{+}20, MPGE24, NN24, PT20, PRW^{+}22, PW23, SCHS24, WC21]. normal-form [PRW^{+}22]. normal-mixture [CDS21]. normalized [HY21a]. normalizing [WZZ24]. normals [CYY23]. norms [GHG24, LL24f, PNS21, HN22b]. NOSAS [YS24b]. notched [SKS22]. note [ANN21, CFB24b, KK21, NS24b, PLZ21, Sad23, Shi21b, Yam21b, YHC20, ZS20b]. nothing [YSBT20]. notice [AA21b, AA21c, GAA21]. Novel [SNMK21, BD23, BEL22, BA22, CG22, CS23, DDT23, DV21b, EWCQ23, GQR23, HJ23, He23, HSGLE19, HCCW20, JGLH20, KK23, LS21, LJL24, MB24, NNCH21, NSL22, ÖHS23, Pat23, RMG21, Ray23, SM22a, SMKV21, SLZD20, SAN^{+}24, SGD20, SFI23, TDD23, WDS22, YLFT21, YY21].

November

[Ano21-34, Ano22-35, Ano23-35, Ano24-32]. nowhere [CTN22]. nozzles [ZGG23]. NP [SD24]. NP-hard [SD24]. NSE [RVX20]. NSFD [DH20]. Nuclear [PN21a, CTGV20, CKT21, JHZD20]. nudging [DS21]. null [ZZ24]. number [BHJB21, BFRV21, CW20, CNYC23, DWZ20, FFT20, GEU24, Gra20, HMS22, Jas22, Jas24, MDY20, Vie21]. numbering [Sug20]. numbers [Ade22, JKKK20, SL22, SLC22]. NUMDIFF [APW^{+}21, AHK^{+}23]. NUMDIFF-15 [APW^{+}21]. NUMDIFF-16 [AHK^{+}23]. numeric [Mik23]. Numerical [ABFSC24, ASB22, Alj22, AA24b, AGMZ22a, AGMZ22b, AGS^{+}23, AVL22, APW^{+}21, AHK^{+}23, ATO22, BJB20, BHRW22, BK24, BFQ22, BDPR23, Bra23, BSZ21, CDD^{+}24, CS24, CLS24, COT21, CHSZ24, CVPM22, DD23, DS20, DMNO24, DYZ24, DG22b, DAM23, DCAA22, EGG23, Fab22, FPS22, FGP24, GCV20, Gao22, GST20, Gol24, GDG^{+}21, HJ20, HAFA22, HJO20, HW21b, HJM24, HLW22, HF21, HMT23, HH20, HC20b, IYA^{+}23, JT20, JQ22, KZW22, KZ23a, KZ24, KM22, Kes23, KV19, KK20b, KLK21, KPS21, LBBB23, LWD20, LCD21, LGY24, MS20b, MJR20, MDD21, NW23, NV20, Oga20, Ois20, ORS23, PZY^{+}24, PO21, PS22a, PV22, PKS23, RTV20, RGA20, SM20, SDV21, SPMB21, SS24a, SBM21, SL24b, Tan20, UV24, Ust21, WK20a, WC21, WWWY23, XLLA21, XH20b, YZ20a, YQZS22, YCG^{+}23, YPLM22, ZH24a, ZHZ24]. numerical [AHK24, AN24, ASA^{+}21, AA24a, Ask21, AA20a, AA20b, AA21b, AA21c, BSdCCGV23, BT23, BBGI24, BLN21, BC22a, BR22, BGV20, BuR24, Bog21, BFRV21, BDR23, BCVV20, CGP22, CMW24, CLL21, Cap20, CCETZ21, CJL^{+}24a, CMS20b, CV20, CSS24, CPP23, Dad20, DDM20, DS22a, DCP22, DC24, DHZ24, DJ22b, DWWZ22, DHQZ24, EM21a, EEE21, FLMR20, FLZ21, FAVM19, FAVM20, FdOPS21, For24, GM22, GLHL24, GM24, GEG23, GeOJD24, GLZ24, GGT24, GWWY21, GLHJ21, HYY21, HM21b, HRT20, HMRZ23, HO22, HZMD22, HWRM24, HPW22, HAD24, IpAmK20, IuR23, Ixa24, JZW22, JNCY23, KAS23, KEKT21, KKA20, Kaz23, KLM22, KZ23b, KXW23, KLMP21, Kri21, KS21, LNSZ22, Lee21, LP23a, LML21, LLYM21, LJ23a, Li23, LX24, LWXL21, LD21, LG22, LXF21, LS22b, MDVR22, MDG22, MDMS22, Mas21, MRG^{+}21, MCL22]. numerical [MZY23, MRC23, MRC24, NNCH21, Naj20, NZA24, Ngo22a, NB23, OCS24, PAAI24, DFMK20, RHM22, RS24a, RS24b, RH23, RG20b, RG22, RR22c, Rou22, RG23, Rou24, RR23b, SuR23, SM23, SKE20, SDC20, SDLZ23, SK23a, SB24a, SSM23, SHV^{+}22, SH21, Svá23, TFR22, TL20, TTLB22, UÖA23, VAMO24, WK20b, WX21, WWJ22, WLH23, XLL23, YTD23, YSFdRdP20, YOCY22, YLS20, YLS23, ZBS22, ZZW^{+}24b, ZZW^{+}24c, ZHWD20, ZCY^{+}20, ZCYP20, ZLG20, ZKT20, ZLS23, ZEW23, ZS20b, ZDF^{+}20, dC20a, BJMR20]. numerics [MT20a]. NUMTA [BSZ21]. NURBS [JYWZ21, Juh21, YJZ23]. Nusselt [FFT20]. nutrient [PA24a]. Nyström [ASML23, DT24, FO22, KFK^{+}23, KFK^{+}24, MRC24, Naj20, ST22]. Nyström-quasilinearization [Naj20].

O157 [BYWA22]. objective [CL24b]. objects [DFK20, RV20]. oblique [GW23]. observation [AST^{+}23, HLR22, LCZ19, LCGW24, Ruj23, XZ22, YTT23]. observations [AK20, DJ22a, FJMP24, HX22, JD22, New24, TN24, YT19]. observed [DNK23, NPP^{+}21, WTL20]. observer [DSCZ24]. obstacle [ACDV22, KP22, KNM24]. obstacles [ETKF23, NYL23, XLY22]. obtain [EHVMM24]. obtained [Bra21]. obtaining [MGG^{+}24]. ocean [AMS23]. October [Ano21f, Ano21k, Ano22f, Ano23f, Ano23l, Ano24e, Ano24j, Ano22l]. ODE [DFMK20, UV24]. ODEs [Bün20, CT20, DSCF24, MT20a, ON22]. off [LMS21]. Offline [AMZ24a]. Ohta [LLZY22, WZY23, ZCYP20]. oil [PCWY20, ZZ21]. Oldroyd [AAA20, GH22]. Oldroyd-B [AAA20]. Oliver [ÁGKM21]. Olkin [BSNL19]. One [HZ23, LMS21, MCY23, MSS^{+}21, Bac19, BT21, BCMP23, BJP24, BC24, BS20, BRLS23, BDF21, Bra23, CT20, DP21a, DDUF22, DNS21, DW20, FDDM24, KLH21, KHAW21, KV19, LZW20, Mag20, MDS20, MQ20, PJA24, PMVB22, SNMK21, SDW24a, SK23a, SL24b, Xin21, YAE23, uDuIZ22, ZTY^{+}20, ZDY20, ZM21a]. One-dimensional [HZ23, LMS21, BS20, BRLS23, BDF21, DDUF22, DW20, FDDM24, KLH21, LZW20, MQ20, PJA24, SNMK21, SDW24a, uDuIZ22]. one-shot [BCMP23, BJP24, BC24]. One-sided [MSS^{+}21]. one-speed [PMVB22]. One-step [MCY23, KHAW21]. Online [AMZ24b, YWW24, AGS^{+}23, DR23b, HCCZ24, SLW^{+}22, SSV23]. only [PK24]. Open [GL24, ZCJL20, MBG21a, ZAC^{+}23]. Open-loop [GL24, ZCJL20]. open-source [MBG21a, ZAC^{+}23]. operation [PKK^{+}22]. Operational [DR23a, ZE23, AEST21, RB19]. operations [DDT23]. Operator [BitH21, CDD20, ABBB23, AKM^{+}24, BCLS24, BBAO23, BA23, CCL23, CQZ22, DMNO24, DL23b, Dom21, FL24, GT23, HM21b, HC21, KWN20, LL24f, LZG24, LV20, MV22, NM24, PSZ22, RKS^{+}22, SN23, Vab22, XSL22, ZWZ^{+}23a]. operators [Alj22, AN24, Ans21, BCV24, BM21, CJP20, CV20, CC20, DDNZ21, Dum24, EK24b, IY20, IuR23, Kad22a, KRT23, Liu20, MMJM22, NRV21, Rab21, RVS^{+}22, SMM22, WYG23, WDPY24, YTA22]. opinion [AKMO21]. optical [ACT22, BSV20, HLLY24, HKU22, RV20, ZAH^{+}24]. optical/SAR [HKU22]. optics [CLM22b]. Optimal [AGH20, AKMO21, AS23, BEK^{+}23, CZHQ23, CBF24, Dad20, DHH24, DWWZ22, ER20a, ER21, FBJ20, FPGSR21, FC23, GSZ22, GDA22, GH20, Hon22, HS20, LLW21, LR22, LMYZ21, LX24, LL24c, LLGC24a, LZ19, LY23a, MULS23, MX20, MFH24, NL23b, PA24a, RK23, SVK20, WCPW20, WMH22, WW23, WY21a, YYC24, ZZK21, ZXK24, AA23, AK22a, Asa24, BCT22b, BBTW24, Bre22, CCJ20, CLHL20, CLH22, CCHW24, DMP23, DL21, DDRS24, Els21b, EY23, EP20, FCS22, FMAV23, GYZ22, HÇYK24, HJL20, HJS21, HCH^{+}23, HS24b, JGW^{+}23, JGZ^{+}24, KLZ^{+}24, KWYN23, LS22a, LMP20, Li23, LY24a, LLZ24a, LCH23, LW24, LF21, LLL22a, LTZ24, Mam23, MMMM20, MM22b, PPSM24, RSS24, Ros23, RR21, SMDVA22, SS20a, SGY22, SZA24, SLZL23, TY23, VZ21, WZ20a, WWS^{+}21, WRZW21, WZZ21a, WWZ24, WZC22b, YSFdRdP20, YZZ22, YWTL19, YW23, YLH23, ZG22]. optimal [ZXK23]. optimal-order [RR21, WZ20a]. optimality [BLyS20, LMM22, PT24, SH24a, XXH^{+}22, YT19]. Optimally [vdMOB22]. optimise [EGM23]. Optimization [DXW21, FPCV^{+}22, PKK^{+}22, PS24b, Zha20a, AF23, AORS22, BAB20, BTC21, Ben24, CDLR24, CWLM22, CZM23, DAMA22, DL20a, ER20b, For24, GS20a, GXL22, GJGDPR23, HAFA22, HH20, HTTO20, IHK20, JHW21, JLY^{+}21, JMLX23, JNCY23, JKW23, KLM21, KLS21, KJO23, KAM^{+}23, La 22, LMM22, LZD21, LZH21, LWG24, LT21, MRI22, NP23, OPK23, OF22, PS22a, PKPD21, RCVA24, Ros23, RPQ24, SNMK21, SCY23, SA23, SZL20, SW23, TFR22, TLY23, UPPZ24, WSLW23, WPX23, WY24b, WZL^{+}23, WLZY21, WZC22b, WSL^{+}23, YZ24a, YYT24, YWL24, ZSLG21, ZSCH21, ZZ24, ZYZ24, ZHW20, Zhu21]. optimized [CAJ24, CV24, DDT23, GNGGL^{+}19, KK23, SSRK23, Xin21, Yam23]. Option [HL21a, KKJ22, MYC22, NW23, AHB20, AHB21, AV21, Boe20, CLS24, CST20, DLS20, GJNM24, HCCW20, HSYZ24, Kaz23, KKKL22, LL24a, MW20, MVB20, MNH21, RG20a, Rou22, SZ23, SYTD24, WSC23, XW22, XSL22, dFG21]. options [AZ22, ACD22, Hai23, JCY20, KYP21, Lee21, LH21, LLH24, MST22, MN20, RG22, Son24, Wan21b, WZ24b, YL21a, YSBT20, YHSK22, YWTL19]. oracles [WPX23]. orbit [KPS24a]. orbits [ADMV22]. Order [EGG22, MFA21, ADZ20, ASML23, AEST21, AAKA24, ALK20, AA23, AAHK24, AAB24, AN24, ADEEY21, AB23, AMT23, AI22b, BJB20, BT21, BHRW22, BJB22, BK21a, Bay22, BK24, BFQ24, BR22, BA22, BBMK22, BCT22a, BCT22b, Bic23, BYE^{+}23, Bos24, BDS20, BIZ20, BP22b, CCS24, CZY22, CZHQ23, Cap20, CLX21, CLW^{+}22, CZ23a, CAJ24, CT20, CHH20, CGW21, CKA21, CWLM22, CJL^{+}24a, CLS24, CV20, CD22, CC20, CNQRR22, CLL20b, DP21b, DS20, DRR22, DL21, DGL24, DVRG21, DG20, DC20b, DCP22, DL23b, DHH24, DW20, DL20b, DW23, DLZ22, DSCZ24, DCAA22, ESA23, Ela23, ER20b, EGG23, EKÇ20, EPL21, EHVM22, FM21a, FZ21, FHY^{+}21, FHM23, DSCF24, FMNT24, FDDM24, FK23, GM22, GR22, GJN20, GJDL20, GPHHA23, GPST23, GWWY21, HWCJ21, HZC^{+}23, mHLZ20, HLLL23, HR22, HOGO23, HE24, HL21b]. order [Hu24, HZL24, HPWM20, HL22c, HC24, HTIA24, HWAT22, HH24c, IMMS21, JWZ21, JWG21, JZZH22, JQ22, JLM24, JCLC20, JVAD22, Kar24, KCKK24, Kaz23, KKJ20, KCK24, KR21, KRT23, KFK^{+}23, KAA22, KD22, KD23, LLFT20, LL20a, LML21, LLZC21, LLZY22, LLX^{+}22a, LLS22, LW23b, Li23, LZ23a, LWW23a, LJ23b, LWWY24, LWWZ24, LVW24, LL23, LZ22, LHMT20, LMLB22, LY22b, LZ23b, LC23, LYL23, LST^{+}24, LXY24, LC24a, LL24e, LP23b, LXF21, LN23, LXYL24, LM21b, MGG^{+}24, MH21, MGD22, MTZ23, MKS21, MMMM20, MWBM22, MM22b, MBZ23, MZY23, MQZ^{+}24, MA19, MJR20, Moi24, MF24, MMPM22, MRC24, MJK20, MB24, NMST21, NS24a, Ngo22a, ON22, OH23, OGR22, OWV24, ÖHS23, PA24a, PL21, PM20b, Pat23, PWL23, Pou24, PI24, QXZG22, QZW23, RVS^{+}22, RR22a, RAN20, Ray23, RB23a, RGA20, RG20a, RG20b, RR22c, RZ20, RR21, RR23a, SM23]. order [SXX21, SLXF24, ST22, SMRN24, SC23b, SWN23, SLA^{+}24, SH24a, SPC21, Sur21, SSVS20, TK24a, TKM21, TH24a, TXS21, TSL^{+}21, THS24, TCGY24, UGG^{+}20, Vab23, VV23, VAMO24, WEEA20, WZ20a, WCPW20, WLL21, WWJ22, WOV22, WSC23, WWW^{+}24, WWZ24, WKN22, WSY20, WLB^{+}21, WY21a, WZZ24, XWXZ22, XLKL23, XLY22, XYAH23, XH20b, YZW21, bYW23, YCXW24, YVXX24, YLFT21, YTD23, YKY24, YZZ22, YOCY22, YZ24b, YF22, ZVT^{+}23, ZBS22, ZCY^{+}20, ZG20a, ZZ22b, ZWY22, ZLJT22, ZW24, ZFW22, ZMFL20, ZPW^{+}24, dA23]. order-preserving [LZ23a]. ordered [New24, SHB24]. orderings [AB22, SHB24]. orders [AM21, LFL20, WJ23, ZM20]. ordinary [ASML23, AA20a, AA20b, AA21b, AA21c, CKL23, FM22, GGRCGG^{+}23]. oriented [AESW22, DC23, FYL^{+}24, FHB20, LY22a, MVY19, Yüc21]. originated [SMDVA22]. Orlicz [BCV24, DKM23a]. Ornstein [HX22, HS23]. orthoexponential [KS21]. Orthogonal [Dju24, GL20, Ver20, Alh24, BD24, BDF21, CGP22, DKSQ22, DMN24, ESA23, FS20, GANR22, KS22a, LMAV24, LLLD20, LGY24, LMLB22, MkCZ20, Ri22, Ri24, SSB20, SPC21, TWQ^{+}23, WYTL19, Xu20, YC20b]. Orthogonality [ZMZ^{+}24, KJGN20, Yak21]. Orthogonality-based [ZMZ^{+}24]. orthogonalization [DV21a, LNP22, ZYZ24]. orthogonalization-free [ZYZ24]. orthonormal [MDE24]. orthotropic [ZUZ20]. Oscillation [KS22c, Bic21]. oscillations [FJ21]. oscillator [CNQRR22, KZ23a, KZ24, WTP^{+}21]. oscillators [FHY^{+}21, LM21a, ZH24a]. oscillatory [AL20, CYC21, CLU23, Gao22, GJ23, Ixa24, KZW22, KWZX23, KZ23a, KZ24, KEKT21, KAD22b, Maj23a, MI20, MSM24, RG21, WK20a, XLL23, uDuIZ22, ZuIH20, ZFM22, ZFW22, ZH24a]. Oseen [MTZ23, QSW24, SM22b]. outcome [VTB21, VBRT24]. Outer [JWS23, DS22a, KL23b, MSS^{+}21]. outflow [GO24]. outliers [NN24, NN22, ZFH20]. output [LCGW24, MDLY23, XDP24]. outputs [Kan24, PN21b]. over-penalized [QS23]. overcoming [BKMA24]. overdetermined [HM21b]. overlapping [LZ24a, PK24, TCGY24, WCK20]. oversampled [MH22]. overtaking [RLGGAV21].

p [TBA24]. p-median [TBA24]. P1 [LW21]. P1-nonconforming [LW21]. P2P [CZ20]. package [LMAV24]. packaging [GJGDPR23]. packed [XLY22]. packets [ZK20]. pad [TZBK21]. Padé [BZH24, BM24b, CHsL22, JC24, MWHC20]. PageRank [HWH^{+}21, JWSG22, MT20b, TZWG21, WHY^{+}21]. pair [Kum22, LHG^{+}24, LLZ24b, RR23a]. pairs [FCS22, JZ24, LP20, RL23, Sad23, ST22, ZMFL20]. pairwise [HK21]. pandemic [Ngo22a]. panel [XL20, ZZM20]. Pantograph [SSS24, ADEEY21, BR22, GLHJ21]. Papers [APW^{+}21, AHK^{+}23]. Parabolic [Slo21, ZHWD20, AGH20, ALNK24, ASB22, ATHW21, AA22, AA24b, AA24c, Ars22, BDF21, BLRZ23, CJ23, ÇÖÖ23, DMZ23, DCY24, FK23, GPHAPR21, GPHHA23, HW21b, HH24a, Hor20, HTIA24, KSVA22, KCRVA22, LS21, LLS22, LZ19, LW24, LCZL24, MS22, NZA24, QS23, RS23, SMDVA22, SS20a, SDC20, SLXF24, SZ24, SHV^{+}22, THS24, UV24, Vab21, VLR23, WMH22, WLJ24, YVXX24, YLD19, ZGJ21, ZZL20]. parabolic-ODE [UV24]. parabolic/mixed [LS21]. paradigm [SQT24]. paradoxical [AN20]. Parallel [LJ23b, PHDD22, ABM^{+}22, BABR20, CZM23, DLC21, ERB^{+}22, FBJ20, IBW^{+}20, JLW23, KV20, LGG^{+}22, LWA^{+}22, LW24, OKE24, ZC21]. parallel-in-time [JLW23, LW24]. parallel-series [FBJ20]. parallelogram [Yam21a]. Parameter [ALK20, CDLR24, HX22, Liu22, SW20, AHK24, ARK20, Bog21, BPR20a, BPR21, CM24a, CDL24, Cho24b, CSS24, DXM^{+}20, Din23, EK24a, GBTS22, GH21, HSM20, JG21b, JZM^{+}23, Khl21, KLS21, KTK21, KSVA22, LHW22, Liu23, LH20, MDLY23, MS24a, NK24, Pou21, RKA20, RJ22, RMM22, SC21, SCK22, SLP20, SVJG24, Six20, XDX24, YWL24, ZuIKA22, ZZ22a]. parameter-dependent [GH21, RMM22]. parameter-free [JG21b]. parameter-uniform [KSVA22]. parameterization [Cho24a, JYWZ21, LHG^{+}24, YW23, ZZ23]. Parameterized [SPKS^{+}24, AAB^{+}20, HLLL23, KAA22, MDY20, WL22]. parameters [BSNL19, Che21, DJ22a, GIV20, JD22, JKJ^{+}23, JKW23, KDO20, KS24, KCRVA22, PKPD21, XHLD21, YSUY24, YLD19, YS24a, ZKT20, dL23a]. Parametric [MN22, PA24b, BPD20, CNQRR22, HK23, LJ23b, MS20b, RHA20, TTK23, VVB21, YLXY23, ZJ24]. parametrized [YF22]. paraorthogonal [BCBM22]. Parareal [Li24a, DMZ23, MM20b]. paraunitary [Ri22, Ri24]. Pareto [FPLZ21, NN22, Pou21, ZG20b]. Parseval [Hu22]. Parsimonious [NN24]. Part [JKW23, Sur21, XLL23, ABFSC24, CK22, CK24a, YZ21c, YZ21b]. Partial [AGST23, AST^{+}23, BM24b, SS23a, ADZ20, ALK20, ASTK24, AS21b, Ars22, Cap20, CW24, Dar20, DO22, GLCL20, IMT20, JHYZ24, JHZD20, JWG21, KA24, LZ20a, LVW24, LMZ24, MW20, MMJ21, RK22, RB19, RG20a, SM22a, SSL22, SW20, SDW24b, TR22, XFW24, Yam21b, YZ21a, ZE23, ZvS23, ZG20a, ZGJ21, ZB22, ZCZ24, ZMZ^{+}24, PP24]. Partially [HL21c, LL24b, SLLP24, AGST23, AST^{+}23, DNK23, GLLZ21, HSRA24, LZ19, LCGW24, NS24c, WCWZ20, WTL20, WSLW23, YH22, ZZM20]. partially-coupled [LCGW24]. participating [BS23, CMRS21]. Particle [GMW24, GS20a, GXL22, HB23, HBPV21, LJL24, SZL20, TK24b, Wan21a, Zha20a]. particle-laden [HBPV21]. particles [BIZ20, CJCDLL22]. particular [KZ23a, KZ24]. partition [CDLR24, PP22]. partitioned [BJF24, MM21]. partitioning [CJL^{+}24b]. partitions [LWWZ24, LWZ24, LW21]. Pascal [DOP21]. Patankar [ÁGKM21]. patch [GMW24, SSL22]. patches [DOP23, LHL19]. patents [CC21b]. path [AV21]. pathway [AS20a]. pattern [HRT20]. Patterns [HCCSCGAM24, AL20, BBSV21, HMS22, JMX24, Kon22, Pul22]. Patterson [dlCYS22]. Patterson-type [dlCYS22]. payment [Pou21, WRZW21]. payment-type [Pou21]. payoff [Ruj23]. PCA [SW24b]. PDE [AK22a, ÉM21b, FS20, GPHHA23, PP24, RG20b]. PDE-convergence [GPHHA23]. PDEs [BCCW22, CUN24, CDL22, CSS24, FGN^{+}23, Hon22, KCRVA22, LZ20a, MkCZ20, Son24, UGG^{+}20, VLR23, vdMOB22]. Peaceman [DL20a, JMLX23]. Peano [FB24]. Peano-Sard [FB24]. Peer [LS22a, SLW21, KHAW21]. penalized [AFJGJJ22, GLLZ21, HSW21, HG22, HL21c, LZ19, PL21, QS23, SW21b, TWQ^{+}23, YLXY23, YLH23, ZCYP20]. penalty [BK21b, EKÇ20, FHC22, GHTC21, GHL21, HWCJ21, HDW20, HHD23, MCY23, TXZ24, WSC23, WPX23]. pencil [SK24b]. pendulum [BCH20]. penetrable [ETKF23]. penetrating [Kov22]. Penrose [AAMR24]. pension [AZZ23, PZL23, WRZW21, ZLL20a]. perfect [Slo21]. perfectly [DL21, HLYZ24]. perforated [CM20, CHP21, GH21, GH23, RL24, TCVC21]. performability [MR20]. Performance [MO20, OPK23, RCVA24, SA24, CWT23, CKL20, COT21, FPCV^{+}22, HC20a, JB21, LHS^{+}24, ZWH^{+}20]. peridynamics [LST^{+}24]. period [CSZP20, Har21, McA21, MER20, YWTL19]. periodic [ADMV22, BE21, FLZ21, GIV20, JMX24, KMN20, Men24, Ois20, RCN21, WMF20, XZ22, XBC22, ZGZW23]. periodical [DH22]. periodically [CJCDLL22, RL24, SH21]. Periods [KV20, LCZ19, MSBSM23]. permafrost [NSL24, PHV24, VP24]. permanent [KKO23, RCN21]. Perron [Miy22]. perspective [DV21b, FMAV23, KAS23, SZWH21, TL21]. perspectives [JA22]. Perturbation [IM24, SL22, SW24b, XHG22, BV22, DMP23, HZ24b, LSY24, NAS22, TL21, Wu24, Xu20, YS24a, YFW20, ZCT21]. perturbations [AC22, FM22, Ver20]. perturbative [PT20]. perturbed [AVL22, Bha21, Bog21, CLX21, CX24, Che21, CJ23, CSS24, DTM21, DZ22, DLZ22, DCAA22, Ela23, ER20b, GO20, GO24, HWRM24, Kim20, KAA22, KSVA22, KCRVA22, KL23a, LC24a, LYZ24, MH22, NMMH23, PSRVA24, RS24a, RS24b, SS23b, SKVA22, WLM24, ZJY22]. Petrov [Ned22, ZEW23, ZR24b]. Petryshyn [DK24]. PGM [TLY23]. PGM-based [TLY23]. PH [WY21b]. Phase [HLA22, LLZY22, YLK24, AS20a, AJR22, Asa24, CZY22, CDCVV22, CS21, GJDL20, GX24, HET24, HYY21, KS23a, LYM22, LLT^{+}21, LML21, LLYM21, LQS23, MWLY24, NY24b, PHV24, PH21, QH23, SRAG23, VP24, WYTL19, WCLY23, WZY23, XLKL23, YLJW24, YOCY22, YLS20, ZC21, ZCYP20, Zha21b, ZLS23, LMSZ21, WC24a]. Phase-Field [HLA22, LLZY22, YLK24, CZY22, HYY21, LYM22, LLT^{+}21, LLYM21, WCLY23, WZY23, XLKL23, YLJW24, YOCY22, ZCYP20, WC24a]. Phase-II [LMSZ21]. phase-type [GX24]. phenomenon [BCA20, BCGAR22, RBR23]. phi [JdlCD20]. phi-functions [JdlCD20]. photoacoustic [Ben24, DBR24]. photonic [HLLW21, LYL23, LLL^{+}22b]. Physical [HCC24, Alh24, RKS^{+}22, Seg23, XYAH23, ZZ23]. physical-property-preserving [XYAH23]. Physics [DSCF24, BTBR21, CS21, LL24b, NY24a, RDG23, ZG24]. physics-induced [NY24a]. Physics-Informed [DSCF24, ZG24]. physics-preserving [CS21]. Picture [BYE20]. Piecewise [HY21a, BCR22, DDUF22, DFK20, GEU24, HDW20, HHM22b, KR21, SSL24, XHLD21, Yam20, ZSL22, ZXK23, ZSL23, ZXK24]. piecewise-linear [DDUF22, Yam20]. piezoelasticity [MJK20]. piezoelectric [MJK20, TZBK21]. piezomagnetoelectric [ANM24]. PINN [GHJM24]. PINNs [DHJ^{+}24, VLR23]. pipes [FFT20]. Pitaevskii [Bro23, EGG22, SRMD22, WW23]. pivoting [EFJK22]. plan [LLW21, PZL23, ZLL20a]. planar [CFJVG21, CJM20, GW23, JKJ^{+}23, JYWZ21, WSYPD24, ZZ23]. Planck [DHQZ24, Wol24, JZ22, LWW^{+}22b, QQSZ24, YSFS22]. plane [KLMP21, Kuz22, PSWZ21]. planning [YP21]. plans [WRZW21]. plant [CCJ20]. plasmonic [MMM20]. plastic [PA24a]. plasticity [DTMJ24]. plate [FS24a, GeOJD24, JO24b, OKW23, PSZ22]. plates [AAA20, MCD24, NNL^{+}24, RL24, ZUZ20]. platform [CZ20, WQ20]. platykurtic [GMK20]. player [OHK23]. Plebański [YT24]. plus [SW24b, ZYZ22]. PMI [Ri22]. PML [HLW22, HLF19, HLLQ22, Kim22]. PNP [DHQZ24]. POC [Yam23]. Poincaré [ABBB23]. Point [Yam23, AS20a, AET23, AAA^{+}23, BZH24, BP21, BLV22, BM21, BG23, BLyS20, CW21a, Cao20, CHXL20, CTN22, DK24, EK24b, ERB^{+}22, FZ20, FHM23, GGPP21, GH20, HLR22, HWLD23, IHK20, JZ24, LZF20, LLLL20, LHL21, LPP24, MK24, NB23, OF20, OHK23, PH24, PS24a, Rum20, Saâ20, SuR20, SRBK23, SLK24, SCZ20, TSG22, TLD^{+}23, TH24b, WL22, WW23, Wen21a, WV21, WSX21, YTA22, YKY24, YL21c, Zha21b, ZCCK24, ZLL^{+}20b, BM24b]. point-based [PS24a]. point-wise [OHK23, WW23]. points [ABM^{+}22, BGR20, BHJB21, CNYC23, EWT24, JYWZ21, RG21, SBKR24, YL21b, YL20]. Pointwise [FK23, KP22, YL20, Ku24, LLZ24a, LTZ24, ZXK24]. Pointwise-in-time [FK23]. Poisson [Abi20, AK20, DHQZ24, GX24, Gua20, HY21b, JZ22, KSG^{+}22, LWW^{+}22b, LLL22a, MÖ21, OK23, OBdSD24, QQSZ24, VCNR20, WYZZ22, YSFS22, ZTY^{+}20, ZB22]. Poisson-disk [KSG^{+}22]. Polarity [FP21]. polarized [CLM20, CLM21]. poles [Che22]. policies [CMRS21, LMYZ21, Tor21]. policy [EP20, LLL22a, MVB20, MFH24, OKE24, SA24]. Pollution [DV21b, PA24a, SMM22]. poly [WYTL19]. poly-phase [WYTL19]. polyadic [BAD^{+}22]. Polyak [Hun21]. polycircular [HNV23]. polycrystallization [HPW22]. polygon [Juh21, VOJS24]. polygonal [AMS23, LLZ24a, LWZ24, LHMT20, LCZL24, PO21]. polygons [BSV20]. polyharmonic [Seg21]. polyhedral [WYZ24]. polymer [BEK^{+}23, SW21a]. Polynomial [BCDF22, MMPP20, PL24a, AFKM22, AA20a, AA20b, AA21b, AA21c, BL20, Bha23, BCKSV24, CDW20, DG21, DWZ20, FGM21, GMCJGM24, GLCL20, GL20, HY21a, HG22, gHtSjT24, KS22a, KRV24, LZLL20, LD21, LXYL24, MPM23, MLYZ21, MSS20, NZ21, PRW^{+}22, PH20b, PS22b, RK22, RCM20, SSB20, SPC21, SH24b, TG21, VZ21, Vie21, ZXK23]. Polynomial-reproducing [BCDF22]. polynomials [Alh24, AIDAJ23, BD24, BWS20, BWS21, BCBM22, BM24b, CST19, DGL24, DR23a, DIAJ^{+}22, DO22, DMN24, DU19, Dju24, Dum24, ESA23, GSZ22, HAFA22, HZC^{+}23, HVM22, JKKK20, JWSG22, KJGN20, KS23a, KK20a, KS21, LMAV24, LL22, LS22b, MARH^{+}23, MSS20, MT20b, ÖÇ23, PL24b, RSS24, Sad23, SS23b, SZA24, UDI20, Ver20, ZD21, ZH23, ZBH24, ZXK24]. polytopal [Gua20, LWWZ24, LW21, LZ23b, WWH21b, YZ20b, YZ21c, YZ21b, YZZ22]. polytopes [McA21]. population [BTBR21, ÇK24b, GeOJD24, GK21b, GS24, SBM21, VAMO24, YSUY24]. populations [BFRV21, JD22, MTP21]. pore [GIV20, PHV24]. pore- [PHV24]. Poroelastic [Kov22, AGST23, ASTK24, RCN21]. poroelasticity [DHZ24, Lee23, SLLP24, TVSC20, TVAS23, WTL21, WCFH21, WWTL24, ZR24a]. poromechanics [ANR20, Ter20]. porosity [BJF24, BMS21, CGO24, WSHW22]. Porous [JSRdS23, ACS22, AJR22, CS21, CRRB21, DRB22, FPS22, GM24, GIV20, HET24, IKL23, KLM21, LZWC21, LQS23, LP23b, MHM23, Moo22, NY20, RH23, RB23b, SEAS22, SVC20, VMP23, WCFH21, YN21, ZZWS20]. porous-elastic [FPS22]. Porous-elasticity [JSRdS23]. Portfolio [MYLL24, SL23, WSZ20, AD21, CSZP20, CL24b, DXW21, hKhWjH21, LLW21, LWZF21, MER20, RCVA24, WZLT23, YZ24a, ZCJL20]. Portfolio-consumption [MYLL24]. portfolios [ZYZ23]. posed [BPR20a, BPR21, BDR23, DH24b, HLY22, Hun22, Jia20, LHL21, LZG24, MV22, NKBJ20, ORS23, Six20, TSL^{+}21, Wan20a, ZW20, ZZ22a]. posedness [Hun21, HK23, JDBT22, OCS24, TZ20]. posets [OHCC22]. position [IM24]. positive [AAP24, AA21a, BAPPF23, BCBM22, Cao20, CLM22a, FS24b, HWD^{+}24, KH21, LGC23, Liu20, LZLL20, LNP22, LM21b, Miy22, MDE24, Sun20, Tan20, TFK23, WL22, vdBS21]. positive-semidefinite [Cao20]. Positivity [DH20, MWW21, DHQZ24, HAFA22, HWD^{+}24, LXY24, MDSRFO24, MPR24, YVXX24, YF22, ZCF^{+}22]. positivity-preserving [DHQZ24, LXY24, YF22, ZCF^{+}22]. positone [LMZ22]. Possibilistic [LS24, SLC22, Tak20]. possibility [Dad20]. possible [GZ24, LMSZ21]. possibly [DMP23]. post [Dad20, PH22]. post-buckling [PH22]. posteriori [ANR20, Bac19, BPR20a, BPR21, CHY20, DCY24, DN20b, FK23, GGVRB22, GS21b, HOOT20, JHZY20, JKW23, KP22, LJCL20, LY22a, LCH23, LWWN22, MVY19, MS22, MR23, MR25, Ngo24, OBdSD24, RS23, TXZ24, WMH22, YSFS22, Yüc21, CE24, DLZ22, CC21a]. postprocessed [DN20a]. postprocessing [YZM23]. potent [CLST20]. Potential [DMST21, DWWZ22, KH24, Mag20, RL24, SH23, WL24b, Xin21, YZ24d, ZM21a]. potentials [FFS24, FY22, RK23]. Poverty [WQ20, MRI22]. Powell [BEIR23, GS21d]. Power [BCMP23, CMP21, GAA19, GAA21, LTL^{+}24, PA21, YCD^{+}20, CGSVBB^{+}24, DRD^{+}23, SMKV21, SN21b, Vab22, WLZY21, HWH^{+}21, WHY^{+}21]. Power-Law [LTL^{+}24, CGSVBB^{+}24]. powered [KKKL22]. powerful [SCHS24]. powers [CV20, Miy21]. Prabhakar [GPST23]. practical [BCR22]. practice [RSS24]. Prager [CHZ21]. Pragmatic [TFK23]. preassigned [BM24b]. precedence [CBF24]. preCICE [ZAC^{+}23]. precipitation [GM24]. precise [ZCY^{+}20]. precision [LGG^{+}22, MO20, RDH21]. Preconditioned [LHL19, WYD20, ABHK20, BDR23, JS21, NT20, SvG22, IBW^{+}20]. preconditioner [Cao20, Gra20, LZF20, LW24, TR22]. Preconditioners [CGMM20, CXS20, FZ20, Hon22, LWT24, NT20, OF22, WL22, WZ24a, ZSLG21, ZC20a, ZG22]. Preconditioning [VTBM21, Abi20, ACEIJH22, DWWZ20, GZ23, HN22a, Kar20, SA23, SFI23, TH24a, WZC^{+}22a, ZYZ22, ZZwS21]. predator [Gök23, HE24, JMX24, KMUR24, MDVR22, NEYZ22, SS23c, YBM23, YAE23]. predator-prey [Gök23, JMX24, KMUR24, MDVR22, YBM23, YAE23]. predator-taxis [JMX24]. predication [GXL22]. predict [AVLMRMGG23, DKM23a, VTB21]. predicting [Ngo22a]. Prediction [New24, SSM23, AGNGG^{+}23, BRZ22, CLS20, PC24, RMG21]. Predictive [BK22]. predictor [ABP24, HVR22a, ZJGZ20]. predictor-corrector [ABP24, HVR22a, ZJGZ20]. predictors [HA20a]. preferences [WSZ20]. Preisach [BGV20]. Preliminary [LYYZ24, MBG21a]. preloaded [Grz21]. premium [LLW21, LPY20]. premium-reserve [LPY20]. premiums [GL20]. prepayment [YHSK22]. prescribed [AM23, KPS24b, LP20]. presence [AAHK24, KLK21, LMYZ21, MMM23, RG21, SLW^{+}22]. Preservation [BCLS24]. Preserving [SYS24, AZD21, AZL21, Bha21, BIJ21, BIZ20, CS21, DH20, DL23b, DHQZ24, DWZ20, DKM^{+}20, ELS21a, FCS22, HCC24, HL21b, HZZ24, Hu24, HWD^{+}24, JWG21, KS22c, LZ23a, LJL24, LXY24, LF24, MDMS22, MWW21, NS24a, RK23, SY24, TCGY24, Wan21a, WSY22, WLDZ24, WLS24, XYAH23, YQZS21, YVXX24, YKY24, YF22, Zag23, ZCF^{+}22, ZSVL20, ZZZ^{+}22, vdBS21]. Pressure [HLMY24, LR24a, AJR22, BMS21, FHH21, LWA^{+}22, LR24b, Par20, PH21, YY24a, ZZW^{+}24b, ZZW^{+}24c]. pressure-correction [LWA^{+}22, YY24a]. Pressure-robust [HLMY24, LR24a, LR24b, ZZW^{+}24b, ZZW^{+}24c]. pressures [IKL23]. prevention [ADKH20]. preventive [LPLP21]. prey [Gök23, HE24, JMX24, KMUR24, MDVR22, NEYZ22, SS23c, YBM23, YAE23]. prey-predator [SS23c]. price [AAk20, CLS20, HCCW20, Rou22, SLK24, Son24]. prices [DKL20, KKKL22]. Pricing [AZ22, BSdCCGV23, BD22, CKZ21, Hai23, JK21a, KYP21, MST22, MCW24, NW23, Wan24, WZ24b, XCY24, YLWW21, AHB21, AF20, AV21, CNS21, CLS24, CST20, DLS20, GJNM24, GKV21, HL21a, HSYZ24, Kaz23, KKJ22, Lee21, LL24a, LLS21, LH21, LLH24, MW20, MYC22, MNH21, MX20, ÖS20, RG22, Ruj23, SZ23, SYTD24, WLL21, WSC23, XSL22, YL21a, YHSK22, dFG21]. primal [CWW23b, CWW23a, GBM23, LWW22a, LWW23a, OBdSD24, PK24, SCY23, Wan20a, WZ21, WWTL24, WZL^{+}23]. primal-dual [CWW23b, LWW22a, SCY23, Wan20a, WZ21, WZL^{+}23]. Prince [NNJ24]. principal [EFJK22, KZW22, MIdMR23]. principle [HZZ24, NS24a, SY24, TCGY24, WSY22, WLS24, YF22, ZWJ22]. principles [KGV23, ZZ24]. prior [FPGSR21, KPS24a, MWLY24]. priori [LJCL20, LLL24a, LCH23, LY23a, MR23, MR25, SV23, CE24]. priority [DXWC22]. priors [ACFM21]. prismatic [NY24b]. Probabilistic [DRB22, AVRLRMGG23, Psa24, Ruj23]. probabilities [CX24, PLZ21]. probability [CMdMRFG22, HQT20, JCLC20, KDLY22, PA21, YHSK22]. problem [ACK23, ASX21, AF20, ACPV24, AVL22, ABA20, AZ20, Bal24, BBAO23, BZH24, BFQ21a, BFQ21b, BFQ22, BFLQ24, BFQ24, BB21b, BBBK24, BP22a, BMS21, BHB21, CW21a, CW20, CC24, CDTFGV22, CE24, CGO24, CHXL20, CLX21, CFS21, CR23, CJM20, CMFMSQ22, CDD20, CCHW24, Che21, CO19, CST20, DG22a, DMTW22, DR23b, EK24b, FZQ22, FAVM19, FAVM20, FRT23, GCV20, GM23, GHG24, GS21c, HÇYK24, HJO20, HLZ20, HO22, HS24b, HZZ21, HSC21, gHtSjT24, HZ24b, HTIA24, HWAT22, IHK20, IKL24, JGW^{+}23, JLZ23, JLM24, JNCY23, KVHC23, KLH21, KP22, KNM24, KP20, KK20b, KAA22, LS21, LSE24, LRV23, LMV24, LL24b, LZZ20, Li20a, LHL^{+}22, Li23, LW23a, LLK23, LLL24a, LX24, LLGC24a, LLZ24a, LW24, LLLL20, LZ20b, LTZ24, LC24a, LYZ24, LDM23, LRQV24, LR24a, LR24b, MWHC20, MTZ23, Ma24a, MYLL24, MÖ21]. problem [MRI22, MWBM22, MM22b, MMF24, MR23, MR25, NNCH21, NT23, NMB22, Ned22, NSL24, NS24c, OBdSD24, OCS24, PS22a, PK24, PKMM24, PW23, PS24b, RK22, Ret24, RR22d, Saâ20, SBC20, SL22, SK24b, SXX21, SSS24, Slo21, SVWC22, SSV23, SP22, SHV^{+}22, SLLP24, SCZ20, TBA24, Ter20, TL21, TBT^{+}20, TZ20, TVSC20, TCVC21, TVAS23, Vab23, VSSV20, WLHJ20, WC21, WZZ21a, WYZZ22, WMZZ23, WHZ24, WL24b, WL24a, Wen21a, XH20a, XBC21, XBC22, YLD19, YPL20, YCH24, YT19, YC21, ZZW^{+}24a, ZSH20, ZCJL20, ZW20, ZZK21, ZMYL23, ZZNL24, ZHMD22, ZWZ23c, ZLC24, ZS20b, ZR24b]. Problems [MJR20, AGH20, ANR20, Ais20, AF23, AVC22, AGST23, AST^{+}23, AORS22, AK22a, AA24c, AM22b, ABHK20, BJ22, BT23, BGH22, BSMM20, BP21, BDF21, Bic23, BB20b, Bre22, BLyS20, BPR20a, BPR21, BDR23, CGS20, CGMM20, CTN22, CLHL20, CLH22, CKY23, CM24b, COT21, CPBG23, CM20, CA23, CPRV22, DMZ23, DMP23, DCY24, DRR22, DAMA22, DDF24, DC20b, DLC21, DCP22, DHF23, DH24a, DC24, Din23, DH24b, DC23, DT24, DV21a, DLZ22, EOT22, EOR21, EFTW21, ETKF23, Fal21, FZ20, Fan22c, FHM23, FMFM20, FC23, FHC22, GLYS21, GW23, GGVRB22, GDA22, GHJM24, GWHZ23, GLZ24, GPHAPR21, GPHHA23, GS21b, GO20, GO24, GJ21, GHTC21, Gu21, GYZ22, GJW23, GH20, GLLZ21, GSWH22, HWCJ21, HMRZ23, HDW20, mHLZ20, HHM22b, HCH^{+}23, HV23, HLLL23, He23, HVR22a, HVR22b, Hes20, HR22, HSM20, HW20]. problems [Hor20, HOOT20, HL24b, Hri22, HWLD23, HZL24, HL21c, HLY22, HC23, HTTO20, HTZ21, Hun22, HWAT22, HH24c, IMT23, IuR23, IKL23, JDBT22, Jia20, JHW21, JGZ^{+}24, JLY^{+}21, JMLX23, JHYZ24, JLW23, JW23, JYYL23, JH24, JWS23, JKW23, JM21, JZ21, Kan20, Kar24, Kar20, KWI24, KS23a, KMZ21, KKJ20, KPW21, KFK^{+}23, KZM21, KJO23, KDLY22, KSVA22, KL23a, KD23, KL23b, La 22, LR22, LMZ22, LCB20, LZF20, LZD21, LCD21, LDY22, LY22a, LWL22, LWW23a, LWWY24, LWG24, LWZ24, LWT24, LZ19, LCH23, LF21, LWWN22, LQS23, LXY24, LCZL24, LHL21, LH23, LT21, LLL^{+}22b, MF21, MHW23, Mag20, MVB20, MN21, MY21, Mam23, MS22, MMMM20, MMM22, MW21, Mik23, MG23, NL23a, Ngo24, NMMH23, NP23, ORS23, OHK23, OMRdQ24, PT21, PC24, PWL23, PK23, PN21a, PSRVA24, Pou24, QS23, QXZG22, QXZF23, RR22a, RYH24, RS23]. problems [RFGGM20, RTT24, RlLZZ21, RSRW24, RPQ24, SN23, SKE20, SMDVA22, SS20a, SGY22, SCY23, SK23a, SL20, Six20, SZA24, Sou22, SVC20, SH24a, SGD20, SD24, Svá23, SCZ20, TBA22, TL20, TL22, Tan20, TLY23, TXZ24, TSL^{+}21, TGLa21, TLD^{+}23, TY23, TGS21, TH24b, UPPZ24, Vab22, Vab24, VS23, Wan20a, WCFH21, WGC22, WDS22, WL22, WT22, WSLJ23, WSC23, WLCH23, WSLW23, WWZ23, WZ23a, WY24b, WWTL24, WYC24, WLM24, WWZ24, WKC20, WZL^{+}23, WDY20, WZC22b, Wu22, WSL^{+}23, WZZ24, XHG22, XLZ20b, XXH^{+}22, XWL24, YKAA21, YHW22, YL22, YZ24a, YMC22, YDC20, YCH^{+}21, YW23, YYZB24, YWL24, Yüc21, ZSLG21, ZSMG20, ZG20a, Zha21a, ZSCH21, ZG22, ZDL22, ZDW24, ZZ24, ZC20b, ZIZ21, ZJY22, ZXZ24, ZLV20, Zho20, ZW22, ZZWX23, ZZL20, ZDF^{+}20, dSFSS24, YL22]. procedure [EHVMM24, IMT23, JYYL23, uDuIZ22]. procedures [RKA20]. process [AKN24, APR24, AS20b, BKMA24, CX24, CES23, EMR21, EHVM22, GJGDPR23, GX24, LMSZ21, LLL22a, MMV20, NKBJ20, OR21, ÖS20, PA21, PA24b, SCHS24, WWC23, Yam23, YPLM22, YWW24, ZRQ22, ZW23, ZJ24, LTL^{+}24]. Processes [GMSR18, GMSR21, AHB20, AHB21, CCCMH24, CCGT22, CS23, DMP23, GJNR21, GHF24, HX22, HS23, JJ24, MLS22, NRV21, PLZ21, SPS21, Wan20b, ZX23b]. processing [Kan20, Kog21, KNOR20, MGV23, RYH24, RU22, SY23]. produced [Mia20]. Product [ZB22, ALKH21, BCV24, CLM22a, Dax20, DEH^{+}24, Dum24, EJHR24, ERB^{+}22, GJ23, Kra22, LCD21, MLY22, RU22, WLDZ24, YLXY23]. Product-of-Curvatures [ZB22]. production [ÁGKM21, LPLP21, WZ23c]. products [Alh24, GS20a, KK20a, LC24b, NP23, XZP22, Yam21b]. Profile [MÖ21]. profiles [BEK^{+}23, Hri22]. profit [AAk20]. programmable [TDD23]. programming [DH24a, DV21a, GM23, HLLZ24, HL24b, LZZ20, MM22a, ME20, OF20, PT24, PS24b, SZZ21, SWW23, SDW24a, TGLa21, Yam20, ZL22, ZZWX23]. programming-type [OF20]. programs [LXQ21, OF20, ZSHD24]. progression [SŁK^{+}22]. Progressive [PM20a, AK20, LWLD23, LHL19, QG20, Saj23, WTL20, YWW24, Zhu23, Zhu24]. progressively [ZG20b, Zhu20b]. Projected [Bha21, WX20, GGO21, La 22, LWG24, dOdO23]. projecting [HL22b]. Projection [LV23, YN21, AZL21, BBTW24, EKÇ20, GEG23, GVT21, IAH24, KWYN23, KLMP21, LMM22, LSC21, LWLW23, LLL24a, LLX^{+}22b, PC24, PWL23, Rab21, RYH24, SZL22, SDL22, SKZ23, SD24, TDAT23, WYTL19, WAH22, WSL^{+}23, Xu20, ZLZL23, ZYL20, dSFSS24]. Projection-based [LV23, YN21, BBTW24, LWLW23, LLX^{+}22b]. projections [BCKSV24, MGL20]. projective [AHPS22, BLV20, GÇSA23]. projector [HHM22a]. projects [JLL21]. prolate [Tia22]. proliferation [GNGGL^{+}19]. Prony [CHsL22]. proofs [CW21a]. propagate [FM22]. propagation [CCJ20, DH20, HH23, HLF19, JA22, JLL21, Kap21, TGJY22, CQBL20]. propellant [FDDM24]. Propensity [CMdMRFG22]. Proper [MkCZ20, FS20, GANR22, SPC21, YC20b]. Properties [OVGI20, RS22a, WJ23, BCLS24, Bha21, BC22b, CZH21, CL24c, DdO22, DNS21, FAUU20, FU19, HL22a, Hor20, HW24b, IYA^{+}23, KUA20, MS24b, TT20, TTLB22, YKY24, ZFwC20]. property [CFB24b, FHM23, Kri21, LC20, Luc23, XYAH23]. proportional [AID24b, YL21a, YQZS22]. proposal [AVRLRMGG23, RLGGAV21]. prostate [MDD21]. Protection [GDG^{+}21]. protein [HCCSCGAM24, TVB21, VTB21]. proteins [VTB21]. Provably [LJ23a]. prove [ZSH20]. Proximal [KCL20, DL20a, JHW21, JW23, Kan20, LL24f, QXZF23, WSLW23, WPX23, YL22, ZSCH21, ZCCK24]. proximity [HLR22, IHK20, VBRT24, WSX21]. PRP [CKY23, WAH22]. PRP-type [CKY23, WAH22]. PRSM [DH24a]. pseudo [AN24, AA22, AA24b, Ant20, CCETZ21, DS22a, DZLC20, EOR21, HTIA24, JH24, LLX^{+}22b, SDL22, SQT24, TV22, TDL^{+}22, WLB^{+}21, XBC21, XBC22]. pseudo-boundaries [DZLC20]. pseudo-differential [AN24]. pseudo-inverse [SQT24]. pseudo-inverses [DS22a]. pseudo-Jacobi [XBC21, XBC22]. pseudo-monotone [JH24, LLX^{+}22b, SDL22, TV22, TDL^{+}22]. pseudo-parabolic [AA22, AA24b, HTIA24]. pseudo-spectral [EOR21]. pseudo-state [WLB^{+}21]. pseudo-symmetric [CCETZ21]. pseudo-symplectic [Ant20]. pseudodistances [FJMP24]. pseudoinverse [SHZ23]. pseudomonotone [TDAT23, TH24b]. pseudorandom [GGPP21, HMS22, LLGZ20]. Pseudospectra [HLW20, KCS21, YHC20]. Pseudospectral [MMPM22, SDV21, dFG21]. pseudospectrum [KCS21]. PSO [DRD^{+}23]. PU [CST20]. pulmonary [ACFM21]. pulse [KM20]. pulsed [ZHWD20]. pursuit [YSFdRdP20]. pursuit-evasion [YSFdRdP20]. put [Lee21, MVB20, MNH21]. PyAlbany [LPH23]. Pythagorean [BL21, FPS21, SS23a]. Pythagorean-hodograph [FPS21]. Python [LPH23, SL21].

QCD [WGM21]. QCQP [JLY^{+}21].

QCQP-based [JLY^{+}21]. QMC [DEH^{+}24]. QNM [SA24]. QTT [BD20]. quad [HZ24b, WWZ23, Tak24]. quad-curl [HZ24b, WWZ23]. Quad-SAV [Tak24]. QUadratic [SYS24, Saâ20, BJB20, BB20b, CF20, CGL22, DV21a, EHVMM24, FP21, HF24, HB22, JHYZ24, KP22, KWYN23, LZZ20, LLL24b, LWX24, MM22a, Nud24, OF20, PS24b, PN21b, Sad23, TLY23, YZLL22, ZSHD24, ZW22, ZZWX23]. quadratically [Ais20]. quadratization [BCCW22, CG22, WC24a, ZZG20].

Quadrature [BBC22, ADDG22, APR24, BPDC20, BEJR22, BD20, BCBM22, CYC21, CDS21, DOR21, DO22, DDRS21, DDRS24, ESS21, EMR21, FMRS24, HS24a, HJL20, HJS21, JJR21, JPS24, LMAV24, LZZ24, LL24e, Ma24b, MMJM22, MSM24, OKW23, OH23, PPSM24, RG21, RS22a, SuR23, TZBK21, Tom24, Ver20, WKC20, Xia22, dlCYS22, vdBS21]. quadratures [FL24]. quadrics [GVT21]. quadrilateral [HW20, MCL22, WTL21]. quadrilaterals [ZMFL20]. qualifications [LMM22]. Qualitative [TTLB22, UÖA23, Hor20]. quality [JYWZ21]. quantification [KS22c, SH24b]. quantifiers [CDMM22]. Quantile [CN24, CYY23, MRI22, SW21b]. quantitative [DBR24]. quantities [Seg23]. quantization [LMP20]. Quantum [ZWX24, dS24, BB20b, DW23, SY23, WTX24, XWL24, ZM21a, FPCV^{+}22]. Quantum-inspired [ZWX24]. quartic [AH23, Bic21, BC22b]. Quasi [GQR23, KJGN20, MSW23, OHCC22, UPPZ24, XW22, BEL22, BEL23a, BEIR23, Bos24, BCBM22, CKL20, Cho24a, DHZ24, FFRS23, Har21, HV23, Hun21, HTZ21, Hun22, HH24c, JLW23, LNSZ22, LYM22, MYZ23, NZA24, RKA20, SXX21, TFM24, TKVA21, XFW24, XHW24, YKAA21, Yan22, ZBH24, ZR24a, ZF24, ZZ20b, Zho20]. quasi-boundary [JLW23, ZF24]. Quasi-closed [OHCC22]. quasi-complementarity [HV23]. quasi-conformal [Cho24a]. quasi-equilibrium [Hun22]. quasi-gas-dynamic [CKL20]. quasi-generalized [Yan22]. Quasi-Geostrophic [GQR23]. quasi-hemivariational [Hun21]. quasi-interpolants [Bos24, LNSZ22]. quasi-interpolating [BEL23a]. quasi-interpolation [BEL22, BEIR23, TFM24]. quasi-linear [RKA20]. Quasi-Monte [XW22, Har21, XFW24, XHW24]. Quasi-Newton [MSW23, UPPZ24, HH24c, YKAA21, ZZ20b, Zho20]. Quasi-orthogonality [KJGN20]. quasi-paraorthogonal [BCBM22]. quasi-reversibility [SXX21, TKVA21]. quasi-stability [ZBH24]. quasi-static [DHZ24, FFRS23, LYM22, ZR24a]. quasi-variational [HTZ21]. quasi-variational-hemivariational [MYZ23]. quasicrystal [NNL^{+}24]. quasilinear [ACC23, GSWH22, LZ20a]. quasilinearization [Naj20, PT21]. quasiseparable [MDY20]. quasistatic [HM23]. quaternion [JLZ23, WJVG24, WLDZ24, YHQW21]. Quaternionic [RDH21, AKP23, WJVG24]. quaternions [RDH21]. queue [MULS23, SA24, SACQ21]. queueing [AGGAV^{+}21, AGGAVGG23, BS24a, CJL^{+}22, SACQ21]. queues [LLZ24b]. quick [CGSVBB^{+}24, XSL22]. Quintic [MA21a, BK24, CVPM22]. quotient [FH24, Mia20].

R [TV22]. R-linear [TV22]. Racah [DKSQ22]. Rachford [DL20a, JMLX23]. Radau [APR21, JJR21, RS24c]. Radau- [RS24c]. radial [AID24a, CDL24, CDLR24, FGPP23, GJ21, KH21, MBG21b, MMV20, NNL^{+}24, PL24b, Seg23, Son24, WKC20, YSBT20, ZH24c]. radiating [BGH22]. radiation [Kim20]. radiative [BS23]. radical [WX21]. radioactive [Kap21]. radius [KZ23b]. radix [PL20]. radix-2 [PL20]. Radon [MA21b, Ust23]. rain [ZLZ^{+}24]. rainbow [Boe20]. Random [AK22a, BAPPF23, JL21, BG21, BS24a, BCVV20, CLW^{+}22, CW22, CMJN22, DMP23, DG20, DYZ24, GM23, GZ24, HA21b, HZS^{+}24, KLK21, LLY21, LN22, MkCZ20, New24, NMF^{+}22, PBR21, PN21b, PNS21, SHB24, SQT24, TBA22, TY23, Tor21, TZTC20, ZTY^{+}24]. randomization [LRP24]. Randomized [CWY21, JZY22, LM24, MP21, PSS24, ZW20, CNQRR22, LCZ19, MW22, NK24, SDW24b, WCMW23, WLC23, WX20, XFW24, YC20b]. range [CGSVBB^{+}24, HTJC24, JLLL23, Kri21, LML21, SW20, SHZ23, TG21]. range-restricted [JLLL23]. rank [AA24c, ACT22, CGMM20, CMFMSQ22, CWY21, CZM23, Dax20, Dax24, DKM^{+}23b, DYM21, GHG24, HLR22, JNCY23, KF24, KCL20, LM24, MHW23, MDE24, MJ24, PN21a, QXZF23, SWN23, WYTL19, WLCH23, WDPY24, XWXX22, YZM^{+}20, YLZ^{+}22, ZWH^{+}20, ZDY20]. rank- [JNCY23]. rank-one [ZDY20]. rank-structured [LM24]. ranked [BK22, New24, RZA^{+}24]. ranked-set [RZA^{+}24]. ranking [dHSR22]. ranks [BD20]. Raphson [Pho22]. rapid [Ngo24]. Rapidly [LC24b, CDD20, ONT21]. ratchet [Wan24]. ratcheting [JO24a]. Rate [FHB20, CMRS21, DWWZ22, EM21a, FPR^{+}24, GGHY24, Gök23, HN22b, LY20, NS24b, OEH23, SB24a, SL24a, Sou22, WCPW20, WZLT23, YDC20, YCD^{+}20, ZLL20a]. rates [CJP20, CHS20, JK21b, gK21a, gKZ21, LGC23, MW20, MG22, WFZ22, ZZ21, ZXK23, Zho22]. ratio [BK22, CMJN22, GHG24, HZS^{+}24, WWWY23, YT20, ZCJL20, Zho24]. Rational [McA21, ABP24, AQ20, AHPS22, BE21, CS24, CCM22, FKV21, GÇSA23, IpAmK20, JK21b, LHH21, Ma24b, MDLY23, SS23a]. rationality [HTOC20, XZP22]. rationals [HAFA22]. RATTLie [HA21a]. Raviart [Nud24, PKE20, ZOK21]. Rayleigh [ABHK20, FH24, IMT23, Mia20, OVGI20, PM20a, ZW21]. RBF [CKA21, CST20, CPS20, HSW21, MDD21, Sha23]. RBF-based [HSW21, Sha23]. RBF-FD [MDD21]. RBFs [AID24b, HAD24]. RC [AAKA24]. re [For24]. reach [YLH23]. reacting [FDDM24]. reaction [AAM20, AVL22, BB21a, BBTW24, BGG^{+}21, BMV23, CJ23, CSS24, CPP23, FJ24, FP24, GJNR21, GM24, GO20, HL24a, JHG^{+}20, JL24b, Kar20, LY22a, LG23, LC24a, MD20, MDG22, MBKK24, PAAI24, PM20b, PLS22, Pul22, RS24a, RlLZZ21, RG23, SMDVA22, SKVA23, Sun20, TBT^{+}20, VSP23, VAMO24, WMF20, WZZ21a, ZVT^{+}23, ZWY22, ZEW23, ZXC22]. reactive [GRS20, HET24, PFG24]. Real [LL24a, NZ21, Sur21, BYE^{+}23, CZ23a, DHF23, DRD^{+}23, HLLZ24, HOOT20, HL24b, LWG24, LM21b, Miy21, NT20, ÖHS23, QXZF23, Rum20, SW24a, SKZ23, TDD23, Uça23, VV23, WX21, ZCCK24]. real-symmetric [HOOT20]. Real-time [Sur21, DRD^{+}23]. real-world [HLLZ24, QXZF23]. realistic [CSZP20]. realizable [AFKM22]. rebates [LH21]. recession [MSBSM23]. reciprocity [YCH^{+}21]. recirculation [FAVM19, FAVM20, FMAV23]. recognition [MGV23, RFB22, ZJ24]. recombining [Wan20b]. recommendation [Tia24]. reconstruct [ALNK24]. reconstructing [HSB24, ZF24]. Reconstruction [AAB24, CL24a, CLL20a, DDF23, DBR24, DL20b, FGL23, GMW24, HHM22a, HJS21, HL22b, NY24a, PT20, RV20, Sur21, WV21, XTL23, ZM21a]. record [JB21, MKS21, WTWZ20]. records [HBMA22, JS22b, New24]. Recovering [MHHV20, CZM23, HLR22, HTIA24, MHW23, QX24, ZZLY23]. Recovery [CHY20, MH21, MA21b, ZWY22, ACT22, Ben24, CK20, DWZ20, Fan22b, GHG24, JHYZ24, JHZD20, JW23, KPS24a, KAM^{+}23, LW23b, LWWN22, LMZ24, MQZ^{+}24, SCY23, SZZM23, WZH^{+}21, WSL^{+}23, ZWH^{+}20, Zha24, Zho24]. recovery-based [LWWN22]. rectangular [APS22, BGR20, GHJM24, KS22b, RL23, TG21, WCWZ20, ZUZ20, ZWJ22]. recurrence [AM22b, SS23b]. recurrent [HPWM20, PK23]. Recursive [XZ22, ZBS22, KH24, LCGW24, MSS20, PL20, SSB20]. redefined [RG22]. reduce [AHPS22, CMFMSQ22]. Reduced [CM24a, WZZ24, ACS22, CAJ24, CWLM22, DEH^{+}24, SPC21, YL22]. Reduced-order [WZZ24, CWLM22, SPC21]. reducible [Ma24b]. reducing [FGM21, FJ21, HL22a]. reduction [CLPZ21, DP21b, GJW23, HLLL23, HL24b, JWS23, LJ23b, LH23, Men24, MB24, PWL23, PL20, SV23, SW20, SB24b, TCVC21, VMP23, WZLX20, ZW20]. reference [CL24b]. referenced [JNCY23]. refined [WZ24a, Wu22]. refinement [FS24a, HLR22, Kog21, KS22b, OA20, OBB23, RSS24, RS23, SMDVA22]. reflectance [DKM23a]. reflected [ISD23, WWC23]. reflecting [WWC23]. reformulating [LH20]. refractive [HSB24]. refuge [SS23c, YBM23]. regime [AHB20, AHB21, AZ22, AZZ23, Cal23, DLS20, GKV21, HSYZ24, KAD22b, KSZ21, LL24a, MW20, Mit20, MCW24, PZL23, WSZ20, ZCZ23]. regime-switching [AHB20, AHB21, AZ22, AZZ23, DLS20, GKV21, HSYZ24, LL24a, MW20, WSZ20]. region [BA20, BKMO21, BM22a, BS24b, TG21, YHW22, YYT24]. regional [LGMC20, LLG^{+}22]. regions [AAB24, GEU24, LZ22, LSZ23]. registration [HW21a, HKU22]. regression [AS20a, AS20b, AK22b, BTC21, BYE20, CEKN20, CYY23, DDUF22, DMN24, GK21b, GBA21, HZ24a, HA20a, HA20b, HA21b, HJC24, HTJC24, JHW21, KMT20, LTF^{+}20, MÖ21, RHA20, SW21b, WZLX20, WCLW20, WXZ21, WM20, XY22, ZSY24, ZJ24, ZZWS22]. regression-based [MÖ21]. regular [AH23, APS22, YC23]. regularisation [WMK22]. Regularity [DJ22b, LWYZ21, Bro23, Kar24, LZ23b, WZ21, ZZW^{+}24a]. regularization [BEJ20, BPR20b, BHB21, CZ23a, DH24b, FZQ22, GWHZ23, HL23, HSM20, HLY22, LCB20, LD22, MHW23, ME20, SL22, Six20, TWQ^{+}23, WYW^{+}24, WLW23, XXL21, YPL20, YZM^{+}20, ZSMG20, ZG20a, Zha21a, ZL22, ZZLY23]. regularization-based [MHW23]. regularize [Bon24]. regularized [BS20, HTZ21, JBSN22, KCL20, KDLY22, LF21, MSW23, MG22, PKMM24, QXZF23, WXZ21, YQZS21, YLZ^{+}22]. regularizers [LRT20]. regularizing [LZG24]. regulatory [LHE24]. reinforced [FFT20]. reinforcing [FFRS23]. Reinsurance [WZJQ21, LWZ21, MSK21, YCX20, YLH23, ZZK21, ZQL24, ZGL20]. Reinsurance-investment [WZJQ21, YCX20, ZQL24]. reinterpreted [DYZ24]. rejection [KFK^{+}24]. relapse [LHES21]. related [AM22b, CL24c, FL24, MSM24, SS23b, ZZWS22, ZCT21, Zhu20a]. relation [AM22b, MK24, SS23b, ZH23]. relations [dS24]. relationship [SLC22, XL20]. Relative [Mas21, YFW20, AAk20, CLM22a, Jea20, MLY22, TL21, Yan22, YLXY23]. relativistic [Cal23]. Relaxation [WL24a, HC23, LDY22, LWL22, LY23a, NK24, SCY23, WDS22, YLJW24, ZSHD24, ZZWX23]. relaxation-strategy-based [WDS22]. relaxation-type [LY23a]. Relaxed [LZF20, JH24, KZM21, MW21, SZZ21, SMRN24, TLD^{+}23, TZWG21, WLZ22, YJJW23]. relaxed-inertial [JH24]. Reliability [DNS21, DAZ23, GMSR18, GMSR21, WDTW20, Zhu22, Zhu24, BLBH21, BZHB23, Bha23, CCCMH24, DG20, GZ24, MS24a, NN22, Ozk21, TE24, Zhu23]. reliable [DL20b, HM21b, HVMS22]. relief [CLM22b]. remainder [CAJ24]. remeshing [CFO22]. Remote [YZM^{+}20]. removal [LRT20, LL23, ZLZ^{+}24]. removals [QG20]. renewable [BSdCCGV23]. renewal [GHF24, ÖS20, SDV21]. Rényi [FJMP24]. repairable [EKCF22, LTL^{+}24]. repeatable [KLZ^{+}24]. replacement [EP20, OKE24]. reporting [SB24a]. representation [ABBB23, Bal24, BEH^{+}23, CZ23b, GS21d, KA21, Vab21]. Representations [HQT20, BC21]. representative [EWT24]. Reproducibility [IBW^{+}20]. reproducing [BCDF22]. reproduction [BFRV21]. repulsive [JZW22]. resampling [MMPP20]. Research [LWZF21, LGMC20, LLG^{+}22, SZWH21]. reserve [LPY20]. reservoir [GNGGL^{+}19]. reservoirs [WCC^{+}19]. Residual [DLZ22, AC22, BJ22, BLZL22, BK20, CD23, Fal21, GW24, La 22, MS24b, SK24a, TA22, XWZ24, YSN24, ZWZ23b, ZH24b, ZZ23]. Residual-based [DLZ22, BLZL22, GW24, YSN24]. residual-driven [BJ22]. residual-time [BK20]. resolution [BFVY23, DKM23a, DBR24, ELTD23, FY22, FAVM19, FAVM20, HZ23, HH24b, KSG^{+}22, LRP24, NMST21, SSVS20]. resolvents [KWN20]. respect [BKMO21, Dju24, MMMM20]. respecting [CVPM22]. response [MCD24, MJK20, SLP20, SS23c, TZBK21, VS21, Xia22]. responses [CLM22a, HA20a, TKM21, TFK23]. responsibility [HC20a]. restart [FRR24, JYYL23, JH24]. Restarted [YJN20, ACR24, JLZ23]. restarting [BK20]. restoration [AKP23, BKMA24, BKM22, BR20b, CKY23, HWG21, JW23, JYYL23, KJO23, LZ24a, LNN20, SCY23, WY24b, WZL^{+}23, hYbJzJ21, YJJW23]. restoration-nonsmooth [KJO23]. Restricted [EHV22, JGW^{+}23, JGZ^{+}24, JLLL23]. restriction [RHA20]. result [SPKS^{+}24]. resulting [BMS22]. results [BCV24, CHLR22, EGG22, HIM24, gHtSjT24, IHK20, KK20b, LRZ23, LX24, MK24, RVS^{+}22, bWLgZbX24]. retirement [Dad20, MYLL24]. Retraction [AA21b, AA21c, GAA21]. retrial [AGGAVGG23, MULS23]. retrieval [MWLY24]. retrospective [SHV^{+}22]. return [LLW21, ZZ21]. reverse [MCY23]. reversibility [SXX21, TKVA21]. reversion [Wu24, ZSZ24]. reverting [AHB20, AHB21, LET23]. Revisited [Gök23, IHK20, LPP24]. Revisiting [PVV24]. revolution [GH23]. Reweighted [PN21a, WHlZH23, Kan20, LMZ24]. Rham [WWYZ21]. rho [BS24b]. ribbons [BRLS23]. Riccati [DL20c, LZL21, LWXL21, Wen21b, ZFwC20, ZKT20]. Richards [PCM21, SSM23]. Ridge [RHA20, BYE20]. Riemann [BBAO23, CM24b, CL24c, DP24, GS21c, Gu21, HF23, HZZ21, Li20a, Li20b, LLLL20, WZ23b, Zha21c, ZS20b]. Riemannian [BFF22, HKM24, SA23, SWN23, WDPY24, ZJY22, ZYZ24]. Riesz [ASADB23, JHG^{+}20, MD20, PAAI24, Ray23, SRMD22]. Riesz-tempered [PAAI24]. right [AZD21, Guo24, MDY20]. right-hand [MDY20]. right-side [Guo24]. right-sided [AZD21]. rigid [CM22, HKU22, HHL20, LR22, LSE24, NT23]. Rigorous [CLM21, MT20a, YTMK19, CLM22b, EN20, CLM20]. ring [NMB22]. RIP [GXC24, ZWH^{+}20]. RIP-based [ZWH^{+}20]. risk [BD22, BVBO23, CZ20, CX24, CCHW24, CY20, DLL21, HL22a, HQL21, KH22, KH24, LMYZ21, LLL22a, Mit20, PLZ21, QGW21, SZWH21, SZL20, SYG19, TBA22, TBA24, VLR23, WRZW21, WWC23, Wan24, WZLT23, XZ22, XHW24, Zha20b, Zha20a, ZCZ23, ZSZ24, ZY22]. risk-reducing [HL22a]. risks [BSNL19, BCMP23, BC24, CQBL20, DNK23, QG20, WTL20, ZZK21, ZZWS22, ZGL20]. Ritz [MR25, IMT23, Mam23, MR23, YJN20]. RK [HB22, Hu24]. RK-GSAV [Hu24]. RKDG [LY22b]. RKHSs [HN22a]. RKN [WW22]. RNA [HW24a]. Robin [Fan22b, SC23a]. robotic [YKAA21]. robots [JKJ^{+}23]. Robust [AS20a, BC24, DTM21, GBA21, HH21a, HZ24b, LPY20, VP24, WRZW21, YLL20, YLH23, ZX23b, Bog21, BDS20, BGIR24, CHZ21, ÇK24b, GMK20, Gra20, GK21b, GS24, HA20b, HA21b, HLMY24, HSC21, KKO23, LR24a, LR24b, MGV23, MER20, Ngo22a, OPK23, PAAI24, Pou21, RMG21, Rou22, SCHS24, SBKR24, TK24a, VPHE23, WW24b, ZSY24, ZZW^{+}24b, ZZW^{+}24c, ZC20a]. robust-adaptive [TK24a]. robustly [MARH^{+}23]. Robustness [CMP21, HTOC20, PVV24]. role [AHK24, KLM22]. rolling [VOJS24]. root [ABO24, CL24b, HP22, JZM^{+}23, Moi22]. root-solvers [HP22]. roots [BCT22b, NZ21, RSS24, SBKR24, Uga24, WX21, ZL23b]. Rosenbrock [PS24c, TGS21]. Rosenbrock- [PS24c]. Rosenzweig [HE24]. Ross [KLM22, ZSY24]. Rotated [HF23, MCL22]. Rotated-hybrid [HF23]. rotating [OGR22]. rotation [RR23b, WW23]. rotational [AHGG20, LWA^{+}22]. rotations [ÇÖÖ23, Nov24]. rotorcraft [MBG21a, RMG21]. Rough [CGCMRP24, Hai23, QGW21]. roughness [GBG23, GHL21]. round [LMS21]. round-off [LMS21]. routing [GM23]. rowwise [ZTY^{+}24]. ruin [CX24, Dad20, LSY24]. rule [CYC21, Den24, FPLZ21, GH22, RG21, SWW23, SDW24a, Xia22, ZZ22a, ZSHD24, vdBS21]. ruled [AQ20, BAIB24]. rules [AVRLRMGG23, APR21, APR24, BPDC20, BEJR22, BPR20a, BPR21, DG21, DDRS21, DDRS24, EMR21, GJ23, Gol24, JPS24, LMAV24, LL24e, Ma24b, Maj23a, Maj23b, MSM24, ORT22, OH23, PPSM24, RS24c, SuR23, Tom24, Ver20, dlCYS22]. Runge [ÁGKM21, ASML23, Bha21, BP22b, ESS21, FZ21, FCS22, GS20c, HR21, HCH20, HW24b, HKL21, KFK^{+}23, KFK^{+}24, LF21, MM21, Moi24, MF24, MRC23, MRC24, NS24a, NNJ24, RAN20, RBR23, SL20, ST22, SW23, WH24, YLK24]. running [WWC23]. runs [Kon22].

S [HCCSCGAM24, YY21]. S-SPAM

[HCCSCGAM24]. Sabin [BEIR23, GS21d]. saddle [BG23, BLyS20, Cao20, FZ20, GH20, LZF20, LPP24, WL22, YL21b]. saddle-point [BG23]. Saddlepoint [ZRQ22]. safe [VLR23]. Sahalia [EM21a]. SAHARA [CNS21]. Sakurai [PI24]. sales [LWC20]. saliency [GRS20]. Salpeter [BP22a]. Sample [KLZ^{+}24, PM20a, CLS20, CMJN22, New24, NMF^{+}22, YL20, YL21c, ZSY24]. sampled [AUAA20]. sampled-data [AUAA20]. samples [CK20, GC19]. sampling [BBTW24, BK22, BCV24, CMP21, CD23, ÇK24b, CBF24, GANU21, HG23, JL21, KSG^{+}22, New24, NL23b, RZA^{+}24, SJZ20]. SAR [HKU22]. Sard [FB24]. Sarmanov [HBMA22]. SARS [BYE^{+}23, HCCSCGAM24, ÖHS23]. SARS-CoV- [BYE^{+}23]. SARS-CoV-2 [HCCSCGAM24, ÖHS23]. satellite [HKU22]. satisfying [CC20]. saturated [AK21, LWX24, RCN21]. saturation [WY24a]. saturation-value [WY24a]. SAV [QH23, CLS21, HJM24, HCC24, LLG24, Tak24, YY24a, YYC24, ZaZWZ23]. SAV-FEM [YY24a, YYC24]. scalable [Gra20]. scalar [AAP24, CCL23, Cap20, FFS24, JS22a, Kri21, MN22, NMST21, QH23]. Scale [PH22, AB22, BTC21, CH20, CYY23, DYM21, DV21a, GIV20, GW24, HOOT20, HL24b, HLY22, JLZ23, KPD23, LZL21, LZD21, LMSZ21, LWLW23, LLX^{+}22b, MGV23, PHV24, RYH24, RH23, WXZ21, WLCH23, WY24b, WAH22, Wen21b, WSL^{+}23, hYbJzJ21, YJJW23, Zha21a, ZLZL23, dOdO23]. scale-invariant [MGV23]. scale-shape [AB22]. scaled [Bra23, LZD21, LLL24a, YWL24]. scales [RFGGM20]. scaling [MIdMR24]. scanning [RV20]. Scattered [KH21, CDD^{+}24, DDNZ21, FGL23, HSW21, ZFH20]. scattering [ABA20, AZ20, Bal24, CLM22b, DT24, ETKF23, KVHC23, Rod21, XLY22, ZCSL21, ZG24, ZZNL24]. scavenger [KK24a]. Schatten [LL24f]. scheduling [FRT23, LPLP21, WZC22b]. scheme [Abi20, AFLP22, AAP24, AHB21, AAHK24, AJR22, AET23, BA22, BBMK22, BCT22b, BDPR23, BB20b, CMW24, CCS24, CZY22, CZHQ23, CHXL20, CGW21, CS21, CG23, CJL^{+}24a, CMS20b, CQZ22, CDL22, CVPM22, DL21, DHKV20, DWWZ22, DHQZ24, DKM^{+}20, EO24, ETKF23, ÉM21b, FSC24, FRT23, GJDL20, GGHY24, GLCL20, GWWY21, HET24, HJM24, HA21a, HVR22a, HVR22b, HH24a, HW20, HPW22, HAD24, HZZ24, HSYZ24, JYY21, JZZH22, KAD22b, KK20c, KK21, KCRVA22, KD23, KRV24, LL20a, LLT^{+}21, LLZY22, LS23, Li23, LLLL20, LD21, LYL23, LST^{+}24, LDM23, LM21b, LR24a, MD20, MD22, MDG22, MBZ23, Mia24, MZY23, MDD21, NSL22, NA21, NBLZ24, OWV24, OMRdQ24, PT21, PAAI24, PM20b, PFG24, QQSZ24, RS22b, RG23, SES20, SRMD22, SRBK23, SDC20, SZZ21, SDLZ23, SY24, SA23, SKVA23, Son24, SS24b, SCZ20, Tak24, VAMO24]. scheme [VP24, WY20, WCPW20, WOV22, WSY22, WCLY23, WZY23, WWW^{+}24, XDX24, YOCY22, YLLA23, YWW24, YY24b, YZ24d, ZVT^{+}23, ZZW^{+}24b, ZZW^{+}24c, ZHWD20, ZCYP20, ZHY20, ZCF^{+}22, ZWZ23c, ZPW^{+}24]. Schemes [For24, TFR22, ABP24, AI22a, ANR20, ADDG22, AS21a, ADEEY21, ÁGKM21, BitH21, Bro23, Cap20, CG22, CV20, DH20, Dar20, DFFM21, DL23b, ELS21a, GZ23, GM22, GANR22, GGT24, HBGS21, HJ23, Han24, HCC24, HVM22, HZS^{+}24, IKL24, JZW22, Kri21, KPS21, LYK23, LY20, LML21, LLYM21, LWA^{+}22, LLS22, LL20b, LY21, LY22b, LF24, LET23, MDVR22, MDS20, NKBJ20, PJA24, QH23, RlLZZ21, RWW22, RK23, WW23, WYJ23, XWXZ22, YLJW24, YKY24, YF22, ZZG20, ZJGZ20, ZOW^{+}21, ZLS23, ZY24, ZW22]. schlichtness [NS20]. Schmidt [MAA^{+}22]. Schmidt-Schwetlick-Kurchatov [MAA^{+}22]. Schnirelman [YL21b]. Schoenberg [FL24]. Scholes [RG20b, ACD22, AF20, CLS24, Kaz23, OT21, RHM22, RG20a]. Scholes-type [AF20]. Schrödinger [BBGI24, Alj22, BV22, BK24, Bra23, Cal23, CD20a, CQZ22, DVRG21, DVRV22, DL23b, DW20, DW23, HR23, HC20b, HC21, HZZ21, KAD22b, LGY24, LY23a, MM20a, PJA24, Ray23, SW24a, WYD20, WZ23b, WL24b, Xin21]. Schrödinger-type [CD20a]. Schur [CMVV20, GBM23, LZ24b, SSB20]. Schwarz [Axe20, WCK20, Xin21, YS24b]. Schwetlick [MAA^{+}22]. Science [VAAR22]. Scientific [BJMR20, ZDF^{+}20]. sclerosis [dPdMQL23]. Score [CMdMRFG22, CECCCR23, VBRT24]. Scott [HRT20, Par20, ZWZ^{+}23a]. scramble [VS21]. SDEs [CG20, EL22, LGC23, MP21, PSS24, RS22b]. sea [USTG23]. search [HCCSCGAM24, JMHZ24, KL23b, LZD21, LH20, Mim20, PH22, YKAA21, Zho20]. searching [Zha21c]. seasonality [CCJ20]. seawater [AET23]. Secant [dSFSS24, EHVMM23, EHVMM24]. Secant-inexact [dSFSS24]. secant-type [EHVMM23]. Second [Ela23, JLM24, JCLC20, MBZ23, MF24, RAN20, ZG20a, ZW24, ASML23, AAHK24, BT21, BHRW22, BBSV21, BA23, BDS20, CZY22, CZHQ23, CLX21, CAJ24, CHH20, CJL^{+}24a, DCP22, DHH24, DL20b, DCAA22, EKÇ20, EPL21, FHY^{+}21, FMNT24, GJDL20, GWWY21, HWCJ21, mHLZ20, HE24, HL22c, HWAT22, HH24c, JQ22, KN20, KCKK24, Kaz23, KXW23, KAA22, KD22, LML21, LLZC21, LLZY22, LLX^{+}22a, LWWY24, LZ23b, LXF21, LXYL24, MGD22, Ma24b, MMMM20, MZY23, Moi24, MRC24, MFA21, Ngo22a, NAS22, OGR22, OWV24, PL21, Pou24, TH24a, UGG^{+}20, WCPW20, WWJ22, WOV22, WWW^{+}24, WWZ24, WSY20, XLKL23, XH20b, YCXW24, YT24, YOCY22, ZCY^{+}20, ZPW^{+}24]. second-kind [Ma24b]. Second-order [RAN20, ZW24, ASML23, AAHK24, BT21, BHRW22, CZY22, CZHQ23, CLX21, DCP22, DHH24, DL20b, EPL21, FHY^{+}21, FMNT24, GJDL20, GWWY21, HWCJ21, HE24, HH24c, KCKK24, KD22, LML21, LLZC21, LLZY22, LLX^{+}22a, LZ23b, LXF21, LXYL24, MGD22, MMMM20, MZY23, Ngo22a, OGR22, OWV24, TH24a, WOV22, WWW^{+}24, WSY20, XLKL23, YCXW24, YOCY22, ZCY^{+}20, ZPW^{+}24]. sections [DFK20]. sectorial [RVS^{+}22]. Secure [GGPP21]. Segel [IYA^{+}23]. segmentation [GRS20]. segments [FPS21]. Seidel [MS24b]. SEIR [SMR23]. seismic [WSX21]. Selected [SC23b, DDT23, PKK^{+}22, PL24b, APW^{+}21, AHK^{+}23]. Selection [FJMP24, AD21, AFJGJJ22, AMRP22, CSZP20, CL24b, CYY23, DOT23, DZLC20, EN20, GBA21, hKhWjH21, LLW21, LMZ24, SL23, Tor21, VTB21, WCLW20, WCPW20, WYJ23, WSZ20, WM20, WMK22, WLZ22, WZLT23, YZ24a, ZMOW23, ZCJL20]. self [CJCDLL22, HWLD23, NY24b, PL20, RH23, SH21, TH24b, ZLZL23, ZL23b]. self-accelerating [ZL23b]. self-adaptive [HWLD23, TH24b, ZLZL23]. self-climb [NY24b]. self-contact [RH23]. self-similar [CJCDLL22]. self-sustained [SH21]. self/completely [PL20]. selling [SYG19]. Semi [AAA^{+}23, GJL23, GHJM24, ZIZ21, AFLP22, AAP24, Bro23, CZHQ23, EEE21, FFT20, FV24, HS24a, Hur22, JZZH22, Liu20, LC23, LM21b, MBZ23, Mim20, MER20, MDD21, NBLZ24, OSK20, OF20, QXZF23, RHA20, WYTL19, WLDZ24, YX20, ZLJT22]. semi-absolute [MER20]. Semi-analytic [GJL23, GHJM24]. semi-analytical [FFT20, FV24]. Semi-decoupling [ZIZ21]. semi-definite [Liu20]. semi-definiteness [LM21b]. semi-discrete [AFLP22, AAP24, Bro23, EEE21]. semi-explicit [Mim20]. semi-implicit [Hur22, JZZH22, LC23, MBZ23, MDD21, YX20]. semi-infinite [HS24a, OSK20, OF20, ZLJT22]. semi-linear [CZHQ23]. Semi-local [AAA^{+}23]. semi-parametric [RHA20]. semi-proximal [QXZF23]. semi-rank [WYTL19]. semi-sharp [NBLZ24]. semi-tensor [WLDZ24]. semiclassical [AV21, WL24b]. semiconductor [LLGC24b, YLS23]. semidefinite [Cao20, LZZ20, MN21]. semidiscrete [YZ21a]. semidiscretization [Kov22]. semigroup [LC20]. semilinear [Cso20, FZ21, LVW24, LCZL24, WLJ24, YL21b]. semilocal [AAB^{+}20]. semiparametric [YLL20, Zha20c]. semipermeable [HW21b]. semipositone [LMZ22]. Semismooth [BFVY23, JG21a, Zha21a]. sense [GGHY24, PPSM24]. sensing [LWLW23, YZM^{+}20]. sensitive [GH23, VS21, ZSY24]. sensitivities [XW22]. Sensitivity [PN21b, PNS21, SLP20, TTK23, Bha23, OPK23]. sentencing [FPGSR21]. Separable [XD23, HLLZ24, HVMS22, SZZ21]. separated [GJW23]. separation [CJLL21, HF21]. September [Ano21-35, Ano22-36, Ano23-36, Ano24-33]. septic [BK24]. sequence [GHC^{+}20, YYT24]. sequences [BG21, BCV24, BRZ22, DP21c, SS23b]. sequencing [HW24a]. Sequential [LMM22, ADMV22, ABM^{+}22, BB20b, HB22, OF20, SZZ21, SBM21, WYW^{+}24]. serially [ZZM20]. series [BABR20, BYE20, CGSVBB^{+}24, CECCCR23, DMST21, DIP^{+}22, EN20, EK24a, FBJ20, GMK20, LWC20, LC24b, LN23, MGG^{+}24, Mil22, PA21, SUW23, Wan24, YCD^{+}20]. series-parallel [FBJ20]. server [SA24, SACQ21]. service [BS24a, SVK20]. set [BK22, Bra23, CGCMRP24, FYL^{+}24, GN24, Juk20, KLZ^{+}24, LR22, LZH21, LL21, New24, PPSM24, PS24b, QGW21, RZA^{+}24, tSjTsJ24, ZZ24]. set-oriented [FYL^{+}24]. set-valued [ZZ24]. sets [BHB21, GGPP21, HLW20, HTTO20, LS24, WTWZ20, WV21, Yam20]. setting [GQR23, RKS^{+}22]. seventh [BA22]. seventh-order [BA22]. Several [TZWG21, JZ24]. SFDEs [LD21]. SGD [WLCH23, YZ24a]. SGFEM [ZZL20, GLZ24]. SGLM [HK21]. Shafer [HH23]. Shakhov [TFR^{+}20]. shale [ZZWS20]. shallow [DL20b, KD20a]. Shape [Ben24, CWLM22, ASX21, AB22, CM24a, Fan22b, KP20, Khl21, LT21, YKY24, ZSVL20, ZHW20, ZLL^{+}20b]. shape-based [ZLL^{+}20b]. shape-preserving [YKY24, ZSVL20]. shaped [DDF24]. shapes [GBG23]. Shapiro [MPGE24]. Shapiro-Wilk [MPGE24]. Shared [QDHW21, GM23]. Shared-memory [QDHW21]. sharing [San21]. Sharp [RlLZZ21, YT20, AET23, DSCF24, NBLZ24, WJ23]. sharp-diffuse [AET23]. Shaw [JGLH20, WWJ22]. shearlets [BEH^{+}23]. shedding [RMG21, RDG23]. sheet [WD20]. shell [KNM24]. shells [GH23, TL22]. Shepard [DDNZ21, DDF24, Dum24]. shift [JL21, KCRVA22, LMSZ21]. shift-invariant [JL21]. Shifted [KKJ20, BVDB20, Bra23, DO22, ESA23, GHC^{+}20, HZC^{+}23, SN21b]. Shishkin [HO22, WLM24]. Shiu [XZ22]. SHNC [CTGV20]. Shock [GHF24, BKMO19, BKMO20, CUN24, Cha24, DAZ23, GX24, KO22, KT24, LNP19, LN22]. shock-capturing [CUN24]. shocks [LN22]. shooting [GJL23, JG21a]. shooting-like [JG21a]. shop [FRT23]. short [CJLL21, Har21, MN20, SYG19]. short-period [Har21]. short-time [CJLL21]. Shortest [GANU21]. Shortest-support [GANU21]. Shortley [SCZ20]. shot [BCMP23, BJP24, BC24]. Shrinkage [SVJG24, AJ21]. shunt [Ma24a]. side [DR23b, Guo24, MDY20]. sided [AZD21, DKMT24, JWZ21, MSS^{+}21, TGLa21, YLFT21, ZEW23]. SIF [JO24b]. sign [APS22, JC24, SK24b, Tom24, YC23]. Signal [Kog21, CJLL21, JHYZ24, JW23, Kan20, RYH24, SY23, SCY23, SZZM23, WTP^{+}21, WSL^{+}23, XTL23, YJJW23, Zho24]. signals [CD20b, CHsL22, DKSQ22, FS24b, ZZLY23]. signature [BLBH21, BZHB23, CCCMH24, YC23]. signed [AL20]. significant [EHVMM23]. Signorini [SP22, WHZ24]. SignReLU [LFZ24]. SIMD [SFI23]. similar [CJCDLL22]. Similarity [PP22, FLMR20]. simple [Alh24, CFO22, FZ20]. simpler [JWSG22, YJN20]. simplices [HAFA22]. Simplicial [AHGG20]. Simplified [CK23]. simulate [ADKH20]. simulated [GBG23, ZMOW23]. simulating [BIZ20, Hur22, LM21a, TGJY22]. Simulation [WAAM20, AGS^{+}23, AGST23, AST^{+}23, AVL22, BHRW22, ELTD23, FS20, FHH21, GR22, GeOJD24, GDG^{+}21, HRT20, HLLY24, HBPV21, HF21, HPW22, KCJ23, LYM22, LLGC24b, MDVR22, MLYZ21, MS20b, MDD21, NMF^{+}22, RDG23, SPMB21, SKS22, SH21, Svá23, WK20b, XW22, XHW24, YCG^{+}23, ZZWS20, ZSVL20, GQR23]. simulations [BB21a, CKT21, GM24, IYA^{+}23, Kog21, LGLC20, LP23b, NGNB23, OBB23, PCMH20, PCM21, PH22, SŁK^{+}22, UÖA23, WCK^{+}22]. Simultaneous [CLL20a, Fan22b, RZ20, HMA24, HSGLE19, LMAV24, PH20b, WLW23]. simultaneously [GHC^{+}20, JA22, QX24]. sinc [Bra23, JLZ22, NAA20, OSK20, QXZG22].

Sinc-collocation [NAA20, QXZG22]. sinc-Galerkin [JLZ22]. sinc-Galerkin/finite-difference [JLZ22]. sine [Bac19]. sine-Gordon [Bac19]. Single [IKL24, CBF24, CLM22a, HET24, HWG21, HW24a, LLL24b, Pou21, SA24, SACQ21, SKS22, TDAT23, ZZM20, ZZQ21]. single-index [CLM22a, ZZM20, ZZQ21]. single-parameter [Pou21]. Single-temperature [IKL24]. Singular [KO22, WLZ23, ADZ20, ADDG22, All24, AN24, AAA20, Bon24, BG23, BPR20b, CZHQ23, CJP20, CS24, CYC21, CDW20, CJL^{+}24b, CHSZ24, DTM20, DTM21, DCP22, DC24, FYL^{+}24, FO22, GSK20, GPK23, GH20, HZ24b, KZ23a, KZ24, KN20, KM22, KEKT21, KXW23, LZ20a, Li20a, Li20b, LQS23, Lu24, Naj20, OA20, PT21, PL24a, PZ21, RB19, RCM20, Rou24, RL23, Shi21b, SHZ23, WSLJ23, WJVG24, XZZ^{+}20, XTL23, YSN24, Yam21b, YS24a, YZL24, YWZ24, ZE23, ZBS22, ZM20, ZLCT22, ZIZ21, ZS22, ZHZ24]. singular-oscillatory-Bessel-type [CYC21]. singularities [KLMP21, SL21]. Singularity [AZD21, Gao22, KWZX23, OWV24, RR23b, SKVA23]. Singularly [GO20, GO24, AVL22, Bog21, CLX21, Che21, CJ23, CSS24, DZ22, DLZ22, DCAA22, Ela23, ER20b, HWRM24, KAA22, KSVA22, KCRVA22, KL23a, LC24a, LYZ24, NMMH23, PSRVA24, RS24a, RS24b, SKVA22, WK20a, WLM24]. sinh [CECCCR23, XZZ^{+}20]. sinh-arcsinh [CECCCR23]. sinks [EFTW21]. SIR [Pro23]. SIS [CGW21, YPLM22]. site [Kap21, NGNB23]. Sitter [MD22]. Sivashinsky [AL20, FC23]. six [FKV21, LC23]. six-order [LC23]. sixth [BJB20, MGG^{+}24, RG20b]. sixth-kind [BJB20]. size [CMJN22, DKSQ22, GLHJ21, KAD22b, MV22, SSRK23, WMH22, ZCCK24]. sizes [New24, NMF^{+}22, SLW21, SZZ21, SBM21]. skew [AB22, HK21, LY24a, MPGE24, WYD20, ZY22]. skew-elliptical [LY24a, ZY22]. skew-Hermitian [WYD20]. skew-normal [AB22]. skewed [AS20b]. slab [PMVB22]. slacks [AJSGL22]. slacks-based [AJSGL22]. slashed [OVGI20]. slashed-Rayleigh [OVGI20]. SLBS [EY23]. sliding [TK24a]. slip [AAA20, WLHJ20, ZOK21]. slope [WCPW20, WYJ23]. slow [ZVS21]. slow-fast [ZVS21]. slugs [BEK^{+}23]. Small [HSRA24, BD20, FY22, Gua20, KPD23, SMDVA22, WWWY23]. small-scale [KPD23]. smart [BD23, WQ20]. Smoluchowski [KK23, LYYZ24]. Smooth [BAIB24, AC22, ACEIJH22, BB20b, Bre22, FP24, GO24, GS21d, KM22, KLMP21, La 22, LHL21, MH21, MMM22, MG23, MDKG20, WZC22b, Zhu21]. Smoothed [HB23]. smoothing [BFVY23, LZ20b, Naj20, SD24, TR22]. smoothness [GSZ22, HC24]. SMVIPs [YL22]. Sobolev [CZHQ23, DKM23a, Kar20, NP23, Rab21, XFW23, YTA22, ZW24]. Sobolev-type [CZHQ23]. SOC [ZL22]. social [HC20a]. SOCP [ZZWX23]. SODES [GGRCGG^{+}23]. soft [LHS^{+}24, ZS20a]. soft-margin [LHS^{+}24]. software [AHK24]. soil [PA24a]. solenoid [OPK23]. solid [CJCDLL22, FDDM24, ZWKS24]. solids [LYM22]. soliton [KAS23, ZAH^{+}24]. solitons [MGA24]. Solow [UV24]. Solution [APW^{+}21, AI22b, NT23, RK22, SM23, tSjTsJ24, TGLa21, ASB22, ASA^{+}21, AA24b, AA24a, AHK^{+}23, Ask21, ALKH21, BJB20, BBTZ20, BBGI24, BLN21, BGV20, Bra23, CK22, CK24a, CS24, CJM20, COT21, CHSZ24, CUUdB24, DT24, DAM23, DCAA22, EGG23, FFS24, FMRS24, GM22, GSK20, GPK23, GEG23, Gol24, HMRZ23, HCH^{+}23, HC20b, HTTO20, IHK20, IpAmK20, KM22, Kaz21, KMN20, KS24, KZM21, KLK21, Kum22, LBBB23, LG23, LWXL21, MK24, MGV23, MJR20, MRC23, MRC24, Naj20, NZA24, ON22, OOO22, PMVB22, PP24, PL24a, RDHA20, Rab21, RO20, RS24b, RFGGM20, RTT24, RGA20, Rou24, SM20, SMM22, SuR20, SS24a, TSA23, TWQ^{+}23, TZTC20, UV24, Vab21, Vab23, Vab24, WEEA20, WW22, WSC23, Yam20, ZBS22, Zha21c, ZS20b, ZWX24]. solution-dependent [WW22]. solutions [ASADB23, AA21a, AEST21, AAKA24, AF20, ATO22, AKP23, BuR24, BGdlR24, BCVV20, CTN22, CT20, CW24, DRR22, DPW24, DZLC20, DV21b, EOT22, FM22, FPLZ21, FP24, Fro24, GLHJ21, Guo24, HLW22, HIM24, Hor20, HM23, HWLD23, HWD^{+}24, gHtSjT24, JM21, KAS23, KLH21, KM22, KF24, LGC23, LZ20a, Li20a, LGY24, LYYZ24, LY24b, LZ20c, Lu24, MH21, MMJ21, MHM23, Mas21, MT20a, MPV23, Mik23, MORO20, Miy22, MP21, MVFAMF21, NL23a, Ois20, ÖKG23, PO21, Pie20, Pou24, PSS24, Pul22, Qiu20, RHM22, SPMB21, SC21, SCK22, SW21a, Sun20, SD24, SL24b, Tan20, TH24b, WXS21, WYW^{+}24, bWLgZbX24, XLZ20b, Yam21a, YC23, YL21b, ZAH^{+}24, ZUZ20, ZB22]. Solvability [BBAO23, DK24, Li20b, MN21, Li20a]. Solve [CM24b, BJB22, CJ23, HO22, HVMS22, HAD24, HL24b, JHYZ24, JYYL23, JNCY23, MD20, MD22, MA21a, MCMR23, NAA20, PH24, RDHA20, Rab21, RPQ24, SRMD22, XHLD21, YZ24a]. solved [KS23a]. solver [BDS20, FHH21, GHG24, HF23, LPH23, LWW^{+}22b, Ngo24, OBB23, SGD20, TL20, VSP23, VP24, WTL21, Yam20, YZL24]. solvers [All24, HP22, KT21, LMZ22, MVY19, Mar19, MAA^{+}22]. Solving [BGG^{+}21, CDTFGV22, CM22, CNQRR22, DAMA22, GGRCGG^{+}23, GS21a, HJ23, HVR22b, LLL^{+}22b, PK23, RR23a, TL22, TSL^{+}21, UGG^{+}20, WCMW23, Wen21b, ZZNL24, ADZ20, AORS22, AID24a, AID24b, BKMA24, BLZL22, BR22, BP22a, BP21, BB20b, Bre22, Bro23, Bün20, CHH20, CV20, CVPM22, DR23a, DO22, DS22b, DHF23, ER20b, FZQ22, GJN20, GBTS22, GLCL20, GGO21, GHC^{+}20, HWCJ21, HVR22a, HVM22, HSMT20, Hua21, HLY22, IAH24, JGW^{+}23, JG21a, JZY22, JWS23, gK23, KKJ20, KDLY22, KL23b, KS21, KRV24, La 22, LWD20, LDY22, LWL22, LZZZ21, LDC20, LET23, LHL21, LH23, MHW23, MPM23, Mam23, MHM23, MV22, MH23, MW22, MG23, Ned22, NAS22, PT21, PRW^{+}22, PAAI24, PKE20, PKMM24, PLS22, RR22a, RB19, RTT24, RRS24, RB23a, Rou24, RR23b, SES20, SDLZ23, SCY23, SK23a, SZA24]. solving [SHV^{+}22, SLW^{+}22, TBA24, TLY23, TH24a, TDL^{+}22, TLD^{+}23, TDAT23, UPPZ24, Vab22, WLHJ20, WCW20, WDS22, WLCH23, WL24b, WAH22, WDY20, WSX21, WLZ22, XHG22, XY24, XFW24, XA21, YTD23, YSFdRdP20, YF22, hYbJzJ21, YJJW23, YCH^{+}21, YZLL22, ZE23, ZTZZ20, ZHY20, Zha21a, ZZ22b, ZZ22a, ZLZL23, ZSHD24, ZYL20, ZS22, ZCCK24, dSFSS24, vdMOB22]. Some [Fro24, KUA20, KWN20, LS22b, NNCH21, NRV21, Pou24, SS20b, TA22, Vab22, YZ24c, AZ22, CPRV22, FU19, GPST23, HZMD22, Hor20, JBSN22, KJGN20, Kar20, KKA20, KDLY22, LC20, MD20, NS20, RSS24, SHB24, Seg23, ZZ24]. Sonine [ARF23]. SOR [GH20, HL22c]. SOR-like [GH20, HL22c]. Source [DZLC20, AA24c, BB21b, CL24a, CNYC23, DC20b, DLC21, DL20b, EFTW21, GPK23, HTIA24, JLW23, LRP24, LYYZ24, MHHV20, MBG21a, NY24a, OHK23, QX24, RS24b, RZ20, Sha23, ZAC^{+}23, ZDL22, ZGZ23, ZMYL23, ZF24, ZZZ^{+}22]. source-free [ZZZ^{+}22]. source-sinks [EFTW21]. sources [Ben24, BM24a, ETKF23, LWYZ21, Tam20, YCH^{+}21]. SP [BS23]. Space [HPW22, ADZ20, AHPS22, Bac19, BT21, BBTZ20, BB21b, BRZ22, CLL20a, CZHQ23, CLH22, CQZ22, DKM23a, DWWZ20, EBVB21, FLT20, FZQ22, GDZN23, HJL20, HM22, HQT20, HL21b, HL24b, HZZ24, JWZ21, JHG^{+}20, JLW23, JWS23, JZ24, KKA20, KWN20, KLMP21, LRP24, LWG24, LL20b, LCW20, LT24b, LY23b, LXF21, MD22, MV22, MB24, MMV20, PH24, PM20b, Pie20, Rab21, Ray23, RFB22, RSRW24, SM23, SRMD22, Sha23, TH24a, TDAT23, TFR^{+}20, VV23, VVB21, WZ20a, WLJ24, WFZ22, XZZ^{+}20, XLLA21, YLFT21, YLLA23, YY24b, YZ24d, ZZ22a, ZGZ23, ZF24, ZS20b]. space-dependent [BB21b, CLL20a, JLW23, Sha23, ZGZ23]. space-fractional [DWWZ20, HL21b, JWZ21, Ray23, SRMD22, TH24a]. Space-time [HPW22, BT21, BBTZ20, CZHQ23, CLH22, LCW20, LXF21, MD22, MB24, RSRW24, SM23, YY24b]. spaced [BE21]. spaces [All24, BGR20, BEL22, BCDF22, BCV24, BM21, DHF23, FO22, GG24, GHTC21, HWLD23, HbX24, HC24, JL21, KK24b, KDLY22, MDE24, PT24, RCM20, tSjTsJ24, TDL^{+}22, TLD^{+}23, TH24b, WWYZ21, ZCCK24, dA23]. spaces-based [GG24]. SPAM [HCCSCGAM24]. spanning [JGW^{+}23, JGZ^{+}24]. Sparse [GMCJGM24, NYL23, SSS24, SH24b, Bha23, JHW21, JHYZ24, JW23, LYB22, LMZ24, MLYZ21, MCY23, QXZF23, SKZ23, SZZM23, SD24, WXZ21, WHlZH23, WZ24a, WYW^{+}24, YZ24a, YJJW23, Zha21a, ZZLY23, Zho24]. Sparsest [DDUF22]. Sparsifiable [ZXZ24]. sparsity [BB20b, DH24b, LLLD20, LZ24a, SCY23, Zho24]. sparsity-induced [SCY23]. Spatial [FY22, BS20, DKM23a, GZ23, GS21c, HS20, HN22b, KPS24b, LL20a, MDVR22, ML23, SWW23, SZ23, UV24, WY21b, XL20, ZYZ22]. spatial-temporal [SZ23]. spatially [CPBG23, Wol24, YZ21a, YCD^{+}20]. Spatio [BYWA22, HCCSCGAM24, BEH^{+}23, FGN^{+}23]. Spatio-temporal [BYWA22, HCCSCGAM24, BEH^{+}23, FGN^{+}23]. spatiotemporal [YO20]. Spearman [BKMO21, BM22a, BS24b, SHZ20]. Special [DFH21, FMRS24, TE24, AKP23, BLV20, BSZ21, ZuIH20]. species [BB20a, KK24a, TKVA21]. specific [Gök23, HWLD23, ÜY20]. specification [HSGLE19]. specifications [ARMP20]. spectra [ABFSC24]. Spectral [ASADB23, ADEEY21, AS21b, CZY22, CLST20, CZH21, ESA23, FSR24, Gu21, MZY23, TN24, YS24b, ZLJT22, ADZ20, AEST21, Alj22, AZD21, AF23, BR20a, CC21a, CLHL20, CLH22, CNRR20, EOR21, ER20b, EPL21, FLMR20, GEG23, HZ23, Hes20, HR22, KKJ20, KCK24, LWLW23, LWG24, LCH23, MH21, Mag20, MDKG20, NM24, PL21, PMVB22, RFGGM20, SW24a, WZZ21a, YX20, YTMK19, YWZ24, YF22, ZVT^{+}23, ZWZ^{+}23a, ZLCT22, ZJBY23, ZXZ24, ZCZ24, WZY23, dLNO21]. spectral-collocation [BR20a]. Spectral-Galerkin [CZY22, MZY23, KKJ20, WZY23]. spectral-like [HZ23]. Spectrally [BBTZ20]. spectrum [Gao22, PT20, dF20]. speed [HF23, JZZH22, Kan24, PMVB22]. speeding [SNMK21]. Speedup [KT21]. SPH [LV20, SLA^{+}24, ZWKS24]. sphere [AGS24, CJP20, HSW21, HG22, IN22, BBC22, Bel23b, BC23]. sphere-Frames [IN22]. Spherical [Seg23, BAIB24, BC23, CTGV20, CMS20b, HH21b, LJ23a, LC24b, TLY23, WWWY23, eBI23]. spherically [PT20]. spike [HCCSCGAM24]. Spillover [PCWY20]. Spin [IN22]. Spline [BEIR23, BEL22, BCDF22, BDF21, BC22b, BL21, Bos24, BGIR24, CF20, CTN22, CGL22, DDF23, DU19, FS24a, GANU21, GS21d, HA20a, HC24, Kra22, LNSZ22, LGY24, LNP22, MA19, MA21a, MJ24, RG22, SLZL23, TFM24, WZNT24, ZM20]. splines [BEL23a, BPDC20, BC21, BCR22, BC22a, Bic21, Bic23, FAUU20, GeOJD24, GS21d, KR21, KRT23, NWD20, PJM21, Seg21, SKVA23, FM21c]. split [ACK23, BEIR23, BP21, BHB21, EK24b, HWLD23, LHHW21, LDQF23, PKMM24, PLS22, RTT24, TH24b, WJVG24, WLDZ24, Wen21a, YPLM22, YZ21d]. split-step [LHHW21, LDQF23, YPLM22, YZ21d]. Splitting [JHYZ24, KSZ21, BitH21, Cao20, CW21b, CJ23, DWWZ20, DL20a, EOR21, Fan22c, HV23, HPW22, Hua21, HC23, ISD23, JLY^{+}21, JMLX23, LYK23, LL24b, LL20a, LDY22, LW22, LWL22, LS23, LW23a, LT24b, NM24, PLS22, PFG24, RVX20, SLXF24, WYD20, WL24a, WZD^{+}24, XSL22, YYZB24, ZWZ^{+}23a, ZWL21, ZCSS24]. spread [AMT23, Hai23, Ngo22a, Zha21b]. spreads [ABO24]. Spurious [Par20, FJ21, HB23]. SQH [Bre22]. SQP [JLY^{+}21]. square [ABO24, BCVV20, HH21c, JZM^{+}23, LHHW21, LHHW23, Moi22, SVJG24]. square-root [ABO24]. Squares [SAL21b, AKP23, BC23, BGIR24, CCL23, CGMM20, CKL23, CUUdB24, DMNO24, Din23, HM21a, Hes20, HSW21, HG22, HH24c, HKM24, JM21, KL23b, LCGW24, MH22, MI20, MDKG20, RR22b, Saj23, SL22, WYZZ22, WD20, YZZ20b, ZZWS22, ZJY22, ZFH20, ZH24c, ZWX24]. SRKCD [SW23]. SSP [HR21]. stabilisation [LSC21]. stabilities [YZ24c]. Stability [AA23, AA24a, AB23, Ask21, ACEIJH22, DG21, GLHJ21, LLS22, LDQF23, LY23b, LP23b, MVFAMF21, SC21, TGJY22, TS20, WKN22, WZD^{+}24, YX20, YBM23, BCT22a, BIJ21, BM22b, CZY22, CG20, CCGT22, CFO22, DH20, DTM21, EBVB21, FCS22, HOGO23, HCH20, HH21c, HW24b, HC24, HTOC20, JL24a, KK22, KCKK24, LHES21, LMZ22, LLYM21, LHHW21, LHHW23, LL20b, LD21, MK24, Mas21, MZY23, MG22, Ngo22b, Qiu20, QWY24, RG23, SCK22, SLZD20, SLK24, SSL24, SZ23, bWLgZbX24, XLKL23, YH22, YY21, Yu22, ZD21, ZH23, ZBH24, ZSH20, ZSL22, ZY20, ZY24, ZVS21, dC20a]. Stabilization [GS21b, BCH20, GANR22, GRVZ23, HH20, MWBM22]. stabilizations [Sha24]. Stabilized [JF22, PL24b, RYH24, TSA23, AAM20, ACPV24, Axe20, CUN24, CK23, JKW23, KLS21, LJCL20, MTIR22, MBZ23, MF24, SW23, YZ24d]. Stabilizer [KD22, MGD22, ATHW21, LLGC24a, YZ20b, YZ21c, YZ21b]. Stabilizer-free [MGD22, LLGC24a, YZ20b]. stabilizing [Hu22]. Stable [BC22a, DKSQ22, ZZL20, BJB22, CD22, DC20b, DHQZ24, Fab22, GHM21, HYY21, HCC24, HWRM24, HZL24, IM24, JZW22, KCRVA22, LLT^{+}21, LML21, LLX^{+}22a, LJ23a, LL24c, LY21, LY22b, LV20, MDSRFO24, MARH^{+}23, QQSZ24, SUW23, SL20, THS24, WY20, WWJ22, WYJ23, WC24a, YKBH23, YLJW24, YLK24, ZJGZ20, ZHY20, ZOW^{+}21, ZCF^{+}22, ZLS23, ZPW^{+}24, ZWKS24]. Stackelberg [XZP22]. stage [CLPZ21, FRR24, GJM^{+}23, YBAE23]. stage-restart [FRR24]. stages [BYE^{+}23]. staggered [CK22, CK24a, Hur22]. standard [gK23, SLC22, Yam23]. standby [DNS21]. star [FLMP24, GGPP21]. star-dagger [FLMP24]. star-discrepancy [GGPP21]. state [AGNGG^{+}23, AK21, BLBH21, BZHB23, Bre22, BM24a, CLHL20, CLH22, DG22a, DAZ23, EBVB21, EGG22, FJ21, GSK20, GDZN23, GDG^{+}21, HMWY23, HOOT20, Kar24, KCKK24, Kon22, KTK21, LTZ24, LNP19, DFMK20, SA24, SS20a, SLP20, WSZ20, WLB^{+}21, YSFS22, ZW24]. state-constrained [Kar24]. state-dependent [KCKK24, SA24, WSZ20]. state-of-charge [AGNGG^{+}23]. state-of-the-art [GDG^{+}21]. state-space [EBVB21, GDZN23]. state-time [DFMK20]. states [GZ24]. static [BB21a, CHLR22, DHZ24, FFRS23, LYM22, NT23, ZR24a]. station [RLGGAV21]. Stationary [BS24a, FMNT24, BDR23, CG23, CV24, DZ22, DW20, DDKD19, GG24, HHD23, KLM21, RG21, WLW23, XH20b, YKY24]. stations [RLGGAV21]. statistic [SAZW24]. Statistical [Ais23, BLBH21, CMdMRFG22, LTL^{+}24, QG20, Zhu20b, KPW21, LLGZ20, SL22, SAZW24, WMK22, XY22, Xia22, ZS20a, ZTY^{+}24]. statistics [AAB24, ARMP20, Bay22, CMP21, DG20, FMNT24, MKS21]. steady [AK21, BM24a, DG22a, FZ20, FJ21, GSK20, Sha21, YSFS22, ZR24b]. steady-state [AK21, BM24a, FJ21, GSK20, YSFS22]. steam [FHH21]. steepest [MG23]. Stefan [KS23a, NS24c]. Steffensen [LP20, OCS24]. Stein [LLH24]. Stein-Stein [LLH24]. Steklov [ABBB23, BT23, CDD20, Dom21]. stencil [DOT23]. step [ASML23, AAB^{+}20, BA20, BJP24, BK21a, DH24a, FHY^{+}21, GGRCGG^{+}23, GLHJ21, HJ23, Hur22, IAH24, ISD23, KSPC22, KHAW21, KAD22b, KFK^{+}24, LYK23, LHHW21, LZZZ21, LLL24b, LSZ22, LDQF23, LST^{+}24, MV22, MCY23, Moi24, MF24, RR22a, RDH21, SLW21, SRBK23, SDC20, SZZ21, SDLZ23, SSRK23, WMH22, WYJ23, WCK^{+}22, YPLM22, YZ21d, ZHWD20, ZJY22, ZCCK24, ZM21b, MGV23]. step-size [SSRK23, WMH22]. step-stress [BJP24]. step-up-down [LLL24b]. stepping [ANR20, BJF24, LLZC21, LN23, QHPL21, RlLZZ21, WSY20, YZM23, ZOW^{+}21]. steps [HZZ24, TDAT23, Yam23]. Stepsize [EN20, RR23a]. stepsizes [TDL^{+}22]. Stiefel [SA23, WDPY24]. Stieltjes [ZH23]. Stiff [SS24b, Fal21, LZL21, LY23b, Mas21, ON22]. Stiff-cut [SS24b]. stiffness [BPDC20, SIS^{+}22]. stirring [VNA22]. stitched [FFRS23]. Stochastic [BAB20, BABR20, BM20, CPS20, KAMW21, LWZ21, ZYZ23, AGS24, Ant20, AS21b, ABO24, Ask21, BTBR21, BJB22, BVDB20, BB21a, BBTW24, CMW24, CKZ21, CLL21, CGW21, CLYS24, CNQRR22, CMRS21, DD23, DFFM21, DHKV20, DG22b, DKM^{+}20, EPL21, ÉM21b, FJ24, FLZ21, FPR^{+}24, GR22, GSZ22, GGHY24, GL24, GMCJGM24, GDZN23, GJW23, GLHJ21, HLW22, HRT20, HCH^{+}23, HQT20, HH21c, HWD^{+}24, IYA^{+}23, JQ22, JL24a, JCLC20, KCKK24, KF24, KYP21, KJO23, KW23, KK20c, KK21, LY20, LPY20, LHHW21, LHW22, Li23, LHHW23, LVW24, LLW22, LG22, LDQF23, LL21, LH23, MYC22, MSW23, MST22, MWW21, MA19, MA21a, Mit20, NZA24, PS24c, RAN20, RB23a, RJ22, RHA20, RWW22, SM20, SPMB21, SSL24, SW23, SYG19, TM24, TS20, TY23, WZC^{+}22a, WC24b, WZLT23, Xia22, XDP24, YWKM19, YZ21a, YH22, YPLM22, YL22, YZ24a]. stochastic [YHF22, YZ24c, YZ21d, ZLL20a, ZLG20, ZGJ21, ZSL22, ZSL23, ZHZ24, ZGL20, ZVS21, YL22]. stock [PCWY20]. stocking [MBKK24]. Stokes [BS20, CHLY21, GLHL24, MR25, ZZW^{+}24b, ACPV24, BJF24, CC24, CRW20, CR23, CJL^{+}24a, CZ23b, CK23, CHP21, DG22b, DN20a, DN20b, FZ20, GANR22, GRVZ23, HH21a, HJM24, Hes20, HLMY24, JF22, JZZH22, JQ22, JHZY20, JZ21, LSC21, LRV23, LZF20, LLZC21, LLZY22, LWA^{+}22, LS23, LLZ24a, Li24b, LWT24, LHMT20, LR24a, LR24b, MR23, MDKG20, NMST21, Ngo22b, Ngo24, OHK23, PS22a, PK24, QHPL21, RR23b, Sha21, Sha24, SGY22, WLHJ20, WZZ20, WSHW22, XW21, XH20b, YY24a, YZL24, YYC24, ZZW^{+}24a, ZZW^{+}24c, ZY20, ZMFL20, ZOK21, ZLC24, ZR24b]. Stokes/Darcy [JF22, LLZC21, LWA^{+}22, QHPL21, XH20b]. stop [GL20, MSK21]. stop-loss [GL20, MSK21]. stopping [ANR20, DMP23]. storage [AGNGG^{+}23, HR21]. Störmer [YCXW24]. Störmer-Cowell [YCXW24]. strain [BFQ24, LYM22]. strain-limiting [LYM22].

Strang [WZD^{+}24]. strategies [AGGAV^{+}21, AS23, HL22a, IBW^{+}20, LLL22a, MULS23, WLZY21, YP21]. strategy [BFF22, Dad20, DV21a, ER20b, GJM^{+}23, HCH^{+}23, JH24, hKhWjH21, KKJ22, LZD21, LCZ19, MQZ^{+}24, NAS22, PVV24, PZL23, TLY23, WRZW21, WDS22, WZ23c, WPX23, WLC23, YKAA21, YCX20, YY21, YYZB24, ZLL20a, ZSMG20, ZCJL20]. streaks [ZLZ^{+}24]. strength [BLBH21, BZHB23, MS24a, NN22, WDTW20, Zhu22, Zhu23, Zhu24]. strengths [BZHB23]. Stress [NN22, ANR20, BLBH21, BZHB23, BJP24, ER21, GGVRB22, JO24b, KK24b, LRV23, Let24, MS24a, WDTW20, Zhu22, Zhu23, Zhu24]. stress-assisted [GGVRB22]. Stress-strength [NN22, BLBH21, BZHB23, MS24a, WDTW20, Zhu22, Zhu23, Zhu24]. stresses [BZHB23, PV22]. Strictly [KH21, APS22, CVPM22, DL20a, DH24a, LZ24b]. string [KPW21]. strip [ZX23a]. Strong [CG23, EL22, HN22b, LGC23, Liu23, YYY21, YZ21d, ZSL23, Ais23, BIJ21, CLL21, CGW21, DC24, DJ22b, FCS22, HTZ21, KAMW21, LZLL20, LG22, NEYZ22, RB23a, SRBK23, Wan21a, Wen21a]. Strongly [HSM20, YZ21a, AORS22, BM21, TV22, WC24b, WSHW22, Zhu20a]. Structural [PLZ21, ZTS21, GPP^{+}20, HTOC20]. structurally [VVB21]. Structure [LG23, ZAC^{+}23, ABFSC24, AESW22, BHPV23, DL23b, DDKD19, ELS21a, LSY24, MCW24, NY24a, SFI23, Svá23, Tor21, VS23, WWS^{+}21, WLDZ24, Wu24, YT24]. structure-preserving [DL23b, ELS21a, WLDZ24]. Structured [MDY20, BFRV21, HH24c, LM24, LJ23b, LC20, LWXL21, RK22, WFZ22, YKAA21, YHC20]. structures [DZ22, MMM20, Men24, OKW23, PH22, Pet20, PS22b, RCN21]. studies [HJO20, RAN20]. Study [CD20b, SBC20, AN24, AMT23, BK24, BYE^{+}23, CK22, CK24a, CLST20, CNQRR22, FPCV^{+}22, FDDM24, GNGGL^{+}19, HRT20, HP22, HWAT22, Jas24, LCD21, Moi22, PZY^{+}24, PKS23, RCVA24, SESH23, SB24a, Yam23, Yam21b, BGG^{+}21, HCCSCGAM24]. studying [AA23, FFT20]. Sturm [IuR23, Mag20]. style [LH21]. SU2 [MBG21a, ZAC^{+}23]. sub [CD23, LLFT20, RFGGM20, WOV22]. sub-diffusion [LLFT20, WOV22]. sub-grid [RFGGM20]. sub-sampling [CD23]. subcomponents [Ozk21]. subdifferential [TTK23]. subdiffusion [HZMD22, Li24a, QWY24, RlLZZ21, TL20, TL22]. subdivision [AI22a, BEL23a, CJL^{+}24b, GS21d, Han24, JYY21, NBLZ24, YKY24]. subgradient [LH23, TDL^{+}22]. Subgrid [CK23]. subject [EKCF22, MG22]. subjected [JO24b]. sublinear [LMZ22, PT24]. subsampling [NYL23]. Subsidence [VTBM21]. Subspace [TGS21, BEJ20, BK20, ERB^{+}22, FMNT24, HJ20, HN22a, Mia20, MW21, SL24a, ZFwC20]. subspaces [IS23, Jia20, JL21, SK24b]. substructures [Men24]. substructuring [GBM21]. Successive [MP24, GW24]. sudden [WAAM20]. Sufficient [YJZ23, Juk20, LDC20, Sun20]. suggestion [Yam23]. Sugiura [PI24]. suitable [SFI23, YL22, ZFH20]. sum [CLYS24, CFL21, HKM24, JHZD20, YSFdRdP20, ZGL20]. summation [CW22, ERB^{+}22, LGG^{+}22]. sums [KK20a, LLY21, TE20, ZTY^{+}24]. sup [Zho22]. sup-norm [Zho22]. Super [CJP20, GS21d, HH24b, SLW21, WZZT20, CG20, RS22b]. Super-convergent [SLW21]. Super-critical [WZZT20]. Super-exponential [CJP20]. super-linear [RS22b]. super-linearly [CG20]. Super-resolution [HH24b]. Super-smooth [GS21d]. SuperAdjoint [HH24b]. Supercloseness [LYZ24]. supercomputers [CKT21]. Superconsistent [DO21]. Superconvergence [DWZ20, SZ24, Tia22, WZH^{+}21, ZC24, KAMW21]. Superconvergent [CAN23, CMZ22, YZM23, YZ24b]. Superlinear [LN21, CMW24, DFFM21]. supersonic [YCG^{+}23, ZGG23]. supersymmetric [CZ23a]. supervision [Fan22a]. SUPG [CUN24]. SUPG-based [CUN24]. supply [CQBL20, GXL22, San21, WQ20]. Support [FAUU20, GANU21, HWG21, YYT24]. support-denoiser-driven [HWG21]. supported [HAD24, SUW23]. supporting [SS20c]. suppressed [HH24a]. surface [CL24a, CLM22b, CFO22, DKM23a, HBPV21, MJ24, MJK20]. surface-higher [MJK20]. surface-relief [CLM22b]. surfaces [AQ20, AM22a, AM23, BAIB24, Cho24a, HF24, JKJ^{+}23, LZW^{+}22, LLX^{+}22a, LHG^{+}24, WZ20b, XLKL23, YLK24]. surfactant [YLJW24]. surplus [SZ20]. surrender [JK21a, ZCZ23, ZSZ24]. surveys [CMdMRFG22, ZSY24]. survival [BLBH21, BZHB23, CCCMH24, VBRT24]. suspension [CJCDLL22]. sustained [SH21]. SV [hKhWjH21]. SVD [YS24a, KTK21, WLC23]. SVD-based [KTK21]. SVEEs [DJ22b]. SVIR [OEH23]. SVM [LHS^{+}24, SL24b]. SVR [GNGGL^{+}19]. SVR-based [GNGGL^{+}19]. swap [ABO24]. Swapping [CMVV20]. swaps [CKZ21]. swarm [GS20a, GXL22, SZL20, Zha20a]. sweep [LDY22, LWL22, LF21, VOJS24]. Swift [CVPM22, Lee20, LYK23, LY22b, LC23, WC24a]. swing [AHB21]. switch [LLL24b]. switched [YY21]. switching [AHB20, AHB21, AZ22, AZZ23, DLS20, GKV21, HSYZ24, KK20c, KK21, LL24a, MW20, Mit20, MCW24, PZL23, SL23, WSZ20, YW23, ZCZ23]. Sylvester [CHSZ24, GBTS22, HSMT20, NKBJ20, YHQW21]. Sylvester-type [YHQW21]. Symbolic [SB24b, WX21]. symbolic-numerical [WX21]. Symmetric [LL24e, AM22b, BEJR22, BLRZ23, BA23, CCETZ21, CUUdB24, FHY^{+}21, GH20, GLLZ21, GW24, HOOT20, Hua21, JW23, JNCY23, Liu20, LZWL23, LM21b, MTP21, MW21, NT20, Ri22, Ri24, SUW23, SHZ23, Ver20, WCWZ20, WL22, WSLW23, XWZ24, ZTZZ20, ZKT20, ZZ20a, Zho20, ZM21b]. symmetrical [GMK20]. Symmetries [BLV22, AQ20, GÇSA23, MGA24]. symmetry [AHGG20, AM22a, BBC22, CW21a]. symmetry-breaking [CW21a]. Symplectic [LF21, Ant20, BCCW22, JLZ23, YZ23]. synchronization [AUAA20, CPRV22]. Synchronous [FS24b, KKO23, XDX24]. Synge [CHZ21]. synthesis [XD23]. synthetic [PMVB22, SKZ23]. System [GDG^{+}21, AA21a, AVRLRMGG23, AGGAV^{+}21, AGGAVGG23, BCLS24, BLBH21, BHRW22, BZHB23, BS20, BGG^{+}21, BMV23, BB20a, Bog21, BS24a, CRW20, CJL^{+}22, CCCMH24, CJL^{+}24a, CL24b, CKL20, CSS24, CST19, DFH21, DdO22, DYM21, DWWZ22, DHQZ24, DS21, EO24, EK24a, EKCF22, FM21a, FGM21, FJ24, FC22, GJ19, GM24, Gök23, HH24a, HCCW20, Jas22, Jas24, JGLH20, KT21, KK24a, LLZC21, LLX^{+}22a, LWA^{+}22, LW22, LGY24, LWW^{+}22b, LY24b, LH24, MDMS22, ML23, NZ21, Ozk21, PNS21, RS24a, RKA20, SES20, SuR20, SRMD22, SMDVA22, SVK20, SACQ21, SS20c, Sun20, Sur21, TSG22, TE24, TGLa21, WWJ22, WCLY23, WCMW23, WWW^{+}24, WZ24a, WTX24, WZZT20, XYAH23, YSN24, YZ20a, YBM23, ZC20a, Zha21c, Zhu22]. systematic [HQL21]. systems [AAP24, ASB22, AS21a, ADMV22, ABM^{+}22, AZD21, Ant20, Asa24, ARK20, ÁGKM21, ALKH21, BTBR21, BA20, BLZL22, BBTZ20, BABR20, BB21a, BA22, BBMK22, BFVY23, BP22a, BL20, BCGAR22, Bha21, BG23, BLyS20, Bre23, Cal23, Cao20, CWW23b, CLMO23, CXS20, CDW20, CLS21, COT21, CJ23, CFdD23, CUUdB24, CNQRR22, DTM20, DTM21, DFH21, DNS21, DJ24, DXM^{+}20, DRD^{+}23, DAZ23, DSCZ24, DKM^{+}20, ESS21, ELTD23, ELS21a, EP20, EPL21, FBJ20, FM22, FGP24, GR22, GZ24, GHC^{+}20, HA21a, HOGO23, Hon22, HH20, Hu22, Hua21, Hun21, Hun22, IM24, JG21b, JZY22, JM21, JVAD22, KA24, gK21a, gK23, KT24, KV20, KS22c, LSC21, LJ23b, LZ21, LZZZ21, LWT24, LCGW24, LTL^{+}24, LNP19, LHE24, LS24, MULS23, Mar19, MM21, MDY20, MW22, MDLY23, Mim20, MORO20, Miy22, Moi22, NNJ24]. systems [NT20, ON22, OKE24, PRW^{+}22, PSWZ21, PN21b, PS24c, RVS^{+}22, SV23, SPMB21, SLZD20, SYS24, SW20, SHZ23, ŚK23b, TK24a, Tak24, TDD23, TS20, Tor21, VSP23, WCW20, WW22, WLB^{+}21, WA21, XY24, XWZ24, XDP24, XD23, XDX24, YC23, YZZ20b, YY21, ZYZ22, ZZ20a, ZZ22b, ZWZ23b, ZH24b, ZM21b, dA23]. Szego [XTL23].

T [HC24, KP20, Khl21, TKM21, WMF20, YY21]. T-lymphocyte [TKM21]. T-meshes [HC24]. T-S [YY21]. T-shape [KP20, Khl21]. table [Har21]. tackle [YSBT20]. tail [CW22, HL22a, ZY22, LY24a]. tailed [NK22, ZZWS22]. tailed-related [ZZWS22]. tails [CX24, ZX23a]. Taiwan [CC21b]. Tamed [DFFM21, EL22, GGHY24, KK20c, ZSL23]. Tanabe [gK21a, gK23]. tangent [BAB20, YJZ23, ZJBY23]. target [WRZW21]. task [MGV23, MFH24]. task-driven [MFH24]. tau [NM24]. Tausworthe [Har21]. taut [KPW21]. taxis [JMX24]. Taylor [Bün20, CAJ24, Dum24, EN20, LBBB23, LG22, LN23, MGG^{+}24, ZW21]. Taylor-like [CAJ24]. TCP [HBLZ21]. tech [LGMC20, LLG^{+}22]. technical [FRT23]. technique [ASADB23, Ais23, AID24b, CFO22, DWWZ20, DHLT24, Els21b, EWCQ23, FNH20, GNGGL^{+}19, HN22a, HMA24, JG21a, JH24, KS22b, NA21, OWV24, Pat23, RR22c, Rou22, SM22a, SZL22, SSL22, SSVS20, Vab24, VTBM21, VMP23, WYW^{+}24, YC20b, YW23, ZYZ22, ZZwS21, Zha24, ZL23b]. Techniques [NW23, AGNGG^{+}23, BPRZ21, BHJB21, CJL^{+}24b, DS21, GMW24, JNCY23, MHW23, NNCH21, PKPD21, PLS22, Ray23, SNMK21, SS24a, VNV23]. technology [CZ20, DLL21, Fan22a, LGMC20, LLG^{+}22, UV24]. technology-induced [UV24]. telegraph [OWV24]. Temam [RVX20]. temperature [CLL20a, COT21, FdOPS21, IKL24, WCL22]. tempered [BuR24, FU19, Liu24, PAAI24, SM23]. Temporal [AZL21, BYWA22, BEH^{+}23, DKM23a, FGN^{+}23, HCCSCGAM24, HN22b, RSRW24, SZ23, WL21]. temporal-spatial [HN22b]. temporally [WWW^{+}24]. tends [BHJB21]. tensile [LYM22]. tension [FGPP23]. Tensor [EJHR24, RMM22, ACT22, Bal24, CZ24, Guo24, HSMT20, gHtSjT24, KMZ21, Kra22, LWD20, LYB22, LZZZ21, LDC20, LM21b, MHW23, MDW23, MQZ^{+}24, Miy22, NKBJ20, RU22, tSjTsJ24, SN21b, SWN23, SD24, TT20, WCW20, WLCH23, WCMW23, WLDZ24, ZWH^{+}20, ZDY20, ZH24b]. Tensor-Krylov [RMM22]. tensor-train [LZZZ21]. tensorial [BBBK24]. tensorizations [UDI20]. tensors [BAD^{+}22, CW22, CZ23a, CWY21, HLW20, HBLZ21, JNCY23, KK24b, LZLL20, MQZ^{+}24, Miy22, MSS^{+}21, SIS^{+}22, WCWZ20]. term [AF23, AM22b, BB21b, CTN22, DL20a, DL20b, FSC24, HSC21, JW23, JYYL23, MP23, QX24, RS24b, RGA20, RZ20, Sha23, WC21, WZC22b, Wu24, XLLA21, YZ24d, ZGZ23, ZF24]. terminal [Gu21, TKVA21]. terminal-boundary [TKVA21]. terms [BD23, BGG^{+}21, gHtSjT24, SMDVA22, SQT24, TY23, WW23, YT24]. ternary [DWWZ22, WZY23, WZD^{+}24]. terrestrial [FMTB21]. test [ARMP20, BJP24, BC24, CMP21, CMJN22, EBVB21, MPGE24, Yam21a, ZG20b, ZZQ21]. testing [AHK24, BCMP23, ER20a, ER21, JS22b, LLGZ20, MPGE24, SAZW24, WHC^{+}20]. tests [CBF24, HM22, LLGZ20]. tetrahedra [PP22]. texture [WW24a]. TFC [WT22]. TFC-based [WT22]. th [Uga24, WLZY21]. their [AM21, BD23, BEL22, BEL23a, BL20, Cho24b, CST19, DRR22, DL23b, DFK20, GM23, Han24, Hun22, HBMA22, JK21b, LL24e, LK21, MPR24, RTT24, Rod21, SL22, SS20b, TGLa21, UPPZ24, WK20a, WLBG22, YS24b, ZG24, ZL23a]. theorem [CAJ24, CC20, DK24, FB24, Hu22, LZW20, NB23, PH24, Psa24, SuR20, TSG22]. theorems [CFB24a, CFB24b, JZ24, LZ20c, SY23]. Theoretical [BDR23, HRT20, HZMD22, RR23b, SRMD22, SB24a, ZLG20]. Theory [BSZ21, GAA19, GAA21, KPW21, MM22a, dS24, AM21, AUAA20, BSH24, BFQ22, CSZP20, CL24b, DRB22, HH23, gK23, LZW20, Li20b, LS22b, MJK20, Oga20, OGR22, Ros23, TL21, VAMO24, YZ23]. thermal [BFQ24, CL24a, DZ22, GDG^{+}21, SEAS22, SIS^{+}22, SH24b, XLKL23]. thermo [DHZ24, FMTB21, LQS23, ZR24a]. thermo-mechanical [FMTB21]. thermo-poroelasticity [DHZ24, ZR24a]. thermo-porous-elasticity [LQS23]. Thermodynamically [BS20, SW21a]. thermodynamics [HQT20]. thermoelastic [BFQ21a, BFQ21b, BFQ22, BFLQ24, FFRS23]. thermoporoelasticity [AVC22]. thermoviscoelastic [JLM24]. thermoviscoelasticity [MQ20]. these [RLGGAV21]. theta [LHHW21, LHHW23, LDQF23, ZSL22].

Thick [ACR24]. Thick-restarted [ACR24]. thin [FS24a, GJNR21, GeOJD24, JO24b, LSE24, Let24, SVWC22, WCPW20, ZUZ20]. thinning [CLW^{+}22, MCW24, ZZK21]. thinning-dependence [MCW24]. third [AID24a, AID24b, EHVM22, JKKK20, LYL23, MH21, MQZ^{+}24, RR22a, RR23a, SWN23]. third-kind [MH21]. third-order [EHVM22, LYL23, MQZ^{+}24, RR22a, RR23a, SWN23]. thirds [GN24]. Thirring [KSZ21]. Thomas [PHDD22, SS24a]. Thompson [BFQ21a, BFQ22, BFLQ24]. threads [FFRS23]. threat [AS23]. Three [AA22, HHD23, AF23, AA24b, AM22b, ABO24, AGGAV^{+}21, ALKH21, BB23, BWS20, BWS21, CGO24, Els21b, GS21d, JYYL23, Kaz21, KK24a, LLL^{+}22b, OK23, OHK23, PSWZ21, SRBK23, SŁK^{+}22, WZ20a, WL22, WZZT20, ZSCH21, ZZwS21]. three-block [ZSCH21]. three-by-three [ALKH21, WL22]. three-component [WZZT20]. three-dimensional [Kaz21, PSWZ21, SŁK^{+}22]. three-directional [GS21d]. three-factor [ABO24]. three-field [CGO24]. three-level [Els21b, ZZwS21]. three-player [OHK23]. three-species [KK24a]. three-step [SRBK23]. three-term [AF23, AM22b, JYYL23]. threshold [AAMR24, HX22, LCZ19, MULS23]. thresholding [SZZM23, dLNO21]. tight [BBSV21]. Tikhonov [BPR20b, FZQ22, HLY22, ME20, QX24, RU22, SL22, WLW23, YPL20, Zha21a]. Tikhonov-type [RU22]. tilings [BCDF22]. timber [SKS22]. Time [Fra21, HL21b, KH22, LWC20, MSK21, PZL23, YCX20, ZGL20, ADZ20, ANR20, ACDV22, AH22, AAHK24, AGS24, AUAA20, AA24b, AGGAV^{+}21, ACG20, BT21, BHRW22, BBTZ20, BYE20, BB21b, BVBO23, BJF24, BDPR23, BK20, CGSVBB^{+}24, CZY22, CZHQ23, CL24a, CTGV20, CMFMSQ22, CLH22, CW24, CJLL21, CECCCR23, DVRV22, DRD^{+}23, EGG23, EPL21, EKCF22, FFS24, FY22, FZQ22, FS24b, FC23, FDDM24, FK23, Fro24, FC22, GLYS21, GBTS22, GMK20, HH21a, HJM24, HYW20, HLZ20, HLLY24, HZMD22, HR23, HPW22, HZZ24, HS20, HPWM20, HSC21, HLYZ24, HTIA24, JS21, JDBT22, JZ22, JF22, JMX24, JLZ22, JLW23, JZM^{+}23, Kaz23, KKJ22, LYK23, LL24a, LPY20, LLZC21, LLZY22, LS23, LJ23b, LVW24, Li24a, LW24, LCW20, Liu24, LST^{+}24, LDM23, LRQV24, LNP19, LXF21, LM21a, LN23, LR24a, MYC22, MGD22, MD22, MMM20, MSD21, Mar19]. time [Mas21, MZY23, MFH24, MDD21, MB24, MLS22, NMST21, NEYZ22, NP22, NZA24, NYL23, Ngo22b, NS24c, Ois20, OWV24, DFMK20, QS23, QHPL21, QWY24, RHM22, RlLZZ21, Ret24, RSRW24, RG23, Rou24, RZ20, Ruj23, SM23, SM22a, SUW23, SN23, Sha23, SLZD20, SXX21, SC23a, SS23c, SKVA23, SZ23, SLW^{+}22, Sur21, TK24a, TDD23, Tam20, TVB21, TTK23, TGS21, TBT^{+}20, TVAS23, VBRT24, Wan20b, WMF20, WXS21, WYH22, WCL22, WW22, WOV22, WYJ23, WCK^{+}22, WSY20, WY21a, WX22, WLW23, WZC22b, WZLT23, XWXZ22, XHLD21, XZL20, XW21, XZ22, XXL21, XLLA21, XDX24, YX20, bYW23, YVXX24, YPL20, YLJW24, YMC22, YZM23, YY21, YW23, YLLA23, YLH23, YY24b, YZ24d, ZVT^{+}23, ZHY20, ZFwC20, ZOW^{+}21, ZZ21, ZWY22, ZLJT22, ZGZ23, ZMYL23, ZF24]. time-accuracy [CZY22]. time-accurate [LLZY22]. time-average [WCL22]. time-between-event [MLS22]. time-changed [KKJ22]. Time-consistent [KH22, PZL23, YCX20, ZGL20]. time-delayed [WMF20, YW23]. time-dependent [CL24a, CTGV20, FC23, FC22, GLYS21, HH21a, HJM24, JZ22, JF22, JLZ22, LDM23, LRQV24, LR24a, MFH24, QS23, RZ20, WX22, XW21, bYW23, YY24b, ZMYL23]. time-domain [DVRV22, HLLY24, HLYZ24]. time-fractional [AH22, AAHK24, AGS24, BB21b, EGG23, FK23, HZMD22, HS20, HSC21, HTIA24, Kaz23, Liu24, MSD21, NZA24, RHM22, RG23, Rou24, RZ20, SC23a, SKVA23, SZ23, WYJ23, WY21a, WX22, WLW23, XXL21, bYW23, YPL20, YMC22, YZ24d, ZVT^{+}23, ZWY22, ZGZ23]. time-harmonic [FFS24, HYW20, MMM20, WYH22, XZL20]. time-homogeneous [Wan20b]. time-implicit [YVXX24]. time-integral [NP22]. time-marching [YX20, YLJW24]. time-space [ADZ20, FZQ22, YZ24d, ZF24]. time-stepping [ANR20, LLZC21, LN23, QHPL21, RlLZZ21, WSY20, ZOW^{+}21]. time-steps [HZZ24]. time-to-event [TVB21]. time-to-failure [EKCF22]. time-variant [JZM^{+}23, XHLD21]. Time-varying [Fra21, AUAA20, GBTS22, Ruj23, SLZD20, SLW^{+}22, Tam20, XHLD21, XDX24, ZZ21]. times [GMSR18, GMSR21, GHF24, LN22, NMB22, PA21, YW23]. tissue [PV22]. tissues [KA21]. TLS [HW22]. Toeplitz [Hon22, JG21b, NT20, ZYZ22, dF20]. tolerant [SLW^{+}22]. tomographic [HHM22a]. tomography [ACFM21, ACT22, DBR24, JMA21, LCB20, Ma24a, SH24b, LW23b]. tool [FdOPS21, SCHS24, ZGG23]. toolbox [Bün20]. toolkit [MBG21a]. Tools [GVT21]. topological [Ben24, HW24a]. topology [LZH21, PS22a, ZSLG21]. Torii [PI24]. Torrey [XLLA21]. Total [KMD21, BBBK24, LRT20, LCB20, LZ24a, MPR24, SL22, WY24a, YZM^{+}20, ZWX24]. totally [BAPPF23]. Totik [GSZ22]. toxicity [BB20a]. trace [KZM21]. Traces [ACEIJH22]. track [RLGGAV21]. tract [ASX21]. trade [SS20c]. trading [HCCW20]. traditional [XCWY20]. traffic [CZM23, Hun22, MQZ^{+}24, SACQ21, TKT^{+}24, XWXX22, YL22]. train [LYB22, LZZZ21]. training [GJM^{+}23, HB22]. tranches [MCW24]. transaction [GS20b, LMYZ21, SS20c, YL21a, dFG21]. transcription [BDPR23]. transfer [AAA21, BS23, CLL20a, DZ22, FAVM19, FAVM20, SSV23, VSSV20]. transform [AI22b, Bel23b, CJLL21, GJ19, HCH^{+}23, KEKT21, LZ20a, MMJ21, Mil22, MA21b, RKS^{+}22, RFB22, Ust23, WGM21, WSX21, Yak21, ZBS22, dLNO21]. transform-based [CJLL21]. transformation [CLM22b, HLLY24, HMT23, KPD23, LP24, Naj20, QXZG22, ZUZ20]. transformations [CT20, CJL^{+}24b, HHL20, RHM22, SPKS^{+}24, TFK23, XZZ^{+}20]. transformed [ÉM21b, LGC23, MDLY23]. Transforms [MCMR23, CYC21, EGG23, KK22, KWZX23, KZ23a, KZ24, KGV23, KS22a, ORT22, Oga20, SQT24, WK20a, WZ24b, ZuIKA22]. Transient [LP23a, CK23, FHB20, YCH^{+}21, ZCSL21]. transition [Asa24, CGSVBB^{+}24, FSR24, HF21, IIS^{+}21, Juh21, LL24d]. transitions [VP24]. Transitory [ÖS20]. translation [AM22a]. Transmission [Evi23, AA23, ÉM21b, KZM21, RL24, SMKV21, vtW24]. transmuted [Hri22]. transparent [EOR21]. Transport [PMVB22, AFLP22, AF20, CTGV20, DRB22, DYM21, DW23, EO24, FdOPS21, GIV20, HET24, LWW22a, LWW23a, LGLC20, LR24a, NMST21, OMRdQ24, PWL23, RB23b, WDPY24, Wen21b, YN21, ZR24a]. Transportation [KMTZ21]. transpose [CCZ23]. transpose-free [CCZ23]. trapezoidal [BLN21, ONT21]. Traveling [AL20, LY24b, SC21, SCK22, WZZT20]. Travelling [BGdlR24]. treatment [ESA23, HM21b, HVMS22, LLS22, OEH23, PP24, Sal21a, SKE20, ZWKS24]. Tree [PCT23, JGW^{+}23, JGZ^{+}24, TBA22, YC20a, YZ23, ZX23b]. trees [AVRLRMGG23, BP22b, Wan20b]. Trefftz [YY24b]. trend [XCWY20]. tri [CL24b, YZ23]. tri-coloured [YZ23]. tri-objective [CL24b]. trials [Kon22]. triangle [AAP24, GJ23]. triangle-based [AAP24]. triangles [LJ23a, LL24e]. triangular [BWS21, CLU23, DOP23, DDNZ21, LZF20, LWT24, LHL19, SH23, WZC^{+}22a, WWWY23, YZ24b, ZZMK23, ZWJ22, ZW22]. triangulation [BEIR23]. triangulations [GS21d, Kop20]. triblock [WZY23]. tridiagonal [AM22b, KT21, PHDD22, dF20]. trigonometric [BE21, BLV22, LCD21, WY21b, ZFW22]. trigonometrically [FHY^{+}21]. trigonometrically-fitted [FHY^{+}21]. triharmonic [WZ20b]. trinomial [Wan20b]. triopoly [LZL^{+}23]. triple [ACK21, MQZ^{+}24]. tripling [EWCQ23]. triptych [OK23]. Triquintic [BB23]. trirectangular [PP22]. Trotter [Lem24]. Truncated [EM21a, LLW22, Pou21, YHW22, ABA19, DDRS21, JL24a, LGC23, MWW21, PT20, TM24, VCNR20, YWKM19, YH22, YPLM22, ZWX24]. truncation [LY20, LMS21, SPKS^{+}24, SV23]. trust [BA20, YHW22, YYT24]. trust-region [BA20]. Tsallis [TA22, dS24]. Tseng [TLD^{+}23]. TT/QTT [BD20]. tubal [ACT22, JHZD20, ZWH^{+}20]. tumor [MDD21, SŁK^{+}22, SCZ20, VORD20]. Tuning [SŁK^{+}22, AVLMRMGG23, CWT23, CDL24, CDLR24, NGNB23, SNMK21, TDD23]. tuples [gHtSjT24]. turbine [Kan24]. turbulent [BO22, FFT20, HF21]. Turing [ML23]. Turkey [Uça23]. TV [BEJ20, DDF23, LL23]. TV-based [DDF23]. TVP [HQL21]. TVP-VAR [HQL21]. twice [MS20c]. Two [FHY^{+}21, IAH24, ISD23, KS23a, LLLD20, LQS23, MDMS22, PH22, RB19, Ray23, TDL^{+}22, WYJ23, YQZS21, YL22, ZKT20, ZXC22, ADZ20, AS20a, ALK20, AHT21, AJR22, AGMZ22a, AGMZ22b, AAA^{+}23, AAA20, BV22, BA20, BZH24, BB20a, Boe20, BitH21, CDCVV22, CHXL20, CS24, CTN22, CS21, Che21, CH20, CSS24, DKM23a, DWWZ20, DR23a, DDF24, DC24, DKMT24, DW23, DN20a, EFTW21, ETKF23, FLT20, GJ19, GHG24, GG24, GVT21, GK21b, GJM^{+}23, HYY21, Hes20, HM22, HC20b, HC21, HZS^{+}24, Hur22, IS23, IIS^{+}21, JA22, JWZ21, JLY^{+}21, KM22, KMUR24, KHAW21, KR21, LLW21, LBBB23, LR22, LSE24, LL20a, LWYZ21, LZH21, LDY22, LWL22, LHW22, LZZ24, LZ21, LZZZ21, LWT24, LLH24, LLLL20, LHH21, LHL21, LXZ23, MTP21, MH22, MS20a, MDS20, MH23, Moi24, MF24, Ngo22a, Ngo22b, Ngo24, NM24, OWV24, PSRVA24]. two [PH21, PLS22, QX24, RDHA20, RR22a, RlLZZ21, RH23, SM20, SH23, Sha21, Sha24, SZK20, SDC20, SLZD20, SDLZ23, SRAG23, SCZ20, TGLa21, WEEA20, WK20a, WCK20, WZJQ21, WYH22, WCLY23, WW23, WHZ24, Wen21b, WQ22, YT20, YLFT21, YMC22, YOCY22, YAE23, YLS20, YWL24, ZHWD20, ZZwS21, ZLJT22, ZLS23, ZZLY23, ZC24, ZJY22, ZEW23, ZMFL20, BM24b]. two-asset [Boe20, BitH21]. two-block [JLY^{+}21]. two-by-two [LZ21, LWT24]. two-component [HYY21, IIS^{+}21]. Two-dimensional [RB19, ADZ20, BV22, CS24, DWWZ20, DR23a, DDF24, DW23, EFTW21, ETKF23, FLT20, GG24, Hes20, HC20b, HC21, KM22, KMUR24, LBBB23, LR22, LSE24, LL20a, LWYZ21, LZZ24, LXZ23, MS20a, MH23, Ngo24, OWV24, PSRVA24, QX24, Ray23, SM20, SH23, SZK20, WEEA20, WYH22, Wen21b, ZZwS21, ZLJT22, ZC24]. two-factor [LLH24]. Two-frequency [FHY^{+}21]. Two-grid [ZXC22, HC20b, HC21, LZH21, OWV24]. two-layer [Hur22]. two-level [DKM23a, Ngo22a, Ngo22b, Sha21, Sha24, WCK20]. two-mode [JA22, ZLS23]. two-parameter [CSS24, LHW22, YWL24]. Two-parameters [ZKT20]. Two-phase [KS23a, AS20a, AJR22, CDCVV22, CS21, PH21, SRAG23, WCLY23, YOCY22, YLS20]. two-point [BZH24, CHXL20, CTN22, LLLL20, LHL21, SCZ20, BM24b]. Two-Scale [PH22, CH20, RH23]. two-sided [DKMT24, JWZ21, TGLa21, YLFT21, ZEW23]. two-species [BB20a]. two-stage [GJM^{+}23]. Two-step [IAH24, ISD23, BA20, KHAW21, LZZZ21, Moi24, MF24, RR22a, SDC20, SDLZ23, ZHWD20, ZJY22]. two-sweep [LDY22, LWL22]. two-variable [AHT21]. Type [PM20a, YWW24, AMS23, AVLMRMGG23, AAKA24, AL16, AF20, AK22b, BD24, BEIR23, BRLS23, BFQ21b, BLN21, BEJR22, BK22, Bic21, BC22b, BFF22, BM21, CMP21, CD20a, CZHQ23, CHY20, CYC21, CLU23, Cso20, Dar20, DGL24, DKM^{+}23b, DP24, DN20b, ER20a, EM21a, EEE21, EHVMM23, FL23, FC23, GSZ22, GST20, GGO21, GX24, GO24, GK21b, GHTC21, HWD^{+}24, Hun21, JMLX23, JHYZ24, gKZ21, KWZX23, gK23, Kes23, KSZ21, KK20c, KK21, Li20a, LZWL23, LY23a, LC24a, LZG24, LH20, LN21, MSM24, MQ20, NMMH23, OF20, PL24a, Pie20, Pou21, PI24, RG21, RU22, RS24c, SM22a, SKE20, SS24a, SS23b, SS23c, SLW^{+}22, SSVS20, TBA22, TBA24, Tom24, WLHJ20, WXZ21, WZY23, WAH22, Wen21a, WQ22, YSFS22, YL21b, YBM23, YAE23, dlCYS22, YHQW21, YLLA23, YWL24, ZVT^{+}23, ZSY24, ZZG20, ZCYP20, ZG20b, Zha21a, ZWZ23b]. type [ZHMD22, Zho20, dOdO23, GC19, LMSZ21, QG20, Tak20, CKY23, ER21, GHG24]. type-1 [BEIR23, Tak20]. type-2 [TBA22, TBA24]. type-I [ER20a, GC19]. type-II [SS23c, LMSZ21, QG20, ER21]. types [Jas22, Jas24, MMM23, TGLa21, ZZLY23].

UBB [YS24a]. UE [FM21c]. UE-splines [FM21c]. Uhlenbeck [HX22, HS23]. Ulm [WSLJ23]. Ulm-like [WSLJ23]. ultra [BT21]. ultra-weak [BT21]. ultracold [ZW24]. ultraconvergence [mHLZ20]. ultrashort [ZHWD20]. ultrashort-pulsed [ZHWD20]. ultraslow [LFL20]. ultrasound [ACFM21]. unbiased [AK22b]. unbounded [Mag20, ZX23a]. Uncertain [CLYS24, GM23, YC20a, DKM^{+}20, GJNM24, GZ24, LLK23, LZ20c, TBA22, XGYY23, YP21, dHSR22]. uncertainties [AB23, FYL^{+}24, FGM21, JCLC20]. uncertainty [CGCMRP24, CSZP20, CL24b, CY20, GJNM24, HH23, KGV23, KS22c, LJ23b, SH24b, WZJQ21, YYBL24]. Unconditional [LCZL24, Ngo22b, LLYM21, MZY23, XLKL23]. Unconditionally [LY21, WC24a, CZY22, CD22, HCC24, LML21, LLX^{+}22a, LL24c, LY22b, QQSZ24, YLJW24, ZHY20, ZOW^{+}21, ZCF^{+}22]. unconstrained [AF23, DAMA22, KAM^{+}23, LZD21, NP23, WY24b, WSL^{+}23, YYT24, YWL24]. uncontrollability [Mim20]. under-expanded [YCG^{+}23]. underground [SMM22]. underlying [BN21, KKKL22]. underpin [AZZ23]. understand [SLA^{+}24]. understanding [dPdMQL23]. undetected [Ngo22a]. unequal [New24]. unfitted [CO19, HS24b, MTZ23, SM22b]. Unfoldings [ZDY20]. unforced [WTP^{+}21]. unidirectional [EEE21, YZM^{+}20]. Unified [PW23, CDJ20, DNK23, FP21, SEAS22]. Uniform [GXC24, Zho22, ABFSC24, GST20, JYY21, KSVA22, LPP24, NK23, SSVS20, YLLA23]. uniform-Laplace [NK23]. Uniformly [Cal23, ZC20a, GHTC21, SMDVA22]. unilateral [ONT21, WC21]. Uniqueness [DDF23, LFL20, DPW24, GSK20, HM23, IKL23, ZM21a]. unit [AGS24, BCA20, BM24b, DNS21, For24, RR22b, TFR22, Ver20, Vie21]. unity [CDLR24]. Univariate [JLLL23, HA20a]. universal [GMCJGM24]. unknown [AA24c, TK24a, YS24a]. unknown-but-bounded [YS24a]. unknowns [PK24]. unmatched [HHM22a]. unmixing [LT24a, WHlZH23]. unobserved [AAB24]. unreliable [SA24]. unsaturated [SVC20]. unstable [BCH20, IM24, ZvS23]. unsteady [DWWZ20, MHM23, QHPL21, SRAG23]. unstructured [DBSS^{+}21, FLT20, PFG24, SRAG23]. Unsupervised [ZLZ^{+}24, WW24b]. untrained [MWLY24]. Unveiling [HMS22]. Update [KLZ^{+}24, LLLD20, LZD21, ZCCK24]. updating [GPP^{+}20, SZZ21, WLZY21]. upgrade [JA22]. upon [BB21a]. upper [BCH20, DDKD19, LZF20, SH23]. upper-lower [LZF20]. upright [GJGDPR23]. upscaled [PCMH20]. upscaling [GCV20, RB23b, YCCZ24, ZC22, ZC20b]. upwind [ATM21, NMMH23, WLM24, YCH24, YLS20, YLS23]. upwind-block-centered [YLS23]. upwind-mixed [YCH24]. Urysohn [BLN21]. Usage [KPD23]. use [ACFM21, Bha23, dPdMQL23]. used [DDT23]. users [FRT23]. Using [AHPS22, CC21b, HM22, MJ24, WCK^{+}22, ABBB23, AKN24, AC22, AA24b, AGS^{+}23, AGST23, AST^{+}23, ASTK24, ANM24, AGNGG^{+}23, AV21, BAB20, ACT22, BLBH21, BZHB23, BYWA22, Ben24, BGG^{+}21, BCGAR22, BK22, Bra23, BEH^{+}23, CQBL20, CUN24, CSLQ22, CLS20, CNYC23, CM24b, CL24b, CCM22, DMST21, DMN24, DXM^{+}20, ESA23, EGG23, FGPP23, GHG24, Gol24, GK21b, GS24, GJW23, GBA21, HHM22b, HH23, HLLL23, HO22, HBPV21, HVM22, HSM20, HWRM24, HC20b, HWAT22, IMT20, IMT23, JS22b, KLM21, KS24, Kra22, LBBB23, LSC21, Let24, LMZ22, LRT20, LMLB22, LP23b, LN21, MGG^{+}24, Ma24b, MMJ21, MÖ21, MLYZ21, MHM23, MI20, MQZ^{+}24, MA21b, ME20, MPGE24, MLS22, NRV21, NY24a, NAS22, NNL^{+}24, PK23, PL24b, PS22b, PN21b, PNS21, QJLQ20, RO20, RS23, RU22, RLGGAV21, RDH21, Ros23, RR23a]. using [Ruj23, RK23, RV20, SM23, SA24, SL21, Seg21, SNMK21, SK24b, SVK20, SZZ21, SWW23, SZA24, SSM23, SHZ23, SKZ23, Svá23, Tam20, TSL^{+}21, Uça23, Uga24, UPPZ24, UGG^{+}20, VTBM21, VTB21, WGM21, WCLY23, WSC23, WZ24b, WYW^{+}24, WLC23, YSUY24, Yam20, Yam21a, YQZS22, YZM^{+}20, ZMOW23, ZSY24, ZAC^{+}23, ZUZ20, Zha21b, ZSHD24, ZZWS22, ZL23b, ZH24c, ZM21b, dS24]. Utility [YL21a, CNS21, GKV21, HCH^{+}23, hKhWjH21, SL23, YWTL19, dFG21]. Utility-indifference [YL21a]. utilizing [HAD24]. Uzawa [CMS20a].

vacation [SVK20]. vaccinated [BYE^{+}23]. Vaccination [MP23, OEH23]. vaccination-age [OEH23]. Valid [NAA20, LZ22, LSZ23]. Validated [CHsL22]. Validation [CTGV20, BR20b, ZGZW23]. Valuation [AZZ23, GJNM24, Ruj23, SPS21, ZSZ24, AHB20, JCY20, Wan21b, YHSK22, ZCZ23]. Value [MJR20, BSMM20, Bic23, BPR20b, CHXL20, CTN22, CM24b, DRR22, DCP22, DC24, Fan22a, FYL^{+}24, GS21c, GJ21, Gu21, HL22c, IHK20, JGW^{+}23, JGZ^{+}24, JLW23, JM21, KZW22, KKJ20, LMZ22, LLLL20, MY21, MGL20, NL23a, OA20, ON22, PT21, PA21, PA24b, Psa24, QXZG22, RR22a, RL23, SV23, Shi21b, Sou22, SCZ20, TBA22, TBA24, TGS21, TBT^{+}20, TZ20, TKVA21, WSLJ23, WY24a, WJVG24, XTL23, YZ20a, YQZS22, YPL20, YS24a, YCD^{+}20, ZvS23, ZSMG20, ZF24, ZWL21, ZS22, ZCSS24]. valued [BEJR22, CLW^{+}22, RPQ24, UPPZ24, ZH23, ZZ24]. values [ABHK20, Boe20, CJP20, JB21, MKS21, MMJM22, MA21b, SL21, WTWZ20, WSLJ23]. Valuing [ZYY20]. Vandermonde [DKM^{+}23b, Yan22]. Vandermonde-type [DKM^{+}23b]. vanishing [WSY20, YSN24]. Vapor [CSLQ22]. Vapor-liquid [CSLQ22]. VAR [HQL21, CY20, Zha20b]. varentropy [SK24a]. variability [HCCSCGAM24]. Variable [BPRZ21, HPWM20, KSG^{+}22, QH23, Abi20, AEST21, AHT21, AMRP22, BJB20, CMS20a, CYY23, EKÇ20, FM21a, FLT20, FPR^{+}24, GZ23, GJN20, GLHL24, GJW23, GLHJ21, HZC^{+}23, HZ23, HYW20, HL24a, HCC24, HR22, HOGO23, HZZ24, JS22a, JK21a, JWZ21, KD20a, KJO23, KS21, LL24c, LC23, LXF21, LN23, NP23, PM20b, QHPL21, Ray23, RR23a, SLW21, SC21, SW21a, SCZ20, TBA22, TDL^{+}22, Tom24, VS21, WMH22, WCLY23, WYJ23, WY21a, WMK22, XZL20, YLFT21, YW23, YLLA23, ZVT^{+}23, ZEW23, ZPW^{+}24]. variable-density [WCLY23]. Variable-order [HPWM20, AEST21, HR22, HOGO23, JWZ21, LXF21, Ray23, WY21a, ZVT^{+}23]. variable-separated [GJW23]. variable-sign [Tom24]. variable-step [WYJ23]. variables [BG21, DG20, GJ19, GK21b, HZ24a, HA21b, HK21, HZS^{+}24, LLY21, ZSY24, ZW23, ZTY^{+}24]. variably [AK21]. variance [CKZ21, CES23, FPR^{+}24, GS24, hKhWjH21, LH23, PA21, PA24b, SYG19, TBA24, WZJQ21, YL22, ZCJL20, ZGZW23, LY24a, MSBSM23]. variance-reduced [YL22]. variant [FW22, GK21b, HWH^{+}21, JZM^{+}23, XHLD21]. variants [DDRS21, GHC^{+}20]. variation [BBBK24, FL24, LRT20, LZ24a, SVJG24, WY24a, YZM^{+}20]. Variational [CFdD23, GMW24, HKU22, Kov22, OBO23, YL22, BP21, BLRZ23, DKM23a, EK24b, HA21a, HTZ21, IHK20, ISD23, JDBT22, JHZY20, LCB20, LX24, LH23, MMM22, MYZ23, PC24, PKMM24, QZW23, SDL22, SLP20, Svá23, TV22, TLD^{+}23, TDAT23, TH24b, WW24a, Wen21a, bWLgZbX24, WZZ24, XL23, YYBL24, ZJ24, ZZ24, ZCCK24]. variational-hemivariational [LX24]. variations [BT23, MMM23]. varieties [BLV20]. variograms [CGSVBB^{+}24]. various [MMM23, RZA^{+}24, YDC20]. varying [AUAA20, CGO24, CLM22a, Fra21, GBTS22, HA20b, LPY20, Ruj23, SLZD20, SLW^{+}22, Tam20, XHLD21, XDX24, ZZ21, Zha20c, ZMZ^{+}24]. varying-coefficient [CLM22a, Zha20c]. varying-parameter [GBTS22]. Vasicek [MN20]. Vector [Han24, BZH24, BRZ22, CST19, CFB24a, CFB24b, DEH^{+}24, Guo24, HTZ21, LMM22, LLLD20, MWHC20, SBC^{+}24, SQT24, SH24a, WDPY24, YYT24, ZJBY23]. vector-wise [LLLD20]. Vectorial [ZY24]. vectorization [SFI23]. vectors [ERB^{+}22, Miy22, SHB24, YJN20]. vegetation [CPP23]. vehicle [TKT^{+}24, TFR^{+}20]. Velocity [BEG^{+}22, BEM^{+}24, FFT20, LRV23, PK24]. velocity-stress [LRV23]. vendor [WZ23c]. verification [ATO22, GHM21, HOOT20, IY20, Tan20]. Verified [IMT20, IMT23, Miy21, Rum20, Miy22, RKA20]. verifying [CDW20]. version [LZW20, MN20, WSY20]. versions [Kan20, YS24b]. versus [AAMR24, MMPM22, SBM21, WL21, WD20]. vertex [WSY22]. vertex-centered [WSY22]. vertical [EFJK22, gHtSjT24, WL24a, YYZB24]. very [BD20, RSS24]. vesicle [CZY22]. vesicles [LLT^{+}21]. VG [YHSK22]. via [AAP24, ACK23, ARF23, APR21, AN24, ASA^{+}21, BEL23a, BPDC20, BBTW24, BLV22, CS24, CMZ22, CZM23, CZ24, DAMA22, FFS24, FS24b, For24, GRVZ23, HJ20, HVR22b, Hu22, JDBT22, JW24, JHZD20, JCLC20, KZM21, LRT20, Li20a, Li20b, LHS^{+}24, LLGZ20, LMZ24, MK24, MMPP20, MPPR23, MT20b, Pat23, PP24, Pie20, QXZG22, RKS^{+}22, SM20, SDV21, SK24b, SLZD20, SN21b, SSVS20, TFR22, TTK23, Ust21, Ust23, Vab21, WZLX20, WCLW20, WCMW23, XHW24, Yak21, YZ20a, YFW20, ZZ24, ZG24, ZZNL24, Zho24, dA23, dHSR22]. vibration [BHPV23, GH21, LMV24, MJK20, SH21, ZUZ20]. vibrations [PSZ22]. video [ZLZ^{+}24]. VIEs [AID24a, AID24b]. Vieta [AEST21]. view [RKS^{+}22]. viewed [RB24]. viewpoint [ABA19]. VIPs [TDL^{+}22]. viral [WFZ22]. Virtual [LLZ24a, AMS23, BSZ21, FFS24, GG24, HY21b, LCZL24, MMF24, QZW23, WWH21a, WWH21b, XL23, MWBM22]. virus [AS23, CCJ20, Evi23]. viruses [DH20]. viscoelastic [DYF24, HJO20, JS22a, Let24, NA21, WGS20, ZW21]. viscoplastic [WC21]. Viscosity [XW21, KCJ23, PK24, RTV20, WCLY23]. viscous [AAA21, BEG^{+}22, BEM^{+}24, LSY23]. visualization [HP22, USTG23]. Vlasov [GGHY24, RS22b]. VMS [RFGGM20]. VNS [LPLP21]. vocal [ASX21, SH21]. Vogelius [Par20]. voids [MQ20]. Voigt [DYF24]. Volatility [CPS20, NW23, CKZ21, GJNM24, HCH^{+}23, KYP21, MYC22, MST22, Mit20, Ruj23, SYG19, YLWW21, YHF22, ZLL20a, ZGL20]. Volterra [FL23, AC21, ABP24, All24, AL16, AA24a, BVDB20, BLN21, BR22, CS24, CST19, DRR22, DDM20, DOR21, DO22, FO22, Fro24, GPST23, KN20, KM20, KM22, KAMW21, KXW23, Kum22, LBBB23, LHHW21, LHW22, LHH21, MH21, Ma24b, MS20a, MK24, MWW21, MPV23, MH23, MA21a, NB23, Pie20, RB19, RGA20, SES20, SESH23, SM22a, SYD21, SKVA22, SL24b, Ust21, WEEA20, WH24, YTMK19, YYY21, YTD23, YWZ24, YZM23, YBM23, uDuIZ22, ZE23, ZBS22, ZLG20, ZLCT22, ZFM22, ZH24a, ZCZ24, ZS22]. Volume [GQR23, ZZZ^{+}22, AJR22, AET23, CZY22, CL22, CLL20b, DKMT24, FLT20, FGPP23, GLYS21, GKV21, HET24, HW20, LJCL20, LZWC21, LLT^{+}21, Mia24, NSL22, PH21, RSRW24, TT20, Ter20, WSY22, WCLY23, WLH23, WLM24, WLS24, YLFT21, YF22, YCH24, YLLA23, YLS20, ZIZ21, ZW22]. volume-conserved [CZY22, LLT^{+}21, WCLY23]. Volume-preserving [ZZZ^{+}22]. volume/element [PH21]. volumes [YJZ23]. Volumetric [LHG^{+}24]. voluntary [MYLL24]. vortex [BHPV23]. vortex-induced [BHPV23]. vortices [PZY^{+}24]. vorticity [LRV23]. vorticity-based [LRV23]. VPNets [ZZZ^{+}22]. Vries [SC21, SCK22, DD23, LF24]. VSI [DFH21]. vuggy [WCC^{+}19]. vulnerable [Wan21b, WZ24b].

W [GPHAPR21, GPHHA23]. Wald [CMP21]. Wald-type [CMP21]. warning [DLL21, HQL21]. Wasserstein [HKM24, KDLY22]. waste [Kap21]. water [BBSG23, DL20b, FAVM19, FAVM20, FMAV23, SMM22, KD20a]. Watson [DJ22a]. Wave [CLM20, CLM21, ASADB23, AL20, AAHK24, BT21, BGdlR24, CW20, CQZ22, CH20, CLM22b, DL23b, DWZ20, ESA23, EGG23, EEE21, FFS24, GEG23, HL21b, HC21, Hu24, HLF19, JS22a, KA24, KD22, LSY23, Liu23, LY24b, MDS20, MHHV20, NA21, PSWZ21, SM23, SC21, SCK22, SDLZ23, SC23a, TXS21, WGS20, WC24b, WKN22, WX22, YZW21, bYW23, YY24b, ZAH^{+}24, ZWY22, ZLJT22, ZGZ23]. wavefield [WSX21]. waveform [YHW22]. wavefronts [BYWA22]. waveguides [Kim20, LYL23, Rod21]. Wavelet [Li24a, PSRVA24, WD20, YL21c, dLNO21, ASA^{+}21, BCP20, GNGGL^{+}19, HMA24, KGV23, LRT20, ÖKG23, Rab21, SSS24, YLXY23, YL20, ZK20, Zho22, SSVS20].

Wavelet-based [Li24a, PSRVA24, WD20]. wavelets [BR22, CF20, DOR21, HR23, IN22, RB19, Ray23, Ri22, SES20]. waves [BGH22, JA22, Lee23, RCN21, WZZT20, ZG24]. way [RK23, SLA^{+}24]. Weak [Gua20, KD22, LWZ24, Mu21, WGS20, WWZ23, WWZ24, ZR23, AAP24, ATM21, ATHW21, BT21, CWW23b, CWW23a, DCY24, DG22a, DYF24, Gao22, GDA22, GSWH22, HWAT22, JZ22, KD23, LGC23, LMP20, LWW23b, LWWY24, LLGC24a, LLGC24b, Li24b, LHMT20, LWWN22, LZ23b, LYZ24, MGD22, PL21, QS23, QSW24, SKVA23, WZZ20, Wan20a, WZ21, WWYZ21, WWH21b, WYZZ22, WMZZ23, WZ23a, WWTL24, WYZ24, Wu24, XFW23, YMC22, YZZ20a, YZ20b, YZ21c, YZ21b, ZTZZ20, ZZ20a, ZZ22b, ZWZ23c, ZWJ22]. Weakly [FO22, ADZ20, All24, CZHQ23, CS24, CJL^{+}24b, CSS24, JG21b, KN20, KM22, KXW23, LW22, Miy22, Naj20, RB19, Rou24, SRAG23, YWZ24, ZE23, ZLCT22, ZS22]. weakly-coupled [CSS24]. Weibull [BSNL19, BC24, JB21, WTL20, YWW24, Zhu20b]. Weierstrass [CFB24b, CFB24a, Ned22]. weighers [GJGDPR23]. weight [CCGT22, Dju24, JPS24, MSM24, OH23, Tom24, VTB21]. Weighted [DP21c, MMJM22, RU22, SK24a, SC24, SRAG23, XTL23, ZY20, All24, AORS22, BPDC20, FO22, IN22, JGW^{+}23, JGZ^{+}24, JZZH22, MSS^{+}21, ORT22, RCM20, SN21a, WCLW20, XXL21, ZTY^{+}24, vdMOB22]. weights [BCA20, JYWZ21, TBA24, vdBS21]. Well [JDBT22, TZ20, BE21, Hun21, Hun22, HK23, Kap21, OCS24, YZ24d]. well-posed [Hun22]. Well-posedness [JDBT22, TZ20, Hun21, HK23]. well-spaced [BE21]. Weller [SCZ20]. Wendroff [LLS22]. WENO [LZ23a]. wet [FHH21]. Whaley [BO22]. which [MDSRFO24]. while [vdBS21]. white [DCP22, DC24, TKVA21]. Whitehead [IYA^{+}23]. Whitney [LK21, RBR23]. Whitt [HMT23]. Whittaker [PP21]. whole [RRS24]. Wick [Yam21b]. wide [ZZNL24]. wide-band [ZZNL24]. width [BGdlR24]. Wiener [HVR22a, HVR22b, SQT24, YPLM22]. Wigner [JLZ22]. Wilbraham [BCA20, BCGAR22]. Wilk [MPGE24]. willow [ZX23b]. wind [AMS23, BLRZ23, Kan24, LRZ23]. wind-driven [AMS23]. wise [LLLD20, OHK23, WW23]. within [SQT24]. without [Dad20, DHF23, HG23, KV19, LMM22, LLX^{+}22b, MGG^{+}24, MMPP20, Shi21b, WYJ23, ZM21b]. WKB [KAD22b]. WKB-based [KAD22b]. Wolf [NRGJ21]. Wong [WC24b]. WOPSIP [ZZW^{+}24b, ZZW^{+}24c]. world [HLLZ24, QXZF23]. wormhole [TGJY22]. Wright [Ans21, CG23, KUA20]. WSGD [LL20b]. Wuytack [BPR22].

XII [QG20]. XOR [KV20]. xorshift128 [HMS22].

Yang [Lu24]. Yosida [RVX20].

Z [SLW^{+}22]. Z-type [SLW^{+}22]. Zakai [WC24b]. Zakharov [BGdlR24, MDMS22, WTX24, XYAH23]. Zernike [MPPR23]. Zero [MARH^{+}23, CLYS24, CFL21, DMP23, GS20c, MN20, MMV20, WSLJ23, YSFdRdP20, ZTY^{+}20, ZJBY23, ZGL20]. zero-and-one [ZTY^{+}20]. zero-coupon [MN20]. zero-mean [MMV20]. zero-sum [CLYS24, CFL21, YSFdRdP20]. zeroing [GBTS22, LWX24]. Zeros [BCBM22, BM24b, Che22, CDW20, DMN24, KJGN20, PH20b, SS23b, Vie21]. zeroth [AN24]. zeroth-order [AN24]. zeta [YT20]. Ziolkowski [HLF19]. ZNN [JZM^{+}23, XHLD21]. ZNN-based [XHLD21]. zonotopal [BCDF22].

References

Atangana:2020:CNN

[AA20a] Abdon Atangana and Seda Igret Araz. Corrigendum to “New numerical method for ordinary differential equations: Newton polynomial” [J. Comput. Appl. Math. 372 (2020) 112622]. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                                        www.sciencedirect.com/
                                                        science/article/pii/S0377042719306739.

See [AA20b] and retraction

notice [AA21c].

Atangana:2020:NNM

[AA20b] Abdon Atangana and Seda Igret Araz. New numerical method for ordinary differential equations: Newton polynomial. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                                        www.sciencedirect.com/
                                                        science/article/pii/S0377042719306272.

See [AA20a, AA21b] and retraction notice [AA21c].

Abualrub:2021:DPS

[AA21a] S. Abualrub and M. Aloqeili. Dynamics of positive solutions of a system of difference equations. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001084.

Atangana:2021:RNC

[AA21b] Abdon Atangana and Seda Igret Araz. Retraction notice to ‘corrigendum to “New numerical method for ordinary differential equations: Newton polynomial”’ [j. comput. appl. math. (2019) 112622]. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001679.

See [AA20a].

Atangana:2021:RNN

[AA21c] Abdon Atangana and Seda Igret Araz. Retraction notice to “New numerical method for ordinary differential equations: Newton polynomial” [J. Comput. Appl. Math. 371 (2020) 112668]. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001667.

See [AA20a].

Amirali:2022:TLD

[AA22] Ilhame Amirali and Gabil M. Amiraliyev. Three layer difference method for linear pseudo-parabolic equation with delay. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004088.

Ali:2023:SOC

[AA23] Hegagi Mohamed Ali and Ismail Gad Ameen. Stability and optimal control analysis for studying the transmission dynamics of a fractional-order MSV epidemic model. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002960.

Amirali:2024:SIN

[AA24a] Ilhame Amirali and Hülya Acar. Stability inequalities and numerical solution for neutral Volterra delay integro-differential equation. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300287X.

Amirali:2024:NSL

[AA24b] Ilhame Amirali and Gabil M.

Amiraliyev. Numerical solution of linear pseudoparabolic equation with time delay using three layer difference method. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003618.

Arjmand:2024:ELR

[AA24c] Doghonay Arjmand and Maksat Ashyraliyev. Efficient low rank approximations for parabolic control problems with unknown heat source. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002097.

Awan:2020:ESF

[AAA20] Aziz Ullah Awan, Mukarram Ali, and Kashif Ali Abro. Electroosmotic slip flow of Oldroyd-B fluid between two plates with nonsingular kernel. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030176X.

Ali:2021:AMF

[AAA21] Rizwan Ali, Muhammad Imran Asjad, and Ali Akgül. An analysis of a mathematical fractional model of hybrid viscous nanofluids and its application in heat and mass transfer. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303873.

Argyros:2023:LSL

[AAA^{+}23] C. Argyros, M. I. Argyros, I. K. Argyros, Á. A. Magreñán, and Í. Sarría. Local and semi-local convergence for Chebyshev two point like methods with applications in different fields. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300016X.

Amat:2020:LSC

[AAB^{+}20] S. Amat, I. Argyros, S. Busquier, M. A. Hernández-Verón, and D. F. Yañez. On the local and semilocal convergence of a parameterized multi-step Newton method. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301345.

Alizadeh:2024:RRU

[AAB24] M. Alizadeh, A. Asgharzadeh, and E. Basiri. Reconstruction regions for unobserved middle order statistics in the exponential distribution. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006398.

Alikhanov:2024:SOD

[AAHK24] Anatoly A. Alikhanov, Mohammad Shahbazi Asl, Chengming Huang, and Aslanbek Khibiev. A second-order difference scheme for the nonlinear time-fractional diffusionwave equation with generalized memory kernel in the presence of time delay. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004594.

Askar:2020:DID

[AAk20] S. S. Askar and A. Alkhedhairi. Dynamic investigations in a duopoly game with price competition based on relative profit and profit maximization. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304674.

AlAhmad:2024:SLN

[AAKA24] Rami AlAhmad, Mohammad Al-Khaleel, and Hasan Almefleh. On solutions of linear and nonlinear fractional differential equations with application to fractional order RC type circuits. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300451X.

Araya:2020:ASM

[AAM20] Rodolfo Araya, Jorge Aguayo, and Santiago Muñoz. An adaptive stabilized method for advection– diffusion–reaction equation. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301497.

Alazemi:2024:CVA

[AAMR24] Abdullah Alazemi, Milica Andelić, Jovana Milenković, and Jovana Nikolov Radenković. Combinatorial versus algebraic formulae for the Moore–Penrose inverse of a Laplacian matrix of a threshold graph. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300657X.

Abreu:2024:TBP

[AAP24] Eduardo Abreu, Jorge Agudelo, and John Pérez. A triangle-based positive semi-discrete Lagrangian– Eulerian scheme via the weak asymptotic method for scalar equations and systems of hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004090.

Amiri:2022:HIH

[AB22] Mehdi Amiri and Narayanaswamy Balakrishnan. Hessian and increasing-Hessian orderings of scale-shape mixtures of multivariate skewnormal distributions and applications. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004234.

Aravind:2023:SCM

[AB23] R. Vijay Aravind and P. Balasubramaniam. Stability criteria for memristorbased delayed fractionalorder Cohen–Grossberg neural networks with uncertainties. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003909.

Acay:2019:NLF

[ABA19] Bahar Acay, Erdal Bas, and Thabet Abdeljawad. Non-local fractional calculus from different viewpoint generated by truncated M-derivative. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304133.

Athanasiadou:2020:IES

[ABA20] E. S. Athanasiadou, N. Bardis, and I. Arkoudis. An inverse electromagnetic scattering problem for a layered ellipsoid. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303115.

Abid:2023:DCA

[ABBB23] Chaima Abid, Amel Ben Abda, Riadh Ben Fatma, and Yosra Boukari. A data completion algorithm using an integral representation of the Steklov– Poincaré operator. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004538.

Adriani:2024:ASL

[ABFSC24] Andrea Adriani, Davide Bianchi, Paola Ferrari, and Stefano Serra-Capizzano. Asymptotic spectra of large (grid) graphs with a uniform local structure, Part II: Numerical applications. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004053.

Aulisa:2020:FVA

[ABHK20] Eugenio Aulisa, Giorgio Bornia, Victoria Howle, and Guoyi Ke. Field-of-values analysis of preconditioned linearized Rayleigh– Bénard convection problems. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305874.

Abide:2020:FDP

[Abi20] Stéphane Abide. Finite difference preconditioning for compact scheme discretizations of the Poisson equation with variable coefficients. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301631.

Aledo:2022:FPG

[ABM^{+}22] Juan A. Aledo, Ali Barzanouni, Ghazaleh Malekbala, Leila Sharifan, and Jose C. Valverde. Fixed points in generalized parallel and sequential dynamical systems induced by a minterm or maxterm Boolean functions. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006294.

Ascione:2024:CDS

[ABO24] Giacomo Ascione, Michele Bufalo, and Giuseppe Orlando. Credit default swap spreads modeling and forecasting with a stochastic square-root three-factor model. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002437.

Abdi:2024:BRP

[ABP24] A. Abdi, J.-P. Berrut, and H. Podhaisky. The barycentric rational predictorcorrector schemes for Volterra integral equations. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005551.

Abdi:2021:IGL

[AC21] A. Abdi and D. Conte. Implementation of general linear methods for Volterra integral equations. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305525.

Alcantara:2022:NCN

[AC22] Jan Harold Alcantara and Jein-Shan Chen. A new class of neural networks for NCPs using smooth perturbations of the natural residual function. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000061.

Arnal:2023:MIL

[ACC23] Ana Arnal, Fernando Casas, and Cristina Chiralt. Magnus integrators for linear and quasilinear delay differential equations. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002170.

Almeida:2022:ADG

[ACD22] Rui M. P. Almeida, Teófilo D. Chihaluca, and José C. M. Duque. Approach to the Delta Greek of nonlinear Black–Scholes equation governing European options. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100412X.

Alberini:2022:TFH

[ACDV22] C. Alberini, R. Capitanelli, M. D’Ovidio, and S. Finzi Vita. On the time fractional heat equation with obstacle. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002266.

Ayala:2022:LMT

[ACEIJH22] Alan Ayala, Xavier Claeys, Paul Escapil-Inchauspé, and Carlos Jerez-Hanckes. Local multiple traces formulation for electromagnetics: Stability and preconditioning for smooth geometries. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001583.

Alsaker:2021:CUP

[ACFM21] Melody Alsaker, Diego Armando Cardona Cárdenas, Sérgio Shiguemi Furuie, and Jennifer L. Mueller. Complementary use of priors for pulmonary imaging with electrical impedance and ultrasound computed tomography. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002119.

Ayala:2020:LTC

[ACG20] Alan Ayala, Xavier Claeys, and Laura Grigori. Lineartime CUR approximation of BEM matrices. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305333.

Alevizakos:2021:TMA

[ACK21] Vasileios Alevizakos, Kashinath Chatterjee, and Christos Koukouvinos. The triple moving average control chart. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304623.

Abubakar:2023:IMS

[ACK23] Jamilu Abubakar, Parin Chaipunya, and Poom Kumam. Iterative method for split equilibrium problem and minimization problem via conjugate gradient method. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001358.

Araya:2024:ASF

[ACPV24] Rodolfo Araya, Cristian Cárcamo, Abner H. Poza, and Eduardo Vino. An adaptive stabilized finite element method for the Stokes–Darcy coupled problem. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000025.

Alvarruiz:2024:TRJ

[ACR24] Fernando Alvarruiz, Carmen Campos, and Jose E. Roman. Thick-restarted joint Lanczos bidiagonalization for the GSVD. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723004508.

Alotaibi:2022:GMF

[ACS22] Manal Alotaibi, Huangxin Chen, and Shuyu Sun. Generalized multiscale finite element methods for the reduced model of Darcy flow in fractured porous media. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001297.

Assoweh:2022:LTR

[ACT22] Mohamed Ibrahim Assoweh, Stéphane Chrétien, and Brahim Tamadazte. Low tubal rank tensor recovery using the Bürer– Monteiro factorisation approach. Application to optical coherence tomography. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000036.

Akbas:2021:HAP

[AD21] Serkan Akbas and Türkan Erbay Dalkiliç. A hybrid algorithm for portfolio selection: an application on the Dow Jones Index (DJI). Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003009.

Aimi:2022:HAQ

[ADDG22] Alessandra Aimi, Giulia Di Credico, Mauro Diligenti, and Chiara Guardasoni. Highly accurate quadrature schemes for singular integrals in energetic BEM applied to elastodynamics. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000619.

Adell:2022:EEC

[Ade22] José A. Adell. Explicit estimates for Comtet numbers of the first kind. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001996.

Alsuyuti:2021:SGS

[ADEEY21] M. M. Alsuyuti, E. H.

Doha, S. S. Ezz-Eldien, and I. K. Youssef. Spectral Galerkin schemes for a class of multi-order fractional pantograph equations. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304489.

Abbaszadeh:2020:AAI

[ADKH20] Mostafa Abbaszadeh, Mehdi Dehghan, Amirreza Khodadadian, and Clemens Heitzinger. Analysis and application of the interpolating element free Galerkin (IEFG) method to simulate the prevention of groundwater contamination with application in fluid flow. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930456X.

Aledo:2022:EPO

[ADMV22] Juan A. Aledo, Luis G. Diaz, Silvia Martinez, and Jose C. Valverde. Enumerating periodic orbits in sequential dynamical systems over graphs. Journal of Computational and Applied Mathematics, 405 (??):??, May 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303757.

Abbaszadeh:2020:CNG

[ADZ20] Mostafa Abbaszadeh, Mehdi Dehghan, and Yong Zhou. Crank–Nicolson/Galerkin spectral method for solving two-dimensional timespace distributed-order weakly singular integropartial differential equation. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300303.

Agarwal:2021:VFO

[AEST21] P. Agarwal, A. A. El-Sayed, and J. Tariboon. Vieta–Fibonacci operational matrices for spectral solutions of variableorder fractional integrodifferential equations. Journal of Computational and Applied Mathematics, 382 (??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272030354X.

Ahuja:2022:MOE

[AESW22] K. Ahuja, B. Endtmayer, M. C. Steinbach, and T. Wick. Multigoaloriented error estimation and mesh adaptivity for fluid-structure interaction. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001340.

Amaziane:2023:CMP

[AET23] Brahim Amaziane, Mustapha El Ossmani, and Khadija Talali. Convergence of a multi-point flux approximation finite volume scheme for a sharp-diffuse interfaces model for seawater intrusion. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003521.

Arabas:2020:DPT

[AF20] Sylwester Arabas and Ahmad Farhat. Derivative pricing as a transport problem: MPDATA solutions to Black–Scholes-type equations. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719302584.

Amini:2023:GCM

[AF23] Keyvan Amini and Parvaneh Faramarzi. Global convergence of a modified spectral three-term CG algorithm for nonconvex unconstrained optimization problems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003156.

Alba-Fernandez:2022:MSB

[AFJGJJ22] M. V. Alba-Fernández, M. D. Jiménez-Gamero, and F. Jiménez-Jiménez. Model selection based on penalized φ-divergences for multinomial data. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304726.

Alldredge:2022:RFI

[AFKM22] Graham Alldredge, Martin Frank, Jonas Kusch, and Ryan McClarren. A realizable filtered intrusive polynomial moment method. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100618X.

Abreu:2022:SDL

[AFLP22] Eduardo Abreu, Jean François, Wanderson Lambert, and John Pérez. A semi-discrete Lagrangian– Eulerian scheme for hyperbolictransport models. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005963.

Atencia:2021:DTQ

[AGGAV^{+}21] Iván Atencia, José Luis Galán-García, Gabriel Aguilera-Venegas, Pedro Rodríguez-Cielos, María Ángeles Galán-García, and Yolanda Padilla-Domínguez. A discrete-time queueing system with three different strategies. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001059.

Atencia:2023:NMR

[AGGAVGG23] Ivan Atencia, María Ángeles Galán-García, Gabriel Aguilera-Venegas, and José Luis Galán-García. A non Markovian retrial queueing system. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002212.

Abdulla:2020:OCC

[AGH20] Ugur G. Abdulla, Jonathan Goldfarb, and Ali Hagverdiyev. Optimal control of coefficients in parabolic free boundary problems modeling laser ablation. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300273.

Avila:2021:EMP

[ÁGKM21] Andrés I. Ávila, Galo Javier González, Stefan Kopecz, and Andreas Meister. Extension of modified Patankar–Runge–Kutta schemes to nonautonomous production– destruction systems based on Oliver’s approach. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306415.

Ammosov:2022:CNM

[AGMZ22a] D. A. Ammosov, A. V.

Grigorev, N. V. Malysheva, and L. S. Zamorshchikova. Corrigendum to “Numerical modeling two natural languages interaction” [j. comput. appl. math. 407 (2022) 114074]. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000504.

Ammosov:2022:NMT

[AGMZ22b] D. A. Ammosov, A. V. Grigorev, N. V. Malysheva, and L. S. Zamorshchikova. Numerical modeling two natural languages interaction. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006312.

See corrigendum [AGMZ22a].

Anton:2023:DDS

[AGNGG^{+}23] J. C. Álvarez Antón, P. J. García-Nieto, E. García-Gonzalo, M. González Vega, and C. Blanco Viejo. Data-driven stateof-charge prediction of a storage cell using ABC/ GBRT, ABC/MLP and LASSO machine learning techniques. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002492.

Ammosov:2023:NSL

[AGS^{+}23] D. A. Ammosov, A. V.

Grigorev, S. P. Stepanov, N. V. Malysheva, and L. S. Zamorshchikova. Numerical simulation of language interactions using online coupled Generalized Multiscale Finite Element Method. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200560X.

Alodat:2024:ATF

[AGS24] Tareq Alodat, Quoc T. Le Gia, and Ian H. Sloan. On approximation for timefractional stochastic diffusion equations on the unit sphere. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001122.

Ammosov:2023:PLUa

[AGST23] Dmitry Ammosov, Aleksandr Grigorev, Sergei Stepanov, and Aleksei Tyrylgin. Partial learning using partially explicit discretization for multicontinuum/multiscale problems. Fractured poroelastic media simulation. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200601X.

Alikhanov:2022:CTF

[AH22] Anatoly A. Alikhanov and Chengming Huang. A class of time-fractional diffusion equations with generalized fractional derivatives. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200200X.

Ahn:2023:HIQ

[AH23] Young Joon Ahn and Christoph Hoffmann. G^{2} Hermite interpolation with quartic regular linear normal curves. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005799.

Ahmadi:2020:LBA

[AHB20] Z. Ahmadi, S. Mohammad Hosseini, and A. Foroush Bastani. A latticebased approach to option and bond valuation under mean-reverting regimeswitching diffusion processes. Journal of Computational and Applied Mathematics, 363(??):156–170, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302742.

Ahmadi:2021:NLB

[AHB21] Z. Ahmadi, S. M. Hosseini, and A. Foroush Bastani. A new lattice-based scheme for swing option pricing under mean-reverting regimeswitching jump-diffusion processes. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304234.

Albertsson:2020:SCA

[AHGG20] Sigurdur Albertsson, Sigurdur Hafstein, Peter Giesl, and Skuli Gudmundsson. Simplicial complex with approximate rotational symmetry:a general class of simplicial complexes. Journal of Computational and Applied Mathematics, 363(??):413–425, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302924.

Arnold:2023:NSD

[AHK^{+}23] Martin Arnold, Maren Hantke, Raphael Kruse, Helmut Podhaisky, Elena Celledoni, Jason Frank, and Jens Lang. Numerical solution of differential and differential–algebraic equations. Selected papers from NUMDIFF-16. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001772.

Alazemi:2024:FPE

[AHK24] Abdullah Alazemi, Tim Hopkins, and Emrah Kiliç. A four parameter extension to the Clement matrix and its role in numerical software testing. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400236X.

Alcazar:2022:UBR

[AHPS22] Juan Gerardo Alcázar, Carlos Hermoso, Sonia Pérez Díaz, and Li-Yong Shen. Using μ-bases to reduce the degree in the computation of projective equivalences between rational curves in n -space. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002813.

Almani:2021:FBM

[AHT21] H. Maleki Almani, S. M.

Hosseini, and M. Tahmasebi. Fractional Brownian motion with twovariable Hurst exponent. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720305537.

Ahanchaou:2022:HIG

[AI22a] Taoufik Ahanchaou and Aziz Ikemakhen. Hyperbolic interpolatory geometric subdivision schemes. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721003381.

Azhar:2022:SFF

[AI22b] Noreen Azhar and Saleem Iqbal. Solution of fuzzy fractional order differential equations by fractional Mellin transform method. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003496.

Aourir:2024:CAS

[AID24a] E. Aourir, N. Izem, and H. Laeli Dastjerdi. A computational approach for solving third kind VIEs by collocation method based on radial basis functions. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005800.

Aourir:2024:EMT

[AID24b] E. Aourir, N. Izem, and H. Laeli Dastjerdi. An efficient meshless technique based on collocation and RBFs for solving nonlinear VIEs of third kind with proportional delays. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002875.

Alonso:2023:EPM

[AIDAJ23] José M. Alonso, J. Ibáñez, E. Defez, and P. Alonso-Jordá. Euler polynomials for the matrix exponential approximation. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000183.

Aishima:2020:QCA

[Ais20]  Kensuke Aishima.  A

quadratically convergent algorithm for inverse generalized eigenvalue problems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304881.

Aishima:2023:SMA

[Ais23] Kensuke Aishima. Statistical modeling and an adaptive averaging technique for strong convergence of the dynamic mode decomposition. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002710.

Alkhidhr:2021:CBM

[AJ21] Hanan Alkhidhr and Qingtang Jiang. Correspondence between multiwavelet shrinkage and nonlinear diffusion. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303654.

Amaziane:2022:CFV

[AJR22] Brahim Amaziane, Mladen Jurak, and Ivana Radisić. Convergence of a finite volume scheme for immiscible compressible two-phase flow in porous media by the concept of the global pressure. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003502.

Arana-Jimenez:2022:FDS

[AJSGL22] Manuel Arana-Jiménez, M. Carmen Sánchez-Gil, and Sebastián Lozano. A fuzzy DEA slacks-based approach. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304714.

Alevizakos:2020:DPM

[AK20] Vasileios Alevizakos and Christos Koukouvinos. A double progressive mean control chart for monitoring Poisson observations. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719301980.

Anuprienko:2021:NCM

[AK21] Denis Anuprienko and Ivan Kapyrin. Nonlinearity continuation method for steady-state groundwater flow modeling in variably saturated conditions. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001217.

Aretaki:2022:RGO

[AK22a] Aikaterini Aretaki and Efthymios N. Karatzas. Random geometries for optimal control PDE problems based on fictitious domain FEMs and cut elements. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001182.

Asar:2022:AUL

[AK22b] Yasin Asar and Merve Korkmaz. Almost unbiased Liu-type estimators in gamma regression model. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004416.

Abubakar:2024:HBL

[AKM^{+}24] A. B. Abubakar, P. Kumam, H. Mohammad, A. H. Ibrahim, T. Seangwattana, and B. A. Hassan. A hybrid BFGS-like method for monotone operator equations with applications. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001067.

Almeida:2021:OLF

[AKMO21] Ricardo Almeida, Rafał Kamocki, Agnieszka B. Malinowska, and Tatiana Odzijewicz. Optimal leader-following consensus of fractional opinion formation models. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302879.

Aghabazaz:2024:DMM

[AKN24] Zeynab Aghabazaz, Iraj Kazemi, and Alireza Nematollahi. Dynamic mixed models with heterogeneous covariance components using multivariate GARCH innovations and the Dirichlet process mixture. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300523X.

Atali:2023:NIR

[AKP23] Gokhan Atali, Hidayet Huda Kosal, and Muge Pekyaman. A new image restoration model associated with special elliptic quaternionic least-squares solutions based on LabVIEW. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000158.

Alvarez:2016:AFV

[AL16] Edgardo Alvarez and Carlos Lizama. Attractivity for functional Volterra integral equations of convolution type. Journal of Computational and Applied Mathematics, 301(??):230–240, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042716300280.

See comment [FL23].

Alama:2020:TWO

[AL20] Yvonne Bronsard Alama and Jean-Philippe Lessard. Traveling wave oscillatory patterns in a signed Kuramoto–Sivashinsky equation with absorption. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306132.

Alhaidari:2024:ESF

[Alh24] A. D. Alhaidari. Exact and simple formulas for the linearization coefficients of products of orthogonal polynomials and physical application. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003126.

Aljawi:2022:NCE

[Alj22] Salma Aljawi. Numerical computation of eigenvalues in spectral gaps of Schrödinger operators. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002199.

Aldoghaither:2020:PDO

[ALK20] Abeer Aldoghaither and Taous-Meriem Laleg-Kirati. Parameter and differentiation order estimation for a two dimensional fractional partial differential equation. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305758.

Axelsson:2021:IPF

[ALKH21] Owe Axelsson, Zhao-Zheng Liang, Jakub Kruzik, and David Horak. Inner product free iterative solution and elimination methods for linear systems of a three-by-three block matrix form. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304088.

Allouch:2024:FAS

[All24] Chafik Allouch. Fast and accurate solvers for weakly singular Volterra integral equations in weighted spaces. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300479X.

Abhishek:2024:CNM

[ALNK24] Anuj Abhishek, Thuy T.

Le, Loc H. Nguyen, and Taufiquar Khan. The Carleman–Newton method to globally reconstruct the initial condition for nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000761.

Ahmadova:2021:LDE

[AM21] Arzu Ahmadova and Nazim I.

Mahmudov. Langevin differential equations with general fractional orders and their applications to electric circuit theory. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305902.

Alcazar:2022:AES

[AM22a] Juan Gerardo Alcázar and Georg Muntingh. Affine equivalences of surfaces of translation and minimal surfaces, and applications to symmetry detection and design. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000759.

Asai:2022:EPC

[AM22b] Nobuyoshi Asai and Yoshinori Miyazaki. Eigenvalue problems for a class of infinite complex symmetric tridiagonal matrices with related three-term recurrence relation. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005677.

Arnal:2023:BSP

[AM23] A. Arnal and J. Monterde.

Bézier surfaces with prescribed diagonals. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006161.

Aragon:2022:ILC

[AMRP22] Roberto G. Aragón, Jesús Medina, and Eloísa Ramírez-Poussa. Impact of local congruences in variable selection from datasets. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721000352.

Adak:2023:MTV

[AMS23] D. Adak, D. Mora, and A. Silgado. A Morley-type virtual element approximation for a wind-driven ocean circulation model on polygonal meshes. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006240.

Arora:2023:FOM

[AMT23] Sugandha Arora, Trilok Mathur, and Kamlesh Tiwari. A fractional-order model to study the dynamics of the spread of crime. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000468.

Ammosov:2024:GMFb

[AMZ24a] D. A. Ammosov, N. V. Malysheva, and L. S. Zamorshchikova. Generalized multiscale finite element method for language competition modeling I: Offline approach. Journal of Computational and Applied Mathematics, 442 (??):??, May 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006738.

Ammosov:2024:GMFa

[AMZ24b] D. A. Ammosov, N. V. Malysheva, and L. S. Zamorshchikova. Generalized multiscale finite element method for language competition modeling II:

Online approach. Journal of Computational and Applied Mathematics, 442 (??):??, May 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300674X.

Arashi:2020:PAA

[AN20] M. Arashi and S. Nadarajah. A paradoxical argument about domination. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306697.

Almonacid:2024:CSF

[AN24] Javier A. Almonacid and Nilima Nigam. Characterization of singular flows of zeroth-order pseudodifferential operators via elliptic eigenfunctions: a numerical study. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004545.

Ammosov:2024:CMM

[ANM24] Dmitry Ammosov, Andrey Nasedkin, and Galina Muratova. A computational macroscopic model of piezomagnetoelectric materials using Generalized Multiscale Finite Element Method. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003643.

Arashi:2021:NLC

[ANN21] M. Arashi, S. Nadarajah, and D. K. Nagar. A note on linearly constrained Bayes estimator in elliptical models. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303794.

Anonymous:2019:EB

[Ano19]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305126.

Anonymous:2020:Da

[Ano20a] Anonymous. 1 December 2020. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2020:Db

[Ano20b] Anonymous. 15 December 2020. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2020:EBa

[Ano20c]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 363(??):ii, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303486.

Anonymous:2020:EBb

[Ano20d]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303887.

Anonymous:2020:EBc

[Ano20e]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304479.

Anonymous:2020:EBd

[Ano20f]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305588.

Anonymous:2020:EBe

[Ano20g]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306478.

Anonymous:2020:EBf

[Ano20h]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306958.

Anonymous:2020:EBg

[Ano20i]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300194.

Anonymous:2020:EBh

[Ano20j]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300418.

Anonymous:2020:EBi

[Ano20k]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300686.

Anonymous:2020:EBj

[Ano20l]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300558.

Anonymous:2020:EBk

[Ano20m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301412.

Anonymous:2020:EBl

[Ano20n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301898.

Anonymous:2020:EBm

[Ano20o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030220X.

Anonymous:2020:EBn

[Ano20p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302478.

Anonymous:2020:EBo

[Ano20q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302661.

Anonymous:2020:EBp

[Ano20r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302946.

Anonymous:2020:EBq

[Ano20s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030337X.

Anonymous:2021:Ab

[Ano21a] Anonymous. 1 August 2021. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Da

[Ano21b] Anonymous. 1 December 2021. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Ja

[Ano21c] Anonymous. 1 January 2021. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Ma

[Ano21d] Anonymous. 1 March 2021.

Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Mc

[Ano21e] Anonymous. 1 May 2021.

Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Oa

[Ano21f] Anonymous. 1 October 2021. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Ac

[Ano21g] Anonymous. 15 August 2021. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Db

[Ano21h] Anonymous. 15 December 2021. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Jb

[Ano21i] Anonymous. 15 January 2021. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Mb

[Ano21j] Anonymous. 15 March 2021. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Ob

[Ano21k] Anonymous. 15 October 2021. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Aa

[Ano21l] Anonymous. April 2021. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:EBa

[Ano21m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303940.

Anonymous:2021:EBb

[Ano21n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304362.

Anonymous:2021:EBc

[Ano21o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304763.

Anonymous:2021:EBd

[Ano21p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030515X.

Anonymous:2021:EBe

[Ano21q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305471.

Anonymous:2021:EBf

[Ano21r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305744.

Anonymous:2021:EBg

[Ano21s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306361.

Anonymous:2021:EBh

[Ano21t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000194.

Anonymous:2021:EBi

[Ano21u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000546.

Anonymous:2021:EBj

[Ano21v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000923.

Anonymous:2021:EBk

[Ano21w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001321.

Anonymous:2021:EBl

[Ano21x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001825.

Anonymous:2021:EBm

[Ano21y]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002065.

Anonymous:2021:EBn

[Ano21z]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002417.

Anonymous:2021:EBo

[Ano21-27]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002545.

Anonymous:2021:EBp

[Ano21-28]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002831.

Anonymous:2021:EBq

[Ano21-29]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003083.

Anonymous:2021:EBr

[Ano21-30]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003587.

Anonymous:2021:F

[Ano21-31]  Anonymous.  February

2021. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Jd

[Ano21-32] Anonymous. July 2021. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:Jc

[Ano21-33] Anonymous. June 2021. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:N

[Ano21-34] Anonymous. November 2021. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2021:S

[Ano21-35] Anonymous. September 2021. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Ab

[Ano22a] Anonymous. 1 August 2022. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Da

[Ano22b] Anonymous. 1 December 2022. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Ja

[Ano22c] Anonymous. 1 January 2022. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Ma

[Ano22d] Anonymous. 1 March 2022.

Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Md

[Ano22e] Anonymous. 1 May 2022. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Oa

[Ano22f] Anonymous. 1 October 2022. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Ac

[Ano22g] Anonymous. 15 August 2022. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Db

[Ano22h] Anonymous. 15 December 2022. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Jb

[Ano22i] Anonymous. 15 January 2022. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Mb

[Ano22j] Anonymous. 15 March 2022. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Mc

[Ano22k] Anonymous. 15 May 2022.

Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Ob

[Ano22l] Anonymous. 15 October 2022. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Aa

[Ano22m] Anonymous. April 2022. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:EBa

[Ano22n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003794.

Anonymous:2022:EBb

[Ano22o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004271.

Anonymous:2022:EBc

[Ano22p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100460X.

Anonymous:2022:EBd

[Ano22q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004817.

Anonymous:2022:EBe

[Ano22r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005409.

Anonymous:2022:EBf

[Ano22s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005793.

Anonymous:2022:EBg

[Ano22t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006269.

Anonymous:2022:EBh

[Ano22u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000188.

Anonymous:2022:EBi

[Ano22v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000541.

Anonymous:2022:EBj

[Ano22w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200070X.

Anonymous:2022:EBk

[Ano22x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000991.

Anonymous:2022:EBl

[Ano22y]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001443.

Anonymous:2022:EBm

[Ano22z]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001753.

Anonymous:2022:EBn

[Ano22-27]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001923.

Anonymous:2022:EBo

[Ano22-28]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002230.

Anonymous:2022:EBp

[Ano22-29]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002503.

Anonymous:2022:EBq

[Ano22-30]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002898.

Anonymous:2022:EBr

[Ano22-31]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003211.

Anonymous:2022:F

[Ano22-32]  Anonymous.  February

2022. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Jd

[Ano22-33] Anonymous. July 2022.

Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:Jc

[Ano22-34] Anonymous. June 2022.

Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:N

[Ano22-35] Anonymous. November 2022. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2022:S

[Ano22-36] Anonymous. September 2022. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Ab

[Ano23a] Anonymous. 1 August 2023. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Da

[Ano23b] Anonymous. 1 December 2023. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Ja

[Ano23c] Anonymous. 1 January 2023. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Ma

[Ano23d] Anonymous. 1 March 2023.

Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Md

[Ano23e] Anonymous. 1 May 2023.

Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Oa

[Ano23f] Anonymous. 1 October 2023. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Ac

[Ano23g] Anonymous. 15 August 2023. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Db

[Ano23h] Anonymous. 15 December 2023. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Jb

[Ano23i] Anonymous. 15 January 2023. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Mb

[Ano23j] Anonymous. 15 March 2023. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Mc

[Ano23k] Anonymous. 15 May 2023. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Ob

[Ano23l] Anonymous. 15 October 2023. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Aa

[Ano23m] Anonymous. April 2023.

Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:EBa

[Ano23n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003806.

Anonymous:2023:EBb

[Ano23o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004253.

Anonymous:2023:EBc

[Ano23p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004587.

Anonymous:2023:EBd

[Ano23q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005039.

Anonymous:2023:EBe

[Ano23r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005271.

Anonymous:2023:EBf

[Ano23s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005866.

Anonymous:2023:EBg

[Ano23t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200646X.

Anonymous:2023:EBh

[Ano23u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000237.

Anonymous:2023:EBi

[Ano23v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000559.

Anonymous:2023:EBj

[Ano23w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300119X.

Anonymous:2023:EBk

[Ano23x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001450.

Anonymous:2023:EBl

[Ano23y]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001917.

Anonymous:2023:EBm

[Ano23z]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002364.

Anonymous:2023:EBn

[Ano23-27]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300256X.

Anonymous:2023:EBo

[Ano23-28]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002698.

Anonymous:2023:EBp

[Ano23-29]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003023.

Anonymous:2023:EBq

[Ano23-30]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003175.

Anonymous:2023:EBr

[Ano23-31]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003333.

Anonymous:2023:F

[Ano23-32]  Anonymous.  February

2023. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Jd

[Ano23-33] Anonymous. July 2023. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:Jc

[Ano23-34] Anonymous. June 2023. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:N

[Ano23-35] Anonymous. November 2023. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2023:S

[Ano23-36] Anonymous. September 2023. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Ab

[Ano24a] Anonymous. 1 August 2024. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Da

[Ano24b] Anonymous. 1 December 2024. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Ma

[Ano24c] Anonymous. 1 March 2024. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Mc

[Ano24d] Anonymous. 1 May 2024. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Oa

[Ano24e] Anonymous. 1 October 2024. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Ac

[Ano24f] Anonymous. 15 August 2024. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Db

[Ano24g] Anonymous. 15 December 2024. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Ja

[Ano24h] Anonymous. 15 January 2024. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Mb

[Ano24i] Anonymous. 15 May 2024.

Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Ob

[Ano24j] Anonymous. 15 October 2024. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Aa

[Ano24k] Anonymous. April 2024.

Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:EBc

[Ano24l]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 435(??):??, ???? 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004673.

Anonymous:2024:EBd

[Ano24m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004892.

Anonymous:2024:EBe

[Ano24n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005356.

Anonymous:2024:EBf

[Ano24o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005605.

Anonymous:2024:EBg

[Ano24p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006349.

Anonymous:2024:EBh

[Ano24q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006805.

Anonymous:2024:EBi

[Ano24r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000281.

Anonymous:2024:EBj

[Ano24s]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272400058X.

Anonymous:2024:EBk

[Ano24t]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000980.

Anonymous:2024:EBl

[Ano24u]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001298.

Anonymous:2024:EBm

[Ano24v]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001572.

Anonymous:2024:EBn

[Ano24w]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001985.

Anonymous:2024:EBo

[Ano24x]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002577.

Anonymous:2024:EBp

[Ano24y]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002796.

Anonymous:2024:EBq

[Ano24z]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003078.

Anonymous:2024:EBr

[Ano24-27]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003753.

Anonymous:2024:EBs

[Ano24-28]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724004229.

Anonymous:2024:F

[Ano24-29]  Anonymous.  February

2024. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Jc

[Ano24-30] Anonymous. July 2024.

Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:Jb

[Ano24-31] Anonymous. June 2024.

Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:N

[Ano24-32] Anonymous. November 2024. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2024:S

[Ano24-33] Anonymous. September 2024. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Ahmed:2020:AAT

[ANR20] Elyes Ahmed, Jan Martin Nordbotten, and Florin Adrian Radu. Adaptive asynchronous time-stepping, stopping criteria, and a posteriori error estimates for fixed-stress iterative schemes for coupled poromechanics problems. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303097.

Ansari:2021:LFE

[Ans21] Alireza Ansari. Logarithmic functions are eigenfunctions of integral operators with M-Wright kernels. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000765.

Anton:2020:EPS

[Ant20] Cristina Anton. Explicit pseudo-symplectic methods based on generating functions for stochastic Hamiltonian systems. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719304364.

Andreani:2022:EDW

[AORS22] R. Andreani, H. Oviedo, M. Raydan, and L. D. Secchin. An extended delayed weighted gradient algorithm for solving strongly convex optimization problems. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002588.

Alahmadi:2021:CEB

[APR21] J. Alahmadi, M. Pranić, and L. Reichel. Computation of error bounds via generalized Gauss– Radau and Gauss–Lobatto rules. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002247.

Almutairi:2024:EEQ

[APR24] Hanan Almutairi, Miroslav Pranić, and Lothar Reichel. Error estimates for quadrature rules based on the Arnoldi process. Journal of Computational and Applied Mathematics, 442 (??):??, May 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006696.

Alonso:2022:ASS

[APS22] P. Alonso, J. M. Peña, and M. L. Serrano. Almost strictly sign regular rectangular matrices. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030412X.

Arnold:2021:NSD

[APW^{+}21] Martin Arnold, Helmut Podhaisky, Rüdiger Weiner, Elena Celledoni, Jason Frank, and Jens Lang. Numerical solution of differential and differential– algebraic equations. Selected papers from NUMDIFF-

15. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306168.

Alcazar:2020:AEI

[AQ20] Juan Gerardo Alcázar and Emily Quintero. Affine equivalences, isometries and symmetries of ruled rational surfaces. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930336X.

Araneda:2020:FMF

[Ara20] Axel A. Araneda. The fractional and mixedfractional CEV model. Journal of Computational and Applied Mathematics, 363(??):106–123, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930278X.

Al-Refai:2023:GFC

[ARF23] Mohammed Al-Refai and Arran Fernandez. Generalising the fractional calculus with Sonine kernels via conjugations. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001036.

Auer:2020:EBP

[ARK20] Ekaterina Auer, Andreas Rauh, and Julia Kersten. Experiments-based parameter identification on the GPU for cooperative systems. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306624.

Alonso-Revenga:2020:NST

[ARMP20] J. M. Alonso-Revenga, N. Martín, and L. Pardo. New statistics to test log– linear modeling hypothesis with no distributional specifications and clusters with homogeneous correlation. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300480.

Arshad:2022:DDE

[Ars22] Muhammad Arshad. Domain decomposition and expanded mixed method for parabolic partial differential equations. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000607.

Acitas:2020:RCP

[AS20a] Sukru Acitas and Birdal Senoglu. Robust change point estimation in twophase linear regression models: an application to metabolic pathway data. Journal of Computational and Applied Mathematics, 363(??):337–349, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302936.

Alodat:2020:GPR

[AS20b] M. T. Alodat and Mohammed K. Shakhatreh. Gaussian process regression with skewed errors. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306703.

AlSayedAli:2021:AIS

[AS21a] Mouhamad Al Sayed Ali and Miloud Sadkane. Acceleration of implicit schemes for large linear systems of differential–algebraic equations. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306555.

Arezoomandan:2021:SCM

[AS21b] Mahdieh Arezoomandan and Ali R. Soheili. Spectral collocation method for stochastic partial differential equations with fractional Brownian motion. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306609.

Avci:2023:OCS

[AS23] Derya Avci and Fatma Soytürk. Optimal control strategies for a computer network under virus threat. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003727.

Amin:2021:EAN

[ASA^{+}21] Rohul Amin, Kamal Shah, Muhammad Asif, Imran Khan, and Faheem Ullah. An efficient algorithm for numerical solution of fractional integrodifferential equations via Haar wavelet. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303198.

Asadi:2024:PTS

[Asa24] Majid Asadi. On the phase transition of k-out-of-n systems with applications to optimal maintenance. Journal of Computational and Applied Mathematics, 435(??):??, ???? 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002303.

Abdelkawy:2023:SSC

[ASADB23] M. A. Abdelkawy, E. M.

Soluma, Ibrahim Al-Dayel, and Dumitru Baleanu. Spectral solutions for a class of nonlinear wave equations with Riesz fractional based on Legendre collocation technique. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005684.

Al-Sultani:2022:NSN

[ASB22] Mohamed Al-Sultani and Igor Boglaev. Numerical solution of nonlinear parabolic systems by block monotone methods. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001170.

Asker:2021:SDN

[Ask21] Hussein K. Asker. Stability in distribution of numerical solution of neutral stochastic functional differential equations with infinite delay. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002478.

Abdulsalam:2023:AMS

[ASML23] Athraa Abdulsalam, Norazak Senu, Zanariah Abdul Majid, and Nik Mohd Asri Nik Long. Adaptive multi-step Runge– Kutta–Nyström methods for general second-order ordinary differential equations. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004721.

Afuecheta:2023:CED

[ASON23] Emmanuel Afuecheta, Mujahida Sayyed, Idika E. Okorie, and Saralees Nadarajah. A compound exponential distribution with application to control charts. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002989.

Ammosov:2023:PLUb

[AST^{+}23] D. A. Ammosov, S. P.

Stepanov, A. A. Tyrylgin, N. V. Malysheva, and L. S. Zamorshchikova. Partial learning using partially explicit discretization for multicontinuum/multiscale problems with limited observation: Language interactions simulation. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200632X.

Ammosov:2024:BDM

[ASTK24] Dmitry Ammosov, Sergei Stepanov, Aleksei Tyrylgin, and Alexander Karandeev. Bayesian decision making using partial data for fractured poroelastic media. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300482X.

Aktosun:2021:IPD

[ASX21] Tuncay Aktosun, Paul Sacks, and Xiao-Chuan Xu. An inverse problem to determine the shape of a human vocal tract. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000960.

Al-Taweel:2021:SFW

[ATHW21] Ahmed Al-Taweel, Saqib Hussain, and Xiaoshen Wang. A stabilizer free weak Galerkin finite element method for parabolic equation. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306646.

Al-Taweel:2021:NUW

[ATM21] Ahmed Al-Taweel and Lin Mu. A new upwind weak Galerkin finite element method for linear hyperbolic equations. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306671.

Asai:2022:NVA

[ATO22] Taisei Asai, Kazuaki Tanaka, and Shin’ichi Oishi. Numerical verification for asymmetric solutions of the Hénon equation on bounded domains. Journal of Computational and Applied Mathematics, 399 (??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721003307.

Alsaedi:2020:FTS

[AUAA20] Ahmed Alsaedi, M. Usha, M. Syed Ali, and Bashir Ahmad. Finite-time synchronization of sampleddata Markovian jump complex dynamical networks with additive time-varying delays based on dissipative theory. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305837.

Araneda:2021:CCO

[AV21] Axel A. Araneda and Marcelo J. Villena. Computing the CEV option pricing formula using the semiclassical approximation of path integral. Journal of Computational and Applied Mathematics, 388 (??):??, May 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305355.

Ammosov:2022:GMF

[AVC22] Dmitry Ammosov, Maria Vasilyeva, and Eric T. Chung. Generalized Multiscale Finite Element Method for thermoporoelasticity problems in heterogeneous and fractured media. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005884.

Argun:2022:NSF

[AVL22] R. L. Argun, V. T. Volkov, and D. V. Lukyanenko. Numerical simulation of front dynamics in a nonlinear singularly perturbed reaction–diffusion problem. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001224.

Aguilera-Venegas:2023:CTM

[AVLMRMGG23] Gabriel Aguilera-Venegas, Amador López-Molina, Gemma Rojo-Martínez, and José Luis Galán-García. Comparing and tuning machine learning algorithms to predict type 2 Diabetes mellitus. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000596.

Aguilera-Venegas:2023:PMD

[AVRLRMGG23] Gabriel Aguilera-Venegas, Eugenio Roanes-Lozano, Gemma Rojo-Martínez, and José Luis Galán-García. A proposal of a mixed diagnostic system based on decision trees and probabilistic experts rules. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000742.

Axelsson:2020:ECF

[Axe20] Owe Axelsson. Extensions of a coarse-fine mesh stabilized Schwarz alternating iteration domain decomposition method. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303383.

Athanasiadou:2020:NFI

[AZ20] E. S. Athanasiadou and S. Zoi. A near-field inverse scattering problem for an elastic ellipsoid. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305345.

Ai:2022:PSL

[AZ22] Meiqiao Ai and Zhimin Zhang. Pricing some life-contingent lookback options under regimeswitching Lévy models. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000012.

Ameen:2021:SPS

[AZD21] Ibrahem G. Ameen, Mahmoud A. Zaky, and Eid H. Doha. Singularity preserving spectral collocation method for nonlinear systems of fractional differential equations with the right-sided Caputo fractional derivative. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100087X.

An:2021:TCA

[AZL21] Rong An, Chao Zhang, and Yuan Li. Temporal convergence analysis of an energy preserving projection method for a coupled magnetohydrodynamics equations. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305276.

Ai:2023:VDU

[AZZ23] Meiqiao Ai, Zhimin Zhang, and Wei Zhong. Valuation of a DB underpin hybrid pension under a regime-switching Lévy model. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003715.

Bahrami:2020:ETS

[BA20] Somayeh Bahrami and Keyvan Amini. An efficient two-step trust-region algorithm for exactly determined consistent systems of nonlinear equations. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930473X.

Behl:2022:CNS

[BA22] Ramandeep Behl and Himani Arora. CMMSE: a novel scheme having seventh-order convergence for nonlinear systems. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305926.

Bozorgnia:2023:ASE

[BA23] Farid Bozorgnia and Avetik Arakelyan. Approximation of the second eigenvalue of the p-Laplace operator in symmetric domains. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002935.

Aribi:2020:SGO

[BAB20] Walid Ben Aribi, Hamadi Ammar, and Mohamed Ben Alaya. Stochastic global optimization using tangent minorants for Lipschitz functions. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304650.

Barmalzan:2020:SCS

[BABR20] Ghobad Barmalzan, Seyed Masih Ayat, Narayanaswamy Balakrishnan, and Rasool Roozegar. Stochastic comparisons of series and parallel systems with dependent heterogeneous extended exponential components under Archimedean copula. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302569.

Baccouch:2019:PEA

[Bac19] Mahboub Baccouch. A posteriori error analysis of the local discontinuous Galerkin method for the sine-Gordon equation in one space dimension. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304352.

Berger:2022:EPD

[BAD^{+}22] Guillaume O. Berger, Pierre-Antoine Absil, Lieven De Lathauwer, Raphaël M. Jungers, and Marc Van Barel. Equivalent polyadic decompositions of matrix multiplication tensors. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005562.

Bellaihou:2024:SMI

[BAIB24] Mohamed Bellaihou, Taoufik Ahanchaou, Aziz Ikemakhen, and Alaa Eddine Bensad. Smooth morphing of immersed spherical curves and application to ruled surfaces. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002231.

Balandin:2024:ISP

[Bal24]  A. L. Balandin.  In-

verse scattering problem for the tensor fields in a coordinate-free representation. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006404.

Baz:2023:GMR

[BAPPF23] Juan Baz, Pedro Alonso, Juan Manuel Peña, and Raúl Pérez-Fernández. Gaussian Markov Random Fields and totally positive matrices. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000420.

Bayramoglu:2022:FID

[Bay22] Ismihan Bayramoglu. Fuzzy improved distribution function and order statistics. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001091.

Biswas:2020:DCT

[BB20a] Milan Biswas and Nandadulal Bairagi. On the dynamic consistency of a twospecies competitive discrete system with toxicity: Local and global analysis. Journal of Computational and Applied Mathematics, 363(??):145–155, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302778.

Breitenbach:2020:SQH

[BB20b] Tim Breitenbach and Alfio Borzì. A sequential quadratic Hamiltonian scheme for solving non-smooth quantum control problems with sparsity. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305886.

Bartmanski:2021:EDD

[BB21a] Bartosz J. Bartmanski and Ruth E. Baker. Effects of different discretisations of the Laplacian upon stochastic simulations of reaction–diffusion systems on both static and growing domains. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001904.

Bazhlekova:2021:ISD

[BB21b] Emilia Bazhlekova and Ivan Bazhlekov. Identification of a space-dependent source term in a nonlocal problem for the general time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305045.

Boateng:2023:TIT

[BB23] Henry A. Boateng and Kyle Bradach. Triquintic interpolation in three dimensions. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300198X.

Baltaeva:2023:SMP

[BBAO23] Umida Baltaeva, Yulduz Babajanova, Praveen Agarwal, and Necati Ozdemir. Solvability of a mixed problem with the integral gluing condition for a loaded equation with the Riemann–Liouville fractional operator. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000109.

Benchettou:2024:CTT

[BBBK24] Oumaima Benchettou, Abdeslem Hafid Bentbib, Abderrahman Bouhamidi, and Karim Kreit. Constrained tensorial total variation problem based on an alternating conditional gradient algorithm. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002681.

Belafhal:2022:AME

[BBBU22] A. Belafhal, H. Benzehoua, A. Balhamri, and T. Usman. An advanced method for evaluating Lommel integral and its application in marine environment. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002990.

Bellet:2022:QSC

[BBC22] Jean-Baptiste Bellet, Matthieu Brachet, and Jean-Pierre Croisille. Quadrature and symmetry on the Cubed Sphere. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000358.

Barletti:2024:RAN

[BBGI24] Luigi Barletti, Luigi Brugnano, Gianmarco Gurioli, and Felice Iavernaro. Recent advances in the numerical solution of the Nonlinear Schrödinger Equation. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400075X.

Behl:2022:EHO

[BBMK22] Ramandeep Behl, Sonia Bhalla, Á. A. Magreñán, and Sanjeev Kumar. An efficient high order iterative scheme for large nonlinear systems with dynamics. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305409.

Bellosta:2023:LEA

[BBSG23] Tommaso Bellosta, Giacomo Baldan, Giuseppe Sirianni, and Alberto Guardone. Lagrangian and Eulerian algorithms for water droplets in in-flight ice accretion. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001747.

Baricz:2021:GMF

[BBSV21] Árpád Baricz, Nitin Bisht, Sanjeev Singh, and V. Antony Vijesh. The generalized Marcum function of the second kind: Monotonicity patterns and tight bounds. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303848.

BenHammouda:2024:AIS

[BBTW24] Chiheb Ben Hammouda, Nadhir Ben Rached, Raúl Tempone, and Sophia Wiechert. Automated importance sampling via optimal control for stochastic reaction networks: a Markovian projectionbased approach. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400102X.

Barletti:2020:SAS

[BBTZ20] Luigi Barletti, Luigi Brugnano, Yifa Tang, and Beibei Zhu. Spectrally accurate space-time solution of Manakov systems. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302090.

Beccari:2021:MRM

[BC21] Carolina Vittoria Beccari and Giulio Casciola. Matrix representations for multi-degree Bsplines. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302983.

Beccari:2022:SNE

[BC22a] Carolina Vittoria Beccari and Giulio Casciola. Stable numerical evaluation of multi-degree Bsplines. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003654.

Bica:2022:CPT

[BC22b] Alexandru Mihai Bica and Diana Curila. The convergence properties of a type II complete deficient quartic spline. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000140.

Bellet:2023:LSS

[BC23] Jean-Baptiste Bellet and Jean-Pierre Croisille. Least squares spherical harmonics approximation on the Cubed Sphere. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001577.

Balakrishnan:2024:RID

[BC24] N. Balakrishnan and E. Castilla. Robust inference for destructive one-shot device test data under Weibull lifetimes and competing risks. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003965.

Berriochoa:2020:GWP

[BCA20] E. Berriochoa, A. Cachafeiro, and J. M. García Amor. Gibbs-Wilbraham phenomenon on Lagrange interpolation based on analytic weights on the unit circle. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303796.

Bultheel:2022:ZQP

[BCBM22] Adhemar Bultheel, Ruymán Cruz-Barroso, and Carlos Díaz Mendoza. Zeros of quasi-paraorthogonal polynomials and positive quadrature. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006105.

Bo:2022:EIE

[BCCW22] Yonghui Bo, Jiaxiang Cai, Wenjun Cai, and Yushun Wang. The exponential invariant energy quadratization approach for general multi-symplectic Hamiltonian PDEs. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100563X.

Barucq:2022:PRS

[BCDF22] Hélène Barucq, Henri Calandra, Julien Diaz, and Stefano Frambati. Polynomial-reproducing spline spaces from fine zonotopal tilings. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004349.

Berriochoa:2022:GWP

[BCGAR22] E. Berriochoa, A. Cachafeiro, J. García-Amor, and H. García Rábade. The Gibbs– Wilbraham phenomenon in the approximation of |x| by using Lagrange interpolation on the Chebyshev– Lobatto nodal systems. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001856.

Banhelyi:2020:ESU

[BCH20] Balázs Bánhelyi, Tibor Csendes, and László Hatvani. On the existence and stabilization of an upper unstable limit cycle of the damped forced pendulum. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307071.

Bialas-Ciez:2024:CAM

[BCKSV24] Leokadia Białas-Ciez, Dimitri Jordan Kenne, Alvise Sommariva, and Marco Vianello. Chebyshev admissible meshes and Lebesgue constants of complex polynomial projections. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000153.

Badia:2024:PLC

[BCLS24] F. G. Badía, J. H. Cha, H. Lee, and C. Sangüesa. Preservation of the log concavity by Bernstein operator with an application to ageing properties of a coherent system. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006908.

Beccari:2020:CLA

[BCM20] Carolina Vittoria Beccari, Giulio Casciola, and Marie-Laurence Mazure. Critical length: an alternative approach. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306089.

Balakrishnan:2023:PDA

[BCMP23] N. Balakrishnan, E. Castilla, N. Martin, and L. Pardo. Power divergence approach for one-shot device testing under competing risks. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003429.

Bruni:2020:FLD

[BCP20] Vittoria Bruni, Mariantonia Cotronei, and Francesca Pitolli. A family of level-dependent biorthogonal wavelet filters for image compression. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304704.

Beccari:2022:PMC

[BCR22] Carolina Vittoria Beccari, Giulio Casciola, and Lucia Romani. A practical method for computing with piecewise Chebyshevian splines. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006166.

Behl:2022:HOF

[BCT22a] Ramandeep Behl, Alicia Cordero, and Juan R. Torregrosa. High order family of multivariate iterative methods: Convergence and stability. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303447.

Behl:2022:NHO

[BCT22b] Ramandeep Behl, Alicia Cordero, and Juan R. Torregrosa. A new higherorder optimal derivative free scheme for multiple roots. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003952.

Boccali:2024:CRO

[BCV24] Lorenzo Boccali, Danilo Costarelli, and Gianluca Vinti. Convergence results in Orlicz spaces for sequences of max-product sampling Kantorovich operators. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002073.

Burgos:2020:MSC

[BCVV20] C. Burgos, J.-C. Cortés, L. Villafuerte, and R.-J. Villanueva. Mean square convergent numerical solutions of random fractional differential equations: approximations of moments and density. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302168.

Bertaccini:2020:CFV

[BD20] D. Bertaccini and F. Durastante. Computing functions of very large matrices with small TT/ QTT ranks by quadrature formulas. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306685.

Barigou:2022:PEL

[BD22] Karim Barigou and Lukasz Delong. Pricing equitylinked life insurance contracts with multiple risk factors by neural networks. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005458.

Banach:2023:NAC

[BD23] Marzena Banach and Rafał Długosz. A novel approach to cities’ assessment in terms of their implementation of smart city idea. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300105X.

Bailey:2024:DTO

[BD24] Rachel Bailey and Maxim Derevyagin. DEK-type orthogonal polynomials and a modification of the Christoffel formula. Journal of Computational and Applied Mathematics, 438 (??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723005058.

Bhal:2021:CNO

[BDF21] Santosh Kumar Bhal, P. Danumjaya, and G. Fairweather. The Crank– Nicolson orthogonal spline collocation method for one-dimensional parabolic problems with interfaces. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304106.

Boatman:2023:NAT

[BDPR23] K. Boatman, L. Davis, F. Pahlevani, and T. Susai Rajan. Numerical analysis of a time filtered scheme for a linear hyperbolic equation inspired by DNA transcription modeling. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000791.

Buccini:2023:TNA

[BDR23] Alessandro Buccini, Marco Donatelli, and Lothar Reichel. Theoretical and numerical aspects of a nonstationary preconditioned iterative method for linear discrete ill-posed problems. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005386.

Brenner:2020:RSS

[BDS20] Susanne C. Brenner, Amanda E.

Diegel, and Li-Yeng Sung. A robust solver for a second order mixed finite element method for the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930319X.

Berrut:2021:BLC

[BE21] Jean-Paul Berrut and Giacomo Elefante. Bounding the Lebesgue constant for a barycentric rational trigonometric interpolant at periodic wellspaced nodes. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002867.

Bakharev:2022:VVF

[BEG^{+}22] F. Bakharev, A. Enin, A. Groman, A. Kalyuzhnyuk, S. Matveenko, Yu. Petrova, I. Starkov, and S. Tikhomirov. Velocity of viscous fingers in miscible displacement: Comparison with analytical models. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004301.

Bubba:2023:ERS

[BEH^{+}23] Tatiana A. Bubba, Glenn Easley, Tommi Heikkilä, Demetrio Labate, and Jose P. Rodriguez Ayllon. Efficient representation of spatio-temporal data using cylindrical shearlets. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001504.

Barrera:2023:SQI

[BEIR23] D. Barrera, S. Eddargani, M. J. Ibáñez, and S. Remogna. Spline quasiinterpolation in the Bernstein basis on the Powell– Sabin 6-split of a type-1 triangulation. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006094.

Bentbib:2020:GMK

[BEJ20] A. H. Bentbib, M. El Guide, and K. Jbilou. A generalized matrix Krylov subspace method for TV regularization. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930408X.

Bentbib:2022:ESB

[BEJR22] A. H. Bentbib, M. El Ghomari, K. Jbilou, and L. Reichel. The extended symmetric block Lanczos method for matrixvalued Gauss-type quadrature rules. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006099.

Bakharev:2023:OPS

[BEK^{+}23] Fedor Bakharev, Aleksandr Enin, Konstantin Kalinin, Yulia Petrova, Nikita Rastegaev, and Sergey Tikhomirov. Optimal polymer slugs injection profiles. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006409.

Barrera:2022:NCB

[BEL22] D. Barrera, S. Eddargani, and A. Lamnii. A novel construction of B-spline-like bases for a family of many knot spline spaces and their application to quasi-interpolation. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003836.

Barrera:2023:CQI

[BEL23a] D. Barrera, S. Eddargani, and A. Lamnii. On C^{2} cubic quasi-interpolating splines and their computation by subdivision via blossoming. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004320.

Bellet:2023:DFT

[Bel23b] Jean-Baptiste Bellet. A discrete Funk transform on the Cubed Sphere. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001498.

Bakharev:2024:VVF

[BEM^{+}24] Fedor Bakharev, Aleksandr Enin, Sergey Matveenko, Dmitry Pavlov, Yulia Petrova, Nikita Rastegaev, and Sergey Tikhomirov. Velocity of viscous fingers in miscible displacement: Intermediate concentration. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400356X.

BenSalah:2024:SLR

[Ben24] Mohamed BenSalah. Shape and location recovery of laser excitation sources in photoacoustic imaging using topological gradient optimization. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002978.

Bortoloti:2022:EDN

[BFF22] M. A. A. Bortoloti, T. A.

Fernandes, and O. P. Ferreira. An efficient damped Newton-type algorithm with globalization strategy on Riemannian manifolds. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004763.

Bazarra:2024:ATP

[BFLQ24] N. Bazarra, J. R. Fernández, L. Liverani, and R. Quintanilla. Analysis of a thermoelastic problem with the Moore–Gibson– Thompson microtemperatures. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005150.

Bazarra:2021:AMG

[BFQ21a] N. Bazarra, J. R. Fernández, and R. Quintanilla. Analysis of a Moore–Gibson– Thompson thermoelastic problem. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303496.

Bazarra:2021:TIT

[BFQ21b] N. Bazarra, J. R. Fernández, and R. Quintanilla. A type III thermoelastic problem with mixtures. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306488.

Bazarra:2022:NAT

[BFQ22] N. Bazarra, J. R. Fernández, and R. Quintanilla. Numerical analysis of a thermoelastic dielectric problem arising in the Moore– Gibson–Thompson theory. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002151.

Bazarra:2024:SGP

[BFQ24] N. Bazarra, J. R. Fernández, and R. Quintanilla. A strain gradient problem with a fourth-order thermal law. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000670.

Breda:2021:ENC

[BFRV21] Dimitri Breda, Francesco Florian, Jordi Ripoll, and Rossana Vermiglio. Efficient numerical computation of the basic reproduction number for structured populations. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304568.

BenGharbia:2023:SSN

[BFVY23] Ibtihel Ben Gharbia, Joëlle Ferzly, Martin Vohralík, and Soleiman Yousef. Semismooth and smoothing Newton methods for nonlinear systems with complementarity constraints: Adaptivity and inexact resolution. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003892.

Bayramoglu:2021:MMM

[BG21] Ismihan Bayramoglu and Omer L. Gebizlioglu. A max-min model of random variables in bivariate random sequences. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305951.

Bradley:2023:EBS

[BG23] Susanne Bradley and Chen Greif. Eigenvalue bounds for saddle-point systems with singular leading blocks. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722005945.

Bruzon:2024:TWS

[BGdlR24] M. S. Bruzón, T. M. Garrido, and R. de la Rosa. Travelling wave solutions for a Zakharov–Kuznetsov modified equal width equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001820.

Benito:2021:SRD

[BGG^{+}21] J. J. Benito, A. García, L. Gavete, M. Negreanu, F. Ureña, and A. M. Vargas. Solving a reaction– diffusion system with chemotaxis and non-local terms using Generalized Finite Difference Method. Study of the convergence. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306166.

Barkhan:2022:ARW

[BGH22] J. Barkhan, M. Ganesh, and S. C. Hawkins. Approximation of radiating waves in the near-field: Error estimates and application to a class of inverse problems. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003915.

Brugnano:2024:EBA

[BGIR24] Luigi Brugnano, Domenico Giordano, Felice Iavernaro, and Giorgia Rubino. An entropy-based approach for a robust least squares spline approximation. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000220.

Bokil:2020:EMM

[BGN^{+}20] V. A. Bokil, N. L. Gibson, S. L. Nguyen, E. A. Thomann, and E. C. Waymire. An Euler– Maruyama method for diffusion equations with discontinuous coefficients and a family of interface conditions. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305503.

Baradol:2020:CFP

[BGR20] Pravin Baradol, Dhananjay Gopal, and Stojan Radenović. Computational fixed points in graphical rectangular metric spaces with application. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300960.

Bermudez:2020:MAN

[BGV20] A. Bermúdez, D. Gómez, and P. Venegas. Mathematical analysis and numerical solution of models with dynamic Preisach hysteresis. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304558.

Bhatt:2021:PER

[Bha21] Ashish Bhatt. Projected exponential Runge–Kutta methods for preserving dissipative properties of perturbed constrained Hamiltonian systems. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100176X.

Bhattacharyya:2023:USB

[Bha23] Biswarup Bhattacharyya. On the use of sparse Bayesian learning-based polynomial chaos expansion for global reliability sensitivity analysis. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004174.

Buong:2021:NIR

[BHB21] Nguyen Buong, Pham Thi Thu Hoai, and Khuat Thi Binh. New iterative regularization methods for the multiple-sets split feasibility problem. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305823.

Braun:2021:LBD

[BHJB21] Phillip Braun, Warren Hare, and Gabriel Jarry-Bolduc. Limiting behavior of derivative approximation techniques as the number of points tends to infinity on a fixed interval in R. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305094.

Bublik:2023:NNB

[BHPV23] Ondrej Bublík, Václav Heidler, Ales Pecka, and Jan Vimmr. Neural-network-based fluid-structure interaction applied to vortexinduced vibration. Journal of Computational and Applied Mathematics, 428 (??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723001140.

Bai:2022:NST

[BHRW22] Xixian Bai, Jian Huang, Hongxing Rui, and Shuang Wang. Numerical simulation for 2D/3D time fractional Maxwell’s system based on a fast secondorder FDTD algorithm. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002941.

Bica:2021:MDO

[Bic21] Alexandru Mihai Bica. On the minimal derivative oscillation for deficient quartic splines of Marsden’s type. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305677.

Bica:2023:OCI

[Bic23] Alexandru Mihai Bica. On the order of convergence of the iterative Bernstein splines method for fractional order initial value problems. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002182.

Bras:2021:NCS

[BIJ21] Michał Braś, Giuseppe Izzo, and Zdzisław Jackiewicz. A new class of strong stability preserving general linear methods. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100234X.

Boen:2021:OSS

[BitH21] Lynn Boen and Karel J.

in ’t Hout. Operator splitting schemes for the two-asset Merton jumpdiffusion model. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930305X.

Brugnano:2020:AHO

[BIZ20] Luigi Brugnano, Felice Iavernaro, and Ruili Zhang. Arbitrarily high-order energypreserving methods for simulating the gyrocenter dynamics of charged particles. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302855.

Ba:2022:RDA

[BJ22] Yuming Ba and Lijian Jiang. A residual-driven adaptive Gaussian mixture approximation for Bayesian inverse problems. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003290.

Babaei:2020:NSV

[BJB20] A. Babaei, H. Jafari, and S. Banihashemi. Numerical solution of variable order fractional nonlinear quadratic integrodifferential equations based on the sixth-kind Chebyshev collocation method. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301990.

Banihashemi:2022:SCA

[BJB22] S. Banihashemi, H. Jafari, and A. Babaei. A stable collocation approach to solve a neutral delay stochastic differential equation of fractional order. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004684.

Bi:2024:IPT

[BJF24] Yuxin Bi, Xiaofeng Jia, and Hui Feng. An improved partitioned time stepping method based on modified characteristic FEM for the evolutionary dual-porosity-Navier– Stokes model. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001870.

Bini:2020:NAS

[BJMR20] Dario Bini, Khalide Jbilou, Marilena Mitrouli, and Lothar Reichel. Numerical Analysis and Scientific Computation (NASCA18). Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271930617X.

Balakrishnan:2024:NDO

[BJP24] Narayanaswamy Balakrishnan, María Jaenada, and Leandro Pardo. Nondestructive one-shot device test under stepstress experiment with lognormal lifetime distribution. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004272.

Botchev:2020:AAR

[BK20] M. A. Botchev and L. A. Knizhnerman. ART: Adaptive residual-time restarting for Krylov subspace matrix exponential evaluations. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303085.

Bauer:2021:EMI

[BK21a] Tobias Peter Bauer and Oswald Knoth. Extended multirate infinitesimal step methods: Derivation of order conditions. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305461.

Brenner:2021:AIP

[BK21b] Susanne C. Brenner and Ellya L. Kawecki. Adaptive C^{0} interior penalty methods for Hamilton–Jacobi– Bellman equations with Cordes coefficients. Journal of Computational and Applied Mathematics, 388 (??):??, May 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030532X.

Bhushan:2022:PEA

[BK22] Shashi Bhushan and Anoop Kumar. Predictive estimation approach using difference and ratio type estimators in ranked set sampling. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000796.

Bayramukov:2024:NSM

[BK24] Alim A. Bayramukov and Nikolay A. Kudryashov. Numerical study of the model described by the fourth order generalized nonlinear Schrödinger equation with cubic-quintic-septic-nonic nonlinearity. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004417.

Birgin:2022:IRD

[BKM22] E. G. Birgin, N. Krejić, and J. M. Martínez. Inexact restoration for derivativefree expensive function minimization and applications. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000656.

Babaie-Kafaki:2024:CCA

[BKMA24] Saman Babaie-Kafaki, Nasrin Mirhoseini, and Zohre Aminifard. A class of CG algorithms overcoming jamming of the iterative solving process and its application in image restoration. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006702.

Bukovsek:2019:NEC

[BKMO19] Damjana Kokol Bukovsek, Tomaz Kosir, Blaz Mojskerc, and Matjaz Omladic. Non-exchangeability of copulas arising from shock models. Journal of Computational and Applied Mathematics, 358(??):61– 83, October 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0377042719301049.

See erratum [BKMO20].

Bukovsek:2020:ENE

[BKMO20] Damjana Kokol Bukovsek, Tomaz Kosir, Blaz Mojskerc, and Matjaz Omladic. Erratum to “Nonexchangeability of copulas arising from shock models” [J. Comput. Appl. Math. 358 (2019) 61– 83]. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719304224.

Bukovsek:2021:SFG

[BKMO21] Damjana Kokol Bukovsek, Tomaz Kosir, Blaz Mojskerc, and Matjaz Omladic. Spearman’s footrule and Gini’s gamma: Local bounds for bivariate copulas and the exact region with respect to Blomqvist’s beta. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000042.

Benterki:2020:CTC

[BL20] Rebiha Benterki and Jaume Llibre. The centers and their cyclicity for a class of polynomial differential systems of degree 7. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304595.

Bizzarri:2021:IHD

[BL21] Michal Bizzarri and Miroslav Lávicka. Interpolation of Hermite data by clamped Minkowski Pythagorean hodograph B-spline curves. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000881.

Belousov:2023:MFD

[BL23] Danila Belousov and Vadim Lisitsa. Mimetic finite differences for boundaries misaligned with grid nodes. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001292.

Bai:2021:SID

[BLBH21] Xuchao Bai, Xiangrong Li, Narayanaswamy Balakrishnan, and Mu He. Statistical inference for dependent stress-strength reliability of multi-state system using generalized survival signature. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306075.

Bazm:2021:AET

[BLN21] Sohrab Bazm, Pedro Lima, and Somayeh Nemati. Analysis of the Euler and trapezoidal discretization methods for the numerical solution of nonlinear functional Volterra integral equations of Urysohn type. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002508.

Boyana:2023:CAS

[BLRZ23] Satyajith Bommana Boyana, Thomas Lewis, Aaron Rapp, and Yi Zhang. Convergence analysis of a symmetric dual-wind discontinuous Galerkin method for a parabolic variational inequality. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005209.

Bizzarri:2020:CPE

[BLV20] Michal Bizzarri, Miroslav Lávicka, and Jan Vrsek. Computing projective equivalences of special algebraic varieties. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304418.

Bizzarri:2022:SDC

[BLV22] Michal Bizzarri, Miroslav Lávicka, and Jan Vrsek. Symmetries of discrete curves and point clouds via trigonometric interpolation. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000255.

Brenner:2020:MMS

[BLyS20] Susanne C. Brenner, Sijing Liu, and Li yeng Sung. Multigrid methods for saddle point problems: optimality systems. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300248.

Bao:2022:CRB

[BLZL22] Wendi Bao, Zhonglu Lv, Feiyu Zhang, and Weiguo Li. A class of residualbased extended Kaczmarz methods for solving inconsistent linear systems. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002606.

Bessaih:2020:SHM

[BM20] Hakima Bessaih and Razvan Florian Maris. Stochastic homogenization of multicontinuum heterogeneous flows. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300376.

Bot:2021:SCK

[BM21] Radu Ioan Bot and Dennis Meier. A strongly convergent Krasnosel’skiĭ–Manntype algorithm for finding a common fixed point of a countably infinite family of nonexpansive operators in Hilbert spaces. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002090.

Bukovsek:2022:ERD

[BM22a] Damjana Kokol Bukovsek and Blaz Mojskerc. On the exact region determined by Spearman’s footrule and Gini’s gamma. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000784.

Buvoli:2022:SEI

[BM22b] Tommaso Buvoli and Michael L.

Minion. On the stability of exponential integrators for non-diffusive equations. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000267.

Bucataru:2024:FBA

[BM24a] Mihai Bucataru and Liviu Marin. FDM-based alternating iterative algorithms for inverse BVPs in 2D steady-state anisotropic heat conduction with heat sources. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003017.

Bultheel:2024:EPU

[BM24b] Adhemar Bultheel and Carlos Díaz Mendoza. Extremal polynomials on the unit circle with preassigned zeros and Twopoint Partial Padé approximation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004247.

Birhanu:2021:DPP

[BMS21] Zerihun Kinfe Birhanu, Tadele Mengesha, and Abner J. Salgado. The Darcy problem with porosity depending exponentially on the pressure. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002648.

Bieniasz:2022:GJI

[BMS22] L. K. Bieniasz, S. McKee, and B. D. Sleeman. A generalized Jaeger I(0, 1; t) integral, resulting from mathematical modelling in electroanalytical chemistry. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200005X.

Bieniasz:2023:RDS

[BMV23] L. K. Bieniasz, S. Mc-Kee, and M. Vynnycky. A reaction–diffusion system arising from electrochemistry. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005593.

Bakar:2021:CMU

[BN21] S. A. Abu Bakar and S. Nadarajah. Composite models with underlying folded distributions. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306427.

Bufalo:2022:IBA

[BO22] Michele Bufalo and Giuseppe Orlando. An improved Barone–Adesi Whaley formula for turbulent markets. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005872.

Boen:2020:ERO

[Boe20] Lynn Boen. European rainbow option values under the two-asset Merton jump-diffusion model. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303413.

Boglaev:2021:PRN

[Bog21] Igor Boglaev. A parameter robust numerical method for a nonlinear system of singularly perturbed elliptic equations. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303083.

Bongiorno:2024:BRS

[Bon24] Christian Bongiorno. Bootstraps regularize singular correlation matrices. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002085.

Bosner:2024:HOA

[Bos24] Tina Bosner. High order approximation by CCCspline quasi-interpolants. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723006581.

Berinde:2021:KFP

[BP21] Vasile Berinde and Madalina Pacurar. Kannan’s fixed point approximation for solving split feasibility and variational inequality problems. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305082.

Benner:2022:EAA

[BP22a] Peter Benner and Carolin Penke. Efficient and accurate algorithms for solving the Bethe–Salpeter eigenvalue problem for crystalline systems. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002727.

Butcher:2022:ITO

[BP22b] John C. Butcher and Helmut Podhaisky. Isomeric trees and the order of Runge–Kutta methods. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200231X.

Blasco:2020:NAC

[BPD20] Angel Blasco and Sonia Pérez-Díaz. A new approach for computing the asymptotes of a parametric curve. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930353X.

Barton:2020:EMS

[BPDC20] Michael Barton, Vladimir Puzyrev, Quanling Deng, and Victor Calo. Efficient mass and stiffness matrix assembly via weighted Gaussian quadrature rules for B-splines. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306314.

Buccini:2020:CPP

[BPR20a] Alessandro Buccini, Yonggi Park, and Lothar Reichel. Comparison of aposteriori parameter choice rules for linear discrete illposed problems. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930069X.

See corrigendum [BPR21].

Buccini:2020:GSV

[BPR20b] Alessandro Buccini, Mirjeta Pasha, and Lothar Reichel. Generalized singular value decomposition with iterated Tikhonov regularization. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719302596.

Buccini:2021:CCP

[BPR21] Alessandro Buccini, Yonggi Park, and Lothar Reichel. Corrigendum to “Comparison of a-posteriori parameter choice rules for linear discrete ill-posed problems” [J. Comput. Appl. Math. 373 (2020) 112138]. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001916.

See [BPR20a].

Bultheel:2022:EML

[BPR22] Adhemar Bultheel, Miodrag Petkovic, and Lucia Romani. Editorial in memoriam Luc Wuytack. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004313.

Bonettini:2021:VMT

[BPRZ21] S. Bonettini, F. Porta, V. Ruggiero, and L. Zanni. Variable metric techniques for forward-backward methods in imaging. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304830.

Benyoussef:2020:ESC

[BR20a] Soufiane Benyoussef and Azedine Rahmoune. Efficient spectral-collocation methods for a class of linear Fredholm integrodifferential equations on the half-line. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301850.

Buccini:2020:MMC

[BR20b] Alessandro Buccini and Lothar Reichel. An ℓ^{p}– ℓ^{q} minimization method with cross-validation for the restoration of impulse noise contaminated images. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301151.

Behera:2022:ENM

[BR22] S. Behera and S. Saha Ray. An efficient numerical method based on Euler wavelets for solving fractional order pantograph Volterra delayintegro-differential equations. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004477.

Braun:2021:CDO

[Bra21] M. Braun. Comparison of densities obtained with competing density functional molecular codes. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002715.

Braun:2023:NSO

[Bra23] Moritz Braun. Numerical solution of the one dimensional Schrödinger equation using a basis set of scaled and shifted sinc functions on a finite interval. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001681.

Breitenbach:2022:SMS

[Bre22] Tim Breitenbach. On the SQH method for solving optimal control problems with non-smooth state cost functionals or constraints. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002539.

Brenner:2023:MCJ

[Bre23] Konstantin Brenner. On the monotone convergence of Jacobi-Newton method for mildly nonlinear systems. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003648.

Bauer:2023:IOD

[BRLS23] Benjamin Bauer, Michael Roller, Joachim Linn, and Bernd Simeon. An isogeometric one-dimensional Kirchhoff–Love type model for developable elastic ribbons. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004514.

Bronsard:2023:EAC

[Bro23] Yvonne Alama Bronsard. Error analysis of a class of semi-discrete schemes for solving the Gross– Pitaevskii equation at low regularity. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003168.

Brezinski:2022:EPS

[BRZ22] Claude Brezinski and Michela Redivo-Zaglia. Extrapolation and prediction of sequences in a vector space. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000462.

Balashov:2020:TCS

[BS20] V. A. Balashov and E. B.

Savenkov. Thermodynamically consistent spatial discretization of the one-dimensional regularized system of the Navier– Stokes–Cahn–Hilliard equations. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300340.

Belhamadia:2023:CEN

[BS23] Youssef Belhamadia and Mohammed Seaid. Computing enhancement of the nonlinear SP_{N} approximations of radiative heat transfer in participating material. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002868.

Bu:2024:SAD

[BS24a] Qihui Bu and Yun Sun.

Stationary analyses for a double-ended queueing system with random service capacity and balking customers. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300417X.

Bukovsek:2024:ERD

[BS24b] Damjana Kokol Bukovsek and Nik Stopar. On the exact region determined by Spearman’s rho and Spearman’s footrule. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004077.

Baamonde-Seoane:2023:PRE

[BSdCCGV23] María A. Baamonde-Seoane, María del Carmen Calvo-Garrido, and Carlos Vázquez. Pricing renewable energy certificates with a Crank–Nicolson Lagrange– Galerkin numerical method. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042722004897.

Barakat:2024:AMF

[BSH24] H. M. Barakat, A. A.

Semida, and M. H. Harpy. Asymptotic maxima of folded distributions with application to the multivariate extreme theory. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000141.

Becerra-Sagredo:2020:GBM

[BSMM20] Julián-Tercero Becerra-Sagredo, Carlos Málaga, and Francisco Mandujano. A GPU-based multi-level algorithm for boundary value problems. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304844.

Bai:2019:IAD

[BSNL19] Xuchao Bai, Yimin Shi, Hon Keung Tony Ng, and Yiming Liu. Inference of accelerated dependent competing risks model for Marshall–Olkin bivariate Weibull distribution with nonconstant parameters. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304017.

Bauman:2020:CAC

[BSV20] B. Bauman, A. Sommariva, and M. Vianello. Compressed algebraic cubature over polygons with applications to optical design. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306636.

Brugnano:2021:FVS

[BSZ21] Luigi Brugnano, Yaroslav D. Sergeyev, and Anatoly Zhigljavsky. Foreword to the virtual special issue dedicated to the 3rd International Conference NUMTA 2019 “Numerical Computations: Theory and Algorithms”. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001643.

Baccouch:2021:HOS

[BT21] Mahboub Baccouch and Helmi Temimi. A highorder space-time ultraweak discontinuous Galerkin method for the secondorder wave equation in one space dimension. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306221.

Bahadir:2023:ANA

[BT23] Eylem Bahadir and Önder Türk. Analytical and numerical assessments of boundary variations in Steklov eigenvalue problems. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004988.

Baccouch:2021:DGM

[BTBR21] Mahboub Baccouch, Helmi Temimi, and Mohamed Ben-Romdhane. A discontinuous Galerkin method for systems of stochastic differential equations with applications to population biology, finance, and physics. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305884.

Bagirov:2021:IDO

[BTC21] Adil M. Bagirov, Sona Taheri, and Emre Cimen. Incremental DC optimization algorithm for largescale clusterwise linear regression. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306142.

Bunger:2020:TMT

[Bün20] Florian Bünger. A Taylor model toolbox for solving ODEs implemented in MATLAB/INTLAB. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305163.

Bibi:2024:NMS

[BuR24] Amna Bibi and Mujeeb ur Rehman. A numerical method for solutions of tempered fractional differential equations. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000219.

Baeyens:2022:ICE

[BV22] Toon Baeyens and Marnix Van Daele. Improvements to the computation of eigenvalues and eigenfunctions of twodimensional Schrödinger equations by constant perturbation based algorithms. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001212.

Barnsley:2023:HFF

[BV23] Michael F. Barnsley and P. Viswanathan. Histopolating fractal functions. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000171.

Bhattacharjee:2023:MRD

[BVBO23] Atanu Bhattacharjee, Gajendra K. Vishwakarma, Souvik Banerjee, and Seng Huat Ong. A modified risk detection approach of biomarkers by frailty effect on multiple time to event data. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003442.

Barikbin:2020:ISC

[BVDB20] M. S. Barikbin, A. R. Vahidi, T. Damercheli, and E. Babolian. An iterative shifted Chebyshev method for nonlinear stochastic Itô–Volterra integral equations. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030203X.

Wu:2024:NES

[bWLgZbX24] Zeng bao Wu, Wei Li, Quan guo Zhang, and Yi bin Xiao. New existence and stability results of mild solutions for fuzzy fractional differential variational inequalities. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001766.

Bourne:2020:CDG

[BWS20] Martin Bourne, Joab R.

Winkler, and Yi Su. The computation of the degree of the greatest common divisor of three Bernstein basis polynomials. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303760.

Bourne:2021:AFT

[BWS21] Martin Bourne, Joab R.

Winkler, and Yi Su. An approximate factorisation of three bivariate Bernstein basis polynomials defined in a triangular domain. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306725.

Bas:2020:PFR

[BYE20] Eren Bas, Ufuk Yolcu, and Erol Egrioglu. Picture fuzzy regression functions approach for financial time series based on ridge regression and genetic algorithm. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306612.

Bilgil:2023:FOM

[BYE^{+}23] Halis Bilgil, Ali Yousef, Ayhan Erciyes, Ümmügülsüm Erdinç, and Zafer Öztürk. A fractional-order mathematical model based on vaccinated and infected compartments of SARS-CoV-2 with a real case study during the last stages of the epidemiological event. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006136.

Yan:2023:IFO

[bYW23] Xiong bin Yan and Ting Wei. Identifying a fractional order and a timedependent coefficient in a time-fractional diffusion wave equation. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005933.

Bani-Yaghoub:2022:STA

[BYWA22] Majid Bani-Yaghoub, Xueying Wang, and Sharif. S. Aly. Spatio-temporal analysis of coinfection using wavefronts of Escherichia coli O157: H7 in a dairy cattle farm. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005537.

Bao:2020:CLM

[BZ20] Lianzhang Bao and Zhengfang Zhou. Continuum and lattice models analysis of the aggregation diffusion cell movement. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306879.

Ban:2024:TPP

[BZH24] Bohui Ban, Xuzhou Zhan, and Yongjian Hu. The twopoint Padé approximation problem and its Hankel vector. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001614.

Bai:2023:ISS

[BZHB23] Xuchao Bai, Jieqiong Zhang, Mu He, and Narayanaswamy Balakrishnan. Inference for stress-strength reliability of multi-state system with dependent stresses and strengths using improved generalized survival signature. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004113.

Chou:2023:CDE

[CA23] So-Hsiang Chou and C. Attanayake. Construction of discontinuous enrichment functions for enriched or generalized FEM’s for interface elliptic problems in 1D. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001243.

Chaskalovic:2024:NSO

[CAJ24] Joël Chaskalovic, Franck Assous, and Hessam Jamshidipour. A new second order Taylorlike theorem with an optimized reduced remainder. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004405.

Calvo:2023:UAI

[Cal23] María Cabrera Calvo. Uniformly accurate integrators for Klein–Gordon– Schrödinger systems from the classical to nonrelativistic limit regime. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042722003855.

Chakraborty:2023:SMG

[CAN23] Samiran Chakraborty, Shivam Kumar Agrawal, and Gnaneshwar Nelakanti. Superconvergent multi-Galerkin method for nonlinear Fredholm– Hammerstein integral equations. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000365.

Cao:2020:BPS

[Cao20]  Yang Cao.  A block

positive-semidefinite splitting preconditioner for generalized saddle point linear systems. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300789.

Capdeville:2020:CHO

[Cap20]  G. Capdeville.  Com-

pact high-order numerical schemes for scalar hyperbolic partial differential equations. Journal of Computational and Applied Mathematics, 363(??):171– 210, January 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S037704271930264X.

Chakraborty:2024:OPT

[CBF24] Niladri Chakraborty, Narayanaswamy Balakrishnan, and Maxim Finkelstein. Optimal precedence tests under single and double-sampling framework. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000542.

Chen:2020:CFC

[CBJ20] Churong Chen, Martin Bohner, and Baoguo Jia. Caputo fractional continuous cobweb models. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030025X.

Corbino:2020:HOM

[CC20] Johnny Corbino and Jose E.

Castillo. High-order mimetic finite-difference operators satisfying the extended Gauss divergence theorem. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303231.

Chaudhry:2021:PEE

[CC21a] Jehanzeb H. Chaudhry and J. B. Collins. A posteriori error estimation for the spectral deferred correction method. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303885.

Chen:2021:UML

[CC21b] Ting-Hsuan Chen and Rong-Cih Chang. Using machine learning to evaluate the influence of FinTech patents: the case of Taiwan’s financial industry. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305069.

Cao:2024:ENF

[CC24] Pei Cao and Jinru Chen.

An extended nonconforming finite element method for the coupled Darcy– Stokes problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003418.

Chang:2024:GSR

[CCCMH24] Miaoxin Chang, Frank P. A. Coolen, Tahani Coolen-Maturi, and Xianzhen Huang. A generalized system reliability model based on survival signature and multiple competing failure processes. Journal of Computational and Applied Mathematics, 435(??):??, ???? 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003084.

Casas:2021:CPS

[CCETZ21] Fernando Casas, Philippe Chartier, Alejandro Escorihuela-Tomàs, and Yong Zhang. Compositions of pseudosymmetric integrators with complex coefficients for the numerical integration of differential equations. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302971.

Chicharro:2022:EMW

[CCGT22] Francisco I. Chicharro, Alicia Cordero, Neus Garrido, and Juan R. Torregrosa. On the effect of the multidimensional weight functions on the stability of iterative processes. Journal of Computational and Applied Mathematics, 405 (??):??, May 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303435.

Chen:2024:OIP

[CCHW24] Xinyue Chen, Peimin Chen, Yong He, and Xiaoyang Wang. The optimal investment problem with inflation and liquidity risk. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005241.

Chen-Charpentier:2020:DIO

[CCJ20] Benito M. Chen-Charpentier and Mark Jackson. Direct and indirect optimal control applied to plant virus propagation with seasonality and delays. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302740.

Cheung:2020:CEM

[CCL20] Siu Wun Cheung, Eric T.

Chung, and Wing Tat Leung. Constraint energy minimizing generalized multiscale discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030251X.

Cai:2023:LSN

[CCL23] Zhiqiang Cai, Jingshuang Chen, and Min Liu. Least-squares neural network (LSNN) method for scalar nonlinear hyperbolic conservation laws: Discrete divergence operator. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300242X.

Cross:2022:CRM

[CCM22] Ben Cross, Robert J. Cripps, and Glen Mullineux. C^{1} and G^{1} continuous rational motions using a conformal geometric algebra. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001169.

Calzola:2024:HOS

[CCS24] Elisa Calzola, Elisabetta Carlini, and Francisco J. Silva. A high-order scheme for mean field games. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000189.

Caliari:2023:BMF

[CCZ23] Marco Caliari, Fabio Cassini, and Franco Zivcovich. BAMPHI: Matrix-free and transpose-free action of linear combinations of ϕfunctions from exponential integrators. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005714.

Cano:2020:NNS

[CD20a] B. Cano and A. Durán.

On nonparaxial nonlinear Schrödinger-type equations. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719301025.

Cicone:2020:SBC

[CD20b] Antonio Cicone and Pietro Dell’Acqua. Study of boundary conditions in the Iterative Filtering method for the decomposition of nonstationary signals. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719302250.

Connors:2022:USH

[CD22] Jeffrey M. Connors and Robert D. Dolan. An unconditionally stable, highorder and flux-conservative fluid–fluid coupling method. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000620.

Cavoretto:2023:ARS

[CD23] Roberto Cavoretto and Alessandra De Rossi. An adaptive residual subsampling algorithm for kernel interpolation based on maximum likelihood estimations. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003338.

Casado-Diaz:2022:MPL

[CDCVV22] Juan Casado-Díaz, Carlos Conca, and Donato Vásquez-Varas. Minimization of the p-Laplacian first eigenvalue for a twophase material. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003447.

Chechkina:2020:OEE

[CDD20] Aleksandra G. Chechkina, Ciro D’Apice, and Umberto De Maio. Operator estimates for elliptic problem with rapidly alternating Steklov boundary condition. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300935.

Cavoretto:2024:NCS

[CDD^{+}24] Roberto Cavoretto, Alessandra De Rossi, Francesco Dell’Accio, Filomena Di Tommaso, Najoua Siar, Alvise Sommariva, and Marco Vianello. Numerical cubature on scattered data by adaptive interpolation. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000426.

Conley:2020:HMU

[CDJ20] Rebecca Conley, Tristan J.

Delaney, and Xiangmin Jiao. A hybrid method and unified analysis of generalized finite differences and Lagrange finite elements. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301539.

Cheng:2022:CGE

[CDL22] Hanz Martin Cheng, Jérôme Droniou, and Kim-Ngan Le. A combined GDM– ELLAM–MMOC scheme for advection dominated PDEs. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100501X.

Cavoretto:2024:BAR

[CDL24] Roberto Cavoretto, Alessandra De Rossi, and Sandro Lancellotti. Bayesian approach for radial kernel parameter tuning. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006593.

Cavoretto:2024:PTR

[CDLR24] Roberto Cavoretto, Alessandra De Rossi, Sandro Lancellotti, and Federico Romaniello. Parameter tuning in the radial kernelbased partition of unity method by Bayesian optimization. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003571.

Cornejo:2022:GQF

[CDMM22] M. Eugenia Cornejo, Juan Carlos Díaz-Moreno, and Jesús Medina. Generalized quantifiers in formal concept analysis. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003940.

Choi:2021:IGQ

[CDS21] Jaehyuk Choi, Yeda Du, and Qingshuo Song. Inverse Gaussian quadrature and finite normalmixture approximation of the generalized hyperbolic distribution. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305938.

Caravantes:2022:SIP

[CDTFGV22] J. Caravantes, G. M. Diaz-Toca, M. Fioravanti, and L. Gonzalez-Vega. Solving the interference problem for ellipses and ellipsoids: New formulae. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006300.

Cheng:2020:NDM

[CDW20] Jin-San Cheng, Xiaojie Dou, and Junyi Wen. A new deflation method for verifying the isolated singular zeros of polynomial systems. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301163.

Caucao:2024:AMF

[CE24] Sergio Caucao and Johann Esparza. An augmented mixed FEM for the convective Brinkman– Forchheimer problem: a priori and a posteriori error analysis. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004612.

Contreras-Espinoza:2023:GAS

[CECCCR23] Sergio Contreras-Espinoza, Christian Caamaño-Carrillo, and Javier E. Contreras-Reyes. Generalized autoregressive score models based on sinh-arcsinh distributions for time series analysis. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005738.

Chesneau:2020:NER

[CEKN20] Christophe Chesneau, Salima El Kolei, Junke Kou, and Fabien Navarro. Nonparametric estimation in a regression model with additive and multiplicative noise. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302624.

Cuomo:2023:KBM

[CES23] Salvatore Cuomo, Wolfgang Erb, and Gabriele Santin. Kernel-based models for influence maximization on graphs based on Gaussian process variance minimization. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005490.

Cerna:2020:GMN

[CF20] Dana Cerná and Václav Finek. Galerkin method with new quadratic spline wavelets for integral and integro-differential equations. Journal of Computational and Applied Mathematics, 363(??):426–443, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303140.

Cotrina:2024:ENE

[CFB24a] John Cotrina and Fabián Flores-Bazán. Existence of Nash equilibria for generalized multiobjective games through the vector extension of Weierstrass and Berge maximum theorems. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006635.

See note [CFB24b].

Cotrina:2024:NFI

[CFB24b] John Cotrina and Fabián Flores-Bazán. A note on the finite intersection property in “Existence of Nash equilibria for generalized multiobjective games through the vector extension of Weierstrass and Berge maximum theorems”. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002218.

See [CFB24a].

Colombo:2023:VIN

[CFdD23] Leonardo Colombo, Manuela Gamonal Fernández, and David Martín de Diego. Variational integrators for non-autonomous Lagrangian systems. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005647.

Canton:2021:CPA

[CFJVG21] A. Cantón, L. Fernández-Jambrina, and M. J. Vázquez-Gallo. Curvature of planar aesthetic curves. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303332.

Cococcioni:2021:NAZ

[CFL21] Marco Cococcioni, Lorenzo Fiaschi, and Luca Lambertini. Non-Archimedean zero-sum games. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001023.

Crawshaw:2022:SHD

[CFO22] Jessica R. Crawshaw, Jennifer A. Flegg, and James M. Osborne. A simple history-dependent remeshing technique to increase finite element model stability in elastic surface deformations. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004994.

Cerroni:2021:CPA

[CFS21] D. Cerroni, L. Formaggia, and A. Scotti. A control problem approach to Coulomb’s friction. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304878.

Chen:2020:CSB

[CG20] Ziheng Chen and Siqing Gan. Convergence and stability of the backward Euler method for jump-diffusion SDEs with super-linearly growing diffusion and jump coefficients. Journal of Computational and Applied Mathematics, 363(??):350–369, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719302912.

Chen:2022:NLS

[CG22] Rui Chen and Shuting Gu. On novel linear schemes for the Cahn–Hilliard equation based on an improved invariant energy quadratization approach. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001868.

Chen:2023:SCS

[CG23] Lin Chen and Siqing Gan. Strong convergence and stationary distribution of an explicit scheme for the Wright– Fisher model. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200615X.

Chacon-Gomez:2024:RSD

[CGCMRP24] Fernando Chacón-Gómez, M. Eugenia Cornejo, Jesús Medina, and Eloísa Ramírez-Poussa. Rough set decision algorithms for modeling with uncertainty. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003576.

Conchin-Gubernati:2022:MQS

[CGL22] Alice Conchin-Gubernati and Paola Lamberti. Multilevel quadratic spline integration. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006191.

Colmenares:2021:AAF

[CGM21] Eligio Colmenares, Gabriel N.

Gatica, and Willian Miranda. Analysis of an augmented fully-mixed finite element method for a bioconvective flows model. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001230.

Cerdan:2020:PRD

[CGMM20] J. Cerdán, D. Guerrero, J. Marín, and J. Mas. Preconditioners for rank deficient least squares problems. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306260.

Caucao:2024:TFM

[CGO24] Sergio Caucao, Gabriel N. Gatica, and Juan P. Ortega. A three-field mixed finite element method for the convective Brinkman– Forchheimer problem with varying porosity. Journal of Computational and Applied Mathematics, 451 (??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400339X.

Caboussat:2022:AAM

[CGP22] Alexandre Caboussat, Dimitrios Gourzoulidis, and Marco Picasso. An anisotropic adaptive method for the numerical approximation of orthogonal maps.

Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005896.

Cangiani:2020:CAD

[CGS20] Andrea Cangiani, Emmanuil H. Georgoulis, and Younis A. Sabawi. Convergence of an adaptive discontinuous Galerkin method for elliptic interface problems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304005.

Cantu-Gonzalez:2024:PLE

[CGSVBB^{+}24] José Roberto Cantú-González, Roberto Soto-Villalobos, Francisco Gerardo Benavides-Bravo, Ángela Gabriela Benavides-Ríos, Pablo Antonio Ramírez-Trejos, and Mario A. Aguirre-López. A power-law exponential model for variograms with quick transition and known range: Construction and application to geostatistical time series. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003497.

Chen:2021:FOS

[CGW21] Lin Chen, Siqing Gan, and Xiaojie Wang. First order strong convergence of an explicit scheme for the stochastic SIS epidemic model. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001011.

Chu:2020:HDF

[CH20] Van Tiep Chu and Viet Ha Hoang. High dimensional finite elements for twoscale Maxwell wave equations. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300479.

Chadjiconstantinidis:2024:MCS

[Cha24] Stathis Chadjiconstantinidis. On mixed censored δ-shock models. Journal of Computational and Applied Mathematics, 435 (??):??, ???? 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002121.

Cheng:2021:LDG

[Che21] Yao Cheng. On the local discontinuous Galerkin method for singularly perturbed problem with two parameters. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001047.

Chen:2022:LZP

[Che22] Haotian Chen. On locating the zeros and poles of a meromorphic function. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004180.

Chen:2020:FMG

[CHH20] Jian Chen, Minfan He, and Yong Huang. A fast multiscale Galerkin method for solving second order linear Fredholm integrodifferential equation with Dirichlet boundary conditions. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303553.

Chouly:2022:NMC

[CHLR22] Franz Chouly, Patrick Hild, Vanessa Lleras, and Yves Renard. Nitsche method for contact with Coulomb friction: Existence results for the static and dynamic finite element formulations. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002746.

Cao:2021:DMC

[CHLY21] Luling Cao, Yinnian He, Jian Li, and Di Yang. Decoupled modified characteristic FEMs for fully evolutionary Navier–Stokes– Darcy model with the Beavers–Joseph interface condition. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304192.

Choi:2024:FEC

[Cho24a] Gary P. T. Choi. Fast ellipsoidal conformal and quasi-conformal parameterization of genus-0 closed surfaces. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001377.

Choudhury:2024:NPA

[Cho24b]  Satwik Choudhury.  A

new parameter of assessment of linear iterative algorithms: On their informative efficiency and law of exponential-dissipation. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001742.

Chung:2021:CCG

[CHP21] Eric Chung, Jiuhua Hu, and Sai-Mang Pun. Convergence of the CEM-GMsFEM for Stokes flows in heterogeneous perforated domains. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030618X.

Cirillo:2020:CRH

[CHS20] Emiliano Cirillo, Kai Hormann, and Jean Sidon. Convergence rates of a Hermite generalization of Floater–Hormann interpolants. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306296.

Cuyt:2022:VAM

[CHsL22] Annie Cuyt, Yuan Hou, and Wen shin Lee. Validated analysis of modulated signals: From de Prony to Padé and beyond. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001534.

Chu:2024:NSS

[CHSZ24] Eric K.-W. Chu, Liangshao Hou, Daniel B. Szyld, and Jieyong Zhou. Numerical solution of singular Sylvester equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003709.

Cen:2020:MID

[CHXL20] Zhongdi Cen, Jian Huang, Aimin Xu, and Anbo Le. A modified integral discretization scheme for a two-point boundary value problem with a Caputo fractional derivative. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304686.

Chen:2020:RTP

[CHY20] Yaoyao Chen, Yunqing Huang, and Nianyu Yi. Recovery type a posteriori error estimation of adaptive finite element method for Allen–Cahn equation. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305795.

Cai:2021:GPS

[CHZ21] Zhiqiang Cai, Cuiyu He, and Shun Zhang. Generalized Prager–Synge identity and robust equilibrated error estimators for discontinuous elements. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002958.

Clavero:2023:MSM

[CJ23] C. Clavero and J. C. Jorge. A multi-splitting method to solve 2D parabolic reaction–diffusion singularly perturbed systems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722002801.

Calvo-Jurado:2022:JDL

[CJCDLL22] Carmen Calvo-Jurado, Juan Casado-Díaz, and Manuel Luna-Laynez. A justification of the Darcy law for a suspension of not self-similar solid particles non-periodically distributed. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000340.

Chai:2022:MMM

[CJL^{+}22] Xudong Chai, Tao Jiang, Linhong Li, Wei Xu, and Liwei Liu. On a manyto-many matched queueing system with flexible matching mechanism and impatient customers. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002825.

Chen:2024:CAS

[CJL^{+}24a] Wenbin Chen, Jianyu Jing, Qianqian Liu, Cheng Wang, and Xiaoming Wang. Convergence analysis of a second order numerical scheme for the Flory– Huggins–Cahn–Hilliard– Navier–Stokes system. Journal of Computational and Applied Mathematics, 450 (??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724002310.

Chi:2024:AES

[CJL^{+}24b] Baotao Chi, Zhichao Jia, Can Li, Qianjian Guo, Wei Yuan, and Chuanming Ju. An adaptive element subdivision method based on the affine transformations and partitioning techniques for evaluating the weakly singular integrals. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002649.

Chui:2021:AAS

[CJLL21] Charles K. Chui, Qingtang Jiang, Lin Li, and Jian Lu. Analysis of an adaptive short-time Fourier transform-based multicomponent signal separation method derived from linear chirp local approximation. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002272.

Chapko:2020:BIS

[CJM20] Roman Chapko, B. Tomas Johansson, and Leonidas Mindrinos. On a boundary integral solution of a lateral planar Cauchy problem in elastodynamics. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304662.

Castro:2020:SED

[CJP20] Mario H. Castro, Thaís Jordão, and Ana P. Peron. Super-exponential decay rates for eigenvalues and singular values of integral operators on the sphere. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303310.

Cheng:2020:MIN

[CK20] Dong Cheng and Kit Ian Kou. Multichannel interpolation of nonuniform samples with application to image recovery. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305059.

Cansiz:2022:CSF

[CK22] Baris Cansiz and Michael Kaliske. A comparative study of fully implicit staggered and monolithic solution methods. Part I: Coupled bidomain equations of cardiac electrophysiology. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721006014.

Chowdhury:2023:SSM

[CK23] Manisha Chowdhury and B. V. Rathish Kumar. Simplified subgrid multiscale stabilized finite element method in the transient framework for Stokes equations. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005544.

Cansiz:2024:CSF

[CK24a] Baris Cansiz and Michael Kaliske. A comparative study of fully implicit staggered and monolithic solution methods. Part II: Coupled excitation–contraction equations of cardiac electromechanics. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006477.

Cetin:2024:NRC

[ÇK24b] Arzu E. Çetin and Nursel Koyuncu. New robust class of estimators for population mean under different sampling designs. Journal of Computational and Applied Mathematics, 441 (??):??, May 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006131.

Chen:2021:KRC

[CKA21] C. S. Chen, Andreas Karageorghis, and Lionel Amuzu. Kansa RBF collocation method with auxiliary boundary centres for high order BVPs. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003022.

Chetverushkin:2020:CQG

[CKL20] B. N. Chetverushkin, Ya.V. Khankhasaeva, and A. E. Lutskii. Compact quasigas-dynamic system and high performance computing. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300832.

Chung:2023:LSF

[CKL23] Matthias Chung, Justin Krueger, and Honghu Liu. Least-squares finite element method for ordinary differential equations. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200334X.

Chernykh:2021:HHC

[CKT21] Igor Chernykh, Igor Kulikov, and Alexander Tutukov. Hydrogen-helium chemical and nuclear galaxy collision: Hydrodynamic simulations on AVX-512 supercomputers. Journal of Computational and Applied Mathematics, 391 (??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721000145.

Chen:2023:MPT

[CKY23] Yu Chen, Kai Kuang, and Xueling Yan. A modified PRP-type conjugate gradient algorithm with complexity analysis and its application to image restoration problems. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000493.

Cao:2021:PVS

[CKZ21] Jiling Cao, Jeong-Hoon Kim, and Wenjun Zhang. Pricing variance swaps under hybrid CEV and stochastic volatility. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305112.

Chen:2022:FVE

[CL22] Shuangshuang Chen and Xiaoli Li. Finite volume element methods for a multi-dimensional fracture model. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006051.

Cao:2024:RTD

[CL24a] K. Cao and D. Lesnic. Reconstruction of the timedependent source in thermal grooving by surface diffusion. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000384.

Chen:2024:UTB

[CL24b] Yuefen Chen and Bo Li. An uncertainty theory based tri-objective behavioral portfolio selection model with loss aversion and reference level using a modified evolutionary root system growth algorithm. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001080.

Cheng:2024:APF

[CL24c] Qingjin Cheng and Chunyan Luo. Analytical properties, fractal dimensions and related inequalities of (k, h)-Riemann– Liouville fractional integrals. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002498.

Chen:2022:EAS

[CLH22] Yanping Chen, Xiuxiu Lin, and Yunqing Huang. Error analysis of spectral approximation for spacetime fractional optimal control problems with control and state constraints. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001200.

Chen:2020:SEA

[CLHL20] Yanping Chen, Xiuxiu Lin, Yunqing Huang, and Qian Lin. hp spectral element approximation for integral state constrained optimal control problems governed by harmonic equations. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300078.

Cao:2020:SRS

[CLL20a] K. Cao, D. Lesnic, and Jijun Liu. Simultaneous reconstruction of spacedependent heat transfer coefficients and initial temperature. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300911.

Cui:2020:HOC

[CLL20b] Ming Cui, Fangxia Li, and Dong Liang. High-order characteristic-finite volume methods for aerosol dynamic equations. Journal of Computational and Applied Mathematics, 370 (??):??, May 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305989.

Cao:2021:SCE

[CLL21] Wanrong Cao, Jia Liang, and Yufen Liu. On strong convergence of explicit numerical methods for stochastic delay differential equations under non-global Lipschitz conditions. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303708.

Civiletti:2020:ARC

[CLM20] Benjamin J. Civiletti, Akhlesh Lakhtakia, and Peter B. Monk. Analysis of the Rigorous Coupled Wave Approach for s-polarized light in gratings. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304819.

Civiletti:2021:ARC

[CLM21] Benjamin J. Civiletti, Akhlesh Lakhtakia, and Peter B. Monk. Analysis of the rigorous coupled wave approach for p-polarized light in gratings. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305264.

Chen:2022:LLP

[CLM22a] Yinjun Chen, Huilan Liu, and Junjie Ma. Local least product relative error estimation for single-index varying-coefficient multiplicative model with positive responses. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002308.

Civiletti:2022:HRC

[CLM22b] B. J. Civiletti, A. Lakhtakia, and P. B. Monk. Hybridization of the rigorous coupled-wave approach with transformation optics for electromagnetic scattering by a surface-relief grating. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001480.

Celledoni:2023:LHC

[CLMO23] Elena Celledoni, Andrea Leone, Davide Murari, and Brynjulf Owren. Learning Hamiltonians of constrained mechanical systems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200303X.

Chung:2021:MSD

[CLPZ21] Eric Chung, Wing Tat Leung, Sai-Mang Pun, and Zecheng Zhang. A multi-stage deep learning based algorithm for multiscale model reduction. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001254.

Chen:2020:BPP

[CLS20] Zheshi Chen, Chunhong Li, and Wenjun Sun. Bitcoin price prediction using machine learning: an approach to sample dimension engineering. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930398X.

Cheng:2021:GSA

[CLS21] Qing Cheng, Chun Liu, and Jie Shen. Generalized SAV approaches for gradient systems. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001515.

Chen:2024:NAF

[CLS24] Jiahao Chen, Xiaofei Li, and Yunze Shao. Numerical analysis of fractional order Black–Scholes option pricing model with band equation method. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002486.

Catral:2020:SSP

[CLST20] M. Catral, L. Lebtahi, J. Stuart, and N. Thome. Spectral study of {R, s + 1, k}- and {R, s + 1, k, ∗}- potent matrices. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304170.

Conti:2023:NOB

[CLU23] Costanza Conti and Sergio López-Ureña. Nonoscillatory butterfly-type interpolation on triangular meshes. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004010.

Chang:2022:FOR

[CLW^{+}22] Leiya Chang, Xiufang Liu, Dehui Wang, Yingchuan Jing, and Chenlong Li. First-order random coefficient mixed-thinning integer-valued autoregressive model. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000838.

Cen:2021:SOA

[CLX21] Zhongdi Cen, Li-Bin Liu, and Aimin Xu. A secondorder adaptive grid method for a nonlinear singularly perturbed problem with an integral boundary condition. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304969.

Chen:2024:USH

[CLYS24] Xin Chen, Ziqiang Lu, Dongmei Yuan, and Yu Shao. Uncertain stochastic hybrid zero-sum games based on forward uncertain difference equations and backward stochastic difference equations. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001444.

Cho:2020:MHD

[CM20] Kanghun Cho and Minam Moon. Multiscale hybridizable discontinuous Galerkin method for elliptic problems in perforated domains. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303437.

Cardoso:2022:SDE

[CM22] João R. Cardoso and Pedro Miraldo. Solving the discrete Euler–Arnold equations for the generalized rigid body motion. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004362.

Caeiro:2024:RBE

[CM24a] Frederico Caeiro and Ayana Mateus. Reduced bias estimation of the shape parameter of the log-logistic distribution. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002911.

Chen:2024:SRL

[CM24b] Liang Chen and Junjie Ma.

Solve Riemann–Liouville boundary value problems using collocation boundary value methods with the graded mesh. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000116.

Castro-Martin:2022:CSM

[CMdMRFG22] Luis Castro-Martín, María del Mar Rueda, and Ramón Ferri-García. Combining statistical matching and propensity score adjustment for inference from non-probability surveys. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000339.

Chavarria-Molina:2022:EIR

[CMFMSQ22] Jeffry Chavarría-Molina, Juan José Fallas-Monge, and Pablo Soto-Quiros. Effective implementation to reduce execution time of a low-rank matrix approximation problem. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100385X.

Coelho:2022:DLR

[CMJN22] Carlos A. Coelho, Filipe J.

Marques, Nadab Jorge, and Célia Nunes. On the distribution of the likelihood ratio test of independence for random sample size — a computational approach. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000133.

Calvino:2021:RMD

[CMP21] Aida Calviño, Nirian Martín, and Leandro Pardo. Robustness of Minimum Density Power Divergence Estimators and Wald-type test statistics in loglinear models with multinomial sampling. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305057.

Costabile:2021:LAE

[CMRS21] Massimo Costabile, Ivar Massabó, Emilio Russo, and Alessandro Staino. A lattice approach to evaluate participating policies in a stochastic interest rate framework. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305033.

Chen:2020:EEG

[CMS20a] Hongtao Chen, Jingjing Mao, and Jie Shen. Error estimate of gauge–Uzawa methods for incompressible flows with variable density. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303188.

Cheng:2020:ENS

[CMS20b] Ting Cheng, Lina Ma, and Jie Shen. An efficient numerical scheme for a 3D spherical dynamo equation. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306338.

Camps:2020:SNB

[CMVV20] Daan Camps, Nicola Mastronardi, Raf Vandebril, and Paul Van Dooren. Swapping 2 × 2 blocks in the Schur and generalized Schur form. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719302572.

Cai:2024:ANS

[CMW24] Yongmei Cai, Xuerong Mao, and Fengying Wei. An advanced numerical scheme for multidimensional stochastic Kolmogorov equations with superlinear coefficients. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004168.

Chen:2022:SHM

[CMZ22] Gang Chen, Peter Monk, and Yangwen Zhang. Superconvergent HDG methods for Maxwell’s equations via the M-decomposition.

Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721004118.

Cui:2024:QHC

[CN24] Bing Cui and Alireza Najafi. Quantile hedging in the complete financial market under the mixed fractional Brownian motion model and the liquidity constraint. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000864.

Cortes:2022:SFR

[CNQRR22] J.-C. Cortés, A. Navarro-Quiles, J.-V. Romero, and M.-D. Roselló. Solving fully randomized firstorder linear control systems: Application to study the dynamics of a damped oscillator with parametric noise under stochastic control. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100008X.

Concas:2020:SMB

[CNRR20] A. Concas, S. Noschese, L. Reichel, and G. Rodriguez. A spectral method for bipartizing a network and detecting a large anticommunity. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930295X.

Chen:2021:IPU

[CNS21] An Chen, Thai Nguyen, and Nils Sørensen. Indifference pricing under SA-HARA utility. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305793.

Chen:2023:DLS

[CNYC23] C. S. Chen, Amir Noorizadegan, D. L. Young, and Chuin-Shan Chen. On the determination of locating the source points of the MFS using effective condition number. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005532.

Chernyshenko:2019:UFE

[CO19] Alexey Y. Chernyshenko and Maxim A. Olshanskii. An unfitted finite element method for the Darcy problem in a fracture network. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304273.

Colakoglu:2023:NPC

[ÇÖÖ23] H. Baris Çolakoglu, Iskender Öztürk, and Mustafa Özdemir. Non-parabolic conical rotations. Journal of Computational and Applied Mathematics, 420 (??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722003880.

Chetverushkin:2021:NSH

[COT21] Boris N. Chetverushkin, Olga G. Olkhovskaya, and Il’ya P. Tsigvintsev. Numerical solution of hightemperature gas dynamics problems on highperformance computing systems. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306658.

Cheung:2023:CAL

[CPBG23] James Cheung, M. Perego, P. Bochev, and M. Gunzburger. A coupling approach for linear elasticity problems with spatially non-coincident discretized interfaces. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006252.

Conte:2023:NFD

[CPP23] Dajana Conte, Giovanni Pagano, and Beatrice Paternoster. Nonstandard finite differences numerical methods for a vegetation reaction–diffusion model. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004022.

Cornilly:2022:SSP

[CPRV22] Dries Cornilly, Giovanni Puccetti, Ludger Rüschendorf, and Steven Vanduffel. On some synchronization problems with multiple instances. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003198.

Cuomo:2020:RMS

[CPS20] S. Cuomo, F. Piccialli, and F. Sica. RBF methods in a stochastic volatility framework for Greeks computation. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302788.

Cai:2020:EFR

[CQBL20] Xin Cai, Yufeng Qian, Qingshan Bai, and Wei Liu. Exploration on the financing risks of enterprise supply chain using Back Propagation neural network. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304601.

Cheng:2022:CEC

[CQZ22] Xiujun Cheng, Hongyu Qin, and Jiwei Zhang. Convergence of an energyconserving scheme for nonlinear space fractional Schrödinger equations with wave operator. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003848.

Cesmelioglu:2023:HDG

[CR23] Aycil Cesmelioglu and Sander Rhebergen. A hybridizable discontinuous Galerkin method for the coupled Navier–Stokes and Darcy problem. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005210.

Cocquet:2021:EAF

[CRRB21] Pierre-Henri Cocquet, Michaël Rakotobe, Delphine Ramalingom, and Alain Bastide. Error analysis for the finite element approximation of the Darcy–Brinkman– Forchheimer model for porous media with mixed boundary conditions. Journal of Computational and Applied Mathematics, 381 (??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720302995.

Cesmelioglu:2020:EHD

[CRW20] Aycil Cesmelioglu, Sander Rhebergen, and Garth N. Wells. An embeddedhybridized discontinuous Galerkin method for the coupled Stokes–Darcy system. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304790.

Chen:2021:NPP

[CS21] Huangxin Chen and Shuyu Sun. A new physicspreserving IMPES scheme for incompressible and immiscible two-phase flow in heterogeneous porous media. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303265.

Cosgun:2023:NMI

[CS23] Tahir Cosgun and Murat Sari. A novel method to investigate nonlinear advection–diffusion processes. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000018.

Chakir:2024:NST

[CS24] Yassine Chakir and Hassan Safouhi. Numerical solution of two-dimensional weakly singular Volterra integral equations of the first kind via bivariate rational approximants. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003229.

Chakraborty:2022:VLE

[CSLQ22] Suman Chakraborty, Yixuan Sun, Guang Lin, and Li Qiao. Vapor-liquid equilibrium estimation of nalkane/nitrogen mixtures using neural networks. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006208.

Csomos:2020:MTI

[Cso20] Petra Csomós. Magnustype integrator for semilinear delay equations with an application to epidemic models. Journal of Computational and Applied Mathematics, 363(??):92–105, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302663.

Clavero:2024:NAT

[CSS24] Carmelo Clavero, Ram Shiromani, and Vembu Shanthi. A numerical approach for a two-parameter singularly perturbed weaklycoupled system of 2-D elliptic convection–reaction– diffusion PDEs. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003667.

Conte:2019:NFL

[CST19] D. Conte, S. Shahmorad, and Y. Talaei. New fractional Lanczos vector polynomials and their application to system of Abel– Volterra integral equations and fractional differential equations. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304121.

Cuomo:2020:GCO

[CST20] Salvatore Cuomo, Federica Sica, and Gerardo Toraldo. Greeks computation in the option pricing problem by means of RBF–PU methods. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301734.

Chang:2020:MPP

[CSZP20] Jinhua Chang, Lin Sun, Bo Zhang, and Jin Peng. Multi-period portfolio selection with mental accounts and realistic constraints based on uncertainty theory. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301837.

Chau:2020:MTA

[CT20] Ngo Lam Xuan Chau and Ha Trong Thi. Möbius transformations on algebraic ODEs of order one and algebraic general solutions of the autonomous equivalence classes. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302909.

Capilla:2020:VST

[CTGV20] M. T. Capilla, C. F. Talavera, D. Ginestar, and G. Verdú. Validation of the SHNC timedependent transport code based on the spherical harmonics method for complex nuclear fuel assemblies. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301059.

Chand:2022:CSF

[CTN22] A. K. B. Chand, K. R.

Tyada, and M. A. Navascués. Cubic spline fractal solutions of two-point boundary value problems with a non-homogeneous nowhere differentiable term. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305586.

Cengizci:2024:SBS

[CUN24] Süleyman Cengizci, Ömür Ugur, and Srinivasan Natesan. SUPG-based stabilized finite element computations of convectiondominated 3D elliptic PDEs using shock-capturing. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002723.

Coria:2024:EAM

[CUUdB24] Ibai Coria, Gorka Urkullu, Haritz Uriarte, and Igor Fernández de Bustos. An efficient algorithm for the minimal least squares solution of linear systems with indefinite symmetric matrices. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000323.

Ciegis:2020:HON

[CV20] Raimondas Ciegis and Petr N. Vabishchevich. High order numerical schemes for solving fractional powers of elliptic operators. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306326.

Chen:2024:NSA

[CV24] Kewang Chen and Cornelis Vuik. Non-stationary Anderson acceleration with optimized damping. Journal of Computational and Applied Mathematics, 451 (??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042724003236.

Coelho:2022:NSS

[CVPM22] D. L. Coelho, E. Vitral, J. Pontes, and N. Mangiavacchi. Numerical scheme for solving the nonuniformly forced cubic and quintic Swift– Hohenberg equations strictly respecting the Lyapunov functional. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005938.

Cao:2020:IMI

[CW20] Haitao Cao and Haijun Wu. IPCDGM and multiscale IPDPGM for the Helmholtz problem with large wave number. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305953.

Cai:2021:CAP

[CW21a] Shuting Cai and Yoshitaka Watanabe. Computerassisted proofs of the existence of a symmetrybreaking bifurcation point for the Kolmogorov problem. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002235.

Chen:2021:MDS

[CW21b] Hao Chen and Ying Wang. Modified Douglas splitting method for differential matrix equations. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304532.

Chang:2022:GTB

[CW22] Shih Yu Chang and Yimin Wei. General tail bounds for random tensors summation: Majorization approach. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200262X.

Cheng:2024:FSC

[CW24] Xiaoyu Cheng and Lizhen Wang. Fundamental solutions and conservation laws for conformable time fractional partial differential equation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003783.

Chen:2022:SOF

[CWLM22] Long Chen, Kai Wang, Baotong Li, and Yicheng Mao. Shape optimization of fluid cooling channel based on Darcy reducedorder isogeometric analysis. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001078.

Carlson:2023:APT

[CWT23] Max Carlson, Jerry Watkins, and Irina Tezaur. Automatic performance tuning for Albany Land Ice. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001668.

Cao:2023:PDWa

[CWW23a] Waixiang Cao, Chunmei Wang, and Junping Wang. An L^{p}–primal–dual weak Galerkin method for convection–diffusion equations. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003533.

Cao:2023:PDW

[CWW23b] Waixiang Cao, Chunmei Wang, and Junping Wang. An L^{p}–primal-dual weak Galerkin method for div– curl systems. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004794.

Che:2021:RAL

[CWY21] Maolin Che, Yimin Wei, and Hong Yan. Randomized algorithms for the low multilinear rank approximations of tensors. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306713.

Chadjiconstantinidis:2024:IBT

[CX24] Stathis Chadjiconstantinidis and Panos Xenos. Improved bounds on tails of convolutions of compound distributions: Application to ruin probabilities for the risk process perturbed by diffusion. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000840.

Chao:2020:PNI

[CXS20] Zhen Chao, Dexuan Xie, and Ahmed H. Sameh. Preconditioners for nonsymmetric indefinite linear systems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930439X.

Cheung:2020:UVI

[CY20] K. C. Cheung and F. L.

Yuen. On the uncertainty of VaR of individual risk. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304716.

Chen:2021:AEQ

[CYC21] Ruyun Chen, Di Yu, and Juan Chen. Asymptotic expansion and quadrature rule for a class of singular-oscillatory-Bessel-type transforms. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304325.

Chu:2023:BSM

[CYY23] Yuanqi Chu, Zhouping Yin, and Keming Yu. Bayesian scale mixtures of normals linear regression and Bayesian quantile regression with big data and variable selection. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300136X.

Cai:2020:ECR

[CZ20] Shousong Cai and Jing Zhang. Exploration of credit risk of P2P platform based on data mining technology. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300091.

Chang:2023:ACR

[CZ23a] Jingya Chang and Zhi Zhu. An adaptive cubic regularization algorithm for computing H- and Zeigenvalues of real evenorder supersymmetric tensors. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001395.

Cheung:2023:ERN

[CZ23b] Lawrence C. Cheung and Tamer A. Zaki. An eigen-representation of the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722005192.

Cui:2024:CDE

[CZ24] Xingbang Cui and Liping Zhang. Computing the dominant eigenpair of an essentially nonnegative tensor via a homotopy method. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005095.

Chen:2021:SPH

[CZH21] Yangyang Chen, Yi Zhao, and Xinyu Han. Spectral properties of hypercubes with applications. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001709.

Cao:2023:OEA

[CZHQ23] Y. Cao, M. A. Zaky, A. S. Hendy, and W. Qiu. Optimal error analysis of spacetime second-order difference scheme for semilinear non-local Sobolevtype equations with weakly singular kernel. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002315.

Chen:2023:PLR

[CZM23] Ruoning Chen, Liping Zhang, and Zhenyu Ming. A parallel low rank matrix optimization method for recovering internet traffic network data via link flow measurement. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002753.

Cao:2022:FDS

[CZY22] Junying Cao, Jun Zhang, and Xiaofeng Yang. Fullydiscrete spectral-Galerkin scheme with second-order time-accuracy and unconditionally energy stability for the volume-conserved phase-field lipid vesicle model. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005835.

deAlmagro:2023:HOI

[dA23] Rodrigo T. Sato Martín de Almagro. High-order integrators for Lagrangian systems on homogeneous spaces via nonholonomic mechanics. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004356.

Dadashi:2020:OIS

[Dad20] Hassan Dadashi. Optimal investment strategy post retirement without ruin possibility: a numerical algorithm. Journal of Computational and Applied Mathematics, 363(??):325– 336, January 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0377042719301001.

Duque:2023:NSC

[DAM23] José C. M. Duque, Rui M. P. Almeida, and Belchior C. X. Mário. Numerical solution for a class of evolution differential equations with p-Laplacian and memory. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000882.

Deepho:2022:SUO

[DAMA22] Jitsupa Deepho, Auwal Bala Abubakar, Maulana Malik, and Ioannis K. Argyros. Solving unconstrained optimization problems via hybrid CD–DY conjugate gradient methods with applications. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004453.

Dareiotis:2020:FDS

[Dar20] Konstantinos Dareiotis.

On finite difference schemes for partial integro-differential equations of Lévy type. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719305928.

Dax:2020:CPA

[Dax20] Achiya Dax. A crossproduct approach for lowrank approximations of large matrices. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305813.

Dax:2024:CHI

[Dax24]  Achiya Dax.  A com-

pact Heart iteration for low-rank approximations of large matrices. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004119.

Doostmoradi:2023:RAN

[DAZ23] Ali Doostmoradi, Mohammad Reza Akhoond, and Mohammad Reza Zadkarami. Reliability analysis of a new family mixed shock models for the multistate systems. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003272.

Dey:2024:HCR

[DBR24] Anwesa Dey, Alfio Borzì, and Souvik Roy. A high contrast and resolution reconstruction algorithm in quantitative photoacoustic tomography. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003157.

Dargaville:2021:CEA

[DBSS^{+}21] S. Dargaville, A. G.

Buchan, R. P. Smedley-Stevenson, P. N. Smith, and C. C. Pain. A comparison of element agglomeration algorithms for unstructured geometric multigrid. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306701.

deCamargo:2020:NSN

[dC20a] André Pierro de Camargo.

On the numerical stability of Newton’s formula for Lagrange interpolation. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303723.

Deng:2020:HOS

[DC20b] Quanling Deng and Victor Calo. Higher order stable generalized finite element method for the elliptic eigenvalue and source problems with an interface in 1D. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305631.

Dolejsi:2023:GOE

[DC23] Vít Dolejsí and Scott Congreve. Goal-oriented error analysis of iterative Galerkin discretizations for nonlinear problems including linearization and algebraic errors. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300078X.

Deng:2024:SCT

[DC24] Nan Deng and Wanrong Cao. On strong convergence of two numerical methods for singular initial value problems with multiplicative white noise. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000505.

Durmaz:2022:NSS

[DCAA22] Muhammet Enes Durmaz, Musa Cakir, Ilhame Amirali, and Gabil M. Amiraliyev. Numerical solution of singularly perturbed Fredholm integrodifferential equations by homogeneous second order difference method. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200139X.

Deng:2022:NMS

[DCP22] Nan Deng, Wanrong Cao, and Guofei Pang. On numerical methods to secondorder singular initial value problems with additive white noise. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002655.

Dai:2024:PEE

[DCY24] Jiajia Dai, Luoping Chen, and Miao Yang. A posteriori error estimates of the weak Galerkin finite element methods for parabolic problems. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000712.

DAmbrosio:2023:NCI

[DD23] Raffaele D’Ambrosio and Stefano Di Giovacchino. Numerical conservation issues for the stochastic Korteweg–de Vries equation. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005659.

Debarre:2023:TBS

[DDF23] Thomas Debarre, Quentin Denoyelle, and Julien Fageot. TV-based spline reconstruction with Fourier measurements: Uniqueness and convergence of gridbased methods. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005350.

DellAccio:2024:MSM

[DDF24] Francesco Dell’Accio, Filomena Di Tommaso, and Elisa Francomano. Multinode Shepard method for two-dimensional elliptic boundary problems on different shaped domains. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001468.

Dudin:2019:EAC

[DDKD19] Sergei Dudin, Alexander Dudin, Olga Kostyukova, and Olga Dudina. Effective algorithm for computation of the stationary distribution of multi-dimensional level-dependent Markov chains with upper block-Hessenberg structure of the generator. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304285.

DeAngelis:2020:NNM

[DDM20] Paolo De Angelis, Roberto De Marchis, and Antonio Luciano Martire. A new numerical method for a class of Volterra and Fredholm integral equations. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302351.

DellAccio:2021:FAS

[DDNZ21] Francesco Dell’Accio, Filomena Di Tommaso, Otheman Nouisser, and Benaissa Zerroudi. Fast and accurate scattered Hermite interpolation by triangular Shepard operators. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303836.

Deif:2022:SCF

[DdO22] Sarah A. Deif and E. Capelas de Oliveira. A system of Cauchy fractional differential equations and new properties of Mittag-Leffler functions with matrix argument. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005744.

Djukic:2021:IGA

[DDRS21] Dusan Lj. Djukić, Rada M. Mutavdzić Djukić, Lothar Reichel, and Miodrag M. Spalević. Internality of generalized averaged Gauss quadrature rules and truncated variants for modified Chebyshev measures of the first kind. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003186.

Djukic:2024:DOA

[DDRS24] Dusan Lj. Djukić, Rada M. Mutavdzić Djukić, Lothar Reichel, and Miodrag M. Spalević. Decompositions of optimal averaged Gauss quadrature rules. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005307.

Dlugosz:2023:NFC

[DDT23] Zofia Długosz, Rafał Długosz, and Tomasz Talaśka. A novel, flexible circuit used to implement selected mathematical operations for AI algorithms optimized for hardware applications. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001048.

Debarre:2022:SPL

[DDUF22] Thomas Debarre, Quentin Denoyelle, Michael Unser, and Julien Fageot. Sparsest piecewise-linear regression of one-dimensional data. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006130.

Dick:2024:FRQ

[DEH^{+}24] Josef Dick, Adrian Ebert, Lukas Herrmann, Peter Kritzer, and Marcello Longo. The fast reduced QMC matrix-vector product. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005861.

Denich:2024:EEG

[Den24] Eleonora Denich. Error estimates for a Gaussian rule involving Bessel functions. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003928.

daFonseca:2020:CST

[dF20] Carlos M. da Fonseca.

Comments on the spectrum of a tridiagonal k-Toeplitz matrix. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300844.

Deng:2021:TES

[DFFM21] Shounian Deng, Chen Fei, Weiyin Fei, and Xuerong Mao. Tamed EM schemes for neutral stochastic differential delay equations with superlinear diffusion coefficients. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305604.

deFrutos:2021:PMO

[dFG21] Javier de Frutos and Víctor Gatón. A pseudospectral method for option pricing with transaction costs under exponential utility. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001606.

Debbouche:2021:SIL

[DFH21] Amar Debbouche, Michal Feckan, and Eduardo Hernández. Special issue: Latest computational methods on fractional dynamic systems “VSI fractional dynamic system”. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306130.

Dyn:2020:AOP

[DFK20] Nira Dyn, Elza Farkhi, and Shirley Keinan. Approximation of 3D objects by piecewise linear geometric interpolants of their 1D cross-sections. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304698.

Pietro:2020:MMS

[DFMK20] Franco Di Pietro, Joaquín Fernández, Gustavo Migoni, and Ernesto Kofman. Mixed-mode state-time discretization in ODE numerical integration. Journal of Computational and Applied Mathematics, 377 (??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042720302028.

Dembinska:2020:MOS

[DG20] Anna Dembińska and Agnieszka Goroncy. Moments of order statistics from DNID discrete random variables with application in reliability. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719307083.

Dick:2021:SLR

[DG21] Josef Dick and Takashi Goda. Stability of lattice rules and polynomial lattice rules constructed by the componentby-component algorithm. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303538.

Dehghan:2022:AWG

[DG22a] Mehdi Dehghan and Zeinab Gharibi. An analysis of weak Galerkin finite element method for a steady state Boussinesq problem. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100604X.

Doghman:2022:NCA

[DG22b] Jad Doghman and Ludovic Goudenège. Numerical and convergence analysis of the stochastic Lagrangian averaged Navier–Stokes equations. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002114.

Dattoli:2024:HHO

[DGL24] Giuseppe Dattoli, Roberto Garra, and Silvia Licciardi. Hermite, higher order Hermite, Laguerre type polynomials and Burgers like equations. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000700.

Dang:2020:PGS

[DH20] Quang A. Dang and Manh Tuan Hoang. Positivity and global stability preserving NSFD schemes for a mixing propagation model of computer viruses. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300443.

Das:2022:CBD

[DH22] Pankaj Kumar Das and Letminthang Haokip. Codes with burst distance and periodical burst errors. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000929.

Deng:2024:MSI

[DH24a] Zhao Deng and Deren Han. Multi-step inertial strictly contractive PRSM algorithms for convex programming problems with applications. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004132.

Ding:2024:SRN

[DH24b] Liang Ding and Weimin Han. αl_{1} −βl_{2} sparsity regularization for nonlinear illposed problems. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002371.

Deng:2023:FIE

[DHF23] Lanmei Deng, Rong Hu, and Ya-Ping Fang. Fast inertial extragradient algorithms for solving non-Lipschitzian equilibrium problems without monotonicity condition in real Hilbert spaces. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005489.

Ding:2024:OEE

[DHH24] Qianqian Ding, Yuanyuan Hou, and Xiaoming He. Optimal error estimates of a second-order fully decoupled finite element method for the nonstationary generalized Boussinesq model. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002516.

Duan:2024:CDI

[DHJ^{+}24] Chenguang Duan, Junjun Huang, Yuling Jiao, Xiliang Lu, and Jerry Zhijian Yang. Current density impedance imaging with PINNs. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003698.

Doan:2020:EMS

[DHKV20] T. S. Doan, P. T. Huong, P. E. Kloeden, and A. M. Vu. Euler–Maruyama scheme for Caputo stochastic fractional differential equations. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302806.

Dong:2024:FEI

[DHLT24] Xiaojing Dong, Yunqing Huang, Meiyun Liu, and Qili Tang. Finite element iterative algorithm based on Anderson acceleration technique for incompressible MHD equations. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001808.

Dong:2024:PPL

[DHQZ24] Lixiu Dong, Dongdong He, Yuzhe Qin, and Zhengru Zhang. A positivitypreserving, linear, energy stable and convergent numerical scheme for the Poisson–Nernst–Planck (PNP) system. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000335.

deHierro:2022:MCD

[dHSR22] A. F. Roldán López de Hierro, M. Sánchez, and C. Roldán. Multi-criteria decision making involving uncertain information via fuzzy ranking and fuzzy aggregation functions. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304295.

Di:2024:LFN

[DHZ24] Yana Di, Wenlong He, and Jiwei Zhang. A lockingfree numerical method for the quasi-static linear thermo-poroelasticity model. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003467.

Defez:2022:BMP

[DIAJ^{+}22] E. Defez, J. Ibáñez, P. Alonso-Jordá, José M. Alonso, and J. Peinado. On Bernoulli matrix polynomials and matrix exponential approximation. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304982.

Ding:2023:LSP

[Din23] Feng Ding. Least squares parameter estimation and multi-innovation least squares methods for linear fitting problems from noisy data. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000511.

Defez:2022:NHS

[DIP^{+}22] E. Defez, J. Ibáñez, J. Peinado, P. Alonso-Jordá, and José M. Alonso. New Hermite series expansion for computing the matrix hyperbolic cosine. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000024.

Dey:2022:IPW

[DJ22a] Santanu Dey and Nabakumar Jana. Inference on parameters of Watson distributions and application to classification of observations. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004702.

Ding:2022:RAS

[DJ22b] Xiao-Li Ding and Yao-Lin Jiang. Regularity analysis for SVEEs with additive fBms and strong error estimates for the numerical approximations. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002278.

Dembinska:2024:LIG

[DJ24] Anna Dembińska and Krzysztof Jasiński. Likelihood inference for geometric lifetimes of components of k-out-of-n systems. Journal of Computational and Applied Mathematics, 435(??):??, ???? 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300211X.

DelaCruzCabrera:2021:NMM

[DJR21] Omar De la Cruz Cabrera, Jiafeng Jin, and Lothar Reichel. New models for multi-class networks. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001874.

Djukic:2024:OPR

[Dju24] Dusan Lj. Djukić. Orthogonal polynomials with respect to the Abel weight. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004296.

Deep:2024:SNL

[DK24] Amar Deep and Manochehr Kazemi. Solvability for 2D non-linear fractional integral equations by Petryshyn’s fixed point theorem. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000463.

Dhaene:2020:CAP

[DKL20] Jan Dhaene, Alexander Kukush, and Daniël Linders. Comonotonic asset prices in arbitragefree markets. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303073.

Durrwachter:2020:HPD

[DKM^{+}20] Jakob Dürrwächter, Thomas Kuhn, Fabian Meyer, Louisa Schlachter, and Florian Schneider. A hyperbolicity-preserving discontinuous stochastic Galerkin scheme for uncertain hyperbolic systems of equations. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306077.

DApice:2023:TLV

[DKM23a] Ciro D’Apice, Peter I. Kogut, and Rosanna Manzo. A two-level variational algorithm in the Sobolev– Orlicz space to predict daily surface reflectance at LANDSAT high spatial resolution and MODIS temporal frequency. Journal of Computational and Applied Mathematics, 434 (??):??, December 15, 2023. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723002832.

Delgado:2023:BDV

[DKM^{+}23b] Jorge Delgado, Plamen Koev, Ana Marco, José-Javier Martínez, Juan Manuel Peña, Per-Olof Persson, and Steven Spasov. Bidiagonal decompositions of Vandermonde-type matrices of arbitrary rank. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000080.

Donatelli:2024:MTS

[DKMT24] Marco Donatelli, Rolf Krause, Mariarosa Mazza, and Ken Trotti. Multigrid for two-sided fractional differential equations discretized by finite volume elements on graded meshes. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000360.

Daoui:2022:SAL

[DKSQ22] Achraf Daoui, Hicham Karmouni, Mhamed Sayyouri, and Hassan Qjidaa. Stable analysis of largesize signals and images by Racah’s discrete orthogonal moments. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004520.

Deng:2020:IPS

[DL20a] Zhao Deng and Sanyang Liu. Inertial proximal strictly contractive Peaceman–Rachford splitting method with an indefinite term for convex optimization. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300637.

Dong:2020:RSO

[DL20b] Jian Dong and Ding Fang Li. A reliable secondorder hydrostatic reconstruction for shallow water flows with the friction term and the bed source term. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030162X.

Dong:2020:CCN

[DL20c] Liqiang Dong and Jicheng Li. A class of complex nonsymmetric algebraic Riccati equations associated with H-matrix. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305722.

Dastour:2021:GOF

[DL21] Hatef Dastour and Wenyuan Liao. A generalized optimal fourth-order finite difference scheme for a 2D Helmholtz equation with the perfectly matched layer boundary condition. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001631.

deLeon:2023:DEC

[dL23a] José Alfredo Sánchez de León.

Determination of eddy currents equipment parame-

ters: [cutting tolerance],

[cutting position]. Jour-

nal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006215.

Deng:2023:HOS

[DL23b] Dingwen Deng and Zhijun Li. High-order structure-preserving Du Fort–Frankel schemes and their analyses for the nonlinear Schrödinger equation with wave operator. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003077.

Deng:2021:PMD

[DLC21] Xiaomao Deng, Zi-Ju Liao, and Xiao-Chuan Cai. A parallel multilevel domain decomposition method for source identification problems governed by elliptic equations. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000601.

Ysern:2022:CPT

[dlCYS22] Bernardo de la Calle Ysern and Miodrag M. Spalević. On the computation of Patterson-type quadrature rules. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004738.

Du:2021:AIR

[DLL21] Guansan Du, Zixian Liu, and Haifeng Lu. Application of innovative risk early warning mode under big data technology in Internet credit financial risk assessment. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305513.

deLoynes:2021:DDT

[dLNO21] Basile de Loynes, Fabien Navarro, and Baptiste Olivier. Data-driven thresholding in denoising with Spectral Graph Wavelet Transform. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306105.

Deelstra:2020:BOP

[DLS20] Griselda Deelstra, Guy Latouche, and Matthieu Simon. On barrier option pricing by Erlangization in a regime-switching model with jumps. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306119.

Du:2022:RBP

[DLZ22] Shaohong Du, Runchang Lin, and Zhimin Zhang. Residual-based a posteriori error estimators for mixed finite element methods for fourth order elliptic singularly perturbed problems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001376.

DellAccio:2024:EMI

[DMN24] Francesco Dell’Accio, Francisco Marcellán, and Federico Nudo. An extension of a mixed interpolation– regression method using zeros of orthogonal polynomials. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002607.

DellAccio:2024:NAF

[DMNO24] Francesco Dell’Accio, Domenico Mezzanotte, Federico Nudo, and Donatella Occorsio. Numerical approximation of Fredholm integral equation by the constrained mock-Chebyshev least squares operator. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001353.

Dai:2023:ABI

[DMP23] Suhang Dai and Olivier Menoukeu-Pamen. An algorithm based on an iterative optimal stopping method for Feller processes with applications to impulse control, perturbation, and possibly zero random discount problems. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004629.

DeLillo:2021:PFM

[DMST21] T. K. DeLillo, J. Mears, and A. Silva-Trujillo. Potential flow in a multiply connected circle domain using series methods. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000649.

Duan:2022:CMF

[DMTW22] Huoyuan Duan, Junhua Ma, Roger C. E. Tan, and Can Wang. A coercive mixed formulation for the generalized Maxwell problem. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100409X.

Dabaghi:2023:HPM

[DMZ23] Jad Dabaghi, Yvon Maday, and Andrea Zoia. A hybrid parareal Monte Carlo algorithm for parabolic problems. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004071.

Durango:2020:GPA

[DN20a] Francisco Durango and Julia Novo. Generalized postprocessed approximations to the Navier– Stokes equations based on two grids. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305217.

Durango:2020:PEE

[DN20b] Francisco Durango and Julia Novo. A posteriori error estimations for mixed finite element approximations to the Navier– Stokes equations based on Newton-type linearization. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304327.

Dutta:2023:IGF

[DNK23] Subhankar Dutta, Hon Keung Tony Ng, and Suchandan Kayal. Inference for a general family of inverted exponentiated distributions under unified hybrid censoring with partially observed competing risks data. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005325.

Dembinska:2021:RPS

[DNS21] Anna Dembińska, Nikolay I. Nikolov, and Eugenia Stoimenova. Reliability properties of k-out-of-n systems with one cold standby unit. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030580X.

DelIsle:2021:SCM

[DO21] François De l’Isle and Robert G. Owens. Superconsistent collocation methods with applications to convection-dominated convection–diffusion equations. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306580.

Dehestani:2022:EAB

[DO22] H. Dehestani and Y. Ordokhani. An efficient approach based on Legendre-Gauss–Lobatto quadrature and discrete shifted Hahn polynomials for solving Caputo-fabrizio fractional Volterra partial integro-differential equations. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100474X.

deOliveira:2023:LCI

[dOdO23] Fabiana R. de Oliveira and Fabrícia R. de Oliveira. A locally convergent inexact projected Levenberg– Marquardt-type algorithm for large-scale constrained nonsmooth equations. Journal of Computational and Applied Mathematics, 427 (??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042723000651.

Dominguez:2021:SEL

[Dom21] Sebastián Domínguez. Steklov eigenvalues for the Lamé operator in linear elasticity. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001783.

Delgado:2021:ABD

[DOP21] J. Delgado, H. Orera, and J. M. Peña. Accurate bidiagonal decomposition and computations with generalized Pascal matrices. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000625.

Delgado:2023:EAB

[DOP23] J. Delgado, H. Orera, and J. M. Peña. An evaluation algorithm for q-Bézier triangular patches formed by convex combinations. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001280.

Dehestani:2021:CLW

[DOR21] H. Dehestani, Y. Ordokhani, and M. Razzaghi. Combination of Lucas wavelets with Legendre– Gauss quadrature for fractional Fredholm–Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303617.

Davydov:2023:ISS

[DOT23] Oleg Davydov, Dang Thi Oanh, and Ngo Manh Tuong. Improved stencil selection for meshless finite difference methods in 3D. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200629X.

Dai:2021:COM

[DP21a] Shibin Dai and Keith Promislow. Codimension one minimizers of highly amphiphilic mixtures. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306117.

DAmbrosio:2021:MCM

[DP21b] Raffaele D’Ambrosio and Beatrice Paternoster. Multivalue collocation methods free from order reduction. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305205.

Dick:2021:WIH

[DP21c] Josef Dick and Friedrich Pillichshammer. Weighted integration over a hyperrectangle based on digital nets and sequences. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100128X.

Santa:2023:DNN

[DP23] Francesco Della Santa and Sandra Pieraccini. Discontinuous neural networks and discontinuity learning. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003430.

Du:2024:HHT

[DP24] Tingsong Du and Yu Peng.

Hermite-Hadamard type inequalities for multiplicative Riemann–Liouville fractional integrals. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005265.

dePaula:2023:UCM

[dPdMQL23] Matheus Avila Moreira de Paula, Bárbara de Melo Quintela, and Marcelo Lobosco. On the use of a coupled mathematical model for understanding the dynamics of multiple sclerosis. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001073.

Difonzo:2024:EUA

[DPW24] Fabio V. Difonzo, Paweł Przybyłowicz, and Yue Wu. Existence, uniqueness and approximation of solutions to Carathéodory delay differential equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003552.

Davaeifar:2023:OMA

[DR23a] Sara Davaeifar and Jalil Rashidinia. Operational matrix approach based on two-dimensional Boubaker polynomials for solving nonlinear two-dimensional integral equations. Journal of Computational and Applied Mathematics, 421 (??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042722004290.

Dzhoha:2023:MAB

[DR23b] Andrii Dzhoha and Iryna Rozora. Multi-armed bandit problem with online clustering as side information. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000766.

Dinariev:2022:PAK

[DRB22] O. Yu. Dinariev, L. A. Pessoa Rego, and P. Bedrikovetsky. Probabilistic averaging in kinetic theory for colloidal transport in porous media. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004635.

Dlugosz:2023:NDP

[DRD^{+}23] Zofia Długosz, Michał Rajewski, Rafał Długosz, Tomasz Talaśka, and Witold Pedrycz. A new deterministic PSO algorithm for real-time systems implemented on lowpower devices. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001693.

Dykes:2021:LBF

[DRR^{+}21] L. Dykes, R. Ramlau, L. Reichel, K. M. Soodhalter, and R. Wagner. Lanczos-based fast blind deconvolution methods. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303587.

Das:2022:ASC

[DRR22] Pratibhamoy Das, Subrata Rana, and Higinio Ramos. On the approximate solutions of a class of fractional order nonlinear Volterra integro-differential initial value problems and boundary value problems of first kind and their convergence analysis. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304076.

Darania:2020:NAH

[DS20] P. Darania and F. Sotoudehmaram. Numerical analysis of a high order method for nonlinear delay integral equations. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300297.

Du:2021:ACC

[DS21] Yi Juan Du and Ming-Cheng Shiue. Analysis and computation of continuous data assimilation algorithms for Lorenz 63 system based on nonlinear nudging techniques. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305379.

Dehghan:2022:FCA

[DS22a] Mehdi Dehghan and Akbar Shirilord. A fast computational algorithm for computing outer pseudoinverses with numerical experiments. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000279.

Dehghan:2022:NAA

[DS22b] Mehdi Dehghan and Akbar Shirilord. A new approximation algorithm for solving generalized Lyapunov matrix equations. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005215.

daSilva:2024:GRC

[dS24] J. L. E. da Silva. Geometric relations in classical and quantum information theory using the Lambert-Tsallis W _{q} ( x ) function. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006088.

Florio:2024:PIN

[DSCF24] Mario De Florio, Enrico Schiassi, Francesco Calabrò, and Roberto Furfaro. Physics-informed neural networks for 2nd order ODEs with sharp gradients. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003400.

Du:2024:FDO

[DSCZ24] Dongsheng Du, Shennan Sun, Vincent Cocquempot, and Huanyu Zhao. H_{∞}/H_{−} fault detection observer design for nonlinear conformable fractionalorder systems. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006544.

daSilva:2024:SIP

[dSFSS24] P. C. da Silva, O. P. Ferreira, L. D. Secchin, and G. N. Silva. Secant-inexact projection algorithms for solving a new class of constrained mixed generalized equations problems. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005824.

Dominguez:2024:NDB

[DT24] Víctor Domínguez and Catalin Turc. Nyström discretizations of boundary integral equations for the solution of 2D elastic scattering problems. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005666.

Dassios:2020:GFC

[DTM20] Ioannis Dassios, Georgios Tzounas, and Federico Milano. Generalized fractional controller for singular systems of differential equations. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302107.

Dassios:2021:RSC

[DTM21] Ioannis Dassios, Georgios Tzounas, and Federico Milano. Robust stability criterion for perturbed singular systems of linearized differential equations. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030323X.

Dassios:2024:DMF

[DTMJ24] Ioannis Dassios, Georgios Tzounas, Federico Milano, and Andrey Jivkov. A discrete model for force-based elasticity and plasticity. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000451.

Disibuyuk:2019:BSA

[DU19]  Çetin Disibüyük and Sule

Ulutas. A B-spline approach to q-Eulerian polynomials. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304303.

Duman:2024:MPS

[Dum24]  Oktay Duman.  Max-

product Shepard operators based on multivariable Taylor polynomials. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004004.

Dostal:2021:AAL

[DV21a] Zdenek Dostál and Oldrich Vlach. An accelerated augmented Lagrangian algorithm with adaptive orthogonalization strategy for bound and equality constrained quadratic programming and its application to large-scale contact problems of elasticity. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001850.

Dwarka:2021:PAS

[DV21b] V. Dwarka and C. Vuik. Pollution and accuracy of solutions of the Helmholtz equation: a novel perspective from the eigenvalues. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001692.

Decleer:2021:NHO

[DVRG21] Pieter Decleer, Arne Van Londersele, Hendrik Rogier, and Dries Vande Ginste. Nonuniform and higher-order FDTD methods for the Schrödinger equation. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303149.

Decleer:2022:ADH

[DVRV22] Pieter Decleer, Arne Van Londersele, Hendrik Rogier, and Dries Vande Ginste. An alternatingdirection hybrid implicit– explicit finite-difference time-domain method for the Schrödinger equation. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005045.

Dong:2020:HOM

[DW20] Bo Dong and Wei Wang.

High-order multiscale discontinuous Galerkin methods for the one-dimensional stationary Schrödinger equation. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302533.

Dong:2023:HOM

[DW23] Bo Dong and Wei Wang. A high-order multiscale discontinuous Galerkin method for two-dimensional Schrödinger equation in quantum transport. Journal of Computational and Applied Mathematics, 418 (??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272200351X.

Dai:2020:EMS

[DWWZ20] Pingfei Dai, Qingbiao Wu, Hong Wang, and Xiangcheng Zheng. An efficient matrix splitting preconditioning technique for two-dimensional unsteady space-fractional diffusion equations. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306788.

Dong:2022:ORC

[DWWZ22] Lixiu Dong, Cheng Wang, Steven M. Wise, and Zhengru Zhang. Optimal rate convergence analysis of a numerical scheme for the ternary Cahn–Hilliard system with a Flory–Huggins– deGennes energy potential. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200228X.

Du:2020:SAL

[DWZ20] Yu Du, Haijun Wu, and Zhimin Zhang. Superconvergence analysis of linear FEM based on polynomial preserving recovery for Helmholtz equation with high wave number. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300224.

Ding:2020:GEA

[DXM^{+}20] Feng Ding, Ling Xu, Dandan Meng, Xue-Bo Jin, Ahmed Alsaedi, and Tasawar Hayat. Gradient estimation algorithms for the parameter identification of bilinear systems using the auxiliary model. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305801.

Deng:2021:OBI

[DXW21] Yulin Deng, Hongfeng Xu, and Jie Wu. Optimization of blockchain investment portfolio under artificial bee colony algorithm. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304908.

Dun:2022:IPC

[DXWC22] Meng Dun, Zhicun Xu, Lifeng Wu, and Yan Chen. The information priority of conformable fractional grey model. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002187.

Duan:2024:DCW

[DYF24] Mengmeng Duan, Yan Yang, and Minfu Feng. A defect correction weak Galerkin finite element method for the Kelvin– Voigt viscoelastic fluid flow model. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003480.

Dong:2021:DLR

[DYM21] Ning Dong, Bo Yu, and Zhao-Yun Meng. Decoupled low-rank iterative methods for a large-scale system of nonlinear matrix equations arising from electron transport of nano materials. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000637.

Ding:2024:NMR

[DYZ24] Huan Ding, Yang Yang, and Xinghui Zhong. Numerical methods for reinterpreted discrete fracture models with random inputs. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001882.

Davydova:2022:MTS

[DZ22] M. A. Davydova and S. A. Zakharova. Multidimensional thermal structures in the singularly perturbed stationary models of heat and mass transfer with a nonlinear thermal diffusion coefficient. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003538.

Dou:2020:SNE

[DZLC20] Fangfang Dou, Li-Ping Zhang, Zi-Cai Li, and C. S. Chen. Source nodes on elliptic pseudo-boundaries in the method of fundamental solutions for Laplace’s equation; selection of pseudo-boundaries. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301527.

Dang:2024:CDG

[DZZ24] Haoning Dang, Qilong Zhai, and Zhongshu Zhao. A conforming discontinuous Galerkin finite element method for Brinkman equations. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005630.

eddineBensad:2023:GCS

[eBI23] Alaa eddine Bensad and Aziz Ikemakhen. A general construction of spherical barycentric coordinates and applications. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200543X.

Elosmani:2021:LST

[EBVB21] Aissa Omar Elosmani, Djillali Bouagada, Paul Van Dooren, and Kamel Benyettou. LMI stability test for multidimensional linear state-space models. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306543.

Erbay:2021:SDN

[EEE21] H. A. Erbay, S. Erbay, and A. Erkip. A semidiscrete numerical method for convolution-type unidirectional wave equations. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304996.

Ebiefung:2022:BPP

[EFJK22] Aniekan A. Ebiefung, Luís M. Fernandes, Joaquim J. Júdice, and Michael M. Kostreva. A block principal pivoting algorithm for vertical generalized LCP with a vertical block P - matrix. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005367.

Eremin:2021:NMI

[EFTW21] Yuri A. Eremin, George Fikioris, Nikolaos L. Tsitsas, and Thomas Wriedt. A new method of internal auxiliary source-sinks (MI-ASS) for two-dimensional interior Dirichlet acoustic problems. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305227.

Engstrom:2022:HOC

[EGG22] C. Engström, S. Giani, and L. Grubisić. Higher order composite DG approximations of Gross– Pitaevskii ground state: Benchmark results and experiments. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002740.

Engstrom:2023:NSD

[EGG23] Christian Engström, Stefano Giani, and Luka Grubisić. Numerical solution of distributed-order timefractional diffusion-wave equations using Laplace transforms. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006331.

Egidi:2023:IHC

[EGM23] Nadaniela Egidi, Josephin Giacomini, and Pierluigi Maponi. Inverse heat conduction to model and optimise a geothermal field. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005556.

Ezquerro:2022:NCC

[EHV22] J. A. Ezquerro and M. A. Hernández-Verón. A new concept of convergence for iterative methods: Restricted global convergence. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303423.

Ezquerro:2022:GCE

[EHVM22] J. A. Ezquerro, M. A.

Hernández-Verón, and Á. A. Magreñán. On global convergence for an efficient third-order iterative process. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000364.

Ezquerro:2023:SIF

[EHVMM23] J. A. Ezquerro, M. A.

Hernández-Verón, Á. A. Magreñán, and A. Moysi. A significant improvement of a family of secanttype methods. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006008.

Ezquerro:2024:POQ

[EHVMM24] J. A. Ezquerro, M. A.

Hernández-Verón, Á. A. Magreñán, and A. Moysi. A procedure to obtain quadratic convergence from the secant method. Journal of Computational and Applied Mathematics, 448 (??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724001626.

ElHachimi:2024:TGK

[EJHR24] A. El Hachimi, K. Jbilou, M. Hached, and A. Ratnani. Tensor Golub–Kahan method based on Einstein product. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272400298X.

Eryilmaz:2024:EPG

[EK24a] Serkan Eryilmaz and Maria Kateri. Estimating the parameter of a geometric distribution from series system data. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002413.

Eslamian:2024:HVI

[EK24b] Mohammad Eslamian and Ahmad Kamandi. Hierarchical variational inequality problem and split common fixed point of averaged operators. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300434X.

Erkmen:2020:SOD

[EKÇ20] Dilek Erkmen, Songul Kaya, and Aytekin Çibik. A second order decoupled penalty projection method based on deferred correction for MHD in Elsässer variable. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306995.

Etminan:2022:ATF

[EKCF22] J. Etminan, H. Kamranfar, M. Chahkandi, and M. Fouladirad. Analysis of time-to-failure data for a repairable system subject to degradation. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000097.

Erdogan:2022:SCG

[EL22] Utku Erdogan and Gabriel J.

Lord. Strong convergence of a GBM based tamed integrator for SDEs and an adaptive implementation. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003265.

ElAsri:2023:CIC

[EL23] Brahim El Asri and Hafid Lalioui. Continuous and impulse controls differential game in finite horizon with Nash-equilibrium and application. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006070.

Elango:2023:SOS

[Ela23] Sekar Elango. Second order singularly perturbed delay differential equations with non-local boundary condition. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002412.

Eidnes:2021:LIS

[ELS21a] Sølve Eidnes, Lu Li, and Shun Sato. Linearly implicit structure-preserving schemes for Hamiltonian systems. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304923.

Elsawah:2021:ATD

[Els21b] A. M. Elsawah. An appealing technique for designing optimal large experiments with three-level factors. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304556.

Eguillon:2023:MER

[ELTD23] Yohan Eguillon, Bruno Lacabanne, and Damien Tromeur-Dervout. MIS-SILES: an efficient resolution of the co-simulation coupling constraint on nearly linear differential systems through a global linear formulation. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006112.

Emmanuel:2021:TEN

[EM21a] Coffie Emmanuel and Xuerong Mao. Truncated EM numerical method for generalised Ait-Sahalia-type interest rate model with delay. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304283.

Etore:2021:TSE

[ÉM21b] Pierre Étoré and Miguel Martinez. A transformed stochastic Euler scheme for multidimensional transmission PDE. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001710.

Eshghi:2021:NMF

[EMR21] Nasim Eshghi, Thomas Mach, and Lothar Reichel. New matrix function approximations and quadrature rules based on the Arnoldi process. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000613.

Ernsthausen:2020:SSR

[EN20] John M. Ernsthausen and Nedialko S. Nedialkov. Stepsize selection in the rigorous defect control of Taylor series methods. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304868.

ElHajj:2024:CSN

[EO24] A. El Hajj and A. Oussaily.

Convergent scheme for a non-local transport system modeling dislocations dynamics. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001791.

Einkemmer:2021:PSS

[EOR21] Lukas Einkemmer, Alexander Ostermann, and Mirko Residori. A pseudospectral splitting method for linear dispersive problems with transparent boundary conditions. Journal of Computational and Applied Mathematics, 385 (??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042720305318.

Ei:2022:MFS

[EOT22] Shin-Ichiro Ei, Hiroyuki Ochiai, and Yoshitaro Tanaka. Method of fundamental solutions for Neumann problems of the modified Helmholtz equation in disk domains. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004179.

Eryilmaz:2020:OAR

[EP20] Serkan Eryilmaz and Mustafa Hilmi Pekalp. On optimal age replacement policy for a class of coherent systems. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301795.

Esquivel:2021:FDS

[EPL21] Hugo Esquivel, Arun Prakash, and Guang Lin. Flow-driven spectral chaos (FSC) method for longtime integration of secondorder stochastic dynamical systems. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100296X.

El-Raheem:2020:ODM

[ER20a] A. M. Abd El-Raheem.

Optimal design of multiple accelerated life testing for generalized halfnormal distribution under type-I censoring. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305448.

Elgindy:2020:HOG

[ER20b] Kareem T. Elgindy and Hareth M. Refat. Highorder Gegenbauer integral spectral element method integrated with an adaptive Chebyshev optimization strategy for solving linear singularly perturbed differential equations. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300133.

El-Raheem:2021:ODM

[ER21] A. M. Abd El-Raheem. Optimal design of multiple constant-stress accelerated life testing for the extension of the exponential distribution under type-II censoring. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030385X.

Evstigneev:2022:CSD

[ERB^{+}22] N. M. Evstigneev, O. I. Ryabkov, A. N. Bocharov, V. P. Petrovskiy, and I. O. Teplyakov. Compensated summation and dot product algorithms for floatingpoint vectors on parallel architectures: Error bounds, implementation and application in the Krylov subspace methods. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002047.

El-Sayed:2023:STF

[ESA23] A. A. El-Sayed and P. Agarwal. Spectral treatment for the fractionalorder wave equation using shifted Chebyshev orthogonal polynomials. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005313.

Egger:2021:MRK

[ESS21] H. Egger, K. Schmidt, and V. Shashkov. Multistep and Runge–Kutta convolution quadrature methods for coupled dynamical systems. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306235.

Eremin:2023:NSD

[ETKF23] Yuri A. Eremin, Nikolaos L. Tsitsas, Minas Kouroublakis, and George Fikioris. New scheme of the discrete sources method for two-dimensional scattering problems by penetrable obstacles. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002734.

Evirgen:2023:TNV

[Evi23] Firat Evirgen. Transmission of Nipah virus dynamics under Caputo fractional derivative. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003314.

Elsawah:2023:NTC

[EWCQ23] A. M. Elsawah, Yi-An Wang, S. M. Celem, and Hong Qin. A novel technique for constructing nonregular nine-level designs: Adjusted multiple tripling technique. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006148.

Elsawah:2024:MER

[EWT24] A. M. Elsawah, Yi-An Wang, and Fatih Tank. Minimum energy representative points. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004703.

Eroglu:2023:CAF

[EY23] Beyza Billur Iskender Eroglu and Dilara Yapiskan. Comparative analysis on fractional optimal control of an SLBS model. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004381.

Fabien:2022:NEA

[Fab22] Maurice S. Fabien. Numerical error analysis for an energy-stable HDG method for the Allen– Cahn equation. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004222.

Faleichik:2021:MRM

[Fal21] Boris V. Faleichik. Minimal residual multistep methods for large stiff nonautonomous linear problems. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305011.

Fan:2022:DAV

[Fan22a] Huiling Fan. The digital asset value and currency supervision under deep learning and blockchain technology. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100621X.

Fang:2022:SRR

[Fan22b] Weifu Fang. Simultaneous recovery of Robin boundary and coefficient for the Laplace equation by shape derivative. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001686.

Fang:2022:CMB

[Fan22c] Ximing Fang. The convergence of modulus-based matrix splitting iteration methods for implicit complementarity problems. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000917.

Fageot:2020:SAP

[FAUU20] Julien Fageot, Shayan Aziznejad, Michael Unser, and Virginie Uhlmann. Support and approximation properties of Hermite splines. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305084.

Fernandez:2019:MAN

[FAVM19] Francisco J. Fernández, Lino J. Alvarez-Vázquez, and Aurea Martínez. Mathematical analysis and numerical resolution of a heat transfer problem arising in water recirculation. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304054.

See corrigendum [FAVM20].

Fernandez:2020:CMA

[FAVM20] Francisco J. Fernández, Lino J. Alvarez-Vázquez, and Aurea Martínez. Corrigendum to “Mathematical analysis and numerical resolution of a heat transfer problem arising in water recirculation” [j. comput. appl. math. 366 (2020) 112402]. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301503.

See [FAVM19].

Fernandez:2024:PST

[FB24] Arran Fernandez and Suzan Cival Buranay.

The Peano-Sard theorem for Caputo fractional derivatives and applications. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006489.

Fang:2020:OGH

[FBJ20] Longxiang Fang, N. Balakrishnan, and Qing Jin. Optimal grouping of heterogeneous components in series-parallel and parallelseries systems under Archimedean copula dependence. Journal of Computational and Applied Mathematics, 377 (??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720302077.

Fu:2022:CNG

[FC22] Yaoyao Fu and Liqun Cao. The Crank–Nicolson Galerkin method and convergence for the timedependent Maxwell–Dirac system under the Lorentz gauge. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100594X.

Fishelov:2023:OCT

[FC23] D. Fishelov and J.-P. Croisille. Optimal convergence for time-dependent linearized Kuramoto–Sivashinsky type problems: a new approach. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001735.

Fekete:2022:EPO

[FCS22] Imre Fekete, Sidafa Conde, and John N. Shadid. Embedded pairs for optimal explicit strong stability preserving Runge– Kutta methods. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001388.

Francois:2024:HOA

[FDDM24] Laurent François, Joël Dupays, Dmitry Davidenko, and Marc Massot. High-order adaptive time discretisation of one-dimensional low-Mach reacting flows: a case study of solid propellant combustion. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000074.

Ferreira:2021:CNE

[FdOPS21] J. A. Ferreira, P. de Oliveira, G. Pena, and E. Silveira. Coupling nonlinear electric fields and temperature to enhance drug transport: an accurate numerical tool. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304180.

Fankina:2023:HQS

[FFRS23] Irina V. Fankina, Alexey I.

Furtsev, Evgeny M. Rudoy, and Sergey A. Sazhenkov. The homogenized quasistatic model of a thermoelastic composite stitched with reinforcing threads. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300290X.

Falletta:2024:VEM

[FFS24] Silvia Falletta, Matteo Ferrari, and Letizia Scuderi. A virtual element method for the solution of 2D time-harmonic elastic wave equations via scalar potentials. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005691.

Fakhar:2020:MMP

[FFT20] Mohammad Hosein Fakhar, Ahmad Fakhar, and Hamidreza Tabatabaei. Mathematical modeling of pipes reinforced by agglomerated CNTs conveying turbulent nanofluid and application of semi-analytical method for studying the instable Nusselt number and fluid velocity. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302363.

Faigenbaum-Golovin:2023:MRD

[FGL23] Shira Faigenbaum-Golovin and David Levin. Manifold reconstruction and denoising from scattered data in high dimension. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004162.

Fazzi:2021:GAA

[FGM21] Antonio Fazzi, Nicola Guglielmi, and Ivan Markovsky. Generalized algorithms for the approximate matrix polynomial GCD of reducing data uncertainties with application to MIMO system and control. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001187.

Flores:2023:STF

[FGN^{+}23] J. Flores, A. García, M. Negreanu, E. Salete, F. Ureña, and A. M. Vargas. A spatio-temporal fully meshless method for hyperbolic PDEs. Journal of Computational and Applied Mathematics, 430 (??):??, October 1, 2023.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042723001383.

Ferreira:2024:NAL

[FGP24] J. A. Ferreira, H. P. Gómez, and L. Pinto. Numerical analysis of lightcontrolled drug delivery systems. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001110.

Fortes:2023:FFD

[FGPP23] M. A. Fortes, P. González, A. Palomares, and M. Pasadas. Fitting and filling of 3D datasets with volume constraints using radial basis functions under tension. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004393.

Ferrandi:2024:HRQ

[FH24] Giulia Ferrandi and Michiel E. Hochstenbach. A homogeneous Rayleigh quotient with applications in gradient methods. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003849.

Farooq:2020:RTA

[FHB20] Umer Farooq, Randy D.

Hazlett, and D. Krishna Babu. Rate transient analysis of arbitrarily-oriented, hydraulically-fractured media. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302570.

Fu:2022:LMM

[FHC22] Zhenwu Fu, Bo Han, and Yong Chen. Levenberg– Marquardt method with general convex penalty for nonlinear inverse problems. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003939.

Furst:2021:PBS

[FHH21] Jirí Fürst, Jan Halama, and Vladimír Hric. A pressure based solver for simulation of non-equilibrium wet steam flows. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002752.

Feng:2023:CPF

[FHM23] Qiwei Feng, Bin Han, and Peter Minev. Compact 9-point finite difference methods with high accuracy order and/or Mmatrix property for elliptic cross-interface problems. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300095X.

Fang:2021:TFT

[FHY^{+}21] Yonglei Fang, Ting Huang, Xiong You, Juan Zheng, and Bin Wang. Twofrequency trigonometricallyfitted and symmetric linear multi-step methods for second-order oscillators. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306038.

Frerichs:2021:RSO

[FJ21] Derk Frerichs and Volker John. On reducing spurious oscillations in discontinuous Galerkin (DG) methods for steady-state convection–diffusion equations. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001060.

Fellner:2024:CSD

[FJ24] Markus Fellner and Ansgar Jüngel. A coupled stochastic differential reaction– diffusion system for angiogenesis. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005149.

Felipe:2024:MSI

[FJMP24] Angel Felipe, Maria Jaenada, Pedro Miranda, and Leandro Pardo. Model selection for independent not identically distributed observations based on Rényi’s pseudodistances. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005745.

Franz:2023:PTP

[FK23] Sebastian Franz and Natalia Kopteva. Pointwisein-time a posteriori error control for higher-order discretizations of timefractional parabolic equations. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000663.

Ferjancic:2021:RMD

[FKV21] Karla Ferjancic, Marjeta Knez, and Vito Vitrih. On C^{2} rational motions of degree six. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306154.

Fang:2023:CAF

[FL23] Qingxiang Fang and Xiaoping Liu. A comment on “Attractivity for functional Volterra integral equations of convolution type”. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000031.

See [AL16].

Fornaca:2024:MSM

[FL24] Elena Fornaca and Paola Lamberti. Multilevel Schoenberg–Marsden variation diminishing operator and related quadratures. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000530.

Ferreyra:2024:SDM

[FLMP24] D. E. Ferreyra, F. E. Levis, Saroj B. Malik, and A. N. Priori. On star-dagger matrices and the core-EP decomposition. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005022.

Favati:2020:CSM

[FLMR20] P. Favati, G. Lotti, O. Menchi, and F. Romani. Construction of the similarity matrix for the spectral clustering method: numerical experiments. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300868.

Feng:2020:UMC

[FLT20] Libo Feng, Fawang Liu, and Ian Turner. An unstructured mesh control volume method for twodimensional space fractional diffusion equations with variable coefficients on convex domains. Journal of Computational and Applied Mathematics, 364 (??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0377042719303164.

Feng:2021:ENA

[FLZ21] Chunrong Feng, Yu Liu, and Huaizhong Zhao. Ergodic numerical approximation to periodic measures of stochastic differential equations. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100323X.

Faghih:2021:NFC

[FM21a] A. Faghih and P. Mokhtary.

A new fractional collocation method for a system of multi-order fractional differential equations with variable coefficients. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304301.

Fallah:2021:CME

[FM21b] Somayeh Fallah and Farshid Mehrdoust. CEV model equipped with the longmemory. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306506.

Fang:2021:IBE

[FM21c] Meie Fang and Weiyin Ma. Isogeometric boundary element analysis based on UEsplines. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003228.

Farooq:2022:HPM

[FM22] Asma Farooq and Stefano Maset. How perturbations in the matrix of linear systems of ordinary differential equations propagate along solutions. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721006129.

Fernandez:2023:CEM

[FMAV23] Francisco J. Fernández, Aurea Martínez, and Lino J. Alvarez-Vázquez. Controlling eutrophication by means of water recirculation: an optimal control perspective. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004848.

Fernandez-Martinez:2020:CDI

[FMFM20] Juan L. Fernández-Martínez and Zulima Fernández-Muñiz. The curse of dimensionality in inverse problems. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930576X.

Flumian:2024:SSA

[FMNT24] Lea Flumian, Markus Matilainen, Klaus Nordhausen, and Sara Taskinen. Stationary subspace analysis based on secondorder statistics. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003230.

Fermo:2024:SIA

[FMRS24] Luisa Fermo, Gradimir V.

Milovanović, Lothar Reichel, and Miodrag M. Spalević. Special issue on advances in quadrature, cubature, and the solution of integral equations with applications. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000621.

Fika:2020:MAA

[FMRT20] Paraskevi Fika, Marilena Mitrouli, Paraskevi Roupa, and Dimitrios Triantafyllou. The e-MoM approach for approximating matrix functionals. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719302122.

Frederick:2021:TMT

[FMTB21] Jennifer Frederick, Alejandro Mota, Irina Tezaur, and Diana Bull. A thermo-mechanical terrestrial model of Arctic coastal erosion. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001527.

Ferrada:2020:FTF

[FNH20] Héctor Ferrada, Cristóbal A.

Navarro, and Nancy Hitschfeld. A filtering technique for fast Convex Hull construction in R^{2}. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302870.

Fermo:2022:WSL

[FO22] Luisa Fermo and Donatella Occorsio. Weakly singular linear Volterra integral equations: a Nyström method in weighted spaces of continuous functions. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721005914.

Forbes:2024:CRS

[For24]  G. W. Forbes.  Com-

ment re: Schemes for cubature over the unit disk found via numerical optimization. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005873.

Fasano:2021:PCQ

[FP21] Giovanni Fasano and Raffaele Pesenti. Polarity and conjugacy for quadratic hypersurfaces: a unified framework with recent advances. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305392.

Ferreira:2024:FFN

[FP24] J. A. Ferreira and G. Pena. FDM/FEM for nonlinear convection–diffusion– reaction equations with Neumann boundary conditions — convergence analysis for smooth and nonsmooth solutions. Journal of Computational and Applied Mathematics, 446 (??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724001158.

Fernandez-Pendas:2022:SPC

[FPCV^{+}22] Mario Fernández-Pendás, Elías F. Combarro, Sofia Vallecorsa, José Ranilla, and Ignacio F. Rúa. A study of the performance of classical minimizers in the Quantum Approximate Optimization Algorithm. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000078.

Fernandez:2021:OLS

[FPGSR21] Arturo J. Fernández, Carlos J. Pérez-González, and M. Mercedes Suárez-Rancel. Optimal lot sentencing based on defective counts and prior acceptability. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306087.

Fernandez:2021:EHS

[FPLZ21] Pascual Fernández, Blas Pelegrín, Algirdas Lancinskas, and Julius Zilinskas. Exact and heuristic solutions of a discrete competitive location model with Pareto–Huff customer choice rule. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030491X.

Franchini:2024:SGM

[FPR^{+}24] Giorgia Franchini, Federica Porta, Valeria Ruggiero, Ilaria Trombini, and Luca Zanni. A stochastic gradient method with variance control and variable learning rate for deep learning. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003327.

Farouki:2021:AMC

[FPS21] Rida T. Farouki, Francesca Pelosi, and Maria Lucia Sampoli. Approximation of monotone clothoid segments by degree 7 Pythagorean-hodograph curves. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304015.

Ferreira:2022:NAP

[FPS22] J. A. Ferreira, L. Pinto, and R. F. Santos. Numerical analysis of a porouselastic model for convection enhanced drug delivery. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003411.

Franses:2021:TVL

[Fra21] Philip Hans Franses. Timevarying lag cointegration. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030563X.

From:2024:SNB

[Fro24] Steven G. From. Some new bounds for the blow-up time of solutions for certain nonlinear Volterra integral equations. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000839.

Fish:2024:IEM

[FRR24] Alex C. Fish, Daniel R. Reynolds, and Steven B. Roberts. Implicit-explicit multirate infinitesimal stagerestart methods. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004788.

Ferraro:2023:FSS

[FRT23] Augusto Ferraro, Daniel Alejandro Rossit, and Adrián Toncovich. Flow shop scheduling problem with non-linear learning effects: a linear approximation scheme for non-technical users. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005817.

Fareed:2020:EAI

[FS20] Hiba Fareed and John R. Singler. Error analysis of an incremental proper orthogonal decomposition algorithm for PDE simulation data. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305308.

Flynn:2021:CAL

[FS21] Darren Flynn and Khodr Shamseddine. On computational applications of the Levi–Civita field. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303320.

Fang:2024:DBA

[FS24a] Lishan Fang and Linda Stals. Data-based adaptive refinement of finite element thin plate spline. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002255.

Finesso:2024:SDA

[FS24b] Lorenzo Finesso and Peter Spreij. Synchronous deautoconvolution algorithm for discrete-time positive signals via I -divergence approximation. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002759.

Fang:2024:CNF

[FSC24] Jun Fang, Zhijun Shen, and Xia Cui. A coupled nonlinear finite element scheme for anisotropic diffusion equation with nonlinear capacity term. Journal of Computational and Applied Mathematics, 438 (??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723004569.

Frank:2024:SCM

[FSR24] Anna-Simone Frank, Alexander Sikorski, and Susanna Röblitz. Spectral clustering of Markov chain transition matrices with complex eigenvalues. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000402.

Fernandez:2019:SAP

[FU19] Arran Fernandez and Ceren Ustaoglu. On some analytic properties of tempered fractional calculus. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304030.

Fragoyiannis:2024:SAA

[FV24] George Fragoyiannis and Panayiotis Vafeas. A semianalytical approach for the computation of ellipsoidal harmonics. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300362X.

Feng:2022:NVA

[FW22] Bo Feng and Gang Wu. On a new variant of Arnoldi method for approximation of eigenpairs. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003629.

Feng:2022:SRD

[FY22] Yue Feng and Jia Yin.

Spatial resolution of different discretizations over long-time for the Dirac equation with small potentials. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001509.

Fan:2024:CSO

[FYL^{+}24] Ziyao Fan, Chen Yang, Bochao Lin, Ying Yang, and Qinghe Shi. Convex set-oriented singular value decomposition with bounded uncertainties. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001924.

Fan:2020:MSP

[FZ20] Hongtao Fan and Bing Zheng. Modified SIM-PLE preconditioners for saddle point problems from steady incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303632.

Fang:2021:HOE

[FZ21] Jinwei Fang and Rui Zhan. High order explicit exponential Runge– Kutta methods for semilinear delay differential equations. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305707.

Feng:2022:TRM

[FZQ22] Xiaoli Feng, Meixia Zhao, and Zhi Qian. A Tikhonov regularization method for solving a backward timespace fractional diffusion problem. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000905.

Gomez-Aguilar:2019:PEL

[GAA19] J. F. Gómez-Aguilar and Abdon Atangana. Power and exponentials laws: Theory and application. Journal of Computational and Applied Mathematics, 354(??):52–65, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704271930010X. See

retraction notice [GAA21].

Gomez-Aguilar:2021:RNP

[GAA21] J. F. Gómez-Aguilar and Abdon Atangana. Retraction notice to “Power and exponentials laws: Theory and application” [j. comput. appl. math. 354 (2019) 52–65]. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002314.

See [GAA19].

Garcia-Archilla:2022:EAP

[GANR22] Bosco García-Archilla, Julia Novo, and Samuele Rubino. Error analysis of proper orthogonal decomposition data assimilation schemes with grad-div stabilization for the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000954.

Goujon:2021:SSM

[GANU21] Alexis Goujon, Shayan Aziznejad, Alireza Naderi, and Michael Unser. Shortestsupport multi-spline bases for generalized sampling. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002326.

Gao:2022:NAS

[Gao22] Jing Gao. Numerical analysis of the spectrum for the highly oscillatory integral equation with weak singularity. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004428.

Guney:2021:RMS

[GBA21] Yesim Güney, Hamparsum Bozdogan, and Olcay Arslan. Robust model selection in linear regression models using information complexity. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003010.

Gallia:2023:ARC

[GBG23] Mariachiara Gallia, Tommaso Bellosta, and Alberto Guardone. Automatic roughness characterization of simulated ice shapes. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000584.

Gbikpi-Benissan:2021:ASM

[GBM21] Guillaume Gbikpi-Benissan and Frédéric Magoulès. Asynchronous substructuring method with alternating local and global iterations. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001503.

Gbikpi-Benissan:2023:AMB

[GBM23] Guillaume Gbikpi-Benissan and Frédéric Magoulès.

Asynchronous multisplittingbased primal Schur method. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000043.

Gerontitis:2022:FVP

[GBTS22] Dimitrios Gerontitis, Ratikanta Behera, Panagiotis Tzekis, and Predrag Stanimirović. A family of varyingparameter finite-time zeroing neural networks for solving time-varying Sylvester equation and its application. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004489.

Gorny:2019:THC

[GC19] Julian Górny and Erhard Cramer. Type-I hybrid censoring of multiple samples. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304078.

Gozutok:2023:NMD

[GÇSA23] Ugur Gözütok, Hüsnü Anil Çoban, Yasemin Sagiroglu, and Juan Gerardo Alcázar. A new method to detect projective equivalences and symmetries of rational 3D curves. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003983.

Ge:2019:IGN

[GCV19] Xiao Ge, Xiao Shan Chen, and Seak-Weng Vong. Inexact generalized Noda iterations for generalized eigenproblems. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304212.

Galvis:2020:NUF

[GCV20] Juan Galvis, Luis F. Contreras, and Carlos Vázquez. Numerical upscaling of the free boundary dam problem in multiscale highcontrast media. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304406.

Gharibi:2022:OEB

[GDA22] Zeinab Gharibi, Mehdi Dehghan, and Mostafa Abbaszadeh. Optimal error bound for immersed weak Galerkin finite element method for elliptic interface problems. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002795.

Gutierrez:2021:NST

[GDG^{+}21] Bárbara Arizmendi Gutiérrez, Alberto Della Noce, Mariachiara Gallia, Tommaso Bellosta, and Alberto Guardone. Numerical simulation of a thermal ice protection system including state-of-the-art liquid film model. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100073X.

Gomez-Deniz:2021:BND

[GDSCO21] E. Gómez-Déniz, J. M.

Sarabia, and E. Calderín-Ojeda. Bimodal normal distribution: Extensions and applications. Journal of Computational and Applied Mathematics, 388 (??):??, May 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305835.

Gu:2023:HMI

[GDZN23] Ya Gu, Wei Dai, Quanmin Zhu, and Hassan Nouri. Hierarchical multi-innovation stochastic gradient identification algorithm for estimating a bilinear statespace model with moving average noise. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004046.

Giani:2023:ASP

[GEG23] Stefano Giani, Christian Engström, and Luka Grubisić. khp-adaptive spectral projection based discontinuous Galerkin method for the numerical solution of wave equations with memory. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001565.

Goligerdian:2024:ATP

[GeOJD24] Arash Goligerdian, Mahmood Khaksar e Oshagh, and Majid Jaberi-Douraki. Applying thin plate splines to the Galerkin method for the numerical simulation of a nonlinear model for population dynamics. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002863.

Goujon:2024:NRP

[GEU24] Alexis Goujon, Arian Etemadi, and Michael Unser. On the number of regions of piecewise linear neural networks. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006118.

Gatica:2024:BSB

[GG24] Gabriel N. Gatica and Zeinab Gharibi. A Banach spaces-based fully mixed virtual element method for the stationary twodimensional Boussinesq equations. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001341.

Gao:2024:CRS

[GGHY24] Shuaibin Gao, Qian Guo, Junhao Hu, and Chenggui Yuan. Convergence rate in L^{p} sense of tamed EM scheme for highly nonlinear neutral multiple-delay stochastic McKean-Vlasov equations. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006258.

Goncalves:2021:IPL

[GGO21] Douglas S. Gonçalves, Max L. N. Gonçalves, and Fabrícia R. Oliveira. An inexact projected LM type algorithm for solving convex constrained nonlinear equations. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000406.

Gomez:2021:SPB

[GGPP21] Ana I. Gómez, Domingo Gómez-Pérez, and Friedrich Pillichshammer. Secure pseudorandom bit generators and point sets with low star-discrepancy. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002211.

Galan-Garcia:2023:SSO

[GGRCGG^{+}23] José Luis Galán-García, Pedro Rodríguez-Cielos, María Ángeles Galán-García, Morgan le Goff, Yolanda Padilla-Domínguez, Pablo Rodríguez-Padilla, Iván Atencia, and Gabriel Aguilera-Venegas. SODES: Solving ordinary differential equations step by step. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000717.

Guillen-Gonzalez:2024:FEN

[GGT24] F. Guillén-González and G. Tierra. Finite element numerical schemes for a chemo-attraction and consumption model. Journal of Computational and Applied Mathematics, 441 (??):??, May 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006209.

Gatica:2022:PEA

[GGVRB22] Gabriel N. Gatica, Bryan Gómez-Vargas, and Ricardo Ruiz-Baier. A posteriori error analysis of mixed finite element methods for stress-assisted diffusion problems. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200036X.

Guo:2020:OAS

[GH20] Xue-Ping Guo and Apostolos Hadjidimos. Optimal accelerated SOR-like (ASOR) method for singular symmetric saddle point problems. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306673.

Giani:2021:EPD

[GH21] Stefano Giani and Harri Hakula. On effects of perforated domains on parameter-dependent free vibration. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100145X.

Guo:2022:CNE

[GH22] Yingwen Guo and Yinnian He. Crank–Nicolson extrapolation and finite element method for the Oldroyd fluid with the midpoint rule. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200214X.

Giani:2023:ECL

[GH23] Stefano Giani and Harri Hakula. On effects of concentrated loads on perforated sensitive shells of revolution. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001097.

Gu:2020:EVC

[GHC^{+}20] Xian-Ming Gu, Ting-Zhu Huang, Bruno Carpentieri, Akira Imakura, Ke Zhang, and Lei Du. Efficient variants of the CMRH method for solving a sequence of multishifted non-Hermitian linear systems simultaneously. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300790.

Goyal:2024:SMB

[GHF24] Dheeraj Goyal, Nil Kamal Hazra, and Maxim Finkelstein. Shock models based on renewal processes with matrix Mittag-Leffler distributed interarrival times. Journal of Computational and Applied Mathematics, 435(??):??, ???? 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000341.

Gao:2024:LRM

[GHG24] Kaixin Gao, Zheng-Hai Huang, and Lulu Guo. Low-rank matrix recovery problem minimizing a new ratio of two norms approximating the rank function then using an ADMM-type solver with applications. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005083.

Gie:2024:SAP

[GHJM24] Gung-Min Gie, Youngjoon Hong, Chang-Yeol Jung, and Tselmuun Munkhjin. Semi-analytic PINN methods for boundary layer problems in a rectangular domain. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002395.

Gui:2021:FME

[GHL21] Wenyong Gui, Rongtan Huang, and X. Sheldon Lin. Fitting multivariate Erlang mixtures to data: a roughness penalty approach. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305070.

Giesl:2021:CVC

[GHM21] Peter Giesl, Sigurdur Hafstein, and Iman Mehrabinezhad. Computation and verification of contraction metrics for exponentially stable equilibria. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306233.

Gu:2021:AKT

[GHTC21] Ruixue Gu, Bo Han, Shanshan Tong, and Yong Chen. An accelerated Kaczmarz type method for nonlinear inverse problems in Banach spaces with uniformly convex penalty. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305021.

Huang:2024:ERS

[gHtSjT24] Yi gong Huang, Tong tong Shang, and Guo ji Tang. Existence results of solutions to a generalized vertical polynomial complementarity problem in terms of vertical block tensor tuples. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003625.

Grigoriev:2020:CIA

[GIV20] Vasiliy V. Grigoriev, Oleg Iliev, and Petr N. Vabishchevich. Computational identification of adsorption and desorption parameters for pore scale transport in periodic porous media. Journal of Computational and Applied Mathematics, 370 (??):??, May 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306661.

Ganie:2019:SQL

[GJ19] Javid Ahmad Ganie and Renu Jain. On a system of q-Laplace transform of two variables with applications.

Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304108.

Gu:2021:AGR

[GJ21] Jiaxi Gu and Jae-Hun Jung. Adaptive Gaussian radial basis function methods for initial value problems: Construction and comparison with adaptive multiquadric radial basis function methods. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303277.

Gao:2023:PIR

[GJ23] Jing Gao and Yaolin Jiang. On product integration rules for highly oscillatory integrals on a triangle. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004733.

Gao:2020:ESO

[GJDL20] Huadong Gao, Lili Ju, Ravindra Duddu, and Hongwei Li. An efficient second-order linear scheme for the phase field model of corrosive dissolution. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304753.

Garcia-Jimenez:2023:BFP

[GJGDPR23]  Rafael  García-Jiménez,

J. Carlos García-Díaz, and Alexander D. Pulido-Rojano. Bicriteria food packaging process optimization in double-layered upright and diagonal multihead weighers. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001127.

Gie:2023:SAS

[GJL23] Gung-Min Gie, Chang-Yeol Jung, and Hoyeon Lee. Semi-analytic shooting methods for Burgers’ equation. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003491.

Guo:2023:JDD

[GJM^{+}23] Yu Guo, Qiyu Jin, Jean-Michel Morel, Tieyong Zeng, and Gabriele Facciolo. Joint demosaicking and denoising benefits from a two-stage training strategy. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002741.

Ganji:2020:NAS

[GJN20] R. M. Ganji, H. Jafari, and S. Nemati. A new approach for solving integrodifferential equations of variable order. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302375.

Gao:2024:CEO

[GJNM24] Jinwu Gao, Ruru Jia, Idin Noorani, and Farshid Mehrdoust. Calibration of European option pricing model in uncertain environment: Valuation of uncertainty implied volatility. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001390.

Gahn:2021:CEE

[GJNR21] Markus Gahn, Willi Jäger, and Maria Neuss-Radu. Correctors and error estimates for reaction–diffusion processes through thin heterogeneous layers in case of homogenized equations with interface diffusion. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304179.

Guan:2023:MMR

[GJW23] Xiaofei Guan, Lijian Jiang, and Yajun Wang. Multiscale model reduction for stochastic elasticity problems using ensemble variable-separated method. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004939.

Kang:2021:CRK

[gK21a] Chuan gang Kang. Convergence rates of the Kaczmarz-tanabe method for linear systems. Journal of Computational and Applied Mathematics, 394 (??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721001977.

Grover:2021:IRT

[GK21b] Lovleen Kumar Grover and Amanpreet Kaur. An improved regression type estimator of population mean with two auxiliary variables and its variant using robust regression method. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303630.

Kang:2023:SFC

[gK23] Chuan gang Kang. The standard forms and convergence theory of the Kaczmarz–Tanabe type methods for solving linear systems. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002777.

Gyulov:2021:FFV

[GKV21] Tihomir B. Gyulov, Miglena N.

Koleva, and Lubin G. Vulkov. Fitted finite volume method for indifference pricing in an exponential utility regimeswitching model. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304960.

Kang:2021:ECR

[gKZ21] Chuan gang Kang and Heng Zhou. The extensions of convergence rates of Kaczmarz-type methods. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303903.

Goffard:2020:OPE

[GL20] Pierre-Olivier Goffard and Patrick J. Laub. Orthogonal polynomial expansions to evaluate stoploss premiums. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306533.

Gao:2024:OLC

[GL24] Wenhui Gao and Yaning Lin. Open-loop and closed-loop Nash equilibria for the LQ stochastic difference game. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300688X.

Ghimire:2020:HCP

[GLCL20] B. Khatri Ghimire, Xinxiang Li, C. S. Chen, and A. R. Lamichhane. Hybrid Chebyshev polynomial scheme for solving elliptic partial differential equations. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303218.

Guo:2021:SNS

[GLHJ21] Ping Guo, Meng Liu, Zhixiong He, and Hongen Jia. Stability of numerical solutions for the stochastic pantograph differential equations with variable step size. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030594X.

Gao:2024:FDN

[GLHL24] Yali Gao, Rui Li, Xiaoming He, and Yanping Lin. A fully decoupled numerical method for Cahn–Hilliard–Navier– Stokes–Darcy equations based on auxiliary variable approaches. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003072.

Guo:2021:EAS

[GLLZ21] Ruchi Guo, Tao Lin, Yanping Lin, and Qiao Zhuang. Error analysis of symmetric linear/ bilinear partially penalized immersed finite element methods for Helmholtz interface problems. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306695.

Gao:2021:NFV

[GLYS21] Yulong Gao, Yonghai Li, Guangwei Yuan, and Zhiqiang Sheng. New finite volume element methods in the ALE framework for time-dependent convection–diffusion problems in moving domains. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001564.

Gong:2024:IEN

[GLZ24] Wenbo Gong, Hengguang Li, and Qinghui Zhang. Improved enrichments and numerical integrations in SGFEM for interface problems. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004843.

Gande:2022:HON

[GM22] Naga Raju Gande and H. Madduri. Higher order numerical schemes for the solution of fractional delay differential equations. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004325.

Gao:2023:URB

[GM23] Rong Gao and Yebao Ma. Uncertain random bilevel programming models and their application to shared capacity routing problem. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005635.

Ghosh:2024:CEN

[GM24] N. Ghosh and H. S.

Mahato. Corrector estimates and numerical simulations of a system of diffusion–reaction– dissolution–precipitation model in a porous medium. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723004454.

Garcia-Merino:2024:SPC

[GMCJGM24] J. C. García-Merino, C. Calvo-Jurado, and E. García-Macías. Sparse polynomial chaos expansion for universal stochastic kriging. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000438.

Ghasami:2020:LPC

[GMK20] Safdar Ghasami, Mohsen Maleki, and Zahra Khodadadi. Leptokurtic and platykurtic class of robust symmetrical and asymmetrical time series models. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300972.

Garcia-Mora:2018:MDI

[GMSR18] B. García-Mora, C. Santamaría, and G. Rubio. Modeling dependence in the inter-failure times. An analysis in reliability models by Markovian arrival processes. Journal of Computational and Applied Mathematics, 343(??):762–770, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042717306428.

See correction [GMSR21].

Garcia-Mora:2021:CMD

[GMSR21] B. García-Mora, C. Santamaría, and G. Rubio. Corrigendum to “Modeling dependence in the interfailure times. An analysis in Reliability models by Markovian Arrival Processes” [J. Comp. Appl. Math. 343, (2018), 762– 770]. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303769.

See [GMSR18].

Gapyak:2024:VMB

[GMW24] Vladyslav Gapyak, Thomas März, and Andreas Weinmann. Variational modelbased reconstruction techniques for multi-patch data in Magnetic Particle Imaging. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002966.

Green:2024:TCH

[GN24] Christopher C. Green and Mohamed M. S. Nasser. Towards computing the harmonic-measure distribution function for the middle-thirds Cantor set. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001535.

Garcia-Nieto:2019:HOW

[GNGGL^{+}19] P. J. García-Nieto, E. García-Gonzalo, F. Sánchez Lasheras, J. R. Alonso Fernández, and C. Díaz Muñiz. A hybrid DE optimized wavelet kernel SVR-based technique for algal atypical proliferation forecast in La Barca reservoir: a case study. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304200.

Gracia:2020:SPR

[GO20] J. L. Gracia and E. O’Riordan.

Singularly perturbed reaction– diffusion problems with discontinuities in the initial and/or the boundary data. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306430.

Gracia:2024:SPE

[GO24] José Luis Gracia and Eugene O’Riordan. Singularly perturbed elliptic problems of convection– diffusion type with nonsmooth inflow/outflow boundary conditions. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003431.

Gokce:2023:DBP

[Gök23] Aytül Gökçe. Dynamical behaviour of a predatorprey system encapsulating the fear affecting death rate of prey and intraspecific competition: Revisited in a fluctuating environment. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004472.

Golshan:2024:NSN

[Gol24] Hamid Mottaghi Golshan. Numerical solution of nonlinear m-dimensional Fredholm integral equations using iterative Newton–Cotes rules. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001675.

Gonzalez-Pinto:2021:AWM

[GPHAPR21] S. González-Pinto, D. Hernández-Abreu, and S. Pérez-Rodríguez. AMFR-W-methods for parabolic problems with mixed derivates. Applications to the Heston model. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305230.

Gonzalez-Pinto:2023:HOP

[GPHHA23] S. González-Pinto, E. Hairer, and D. Hernández-Abreu. High order PDE-convergence of AMF-W methods for 2D-linear parabolic problems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003247.

Ghosh:2023:ESS

[GPK23] Debdulal Ghosh, Jayanta Paul, and Jitendra Kumar. On equilibrium solution to a singular coagulation equation with source and efflux. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005076.

Girardi:2020:FEM

[GPP^{+}20] Maria Girardi, Cristina Padovani, Daniele Pellegrini, Margherita Porcelli, and Leonardo Robol. Finite element model updating for structural applications. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306806.

Gorska:2023:VPF

[GPST23] K. Górska, T. Pietrzak, T. Sandev, and Z. Tomovski. Volterra–Prabhakar function of distributed order and some applications. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002509.

Girfoglio:2023:NLE

[GQR23] Michele Girfoglio, Annalisa Quaini, and Gianluigi Rozza. A novel Large Eddy Simulation model for the quasi-geostrophic equations in a finite volume setting. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003326.

Ganesh:2022:EML

[GR22] M. Ganesh and B. Reyes.

An efficient multi-level high-order algorithm for simulation of a class of Allen–Cahn stochastic systems. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003873.

Gratien:2020:RSM

[Gra20] Jean-Marc Gratien. A robust and scalable multilevel domain decomposition preconditioner for multi-core architecture with large number of cores. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719306193.

Galiano:2020:NCN

[GRS20] G. Galiano, I. Ramírez, and E. Schiavi. Non-convex non-local reactive flows for saliency detection and segmentation. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301643.

Geredeli:2023:ICA

[GRVZ23] Pelin G. Geredeli, Leo G. Rebholz, Duygu Vargun, and Ahmed Zytoon. Improved convergence of the Arrow–Hurwicz iteration for the Navier–Stokes equation via grad-div stabilization and Anderson acceleration. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005180.

Grzejda:2021:FEM

[Grz21] Rafał Grzejda. Finite element modeling of the contact of elements preloaded with a bolt and externally loaded with any force. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001539.

Gao:2020:AEF

[GS20a] Wenyou Gao and Chang Su. Analysis of earnings forecast of blockchain financial products based on particle swarm optimization. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300157.

Gao:2020:ABC

[GS20b] Wenyou Gao and Chang Su. Analysis on block chain financial transaction under artificial neural network of deep learning. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030282X.

Giri:2020:NCD

[GS20c] Subhajit Giri and Shuvam Sen. A new class of diagonally implicit Runge– Kutta methods with zero dissipation and minimized dispersion error. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301321.

Giani:2021:SEE

[GS21a] Stefano Giani and Pavel Solin. Solving elliptic eigenproblems with adaptive multimesh hp-FEM. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001473.

Gonzalez:2021:SPE

[GS21b] María González and Magdalena Strugaru. Stabilization and a posteriori error analysis of a mixed FEM for convection–diffusion problems with mixed boundary conditions. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030306X.

Gracia:2021:FDM

[GS21c] José Luis Gracia and Martin Stynes. A finite difference method for an initialboundary value problem with a Riemann–Liouville-Caputo spatial fractional derivative. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303113.

Groselj:2021:SSC

[GS21d] Jan Groselj and Hendrik Speleers. Super-smooth cubic Powell–Sabin splines on three-directional triangulations: B-spline representation and subdivision. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305367.

Gu:2022:GHC

[GS22] Ying Gu and Yongzhong Song. Global Hessenberg and CMRH methods for a class of complex matrix equations. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100491X.

Grover:2024:IEP

[GS24] Lovleen Kumar Grover and Anchal Sharma. Improved estimators of population variance using robust measures in case of missing data. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006714.

Ghosh:2020:EUS

[GSK20] Debdulal Ghosh, Jitraj Saha, and Jitendra Kumar. Existence and uniqueness of steady-state solution to a singular coagulationfragmentation equation. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302831.

Gil:2020:NEA

[GST20] A. Gil, J. Segura, and N. M. Temme. Numerical evaluation of Airy-type integrals arising in uniform asymptotic analysis. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030008X.

Guo:2022:MWG

[GSWH22] Liming Guo, Qiwei Sheng, Cheng Wang, and Ziping Huang. A modified weak Galerkin finite element method for nonmonotone quasilinear elliptic problems. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005483.

Gao:2022:OSB

[GSZ22] Qinjiao Gao, Xingping Sun, and Shenggang Zhang. Optimal stochastic Bernstein polynomials in Ditzian– Totik type modulus of smoothness. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005112.

Groh:2023:EAG

[GT23] Dennis Groh and Caren Tischendorf. Error analysis for Galerkin–BDF discretizations of DAEs with elliptic operator constraints. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005441.

Gu:2021:SCM

[Gu21] Zhendong Gu. Spectral collocation method for nonlinear Riemann–Liouville fractional terminal value problems. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002624.

Guan:2020:WGF

[Gua20] Qingguang Guan. Weak Galerkin finite element method for Poisson’s equation on polytopal meshes with small edges or faces. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719305898.

Guo:2024:CEN

[Guo24] Chun-Hua Guo. Computing the extremal nonnegative solutions of the M-tensor equation with a nonnegative right-side vector. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005034.

Gonzalez-Vega:2021:TAI

[GVT21] Laureano Gonzalez-Vega and Alexandre Trocado. Tools for analyzing the intersection curve between two quadrics through projection and lifting. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001412.

Gao:2023:FEA

[GW23] Guangwei Gao and Shuonan Wu. A C^{0} finite element approximation of planar oblique derivative problems in non-divergence form. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000900.

Guo:2024:EMR

[GW24] Jie Guo and Zhong Wan. An efficient modified residualbased algorithm for large scale symmetric nonlinear equations by approximating successive iterated gradients. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300496X.

Gong:2023:CBG

[GWHZ23] Rongfang Gong, Min Wang, Qin Huang, and Ye Zhang. A CCBMbased generalized GKB iterative regularization algorithm for inverse Cauchy problems. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002261.

Guo:2021:IEA

[GWWY21] Jing Guo, Cheng Wang, Steven M. Wise, and Xingye Yue. An improved error analysis for a second-order numerical scheme for the Cahn– Hilliard equation. Journal of Computational and Applied Mathematics, 388 (??):??, May 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305914.

Goyal:2024:CPP

[GX24] Dheeraj Goyal and Min Xie. On the compound Poisson phase-type process and its application in shock models. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001018.

Ge:2024:URA

[GXC24] Huanmin Ge, Yujia Xie, and Wengu Chen. Uniform RIP analysis for the l_{p}–ωl_{q} minimization. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003492.

Gao:2022:ACD

[GXL22] Qian Gao, Hui Xu, and Aijun Li. The analysis of commodity demand predication in supply chain network based on particle swarm optimization algorithm. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003824.

Guan:2022:LFN

[GYZ22] Hongbo Guan, Yong Yang, and Huiqing Zhu. A locking-free nonconforming FEM for optimal control problems governed by linear elasticity equations. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001261.

Gan:2023:EAS

[GZ23] Di Gan and Guo-Feng Zhang. Efficient ADI schemes and preconditioning for a class of highdimensional spatial fractional diffusion equations with variable diffusion coefficients. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005362.

Gao:2024:CRM

[GZ24] Rong Gao and Shijie Zhang. Computing the reliability of mixed uncertain random k-out-of-n systems with multiple possible states. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005113.

Hesamian:2020:FSU

[HA20a] Gholamreza Hesamian and Mohammad Ghasem Akbari. Fuzzy spline univariate regression with exact predictors and fuzzy responses. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300947.

Hesamian:2020:RVC

[HA20b] Gholamreza Hesamian and Mohammad Ghasem Akbari. A robust varying coefficient approach to fuzzy multiple regression model. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307095.

Hante:2021:RVL

[HA21a] Stefan Hante and Martin Arnold. RATTLie: a variational Lie group integration scheme for constrained mechanical systems. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304959.

Hesamian:2021:RMR

[HA21b] Gholamreza Hesamian and Mohammad Ghasem Akbari. A robust multiple regression model based on fuzzy random variables. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305616.

Hosseinian:2024:ENS

[HAD24] Alireza Hosseinian, Pouria Assari, and Mehdi Dehghan. An efficient numerical scheme to solve generalized Abel’s integral equations with delay arguments utilizing locally supported RBFs. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400116X.

Hamadneh:2022:NOP

[HAFA22] Tareq Hamadneh, Ibraheem Abu-Falahah, and Mohammad Alqudah. Numerical optimization and positivity certificates for polynomials and rationals over simplices. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002035.

Hainaut:2023:PSE

[Hai23] Donatien Hainaut. Pricing of spread and exchange options in a rough jumpdiffusion market. Journal of Computational and Applied Mathematics, 419 (??):??, February 2023.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722003831.

Han:2024:VSS

[Han24] Bin Han. Vector subdivision schemes and their convergence for arbitrary matrix masks. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004223.

Harase:2021:TSP

[Har21]  Shin Harase.  A ta-

ble of short-period Tausworthe generators for Markov chain quasi-Monte Carlo. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304271.

Hofmann:2022:SQH

[HB22] S. Hofmann and A. Borzì. A sequential quadratic Hamiltonian algorithm for training explicit RK neural networks. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005574.

Halada:2023:EBC

[HB23] Tomás Halada and Ludek Benes. Effect of boundary conditions on particle spurious movement in Smoothed Particle Hydrodynamics method. Journal of Computational and Applied Mathematics, 427 (??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723000778.

Hachtel:2021:MIE

[HBGS21] Christoph Hachtel, Andreas Bartel, Michael Günther, and Adrian Sandu. Multirate implicit Euler schemes for a class of differential–algebraic equations of index-1. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305023.

He:2021:IDA

[HBLZ21] Hongjin He, Xueli Bai, Chen Ling, and Guanglu Zhou. An index detecting algorithm for a class of TCP(A, q) equipped with nonsingular M-tensors. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001680.

Husseiny:2022:IMR

[HBMA22] I. A. Husseiny, H. M. Barakat, G. M. Mansour, and M. A. Alawady. Information measures in records and their concomitants arising from Sarmanov family of bivariate distributions. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000231.

Heidler:2021:ELE

[HBPV21] Václav Heidler, Ondrej Bublík, Ales Pecka, and Jan Vimmr. Eulerian– Lagrangian and Eulerian– Eulerian approaches for the simulation of particleladen free surface flows using the lattice Boltzmann method. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002946.

Hu:2024:GGF

[HbX24] Rong Hu and Yi bin Xiao. Generalized Dgap functions for hemivariational inequalities in Hilbert spaces. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006660.

Hsu:2020:DCS

[HC20a] Feng-Jui Hsu and Sheng-Hung Chen. Does corporate social responsibility drive better performance by adopting IFRS? Evidence from emerging market. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306363.

Hu:2020:NST

[HC20b] Hanzhang Hu and Yanping Chen. Numerical solution of two-dimensional nonlinear Schrödinger equation using a new twogrid finite element method. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303309.

Hu:2021:AFE

[HC21] Hanzhang Hu and Yanping Chen. Analysis of finite element two-grid algorithms for two-dimensional nonlinear Schrödinger equation with wave operator. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002697.

Huang:2023:DRM

[HC23] Zhengge Huang and Jingjing Cui. The double-relaxation modulus-based matrix splitting iteration method for linear complementarity problems. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000821.

Huang:2024:SDS

[HC24] Bingru Huang and Falai Chen. On the stability of the dimensions of spline spaces with highest order of smoothness over t-meshes. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006246.

He:2024:PFP

[HCC24] Yuyu He, Hongtao Chen, and Hang Chen. Physical feature preserving and unconditionally stable SAV fully discrete finite element schemes for incompressible flows with variable density. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000773.

He:2021:GMA

[HCCC21] Zhengkang He, Eric T.

Chung, Jie Chen, and Zhangxin Chen. Generalized multiscale approximation of a multipoint flux mixed finite element method for Darcy– Forchheimer model. Journal of Computational and Applied Mathematics, 391 (??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721000856.

Hernandez-Cabrera:2024:SSM

[HCCSCGAM24] Francisco Hernández-Cabrera, Cynthia Elizabeth Castillo-Silva, José Roberto Cantú-González, and Francisco Javier Almaguer-Martínez. S-SPAM; a search method for spatio-temporal patterns of mutations. Study case: a variability analysis in the spike protein of SARS-CoV-2. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004521.

Huang:2020:NIO

[HCCW20] Shian-Chang Huang, Chei-Chang Chiou, Jui-Te Chiang, and Cheng-Feng Wu. A novel intelligent option price forecasting and trading system by multiple kernel adaptive filters. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305655.

He:2024:OGM

[HCCZ24] Zhengkang He, Jie Chen, Zhangxin Chen, and Tong Zhang. An online generalized multiscale approximation of the multipoint flux mixed finite element method. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004429.

Hu:2020:DDS

[HCH20] Xiulin Hu, Yuhao Cong, and Guang-Da Hu. Delaydependent stability of Runge–Kutta methods for linear delay differential– algebraic equations. Journal of Computational and Applied Mathematics, 363 (??):300–311, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0377042719302894.

He:2023:DHL

[HCH^{+}23] Yong He, Peimin Chen, Lin He, Kaili Xiang, and Chunchi Wu. A dynamic Heston localstochastic volatility model and Legendre transform dual-asymptotic solution for optimal investment strategy problems with CARA utility. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200591X.

Hacat:2024:OCP

[HÇYK24] Gülnur Haçat, Aytekin Çibik, Fikriye Yilmaz, and Songül Kaya. On an optimal control problem of the leray-α model. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003631.

He:2020:IPF

[HDW20] Xiaoxiao He, Weibing Deng, and Haijun Wu. An interface penalty finite element method for elliptic interface problems on piecewise meshes. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304765.

He:2023:NMM

[He23]  Yunhui He.  A novel

multigrid method for elliptic distributed control problems. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003922.

Hoang:2024:SON

[HE24] Manh Tuan Hoang and Matthias Ehrhardt. A second-order nonstandard finite difference method for a general Rosenzweig– MacArthur predator–prey model. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000013.

Hessari:2020:LSS

[Hes20] Peyman Hessari. Least squares spectral method for the two-dimensional Stokes interface problems. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930322X.

Hamidi:2024:MFV

[HET24] Sadiq Hamidi, Mustapha El Ossmani, and Abdelaziz Taakili. A monotone finite volume scheme for single phase flow with reactive transport in anisotropic porous media. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005708.

Holman:2021:NSS

[HF21] Jirí Holman and Jirí Fürst. Numerical simulation of separation induced laminar to turbulent transition over an airfoil. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001497.

Holman:2023:RHR

[HF23] Jirí Holman and Jirí Fürst. Rotated-hybrid Riemann solver for all-speed flows. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000730.

Hao:2024:CBS

[HF24] Yong-Xia Hao and Wen-Qing Fei. Construction of Bézier surfaces with minimal quadratic energy for given diagonal curves. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001031.

Hesse:2022:EEP

[HG22] Kerstin Hesse and Quoc Thong Le Gia. L_{2} error estimates for polynomial discrete penalized leastsquares approximation on the sphere from noisy data. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200022X.

Hironaka:2023:EEN

[HG23] Tomohiko Hironaka and Takashi Goda. An efficient estimation of nested expectations without conditional sampling. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004125.

Hu:2020:NOF

[HH20] Guang-Da Hu and Renhong Hu. Numerical optimization for feedback stabilization of linear systems with distributed delays. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719307113.

Han:2021:REA

[HH21a] Yongbin Han and Yanren Hou. Robust error analysis of H (div)-conforming DG method for the timedependent incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306567.

Heitsch:2021:EFS

[HH21b] Holger Heitsch and René Henrion. An enumerative formula for the spherical cap discrepancy. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000285.

Hu:2021:DDA

[HH21c] Peng Hu and Chengming Huang. Delay dependent asymptotic mean square stability analysis of the stochastic exponential Euler method. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303599.

He:2023:MAE

[HH23] Yanyan He and M. Yousuff Hussaini. Mixed aleatory and epistemic uncertainty propagation using Dempster– Shafer theory. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001784.

Herrero-Hervas:2024:EIG

[HH24a] Federico Herrero-Hervás. An explicit–implicit generalized finite difference scheme for a parabolic– elliptic density-suppressed motility system. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001249.

Hunter:2024:SSR

[HH24b] Thomas P. Hunter and Steven J. Hulshoff. Super-Adjoint: Super-resolution neural networks in adjointbased error estimation. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723006659.

Huynh:2024:ASQ

[HH24c] Duc Quoc Huynh and Feng-Nan Hwang. An accelerated structured quasi-Newton method with a diagonal second-order Hessian approximation for nonlinear least squares problems. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006611.

Huang:2023:TIP

[HHD23] Huayi Huang, Yunqing Huang, and Xiaojing Dong. Three interior penalty DG methods for stationary incompressible magnetohydrodynamics. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006288.

Hua:2020:FDI

[HHL20] Hao Hua, Ludger Hovestadt, and Biao Li. Fractional derivative for interpolation in Rn and SO(n) applications in functionally graded materials and rigid body transformations. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302284.

Hansen:2022:GMT

[HHM22a] Per Christian Hansen, Ken Hayami, and Keiichi Morikuni. GMRES methods for tomographic reconstruction with an unmatched back projector. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200156X.

He:2022:MFM

[HHM22b] Cuiyu He, Xiaozhe Hu, and Lin Mu. A meshfree method using piecewise deep neural network for elliptic interface problems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001595.

Hoang:2024:MFS

[HIM24] Hieu T. Hoang, Ferenc Izsák, and Gábor Maros. Method of fundamental solutions: New approximation results and applications. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001845.

Hached:2020:NMD

[HJ20] M. Hached and K. Jbilou.

Numerical methods for differential linear matrix equations via Krylov subspace methods. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930679X.

Han:2023:SBB

[HJ23] Qiang Han and Shaolin Ji. Solving BSDEs based on novel multi-step schemes and multilevel Monte Carlo. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002679.

Hesamian:2024:EFL

[HJC24]  Gholamreza  Hesamian,

Arne Johannssen, and Nataliya Chukhrova. An explainable fused lasso regression model for handling high-dimensional fuzzy data. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006647.

Hayotov:2020:OQF

[HJL20] Abdullo R. Hayotov, Soomin Jeon, and Chang-Ock Lee. On an optimal quadrature formula for approximation of Fourier integrals in the space L_{2}(^{1}). Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300042.

Han:2024:NAS

[HJM24] Wei-Wei Han, Yao-Lin Jiang, and Zhen Miao. Numerical analysis of the SAV scheme for the EMAC formulation of the timedependent Navier–Stokes equations. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004582.

Han:2020:NSH

[HJO20] Weimin Han, Michal Jureczka, and Anna Ochal. Numerical studies of a hemivariational inequality for a viscoelastic contact problem with damage. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301771.

Hayotov:2021:AOQ

[HJS21] Abdullo R. Hayotov, Soomin Jeon, and Kholmat M. Shadimetov. Application of optimal quadrature formulas for reconstruction of CT images. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030604X.

Hosseini:2021:APL

[HK21] Fatemeh Hosseini and Omid Karimi. Approximate pairwise likelihood inference in SGLM models with skew normal latent variables. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003149.

Hung:2023:GWP

[HK23] Nguyen Van Hung and André A. Keller. Generalized well-posedness for parametric fuzzy generalized multiobjective games. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722005155.

Kang:2021:ESM

[hKhWjH21] Jian hao Kang, Ming hui Wang, and Nan jing Huang. Equilibrium strategy for mean-variance-utility portfolio selection under Heston’s SV model. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001096.

Huang:2021:CIR

[HKL21] Weizhang Huang, Lennard Kamenski, and Jens Lang. Conditioning of implicit Runge–Kutta integration for finite element approximation of linear diffusion equations on anisotropic meshes. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930500X.

Hwang:2024:NHR

[HKM24] Jinmi Hwang, Sejong Kim, and Vatsalkumar N. Mer. Non-homogeneous Riemannian gradient equations for sum of squares of Bures–Wasserstein metric. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004995.

Hnatushenko:2022:VAR

[HKU22] Volodymyr Hnatushenko, Peter Kogut, and Mykola Uvarov. Variational approach for rigid co-registration of optical/SAR satellite images in agricultural areas. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003642.

Hainaut:2021:OPI

[HL21a] Donatien Hainaut and Nikolai Leonenko. Option pricing in illiquid markets: a fractional jump-diffusion approach. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302867.

Hou:2021:TFO

[HL21b] Baohui Hou and Dong Liang. Time fourth-order energy-preserving AVF finite difference method for nonlinear space-fractional wave equations. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305185.

Huang:2021:PPI

[HL21c] Peiqi Huang and Zhilin Li. Partially penalized IFE methods and convergence analysis for elasticity interface problems. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303502.

Hanbali:2022:MTF

[HL22a] Hamza Hanbali and Daniël Linders. Monotone tail functions: Definitions, properties, and application to risk-reducing strategies. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002345.

Hou:2022:IID

[HL22b] Ruizhi Hou and Fang Li.

IDPCNN: Iterative denoising and projecting CNN for MRI reconstruction. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005719.

Huang:2022:MSL

[HL22c] Baohua Huang and Wen Li. A modified SORlike method for absolute value equations associated with second order cones. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003678.

Han:2023:IRL

[HL23] Jihun Han and Yoonsang Lee. Inhomogeneous regularization with limited and indirect data. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001371.

He:2024:CAF

[HL24a] Mingyu He and Wenyuan Liao. A compact ADI finite difference method for 2D reaction–diffusion equations with variable diffusion coefficients. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003448.

Hou:2024:EIS

[HL24b] Zhisong Hou and Sanyang Liu. An efficient image space branch-reduction-bound algorithm to globally solve generalized fractional programming problems for large-scale real applications. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003200.

Huang:2022:CCP

[HLA22] Ziyang Huang, Guang Lin, and Arezoo M. Ardekani. A consistent and conservative phase-field method for multiphase incompressible flows. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000218.

Huang:2019:MAZ

[HLF19] Yunqing Huang, Jichun Li, and Zhiwei Fang. Mathematical analysis of Ziolkowski’s PML model with application for wave propagation in metamaterials. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304376.

He:2023:MOR

[HLLL23] Xiao-Feng He, Liang Li, Stéphane Lanteri, and Kun Li. Model order reduction for parameterized electromagnetic problems using matrix decomposition and deep neural networks. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002157.

Huang:2022:DGD

[HLLQ22] Yunqing Huang, Jichun Li, Chanjie Li, and Kai Qu. Discontinuous Galerkin discretizations and analysis for the Cohen–Monk PML model. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006063.

Huang:2021:ECI

[HLLW21] Tsung-Ming Huang, Weichien Liao, Wen-Wei Lin, and Weichung Wang. An efficient contour integral based eigensolver for 3D dispersive photonic crystal. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002016.

He:2024:DTD

[HLLY24] Bin He, Fuhao Liu, Xin Liu, and Wei Yang. Design and time-domain finite element simulation of multifunctional transformation optical device. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002309.

He:2024:LGA

[HLLZ24] Jian He, Jinlin Li, Zhenrong Lu, and Bangzhong Zhang. Linearized generalized ADMM-based algorithm for multi-block linearly constrained separable convex programming in real-world applications. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005769.

Hu:2024:PRE

[HLMY24] Xiaozhe Hu, Seulip Lee, Lin Mu, and Son-Young Yi. Pressure-robust enriched Galerkin methods for the Stokes equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300393X.

Hu:2022:FPP

[HLR22] Wenyu Hu, Yao Lu, and Jin Ren. A fixed-point proximity algorithm for recovering low-rank components from incomplete observation data with application to motion capture data refinement. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200084X.

He:2020:PLS

[HLW20] Jun He, Chaoqian Li, and Yimin Wei. Pseudospectra localization sets of tensors with applications. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305850.

Hao:2022:NSH

[HLW22] Yongle Hao, Siyu Liu, and Lin Wang. Numerical solutions for Helmholtz equation with stochastic interface based on PML method. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005008.

Huang:2022:TRM

[HLY22] Guangxin Huang, Yuanyuan Liu, and Feng Yin. Tikhonov regularization with MTRSVD method for solving largescale discrete ill-posed problems. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005690.

Huang:2024:AAT

[HLYZ24] Yunqing Huang, Jichun Li, Xuancen Yi, and Haoke Zhao. Analysis and application of a time-domain finite element method for the Drude metamaterial perfectly matched layer model. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005198.

He:2020:CPD

[HLZ20] Jia Wei He, Carlos Lizama, and Yong Zhou. The Cauchy problem for discrete time fractional evolution equations. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306880.

Hanke:2021:CAL

[HM21a] Michael Hanke and Roswitha März. Convergence analysis of least-squares collocation methods for nonlinear higher-index differential– algebraic equations. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305199.

Hanke:2021:RDN

[HM21b] Michael Hanke and Roswitha März. A reliable direct numerical treatment of differential–algebraic equations by overdetermined collocation: an operator approach. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305254.

Heuchenne:2022:USF

[HM22] Cédric Heuchenne and Gilles Mordant. Using space filling curves to compare two multivariate distributions with distribution-free tests. Journal of Computational and Applied Mathematics, 416 (??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722002394.

Hu:2023:EUS

[HM23] Yu Hu and A. J. Meir. Existence and uniqueness of solutions of the equations of quasistatic electroporoelasticity. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002005.

Hazarika:2024:AGH

[HMA24] Bipan Hazarika, Giriraj Methi, and Rupal Aggarwal. Application of generalized Haar wavelet technique on simultaneous delay differential equations. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002279.

Hayami:2023:RAN

[HMRZ23] Ken Hayami, Keiichi Morikuni, Lothar Reichel, and Ning Zheng. Recent advances in the numerical solution of matrix problems.

Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001838.

Haramoto:2022:UPX

[HMS22] Hiroshi Haramoto, Makoto Matsumoto, and Mutsuo Saito. Unveiling patterns in xorshift128+ pseudorandom number generators. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004131.

Horvath:2023:NIL

[HMT23] Illés Horváth, András Mészáros, and Miklós Telek. Numerical inverse Laplace transformation beyond the Abate–Whitt framework. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003296.

Hantke:2023:NMD

[HMWY23] Maren Hantke, Christoph Matern, Gerald Warnecke, and Hazem Yaghi. A new method to discretize a model for isothermal flow with a multi-component equation of state. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004745.

Hashimoto:2022:PTK

[HN22a] Yuka Hashimoto and Takashi Nodera. A preconditioning technique for Krylov subspace methods in RKHSs. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002370.

Hutzenthaler:2022:SCR

[HN22b] Martin Hutzenthaler and Tuan Anh Nguyen. Strong convergence rate of Euler– Maruyama approximations in temporal-spatial Höldernorms. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001790.

Hakula:2023:CCP

[HNV23] Harri Hakula, Mohamed M. S. Nasser, and Matti Vuorinen. Conformal capacity and polycircular domains. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004083.

Hegarty:2022:NAC

[HO22] A. F. Hegarty and E. O’Riordan.

A numerical algorithm to computationally solve the hemker problem using Shishkin meshes. Journal of Computational and Applied Mathematics, 409 (??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722000413.

Hioual:2023:NNV

[HOGO23] Amel Hioual, Adel Ouannas, Giuseppe Grassi, and Taki-Eddine Oussaeif. Nonlinear nabla variableorder fractional discrete systems: Asymptotic stability and application to neural networks. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005374.

Hon:2022:OBC

[Hon22] Sean Hon. Optimal block circulant preconditioners for block Toeplitz systems with application to evolutionary PDEs. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005689.

Hoshi:2020:PVM

[HOOT20] Takeo Hoshi, Takeshi Ogita, Katsuhisa Ozaki, and Takeshi Terao. An a posteriori verification method for generalized real-symmetric eigenvalue problems in large-scale electronic state calculations. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301217.

Horvath:2020:SDQ

[Hor20]  Róbert Horváth.  On

some discrete qualitative properties of implicit finite difference solutions of nonlinear parabolic problems. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303279.

Herceg:2022:CVD

[HP22] D. Herceg and I. Petković.

Computer visualization and dynamic study of new families of rootsolvers. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003976.

Hoppe:2022:STA

[HPW22] R. H. W. Hoppe, B. Pahari, and J. J. Winkle. Space-time adaptive splitting scheme for the numerical simulation of polycrystallization. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005057.

Huang:2020:VOF

[HPWM20] Lan-Lan Huang, Ju H. Park, Guo-Cheng Wu, and Zhi-Wen Mo. Variableorder fractional discretetime recurrent neural networks. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306387.

Huang:2021:ADL

[HQL21] Anzhong Huang, Lening Qiu, and Zheng Li. Applying deep learning method in TVP-VAR model under systematic financial risk monitoring and early warning. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303563.

Hu:2020:CBF

[HQS20] Kaibo Hu, Weifeng Qiu, and Ke Shi. Convergence of a B–E based finite element method for MHD models on Lipschitz domains. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304807.

Hong:2020:RDS

[HQT20] Liu Hong, Hong Qian, and Lowell F. Thompson. Representations and divergences in the space of probability measures and stochastic thermodynamics. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301333.

Higueras:2021:ESL

[HR21] I. Higueras and T. Roldán.

Efficient SSP low-storage Runge–Kutta methods. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305035.

Heydari:2022:JSM

[HR22] M. H. Heydari and M. Razzaghi. Jacobi spectral method for variableorder fractional Benney-Lin equation arising in falling film problems. Journal of Computational and Applied Mathematics, 402 (??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721004350.

Heydari:2023:HMB

[HR23] M. H. Heydari and M. Razzaghi. A hybrid method based on the Chebyshev cardinal functions/wavelets for time fractional coupled Klein–Gordon-Schrödinger equations. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000869.

Hristov:2022:IBM

[Hri22] Jordan Hristov. Integralbalance method with transmuted profiles: Concept, examples, and emerging problems. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002692.

Hausenblas:2020:TSN

[HRT20] Erika Hausenblas, Tsiry Avisoa Randrianasolo, and Mechtild Thalhammer. Theoretical study and numerical simulation of pattern formation in the deterministic and stochastic Gray–Scott equations. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303322.

Huang:2020:OSN

[HS20] Chaobao Huang and Martin Stynes. Optimal spatial H_{1}-norm analysis of a finite element method for a time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304388.

Hu:2023:EED

[HS23] Yaozhong Hu and Neha Sharma. Ergodic estimators of double exponential Ornstein–Uhlenbeck processes. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300273X.

Hasegawa:2024:AQM

[HS24a] Takemitsu Hasegawa and Hiroshi Sugiura. An automatic quadrature method for semi-infinite integrals of exponentially decaying functions and its Matlab code. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003941.

Henriquez:2024:UHM

[HS24b] Esteban Henríquez and Manuel Solano. An unfitted HDG method for a distributed optimal control problem. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006465.

Hawkins:2024:EIM

[HSB24] Stuart C. Hawkins, Linda Stals, and Sherwin Bagheri. An efficient iterative method for reconstructing the refractive index in complex domains from far field data. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005174.

Huang:2021:RFE

[HSC21] Chaobao Huang, Martin Stynes, and Hu Chen. An α-robust finite element method for a multiterm time-fractional diffusion problem. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306257.

Hernandez-Sanjaime:2019:MSE

[HSGLE19] Rocío Hernández-Sanjaime, Martín González, and Jose J. López-Espín. Multilevel simultaneous equation model: a novel specification and estimation approach. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303814.

Hieu:2020:SCA

[HSM20] Dang Van Hieu, Jean Jacques Strodiot, and Le Dung Muu. Strongly convergent algorithms by using new adaptive regularization parameter for equilibrium problems. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301357.

Heyouni:2020:TFG

[HSMT20] Mohammed Heyouni, Farid Saberi-Movahed, and Azita Tajaddini. A tensor format for the generalized Hessenberg method for solving Sylvester tensor equations. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301692.

Hosseini:2024:SAE

[HSRA24] Seyede Elahe Hosseini, Davood Shahsavani, Mohammad Reza Rabiei, and Mohammad Arashi. Small area estimation with partially linear mixed-t model with measurement error. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001201.

Hesse:2021:LRB

[HSW21] Kerstin Hesse, Ian H.

Sloan, and Robert S. Womersley. Local RBFbased penalized leastsquares approximation on the sphere with noisy scattered data. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303526.

Huang:2024:EAF

[HSYZ24] Cunxin Huang, Haiming Song, Jinda Yang, and Bocheng Zhou. Error analysis of finite difference scheme for American option pricing under regime-switching with jumps. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004284.

Ho:2021:DBA

[HTD21] Vinh Thanh Ho, Hoai An Le Thi, and Tao Pham Dinh. DCA-based algorithms for DC fitting. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306440.

Huntul:2024:IPR

[HTIA24] M. J. Huntul, I. Tekin, Muhammad Kashif Iqbal, and Muhammad Abbas. An inverse problem of recovering the heat source coefficient in a fourth-order time-fractional pseudoparabolic equation. Journal of Computational and Applied Mathematics, 442 (??):??, May 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006556.

Hesamian:2024:FNR

[HTJC24]  Gholamreza  Hesamian,

Faezeh Torkian, Arne Johannssen, and Nataliya Chukhrova. A fuzzy nonparametric regression model based on an extended center and range method. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003217.

Hung:2020:NCG

[HTOC20] Nguyen Van Hung, Vo Minh Tam, Donal O’Regan, and Yeol Je Cho. A new class of generalized multiobjective games in bounded rationality with fuzzy mappings: structural (λ, ϵ)-stability and (λ, ϵ)-robustness to ϵ-equilibria. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300261.

Hung:2020:CAS

[HTTO20] Nguyen Van Hung, Vo Minh Tam, Nguyen Huy Tuan, and Donal O’Regan. Convergence analysis of solution sets for fuzzy optimization problems. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930620X.

Hung:2021:NCS

[HTZ21] Nguyen Van Hung, Vo Minh Tam, and Yong Zhou. A new class of strong mixed vector GQVIP-generalized quasi-variational inequality problems in fuzzy environment with regularized gap functions based error bounds. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303460.

Hu:2022:ADS

[Hu22] Guang-Da Hu. An algorithm to design a stabilizing feedback controller for linear delay systems via Parseval ’s theorem. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002436.

Hu:2024:FDH

[Hu24] Dongdong Hu. Fully decoupled and high-order linearly implicit energypreserving RK-GSAV methods for the coupled nonlinear wave equation. Journal of Computational and Applied Mathematics, 445 (??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724000852.

Huang:2021:EBS

[Hua21] Zheng-Ge Huang. Efficient block splitting iteration methods for solving a class of complex symmetric linear systems. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001941.

Hung:2021:GLP

[Hun21] Nguyen Van Hung. Generalized Levitin–Polyak wellposedness for controlled systems of FMQHI-fuzzy mixed quasi-hemivariational inequalities of Minty type. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042720305549.

Hung:2022:LWP

[Hun22] Nguyen Van Hung. LP well-posed controlled systems for bounded quasiequilibrium problems and their application to traffic networks. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004143.

Hurisse:2022:SIF

[Hur22] Olivier Hurisse. A semiimplicit fractional step algorithm on staggered meshes for simulating a compressible two-layer mixed-flow model. Journal of Computational and Applied Mathematics, 412 (??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722001364.

He:2023:IMB

[HV23] Jiewen He and Seakweng Vong. Improved modulusbased matrix splitting iteration methods for quasicomplementarity problems. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300225X.

Hernandez-Veron:2022:ISS

[HVM22] M. A. Hernández-Verón and Eulalia Martínez. Iterative schemes for solving the Chandrasekhar Hequation using the Bernstein polynomials. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000108.

Hernandez-Veron:2022:RTS

[HVMS22] M. A. Hernández-Verón, Eulalia Martínez, and Sukhjit Singh. A reliable treatment to solve nonlinear Fredholm integral equations with nonseparable kernel. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304064.

Hernandez-Veron:2022:EPC

[HVR22a] M. A. Hernández-Verón and N. Romero. An efficient predictor-corrector iterative scheme for solving Wiener–Hopf problems. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001746.

Hernandez-Veron:2022:SWH

[HVR22b] M. A. Hernández-Verón and N. Romero. Solving Wiener–Hopf problems via an efficient iterative scheme. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303745.

Hong:2020:FVE

[HW20] Qi Hong and Jiming Wu. A Q_{1}-finite volume element scheme for anisotropic diffusion problems on general convex quadrilateral mesh. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300236.

Han:2021:FMG

[HW21a] Huan Han and Andong Wang. A fast multi grid algorithm for 2D diffeomorphic image registration model. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001965.

Han:2021:NAP

[HW21b] Weimin Han and Cheng Wang. Numerical analysis of a parabolic hemivariational inequality for semipermeable media. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306178.

Han:2022:TEA

[HW22] Feiyang Han and Yimin Wei. TLS–EM algorithm of Mixture Density Models for exponential families. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004519.

Hozumi:2024:ASC

[HW24a] Yuta Hozumi and Guo-Wei Wei. Analyzing single cell RNA sequencing with topological nonnegative matrix factorization. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000918.

Hu:2024:MRK

[HW24b] Guang-Da Hu and Zheng Wang. A modified Runge– Kutta method for increasing stability properties. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006416.

Hussain:2022:SMP

[HWAT22] Saqib Hussain, Xiaoshen Wang, and Ahmed Al-Taweel. A study of mixed problem for second order elliptic problems using modified weak Galerkin finite element method. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003927.

Hao:2021:GLO

[HWCJ21] Zijun Hao, Zhongping Wan, Xiaoni Chi, and Zheng-Fen Jin. Generalized lower-order penalty algorithm for solving secondorder cone mixed complementarity problems. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304593.

Hu:2024:PPM

[HWD^{+}24] Xingwei Hu, Mengjie Wang, Xinjie Dai, Yanyan Yu, and Aiguo Xiao. A positivity preserving Milstein-type method for stochastic differential equations with positive solutions. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002139.

He:2021:SDD

[HWG21] Liangtian He, Yilun Wang, and Shaobing Gao. A support-denoiser-driven framework for single image restoration. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100114X.

Hu:2021:VPA

[HWH^{+}21] Qian-Ying Hu, Chun Wen, Ting-Zhu Huang, Zhao-Li Shen, and Xian-Ming Gu. A variant of the Power–Arnoldi algorithm for computing PageRank. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303253.

Hu:2023:ISA

[HWLD23] Shaotao Hu, Yuanheng Wang, Liya Liu, and Qiao-Li Dong. An inertial self-adaptive iterative algorithm for finding the common solutions to split feasibility and fixed point problems in specific Banach spaces. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006082.

Hogeme:2024:SNM

[HWRM24] Mohammed Sumebo Hogeme, Mesfin Mekuria Woldaregay, Laxmi Rathour, and Vishnu Narayan Mishra. A stable numerical method for singularly perturbed Fredholm integro differential equation using exponentially fitted difference method. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006520.

Hu:2022:PET

[HX22] Yaozhong Hu and Yuejuan Xi. Parameter estimation for threshold Ornstein–Uhlenbeck processes from discrete observations. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200108X.

Han:2021:PPC

[HY21a] Xuli Han and Jing Yang. Piecewise polynomial curves with normalized derivatives. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305811.

Huang:2021:MEA

[HY21b] Jianguo Huang and Yue Yu. A medius error analysis for nonconforming virtual element methods for Poisson and biharmonic equations. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305203.

Yin:2021:GHC

[hYbJzJ21] Jiang hua Yin, Jin bao Jian, and Xian zhen Jiang. A generalized hybrid CGPM-based algorithm for solving large-scale convex constrained equations with applications to image restoration. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100042X.

He:2020:CAA

[HYW20] Bin He, Wei Yang, and Hao Wang. Convergence analysis of adaptive edge finite element method for variable coefficient timeharmonic Maxwell’s equations. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301515.

Han:2021:HES

[HYY21] Shizhuan Han, Qiongwei Ye, and Xiaofeng Yang. Highly efficient and stable numerical algorithm for a two-component phasefield crystal model for binary alloys. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306622.

Hasan:2023:CCH

[HZ23] Md Mahmudul Hasan and Xianyi Zeng. A central compact hybrid-variable method with spectrallike resolution: Onedimensional case. Journal of Computational and Applied Mathematics, 421 (??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722004927.

Han:2024:NMR

[HZ24a] Zhong-Cheng Han and Yan-Yong Zhao. Nonparametric modal regression with mixed variables and application to analyze the GDP data. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000906.

Huang:2024:RMF

[HZ24b] Xuehai Huang and Chao Zhang. Robust mixed finite element methods for a quad-curl singular perturbation problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003662.

Hao:2023:DAV

[HZC^{+}23] Yajuan Hao, Meihua Zhang, Yuhuan Cui, Gang Cheng, Jiaquan Xie, and Yiming Chen. Dynamic analysis of variable fractional order cantilever beam based on shifted Legendre polynomials algorithm. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005507.

Hu:2024:HOS

[HZL24] Wenkai Hu, Jicheng Zhang, and Xin Li. Higher order stable generalized isogeometric analysis for interface problems. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000414.

Hendy:2022:DSC

[HZMD22] A. S. Hendy, Mahmoud A.

Zaky, and J. E. Macías-Díaz. On the dissipativity of some Caputo time-fractional subdiffusion models in multiple dimensions: Theoretical and numerical investigations. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003708.

Hu:2024:EDD

[HZS^{+}24] Xuelong Hu, Yitian Zhao, Guan Sun, Jiujun Zhang, Chao Jiang, and Xingfang Huang. Evaluation of different DEWMA schemes for monitoring the ratio of two normal random variables. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002449.

Hu:2021:RHP

[HZZ21] Beibei Hu, Ling Zhang, and Ning Zhang. On the Riemann–Hilbert problem for the mixed Chen– Lee–Liu derivative nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000121.

Hu:2024:CAM

[HZZ24] Bingqing Hu, Wei Zhang, and Xuan Zhao. Convergence analysis of the maximum principle preserving BDF2 scheme with variable time-steps for the space fractional Allen– Cahn equation. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002012.

Ibrahim:2024:TSI

[IAH24] Abdulkarim Hassan Ibrahim and Suliman Al-Homidan. Two-step inertial derivativefree projection method for solving nonlinear equations with application. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003212.

Ibrahimoglu:2020:FAC

[Ibr20] B. Ali Ibrahimoglu. A fast algorithm for computing the mock-Chebyshev nodes. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303334.

Iakymchuk:2020:RSP

[IBW^{+}20] Roman Iakymchuk, Maria Barreda, Matthias Wiesenberger, José I. Aliaga, and Enrique S. Quintana-Ortí. Reproducibility strategies for parallel Preconditioned Conjugate Gradient. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719307022.

Iqbal:2020:ESV

[IHK20] Iram Iqbal, Nawab Hussain, and Marwan A. Kutbi. Existence of the solution to variational inequality, optimization problem, and elliptic boundary value problem through revisited best proximity point results. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300959.

Iudin:2021:NIK

[IIS^{+}21] Dmitry I. Iudin, Fedor D.

Iudin, Artem A. Syssoev, Vitaly Yu. Klimashov, and Alexey A. Emelyanov. Noise-induced kinetic transition in two-component environment. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305598.

Ivanov:2023:NUP

[IKL23] Maksim I. Ivanov, Igor A.

Kremer, and Yuri M. Laevsky. On non-uniqueness of pressures in problems of fluid filtration in fracturedporous media. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006501.

Ivanov:2024:EIS

[IKL24] Maksim I. Ivanov, Igor A.

Kremer, and Yuri M. Laevsky. Explicit-implicit schemes for non-isothermal filtration problem: Singletemperature model. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005836.

Ivicic:2024:PBS

[IM24] Ivana Kuzmanović Ivicić and Suzana Miodragović. Perturbation bounds for stable gyroscopic systems in motion about an unstable equilibrium position. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400311X.

Iavernaro:2021:CHO

[IMMS21] F. Iavernaro, F. Mazzia, M. S. Mukhametzhanov, and Ya. D. Sergeyev. Computation of higher order Lie derivatives on the infinity computer. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030426X.

Imakura:2020:VPE

[IMT20] Akira Imakura, Keiichi Morikuni, and Akitoshi Takayasu. Verified partial eigenvalue computations using contour integrals for Hermitian generalized eigenproblems. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305485.

Imakura:2023:VEE

[IMT23] Akira Imakura, Keiichi Morikuni, and Akitoshi Takayasu. Verified eigenvalue and eigenvector computations using complex moments and the Rayleigh–Ritz procedure for generalized Hermitian eigenvalue problems. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005921.

Iglewska-Nowak:2022:SWW

[IN22] I. Iglewska-Nowak. Spin weighted wavelets on the sphere-frames. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006336.

Irandoust-pakchin:2020:CNG

[IpAmK20] Safar Irandoust-pakchin, Somayeh Abdi-mazraeh, and Hossein Kheiri. Construction of new generating function based on linear barycentric rational interpolation for numerical solution of fractional differential equations. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030090X.

Imakura:2023:CMB

[IS23] Akira Imakura and Tetsuya Sakurai. Complex moment-based eigensolver coupled with two Krylov subspaces. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002273.

Izuchukwu:2023:TSI

[ISD23] Chinedu Izuchukwu, Yekini Shehu, and Qiao-Li Dong. Two-step inertial forwardreflected-backward splitting based algorithm for nonconvex mixed variational inequalities. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000377.

Istafa:2023:NMF

[IuR23] Ghafirlia Istafa and Mujeeb ur Rehman. A numerical method for fractional Sturm–Liouville problems involving the Cauchy-Euler operators. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001656.

Ixaru:2024:EFI

[Ixa24] L. Gr. Ixaru. Exponential fitting for interpolation of oscillatory functions. A numerical approach. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004235.

Ishii:2020:CAV

[IY20] Daisuke Ishii and Tomohito Yabu. Computerassisted verification of four interval arithmetic operators. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301849.

Iqbal:2023:NSN

[IYA^{+}23] Muhammad Sajid Iqbal, Muhammad Waqas Yasin, Nauman Ahmed, Ali Akgül, Muhammad Rafiq, and Ali Raza. Numerical simulations of nonlinear stochastic Newell– Whitehead–Segel equation and its measurable properties. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003089.

Jaradat:2022:GPT

[JA22] Imad Jaradat and Marwan Alquran. Geometric perspectives of the twomode upgrade of a generalized Fisher–Burgers equation that governs the propagation of two simultaneously moving waves. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005318.

Jasinski:2022:NFC

[Jas22] Krzysztof Jasiński. On the number of failed components in a coherent system consisting of multiple types of components. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000632.

Jasinski:2024:SNF

[Jas24] Krzysztof Jasiński. A study on the number of failed components in a failed coherent system consisting of different types of components. Journal of Computational and Applied Mathematics, 435(??):??, ???? 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200437X.

Jafari:2021:IPI

[JB21] A. A. Jafari and S. Bafekri. Inferences on the performance index of Weibull distribution based on k-record values. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303514.

Jajarmi:2022:ASA

[JBSN22] Amin Jajarmi, Dumitru Baleanu, Samaneh Sadat Sajjadi, and Juan J. Nieto. Analysis and some applications of a regularized Ψ-Hilfer fractional derivative. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002291.

Jung:2024:GAI

[JC24] Dohee Jung and Changbum Chun. A general approach for improving the Padé iterations for the matrix sign function. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002923.

Jornet:2020:SOL

[JCLC20] M. Jornet, J. Calatayud, O. P. Le Maître, and J.- C. Cortés. Second order linear differential equations with analytic uncertainties: stochastic analysis via the computation of the probability density function. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300613.

Jeon:2020:AVE

[JCY20] Junkee Jeon, Sun-Yong Choi, and Ji-Hun Yoon. Analytic valuation of European continuous-installment barrier options. Journal of Computational and Applied Mathematics, 363(??):392– 412, January 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0377042719302948.

Jana:2022:EPM

[JD22] Nabakumar Jana and Santanu Dey. Estimating parameters of mixtures of multivariate t-populations and application to classification of observations. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002667.

Jha:2022:WPM

[JDBT22] Shalini Jha, Prasun Das, Sanghamitra Bandhyopadhyay, and Savin Treanta. Well-posedness for multitime variational inequality problems via generalized monotonicity and for variational problems with multi-time variational inequality constraints. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721006075.

Jimenez:2020:ECP

[JdlCD20] J. C. Jimenez, H. de la Cruz, and P. A. De Maio. Efficient computation of phi-functions in exponential integrators. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300492.

Jeannerod:2020:RAX

[Jea20] Claude-Pierre Jeannerod.

The relative accuracy of (x + y) ∗ (x − y). Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306181.

Jia:2022:DSC

[JF22] Xiaofeng Jia and Hui Feng. Decoupled stabilized Crank–Nicolson LeapFrog method for timedependent Navier–Stokes/ Darcy model. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004155.

Jiang:2021:SNM

[JG21a] H. Jiang and N. L. Gibson. Semismooth Newton methods with a shootinglike technique for solving a constrained free-boundary HJB equation. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000479.

Jiang:2021:NPF

[JG21b] Meng-Jiao Jiang and Xue-Ping Guo. A new parameter-free method for Toeplitz systems of weakly nonlinear equations. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001242.

Jia:2020:ANF

[JGLH20] Hongen Jia, Yayu Guo, Jichun Li, and Yunqing Huang. Analysis of a novel finite element method for a modified Cahn–Hilliard–Hele–Shaw system. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301370.

Jia:2023:CAS

[JGW^{+}23] Junhua Jia, Xiucui Guan, Hui Wang, Binwu Zhang, and Panos M. Pardalos. Combinatorial algorithms for solving the restricted bounded inverse optimal value problem on minimum spanning tree under weighted l_{∞} norm. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003843.

Jia:2024:CAR

[JGZ^{+}24] Junhua Jia, Xiucui Guan, Binwu Zhang, Xinqiang Qian, and Panos M. Pardalos. Combinatorial algorithms for restricted inverse optimal value problems on minimum spanning tree under weighted l_{1} norm. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003595.

Jiang:2024:ARI

[JH24] Xianzhen Jiang and Zefeng Huang. An accelerated relaxed-inertial strategy based CGP algorithm with restart technique for constrained nonlinear pseudomonotone equations to image de-blurring problems. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001365.

Jian:2020:FII

[JHG^{+}20] Huan-Yan Jian, Ting-Zhu Huang, Xian-Ming Gu, Xi-Le Zhao, and Yong-Liang Zhao. Fast implicit integration factor method for nonlinear space Riesz fractional reaction–diffusion equations. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302260.

Jia:2021:IAP

[JHW21] Zehui Jia, Jieru Huang, and Zhongming Wu. An incremental aggregated proximal ADMM for linearly constrained nonconvex optimization with application to sparse logistic regression problems. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000030.

Jian:2024:SAL

[JHYZ24] Jinbao Jian, Qiongxuan Huang, Jianghua Yin, and Wei Zhang. Splitting augmented Lagrangian-type algorithms with partial quadratic approximation to solve sparse signal recovery problems. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400222X.

Jiang:2020:MDI

[JHZD20] Tai-Xiang Jiang, Ting-Zhu Huang, Xi-Le Zhao, and Liang-Jian Deng. Multidimensional imaging data recovery via minimizing the partial sum of tubal nuclear norm. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306855.

Jing:2020:APE

[JHZY20] Feifei Jing, Weimin Han, Yongchao Zhang, and Wenjing Yan. Analysis of an a posteriori error estimator for a variational inequality governed by the Stokes equations. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300121.

Jia:2020:AAK

[Jia20] Zhongxiao Jia. Approximation accuracy of the Krylov subspaces for linear discrete ill-posed problems. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300777.

Joseph:2024:NNB

[JJ24] Sobin Joseph and Shashi Jain. A neural network based model for multi-dimensional nonlinear Hawkes processes. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001389.

Jagels:2021:EGL

[JJR21] Carl Jagels, Khalide Jbilou, and Lothar Reichel. The extended global Lanczos method, Gauss–Radau quadrature, and matrix function approximation. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720303186.

Jeon:2021:PVA

[JK21a] Junkee Jeon and Minsuk Kwak. Pricing variable annuity with surrender guarantee. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001278.

Jing:2021:CRF

[JK21b] Ke Jing and Ning Kang.

Convergence rates of a family of barycentric rational Hermite interpolants and their derivatives. Journal of Computational and Applied Mathematics, 395 (??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721001898.

Jean:2023:MAM

[JKJ^{+}23] Koua Kadio Ettien Jean, Jirí Kuthan, Martin Jurík, Martin Vítek, and Frantisek Mach. Magnetic actuation of miniature robots on planar surfaces: Mathematical modeling and estimation of parameters. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300167X.

Jang:2020:DDP

[JKKK20] Lee-Chae Jang, Wonjoo Kim, Hyuck-In Kwon, and Taekyun Kim. On degenerate daehee polynomials and numbers of the third kind. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303401.

John:2023:POP

[JKW23] Volker John, Petr Knobloch, and Ulrich Wilbrandt. A posteriori optimization of parameters in stabilized methods for convection– diffusion problems — part II. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001115.

Jiang:2021:RSM

[JL21] Yingchun Jiang and Wan Li. Random sampling in multiply generated shiftinvariant subspaces of mixed Lebesgue spaces L^{p,q}(R × R^{d}). Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305288.

Jiang:2024:CES

[JL24a] Yu Jiang and Guangqiang Lan. Convergence and exponential stability of modified truncated Milstein method for stochastic differential equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002327.

Jin:2024:MFE

[JL24b] Xinran Jin and Jeonghun J. Lee. Mixed finite element methods for nonlinear reaction–diffusion equations with interfaces. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000050.

Jin:2021:ICB

[JLL21] Xin Jin, Qian Liu, and Huizhen Long. Impact of cost-benefit analysis on financial benefit evaluation of investment projects under back propagation neural network. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304635.

Jiang:2023:URR

[JLLL23] Fushuai Jiang, Chen Liang, Yutong Liang, and Garving K. Luli. Univariate range-restricted C^{2} interpolation algorithms. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006380.

Jiao:2024:SOH

[JLM24] Jinkai Jiao, Zhenhai Liu, and Stanislaw Migórski. Second order hemivariational inequality driven by evolution differential inclusion to a dynamic thermoviscoelastic contact problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003108.

Jiang:2023:DPT

[JLW23] Yi Jiang, Jun Liu, and Xiang-Sheng Wang. A direct parallel-in-time quasiboundary value method for inverse space-dependent source problems. Journal of Computational and Applied Mathematics, 423 (??):??, May 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005568.

Jian:2021:QBS

[JLY^{+}21] Jinbao Jian, Pengjie Liu, Jianghua Yin, Chen Zhang, and Miantao Chao. A QCQP-based splitting SQP algorithm for two-block nonconvex constrained optimization problems with application. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306592.

Jiang:2022:HSG

[JLZ22] Haiyan Jiang, Tiao Lu, and Weitong Zhang. A hybrid sinc-Galerkin/finite-difference method for the time-dependent Wigner equation. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000401.

Jia:2023:IRM

[JLZ23] Zhigang Jia, Xuan Liu, and Meixiang Zhao. The implicitly restarted multisymplectic block-Lanczos method for large-scale Hermitian quaternion matrix eigenvalue problem and applications. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003363.

Johnston:2021:LSS

[JM21] Hunter Johnston and Daniele Mortari. Least-squares solutions of boundary-value problems in hybrid systems. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001436.

Jerez:2021:ACA

[JMA21] Andrés Jerez, Miguel Márquez, and Henry Arguello. Adaptive coded aperture design for compressive computed tomography. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304659.

Jia:2024:DID

[JMHZ24] Fan Jia, Shen Mao, Zijian Huang, and Tieyong Zeng. Detachable image decomposition and illumination mapping search for low-light image enhancement. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003795.

Jian:2023:CAI

[JMLX23] Jinbao Jian, Guodong Ma, Pengjie Liu, and Jiawei Xu. Convergence analysis of an improved Bregmantype Peaceman–Rachford splitting algorithm for nonconvex nonseparable linearly constrained optimization problems. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000304.

Jia:2024:NPP

[JMX24] Caijuan Jia, Yan Meng, and Jiaxin Xiao. Nonhomogeneous periodic patterns in a predator-prey model with time delay and predator-taxis. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003686.

Jing:2023:AAM

[JNCY23] Jiayin Jing, Guyan Ni, Ciwen Chen, and Bo Yang. Alternate algorithms to most referenced techniques of numerical optimization to solve the symmetric rank-R approximation problem of symmetric tensors. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004034.

Jeon:2024:DAA

[JO24a] Junkee Jeon and Jehan Oh.

Dynamic asset allocation and consumption ratcheting with costs. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002164.

Jeong:2024:ESI

[JO24b] Jaewoo Jeong and Hae-Soo Oh. Extraction of stress intensity factors of biharmonic equations with free boundary conditions on crack faces (III): SIF of a thin plate subjected to the bending moment. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003364.

Jandrlic:2024:EBG

[JPS24] D. R. Jandrlić, A. V. Pejcev, and M. M. Spalević. Error bound of Gaussian quadrature rules for certain Gegenbauer weight functions. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006052.

Jiang:2022:NAS

[JQ22] Nan Jiang and Changxin Qiu. Numerical analysis of a second order ensemble algorithm for numerical approximation of stochastic Stokes–Darcy equations. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005525.

Jeng:2021:CPI

[JS21] B.-W. Jeng and Sirilak Sriburadet. Continuation and preconditioned imaginary time evolution methods for boson-fermion mixtures. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303101.

Jang:2022:FEA

[JS22a] Yongseok Jang and Simon Shaw. Finite element approximation and analysis of a viscoelastic scalar wave equation with internal variable formulations. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001492.

Jose:2022:CED

[JS22b] Jitto Jose and E. I. Abdul Sathar. Characterization of exponential distribution using extropy based on lower k-records and its application in testing exponentiality. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004386.

Junior:2023:PEE

[JSRdS23] D. S. Almeida Júnior, M. L. Santos, J. E. Muñoz Rivera, and M. J. dos Santos. Porous-elasticity equations with indefinite damping. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004885.

Jeon:2020:NAI

[JT20] Youngmok Jeon and Mai Lan Tran. Numerical analysis of interface hybrid difference methods for elliptic interface equations. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301606.

Juhasz:2021:NTB

[Juh21] Imre Juhász. A NURBS transition between a Bézier curve and its control polygon. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100248X.

Jukic:2020:NSC

[Juk20] Dragan Jukić. A necessary and sufficient criterion for the existence of the global minima of a continuous lower bounded function on a noncompact set. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300820.

Jothilakshmi:2022:CAM

[JVAD22] G. Jothilakshmi, B. S.

Vadivoo, Y. Almalki, and A. Debbouche. Controllability analysis of multiple fractional order integrodifferential damping systems with impulsive interpretation. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000747.

Jiang:2023:ISA

[JW23] Fan Jiang and Zhongming Wu. An inexact symmetric ADMM algorithm with indefinite proximal term for sparse signal recovery and image restoration problems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003144.

Ji:2024:CMB

[JW24] Jun Ji and Long Wang. Construction of Mercedes-Benz frames via QR factorization. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003734.

Jiang:2021:EHO

[JWG21] Chaolong Jiang, Yushun Wang, and Yuezheng Gong. Explicit high-order energy-preserving methods for general Hamiltonian partial differential equations. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305896.

Jiao:2023:OSB

[JWS23] Hongwei Jiao, Wenjie Wang, and Youlin Shang. Outer space branch-reduction-bound algorithm for solving generalized affine multiplicative problems. Journal of Computational and Applied Mathematics, 419 (??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722003995.

Jin:2022:SGA

[JWSG22] Yu Jin, Chun Wen, Zhao-Li Shen, and Xian-Ming Gu. A simpler GM-RES algorithm accelerated by Chebyshev polynomials for computing PageRank. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001819.

Jia:2021:FCA

[JWZ21] Jinhong Jia, Hong Wang, and Xiangcheng Zheng. A fast collocation approximation to a twosided variable-order spacefractional diffusion equation and its analysis. Journal of Computational and Applied Mathematics, 388 (??):??, May 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305252.

Ji:2021:CHQ

[JYWZ21] Ye Ji, Ying-Ying Yu, Meng-Yun Wang, and Chun-Gang Zhu. Constructing high-quality planar NURBS parameterization for isogeometric analysis by adjustment control points and weights. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002375.

Jeong:2021:NUC

[JYY21] Byeongseon Jeong, Hyoseon Yang, and Jungho Yoon. A non-uniform cornercutting subdivision scheme with an improved accuracy. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000650.

Jiang:2023:TTC

[JYYL23] Xianzhen Jiang, Huihui Yang, Jianghua Yin, and Wei Liao. A three-term conjugate gradient algorithm with restart procedure to solve image restoration problems. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006185.

Jones:2021:CNI

[JZ21] Derrick Jones and Xu Zhang.

A class of nonconforming immersed finite element methods for Stokes interface problems. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001126.

Ji:2022:WGF

[JZ22] Guanghua Ji and Wanwan Zhu. A weak Galerkin finite element method for time-dependent Poisson– Nernst–Planck equations. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002771.

Ju:2024:CFP

[JZ24] Yingying Ju and Chengbo Zhai. Common fixed point theorems for several pairs of generalized enriched contractions in a complete generalized convex metric space. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003145.

Jiang:2023:NZM

[JZM^{+}23] Chao Jiang, Yixiang Zhang, Chao Mou, Bin Li, Xiaobing Sun, and Yang Shi. A new ZNN model for finding discrete time-variant matrix square root: From model design to parameter analysis. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002042.

Jiang:2022:LES

[JZW22] Maosheng Jiang, Jia Zhao, and Qi Wang. Linear energy stable numerical schemes for a general chemo-repulsive model. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002060.

Jiang:2022:RBK

[JZY22] Xiang-Long Jiang, Ke Zhang, and Jun-Feng Yin. Randomized block Kaczmarz methods with k-means clustering for solving large linear systems. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004507.

Jiang:2022:HOA

[JZZH22] Yan-Qun Jiang, Shu-Guang Zhou, Xu Zhang, and Ying-Gang Hu. High order all-speed semi-implicit weighted compact nonlinear scheme for the isentropic Navier–Stokes equations. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001121.

Krupa:2021:DRT

[KA21] Artur Krupa and Izabella Antoniuk. Digital representation of tissues in high compute bioelectromagnetics. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100265X.

Kalachikova:2024:AWE

[KA24] Uygulaana Kalachikova and Dmitry Ammosov. Advancing wave equation analysis in dual-continuum systems: a partial learning approach with discrete empirical interpolation and deep neural networks. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000049.

Kudu:2022:SOA

[KAA22] Mustafa Kudu, Ilhame Amirali, and Gabil M. Amiraliyev. A second order accurate method for a parameterized singularly perturbed problem with integral boundary condition. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005173.

Kadak:2022:MNN

[Kad22a] Ugur Kadak. Multivariate neural network interpolation operators. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002011.

Korner:2022:WBS

[KAD22b] Jannis Körner, Anton Arnold, and Kirian Döpfner. WKB-based scheme with adaptive step size control for the Schrödinger equation in the highly oscillatory regime. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005288.

Kumam:2023:HHL

[KAM^{+}23] Poom Kumam, Auwal Bala Abubakar, Maulana Malik, Abdulkarim Hassan Ibrahim, Nuttapol Pakkaranang, and Bancha Panyanak. A hybrid HS-LS conjugate gradient algorithm for unconstrained optimization with applications in motion control and image recovery. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002480.

Khalaf:2021:SVI

[KAMW21] Anas Dheyab Khalaf, Mahmoud Abouagwa, Almushaira Mustafa, and Xiangjun Wang. Stochastic Volterra integral equations with jumps and the strong superconvergence of the Euler–Maruyama approximation. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303629.

Kang:2020:AVP

[Kan20] Myeongmin Kang. Approximate versions of proximal iteratively reweighted algorithms including an extended IP–ICMM for signal and image processing problems. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030128X.

Kan:2024:AWT

[Kan24] Cihangir Kan. Analyzing wind turbine and wind farm outputs under dependence between wind speed and wind turbine availability. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003273.

Kapyrin:2021:ADD

[Kap21] I. V. Kapyrin. Assessment of density driven convection effect on the dynamics of contaminant propagation on a deep well radioactive waste injection disposal site. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000443.

Karatson:2020:SGP

[Kar20] J. Karátson. Sobolev gradient preconditioning for elliptic reaction–diffusion problems with some nonsmooth nonlinearities. Journal of Computational and Applied Mathematics, 363 (??):223–233, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042719302808.

Karamzin:2024:RCH

[Kar24] D. Yu. Karamzin. On regularity conditions in higherorder state-constrained control problems. Journal of Computational and Applied Mathematics, 452 (??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003534.

Karakoc:2023:NPA

[KAS23] Seydi Battal Gazi Karakoc, Khalid K. Ali, and Derya Yildirim Sucu. A new perspective for analytical and numerical soliton solutions of the Kaup–Kupershmidt and Ito equations. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004484.

Kazemi:2021:AST

[Kaz21] Manochehr Kazemi. Approximating the solution of three-dimensional nonlinear Fredholm integral equations. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002107.

Kazmi:2023:SON

[Kaz23] Kamran Kazmi. A second order numerical method for the time-fractional Black– Scholes European option pricing model. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003284.

Kuczynski:2022:MEC

[KBK^{+}22] M. D. Kuczyński, M. Borchardt, R. Kleiber, A. Könies, and C. Nührenberg. Magnetohydrodynamic eigenfunction classification with a neural network. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005124.

Katz:2023:IVM

[KCJ23] Sarah Katz, Alfonso Caiazzo, and Volker John. Impact of viscosity modeling on the simulation of aortic blood flow. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006343.

Kolev:2024:BSF

[KCK24] Vasil Kolev, Todor Cooklev, and Fritz Keinert. Bauer’s spectral factorization method for low order multiwavelet filter design. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006568.

Kasinathan:2024:ESN

[KCKK24] Dhanalakshmi Kasinathan, Dimplekumar Chalishajar, Ramkumar Kasinathan, and Ravikumar Kasinathan. Exponential stability of noninstantaneous impulsive second-order fractional neutral stochastic differential equations with statedependent delay. Journal of Computational and Applied Mathematics, 451 (??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724002620.

Kuang:2020:MPA

[KCL20] Shenfen Kuang, Hongyang Chao, and Qia Li. Majorized Proximal Alternating Imputation for regularized rank constrained matrix completion. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306843.

Kumar:2022:SFD

[KCRVA22] Kamalesh Kumar, P. Pramod Chakravarthy, Higinio Ramos, and Jesús Vigo-Aguiar. A stable finite difference scheme and error estimates for parabolic singularly perturbed PDEs with shift parameters. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303411.

Kostic:2021:PDB

[KCS21] V. R. Kostić, Lj. Cvetković, and E. Sanca. From pseudospectra of diagonal blocks to pseudospectrum of a full matrix. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305562.

Kounadis:2020:GFE

[KD20a] G. Kounadis and V. A.

Dougalis. Galerkin finite element methods for the Shallow Water equations over variable bottom. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303127.

Kravetc:2020:FEA

[KD20b] Tatiana Kravetc and Rune Dalmo. Finite element application of ERBS extraction. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302387.

Kumar:2022:DSF

[KD22] Naresh Kumar and Bhupen Deka. Developing stabilizer free weak Galerkin finite element method for second-order wave equation. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002175.

Kumar:2023:HOW

[KD23] Raman Kumar and Bhupen Deka. High-order weak Galerkin scheme for H(div)-elliptic interface problems. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002133.

Kum:2022:GBA

[KDLY22] S. Kum, M. H. Duong, Y. Lim, and S. Yun. A GPM-based algorithm for solving regularized Wasserstein barycenter problems in some spaces of probability measures. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200293X.

Katrutsa:2020:BBL

[KDO20] Alexandr Katrutsa, Talgat Daulbaev, and Ivan Oseledets. Black-box learning of multigrid parameters. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305291.

Kechriniotis:2024:CMH

[KDOT24] Aristides I. Kechriniotis, Konstantinos K. Delibasis, Iro Oikonomou, and Georgios N. Tsigaridas. Classical multivariate Hermite coordinate interpolation on n-dimensional grids. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002127.

Kayijuka:2021:CCA

[KEKT21] I. Kayijuka, S. M. Ege, A. Konuralp, and F. S. Topal. Clenshaw–Curtis algorithms for an efficient numerical approximation of singular and highly oscillatory Fourier transform integrals. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304921.

Keslerova:2023:NMG

[Kes23]  Radka Keslerová.  Nu-

merical modeling of generalized Newtonian fluids flow in S-type geometry of bypass. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001814.

Kaya:2024:LRS

[KF24] Adem Kaya and Melina Freitag. Low-rank solutions to the stochastic Helmholtz equation. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001754.

Kovalnogov:2023:RKN

[KFK^{+}23] V. N. Kovalnogov, R. V. Fedorov, M. T. Karpukhina, M. I. Kornilova, T. E. Simos, and Ch. Tsitouras. Runge–Kutta– Nyström methods of eighth order for addressing linear inhomogeneous problems. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200396X.

Kovalnogov:2024:EHR

[KFK^{+}24] V. N. Kovalnogov, R. V.

Fedorov, M. T. Karpukhina, M. I. Kornilova, T. E. Simos, and Ch. Tsitouras. Economical handling of Runge–Kutta– Nyström step rejection. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723004727.

Kaur:2023:MFW

[KGV23] Navneet Kaur, Bivek Gupta, and Amit K. Verma. Multidimensional fractional wavelet transforms and uncertainty principles. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001942.

Kazem:2021:SDI

[KH21] Saeed Kazem and A. Hatam.

Scattered data interpolation: Strictly positive definite radial basis/cardinal functions. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002004.

Ketelbuters:2022:TCE

[KH22] John-John Ketelbuters and Donatien Hainaut. Timeconsistent evaluation of credit risk with contagion. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004714.

Ketelbuters:2024:RMF

[KH24] John-John Ketelbuters and Donatien Hainaut. A recursive method for fractional Hawkes intensities and the potential approach of credit risk. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001456.

Klinge:2021:COS

[KHAW21] M. Klinge, D. Hernández-Abreu, and R. Weiner. A comparison of one-step and two-step W -methods and peer methods with approximate matrix factorization. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305242.

Khludnev:2021:SIE

[Khl21] Alexander Khludnev. Tshape inclusion in elastic body with a damage parameter. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100159X.

Keller:2024:E

[KHYL24] André A. Keller, Nguyen Van Hung, David W. K. Yeung, and Luca Lambertini. Editorial. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000244.

Kim:2020:ACR

[Kim20] Seungil Kim. Application of a complete radiation boundary condition for the Helmholtz equation in locally perturbed waveguides. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304613.

Kim:2022:HAB

[Kim22] Seungil Kim. Hybrid absorbing boundary conditions of PML and CRBC. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721003356.

Kar:2020:QOZ

[KJGN20] P. P. Kar, K. Jordaan, P. Gochhayat, and M. Kenfack Nangho. Quasiorthogonality and zeros of some _{2}φ_{2} and _{3}φ_{2} polynomials. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302016.

Krejic:2023:IRN

[KJO23] Natasa Krejić, Natasa Krklec Jerinkić, and Tijana Ostojić. An inexact restorationnonsmooth algorithm with variable accuracy for stochastic nonsmooth convex optimization problems in machine learning and stochastic linear complementarity problems. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005416.

Kim:2020:SFP

[KK20a] Dae San Kim and Taekyun Kim. On sums of finite products of balancing polynomials. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302041.

Kramarenko:2020:NAE

[KK20b] V. K. Kramarenko and Yu. A. Kuznetsov. New asymptotic expansion for the diffusion problem with high-contrast inclusions: Numerical results. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303668.

Kumar:2020:ETM

[KK20c] Chaman Kumar and Tejinder Kumar. On explicit tamed Milsteintype scheme for stochastic differential equation with Markovian switching. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302089.

Kumar:2021:NEM

[KK21] Chaman Kumar and Tejinder Kumar. A note on explicit Milstein-type scheme for stochastic differential equation with Markovian switching. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002144.

Kang:2022:FTS

[KK22] Dongseung Kang and Hoewoon B. Kim. Fourier transforms and L^{2}-stability of diffusion equations. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722000590.

Kaushik:2023:NOD

[KK23] Sonali Kaushik and Rajesh Kumar. A novel optimized decomposition method for Smoluchowski’s aggregation equation. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003594.

Khan:2024:DATb

[KK24a] Abdul Qadeer Khan and Syed Saqlain Kazmi. Dynamical analysis of a threespecies discrete biological system with scavenger. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005885.

Korotov:2024:DFF

[KK24b] Sergey Korotov and Michal Krízek. Divergence-free finite element spaces for stress tensors. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004818.

Kayvanloo:2020:FMN

[KKA20] Hojjatollah Amiri Kayvanloo, Mahnaz Khanehgir, and Reza Allahyari. A family of measures of noncompactness in the Hölder space C^{n,γ}(R_{+}) and its application to some fractional differential equations and numerical methods. Journal of Computational and Applied Mathematics, 363(??):256–272, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719302857.

Kim:2020:SJS

[KKJ20] Hyunju Kim, Keon Ho Kim, and Bongsoo Jang. Shifted Jacobi spectral-Galerkin method for solving fractional order initial value problems. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030279X.

Kim:2022:OPU

[KKJ22] Kyong-Hui Kim, Su-Hyang Kim, and Ho-Bom Jo. Option pricing under mixed hedging strategy in timechanged mixed fractional Brownian model. Journal of Computational and Applied Mathematics, 416 (??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722002400.

Kim:2022:CPO

[KKKL22] Jerim Kim, Bara Kim, Jeongsim Kim, and Sungji Lee. Computation of powered option prices under a general model for underlying asset dynamics. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005902.

Katona:2023:ARD

[KKO23] Mihály Katona, Miklós Kuczmann, and Tamás Orosz. Accuracy of the robust design analysis for the flux barrier modelling of an interior permanent magnet synchronous motor. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001723.

Kumar:2023:CDS

[KL23a] Vivek Kumar and Günter Leugering. Convection dominated singularly perturbed problems on a metric graph. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000067.

Kuo:2023:ISM

[KL23b] Yueh-Cheng Kuo and Ching-Sung Liu. An index search method based inner-outer iterative algorithm for solving nonnegative least squares problems. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005520.

Karamollahi:2021:CMF

[KLH21] Nasibeh Karamollahi, Ghasem Barid Loghmani, and Mohammad Heydari. A computational method to find dual solutions of the onedimensional Bratu problem. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306002.

Kulikov:2021:NSN

[KLK21] G. Yu. Kulikov, P. M.

Lima, and M. V. Kulikova. Numerical solution of the neural field equation in the presence of random disturbance. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305680.

Kadiri:2021:HDO

[KLM21] Mostafa Kadiri, Mohammed Louaked, and Houari Mechkour. Hydrodynamic design optimization using non stationary porous media model. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304842.

Kelly:2022:RAN

[KLM22] Cónall Kelly, Gabriel Lord, and Heru Maulana. The role of adaptivity in a numerical method for the Cox–Ingersoll–Ross model. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000760.

Krait:2021:CNA

[KLMP21] George Krait, Sylvain Lazard, Guillaume Moroz, and Marc Pouget. Certified numerical algorithm for isolating the singularities of the plane projection of generic smooth space curves. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001734.

Knobloch:2021:IPO

[KLS21] Petr Knobloch, Petr Lukás, and Pavel Solin. Importance of parameter optimization in a nonlinear stabilized method adding a crosswind diffusion. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001461.

Kim:2024:IUS

[KLZ^{+}24] Jichul Kim, Byungjoon Lee, Michael R. Zanetti, Kyle A. Miller, and Seongjai Kim. Iterative Update Sample Consensus (IUSAC): a repeatable algorithm for optimal consensus set. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003679.

Katani:2020:HLB

[KM20] R. Katani and S. Mckee.

A hybrid Legendre blockpulse method for mixed Volterra–Fredholm integral equations. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301588.

Katani:2022:NST

[KM22] R. Katani and S. Mc-Kee. Numerical solution of two-dimensional weakly singular Volterra integral equations with non-smooth solutions. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004015.

Kumar:2021:TCN

[KMD21] Vipin Kumar, Muslim Malik, and Amar Debbouche. Total controllability of neutral fractional differential equation with non-instantaneous impulsive effects. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304490.

Kimura:2020:CEE

[KMN20] Takuma Kimura, Teruya Minamoto, and Mitsuhiro T. Nakao. Constructive error estimates for full discrete approximation of periodic solution for heat equation. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305151.

Koukouvinos:2020:EER

[KMT20] Christos Koukouvinos, Marilena Mitrouli, and Ondrej Turek. Efficient estimates in regression models with highly correlated covariates. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304194.

Kuang:2021:TNE

[KMTZ21] Zhonghong Kuang, Vladimir V.

Mazalov, Xindi Tang, and Jie Zheng. Transportation network with externalities. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303824.

Khan:2024:DATa

[KMUR24] Abdul Qadeer Khan, Atifa Maqbool, Md. Jasim Uddin, and Sarker Md. Sohel Rana. Dynamical analysis of a two-dimensional discrete predator-prey model. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042723005228.

Ke:2021:MLA

[KMZ21] Yifen Ke, Changfeng Ma, and Huai Zhang. A modified LM algorithm for tensor complementarity problems over the circular cone. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003113.

Kant:2020:AMS

[KN20] Kapil Kant and Gnaneshwar Nelakanti. Approximation methods for second kind weakly singular Volterra integral equations. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305369.

Khenfar:2024:FEA

[KNM24] Sokina Khenfar, Serge Nicaise, and Ismail Merabet. On the finite element approximation of the obstacle problem of a Naghdi shell. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006143.

Kumkov:2020:IPE

[KNOR20] Sergey I. Kumkov, Vyacheslav S. Nikitin, Tatyana N. Ostanina, and Valentin M. Rudoy. Interval processing of electrochemical data. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302521.

Kosir:2022:SCS

[KO22] Tomaz Kosir and Matjaz Omladic. Singular components of shock model copulas. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100371X.

Koga:2021:SPA

[Kog21] Kazuki Koga. Signal processing approach to mesh refinement in simulations of axisymmetric droplet dynamics. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304222.

Kong:2022:MCR

[Kon22] Yong Kong. Multiple consecutive runs of multistate trials: Distributions of (k_{1}, k_{2}, …, k_{l}) patterns. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004696.

Kopteva:2020:HAF

[Kop20] Natalia Kopteva. How accurate are finite elements on anisotropic triangulations in the maximum norm? Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303139.

Kovtunenko:2022:PMN

[Kov22] Victor A. Kovtunenko.

Poroelastic medium with non-penetrating crack driven by hydraulic fracture: Variational inequality and its semidiscretization. Journal of Computational and Applied Mathematics, 405 (??):??, May 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005628.

Khludnev:2020:EPE

[KP20] Alexander Khludnev and Tatyana Popova. Equilibrium problem for elastic body with delaminated Tshape inclusion. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301618.

Khandelwal:2022:PPE

[KP22] Rohit Khandelwal and Kamana Porwal. Pointwise a posteriori error analysis of quadratic finite element method for the elliptic obstacle problem. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001625.

Kotlan:2023:UCC

[KPD23] Vaclav Kotlan, Iveta Petrasova, and Ivo Dolezel. Usage of continuous cooling transformation (CCT) diagrams for laser and small-scale induction hardening. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000675.

Kumar:2021:NSC

[KPS21] Kamlesh Kumar, Rajesh K. Pandey, and Farheen Sultana. Numerical schemes with convergence for generalized fractional integro-differential equations. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306099.

Khoo:2024:DNN

[KPS24a] Yuehaw Khoo, Sounak Paul, and Nir Sharon. Deep neural-network prior for orbit recovery from method of moments. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000311.

Knez:2024:CSI

[KPS24b] Marjeta Knez, Francesca Pelosi, and Maria Lucia Sampoli. Construction of G^{2} spatial interpolants with prescribed arc lengths. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006271.

Kim:2021:TSA

[KPW21] Sangjin Kim, Michael Pokojovy, and Xiang Wan. The taut string approach to statistical inverse problems: Theory and applications. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303897.

Kounchev:2021:EEI

[KR21] O. Kounchev and H. Render. Error estimates for interpolation with piecewise exponential splines of order two and four. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000832.

Kravetc:2022:IAU

[Kra22] Tatiana Kravetc. Isogeometric analysis using a tensor product blending spline construction. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002072.

Kriel:2021:RDP

[Kri21]  A. J. Kriel.  On the

range diminishing property of numerical schemes for scalar conservation laws. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303046.

Kounchev:2023:FAI

[KRT23] Ognyan Kounchev, Hermann Render, and Tsvetomir Tsachev. Fast algorithms for interpolation with L-splines for differential operators L of order 4 with constant coefficients. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004964.

Kuzenov:2024:DMS

[KRV24] Victor V. Kuzenov, Sergei V.

Ryzhkov, and Aleksey Yu Varaksin. Development of a method for solving elliptic differential equations based on a nonlinear compact-polynomial scheme. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003479.

Kurkcu:2021:DCN

[KS21] Ömür Kivanç Kürkçü and Mehmet Sezer. A directly convergent numerical method based on orthoexponential polynomials for solving integrodifferential-delay equations with variable coefficients and infinite boundary on half-line. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305410.

Klippenstein:2022:FAC

[KS22a] Brock Klippenstein and Richard Mikaël Slevinsky. Fast associated classical orthogonal polynomial transforms. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004532.

Kolbe:2022:ARM

[KS22b] Niklas Kolbe and Nikolaos Sfakianakis. An adaptive rectangular mesh administration and refinement technique with application in cancer invasion models. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002096.

Kusch:2022:OMH

[KS22c] Jonas Kusch and Louisa Schlachter. Oscillation mitigation of hyperbolicitypreserving intrusive uncertainty quantification methods for systems of conservation laws. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003368.

Kassabek:2023:TPI

[KS23a] Samat A. Kassabek and Durvudkhan Suragan. Twophase inverse Stefan problems solved by heat polynomials method. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004526.

Kheirandish:2023:GBI

[KS23b] Ehsan Kheirandish and Abbas Salemi. Generalized bilateral inverses. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300081X.

Korman:2024:CGS

[KS24] Philip Korman and Dieter S. Schmidt. Continuation of global solution curves using global parameters. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005654.

Krotz:2022:VRP

[KSG^{+}22] Johannes Krotz, Matthew R. Sweeney, Carl W. Gable, Jeffrey D. Hyman, and Juan M. Restrepo. Variable resolution Poissondisk sampling for meshing discrete fracture networks.

Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000073.

Kan:2022:ECS

[KSPC22] Ziyun Kan, Ningning Song, Haijun Peng, and Biaosong Chen. Extension of complex step finite difference method to Jacobian-free Newton– Krylov method. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100354X.

Kumar:2022:PUG

[KSVA22] Sunil Kumar, Sumit, and Jesus Vigo-Aguiar. A parameter-uniform grid equidistribution method for singularly perturbed degenerate parabolic convection– diffusion problems. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305641.

Kramer:2021:SMN

[KSZ21] Patrick Krämer, Katharina Schratz, and Xiaofei Zhao. Splitting methods for nonlinear Dirac equations with Thirring type interaction in the nonrelativistic limit regime. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304972.

Kacala:2021:STS

[KT21] Viliam Kacala and Csaba Török. Speedup of tridiagonal system solvers. Journal of Computational and Applied Mathematics, 381 (??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042720302880.

Koutras:2024:CDF

[KT24] Markos V. Koutras and Ioannis S. Triantafyllou. Consecutive Dual Failure Mode shock systems. Journal of Computational and Applied Mathematics, 435 (??):??, ???? 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002625.

Kulikova:2021:SBS

[KTK21] M. V. Kulikova, J. V.

Tsyganova, and G. Yu. Kulikov. SVD-based state and parameter estimation approach for generalized Kalman filtering with application to GARCH-in-mean estimation. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930490X.

Ku:2024:LPE

[Ku24]  JaEun Ku.  Localized

pointwise error estimates for hybrid finite element methods. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003291.

Khan:2020:SPC

[KUA20] N. U. Khan, T. Usman, and M. Aman. Some properties concerning the analysis of generalized Wright function. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030131X.

Kumar:2022:CSP

[Kum22] D. Ramesh Kumar. Common solution to a pair of nonlinear Fredholm and Volterra integral equations and nonlinear fractional differential equations. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005306.

Kuznetsov:2022:CBF

[Kuz22] Alexey Kuznetsov. Computing the Barnes Gfunction and the gamma function in the entire complex plane. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200111X.

Koleva:2019:NAO

[KV19] Miglena N. Koleva and Lubin G. Vulkov. Numerical analysis of one dimensional motion of magma without mass forces. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303358.

Kozerenko:2020:PXP

[KV20] Sergiy Kozerenko and Jose C. Valverde. Periods in XOR parallel dynamical systems over directed dependency graphs. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303103.

Kalachikova:2023:GMF

[KVHC23] Uygulaana Kalachikova, Maria Vasilyeva, Isaac Harris, and Eric T. Chung. Generalized Multiscale Finite Element Method for scattering problem in heterogeneous media. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005751.

Kruse:2023:BMM

[KW23] Raphael Kruse and Rico Weiske. The BDF2– Maruyama method for the stochastic Allen–Cahn equation with multiplicative noise. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200317X.

Karnik:2024:NNA

[KWI24] Santhosh Karnik, Rongrong Wang, and Mark Iwen. Neural network approximation of continuous functions in high dimensions with applications to inverse problems. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005010.

Kinoshita:2020:SLB

[KWN20] Takehiko Kinoshita, Yoshitaka Watanabe, and Mitsuhiro T. Nakao. Some lower bound estimates for resolvents of a compact operator on an infinitedimensional Hilbert space. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719305667.

Kinoshita:2023:IMO

[KWYN23] Takehiko Kinoshita, Yoshitaka Watanabe, Nobito Yamamoto, and Mitsuhiro T. Nakao. Inclusion method of optimal constant with quadratic convergence for H_{0}^{1}-projection error estimates and its applications. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002564.

Kang:2023:ECO

[KWZX23] Hongchao Kang, Ruoxia Wang, Meijuan Zhang, and Chunzhi Xiang. Efficient computation of oscillatory Bessel transforms with a singularity of Cauchy type. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001644.

Kong:2023:ENM

[KXW23] Desong Kong, Shuhuang Xiang, and Hongyu Wu. An efficient numerical method for Volterra integral equation of the second kind with a weakly singular kernel. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000456.

Kim:2021:PEB

[KYP21] Donghyun Kim, Ji-Hun Yoon, and Chang-Rae Park. Pricing external barrier options under a stochastic volatility model. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001758.

Kang:2023:NEA

[KZ23a] Hongchao Kang and Meijuan Zhang. Numerical evaluation and analysis of highly oscillatory singular Bessel transforms with a particular oscillator. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004332.

See erratum [KZ24].

Kittaneh:2023:BNR

[KZ23b] Fuad Kittaneh and Ali Zamani. Bounds for Anumerical radius based on an extension of A-Buzano inequality. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000146.

Kang:2024:ENE

[KZ24] Hongchao Kang and Meijuan Zhang. Erratum to “Numerical evaluation and analysis of highly oscillatory singular Bessel transforms with a particular oscillator” [J. Comput. Appl. Math. 420 (2023) 114835]. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003254.

See [KZ23a].

Kravcenko:2021:DSH

[KZM21] Michal Kravcenko, Jan Zapletal, and Michal Merta. Distributed solution of the Helmholtz transmission problems via relaxed local multi-trace formulation and adaptive cross approximation. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002661.

Kang:2022:NMC

[KZW22] Hongchao Kang, Meijuan Zhang, and Ruoxia Wang. Numerical methods for Cauchy principal value integrals of oscillatory Bessel functions. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000802.

LaCruz:2022:EPR

[La 22] William La Cruz. An extended projected residual algorithm for solving smooth convex optimization problems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001558.

Laib:2023:NST

[LBBB23] Hafida Laib, Aissa Boulmerka, Azzeddine Bellour, and Fouzia Birem. Numerical solution of twodimensional linear and nonlinear Volterra integral equations using Taylor collocation method. Journal of Computational and Applied Mathematics, 417 (??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722002643.

Lin:2020:SPS

[LC20] Matthew M. Lin and Chun-Yueh Chiang. On the semigroup property for some structured iterations. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300595.

Liu:2023:ESI

[LC23] Zhengguang Liu and Chuanjun Chen. On efficient semi-implicit auxiliary variable methods for the six-order Swift– Hohenberg model. Journal of Computational and Applied Mathematics, 419 (??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722003697.

Liu:2024:LDG

[LC24a] Yanhua Liu and Yao Cheng. Local discontinuous Galerkin method for a singularly perturbed fourth-order problem of reaction–diffusion type. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300585X.

Lovat:2024:RCS

[LC24b] Giampiero Lovat and Salvatore Celozzi. Rapidly convergent series and closedform expressions for a class of integrals involving products of spherical Bessel functions. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000517.

Li:2020:NTV

[LCB20] Li Li, Chen Chen, and Bo Bi. A new total variational regularization method for nonlinear inverse problems in fluorescence molecular tomography. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930411X.

Li:2021:NSI

[LCD21] R. P. Li, X. B. Chen, and W. Y. Duan. Numerical study on integrals involving the product of Bessel functions and a trigonometric function arising in hydrodynamic problems. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304519.

Liu:2024:DIB

[LCGW24] Qinyao Liu, Feiyan Chen, Qian Guo, and Xuchen Wang. Distributed identification based partiallycoupled recursive generalized extended least squares algorithm for multivariate input-output-error systems with colored noises from observation data. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002267.

Lin:2023:PPE

[LCH23] Xiuxiu Lin, Yanping Chen, and Yunqing Huang. A priori and a posteriori error analysis of hp spectral element discretization for optimal control problems with elliptic equations. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005581.

Liu:2020:FGF

[LCW20] Huan Liu, Aijie Cheng, and Hong Wang. A fast Galerkin finite element method for a spacetime fractional Allen–Cahn equation. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304856.

Liu:2019:DUT

[LCZ19] YongGe Liu, Xu Chen, and WenYan Zhuo. Dividends under threshold dividend strategy with randomized observation periods and capital-exchange agreement. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304297.

Liu:2024:UEA

[LCZL24] Wanxiang Liu, Yanping Chen, Jianwei Zhou, and Qin Liang. Unconditional error analysis of linearized BDF2 mixed virtual element method for semilinear parabolic problems on polygonal meshes. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001134.

Liu:2021:CBE

[LD21] Linna Liu and Feiqi Deng. Complete backward Euler numerical scheme for general SFDEs with exponential stability under the polynomial growth condition. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305331.

Li:2022:GCG

[LD22] Hailong Li and Liang Ding. Generalized conditional gradient method for elastic-net regularization. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004957.

Liu:2020:SDN

[LDC20] Jiankun Liu, Shouqiang Du, and Yuanyuan Chen. A sufficient descent nonlinear conjugate gradient method for solving Mtensor equations. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307149.

Long:2023:EAC

[LDM23] Xiaonian Long, Qianqian Ding, and Shipeng Mao. Error analysis of a conservative finite element scheme for time-dependent inductionless MHD problem. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003685.

Liu:2023:SAS

[LDQF23] Linna Liu, Feiqi Deng, Boyang Qu, and Jianyin Fang. Stability analysis of split-step theta method for neutral stochastic delayed neural networks. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002631.

Li:2022:MRT

[LDY22] Nan Li, Jian Ding, and Jun-Feng Yin. Modified relaxation two-sweep modulus-based matrix splitting iteration method for solving a class of implicit complementarity problems. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001650.

Lee:2020:NCS

[Lee20]  Hyun Geun Lee.  A

new conservative Swift– Hohenberg equation and its mass conservative method. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301060.

Lee:2021:ENM

[Lee21] Jung-Kyung Lee. An efficient numerical method for pricing American put options under the CEV model. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306026.

Lee:2023:AFE

[Lee23] Jeonghun J. Lee. Analysis of finite element methods for dynamic poroelasticity: Low frequency waves. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003624.

Lemle:2024:LTA

[Lem24] Ludovic Dan Lemle. On the Lie–Trotter approximation formula. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004107.

Llamazares-Elias:2023:MRS

[LET23] S. Llamazares-Elias and A. Tocino. Mean-reverting schemes for solving the CIR model. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002984.

Letoufa:2024:ESF

[Let24]  Yassine Letoufa.  Esti-

mating the stress and fatigue damage of a thin film with non-linear viscoelastic using asymptotic analysis. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300660X.

Liu:2021:SRK

[LF21] Xin Liu and Jason Frank. Symplectic Runge–Kutta discretization of a regularized forward-backward sweep iteration for optimal control problems. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304246.

Liu:2024:LFE

[LF24] Kai Liu and Ting Fu. Linearly-fitted energy-mass-preserving schemes for Korteweg–de Vries equations. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400164X.

Li:2020:UID

[LFL20] Zhiyuan Li, Kenichi Fujishiro, and Gongsheng Li. Uniqueness in the inversion of distributed orders in ultraslow diffusion equations. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305692.

Li:2024:SNN

[LFZ24] Jianfei Li, Han Feng, and Ding-Xuan Zhou. Sign-ReLU neural network and its approximation ability. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004958.

Liu:2022:FFI

[LG22] Kai Liu and Guiding Gu.

A family of fully implicit strong Itô–Taylor numerical methods for stochastic differential equations. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100546X.

Li:2023:DBF

[LG23] Yulong Li and Victor Ginting. On the Dirichlet BVP of fractional diffusion advection reaction equation in bounded interval: Structure of solution, integral equation and approximation. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000419.

Lei:2023:SWC

[LGC23] Ziyi Lei, Siqing Gan, and Ziheng Chen. Strong and weak convergence rates of logarithmic transformed truncated EM methods for SDEs with positive solutions. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003867.

Lei:2022:FPH

[LGG^{+}22] Xiaojun Lei, Tongxiang Gu, Stef Graillat, Hao Jiang, and Jin Qi. A fast parallel high-precision summation algorithm based on AccSumK. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004490.

Liu:2020:MLT

[LGLC20] Jiangguo Liu, Jianli Gu, Huishu Li, and Kenneth H. Carlson. Machine learning and transport simulations for groundwater anomaly detection. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302739.

Liu:2020:RRD

[LGMC20] Chunyang Liu, Xingyu Gao, Wanli Ma, and Xiangtuo Chen. Research on regional differences and influencing factors of green technology innovation efficiency of China’s hightech industry. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306028.

See corrigendum [LLG^{+}22].

Liao:2024:NSS

[LGY24] Feng Liao, Fazhan Geng, and Lingxing Yao. Numerical solutions of Schrödinger– Boussinesq system by orthogonal spline collocation method. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002346.

Lotfi:2020:EDL

[LH20] Mina Lotfi and S. Mohammad Hosseini. An efficient Dai–Liao type conjugate gradient method by reformulating the CG parameter in the search direction equation. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719307137.

Lin:2021:NIE

[LH21] Sha Lin and Xin-Jiang He. A new integral equation approach for pricing American-style barrier options with rebates. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303988.

Long:2023:FSA

[LH23] Xian-Jun Long and Yue-Hong He. A fast stochastic approximation-based subgradient extragradient algorithm with variance reduction for solving stochastic variational inequality problems. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004009.

Lu:2024:MAE

[LH24] Zibo Lu and Jizu Huang.

A multiscale asymptotic expansion for combustion system with composite materials. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006192.

Luan:2024:EEM

[LHE24] Vu Thai Luan, Nguyen Van Hoang, and Julius O. Ehigie. Efficient exponential methods for genetic regulatory systems. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003680.

Lahrouz:2021:MLS

[LHES21] Aadil Lahrouz, Riane Hajjami, Mustapha El Jarroudi, and Adel Settati. Mittag-Leffler stability and bifurcation of a nonlinear fractional model with relapse. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305380.

Li:2024:DVP

[LHG^{+}24] Xiaowei Li, Qinghua Hu, Renshu Gu, Jinlan Xu, Haiyan Wu, and Gang Xu. DiagPara: Volumetric parameterization with energy-minimizing pair of diagonal surfaces from given boundaries. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001936.

Liu:2021:BBR

[LHH21] Hongyan Liu, Jin Huang, and Xiaoming He. Bivariate barycentric rational interpolation method for two dimensional fractional Volterra integral equations. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306300.

Li:2021:MSS

[LHHW21] Min Li, Chengming Huang, Peng Hu, and Jiao Wen. Mean-square stability and convergence of a split-step theta method for stochastic Volterra integral equations. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030368X.

Li:2023:MSS

[LHHW23] Min Li, Yaozhong Hu, Chengming Huang, and Xiong Wang. Mean square stability of stochastic theta method for stochastic differential equations driven by fractional Brownian motion. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004095.

Liu:2019:PPI

[LHL19] Chengzhi Liu, Xuli Han, and Juncheng Li. Preconditioned progressive iterative approximation for triangular Bézier patches and its application. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303929.

Long:2021:FTP

[LHL21] Haie Long, Bo Han, and Li Li. A fast two-point gradient method for solving non-smooth nonlinear illposed problems. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304052.

Li:2022:CBC

[LHL^{+}22] Ao Li, Jian Huang, Wei Liu, Huayi Wei, and Nianyu Yi. A characteristic block-centered finite difference method for Darcy–Forchheimer compressible miscible displacement problem. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001273.

Liu:2020:LOW

[LHMT20] Jiangguo Liu, Graham Harper, Nolisa Malluwawadu, and Simon Tavener. A lowest-order weak Galerkin finite element method for Stokes flow on polygonal meshes. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304820.

Liu:2024:NKS

[LHS^{+}24] Ju Liu, Ling-Wei Huang, Yuan-Hai Shao, Wei-Jie Chen, and Chun-Na Li. A nonlinear kernel SVM classifier via L_{0/1} soft-margin loss with classification performance. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004156.

Li:2022:TPM

[LHW22] Min Li, Chengming Huang, and Jiao Wen. A twoparameter Milstein method for stochastic Volterra integral equations. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004933.

Li:2020:SES

[Li20a] Pingrun Li. The solvability and explicit solutions of singular integral– differential equations of non-normal type via Riemann– Hilbert problem. Journal of Computational and Applied Mathematics, 374 (??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720300509.

Li:2020:STC

[Li20b] Pingrun Li. Solvability theory of convolution singular integral equations via Riemann–Hilbert approach. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306065.

Li:2023:HON

[Li23] Yunzhang Li. A highorder numerical scheme for stochastic optimal control problem. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001024.

Li:2024:WBE

[Li24a] Guanglian Li. Waveletbased edge multiscale parareal algorithm for subdiffusion equations with heterogeneous coefficients in a large time domain. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005526.

Li:2024:WGMa

[Li24b]  Xiaolin Li.  A weak

Galerkin meshless method for incompressible Navier– Stokes equations. Journal of Computational and Applied Mathematics, 445 (??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724000724.

Liu:2020:EEB

[Liu20]  Xuefeng Liu.  Explicit

eigenvalue bounds of differential operators defined by symmetric positive semidefinite bilinear forms. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306715.

Liu:2022:PEM

[Liu22] Tao Liu. Parameter estimation with the multigridhomotopy method for a nonlinear diffusion equation. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001807.

Liu:2023:SAF

[Liu23] Xing Liu. Strong approximation for fractional wave equation forced by fractional Brownian motion with Hurst parameter H ∈ (0, 12). Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002297.

Liu:2024:EAF

[Liu24] Xing Liu. Error analysis of a fully discrete method for time-fractional diffusion equations with a tempered fractional Gaussian noise. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002036.

Li:2023:APA

[LJ23a] Yipeng Li and Xiangmin Jiao. ARPIST: Provably accurate and stable numerical integration over spherical triangles. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004204.

Li:2023:PMO

[LJ23b] Zhen Li and Yao-Lin Jiang. Parallel model order reduction methods based on structured matrix analysis for discrete-time systems with parametric uncertainty. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000328.

Li:2020:PPE

[LJCL20] Jian Li, Feifei Jing, Zhangxin Chen, and Xiaolin Lin. A priori and a posteriori estimates of stabilized mixed finite volume methods for the incompressible flow arising in arteriosclerosis. Journal of Computational and Applied Mathematics, 363(??):35–52, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302407.

Li:2024:NDE

[LJL24] Yexin Li, Ping Jiang, and Haochen Li. A novel directly energy-preserving method for charged particle dynamics. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001183.

Lohi:2021:WFT

[LK21] Jonni Lohi and Lauri Kettunen. Whitney forms and their extensions. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001394.

Lee:2020:NME

[LKM20] Hosoo Lee, Hyun-Min Kim, and Jie Meng. On the nonlinear matrix equation X^{p} = A + M^{T} (X♯B)M. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303838.

Li:2020:SFO

[LL20a] W. Li and D. Liang. The spatial fourth-order compact splitting FDTD scheme with modified energy-conserved identity for two-dimensional Lorentz model. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304315.

Lin:2020:ASC

[LL20b] Fu-Rong Lin and Wei-Dong Liu. The accuracy and stability of CN-WSGD schemes for space fractional diffusion equation. Journal of Computational and Applied Mathematics, 363(??):77–91, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930281X.

Livinska:2021:AMN

[LL21] H. Livinska and E. Lebedev. On asymptotic merging of nodes set for multichannel stochastic networks. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001540.

Lu:2022:LID

[LL22] Lizheng Lu and Lin Li.

On the linear independence of the derivatives of Bernstein polynomials. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000115.

Lian:2023:NCF

[LL23] Wenhui Lian and Xinwu Liu. Non-convex fractionalorder TV model for impulse noise removal. Journal of Computational and Applied Mathematics, 417 (??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722003065.

Lee:2024:ROP

[LL24a] Sunju Lee and Younhee Lee. Real option pricing under the regimeswitching model with jumps on a finite time horizon. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001432.

Leung:2024:PES

[LL24b] Wing Tat Leung and Wenyuan Li. Partially explicit splitting method for a multi-physics problem. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005721.

Li:2024:OEA

[LL24c] Shiren Li and Yuan Li. Optimal error analysis of an unconditionally stable BDF2 finite element approximation for the 3D incompressible MHD equations with variable density. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000736.

Lim:2024:CIT

[LL24d] Hyuncheul Lim and Sungchul Lee. l_{1} -constrained implied transition densities. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004314.

Liu:2024:SAG

[LL24e] Cuiyun Liu and Bo Liu.

Symmetric and asymmetric Gauss and Gauss–Lobatto quadrature rules for triangles and their applications to high-order finite element analyses. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003953.

Liu:2024:BMC

[LL24f] Yulan Liu and Rongrong Lin. A bisection method for computing the proximal operator of the l_{p}- norm for any 0 < p < 1 with application to Schatten p-norms. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400147X.

Li:2020:GFE

[LLFT20] Lang Li, Fawang Liu, Libo Feng, and Ian Turner. A Galerkin finite element method for the modified distributed-order anomalous sub-diffusion equation. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305941.

Lyu:2022:CRR

[LLG^{+}22] Yanwei Lyu, Chunyang Liu, Xingyu Gao, Yang Liu, Wanli Ma, and Xiangtuo Chen. Corrigendum to “Research on regional differences and influencing factors of green technology innovation efficiency of China’s hightech industry” [J. Comput. Appl. Math. 369 (2020) 112597]. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004568.

See [LGMC20].

Li:2024:SFE

[LLG24] Na Li, Ping Lin, and Fuzheng Gao. A SAV finite element method for the Cahn–Hilliard equation with dynamic boundary conditions. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005289.

Li:2024:OEE

[LLGC24a] Wenjuan Li, Yunxian Liu, Fuzheng Gao, and Jintao Cui. Optimal L^{2} error estimates of stabilizerfree weak Galerkin finite element method for the drift-diffusion problem. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002334.

Li:2024:WGF

[LLGC24b] Wenjuan Li, Yunxian Liu, Fuzheng Gao, and Jintao Cui. A weak Galerkin finite element method for 1D semiconductor device simulation models. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004752.

Lorek:2020:TPG

[LLGZ20] Paweł Lorek, Grzegorz Loś, Karol Gotfryd, and Filip Zagórski. On testing pseudorandom generators via statistical tests based on the arcsine law. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302594.

Lin:2024:APE

[LLH24] Sha Lin, Xuanmeng Lin, and Xin-Jiang He. Analytically pricing European options with a two-factor Stein-Stein model. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006064.

Li:2023:MMA

[LLK23] Yuchen Li, Dan Liu, and Ibrahim Kucukkoc. Mixedmodel assembly line balancing problem considering learning effect and uncertain demand. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004216.

Liu:2022:PIA

[LLL22a] Guoxin Liu, Xiaoying Liu, and Zhaoyang Liu. The policy iteration algorithm for a compound Poisson process applied to optimal dividend strategies under a Cramér–Lundberg risk model. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001649.

Lyu:2022:SME

[LLL^{+}22b] Xing-Long Lyu, Tiexiang Li, Jia-Wei Lin, Tsung-Ming Huang, Wen-Wei Lin, and Heng Tian. Solving Maxwell eigenvalue problems for three dimensional isotropic photonic crystals with fourteen Bravais lattices. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000826.

Li:2024:IMP

[LLL24a] Da Li, Michael P. Lamoureux, and Wenyuan Liao. Incorporating multiple a priori information for inverse problem by inexact scaled gradient projection. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004041.

Lica:2024:NSQ

[LLL24b] Septimiu Lica, Dan Lascu, and Evelyn-Astrid Lovasz. A new step-up-down quadratic dc–dc converter with a single active switch. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003060.

Li:2020:TFV

[LLLD20] Wenbo Li, Jicheng Li, Xuenian Liu, and Liqiang Dong. Two fast vectorwise update algorithms for orthogonal nonnegative matrix factorization with sparsity constraint. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300765.

Liu:2020:CAF

[LLLL20] Li-Bin Liu, Zhifang Liang, Guangqing Long, and Ying Liang. Convergence analysis of a finite difference scheme for a Riemann– Liouville fractional derivative two-point boundary value problem on an adaptive grid. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030100X.

Lv:2021:BMD

[LLLY21] Xiao-Guang Lv, Jun Liu, Fang Li, and Xuan-Liang Yao. Blind motion deconvolution for binary images. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001199.

Li:2021:DPG

[LLS21] Danping Li, Bin Li, and Yang Shen. A dynamic pricing game for general insurance market. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306403.

Li:2022:SAI

[LLS22] Tingting Li, Jianfang Lu, and Chi-Wang Shu. Stability analysis of inverse Lax–Wendroff boundary treatment of high order compact difference schemes for parabolic equations. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003332.

Li:2021:EES

[LLT^{+}21] Xi Li, Tongmao Li, Rungting Tu, Kejia Pan, and Chuanjun Chen Xiaofeng Yang. Efficient energy stable scheme for volumeconserved phase-field elastic bending energy model of lipid vesicles. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304684.

Lai:2021:OPS

[LLW21] Chong Lai, Shican Liu, and Yonghong Wu. Optimal portfolio selection for a defined-contribution plan under two administrative fees and return of premium clauses. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003162.

Liao:2022:TMM

[LLW22] Juan Liao, Wei Liu, and Xiaoyan Wang. Truncated Milstein method for nonautonomous stochastic differential equations and its modification. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004398.

Li:2022:FSO

[LLX^{+}22a] Yibao Li, Rui Liu, Qing Xia, Chenxi He, and Zhong Li. First- and second-order unconditionally stable direct discretization methods for multi-component Cahn–Hilliard system on surfaces. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004003.

Liu:2022:EPB

[LLX^{+}22b] J. K. Liu, Z. L. Lu, J. L. Xu, S. Wu, and Z. W. Tu. An efficient projection-based algorithm without Lipschitz continuity for large-scale nonlinear pseudo-monotone equations. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004441.

Li:2021:CBA

[LLY21] Zihao Li, Ji Luo, and Jing Yao. Convex bound approximations for sums of random variables under multivariate log-generalized hyperbolic distribution and asymptotic equivalences. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000789.

Li:2021:HEL

[LLYM21] Qi Li, Xi Li, Xiaofeng Yang, and Liquan Mei. Highly efficient and linear numerical schemes with unconditional energy stability for the anisotropic phase-field crystal model. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304131.

Li:2024:VED

[LLZ24a] Yanwei Li, Huipo Liu, and Zhaojie Zhou. Virtual element discretization method to optimal control problem governed by Stokes equations with pointwise control constraint on arbitrary polygonal meshes. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002528.

Liu:2024:MQM

[LLZ24b] Heng-Li Liu, Quan-Lin Li, and Chi Zhang. Matched queues with matching batch pair (m, n). Journal of Computational and Applied Mathematics, 449 (??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724002280.

Li:2021:SOF

[LLZC21] Jian Li, Rui Li, Xin Zhao, and Zhangxin Chen. A second-order fractional time-stepping method for a coupled Stokes/Darcy system. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306208.

Li:2022:EFD

[LLZY22] Tongmao Li, Peng Liu, Jun Zhang, and Xiaofeng Yang. Efficient fully decoupled and second-order time-accurate scheme for the Navier–Stokes coupled Cahn–Hilliard Ohta– Kawaski phase-field model of diblock copolymer melt. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004660.

Luan:2021:EET

[LM21a] Vu Thai Luan and Dominik L. Michels. Efficient exponential time integration for simulating nonlinear coupled oscillators. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000480.

Lv:2021:ISI

[LM21b] Changqing Lv and Changfeng Ma. An iterative scheme for identifying the positive semi-definiteness of evenorder real symmetric Htensor. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001175.

Levitt:2024:RCR

[LM24] James Levitt and Per-Gunnar Martinsson. Randomized compression of rank-structured matrices accelerated with graph coloring. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002942.

Laudadio:2024:MPC

[LMAV24] Teresa Laudadio, Nicola Mastronardi, Walter Van Assche, and Paul Van Dooren. A Matlab package computing simultaneous Gaussian quadrature rules for multiple orthogonal polynomials. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003583.

Li:2021:ESO

[LML21] Qi Li, Liquan Mei, and Yibao Li. Efficient secondorder unconditionally stable numerical schemes for the modified phase field crystal model with longrange interaction. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306269.

Liu:2022:FOL

[LMLB22] Xiaodong Liu, Nathaniel R.

Morgan, Evan J. Lieberman, and Donald E. Burton. A fourth-order Lagrangian discontinuous Galerkin method using a hierarchical orthogonal basis on curvilinear grids. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005136.

Lai:2022:MAG

[LMM22] Kin Keung Lai, J. K. Maurya, and S. K. Mishra. Multiobjective approximate gradient projection method for constrained vector optimization: Sequential optimality conditions without constraint qualifications. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000243.

Lemaire:2020:NWE

[LMP20] Vincent Lemaire, Thibaut Montes, and Gilles Pagès. New weak error bounds and expansions for optimal quantization. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306752.

Liu:2021:BTR

[LMS21] Jie Liu, Matthias Möller, and Henk M. Schuttelaars. Balancing truncation and round-off errors in FEM: One-dimensional analysis. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305100.

Li:2021:PIM

[LMSZ21] Qi Li, Amitava Mukherjee, Zhi Song, and Jiujun Zhang. Phase-II monitoring of exponentially distributed process based on Type-II censored data for a possible shift in location-scale. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306063.

Lepe:2024:DGM

[LMV24] Felipe Lepe, David Mora, and Jesus Vellojin. Discontinuous Galerkin methods for the acoustic vibration problem. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300643X.

Li:2021:ODR

[LMYZ21] Peng Li, Qingbin Meng, Kam C. Yuen, and Ming Zhou. Optimal dividend and risk control policies in the presence of a fixed transaction cost. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305628.

Lewis:2022:CSA

[LMZ22] Thomas Lewis, Quinn Morris, and Yi Zhang. Convergence, stability analysis, and solvers for approximating sublinear positone and semipositone boundary value problems using finite difference methods. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005033.

Lu:2024:MSR

[LMZ24] Xiaofan Lu, Huimei Ma, and Linan Zhang. Model selection via reweighted partial sparse recovery. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042723004971.

Lui:2021:SCB

[LN21] S. H. Lui and Sarah Nataj. Superlinear convergence of Broyden’s method and BFGS algorithm using Kantorovich-type assumptions. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304957.

Lorvand:2022:GMS

[LN22] H. Lorvand and A. R.

Nematollahi. Generalized mixed δ-shock models with random interarrival times and magnitude of shocks. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004544.

Lundgren:2023:HOA

[LN23] Lukas Lundgren and Murtazo Nazarov. A highorder artificial compressibility method based on Taylor series time-stepping for variable density flow. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004447.

Liu:2020:NMI

[LNN20] Jingjing Liu, Anqi Ni, and Guoxi Ni. A nonconvex l_{1}(l_{1}-l_{2}) model for image restoration with impulse noise. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302259.

Lorvand:2019:AGD

[LNP19] H. Lorvand, A. R. Nematollahi, and M. H. Poursaeed. Assessment of a generalized discrete time mixed δ-shock model for the multi-state systems. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304182.

Liu:2022:DDP

[LNP22] Xijia Liu, Hiba Nassar, and Krzysztof Podgórski. Dyadic diagonalization of positive definite band matrices and efficient Bspline orthogonalization. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002102.

Lamnii:2022:GSQ

[LNSZ22] A. Lamnii, M. Y. Nour, D. Sbibih, and A. Zidna. Generalized spline quasiinterpolants and applications to numerical analysis. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000103.

Liu:2020:FPP

[LP20] Gang Liu and Saminathan Ponnusamy. Finite pairs of prescribed cycles of König’s and Steffensen’s methods for entire functions. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305540.

Leonenko:2023:TNA

[LP23a] G. Leonenko and T. N.

Phillips. Transient numerical approximation of hyperbolic diffusions and beyond. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004915.

Los:2023:SNL

[LP23b] Marcin Loś and Maciej Paszyński. Stability of non-linear flow in heterogeneous porous media simulations using higher order and continuity basis functions. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000857.

Li:2024:CTM

[LP24] Zhilin Li and Kejia Pan. Coupled transformation methods and analysis for BVPs on infinite domains. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000207.

Liegeois:2023:PPI

[LPH23] Kim Liegeois, Mauro Perego, and Tucker Hartland. PyAlbany: a Python interface to the C++ multiphysics solver Albany. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006355.

Lu:2021:HDV

[LPLP21] Shaojun Lu, Jun Pei, Xinbao Liu, and Panos M. Pardalos. A hybrid DBH-VNS for high-end equipment production scheduling with machine failures and preventive maintenance activities. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304866.

Lopez:2023:CAL

[LPP23] José L. López, Pedro J.

Pagola, and Pablo Palacios. A convergent and asymptotic Laplace method for integrals. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004952.

Lopez:2024:UAM

[LPP24] José L. López, Pedro J. Pagola, and Pablo Palacios. The uniform asymptotic method “saddle point near an end point” revisited. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272400013X.

Li:2020:RAP

[LPY20] Rong Li, Athanasios A. Pantelous, and Lin Yang. Robust analysis for premiumreserve models in a stochastic nonlinear discrete-time varying framework. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305977.

Liu:2023:TSP

[LQS23] Z. Liu, R. Quintanilla, and M. Summers. Two singular problems of dualphase-lag thermo-porous-elasticity with microtemperatures. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006276.

Lazarev:2022:OLF

[LR22] N. Lazarev and E. Rudoy.

Optimal location of a finite set of rigid inclusions in contact problems for inhomogeneous two-dimensional bodies. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003320.

Lv:2024:PRF

[LR24a] Deyong Lv and Hongxing Rui. Pressure-robust finite element scheme for the time-dependent fully coupled Stokes–Darcy-transport problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003388.

Lv:2024:PRM

[LR24b] Deyong Lv and Hongxing Rui. A pressurerobust mixed finite element method for the coupled Stokes–Darcy problem. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003886.

Lahtinen:2024:ESS

[LRP24] Joonas Lahtinen, Atena Rezaei, and Sampsa Pursiainen. Effects of source space resolution, randomization and averaging in focal EEG source localization: Estimation of forward and inverse errors. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001729.

Lopez-Rodriguez:2024:FEA

[LRQV24] Bibiana López-Rodríguez, José Querales, and Pablo Venegas. Finite element approximation for an axisymmetric timedependent acoustic problem. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001900.

Li:2020:MNR

[LRT20] Chunyan Li, Zemin Ren, and Liming Tang. Multiplicative noise removal via using nonconvex regularizers based on total variation and wavelet frame. Journal of Computational and Applied Mathematics, 370 (??):??, May 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306892.

Lepe:2023:EEV

[LRV23] Felipe Lepe, Gonzalo Rivera, and Jesus Vellojin. Error estimates for a vorticity-based velocitystress formulation of the Stokes eigenvalue problem. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200406X.

Lewis:2023:CRD

[LRZ23] Tom Lewis, Aaron Rapp, and Yi Zhang. Consistency results for the dual-wind discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002017.

Lan:2021:NAL

[LS21] Rihui Lan and Pengtao Sun. A novel arbitrary Lagrangian–Eulerian finite element method for a parabolic/mixed parabolic moving interface problem. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304167.

Lang:2022:DAI

[LS22a] Jens Lang and Bernhard A. Schmitt. Discrete adjoint implicit peer methods in optimal control. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002977.

Lubinsky:2022:SBP

[LS22b] D. S. Lubinsky and A. Sidi.

Some biorthogonal polynomials arising in numerical analysis and approximation theory. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004659.

Li:2023:EAF

[LS23] Xiaoli Li and Jie Shen.

Error analysis of a fully discrete consistent splitting MAC scheme for time dependent Stokes equations. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004903.

Luukka:2024:PMB

[LS24] Pasi Luukka and Jan Stoklasa. Possibilistic mean based defuzzification for fuzzy expert systems and fuzzy control-LSD for general fuzzy sets. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006076.

Lamichhane:2021:LPS

[LSC21] Bishnu P. Lamichhane and Jordan A. Shaw-Carmody. A local projection stabilisation finite element method for the Stokes equations using biorthogonal systems. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001618.

Lazarev:2024:EPI

[LSE24] N. Lazarev, G. Semenova, and E. Efimova. Equilibrium problem for an inhomogeneous twodimensional elastic body with two interacting thin rigid inclusions. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004831.

Liu:2024:HOM

[LST^{+}24] Chenguang Liu, Jie Sun, Hao Tian, Wai Sun Don, and Lili Ju. A high-order multi-time-step scheme for bond-based peridynamics. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002188.

Ling:2023:LDG

[LSY23] Dan Ling, Chi-Wang Shu, and Wenjing Yan. Local discontinuous Galerkin methods for diffusiveviscous wave equations. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200348X.

Li:2024:IRM

[LSY24] Zhong Li, Kristina P.

Sendova, and Chen Yang. On an insurance ruin model with a causal dependence structure and perturbation. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002206.

Liu:2022:KFF

[LSZ22] Wei Liu, Peng Shi, and Huiyan Zhang. Kalman filtering with finite-step autocorrelated measurement noise. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000334.

Lin:2023:FEV

[LSZ23] Wei Lin, Kangli Shen, and Jin E. Zhang. Further exploration into the valid regions of Gram–Charlier densities. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001759.

Lujano:2021:SOM

[LT21] John Lujano and Johannes Tausch. A shape optimization method for moving interface problems governed by the heat equation. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305574.

Li:2024:NCM

[LT24a] Yanyan Li and Tao Tan.

A new convex model for linear hyperspectral unmixing. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006519.

Liu:2024:MSM

[LT24b] Hongliang Liu and Tan Tan. Midpoint splitting methods for nonlinear space fractional diffusion equations. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001109.

Liu:2020:FNR

[LTF^{+}20] Xuanyu Liu, Guo-Liang Tian, Yu Fei, Lianjie Shu, and Qiang Zhao. Folded normal regression models with applications in biomedicine. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302326.

Lopes:2024:SIG

[LTL^{+}24] Tito Lopes, Vera L. D. Tomazella, Jeremias Leão, Pedro L. Ramos, and Francisco Louzada. Statistical inference for Generalized Power-Law Process in repairable systems. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000487.

Liu:2024:DGM

[LTZ24] Sijing Liu, Zhiyu Tan, and Yi Zhang. Discontinuous Galerkin methods for an elliptic optimal control problem with a general state equation and pointwise state constraints. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004387.

Lu:2024:CSY

[Lu24] Linzhang Lu. Constructing solutions of the Yang– Baxter-like matrix equation for singular matrices. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003473.

Luckraz:2023:CLP

[Luc23] Shravan Luckraz. On a contraction-like property of dismantlable graphs. Journal of Computational and Applied Mathematics, 423 (??):??, May 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005908.

Lukyanov:2020:SSD

[LV20] Alexander A. Lukyanov and Cornelis Vuik. A stable SPH discretization of the elliptic operator with heterogeneous coefficients. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300364.

Liu:2023:PBG

[LV23] Xuefeng Liu and Tomás Vejchodský. Projectionbased guaranteed L^{2} error bounds for finite element approximations of Laplace eigenfunctions. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001085.

Li:2024:HOT

[LVW24] Yukun Li, Liet Vo, and Guanqian Wang. Higher order time discretization method for a class of semilinear stochastic partial differential equations with multiplicative noise. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003862.

Liu:2021:EPN

[LW21] Yujie Liu and Junping Wang. An extended P1nonconforming finite element method on general polytopal partitions. Journal of Computational and Applied Mathematics, 381 (??):??, January 1, 2021.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720303125.

Li:2022:NBM

[LW22] Cui-Xia Li and Shi-Liang Wu. Newton-based matrix splitting iteration methods for the weakly nonlinear system. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000863.

Li:2023:MBM

[LW23a] Cui-Xia Li and Shi-Liang Wu. Modulus-based matrix splitting methods for complex linear complementarity problem. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000833.

Li:2023:HOD

[LW23b] Xiaosheng Li and Wei Wang. A high order discontinuous Galerkin method for the recovery of the conductivity in Electrical Impedance Tomography. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002881.

Lin:2024:PTP

[LW24] Xue-Lei Lin and Shu-Lin Wu. A parallel-in-time preconditioner for Crank– Nicolson discretization of a parabolic optimal control problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003558.

Li:2022:LPE

[LWA^{+}22] Jian Li, Xue Wang, Md. Abdullah Al Mahbub, Haibiao Zheng, and Zhangxin Chen. Local and parallel efficient BDF2 and BDF3 rotational pressurecorrection schemes for a coupled Stokes/Darcy system. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001406.

Li:2020:TMD

[LWC20] Hailin Li, Yenchun Jim Wu, and Yewang Chen. Time is money: Dynamicmodel-based time series data-mining for correlation analysis of commodity sales. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306648.

Li:2020:NASa

[LWD20] Tao Li, Qing-Wen Wang, and Xue-Feng Duan. Numerical algorithms for solving discrete Lyapunov tensor equation. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306818.

Li:2024:NNS

[LWG24] Ting Li, Zhong Wan, and Jie Guo. A new nonmonotone spectral projected gradient algorithm for box-constrained optimization problems in m × n real matrix space with application in image clustering. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005071.

Li:2022:RGT

[LWL22] Dong-Kai Li, Li Wang, and Yu-Ying Liu. A relaxation general twosweep modulus-based matrix splitting iteration method for solving linear complementarity problems. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000346.

Lin:2023:EMM

[LWLD23] Huizhong Lin, Liang Wang, Yuhlong Lio, and Sanku Dey. Estimation of Matusita measure between generalized inverted exponential distributions under progressive first-failure censored data. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004344.

Li:2023:MSG

[LWLW23] Dandan Li, Jiaqi Wu, Yong Li, and Songhua Wang. A modified spectral gradient projection-based algorithm for large-scale constrained nonlinear equations with applications in compressive sensing. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006045.

Liang:2024:ABT

[LWT24] Zhao-Zheng Liang, Hong-Yi Wan, and Jun-Lin Tian. Additional block triangular preconditioners for coupled block two-by-two linear systems arising from Stokes control problems. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001717.

Li:2022:PDF

[LWW22a] Dan Li, Chunmei Wang, and Junping Wang. A primal-dual finite element method for transport equations in non-divergence form. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001339.

Liu:2022:ISP

[LWW^{+}22b] Chun Liu, Cheng Wang, Steven M. Wise, Xingye Yue, and Shenggao Zhou. An iteration solver for the Poisson–Nernst–Planck system and its convergence analysis. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005999.

Li:2023:PDF

[LWW23a] Dan Li, Chunmei Wang, and Junping Wang. An L^{p}-primal–dual finite element method for firstorder transport problems. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002893.

Li:2023:GWG

[LWW23b] Dan Li, Chunmei Wang, and Junping Wang. Generalized weak Galerkin finite element methods for biharmonic equations. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002972.

Liu:2022:RBP

[LWWN22] Ying Liu, Gang Wang, Mengyao Wu, and Yufeng Nie. A recovery-based a posteriori error estimator of the weak Galerkin finite element method for elliptic problems. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005471.

Li:2024:GWG

[LWWY24] Dan Li, Chunmei Wang, Junping Wang, and Xiu Ye. Generalized weak Galerkin finite element methods for second order elliptic problems. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000827.

Li:2024:HOM

[LWWZ24] Dan Li, Chunmei Wang, Junping Wang, and Shangyou Zhang. High order Morley elements for biharmonic equations on polytopal partitions. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000062.

Luo:2024:NZN

[LWX24] Tangtao Luo, Guancheng Wang, and Xiuchun Xiao. New zeroing NN models with nonconvex saturated activation functions in noisy environments for quadratic minimization dynamics and control. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400133X.

Liu:2021:ANS

[LWXL21] Changli Liu, Wei-Guo Wang, Jungong Xue, and Ren-Cang Li. Accurate numerical solution for structured M-matrix algebraic Riccati equations. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002363.

Li:2021:RFE

[LWYZ21] Hengguang Li, Xiang Wan, Peimeng Yin, and Lewei Zhao. Regularity and finite element approximation for two-dimensional elliptic equations with line Dirac sources. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001370.

Luo:2021:SDR

[LWZ21] Shangzhen Luo, Mingming Wang, and Wei Zhu. Stochastic differential reinsurance games in diffusion approximation models. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305434.

Li:2024:WGMb

[LWZ24] Dan Li, Chunmei Wang, and Shangyou Zhang. Weak Galerkin methods for elliptic interface problems on curved polygonal partitions. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002450.

Li:2021:RPM

[LWZF21] Jia Li, Xu Wu, Linlin Zhang, and Qianying Feng. Research on the portfolio model based on mean-MF– DCCA under multifractal feature constraint. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305550.

Li:2024:OCE

[LX24] Yujie Li and Chuanju Xu. Optimal control of an evolutionary variationalhemivariational inequality model with new results in numerical analysis and control of the dynamic frictional contact problem. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001912.

Lu:2021:SON

[LXF21] Shujuan Lü, Tao Xu, and Zhaosheng Feng. A secondorder numerical method for space-time variable-order diffusion equation. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030649X.

Li:2023:DDD

[LXLL23] Sen Li, Yingzhi Xia, Yu Liu, and Qifeng Liao. A deep domain decomposition method based on Fourier features. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005611.

Li:2021:FLD

[LXQ21] Gang Li, Yinghong Xu, and Zhenhua Qin. Fenchel– Lagrange duality for DC infinite programs with inequality constraints. Journal of Computational and Applied Mathematics, 391 (??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721000455.

Liu:2024:HOP

[LXY24] Xinyuan Liu, Tao Xiong, and Yang Yang. High order positivity-preserving nodal discontinuous Galerkin methods for anisotropic diffusion problems. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300571X.

Luo:2024:CPA

[LXYL24] Qian Luo, Aiguo Xiao, Xiaoqiang Yan, and Guidong Liu. A class of polynomial approximation methods to second-order delay differential equations. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003340.

Lu:2023:EEL

[LXZ23] Jinyang Lu, Yan Xu, and Chao Zhang. Error estimates of the local discontinuous Galerkin methods for two-dimensional (μ)-Camassa–Holm equations. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200365X.

Li:2020:EMS

[LY20] Xiaoyue Li and George Yin. Explicit Milstein schemes with truncation for nonlinear stochastic differential equations: convergence and its rate. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300625.

Liu:2021:UES

[LY21] Hailiang Liu and Peimeng Yin. Unconditionally energy stable discontinuous Galerkin schemes for the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030666X.

Li:2022:PEE

[LY22a] Fei Li and Nianyu Yi. A posteriori error estimates of goal-oriented adaptive finite element methods for nonlinear reaction–diffusion problems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001613.

Liu:2022:HOU

[LY22b] Hailiang Liu and Peimeng Yin. High order unconditionally energy stable RKDG schemes for the Swift–Hohenberg equation. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005987.

Liu:2023:OPE

[LY23a] Huini Liu and Nianyu Yi.

Optimal a priori error estimate of relaxation-type linear finite element method for nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000912.

Long:2023:SAI

[LY23b] Tao Long and Yuexin Yu. Stability analysis of implicit-explicit θ-methods for composite stiff neutral functional differential equations in Banach space. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003879.

Li:2024:TMV

[LY24a] Pingyun Li and Chuancun Yin. The Tail Mean-Variance optimal capital allocation under the extended skew-elliptical distribution. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002152.

Liu:2024:DAE

[LY24b] Hanze Liu and Adilai Yusupu. Dynamical analysis and explicit traveling wave solutions to the higher-dimensional generalized nonlinear wave system. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000748.

Li:2022:FTT

[LYB22] Lingjie Li, Wenjian Yu, and Kim Batselier. Faster tensor train decomposition for sparse data. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005720.

Lee:2023:ETS

[LYK23] Seunggyu Lee, Sungha Yoon, and Junseok Kim. Effective time step analysis of convex splitting schemes for the Swift–Hohenberg equation. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003600.

Liu:2023:FSA

[LYL23] Fuhao Liu, Wei Yang, and Jichun Li. A FETD scheme and analysis for photonic crystal waveguides comprising thirdorder nonlinear and linear materials. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006033.

Liu:2024:ABF

[LYL^{+}24] Xuelong Liu, Guoju Ye, Wei Liu, Yating Guo, and Fangfang Shi. On Atangana–Baleanu fractional granular calculus and its applications to fuzzy economic models in market equilibrium. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272400267X.

Lee:2022:FES

[LYM22] Sanghyun Lee, Hyun Chul Yoon, and S. M. Mallikarjunaiah. Finite element simulation of quasi-static tensile fracture in nonlinear strain-limiting solids with the phase-field approach. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100337X.

Lin:2024:PGC

[LYYZ24] Fubiao Lin, Yang Yang, Xinxia Yang, and Qianhong Zhang. Preliminary group classification and exact solutions of Smoluchowski equation with a source. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000190.

Liu:2024:SWG

[LYZ24] Xiaowei Liu, Min Yang, and Jin Zhang. Supercloseness of weak Galerkin method for a singularly perturbed convection–diffusion problem in 2D. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003485.

Lin:2019:OEB

[LZ19] Tao Lin and Qiao Zhuang.

Optimal error bounds for partially penalized immersed finite element methods for parabolic interface problems. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304042.

Li:2020:NASb

[LZ20a] Zhaoxiang Li and Jianxin Zhou. A new augmented singular transform and its partial Newtoncorrection method for finding more solutions to nonvariational quasilinear elliptic PDEs. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301126.

Liu:2020:SIM

[LZ20b] J. K. Liu and L. Zheng. A smoothing iterative method for the finite minimax problem. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300327.

Lu:2020:CTD

[LZ20c] Qinyun Lu and Yuanguo Zhu. Comparison theorems and distributions of solutions to uncertain fractional difference equations. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301758.

Liang:2021:CAI

[LZ21] Zhao-Zheng Liang and Guo-Feng Zhang. On Chebyshev accelerated iteration methods for twoby-two block linear systems. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000686.

Lin:2022:VRG

[LZ22] Wei Lin and Jin E.

Zhang. The valid regions of Gram–Charlier densities with high-order cumulants. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005586.

Li:2023:LOP

[LZ23a] Ruo Li and Wei Zhong.

Locally order-preserving mapping for WENO methods. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006021.

Liu:2023:AWG

[LZ23b] Kaifang Liu and Peng Zhu. Analysis of a weak Galerkin method for second-order elliptic equations with minimal regularity on polytopal meshes. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003752.

Li:2024:GNN

[LZ24a] Rong Li and Bing Zheng. Generalized nonconvex nonsmooth four-directional total variation with overlapping group sparsity for image restoration. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002954.

Liu:2024:DDS

[LZ24b] Jianzhou Liu and Wenlong Zeng. The dominant degree for Schur complement of S-strictly diagonally dominant matrix and its applications. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003169.

Li:2021:NCA

[LZD21] Xiangli Li, Wenjuan Zhao, and Xiaoliang Dong. A new CG algorithm based on a scaled memoryless BFGS update with adaptive search strategy, and its application to large-scale unconstrained optimization problems. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002922.

Li:2020:RBU

[LZF20] Ya-Jing Li, Xin-Yun Zhu, and Hong-Tao Fan. Relaxed block upper-lower triangular preconditioner for generalized saddle point problems from the incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303267.

Long:2024:AIN

[LZG24] Haie Long, Ye Zhang, and Guangyu Gao. An accelerated inexact Newtontype regularizing algorithm for ill-posed operator equations. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003029.

Li:2021:TGM

[LZH21] Yixin Li, Bangjian Zhou, and Xianliang Hu. A two-grid method for levelset based topology optimization with GPUacceleration. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306270.

Li:2021:EIL

[LZL21] Dongping Li, Xiuying Zhang, and Renyun Liu. Exponential integrators for large-scale stiff Riccati differential equations. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306518.

Li:2023:BAC

[LZL^{+}23] Bo Li, Yue Zhang, Xiaoliang Li, Zohreh Eskandari, and Qizhi He. Bifurcation analysis and complex dynamics of a Kopel triopoly model. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300033X.

Liu:2020:IAB

[LZLL20] Qilong Liu, Jianxing Zhao, Chaoqian Li, and Yaotang Li. An iterative algorithm based on strong Htensors for identifying positive definiteness of irreducible homogeneous polynomial forms. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305862.

Li:2020:COD

[LZW20] Jiu Li, Hongyan Zang, and Xinyuan Wei. On the construction of onedimensional discrete chaos theory based on the improved version of Marotto’s theorem. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302430.

Li:2022:CBS

[LZW^{+}22] Xiaowei Li, Yufan Zhu, Haiyan Wu, Jinlan Xu, Ran Ling, Xiangyang Wu, and Gang Xu. Construction of Bézier surfaces with energy-minimizing diagonal curves from given boundary. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001716.

Li:2021:DFV

[LZWC21] Rui Li, Yongchao Zhang, Jianhua Wu, and Zhangxin Chen. Discontinuous finite volume element method for Darcy flows in fractured porous media. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303162.

Liu:2023:AGT

[LZWL23] Jinkui Liu, Ning Zhang, Jing Wang, and Zuliang Lu. An approximate gradient-type method for nonlinear symmetric equations with convex constraints. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002030.

Li:2020:ISQ

[LZZ20] Lidan Li, Liwei Zhang, and Hongwei Zhang. Inverse semidefinite quadratic programming problem with l_{1} norm measure. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301291.

Li:2024:GQC

[LZZ24] Jin Li, Yuxin Zhang, and Xiaolei Zhang. Gaussian quadrature for certain two-dimensional hypersingular integrals. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003510.

Liang:2021:TSA

[LZZZ21] Maolin Liang, Bing Zheng, Yutao Zheng, and Ruijuan Zhao. A two-step accelerated Levenberg– Marquardt method for solving multilinear systems in tensor-train format. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303605.

Mirzaee:2019:CBS

[MA19] Farshid Mirzaee and Sahar Alipour. Cubic Bspline approximation for linear stochastic integrodifferential equation of fractional order. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304431.

Mirzaee:2021:QBS

[MA21a] Farshid Mirzaee and Sahar Alipour. Quintic B-spline collocation method to solve n-dimensional stochastic Itô–Volterra integral equations. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304441.

Mnatsakanov:2021:RBF

[MA21b] Robert M. Mnatsakanov and Rafik H. Aramyan. Recovery of bivariate functions from the values of its Radon transform using Laplace inversion. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001771.

Ma:2024:MFE

[Ma24a] Erfang Ma. A mixed finite element method for the forward problem of electrical impedance tomography with the shunt model. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003443.

Ma:2024:CRQ

[Ma24b]  Junjie Ma.  A class

of reducible quadrature rules for the second-kind Volterra integral equations using barycentric rational interpolation. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000529.

Moysi:2022:LCC

[MAA^{+}22] A. Moysi, M. Argyros, I. K. Argyros, Á.A. Magreñán, Í. Sarría, and D. González. Local convergence comparison between frozen Kurchatov and Schmidt-schwetlick-Kurchatov solvers with applications. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100011X.

Magherini:2020:CSM

[Mag20] Cecilia Magherini. A corrected spectral method for Sturm–Liouville problems with unbounded potential at one endpoint. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303528.

Majidian:2023:AFR

[Maj23a] Hassan Majidian. Adaptive FCC+ rules for oscillatory integrals. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006100.

Majidian:2023:ECF

[Maj23b] Hassan Majidian. Efficient construction of FCC+ rules. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002953.

Mamehrashi:2023:RAM

[Mam23] Kamal Mamehrashi. Ritz approximate method for solving delay fractional optimal control problems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003028.

Marrero:2019:LTS

[Mar19] J. Abderramán Marrero. On linear-time solvers for comrade linear systems. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304248.

Martinez:2023:ZDC

[MARH^{+}23] Noé Martínez, Alejandro Arceo, Andrea Rodríguez-Hernández, Luis E. Garza, and Gerardo Romero. Zero dynamics for a class of robustly stable polynomials. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005702.

Maset:2021:RES

[Mas21] Stefano Maset. Relative error stability and instability of matrix exponential approximations for stiff numerical integration of long-time solutions. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000066.

Mueller:2024:MOR

[MB24] Nicholas Mueller and Santiago Badia. Model order reduction with novel discrete empirical interpolation methods in spacetime. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000165.

Morelli:2021:DPA

[MBG21a] Myles Morelli, Tommaso Bellosta, and Alberto Guardone. Development and preliminary assessment of the open-source CFD toolkit SU2 for rotorcraft flows. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306312.

Morelli:2021:ERB

[MBG21b] Myles Morelli, Tommaso Bellosta, and Alberto Guardone. Efficient radial basis function mesh deformation methods for aircraft icing. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001114.

Mohebujjaman:2024:DAN

[MBKK24] Muhammad Mohebujjaman, Clarisa Buenrostro, Md. Kamrujjaman, and Taufiquar Khan. Decoupled algorithms for nonlinearly coupled reaction– diffusion competition model with harvesting and stocking. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003655.

Meng:2023:SOS

[MBZ23] Xiangjun Meng, Xuelian Bao, and Zhengru Zhang. Second order stabilized semi-implicit scheme for the Cahn–Hilliard model with dynamic boundary conditions. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000894.

McAllister:2021:RPE

[McA21] Tyrrell B. McAllister.

Rational polytopes with Ehrhart coefficients of arbitrary period. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304465.

Ma:2024:ADR

[MCD24] Huanhuan Ma, Jingrun Chen, and Jiansong Deng. Analysis of the dynamic response for Kirchhoff plates by the elementfree Galerkin method. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400342X.

Meng:2022:ENI

[MCL22] Zhaoliang Meng, Jintao Cui, and Zhongxuan Luo. Effect of numerical integration on a new rotated nonconforming quadrilateral element. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004209.

Morales-Castillo:2023:FSL

[MCMR23] Javier Morales-Castillo and Eva M. Morales-Rodriguez. A formula to solve Laplace and Fourier transforms. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001796.

Mu:2024:PCI

[MCW24] Wanrong Mu, Sung Nok Chiu, and Guojing Wang. Pricing CDS index tranches under thinning-dependence structure with regime switching. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003297.

Ming:2023:OSS

[MCY23] Hao Ming, Yinjun Chen, and Hu Yang. Onestep sparse estimates in the reverse penalty for high-dimensional correlated data. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000638.

Macias-Diaz:2020:DCE

[MD20] J. E. Macías-Díaz. A dynamically consistent exponential scheme to solve some advection–reaction equations with Riesz anomalous diffusion. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302119.

Macias-Diaz:2022:DAD

[MD22] J. E. Macías-Díaz. Design and analysis of a dissipative scheme to solve a generalized multi-dimensional Higgs boson equation in the de Sitter spacetime. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304118.

Mohammadi:2021:NSP

[MDD21] Vahid Mohammadi, Mehdi Dehghan, and Stefano De Marchi. Numerical simulation of a prostate tumor growth model by the RBF-FD scheme and a semi-implicit time discretization. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306051.

Mohammadi:2024:FRO

[MDE24] Maryam Mohammadi, Stefano De Marchi, and Mohammad Karimnejad Esfahani. Full-rank orthonormal bases for conditionally positive definite kernelbased spaces. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000104.

Macias-Diaz:2022:DNA

[MDG22] J. E. Macías-Díaz and A. Gallegos. Design and numerical analysis of a logarithmic scheme for nonlinear fractional diffusion– reaction equations. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030409X.

Mohapatra:2020:NCL

[MDKG20] S. Mohapatra, P. Dutt, B. V. Rathish Kumar, and Marc I. Gerritsma. Nonconforming least-squares spectral element method for Stokes equations on non-smooth domains. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307010.

Miao:2023:IPI

[MDLY23] Guangqin Miao, Feng Ding, Qinyao Liu, and Erfu Yang. Iterative parameter identification algorithms for transformed dynamic rational fraction inputoutput systems. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002418.

Macias-Diaz:2022:TEP

[MDMS22] Jorge E. Macías-Díaz, Romeo Martínez, and Qin Sheng. Two energypreserving numerical models for a multi-fractional extension of the Klein– Gordon–Zakharov system. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721006026.

Maurya:2020:MSO

[MDS20] Rahul Kumar Maurya, Vinita Devi, and Vineet Kumar Singh. Multistep schemes for one and two dimensional electromagnetic wave models based on fractional derivative approximation. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302764.

Macias-Diaz:2024:SCF

[MDSRFO24] J. E. Macías-Díaz, Adán J.

Serna-Reyes, and Luis A. Flores-Oropeza. A stable and convergent finitedifference model which conserves the positivity and the dissipativity of Gibbs’ free energy for a nonlinear combustion equation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004363.

Macias-Diaz:2022:ASN

[MDVR22] J. E. Macías-Díaz and Héctor Vargas-Rodríguez. Analysis and simulation of numerical schemes for nonlinear hyperbolic predatorprey models with spatial diffusion. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002582.

Ma:2023:NIC

[MDW23] Wanli Ma, Weiyang Ding, and Yimin Wei. Noda iteration for computing generalized tensor eigenpairs. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002285.

Meng:2020:SCN

[MDY20] Qingle Meng, Huaian Diao, and Qinghua Yu. Structured condition number for multiple right-hand side linear systems with parameterized quasiseparable coefficient matrix. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305321.

Mohammady:2020:ETR

[ME20] Somaieh Mohammady and M. R. Eslahchi. Extension of Tikhonov regularization method using linear fractional programming. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930682X.

Mencik:2024:IMR

[Men24] Jean-Mathieu Mencik. Improved model reduction with basis enrichment for dynamic analysis of nearly periodic structures including substructures with geometric changes. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000931.

Moghadam:2020:CMP

[MER20] Mostafa Abdolahi Moghadam, Seyed Babak Ebrahimi, and Donya Rahmani. A constrained multi-period robust portfolio model with behavioral factors and an interval semi-absolute deviation. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300339.

Ma:2021:HDG

[MF21] Weijun Ma and Shubin Fu. A hybridizable discontinuous Galerkin generalized multiscale finite element method for highly heterogeneous linear elasticity problems. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304155.

Moisa:2024:SOS

[MF24] Andrew Moisa and Boris Faleichik. Second order stabilized two-step Runge– Kutta methods. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004089.

Moradi:2021:OCS

[MFA21] A. Moradi, J. Farzi, and A. Abdi. Order conditions for second derivative general linear methods. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304911.

Misaii:2024:OTD

[MFH24] Hasan Misaii, Mitra Fouladirad, and Firoozeh Haghighi. Optimal task-driven timedependent covariate-based maintenance policy. Journal of Computational and Applied Mathematics, 435 (??):??, ???? 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002595.

Mittal:2022:CRI

[MG22] Gaurav Mittal and Ankik Kumar Giri. Convergence rates for iteratively regularized Gauss–Newton method subject to stability constraints. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003666.

Mittal:2023:MSD

[MG23] Gaurav Mittal and Ankik Kumar Giri. A modified steepest descent method for solving non-smooth inverse problems. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005957.

Marquez:2024:CLS

[MGA24] Almudena P. Márquez, María L. Gandarias, and Stephen C. Anco. Conservation laws, symmetries, and line solitons of a Kawahara–KP equation. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003564.

Ma:2022:SFW

[MGD22] Jie Ma, Fuzheng Gao, and Ning Du. Stabilizer-free weak Galerkin finite element method with secondorder accuracy in time for the time fractional diffusion equation. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200187X.

M:2024:OCO

[MGG^{+}24] Muniyasamy M., Chandhini G., Santhosh George, Indra Bate, and Kedarnath Senapati. On obtaining convergence order of a fourth and sixth order method of Hueso et al. without using Taylor series expansion. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003856.

Marion:2020:ACV

[MGL20] Pierre Marion, Maxime Godin, and Pierre L’Ecuyer. An algorithm to compute the t-value of a digital net and of its projections. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306740.

Miles:2023:ASF

[MGV23] Victoria Miles, Stefano Giani, and Oliver Vogt. Approaching STEP file analysis as a language processing task: a robust and scale-invariant solution for machining feature recognition. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001103.

Ma:2021:RHO

[MH21] Xiaohua Ma and Chengming Huang. Recovery of high order accuracy in spectral collocation method for linear Volterra integral equations of the third-kind with non-smooth solutions. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000777.

Maierhofer:2022:ALS

[MH22] Georg Maierhofer and Daan Huybrechs. An analysis of least-squares oversampled collocation methods for compactly perturbed boundary integral equations in two dimensions. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002424.

Mi:2023:CMS

[MH23] Jian Mi and Jin Huang. Collocation method for solving two-dimensional nonlinear Volterra–Fredholm integral equations with convergence analysis. Journal of Computational and Applied Mathematics, 428 (??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723001322.

Minh:2020:RIW

[MHHV20] Triet Le Minh, Quan Pham Hoang, Phong Luu Hong, and Canh Vo Van. Recovering the initial wave amplitude for nonlinear elliptic equation with locally Lipschitz source in multiple-dimensional domain. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301680.

He:2020:LUL

[mHLZ20] Wen ming He, Runchang Lin, and Zhimin Zhang. Local ultraconvergence of linear and bilinear finite element method for second order elliptic problems. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300066.

Marinca:2023:ASS

[MHM23] Bogdan Marinca, Nicolae Herisanu, and Vasile Marinca. Analytical solutions for solving unsteady flow of gas through a porous medium by using auxiliary functions method. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002406.

Ma:2023:ELR

[MHW23] Xueshuang Ma, Shenglong Hu, and Jie Wang. Efficient low-rank regularizationbased algorithms combining advanced techniques for solving tensor completion problems with application to color image recovering. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005453.

Matsue:2017:CLF

[MHY17] Kaname Matsue, Tomohiro Hiwaki, and Nobito Yamamoto. On the construction of Lyapunov functions with computer assistance. Journal of Computational and Applied Mathematics, 319(??):385–412, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042717300067.

See errata [MHY21].

Matsue:2021:ECL

[MHY21] Kaname Matsue, Tomohiro Hiwaki, and Nobito Yamamoto. Errata to “On the construction of Lyapunov functions with computer assistance” [j. comp. appl. math. 319 (2017) 385–412]. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304660.

See [MHY17].

Mashayekhi:2020:LSF

[MI20] S. Mashayekhi and L. Gr.

Ixaru. The least-squares fit of highly oscillatory functions using Eta-based functions. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301308.

Miao:2020:CEH

[Mia20] Cun-Qiang Miao. Computing eigenpairs of Hermitian matrices in augmented Krylov subspace produced by Rayleigh quotient iterations. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301138.

Miao:2024:MIF

[Mia24] Shuai Miao. A mimetic interpolation-free cell-centered finite volume scheme for the 2D and 3D heterogeneous anisotropic diffusion equations. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000098.

Martinez:2023:DFE

[MIdMR23] Sergio Martínez, María D. Illescas, and María del Mar Rueda. Distribution function estimation with calibration on principal components. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001334.

Martinez:2024:CED

[MIdMR24] Sergio Martínez, María D.

Illescas, and María del Mar Rueda. Calibration estimation of distribution function based on multidimensional scaling of auxiliary information. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001250.

Miklavcic:2023:ANS

[Mik23] M. Miklavcic. Analytic and numeric solutions of moving boundary problems. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002145.

Miles:2022:LTL

[Mil22]  Justin Miles.  On the

Laplace transform of the lognormal distribution: Analytic continuation and series approximations. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005276.

Miminis:2020:FSA

[Mim20] George Miminis. The feedback search algorithm: estimating the distance of semi-explicit systems to Runcontrollability. Journal of Computational and Applied Mathematics, 373 (??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042719303048.

Mincsovics:2020:DCC

[Min20] Miklós E. Mincsovics. Discrete C^{1} convergence of linear multistep methods. Journal of Computational and Applied Mathematics, 363(??):234–240, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719302730.

Mitra:2020:DRM

[Mit20]  Sovan Mitra.  Down-

side risk measurement in regime switching stochastic volatility. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301369.

Miyajima:2021:VCR

[Miy21] Shinya Miyajima. Verified computation of real powers of matrices. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000509.

Miyajima:2022:FVC

[Miy22] Shinya Miyajima. Fast verified computation for positive solutions to M-tensor multi-linear systems and Perron vectors of a kind of weakly irreducible nonnegative tensors. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005495.

Masson:2020:BHI

[MJ20] Yannick Masson and Bert Jüttler. Bivariate Hermite interpolation by a limiting case of the cross approximation algorithm. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306399.

Mokris:2024:ULR

[MJ24] Dominik Mokris and Bert Jüttler. Using low-rank approximations of gridded data for spline surface fitting. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004636.

Motezaker:2020:ADC

[MJK20] Mohsen Motezaker, Majid Jamali, and Reza Kolahchi. Application of differential cubature method for nonlocal vibration, buckling and bending response of annular nanoplates integrated by piezoelectric layers based on surface-higher order nonlocal-piezoelasticity theory. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306302.

Modebei:2020:BHM

[MJR20] Mark I. Modebei, S. N. Jator, and Higinio Ramos. Block hybrid method for the numerical solution of fourth order boundary value problems. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301679.

Malhotra:2024:ESS

[MK24] Astha Malhotra and Deepak Kumar. Existence and stability of solution for a nonlinear Volterra integral equation with binary relation via fixed point results. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006295.

Ma:2020:POD

[MkCZ20] Dingjiong Ma, Wai ki Ching, and Zhiwen Zhang. Proper orthogonal decomposition method for multiscale elliptic PDEs with random coefficients. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306405.

Mandal:2020:CAD

[MKN20] Moumita Mandal, Kapil Kant, and Gnaneshwar Nelakanti. Convergence analysis for derivative dependent Fredholm–Hammerstein integral equations with Green’s kernel. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306041.

Makouei:2021:MOS

[MKS21] Roghaye Makouei, Hossein Jabbari Khamnei, and Mahdi Salehi. Moments of order statistics and k-record values arising from the complementary beta distribution with application. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000054.

Mu:2023:HTB

[ML23] Yu Mu and Wing-Cheong Lo. Hopf and Turing bifurcation for a competition and cooperation system with spatial diffusion effect. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005222.

Mukherjee:2022:AEU

[MLS22] Amitava Mukherjee, Qi Li, and Zhi Song. An assessment of the effect of using different mappings and Minkowski distances in joint monitoring of the time-between-event processes. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003988.

Ming:2022:LPR

[MLY22] Hao Ming, Huilan Liu, and Hu Yang. Least product relative error estimation for identification in multiplicative additive models. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005094.

Man:2021:GMF

[MLYZ21] Jun Man, Guang Lin, Yijun Yao, and Lingzao Zeng. A generalized multi-fidelity simulation method using sparse polynomial chaos expansion. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002351.

Mackenzie:2020:HAM

[MM20a] J. A. Mackenzie and W. R. Mekwi. An hr-adaptive method for the cubic nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303176.

Maday:2020:APA

[MM20b] Y. Maday and O. Mula. An adaptive parareal algorithm. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302065.

Matsuda:2021:GPR

[MM21] Takeru Matsuda and Yuto Miyatake. Generalization of partitioned Runge– Kutta methods for adjoint systems. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305999.

Mai:2022:TFC

[MM22a] Tina Mai and Daniele Mortari. Theory of functional connections applied to quadratic and nonlinear programming under equality constraints. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005355.

Meng:2022:OOC

[MM22b] Jian Meng and Liquan Mei. The optimal order convergence for the lowest order mixed finite element method of the biharmonic eigenvalue problem. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004052.

Meng:2024:CVE

[MMF24] Jian Meng, Liquan Mei, and Mingfa Fei. H^{1}- conforming virtual element method for the Laplacian eigenvalue problem in mixed form. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003394.

Mahor:2021:ASL

[MMJ21] Teekam Chand Mahor, Rajshree Mishra, and Renu Jain. Analytical solutions of linear fractional partial differential equations using fractional Fourier transform. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304933.

Milovanovic:2022:WNQ

[MMJM22] Gradimir V. Milovanović, Mohammad Masjed-Jamei, and Zahra Moalemi. Weighted nonstandard quadrature formulas based on values of linear differential operators. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000450.

Maier:2020:HTH

[MMM20] Matthias Maier, Dionisios Margetis, and Antoine Mellet. Homogenization of time-harmonic Maxwell’s equations in nonhomogeneous plasmonic structures. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302004.

Mardanov:2022:NCE

[MMM22] Misir J. Mardanov, Telman K. Melikov, and Samin T. Malik. Necessary conditions for the extremum in non-smooth problems of variational calculus. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003459.

Mardanov:2023:NCM

[MMM23] Misir J. Mardanov, Telman K. Melikov, and Samin T. Malik. Necessary conditions for a minimum in classical calculus of variations in the presence of various types of degenerations. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003387.

Mardanov:2020:FSO

[MMMM20] Misir J. Mardanov, Telman K. Melikov, Samin T. Malik, and Kamran Malikov. First- and secondorder necessary conditions with respect to components for discrete optimal control problems. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303395.

Molina-Meyer:2022:PVG

[MMPM22] Marcela Molina-Meyer and Frank Richard Prieto-Medina. Pseudospectral versus Galerkin methods: Fourth order equations. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001546.

Marchi:2020:PIM

[MMPP20] S. De Marchi, F. Marchetti, E. Perracchione, and D. Poggiali. Polynomial interpolation via mapped bases without resampling. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303449.

Makarova:2023:CMC

[MMS^{+}23] Kseniia Makarova, Aleksandr Makarov, Vladislav Strongin, Iuliia Titovets, Yuriy Shevchenko, Vitalii Kapitan, Alexey Rybin, Dmitrii Kapitan, Alena Korol, Egor Vasiliev, Pavel Ovchinnikov, Konstantin Soldatov, Viacheslav Trukhin, and Konstantin Nefedev. Canonical Monte Carlo multispin cluster method. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000973.

Mulder:2020:IRB

[MMV20] Wim De Mulder, Geert Molenberghs, and Geert Verbeke. An interpretation of radial basis function networks as zeromean Gaussian process emulators in cluster space. Journal of Computational and Applied Mathematics, 363(??):249–255, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302845.

Mehrdoust:2020:SMV

[MN20] Farshid Mehrdoust and Ali Reza Najafi. A short memory version of the Vasicek model and evaluating European options on zerocoupon bonds. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030087X.

Mahalik:2021:SIS

[MN21] K. Mahalik and C. Nahak. Solvability of implicit semidefinite and implicit copositive complementarity problems. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303642.

McGregor:2022:PIF

[MN22] Geoffrey McGregor and Jean-Christophe Nave. Parametric interpolation framework for scalar conservation laws. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005148.

Mehrdoust:2021:CDH

[MNH21] Farshid Mehrdoust, Idin Noorani, and Abdelouahed Hamdi. Calibration of the double Heston model and an analytical formula in pricing American put option. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000418.

Mukunoki:2020:PEC

[MO20] Daichi Mukunoki and Takeshi Ogita. Performance and energy consumption of accurate and mixed-precision linear algebra kernels on GPUs. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930706X.

Mammadova:2021:PMC

[MÖ21] Ulduz Mammadova and M. Revan Özkale. Profile monitoring for count data using Poisson and Conway–Maxwell–Poisson regression-based control charts under multicollinearity problem. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305665.

Moir:2022:SSR

[Moi22] T. J. Moir. A study on square root controlsystems. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005549.

Moisa:2024:FTS

[Moi24] Andrew Moisa. A family of two-step second order Runge–Kutta–Chebyshev methods. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001171.

Moon:2022:GMH

[Moo22] Minam Moon. Generalized multiscale hybridizable discontinuous Galerkin (GMsHDG) method for flows in nonlinear porous media. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002084.

Minamihata:2020:MEB

[MORO20] Atsushi Minamihata, Takeshi Ogita, Siegfried M. Rump, and Shin’ichi Oishi. Modified error bounds for approximate solutions of dense linear systems. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305515.

Morkisz:2021:RDF

[MP21] Paweł M. Morkisz and Paweł Przybyłowicz. Randomized derivative-free Milstein algorithm for efficient approximation of solutions of SDEs under noisy information. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304039.

M:2023:VCM

[MP23] Vijayalakshmi G. M. and Roselyn Besi P. Vaccination control measures of an epidemic model with long-term memristive effect. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003740.

Marynets:2024:SAI

[MP24] Kateryna Marynets and Dona Pantova. Successive approximations and interval halving for fractional BVPs with integral boundary conditions. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003059.

Monter-Pozos:2024:TSN

[MPGE24] Aurora Monter-Pozos and Elizabeth González-Estrada. On testing the skew normal distribution by using Shapiro-wilk test. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005939.

Macias:2023:EPG

[MPM23] E. M. Macías, R. Pérez, and H. J. Martínez. An explicit polynomial to globalize algorithms for solving matrix polynomial equations. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004101.

Marriaga:2023:AGB

[MPPR23] Misael E. Marriaga, Teresa E. Pérez, Miguel A. Piñar, and Marlon J. Recarte. Approximation via gradients on the ball. The Zernike case. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002029.

Mainar:2024:TPB

[MPR24] E. Mainar, J. M. Peña, and B. Rubio. On the total positivity of q-Bernstein mass matrices and their accurate computations. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400284X.

Messina:2023:ASN

[MPV23] Eleonora Messina, Mario Pezzella, and Antonia Vecchio. Asymptotic solutions of non-linear implicit Volterra discrete equations. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000122.

Miranville:2020:EDO

[MQ20] Alain Miranville and Ramón Quintanilla. Exponential decay in one-dimensional type II thermoviscoelasticity with voids. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305783.

Ming:2024:NTR

[MQZ^{+}24] Zhenyu Ming, Zhenzhi Qin, Liping Zhang, Yanwei Xu, and Liqun Qi. Network traffic recovery from linkload measurements using tensor triple decomposition strategy for third-order traffic tensors. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001511.

Masetti:2020:CPM

[MR20] G. Masetti and L. Robol. Computing performability measures in Markov chains by means of matrix functions. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305394.

Minakowski:2023:PPE

[MR23] P. Minakowski and T. Richter. A priori and a posteriori error estimates for the Deep Ritz method applied to the Laplace and Stokes problem. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004435.

See erratum [MR25].

Minakowski:2025:EPP

[MR25] P. Minakowski and T. Richter.

Erratum to “A priori and a posteriori error estimates for the Deep Ritz method applied to the Laplace and Stokes problem” [J. Comput. Appl. Math. 421 (2023) 114845]. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400654X.

See [MR23].

Montijano:2023:ERK

[MRC23] J. I. Montijano, L. Rández, and M. Calvo. Explicit Runge–Kutta methods for the numerical solution of linear inhomogeneous IVPs. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000274.

Montijano:2024:ERK

[MRC24] J. I. Montijano, L. Rández, and M. Calvo. Explicit Runge–Kutta–Nyström methods for the numerical solution of second order linear inhomogeneous IVPs. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004776.

Mazzoleni:2021:MMN

[MRG^{+}21] Stefano Mazzoleni, Lucia Russo, Francesco Giannino, Gerardo Toraldo, and Constantinos Siettos. Mathematical modelling and numerical bifurcation analysis of inbreeding and interdisciplinarity dynamics in academia. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304854.

Martinez:2022:OPQ

[MRI22] S. Martínez, M. Rueda, and M. Illescas. The optimization problem of quantile and poverty measures estimation based on calibration. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303459.

Maleknejad:2020:CAH

[MS20a] K. Maleknejad and E. Saeedipoor. Convergence analysis of hybrid functions method for two-dimensional nonlinear Volterra–Fredholm integral equations. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305382.

Maulik:2020:NAP

[MS20b] Romit Maulik and Omer San. Numerical assessments of a parametric implicit large eddy simulation model. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301576.

Mohammed:2020:GFI

[MS20c] Pshtiwan Othman Mohammed and Mehmet Zeki Sarikaya. On generalized fractional integral inequalities for twice differentiable convex functions. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300315.

Manohar:2022:LPE

[MS22] Ram Manohar and Rajen Kumar Sinha. Local a posteriori error estimates for boundary control problems governed by nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000371.

Milosevic:2024:EFS

[MS24a] Bojana Milosević and Jelena Stanojević. On the estimation of fuzzy stressstrength reliability parameter. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004806.

Mustafa:2024:MRE

[MS24b] Ashif Mustafa and Manideepa Saha. Maximal residual extended Kaczmarz and Gauss–Seidel methodsconvergence properties and applications. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001705.

Martinez-Sanchez:2023:MMV

[MSBSM23] Julio César Martínez-Sánchez, Arturo Berrones-Santos, and Javier Almaguer Martínez. The Markowitz’s Mean-Variance Interpretation under the efficient market hypothesis in the context of critical recession periods. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001711.

Manimaran:2021:FEE

[MSD21] J. Manimaran, L. Shangerganesh, and Amar Debbouche. Finite element error analysis of a timefractional nonlocal diffusion equation with the Dirichlet energy. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303575.

Mert:2021:TDS

[MSK21] Ozenc Murat Mert and A. Sevtap Selcuk-Kestel. Time dependent stop-loss reinsurance and exposure curves. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306397.

Milovanovic:2024:GTQ

[MSM24] Gradimir V. Milovanović, Marija P. Stanić, and Tatjana V. Tomović Mladenović. Gaussian type quadrature rules related to the oscillatory modification of the generalized Laguerre weight functions. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300420X.

Messaoudi:2020:MRP

[MSS20] A. Messaoudi, R. Sadaka, and H. Sadok. Matrix recursive polynomial interpolation algorithm: an algorithm for computing the interpolation polynomials. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304741.

Mosic:2021:OSW

[MSS^{+}21] Dijana Mosić, Predrag S.

Stanimirović, Jajati Keshari Sahoo, Ratikanta Behera, and Vasilios N. Katsikis. One-sided weighted outer inverses of tensors. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305847.

Malhotra:2022:PGA

[MST22] Gifty Malhotra, R. Srivastava, and H. C. Taneja. Pricing of the geometric Asian options under a multifactor stochastic volatility model. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005823.

Makmuang:2023:RSN

[MSW23] Dawrawee Makmuang, Siwakon Suppalap, and Rabian Wangkeeree. The regularized stochastic Nesterov’s accelerated quasi-Newton method with applications. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001346.

Matsue:2020:RNB

[MT20a] Kaname Matsue and Akitoshi Takayasu. Rigorous numerics of blowup solutions for ODEs with exponential nonlinearity. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306120.

Miao:2020:AAM

[MT20b] Cun-Qiang Miao and Xue-Yuan Tan. Accelerating the Arnoldi method via Chebyshev polynomials for computing PageRank. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301825.

Medina:2022:SHD

[MTIR22] Emmanuel Y. Medina, Elson M. Toledo, Iury Igreja, and Bernardo M. Rocha. A stabilized hybrid discontinuous Galerkin method for the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006038.

Mahmoudi:2021:KTS

[MTP21] Mohammad Reza Mahmoudi, Bui Anh Tuan, and Kim-Hung Pho. On kurtoses of two symmetric or asymmetric populations. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306610.

Ma:2023:HOU

[MTZ23] Chuwen Ma, Tian Tian, and Weiying Zheng. Highorder unfitted characteristic finite element methods for moving interface problem of Oseen equations. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006264.

Mu:2021:WGF

[Mu21] Lin Mu. Weak Galerkin finite element with curved edges. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303290.

Makushenko:2023:OTS

[MULS23] Igor Makushenko, Iryna Usar, Hanna Livinska, and Michael Sharapov. Optimal threshold strategies for retrial systems with queue. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000808.

Matysik:2022:ASS

[MV22] Oleg Matysik and Marc M.

Van Hulle. Alternating step size method for solving ill-posed linear operator equations in energetic space. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002722.

Magirou:2020:PIA

[MVB20] Evangelos F. Magirou, Paraskevas Vassalos, and Nikolaos Barakitis. A policy iteration algorithm for the American put option and free boundary control problems. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305497.

Munoz-Vazquez:2021:SAC

[MVFAMF21] Aldo Jonathan Muñoz-Vázquez, Guillermo Fernández-Anaya, and Oscar Martínez-Fuentes. Stability analysis of a class of integral equations with not necessarily differentiable solutions. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003241.

Mallik:2019:GOP

[MVY19] Gouranga Mallik, Martin Vohralík, and Soleiman Yousef. Goal-oriented a posteriori error estimation for conforming and nonconforming approximations with inexact solvers. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930370X.

Ma:2020:CRM

[MW20] Jingtang Ma and Han Wang. Convergence rates of moving mesh methods for moving boundary partial integro-differential equations from regimeswitching jump-diffusion Asian option pricing. Journal of Computational and Applied Mathematics, 370 (??):??, May 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930603X.

Miao:2021:RFK

[MW21] Cun-Qiang Miao and Wen-Ting Wu. On relaxed filtered Krylov subspace method for nonsymmetric eigenvalue problems. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003204.

Miao:2022:GRA

[MW22] Cun-Qiang Miao and Wen-Ting Wu. On greedy randomized average block Kaczmarz method for solving large linear systems. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001662.

Meng:2022:LOF

[MWBM22] Jian Meng, Xue Wang, Linlin Bu, and Liquan Mei. A lowest-order freestabilization Virtual Element Method for the Laplacian eigenvalue problem. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005975.

Ma:2020:MPA

[MWHC20] Yanzhun Ma, Wei Wei, Yongjian Hu, and Gongning Chen. The multipoint Padé approximation problem and its Hankel vector. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300984.

Ma:2024:ABF

[MWLY24] Liyuan Ma, Hongxia Wang, Ningyi Leng, and Ziyang Yuan. ADMM based Fourier phase retrieval with untrained generative prior. Journal of Computational and Applied Mathematics, 444(??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000359.

Mao:2021:PPT

[MWW21] Xuerong Mao, Fengying Wei, and Teerapot Wiriyakraikul. Positivity preserving truncated Euler–Maruyama method for stochastic Lotka–Volterra competition model. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001862.

Mi:2020:OID

[MX20] Hui Mi and Lixia Xu.

Optimal investment with derivatives and pricing in an incomplete market. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305278.

Mahmudov:2021:NBV

[MY21] Elimhan N. Mahmudov and Shakir Sh. Yusubov. Nonlocal boundary value problems for hyperbolic equations with a Caputo fractional derivative. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003319.

Ma:2022:CAC

[MYC22] Jingtang Ma, Wensheng Yang, and Zhenyu Cui. Convergence analysis for continuous-time Markov chain approximation of stochastic local volatility models: Option pricing and Greeks. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005240.

Mai:2024:PCC

[MYLL24] Ruifeng Mai, Zhou Yang, Yingyi Lai, and Jianwei Lin. Portfolioconsumption choice problem with voluntary retirement and consumption constraints. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000888.

Migorski:2023:CEQ

[MYZ23] Stanisław Migórski, Jen-Chih Yao, and Shengda Zeng. A class of elliptic quasi-variational-hemivariational inequalities with applications. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004691.

Min:2023:FDS

[MZY23] Xilin Min, Jun Zhang, and Xiaofeng Yang. Fullydiscrete spectral-Galerkin numerical scheme with second-order time accuracy and unconditional energy stability for the anisotropic Cahn–Hilliard model. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002965.

Nikan:2021:CCN

[NA21] O. Nikan and Z. Avazzadeh. Coupling of the Crank–Nicolson scheme and localized meshless technique for viscoelastic wave model in fluid flow. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003174.

Noeiaghdam:2020:VIS

[NAA20] Samad Noeiaghdam, Mohammad Ali Fariborzi Araghi, and Saeid Abbasbandy. Valid implementation of Sinc-collocation method to solve the fuzzy Fredholm integral equation. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306375.

Najafi:2020:NQM

[Naj20] Esmaeil Najafi. Nyströmquasilinearization method and smoothing transformation for the numerical solution of nonlinear weakly singular Fredholm integral equations. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305436.

Noeiaghdam:2022:DSH

[NAS22] Samad Noeiaghdam, Mohammad Ali Fariborzi Araghi, and Denis Sidorov. Dynamical strategy on homotopy perturbation method for solving second kind integral equations using the CESTAC method. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000851.

Nwaigwe:2023:GBF

[NB23] Chinedu Nwaigwe and Deborah Ngochinma Benedict. Generalized Banach fixed-point theorem and numerical discretization for nonlinear Volterra– Fredholm equations. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006173.

Nour:2024:MSS

[NBLZ24] M.-Y. Nour, D. Barrera, A. Lamnii, and A. Zidna. MTH subdivision scheme with sharp and semi-sharp features. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006684.

Nedzhibov:2022:WIM

[Ned22] Gyurhan H. Nedzhibov. The Weierstrass iterative method as a Petrov– Galerkin method for solving eigenvalue problem. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721005665.

Newer:2024:PFO

[New24] Haidy A. Newer. Prediction of future observations based on ordered extreme k-records ranked set sampling with unequal fixed and random sample sizes. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000475.

Naik:2022:CDD

[NEYZ22] Parvaiz Ahmad Naik, Zohreh Eskandari, Mehmet Yavuz, and Jian Zu. Complex dynamics of a discrete-time Bazykin– Berezovskaya prey–predator model with a strong Allee effect. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001844.

Newcome:2023:TAT

[NGNB23] Samuel James Newcome, Fabio Alexander Gratl, Philipp Neumann, and Hans-Joachim Bungartz. Towards auto-tuning Multi-Site Molecular Dynamics simulations with AutoPas. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002224.

Ngondiep:2022:RNT

[Ngo22a] Eric Ngondiep. A robust numerical two-level second-order explicit approach to predicting the spread of covid-2019 pandemic with undetected infectious cases. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004751.

Ngondiep:2022:USL

[Ngo22b]  Eric Ngondiep.  Un-

conditional stability over long time intervals of a two-level coupled Mac-Cormack/Crank–Nicolson method for evolutionary mixed Stokes–Darcy model. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000383.

Ngondiep:2024:PEE

[Ngo24] Eric Ngondiep. A posteriori error estimate of MacCormack rapid solver method for two-dimensional incompressible Navier– Stokes problems. Journal of Computational and Applied Mathematics, 438 (??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042723005137.

Nadarajah:2022:HTM

[NK22] Saralees Nadarajah and Charles Kwofie. Heavy tailed modeling of automobile claim data from Ghana. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005598.

Natido:2023:ULM

[NK23] Amos Natido and Tomasz J.

Kozubowski. A uniform-Laplace mixture distribution. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001802.

Nikazad:2024:CRP

[NK24] T. Nikazad and M. Khakzad.

Choosing relaxation parameter in randomized Kaczmarz method. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000396.

Najafi-Kalyani:2020:GIS

[NKBJ20] Mehdi Najafi-Kalyani, Fatemeh Panjeh Ali Beik, and Khalide Jbilou. On global iterative schemes based on Hessenberg process for (ill-posed) Sylvester tensor equations. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719301761.

Nakano:2023:GLE

[NL23a] Taiga Nakano and Xuefeng Liu. Guaranteed local error estimation for finite element solutions of boundary value problems. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000055.

Niu:2023:OSA

[NL23b] Chengmei Niu and Hanyu Li. Optimal sampling algorithms for block matrix multiplication. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000079.

Niu:2024:OSL

[NM24] Cuixia Niu and Heping Ma. An operator splitting Legendre-tau spectral method for Maxwell’s equations with nonlinear conductivity in two dimensions. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004430.

Navarro:2022:ADT

[NMB22] Juan F. Navarro and M. C.

Martínez-Belda. Analysis of the distribution of times of escape in the Nbody ring problem. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000157.

Nunes:2022:ASM

[NMF^{+}22] Célia Nunes, Anacleto Mário, Dário Ferreira, Elsa M. Moreira, Sandra S. Ferreira, and João T. Mexia. An algorithm for simulation in mixed models with crossed factors considering the sample sizes as random. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000820.

Nhan:2023:NUD

[NMMH23] Thái Anh Nhan, Vinh Quang Mai, Jugal Mohapatra, and Zakia Hammouch. A new upwind difference analysis of an exponentially graded bakhvalov-type mesh for singularly perturbed elliptic convection–diffusion problems. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003119.

NGuessan:2021:HOT

[NMST21] Marc-Arthur N’Guessan, Marc Massot, Laurent Séries, and Christian Tenaud. High order time integration and mesh adaptation with error control for incompressible Navier– Stokes and scalar transport resolution on dual grids. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305473.

Nooghabi:2022:SSR

[NN22] Mehdi Jabbari Nooghabi and Mehrdad Naderi. Stress-strength reliability inference for the Pareto distribution with outliers. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005343.

Naderi:2024:CAD

[NN24] Mehrdad Naderi and Mehdi Jabbari Nooghabi. Clustering asymmetrical data with outliers: Parsimonious mixtures of contaminated mean-mixture of normal distributions. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003771.

Nachaoui:2021:SNN

[NNCH21] A. Nachaoui, M. Nachaoui, A. Chakib, and M. A. Hilal. Some novel numerical techniques for an inverse Cauchy problem. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303216.

Naranjo-Noda:2024:JFL

[NNJ24] F. S. Naranjo-Noda and J. C. Jimenez. Jacobianfree locally linearized Runge– Kutta method of Dormand and Prince for large systems of differential equations. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002243.

Noorizadegan:2024:BAQ

[NNL^{+}24] Amir Noorizadegan, Ahmed Naji, Tsung-Lin Lee, Roberto Cavoretto, and D. L. Young. Bending analysis of quasicrystal plates using adaptive radial basis function method. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002401.

Nakao:2021:F

[NOR21] Mitsuhiro T. Nakao, Takeshi Ogita, and Siegfried M.

Rump. Foreword. Journal of Computational and Applied Mathematics, 382 (??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720303691.

Novakovic:2024:ACJ

[Nov24] Vedran Novaković. Accurate complex Jacobi rotations. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400253X.

Nandori:2022:DTI

[NP22] Péter Nándori and Dan Pirjol. On the distribution of the time-integral of the geometric Brownian motion. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004404.

Novruzi:2023:ASG

[NP23] Arian Novruzi and Bartosz Protas. An accelerated Sobolev gradient method for unconstrained optimization problems based on variable inner products. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004319.

Nichol:2021:MLF

[NPP^{+}21] J. Jake Nichol, Matthew G. Peterson, Kara J. Peterson, G. Matthew Fricke, and Melanie E. Moses. Machine learning feature analysis illuminates disparity between E3SM climate models and observed climate change. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000704.

Nino-Ruiz:2021:EKF

[NRGJ21] Elias D. Nino-Ruiz, Luis Guzman, and Daladier Jabba. An ensemble Kalman filter implementation based on the Ledoit and Wolf covariance matrix estimator. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304544.

Namba:2021:SCU

[NRV21] T. Namba, P. Rybka, and V. R. Voller. Some comments on using fractional derivative operators in modeling non-local diffusion processes. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303319.

Nunokawa:2020:SCS

[NS20] Mamoru Nunokawa and Janusz Sokół. On some conditions for schlichtness of analytic functions. Journal of Computational and Applied Mathematics, 363(??):241–248, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302821.

Nan:2024:NHO

[NS24a] Caixia Nan and Huailing Song. A new highorder maximum-principle-preserving explicit Runge– Kutta method for the nonlocal Allen–Cahn equation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004442.

Navarro:2024:NLB

[NS24b] Jorge Navarro and José María Sarabia. A note on the limiting behaviour of hazard rate functions of generalized mixtures. Journal of Computational and Applied Mathematics, 435 (??):??, ???? 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003302.

Nikiforov:2024:MMMb

[NS24c] Djulustan Nikiforov and Sergei Stepanov. Meshfree multiscale method with partially explicit time discretization for nonlinear Stefan problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400270X.

Nie:2022:NMF

[NSL22] Cunyun Nie, Shi Shu, and Menghuan Liu. A novel monotone finite volume element scheme for diffusion equations. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002163.

Nikiforov:2024:MMMa

[NSL24] Djulustan Nikiforov, Sergei Stepanov, and Nyurgun Lazarev. Meshfree multiscale method for the infiltration problem in permafrost. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002383.

Noutsos:2020:BTP

[NT20] D. Noutsos and G. Tachyridis.

Band Toeplitz preconditioners for non-symmetric real Toeplitz systems by preconditioned GMRES method. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719302274.

Namm:2023:SSC

[NT23] R. V. Namm and G. I. Tsoy. Solution of the static contact problem with Coulomb friction between an elastic body and a rigid foundation. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003661.

Nudo:2024:GQE

[Nud24] Federico Nudo. A general quadratic enrichment of the Crouzeix–Raviart finite element. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003613.

Nasser:2020:NCC

[NV20] Mohamed M. S. Nasser and Matti Vuorinen. Numerical computation of the capacity of generalized condensers. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301564.

Nabubie:2023:NTD

[NW23] Bashiruddin Nabubie and Song Wang. Numerical techniques for determining implied volatility in option pricing. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005118.

Ni:2020:MBF

[NWD20] Qian Ni, Xuhui Wang, and Jiansong Deng. Modified basis functions for MPHT-splines. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301084.

Nikitin:2020:MED

[NY20] Kirill D. Nikitin and Ruslan M. Yanbarisov. Monotone embedded discrete fractures method for flows in porous media. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303565.

Ng:2024:DLB

[NY24a] Michael Ng and He Ming Yao. Deep learning based source reconstruction method using asymmetric encoder–decoder structure and physicsinduced loss. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004478.

Niu:2024:PFM

[NY24b] Xiaohua Niu and Xiaodong Yan. A phase field model for the motion of prismatic dislocation loops by both climb and self-climb. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005162.

Ng:2023:SSF

[NYL23] Yu-Keung Ng, Guoqiao You, and Shingyu Leung. Sparse subsampling of flow measurements for finitetime Lyapunov exponent in domains with obstacles. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001991.

Novak:2021:RFC

[NZ21] Tina Novak and Janez Zerovnik. Real forms of the complex Neumann system: a method for finding real roots of polynomial U_{S}(λ). Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306531.

Nasiri:2024:NSQ

[NZA24] T. Nasiri, A. Zakeri, and A. Aminataei. A numerical solution for a quasi solution of the time-fractional stochastic backward parabolic equation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003850.

Ogita:2020:IRS

[OA20] Takeshi Ogita and Kensuke Aishima. Iterative refinement for singular value decomposition based on matrix multiplication. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305175.

Orlando:2023:IDS

[OBB23] Giuseppe Orlando, Tommaso Benacchio, and Luca Bonaventura. An IMEX-DG solver for atmospheric dynamics simulations with adaptive mesh refinement. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000687.

Oliari:2024:PEE

[OBdSD24] Victor B. Oliari, Paulo Rafael Bösing, Denise de Siqueira, and Philippe R. B. Devloo. A posteriori error estimates for primal hybrid finite element methods applied to Poisson problem. Journal of Computational and Applied Mathematics, 441 (??):??, May 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006155.

Ober-Blobaum:2023:VLE

[OBO23] Sina Ober-Blöbaum and Christian Offen. Variational learning of Euler– Lagrange dynamics from data. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003971.

Ozarslan:2023:CAP

[ÖÇ23] Mehmet Ali Özarslan and Bayram Çekim. Confluent Appell polynomials. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005829.

Ouaissa:2024:WPS

[OCS24] H. Ouaissa, A. Chakib, and A. Sadik. On the well posedness and Steffensen’s based numerical approximation of an inverse Cauchy problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002905.

Ouakka:2023:GDB

[OEH23] Abdellah Ouakka, Abdelhai El Azzouzi, and Zakia Hammouch. Global dynamic behavior of a vaccination-age SVIR model with treatment and general nonlinear incidence rate. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004460.

Okuno:2020:IPS

[OF20] Takayuki Okuno and Masao Fukushima. An interior point sequential quadratic programmingtype method for logdeterminant semi-infinite programs. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300753.

Oseledets:2022:DOB

[OF22] Ivan Oseledets and Vladimir Fanaskov. Direct optimization of BPX preconditioners. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004337.

Ogata:2020:NCF

[Oga20] Hidenori Ogata. Numerical calculation of Fourier transforms based on hyperfunction theory. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302120.

Orti:2022:AMC

[OGR22] José A. López Ortí, Manuel Forner Gumbau, and Miguel Barreda Rochera. An alternative method to construct a consistent second-order theory on the equilibrium figures of rotating celestial bodies. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305963.

Opsomer:2023:HOA

[OH23] Peter Opsomer and Daan Huybrechs. High-order asymptotic expansions of Gaussian quadrature rules with classical and generalized weight functions. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002613.

Ojeda-Hernandez:2022:QCE

[OHCC22] Manuel Ojeda-Hernández, Inma P. Cabrera, and Pablo Cordero. Quasiclosed elements in fuzzy posets. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000091.

Ouni:2023:TPN

[OHK23] Marwa Ouni, Abderrahmane Habbal, and Moez Kallel. A three-player Nash game for pointwise source identification in Cauchy-Stokes problems. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003053.

Ozkose:2023:NFO

[ÖHS23] Fatma Özköse, Rafla Habbireeh, and M. Tamer Senel. A novel fractional order model of SARS-CoV-2 and cholera disease with real data. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005672.

Oishi:2020:NIE

[Ois20] Shin’ichi Oishi. Numerical inclusion of exact periodic solutions for time delay Duffing equation. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306259.

Ohemeng:2023:TTC

[OK23] Matthew A. Ohemeng and Tomasz J. Kozubowski. A triptych on three continuous analogs of the Poisson, binomial, and negative binomial distributions. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002194.

Ozdemir:2024:ARP

[OKE24] Irem Bulanik Ozdemir, Sevval Kilicoglu, and Serkan Eryilmaz. Age replacement policy for heterogeneous parallel systems. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300506X.

Ozaltun:2023:GWS

[ÖKG23] Gökçe Özaltun, Ali Konuralp, and Sevin Gümgüm. Gegenbauer wavelet solutions of fractional integrodifferential equations. Journal of Computational and Applied Mathematics, 420 (??):??, March 1, 2023.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722004289.

Ojo:2023:CBI

[OKW23] Saheed O. Ojo, Hasan M.

Khalid, and Paul M. Weaver. Convergence behaviour of inverse differential quadrature method for analysis of beam and plate structures. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002856.

Oviedo:2024:BSB

[OMRdQ24] P. C. R. Oviedo, M. T.

Magalhães, B. M. Rocha, and R. A. B. de Queiroz. A bounded scheme based on Bézier curves for convection-dominated transport problems. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004466.

Okor:2022:HOE

[ON22] T. Okor and G. C.

Nwachukwu. High order extended boundary value methods for the solution of stiff systems of ODEs. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003721.

Okayama:2021:NCM

[ONT21] Tomoaki Okayama, Tomoki Nomura, and Saki Tsuruta. New conformal map for the trapezoidal formula for infinite integrals of unilateral rapidly decreasing functions. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306452.

Okundalaye:2022:TEA

[OOO22] O. O. Okundalaye, W. A. M. Othman, and A. S. Oke. Toward an efficient approximate analytical solution for 4-compartment COVID-19 fractional mathematical model. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200245X.

Orosz:2023:PAR

[OPK23] Tamás Orosz, David Pánek, and Miklós Kuczmann. Performance analysis of a robust design optimization of a solenoid with different sensitivity metrics. Journal of Computational and Applied Mathematics, 424 (??):??, May 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006197.

Opheim:2021:ICM

[OR21] Timothy Opheim and Anuradha Roy. Inverse of the covariance matrix of an MA(2) process. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002491.

Onisk:2023:NCB

[ORS23] Lucas Onisk, Lothar Reichel, and Hassane Sadok. Numerical considerations of block GMRES methods when applied to linear discrete ill-posed problems. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002066.

Occorsio:2022:FIR

[ORT22] Donatella Occorsio, Maria Grazia Russo, and Woula Themistoclakis. Filtered integration rules for finite weighted Hilbert transforms. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000474.

Ozen:2020:TMJ

[ÖS20] Selin Özen and Sule Sahin. Transitory mortality jump modeling with renewal process and its impact on pricing of catastrophic bonds. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301205.

Okayama:2020:NCM

[OSK20] Tomoaki Okayama, Yuya Shintaku, and Eisuke Katsuura. New conformal map for the Sinc approximation for exponentially decaying functions over the semiinfinite interval. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303619.

Orlando:2021:ATA

[OT21] Giuseppe Orlando and Giovanni Taglialatela. On the approximation of the Black and Scholes call function. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304453.

Olmos:2020:CHS

[OVGI20] Neveka M. Olmos, Osvaldo Venegas, Yolanda M. Gómez, and Yuri A. Iriarte. Confluent hypergeometric slashed-Rayleigh distribution: Properties, estimation and applications. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305539.

Ou:2024:SOF

[OWV24] Caixia Ou, Zhibo Wang, and Seakweng Vong. A second-order fitted scheme combined with time twogrid technique for twodimensional nonlinear time fractional telegraph equations involving initial singularity. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001869.

Ozdemir:2023:F

[ÖYT23] Necati Özdemi̇r, Ramazan Yaman, and Fatih Tank. Foreword. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001826.

Ozkut:2021:CSH

[Ozk21] Murat Ozkut. The (n − k + 1)-out-of-n concomitant system having m subcomponents and its reliability. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305422.

Pekalp:2021:PSE

[PA21] Mustafa Hilmi Pekalp and Halil Aydogdu. Power series expansions for the probability distribution, mean value and variance functions of a geometric process with gamma interarrival times. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305781.

P:2024:OCA

[PA24a] Priya P. and Sabarmathi A. Optimal control on ABC fractal fractional order model of micro-plastic pollution in soil and its effect on the nutrient cycle. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002474.

Pekalp:2024:PEM

[PA24b] Mustafa Hilmi Pekalp and Halil Aydogdu. Parametric estimations of the mean value and variance functions in geometric process. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272400219X.

Partohaghighi:2024:RNS

[PAAI24] Mohammad Partohaghighi, Emmanuel Asante-Asamani, and Olaniyi S. Iyiola. A robust numerical scheme for solving Riesz-tempered fractional reaction–diffusion equations. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002425.

Park:2020:SPS

[Par20] Chunjae Park. Spurious pressure in Scott–Vogelius elements. Journal of Computational and Applied Mathematics, 363(??):370– 391, January 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0377042719302791.

Patra:2023:EMI

[Pat23]  Asim Patra.  An epi-

demiology model involving high-order linear Fredholm integro-differential-difference equations via a novel balancing collocation technique. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004496.

Popovic:2021:NLR

[PBR21] Predrag M. Popović, Hassan S. Bakouch, and Miroslav M. Ristić. A nonlinear random environment INAR(1) model. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000273.

Peeyada:2024:NPA

[PC24] Pronpat Peeyada and Watcharaporn Cholamjiak. A new projection algorithm for variational inclusion problems and its application to cervical cancer disease prediction. Journal of Computational and Applied Mathematics, 441 (??):??, May 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006453.

Park:2021:MSM

[PCM21] Jun Sur Richard Park, Siu Wun Cheung, and Tina Mai. Multiscale simulations for multicontinuum Richards equations. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002703.

Park:2020:MSU

[PCMH20] Jun Sur Richard Park, Siu Wun Cheung, Tina Mai, and Viet Ha Hoang. Multiscale simulations for upscaled multi-continuum flows. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030073X.

Palczynski:2023:FGD

[PCT23] Krzystof Pałczyński, Magda Czyzewska, and Tomasz Talaśka. Fuzzy Gaussian decision tree. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006367.

Peng:2020:SEG

[PCWY20] Yufang Peng, Weidong Chen, Pengbang Wei, and Guanyi Yu. Spillover effect and Granger causality investigation between China’s stock market and international oil market: a dynamic multiscale approach. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304637.

Perez-Diaz:2024:EAC

[PDBdS24] Sonia Pérez-Díaz, Rafael Magdalena Benedicto, and Marian Fernández de Sevilla. An effective algorithm for computing the asymptotes of an implicit curve. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004120.

Petridi:2020:MSD

[Pet20] Constantin M. Petridi. Mathematical structures defined by identities. Journal of Computational and Applied Mathematics, 373 (??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042719305527.

Pries:2024:ISS

[PFG24] Michael Pries, Andreas Fiolitakis, and Peter Gerlinger. An implicit splitting scheme with characteristic boundary conditions for compressible reactive flows on unstructured grids. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003904.

Papadimitriou:2020:ECM

[PGG20] Theophilos Papadimitriou, Periklis Gogas, and Fotios Gkatzoglou. The evolution of the cryptocurrencies market: a complex networks approach. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301229.

Park:2020:HMF

[PH20a] Jun Sur Richard Park and Viet Ha Hoang. Hierarchical multiscale finite element method for multicontinuum media. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930593X.

Petkovic:2020:CMC

[PH20b] I. Petković and D. Herceg.

Computer methodologies for comparison of computational efficiency of simultaneous methods for finding polynomial zeros. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305187.

Presho:2021:CGM

[PH21] Michael Presho and Michael Hill. A conservative generalized multiscale finite volume/element method for modeling two-phase flow with capillary pressure. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303174.

Pena:2022:INE

[PH22] Larry Peña and Jorge Hinojosa. Implementation of a new expression for the search direction in simulations of structures with buckling and post-buckling: “Two-Scale Impedance”. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004210.

Pakhira:2024:NFP

[PH24] Samik Pakhira and Sk Monowar Hossein. A new fixed point theorem in G_{b}-metric space and its application to solve a class of nonlinear matrix equations. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004181.

Parker:2022:PTM

[PHDD22] J. T. Parker, P. A.

Hill, D. Dickinson, and B. D. Dudson. Parallel tridiagonal matrix inversion with a hybrid multigrid-Thomas algorithm method. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003289.

Pho:2022:INR

[Pho22] Kim-Hung Pho. Improvements of the Newton– Raphson method. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000139.

Peszynska:2024:CFE

[PHV24] M. Peszynska, Z. Hilliard, and N. Vohra. Coupled flow and energy models with phase change in permafrost from pore- to Darcy scale: Modeling and approximation. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002140.

Proinov:2024:NFS

[PI24] Petko D. Proinov and Stoil I. Ivanov. A new family of Sakurai–Torii– Sugiura type iterative methods with high order of convergence. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003722.

Pietkun:2020:ESC

[Pie20] Radosław Pietkun. Existence of solutions for a class of multivalued functional integral equations of Volterra type via the measure of nonequicontinuity on the Fréchet space C(Ω, E). Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302612.

Pang:2024:CAD

[PJA24] Gang Pang, Songsong Ji, and Xavier Antoine. Construction and analysis of discretization schemes for one-dimensional nonlocal Schrödinger equations with exact absorbing boundary conditions. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005678.

Pan:2021:EMA

[PJM21] Maodong Pan, Bert Jüttler, and Angelos Mantzaflaris. Efficient matrix assembly in isogeometric analysis with hierarchical B-splines. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305690.

Petrasova:2023:SEP

[PK23] Iveta Petrásová and Pavel Karban. Solving evolutionary problems using recurrent neural networks. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000353.

Park:2024:NOD

[PK24] Eun-Hee Park and Hyea Hyun Kim. Non-overlapping domain decomposition algorithms with only primal velocity unknowns for the discontinuous viscosity Stokes problem. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005848.

Parvizi:2020:ACR

[PKE20] Maryam Parvizi, Amirreza Khodadadian, and M. R. Eslahchi. Analysis of Ciarlet–Raviart mixed finite element methods for solving damped Boussinesq equation. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301096.

Petrasova:2022:OSO

[PKK^{+}22] Iveta Petrasova, Pavel Karban, Petr Kropik, David Panek, and Ivo Dolezel. Optimization of selected operation characteristics of array antennas. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003484.

Phairatchatniyom:2024:NIM

[PKMM24] Pawicha Phairatchatniyom, Poom Kumam, and Juan Martinez-Moreno. New inertial modification of regularized algorithms for solving split variational inclusion problem. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004326.

Petrasova:2021:MDM

[PKPD21] Iveta Petrasova, Vaclav Kotlan, David Panek, and Ivo Dolezel. Methodology of determining material parameters based on optimization techniques. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002934.

Polekhina:2023:NSM

[PKS23] R. R. Polekhina, B. A. Korneev, and E. B. Savenkov. Numerical study of multiphase hyperbolic models. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005234.

Perera:2020:CRS

[PL20] Sirani M. Perera and Jianhua Liu. Complexity reduction, self/completely recursive, radix-2 DCT I/ IV algorithms. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302272.

Pan:2021:PWG

[PL21] Jiajia Pan and Huiyuan Li. A penalized weak Galerkin spectral element method for second order elliptic equations. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305197.

Pi:2024:PSC

[PL24a] ZhiPeng Pi and Xin Lai. Polynomial solution of Cauchy-type singular integro-differential equations with bivariate kernels. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002887.

Pooladi:2024:SIU

[PL24b] Fatemeh Pooladi and Elisabeth Larsson. Stabilized interpolation using radial basis functions augmented with selected radial polynomials. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004260.

Preuss:2022:WHS

[PLS22] Adam Preuss, Jessica Lipoth, and Raymond J. Spiteri. When and how to split? A comparison of two IMEX splitting techniques for solving advection– diffusion–reaction equations. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001960.

Pasricha:2021:NCD

[PLZ21] Puneet Pasricha, Xiaoping Lu, and Song-Ping Zhu. A note on the calculation of default probabilities in “Structural credit risk modeling with Hawkes jump-diffusion processes”. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720303289.

Panahi:2020:EIE

[PM20a] Hanieh Panahi and Nasrin Moradi. Estimation of the inverted exponentiated Rayleigh distribution based on adaptive type II progressive hybrid censored sample. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303425.

Patel:2020:FOC

[PM20b] Kuldip Singh Patel and Mani Mehra. Fourth order compact scheme for space fractional advection– diffusion reaction equations with variable coefficients. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302545.

Patel:2022:TSA

[PMVB22] J. K. Patel, L. R. C.

Moraes, R. Vasques, and R. C. Barros. Transport synthetic acceleration for the solution of the onespeed nonclassical spectral S _{N} equations in slab geometry. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003903.

Phan:2021:AII

[PN21a] Duy Nhat Phan and Thuy Ngoc Nguyen. An accelerated IRNN-Iteratively Reweighted Nuclear Norm algorithm for nonconvex nonsmooth low-rank minimization problems. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002223.

Pulch:2021:SARa

[PN21b] Roland Pulch and Akil Narayan. Sensitivity analysis of random linear dynamical systems using quadratic outputs. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304947.

Pulch:2021:SARb

[PNS21] Roland Pulch, Akil Narayan, and Tatjana Stykel. Sensitivity analysis of random linear differential–algebraic equations using system norms. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002880.

Palta:2021:NSB

[PO21] Birce Palta and Hae-Soo Oh. Numerical solutions of biharmonic equations on non-convex polygonal domains. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303137.

Poudyal:2021:TCA

[Pou21] Chudamani Poudyal. Truncated, censored, and actuarial payment-type moments for robust fitting of a single-parameter Pareto distribution. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306014.

Pousin:2024:SES

[Pou24] Jerome Pousin. Some elliptic second order problems and neural network solutions: Existence and error estimates. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003424.

Polunchenko:2021:EII

[PP21] Aleksey S. Polunchenko and Andrey Pepelyshev. On the evaluation of an integral involving the Whittaker W function. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304027.

Padron:2022:SCG

[PP22] Miguel A. Padrón and Ángel Plaza. Similarity classes generated by the 8T-LE partition applied to trirectangular tetrahedra. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000395.

Peet:2024:NTB

[PP24] Yulia T. Peet and Matthew M. Peet. A new treatment of boundary conditions in PDE solution with Galerkin methods via Partial Integral Equation framework. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006179.

Petrovic:2024:SAG

[PPSM24] Nevena Z. Petrović, Miroslav S.

Pranić, Marija P. Stanić, and Tatjana V. Tomović Mladenović. The set of anti-Gaussian quadrature rules for the optimal set of quadrature rules in Borges’ sense. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006751.

Picklo:2024:LMN

[PR24] Matthew J. Picklo and Jennifer K. Ryan. LSIAC-MRA for nonuniform meshes and applications to mesh adaptivity. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003406.

Prodanov:2023:AAS

[Pro23] Dimiter Prodanov. Asymptotic analysis of the SIR model and the Gompertz distribution. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200499X.

Parkinson:2022:ANF

[PRW^{+}22] Suzanna Parkinson, Hayden Ringer, Kate Wall, Erik Parkinson, Lukas Erekson, Daniel Christensen, and Tyler J. Jarvis. Analysis of normal-form algorithms for solving systems of polynomial equations. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000899.

Peszynska:2021:AHF

[PS21] Małgorzata Peszynska and Ralph E. Showalter. Approximation of hysteresis functional. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306476.

Papadopoulos:2022:NAT

[PS22a] I. P. A. Papadopoulos and E. Süli. Numerical analysis of a topology optimization problem for Stokes flow. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200125X.

Potts:2022:LMF

[PS22b] D. Potts and M. Schmischke. Learning multivariate functions with low-dimensional structures using polynomial bases. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100443X.

Pandiya:2024:IPB

[PS24a] Ridwan Pandiya and Salmah.

Inflection point-based auxiliary function algorithm for finding global minima of coercive functions. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400205X.

Prerna:2024:OQP

[PS24b] Prerna and Vikas Sharma. Optimization of a quadratic programming problem over an integer efficient set. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723005952.

Pulch:2024:RMS

[PS24c] Roland Pulch and Adrian Sandu. Rosenbrock-W methods for stochastic Galerkin systems. Journal of Computational and Applied Mathematics, 438 (??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723004715.

Psarrakos:2024:PMV

[Psa24] Georgios Psarrakos. On probabilistic mean value theorem and covariance identities. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002048.

Podila:2024:WBA

[PSRVA24] Pramod Chakravarthy Podila, Vishwas Sundrani, Higinio Ramos, and Jesús Vigo-Aguiar. Waveletbased approximation for two-dimensional singularly perturbed elliptic problems. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003194.

Przybylowicz:2024:RMA

[PSS24] Paweł Przybyłowicz, Verena Schwarz, and Michaela Szölgyenyi. Randomized Milstein algorithm for approximation of solutions of jump-diffusion SDEs. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005757.

Peng:2021:AMB

[PSWZ21] Jie Peng, Shi Shu, Junxian Wang, and Liuqiang Zhong. Adaptive-multilevel BDDC algorithm for threedimensional plane wave Helmholtz systems. Journal of Computational and Applied Mathematics, 381 (??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042720303022.

Pang:2022:FHO

[PSZ22] Xiangying Pang, Jiguang Sun, and Zhimin Zhang. FE-holomorphic operator function method for nonlinear plate vibrations with elastically added masses. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000425.

Palombi:2020:PAR

[PT20] Filippo Palombi and Simona Toti. A perturbative approach to the reconstruction of the eigenvalue spectrum of a normal covariance matrix from a spherically truncated counterpart. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303371.

Pandey:2021:ESS

[PT21] R. K. Pandey and Saurabh Tomar. An effective scheme for solving a class of nonlinear doubly singular boundary value problems through quasilinearization approach. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000303.

Pokharna:2024:SFB

[PT24] Nisha Pokharna and Indira P. Tripathi. A sublinear functional based approximated equivalence to optimality and duality for multiobjective programming in complex spaces. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001092.

Pulch:2022:IAP

[Pul22] Roland Pulch. Image analysis for patterns in solutions of reaction–diffusion equations. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002023.

Peng:2022:NMC

[PV22] Q. Peng and F. J. Vermolen. Numerical methods to compute stresses and displacements from cellular forces: Application to the contraction of tissue. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100515X.

Paredes:2024:RRM

[PVV24] Diego Paredes, Frédéric Valentin, and Henrique M. Versieux. Revisiting the robustness of the multiscale hybrid-mixed method: the face-based strategy. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300359X.

Porwal:2023:UAD

[PW23] Kamana Porwal and Tanvi Wadhawan. Unified analysis of discontinuous Galerkin methods for frictional contact problem with normal compliance. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002947.

Peng:2023:LBP

[PWL23] Zhichao Peng, Min Wang, and Fengyan Li. A learning-based projection method for model order reduction of transport problems. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002758.

Pronzato:2021:MEM

[PZ21] Luc Pronzato and Anatoly Zhigljavsky. Minimumenergy measures for singular kernels. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303800.

Peng:2023:TCI

[PZL23] Xingchun Peng, Hao Zhou, and Liuling Luo. Timeconsistent investment strategy for a DC pension plan with hidden Markov regime switching. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300002X.

Palakurthy:2024:NSE

[PZY^{+}24] Seshendra Palakurthy, Anup Zope, Yonghua Yan, Eric Collins, and Shanti Bhushan. Numerical study of the effect of micro vortices on chaotic flutter. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300345X.

Quell:2021:SMB

[QDHW21] Michael Quell, Georgios Diamantopoulos, Andreas Hössinger, and Josef Weinbub. Shared-memory block-based fast marching method for hierarchical meshes. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001072.

Qin:2020:SIB

[QG20] Xinyan Qin and Wenhao Gui. Statistical inference of Burr-XII distribution under progressive Type-II censored competing risks data with binomial removals. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302132.

Qi:2021:IFR

[QGW21] Meng Qi, Yunfan Gu, and Qiong Wang. Internet financial risk management and control based on improved rough set algorithm. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304702.

Qi:2023:EES

[QH23] Longzhao Qi and Yanren Hou. Error estimates for the Scalar Auxiliary Variable (SAV) schemes to the modified phase field crystal equation. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002850.

Qin:2021:VTS

[QHPL21] Yi Qin, Yanren Hou, Wenlong Pei, and Jian Li. A variable timestepping algorithm for the unsteady Stokes/Darcy model. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001400.

Qiu:2020:GSL

[Qiu20] Hua Qiu. Global stability of large solutions to the 3D nonhomogeneous incompressible MHD equations. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301047.

Qi:2020:EIF

[QJLQ20] Songqiao Qi, Kaijun Jin, Baisong Li, and Yufeng Qian. The exploration of Internetg finance by using neural network. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306351.

Qiao:2024:UES

[QQSZ24] Tian Qiao, Zhonghua Qiao, Shuyu Sun, and Shenggao Zhou. An unconditionally energy stable linear scheme for Poisson– Nernst–Planck equations. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000086.

Qi:2023:PWG

[QS23] Wenya Qi and Lunji Song. An over-penalized weak Galerkin method for parabolic interface problems with time-dependent coefficients. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004812.

Qi:2024:GWG

[QSW24] Wenya Qi, Padmanabhan Seshaiyer, and Junping Wang. A generalized weak Galerkin method for Oseen equation. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004557.

Quan:2024:LTS

[QWY24] Chaoyu Quan, Xu Wu, and Jiang Yang. Long time H^{1}-stability of fast L2-1_{σ} method on general nonuniform meshes for subdiffusion equations. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005915.

Qiao:2024:IFT

[QX24] Yu Qiao and Xiangtuan Xiong. Iterated fractional Tikhonov method for recovering the source term and initial data simultaneously in a twodimensional diffusion equation. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003121.

Qu:2023:ESP

[QXZF23] Wentao Qu, Xianchao Xiu, Haifei Zhang, and Jun Fan. An efficient semi-proximal ADMM algorithm for lowrank and sparse regularized matrix minimization problems with real-world applications. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006057.

Qiu:2022:ESC

[QXZG22] Wenlin Qiu, Da Xu, Jun Zhou, and Jing Guo. An efficient Sinc-collocation method via the DE transformation for eighth-order boundary value problems. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000322.

Qiu:2023:NVE

[QZW23] Jiali Qiu, Jikun Zhao, and Fei Wang. Nonconforming virtual element methods for the fourth-order variational inequalities of the first kind. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006239.

Rabbani:2021:COE

[Rab21] Mohsen Rabbani. Compact operators for existence of solution and projection method with multi-wavelet bases to solve (F.IES) and error analysis in Sobolev space. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303812.

Rathinasamy:2020:SOB

[RAN20] Anandaraman Rathinasamy, Davood Ahmadian, and Priya Nair. Secondorder balanced stochastic Runge–Kutta methods with multi-dimensional studies. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301813.

Ray:2023:TCN

[Ray23] S. Saha Ray. Two competent novel techniques based on two-dimensional wavelets for nonlinear variable-order Riesz spacefractional Schrödinger equations. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005696.

Ray:2019:TDW

[RB19] S. Saha Ray and S. Behera.

Two-dimensional wavelets operational method for solving Volterra weakly singular partial integrodifferential equations. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304145.

Ren:2023:BIM

[RB23a] Quanwei Ren and Huiqun Bai. Balanced implicit methods with strong order 1.5 for solving stochastic differential equations. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000134.

Russell:2023:CTA

[RB23b] Thomas Russell and Pavel Bedrikovetsky. Colloidal transport in anisotropic porous media: Kinetic equation and its upscaling. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004940.

Rochera:2024:ICI

[RB24] David Rochera and Michael Barton. On inverse construction of isoptics and isochordal-viewed curves. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300376X.

Rodriguez:2023:WEE

[RBR23] Ana Alonso Rodríguez, Ludovico Bruni Bruno, and Francesca Rapetti. Whitney edge elements and the Runge phenomenon. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000614.

Rosa:2020:PCM

[RCM20] M. A. Rosa, J. A. Cuminato, and S. McKee. A polynomial collocation method for singular integro-differential equations in weighted spaces. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930531X.

Rohan:2021:MAW

[RCN21] Eduard Rohan, Robert Cimrman, and Salah Naili. Modelling of acoustic waves in homogenized fluidsaturated deforming poroelastic periodic structures under permanent flow. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001552.

Ranilla-Cortina:2024:PET

[RCVA24] Sandra Ranilla-Cortina and Jesús Vigo-Aguiar. Performance enhancement through portfolio optimization of delayed insider information: an analysis and implementation study. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001043.

Rausa:2023:MPS

[RDG23] Andrea Rausa, Alessandro Donizetti, and Alberto Guardone. Multi-physics simulation of 3D in-flight ice-shedding. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300170X.

Roelfs:2021:QSD

[RDH21] Martin Roelfs, David Dudal, and Daan Huybrechs. Quaternionic step derivative: Machine precision differentiation of holomorphic functions using complex quaternions. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003216.

Rabbani:2020:EST

[RDHA20] Mohsen Rabbani, Anupam Das, Bipan Hazarika, and Reza Arab. Existence of solution for two dimensional nonlinear fractional integral equation by measure of noncompactness and iterative algorithm to solve it. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306594.

Rettieva:2024:CMD

[Ret24] Anna Rettieva. Cooperation maintenance in dynamic discrete-time multicriteria games with application to bioresource management problem. Journal of Computational and Applied Mathematics, 441 (??):??, May 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006428.

Romanengo:2022:HTB

[RFB22] Chiara Romanengo, Bianca Falcidieno, and Silvia Biasotti. Hough transform based recognition of space curves. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002448.

Rabin:2023:MEE

[RFF23] Neta Rabin, Ángela Fernández, and Dalia Fishelov. Multiscale extensions for enhancing coarse grid computations. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000602.

Rebollo:2020:AVS

[RFGGM20] Tomás Chacón Rebollo, Soledad Fernández-García, and Macarena Gómez-Mármol. Anisotropic VMS solution of advection– diffusion problems by spectral approximation of subgrid scales. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302508.

Roul:2020:NHO

[RG20a] Pradip Roul and V. M.

K. Prasad Goura. A new higher order compact finite difference method for generalised black–Scholes partial differential equation: European call option. Journal of Computational and Applied Mathematics, 363(??):464–484, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302882.

Roul:2020:SON

[RG20b] Pradip Roul and V. M. K. Prasad Goura. A sixth order numerical method and its convergence for generalized Black–Scholes PDE. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301722.

Ranjbar:2021:GQR

[RG21] H. Ranjbar and F. Ghoreishi. A Gaussian quadrature rule for Fourier-type highly oscillatory integrals in the presence of stationary points. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002120.

Roul:2022:ENM

[RG22] Pradip Roul and V. M. K. Prasad Goura. An efficient numerical method based on redefined cubic B-spline basis functions for pricing Asian options. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003964.

Roul:2023:ENS

[RG23] Pradip Roul and V. M.

K. Prasad Goura. An efficient numerical scheme and its stability analysis for a time-fractional reaction diffusion model. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005167.

Roohollahi:2020:NSM

[RGA20] A. Roohollahi, B. Ghazanfari, and S. Akhavan. Numerical solution of the mixed Volterra–Fredholm integro-differential multiterm equations of fractional order. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301199.

Rohan:2023:HNA

[RH23] Eduard Rohan and Jan Heczko. Homogenization and numerical algorithms for two-scale modeling of porous media with self-contact in micropores. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002200.

Roozbeh:2020:RES

[RHA20] M. Roozbeh, G. Hesamian, and M. G. Akbari. Ridge estimation in semi-parametric regression models under the stochastic restriction and correlated elliptically contoured errors. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302314.

Rashidi:2022:GFL

[RHM22] Saeede Rashidi, S. Reza Hejazi, and Fatemeh Mohammadizadeh. Group formalism of Lie transformations, conservation laws, exact and numerical solutions of non-linear timefractional Black–Scholes equation. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004866.

Ri:2022:LFB

[Ri22] ChiWon Ri. Lattice factorization based symmetric PMI paraunitary matrix extension and construction of symmetric orthogonal wavelets. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000577.

Ri:2024:LFB

[Ri24] ChiWon Ri. Lattice factorization based causal symmetric paraunitary matrix extension and construction of symmetric orthogonal multiwavelets. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001213.

Rios:2022:LSG

[RJ22] Dany Rios and Bert Jüttler. LSPIA, (stochastic) gradient descent, and parameter correction. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005446.

Rakshit:2022:SLS

[RK22] Suman Rakshit and S. R. Khare. Solution of the linearly structured partial polynomial inverse eigenvalue problem. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000930.

Ruppenthal:2023:OCU

[RK23] Falko Ruppenthal and Dmitri Kuzmin. Optimal control using flux potentials: a way to construct bound-preserving finite element schemes for conservation laws. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002959.

Rauh:2020:IMC

[RKA20] Andreas Rauh, Julia Kersten, and Harald Aschemann. Interval methods and contractor-based branch-and-bound procedures for verified parameter identification of quasilinear cooperative system models. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930487X.

Rashid:2022:AAV

[RKS^{+}22] Saima Rashid, Khadija Tul Kubra, Sobia Sultana, Praveen Agarwal, and M. S. Osman. An approximate analytical view of physical and biological models in the setting of Caputo operator via Elzaki transform decomposition method. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001698.

Rump:2023:FCE

[RL23] Siegfried M. Rump and Marko Lange. Fast computation of error bounds for all eigenpairs of a Hermitian and all singular pairs of a rectangular matrix with emphasis on eigenand singular value clusters. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002765.

Rohan:2024:HAT

[RL24] E. Rohan and V. Lukes. Homogenization of the acoustic transmission on periodically perforated plates interacting with potential mean flow. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004533.

Roanes-Lozano:2021:CAO

[RLGGAV21] Eugenio Roanes-Lozano, José Luis Galán-García, and Gabriel Aguilera-Venegas. A computer approach to overtaking station track layout diagram design using graphs. An alternative track diagram proposal for these stations. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000741.

Ren:2021:SNE

[RlLZZ21] Jincheng Ren, Hong lin Liao, Jiwei Zhang, and Zhimin Zhang. Sharp H^{1}- norm error estimates of two time-stepping schemes for reaction–subdiffusion problems. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306439.

Rausa:2021:NMR

[RMG21] Andrea Rausa, Myles Morelli, and Alberto Guardone. A novel method for robust and efficient prediction of ice shedding from rotorcraft blades. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000716.

Ruymbeek:2022:TKM

[RMM22] Koen Ruymbeek, Karl Meerbergen, and Wim Michiels. Tensor-Krylov method for computing eigenvalues of parameterdependent matrices. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004921.

Rahimkhani:2020:ASN

[RO20] Parisa Rahimkhani and Yadollah Ordokhani. Approximate solution of nonlinear fractional integrodifferential equations using fractional alternative Legendre functions. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303681.

Roddick:2021:CSM

[Rod21] Greg Roddick. Computation of scattering matrices and their derivatives for waveguides. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000728.

Ross:2023:GNA

[Ros23] I. M. Ross. Generating Nesterov’s accelerated gradient algorithm by using optimal control theory for optimization. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005660.

Roul:2022:RNT

[Rou22] Pradip Roul. A robust numerical technique and its analysis for computing the price of an Asian option. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200259X.

Roul:2024:ENA

[Rou24]  Pradip Roul.  Efficient

numerical algorithms for solving a time-fractional diffusion equation with weakly singular solution. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002838.

Roy:2024:GBD

[RPQ24] Priyanka Roy, Geetanjali Panda, and Dong Qiu. Gradient-based descent linesearch to solve interval-valued optimization problems under gHdifferentiability with application to finance. Journal of Computational and Applied Mathematics, 436 (??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723003461.

Ruas:2021:ENF

[RR21] Vitoriano Ruas and Marco Antonio Silva Ramos. Efficiency of nonparametric finite elements for optimal-order enforcement of Dirichlet conditions on curvilinear boundaries. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001424.

Ramos:2022:TSH

[RR22a] Higinio Ramos and Mufutau Ajani Rufai. A twostep hybrid block method with fourth derivatives for solving third-order boundary value problems. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000388.

Ribeiro:2022:MLS

[RR22b] Luana L. Silva Ribeiro and A. Sri Ranga. A modified least squares method: Approximations on the unit circle and on (−1, 1). Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000486.

Roul:2022:HON

[RR22c] Pradip Roul and Vikas Rohil. A high order numerical technique and its analysis for nonlinear generalized Fisher’s equation. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721006142.

Rukavishnikov:2022:EEF

[RR22d] Viktor A. Rukavishnikov and Elena I. Rukavishnikova. Error estimate FEM for the nikol’skij-Lizorkin problem with degeneracy. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004647.

Rufai:2023:STO

[RR23a] Mufutau Ajani Rufai and Higinio Ramos. Solving third-order Lane–Emden– Fowler equations using a variable stepsize formulation of a pair of block methods. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003958.

Rukavishnikov:2023:TAC

[RR23b] Viktor A. Rukavishnikov and Alexey V. Rukavishnikov. Theoretical analysis and construction of numerical method for solving the Navier–Stokes equations in rotation form with corner singularity. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001620.

Remili:2024:MGF

[RRS24] Walid Remili, Azedine Rahmoune, and Abdselam Silem. Mapped Gegenbauer functions for solving Hammerstein generalized integral equations with Green’s kernels on the whole line. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400089X.

Reichel:2022:AGQ

[RS22a] Lothar Reichel and Miodrag M. Spalević. Averaged Gauss quadrature formulas: Properties and applications. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000887.

Reisinger:2022:AEM

[RS22b] Christoph Reisinger and Wolfgang Stockinger. An adaptive Euler–Maruyama scheme for McKean–Vlasov SDEs with super-linear growth and application to the mean-field FitzHugh– Nagumo model. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003472.

Ray:2023:PEE

[RS23] Tanushree Ray and Rajen Kumar Sinha. A posteriori error estimation and adaptive mesh refinement for parabolic interface problems using non-conforming immersed finite element method. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003612.

Rathore:2024:NMS

[RS24a] Ajay Singh Rathore and Vembu Shanthi. A numerical method for a system of singularly perturbed Fredholm integro-differential reaction–diffusion equation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004016.

Rathore:2024:NSS

[RS24b] Ajay Singh Rathore and Vembu Shanthi. A numerical solution of singularly perturbed Fredholm integro-differential equation with discontinuous source term. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001079.

Reichel:2024:RLT

[RS24c] Lothar Reichel and Miodrag M. Spalević. Radauand Lobatto-type averaged Gauss rules. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004193.

Roth:2024:MST

[RSRW24] Julian Roth, Martyna Soszyńska, Thomas Richter, and Thomas Wick. A monolithic space-time temporal multirate finite element framework for interface and volume coupled problems. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000803.

Randig:2024:NMP

[RSS24] Marvin Randig, Dierk Schleicher, and Robin Stoll. Newton’s method in practice, II: the iterated refinement Newton method and near-optimal complexity for finding all roots of some polynomials of very large degrees. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003710.

Reich:2024:NIA

[RTT24] Simeon Reich, Truong Minh Tuyen, and Nguyen Thi Trang. New iterative algorithms for solving a class of split common solution problems and their applications. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005812.

Rodriguez:2020:NBN

[RTV20] J. M. Rodríguez and R. Taboada-Vázquez. Numerical behaviour of a new LES model with nonlinear viscosity. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030159X.

Reichel:2022:WTG

[RU22] Lothar Reichel and Ugochukwu O.

Ugwu. Weighted tensor Golub–Kahan–Tikhonovtype methods applied to image processing using a t-product. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002369.

Rujivan:2023:VVD

[Ruj23] Sanae Rujivan. Valuation of volatility derivatives with time-varying volatility: an analytical probabilistic approach using a mixture distribution for pricing nonlinear payoff volatility derivatives in discrete observation case. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003405.

Rump:2020:VBD

[Rum20] Siegfried M. Rump. Verified bounds for the determinant of real or complex point or interval matrices. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306156.

Ruymbeek:2020:ARD

[RV20] Koen Ruymbeek and Wim Vanroose. Algorithm for the reconstruction of dynamic objects in CTscanning using optical flow. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304625.

Raja:2022:ACR

[RVS^{+}22] M. Mohan Raja, V. Vijayakumar, Anurag Shukla, Kottakkaran Sooppy Nisar, and Haci Mehmet Baskonus. On the approximate controllability results for fractional integrodifferential systems of order 1 < r < 2 with sectorial operators. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002382.

Rebholz:2020:AAC

[RVX20] Leo G. Rebholz, Alex Viguerie, and Mengying Xiao. Analysis of algebraic Chorin temam splitting for incompressible NSE and comparison to Yosida methods. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303693.

Ruan:2022:EDG

[RWW22] Jialin Ruan, Lijin Wang, and Pengjun Wang. Exponential discrete gradient schemes for a class of stochastic differential equations. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004192.

Rao:2024:SBP

[RYH24] Jiayun Rao, Chaozhi Yu, and Na Huang. Stabilized BB projection algorithm for large-scale convex constrained nonlinear monotone equations to signal and image processing problems. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001663.

Ruan:2020:SIT

[RZ20] Zhousheng Ruan and Sen Zhang. Simultaneous inversion of time-dependent source term and fractional order for a time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305710.

Rasheed:2024:DED

[RZA^{+}24] Zahid Rasheed, Hongying Zhang, Syed Masroor Anwar, Muhammad Noor ul Amin, Nurudeen A. Adegoke, and Saddam Akber Abbasi. Designing efficient dispersion control charts under various ranked-set sampling approaches. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006234.

Shustin:2023:ROP

[SA23] Boris Shustin and Haim Avron. Riemannian optimization with a preconditioning scheme on the generalized Stiefel manifold. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005519.

Sanga:2024:PAC

[SA24] Sudeep Singh Sanga and Khushbu S. Antala. Performance analysis and cost investigations for statedependent single unreliable server finite queue under F -policy using GA and QNM. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006222.

Saadaoui:2020:QEA

[Saâ20] Foued Saâdaoui. Quadratic extrapolation for accelerating convergence of the EM fixed point problem. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305825.

Singh:2021:ETI

[SACQ21] Saroja Kumar Singh, Sarat Kumar Acharya, Frederico R. B. Cruz, and Roberto C. Quinino. Estimation of traffic intensity from queue length data in a deterministic single server queueing system. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003150.

Sadkane:2023:NCI

[Sad23]  Miloud Sadkane.  A

note on the computation of invariant pairs of quadratic matrix polynomials. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002789.

Sajavicius:2023:HLS

[Saj23] Svajunas Sajavicius. Hyperpower least squares progressive iterative approximation. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004861.

Salman:2021:MMC

[Sal21a] Sanaa Moussa Salman.

Memory and media coverage effect on an HIV/ AIDS epidemic model with treatment. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304945.

Sober:2021:AFM

[SAL21b] Barak Sober, Yariv Aizenbud, and David Levin. Approximation of functions over manifolds: a moving least-squares approach. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304313.

Sang:2021:AGA

[San21] Bin Sang. Application of genetic algorithm and BP neural network in supply chain finance under information sharing. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304611.

Sriram:2024:DNC

[SAN^{+}24] Gokulakrishnan Sriram, Ahmed M. Ali Ali, Hayder Natiq, Atefeh Ahmadi, Karthikeyan Rajagopal, and Sajad Jafari. Dynamics of a novel chaotic map. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003977.

Skowronek:2024:MGS

[SAZW24] Katarzyna Skowronek, Marek Arendarczyk, Radosław Zimroz, and Agnieszka Wyłomańska. Modified Greenwood statistic and its application for statistical testing. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003716.

Smirnova:2024:TNS

[SB24a] Alexandra Smirnova and Mona Baroonian. Theoretical and numerical study of case reporting rate with application to epidemiology. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003546.

Stocco:2024:SMF

[SB24b] Davide Stocco and Enrico Bertolazzi. Symbolic matrix factorization for differential-algebraic equations index reduction. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001481.

Saifia:2020:SMF

[SBC20] O. Saifia, D. Boucenna, and A. Chidouh. Study of Mainardi’s fractional heat problem. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030234X.

Sirakov:2024:EVF

[SBC^{+}24] N. M. Sirakov, A. Bowden, M. Chen, L. H. Ngo, and M. Luong. Embedding vector field into image features to enhance classification. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006283.

Sharma:2024:RIF

[SBKR24] Himani Sharma, Ramandeep Behl, Munish Kansal, and Higinio Ramos. A robust iterative family for multiple roots of nonlinear equations: Enhancing accuracy and handling critical points. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400044X.

Silva:2021:NVA

[SBM21] Ivair R. Silva, Debanjan Bhattacharjee, and Nitis Mukhopadhyay. Numerical versus asymptotic sequential interval estimation of population sizes. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100340X.

Shargatov:2021:SAT

[SC21] V. A. Shargatov and A. P. Chugainova. Stability analysis of traveling wave solutions of a generalized Korteweg–de Vries– Burgers equation with variable dissipation parameter. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002764.

Shi:2023:IRC

[SC23a] Chengxin Shi and Hao Cheng. Identify the Robin coefficient in an inhomogeneous time-fractional diffusion-wave equation. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002819.

Solin:2023:SAC

[SC23b] Pavel Solin and Jakub Cerveny. Selected aspects of constrained conforming approximation in higherorder FEM. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000626.

Shahsanaei:2024:WCD

[SC24] Fatemeh Shahsanaei and Rahim Chinipardaz. Weighted circular distributions. Journal of Computational and Applied Mathematics, 436 (??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723003205.

Sanaullah:2024:ECC

[SCHS24] Aamir Sanaullah, Aamir Majeed Chaudhary, Muhammad Hanif, and Prayas Sharma. EWMA control charts based on robust estimators: a powerful tool for monitoring a process with a non-normal distribution. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001699.

Shargatov:2022:GST

[SCK22] V. A. Shargatov, A. P.

Chugainova, and G. V. Kolomiytsev. Global stability of traveling wave solutions of generalized Korteweg–de Vries–Burgers equation with non-constant dissipation parameter. Journal of Computational and Applied Mathematics, 412 (??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722001571.

Shen:2023:CRF

[SCY23] Zhengwei Shen, Qian Chen, and Fan Yang. A convex relaxation framework consisting of a primaldual alternative algorithm for solving l_{0} sparsityinduced optimization problems with application to signal recovery based image restoration. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004769.

Sweidan:2020:SWS

[SCZ20] Mohyeedden Sweidan, Xiaojun Chen, and Xiaoming Zheng. The Shortley– Weller scheme for variable coefficient two-point boundary value problems and its application to tumor growth problem with heterogeneous microenvironment. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301655.

Sun:2021:IAN

[SD21] D. Y. Sun and C. Y. Dong.

Isogeometric analysis of the new integral formula for elastic energy change of heterogeneous materials. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303976.

Sun:2024:ESN

[SD24] Jingjing Sun and Shouqiang Du. An effective smoothing Newton projection algorithm for finding sparse solutions to NP-hard tensor complementarity problems. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003248.

Shen:2020:FPT

[SDC20] Shujun Shen, Weizhong Dai, and Jinfa Cheng. Fractional parabolic twostep model and its accurate numerical scheme for nanoscale heat conduction. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301035.

Shehu:2022:FAI

[SDL22] Yekini Shehu, Qiao-Li Dong, and Lu-Lu Liu. Fast alternated inertial projection algorithms for pseudomonotone variational inequalities. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002540.

Shen:2023:ANS

[SDLZ23] Shujun Shen, Weizhong Dai, Qingxia Liu, and Pinghui Zhuang. Accurate numerical scheme for solving fractional diffusionwave two-step model for nanoscale heat conduction. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003636.

Scarabel:2021:NBA

[SDV21] Francesca Scarabel, Odo Diekmann, and Rossana Vermiglio. Numerical bifurcation analysis of renewal equations via pseudospectral approximation. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002338.

Shen:2024:ODB

[SDW24a] Peiping Shen, Yaping Deng, and Yafei Wang. A one-dimensional branching rule based branch-and-bound algorithm for minimax linear fractional programming. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272400150X.

Sun:2024:LRN

[SDW24b] Jingbo Sun, Suchuan Dong, and Fei Wang. Local randomized neural networks with discontinuous Galerkin methods for partial differential equations. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000797.

Salhi:2022:GCU

[SEAS22] Loubna Salhi, Mofdi El-Amrani, and Mohammed Seaid. A Galerkincharacteristic unified finite element method for moving thermal fronts in porous media. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304507.

Segeth:2021:MDF

[Seg21] Karel Segeth. Multivariate data fitting using polyharmonic splines. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002739.

Segeth:2023:SRB

[Seg23]  Karel Segeth.  Spheri-

cal radial basis function approximation of some physical quantities measured. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000729.

Saemi:2020:ESS

[SES20] Fereshteh Saemi, Hamideh Ebrahimi, and Mahmoud Shafiee. An effective scheme for solving system of fractional Volterra– Fredholm integro-differential equations based on the Müntz–Legendre wavelets. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300649.

Saemi:2023:DSV

[SESH23] Fereshteh Saemi, Hamideh Ebrahimi, Mahmoud Shafiee, and Kamyar Hosseini. A detailed study on 2D Volterra–Fredholm integrodifferential equations involving the Caputo fractional derivative. Journal of Computational and Applied Mathematics, 420 (??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722004186.

Suzuki:2023:NIP

[SFI23] Kengo Suzuki, Takeshi Fukaya, and Takeshi Iwashita. A novel ILU preconditioning method with a block structure suitable for SIMD vectorization. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003478.

Simoes:2023:NNM

[SFMdD23] Alexandre Anahory Simoes, Sebastián J. Ferraro, Juan Carlos Marrero, and David Martín de Diego. A nonholonomic Newmark method. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200471X.

Sun:2020:NFD

[SGD20] F. L. Sun, Y. P. Gong, and C. Y. Dong. A novel fast direct solver for 3D elastic inclusion problems with the isogeometric boundary element method. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301953.

Shen:2022:CAN

[SGY22] Yue Shen, Wei Gong, and Ningning Yan. Convergence of adaptive nonconforming finite element method for Stokes optimal control problems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001194.

Svacek:2021:FNS

[SH21] Petr Svácek and Jaromír Horácek. FE numerical simulation of incompressible airflow in the glottal channel periodically closed by self-sustained vocal folds vibration. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001485.

Schulze-Halberg:2023:DAT

[SH23] Axel Schulze-Halberg. Darboux algorithms for twodimensional Dirac equations with upper triangular potential matrix. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000870.

Su:2024:HOK

[SH24a] Tran Van Su and Dinh Dieu Hang. Higher-order KKT optimality conditions through contingent derivatives for constrained nonsmooth vector equilibrium problems. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001651.

Sykora:2024:SPC

[SH24b] Jan Sýkora and Jan Havelka. Sparse polynomial chaos expansions for uncertainty quantification in thermal tomography. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003503.

Shang:2021:NTL

[Sha21] Yueqiang Shang. A new two-level defect-correction method for the steady Navier–Stokes equations. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720303009.

Shayegan:2023:CRB

[Sha23] Amir Hossein Salehi Shayegan.

Coupling RBF-based meshless method and Landweber iteration algorithm for approximating a spacedependent source term in a time fractional diffusion equation. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002618.

Shang:2024:TLF

[Sha24] Yueqiang Shang. A twolevel finite element method with grad-div stabilizations for the incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001146.

Sahoo:2024:MOS

[SHB24] Tanmay Sahoo, Nil Kamal Hazra, and Narayanaswamy Balakrishnan. On multivariate orderings of some general ordered random vectors. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003509.

Shi:2021:IDN

[Shi21a] Kehan Shi. Image denoising by nonlinear nonlocal diffusion equations. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002259.

Shirokov:2021:NHS

[Shi21b] D. S. Shirokov. A note on the hyperbolic singular value decomposition without hyperexchange matrices. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000698.

Su:2022:NMS

[SHV^{+}22] Lingde Su, Jian Huang, V. I. Vasil’ev, Ao Li, and A. M. Kardashevsky. A numerical method for solving retrospective inverse problem of fractional parabolic equation. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001637.

Sun:2020:MDA

[SHZ20] Weiwei Sun, Xiang Hu, and Lianzeng Zhang. Moments of discounted aggregate claims with dependence based on Spearman copula. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301801.

Sugihara:2023:GUP

[SHZ23] Kota Sugihara, Ken Hayami, and Liao Zeyu. GM-RES using pseudoinverse for range symmetric singular systems. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004630.

Sastre:2023:BFE

[SI23] Jorge Sastre and Javier Ibáñez. On the backward and forward error of approximations of analytic functions and applications to the computation of matrix functions. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003570.

Shurina:2022:DTS

[SIS^{+}22] E. P. Shurina, N. B.

Itkina, N. V. Shtabel, E. I. Shtanko, A. Yu. Kutishcheva, S. I. Markov, and D. V. Dobrolubova. Determination of thermal, stiffness and electrical effective tensors in composite media. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005951.

Sixou:2020:CRM

[Six20] Bruno Sixou. Convergence of regularization methods with filter functions for a regularization parameter chosen with GSURE and mildly ill-posed inverse problems. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302296.

Song:2020:ISH

[SJZ20] Xiaoyan Song, Lijian Jiang, and Guang-Hui Zheng. Implicit sampling for hierarchical Bayesian inversion and applications in fractional multiscale diffusion models. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301175.

Shilnikov:2023:OAN

[SK23a] K. E. Shilnikov and M. B. Kochanov. On one approach for the numerical solving of hyperbolic initial-boundary problems on an adaptive moving grids. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004824.

Swietlicka:2023:HEM

[ŚK23b] Aleksandra Świetlicka and Krzysztof Kolanowski. Homogeneous ensemble model built from artificial neural networks for fault detection in navigation systems. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002236.

Saha:2024:WRV

[SK24a] Shital Saha and Suchandan Kayal. Weighted (residual) varentropy and its applications. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006532.

Sharma:2024:EDS

[SK24b] Pallvi Sharma and Munish Kansal. Extraction of deflating subspaces using disk function of a matrix pencil via matrix sign function with application in generalized eigenvalue problem. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006726.

Seydaoglu:2020:ENT

[SKE20] Muaz Seydaoglu, Hüseyin Koçak, and Utku Erdogan. An efficient numerical treatment for the asymptotic behaviour of the nonlinear Airy-type problems. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301242.

Smidova:2022:FES

[SKS22] Eliska Smídová, Petr Kabele, and Michal Sejnoha. Finite element simulation of single edge notched timber arch. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002983.

Sumit:2022:ANS

[SKVA22] Sumit, Sunil Kumar, and Jesus Vigo-Aguiar. Analysis of a nonlinear singularly perturbed Volterra integrodifferential equation. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000297.

Singh:2023:FDS

[SKVA23] Anshima Singh, Sunil Kumar, and Jesus Vigo-Aguiar. A fully discrete scheme based on cubic splines and its analysis for time-fractional reaction– diffusion equations exhibiting weak initial singularity. Journal of Computational and Applied Mathematics, 434(??):??, December 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002820.

Sun:2023:GPN

[SKZ23] Jun Sun, Lingchen Kong, and Shenglong Zhou. Gradient projection Newton algorithm for sparse collaborative learning using synthetic and real datasets of applications. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004708.

Shin:2020:ESR

[SL20] Jaemin Shin and June-Yub Lee. An energy stable Runge–Kutta method for convex gradient problems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304583.

Schwarz:2021:ICC

[SL21] Diana Estévez Schwarz and René Lamour. InitDAE: Computation of consistent values, index determination and diagnosis of singularities of DAEs using automatic differentiation in Python. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304893.

Samar:2022:PAC

[SL22] Mahvish Samar and Fu-Rong Lin. Perturbation analysis and condition numbers for the Tikhonov regularization of total least squares problem and their statistical estimation. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000875.

Shin:2023:PSJ

[SL23] Yong Hyun Shin and Ho-Seok Lee. Portfolio selection and job switching with CARA utility. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300064X.

Song:2024:EIS

[SL24a] Taiyong Song and Zexian Liu. An efficient inertial subspace minimization CG algorithm with convergence rate analysis for constrained nonlinear monotone equations. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001225.

Sun:2024:NSO

[SL24b] Hongli Sun and Yanfei Lu. Numerical solutions to one dimensional linear Volterra–Fredholm integral equations based on LS-SVM model. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002632.

Stoyanovskaya:2024:DAS

[SLA^{+}24] O. P. Stoyanovskaya, V. V. Lisitsa, S. A. Anoshin, T. A. Savvateeva, and T. V. Markelova. Dispersion analysis of SPH as a way to understand its order of approximation. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004399.

Stoklasa:2022:RBP

[SLC22] Jan Stoklasa, Pasi Luukka, and Mikael Collan. On the relationship between possibilistic and standard moments of fuzzy numbers. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001145.

Siwik:2022:TTD

[SŁK^{+}22] Leszek Siwik, Marcin Łoś, Adrian Kłusek, Anna Paszyńska, Keshav Pingali, Witold Dzwinel, and Maciej Paszyński. Tuning three-dimensional tumor progression simulations on a cluster of GPG-PUs. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001303.

Sheng:2024:CPA

[SLK24] Danshu Sheng, Chang Liu, and Yao Kang. Changepoint analysis for binomial autoregressive model with application to price stability counts. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003285.

Su:2024:PEG

[SLLP24] Xin Su, Wing Tat Leung, Wenyuan Li, and Sai-Mang Pun. Partially explicit generalized multiscale finite element methods for poroelasticity problem. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001857.

Slodicka:2021:PPM

[Slo21] Marian Slodicka. Parabolic problem for moving/evolving body with perfect contact to neighborhood. Journal of Computational and Applied Mathematics, 391 (??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721000807.

Shutyaev:2020:SRF

[SLP20] V. Shutyaev, F.-X. Le Dimet, and E. Parmuzin. Sensitivity of response functions in variational data assimilation for joint parameter and initial state estimation. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303711.

Schneider:2021:SCI

[SLW21] Moritz Schneider, Jens Lang, and Rüdiger Weiner. Super-convergent implicit– explicit peer methods with variable step sizes. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305047.

Sun:2022:DTN

[SLW^{+}22] Zhongbo Sun, Yongbai Liu, Gang Wang, Yufeng Lian, Keping Liu, and Long Jin. Discrete-time noise-tolerant z-type model for online solving nonlinear time-varying equations in the presence of noises. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004465.

Shi:2024:FOC

[SLXF24] Yilei Shi, Dong Liang, Shusen Xie, and Kai Fu. Fourth-order compact block-centered splitting domain decomposition method for parabolic equations. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001833.

Shen:2020:ISC

[SLZD20] Cuifeng Shen, Yan Li, Xuelai Zhu, and Wenyong Duan. Improved stability criteria for linear systems with two additive time-varying delays via a novel Lyapunov functional. Journal of Computational and Applied Mathematics, 363(??):312–324, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719302833.

Sun:2023:CSI

[SLZL23] Meng Sun, Lin Lan, Chun-Gang Zhu, and Fengchun Lei. Cubic spline interpolation with optimal end conditions. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006379.

Samadyar:2020:NST

[SM20] Nasrin Samadyar and Farshid Mirzaee. Numerical solution of twodimensional stochastic Fredholm integral equations on hypercube domains via meshfree approach. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301667.

Santra:2022:NFD

[SM22a] S. Santra and J. Mohapatra. A novel finite difference technique with error estimate for time fractional partial integrodifferential equation of Volterra type. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100368X.

Solano:2022:UHM

[SM22b] Manuel Solano and Felipe Vargas M. An unfitted HDG method for Oseen equations. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003435.

Saffarian:2023:SST

[SM23] Marziyeh Saffarian and Akbar Mohebbi. Solution of space-time tempered fractional diffusion-wave equation using a high-order numerical method. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005337.

Shakti:2022:MMR

[SMDVA22] Deepti Shakti, Jugal Mohapatra, Pratibhamoy Das, and Jesus Vigo-Aguiar. A moving mesh refinement based optimal accurate uniformly convergent computational method for a parabolic system of boundary layer originated reaction–diffusion problems with arbitrary small diffusion terms. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304581.

Sereeter:2021:NLP

[SMKV21] B. Sereeter, A. S. Markensteijn, M. E. Kootte, and C. Vuik. A novel linearized power flow approach for transmission and distribution networks. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001928.

Sayevand:2022:ADB

[SMM22] K. Sayevand, J. Tenreiro Machado, and I. Masti. Analysis of dual Bernstein operators in the solution of the fractional convection– diffusion equation arising in underground water pollution. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003514.

Sun:2023:ACJ

[SMR23] Q. Sun, T. Miyoshi, and S. Richard. Analysis of COVID-19 in Japan with extended SEIR model and ensemble Kalman filter. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003910.

Sky:2024:HOB

[SMRN24] Adam Sky, Ingo Muench, Gianluca Rizzi, and Patrizio Neff. Higher order Bernstein–Bézier and Nédélec finite elements for the relaxed micromorphic model. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005125.

Sathar:2021:DWE

[SN21a] E. I. Abdul Sathar and R. Dhanya Nair. On dynamic weighted extropy. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001266.

Sheng:2021:CTE

[SN21b] Zhou Sheng and Qin Ni. Computing tensor Z-eigenvalues via shifted inverse power method. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003393.

Schumann:2023:LMD

[SN23] Yannis Schumann and Philipp Neumann. On linear models for discrete operator inference in time dependent problems. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006203.

Sergeyev:2021:NLT

[SNMK21] Yaroslav D. Sergeyev, Maria Chiara Nasso, Marat S. Mukhametzhanov, and Dmitri E. Kvasov. Novel local tuning techniques for speeding up onedimensional algorithms in expensive global optimization using Lipschitz derivatives. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304258.

Solin:2024:E

[Sol24]  Pavel Solin.  Editorial.

Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001602.

Song:2024:ERB

[Son24] Yanlai Song. An efficient radial basis function generated finite difference meshfree scheme to price multidimensional PDEs in financial options. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003266.

Sousa:2022:CRD

[Sou22] Ercília Sousa. The convergence rate for difference approximations to fractional boundary value problems. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002357.

Su:2022:MMH

[SP22] Xin Su and Sai-Mang Pun. A multiscale method for the heterogeneous Signorini problem. Journal of Computational and Applied Mathematics, 409(??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000449.

Sun:2021:NIF

[SPC21] Xiang Sun, Xiaomin Pan, and Jung-Il Choi. Nonintrusive framework of reduced-order modeling based on proper orthogonal decomposition and polynomial chaos expansion. Journal of Computational and Applied Mathematics, 390(??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306634.

Saminger-Platz:2024:PTT

[SPKS^{+}24] Susanne Saminger-Platz, Anna Kolesárová, Adam Seliga, Radko Mesiar, and Erich Peter Klement. Parameterized transformations and truncation: When is the result a copula? Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002844.

Shahnazi-Pour:2021:NSH

[SPMB21] A. Shahnazi-Pour, B. Parsa Moghaddam, and A. Babaei. Numerical simulation of the Hurst index of solutions of fractional stochastic dynamical systems driven by fractional Brownian motion. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030501X.

Sharma:2021:VEI

[SPS21] Nitu Sharma, Puneet Pasricha, and Dharmaraja Selvamuthu. Valuation of equity-indexed annuities under correlated jumpdiffusion processes. Journal of Computational and Applied Mathematics, 395 (??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721001953.

Soto-Quiros:2024:FRV

[SQT24] Pablo Soto-Quiros and Anatoli Torokhti. Fast random vector transforms in terms of pseudo-inverse within the Wiener filtering paradigm. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001778.

Sirianni:2023:MWI

[SRAG23] Giuseppe Sirianni, Barbara Re, Remi Abgrall, and Alberto Guardone. Momentum Weighted Interpolation for unsteady weakly compressible twophase flows on unstructured meshes. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300153X.

Sharma:2023:NTS

[SRBK23] Puneet Sharma, Higinio Ramos, Ramandeep Behl, and Vinay Kanwar. A new three-step fixed point iteration scheme with strong convergence and applications. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001863.

Serna-Reyes:2022:TAC

[SRMD22] Adán J. Serna-Reyes and J. E. Macías-Díaz. Theoretical analysis of a conservative finite-difference scheme to solve a Riesz space-fractional Gross– Pitaevskii system. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000327.

Sarshar:2021:ADI

[SRS21] Arash Sarshar, Steven Roberts, and Adrian Sandu. Alternating directions implicit integration in a general linear method framework. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306247.

Shakya:2020:FEM

[SS20a] Pratibha Shakya and Rajen Kumar Sinha. Finite element method for parabolic optimal control problems with a bilinear state equation. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304340.

Singh:2020:SCH

[SS20b] Birendra Singh and Uaday Singh. Some characterizations of Hausdorff matrices and their application to Fourier approximation. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304534.

Srichanthamit:2020:MMF

[SS20c] Tanapun Srichanthamit and Hidetsugu Suto. A mathematical model for a framework of a supporting system for international trade transaction. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301011.

Schrocker:2023:PFD

[SS23a] Hans-Peter Schröcker and Zbynek Sír. Partial fraction decomposition for rational Pythagorean hodograph curves. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001401.

Shukla:2023:CSZ

[SS23b] Vinay Shukla and A. Swaminathan. Chain sequences and zeros of polynomials related to a perturbed R_{II} type recurrence relation. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005143.

Singh:2023:BCC

[SS23c] Anuraj Singh and Vijay Shankar Sharma. Bifurcations and chaos control in a discrete-time prey-predator model with Holling type-II functional response and prey refuge. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003375.

Shahni:2024:NSE

[SS24a] Julee Shahni and Randhir Singh. Numerical solution and error analysis of the Thomas–Fermi type equations with integral boundary conditions by the modified collocation techniques. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006441.

Sun:2024:SCL

[SS24b] Tao Sun and Hai-Wei Sun. Stiff-cut leapfrog scheme for fractional Laplacian diffusion equations. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002711.

Sidki:2020:CRO

[SSB20] S. Sidki, R. Sadaka, and A. Benazzouz. Computing recursive orthogonal polynomial with Schur complements. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304091.

Singh:2022:NPT

[SSL22] Swarn Singh, Suruchi Singh, and Zhilin Li. A new patch up technique for elliptic partial differential equation with irregularities. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005732.

Shi:2024:CSE

[SSL24] Hongling Shi, Minghui Song, and Mingzhu Liu. Convergence and stability of an explicit method for nonlinear stochastic differential equations with piecewise continuous arguments. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004934.

Stepanov:2023:PNH

[SSM23] Sergei Stepanov, Denis Spiridonov, and Tina Mai. Prediction of numerical homogenization using deep learning for the Richards equation. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005787.

Singla:2023:EOA

[SSRK23] Rajat Singla, Gurjinder Singh, Higinio Ramos, and V. Kanwar. An efficient optimized adaptive step-size hybrid block method for integrating w′′ = f(t, w, w′) directly. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004368.

Shi:2024:SWG

[SSS24] Lei Shi, Behzad Nemati Saray, and Fazlollah Soleymani. Sparse wavelet Galerkin method: Application for fractional pantograph problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003303.

Spiridonov:2023:OGM

[SSV23] Denis Spiridonov, Sergei Stepanov, and Vasil’ev Vasiliy. An Online Generalized Multiscale finite element method for heat and mass transfer problem with artificial ground freezing. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200276X.

Swati:2020:HOE

[SSVS20] Swati, Karanjeet Singh, Amit K. Verma, and Mandeep Singh. Higher order Emden–Fowler type equations via uniform Haar Wavelet resolution technique. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301278.

Simos:2022:HOR

[ST22] T. E. Simos and Ch. Tsitouras. On high order Runge–Kutta–Nyström pairs. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003757.

Sugrue:2020:IAN

[Sug20]  Darren Sugrue.  Intro-

ducing article numbering to Journal of Computational and Applied Mathematics. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303899.

Sun:2020:NSC

[Sun20] Yan Sun. Necessary and sufficient condition for the existence of positive solutions of a coupled system for reaction–diffusion equations. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303255.

Seemab:2020:ESI

[SuR20] Arjumand Seemab and Mujeeb ur Rehman. Existence of solution of an infinite system of generalized fractional differential equations by Darbo’s fixed point theorem. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303589.

Surkov:2021:RTR

[Sur21] P. G. Surkov. Real-time reconstruction of external impact on fractional order system under measuring a part of coordinates. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303307.

Sabir:2023:NMB

[SuR23] Aneela Sabir and Mujeeb ur Rehman. A numerical method based on quadrature rules for ψfractional differential equations. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003466.

Sathe:2023:FSS

[SUW23] Aastha M. Sathe, Neelesh S. Upadhye, and Agnieszka Wyłomańska. Forecasting of symmetric αstable autoregressive models by time series approach supported by artificial neural networks. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006495.

Schroder:2023:BTM

[SV23] Christian Schröder and Matthias Voigt. Balanced truncation model reduction with a priori error bounds for LTI systems with nonzero initial value. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003582.

Spiridonov:2024:NLM

[SV24] Denis Spiridonov and Maria Vasilyeva. Non-local multi-continuum method (NLMC) for Darcy–Forchheimer flow in fractured media. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005186.

Svacek:2023:NSF

[Svá23] Petr Svácek. On numerical simulation of fluid– structure interaction problems using variational multiscale methods. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000699.

Spiridonov:2020:GMF

[SVC20] Denis Spiridonov, Maria Vasilyeva, and Eric T. Chung. Generalized Multiscale Finite Element method for multicontinua unsaturated flow problems in fractured porous media. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305990.

Shan:2022:DPC

[SvG22] Xiujie Shan and Martin B. van Gijzen. Deflated preconditioned Conjugate Gradient methods for noise filtering of low-field MR images. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003526.

Singh:2024:SES

[SVJG24] Housila P. Singh, Gajendra K. Vishwakarma, Harshada Joshi, and Shubham Gupta. Shrinkage estimation for square of location parameter of the exponential distribution with known coefficient of variation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004338.

Shekhar:2020:OCS

[SVK20] Chandra Shekhar, Shreekant Varshney, and Amit Kumar. Optimal control of a service system with emergency vacation using bat algorithm. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303292.

Spiridonov:2022:MGM

[SVWC22] Denis Spiridonov, Maria Vasilyeva, Min Wang, and Eric T. Chung. Mixed Generalized Multiscale Finite Element Method for flow problem in thin domains. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002849.

Skelton:2020:PRR

[SW20] Andrew Skelton and Allan R. Willms. Parameter range reduction from partial data in systems of differential algebraic equations. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307034.

Shen:2021:TCA

[SW21a] Xiaowen Shen and Qi Wang.

Thermodynamically consistent algorithms for models of incompressible multiphase polymer solutions with a variable mobility. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100193X.

Su:2021:ENP

[SW21b] Meihong Su and Wenjian Wang. Elastic net penalized quantile regression model. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000819.

Stillfjord:2023:SSR

[SW23] Tony Stillfjord and Måns Williamson. SRKCD: a stabilized Runge–Kutta method for stochastic optimization. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002837.

Shen:2024:ESM

[SW24a] Mengxia Shen and Haiyong Wang. An efficient spectral method for the fractional Schrödinger equation on the real line. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000232.

Shi:2024:PAP

[SW24b] Wenya Shi and Gang Wu. Perturbation analysis on PCA plus graph embedding methods and PCA plus exponential graph embedding methods. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000372.

Song:2023:RCG

[SWN23] Guang-Jing Song, Xue-Zhong Wang, and Michael K. Ng. Riemannian conjugate gradient descent method for fixed multi rank third-order tensor completion. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004642.

Shen:2023:ESB

[SWW23] Peiping Shen, Dianxiao Wu, and Yafei Wang. An efficient spatial branchand-bound algorithm using an adaptive branching rule for linear multiplicative programming. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000444.

Shi:2021:FOQ

[SXX21] Wanxia Shi, Xiangtuan Xiong, and Xuemin Xue. A fractional-order quasireversibility method to a backward problem for the time fractional diffusion equation. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001722.

Sarkar:2023:DTA

[SY23] Rahul Sarkar and Theodore J. Yoder. Density theorems with applications in quantum signal processing. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001875.

Sheng:2024:NSP

[SY24] Zhiqiang Sheng and Guangwei Yuan. A nonlinear scheme preserving maximum principle for heterogeneous anisotropic diffusion equation. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003825.

Song:2021:CMC

[SYD21] Huiming Song, Zhanwen Yang, and Teresa Diogo. Collocation methods for cordial Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306129.

Sun:2019:BAC

[SYG19] Zhongyang Sun, Kam Chuen Yuen, and Junyi Guo. A BSDE approach to a class of dependent risk model of mean-variance insurers with stochastic volatility and no-short selling. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304169.

Sim:2024:GEQ

[SYS24] Chol Sim, Nam Yun, and Kwang Sonwu. A Generalized Energy and QUadratic Invariant Preserving (GEQUIP) method for Hamiltonian systems with multiple invariants. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003339.

Song:2024:CFD

[SYTD24] Lina Song, Wang Yu, Yousheng Tan, and Ke Duan. Calculations of fractional derivative option pricing models based on neural network. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004065.

Sendova:2020:MSC

[SZ20] Kristina P. Sendova and Ruixi Zhang. Maximum surplus and R_{n} class of distributions with an application to dividends. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305734.

Soleymani:2023:ESE

[SZ23] Fazlollah Soleymani and Shengfeng Zhu. Error and stability estimates of a time-fractional option pricing model under fully spatial-temporal graded meshes. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000195.

Shi:2024:SAA

[SZ24] Dongyang Shi and Sihui Zhang. Superconvergence analysis of an anisotropic nonconforming FEM for parabolic equation. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002292.

Soradi-Zeid:2024:CMU

[SZA24] Samaneh Soradi-Zeid and Maryam Alipour. A collocation method using generalized Laguerre polynomials for solving nonlinear optimal control problems governed by integro-differential equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003540.

Shaydurov:2020:ATD

[SZK20] V. Shaydurov, S. Zhang, and V. Kornienko. Approximations of two-dimensional Mean Field Games with nonsymmetric controls. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304649.

Song:2020:RBC

[SZL20] Yunan Song, Fengrui Zhang, and Congchong Liu. The risk of block chain financial market based on particle swarm optimization. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306727.

Shao:2022:EMT

[SZL22] Wei Shao, Yijun Zuo, and June Luo. Employing the MCMC technique to compute the projection depth in high dimensions. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001157.

Sheng:2021:RRM

[SZWH21] Chunguang Sheng, Degang Zhang, Guangyu Wang, and Yingli Huang. Research on risk mechanism of China’s carbon financial market development from the perspective of ecological civilization. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302818.

Shen:2021:FGA

[SZZ21] Yuan Shen, Yannian Zuo, and Xiayang Zhang. A faster generalized ADMMbased algorithm using a sequential updating scheme with relaxed step sizes for multiple-block linearly constrained separable convex programming. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001229.

Sun:2023:HBB

[SZZM23] Zhong-Feng Sun, Jin-Chuan Zhou, Yun-Bin Zhao, and Nan Meng. Heavy-ball-based hard thresholding algorithms for sparse signal recovery. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300208X.

Toomaj:2022:SNF

[TA22] Abdolsaeed Toomaj and Habibollah Agh Atabay. Some new findings on the cumulative residual Tsallis entropy. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002910.

Tak:2020:TPF

[Tak20] Nihat Tak. Type-1 possibilistic fuzzy forecasting functions. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306582.

Takhirov:2024:QSS

[Tak24]  Aziz Takhirov.  Quad-

SAV scheme for gradient systems. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000177.

Tamminen:2020:DTV

[Tam20] Janne P. Tamminen. Detection of time-varying heat sources using an analytic forward model. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300923.

Tanaka:2020:NVM

[Tan20] Kazuaki Tanaka. Numerical verification method for positive solutions of elliptic problems. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306521.

Taghikhani:2022:GIT

[TBA22] Sepideh Taghikhani, Fahimeh Baroughi, and Behrooz Alizadeh. A generalized interval type-2 fuzzy random variable based algorithm under mean chance value at risk criterion for inverse 1-median location problems on tree networks with uncertain costs. Journal of Computational and Applied Mathematics, 408 (??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000127.

Taghikhani:2024:MVV

[TBA24] Sepideh Taghikhani, Fahimeh Baroughi, and Behrooz Alizadeh. Mean-variance value at risk criterion for solving a p-median location problem on networks with type-2 intuitionistic fuzzy weights. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004259.

Tuan:2020:FVP

[TBT^{+}20] Nguyen Huy Tuan, Dumitru Baleanu, Tran Ngoc Thach, Donal O’Regan, and Nguyen Huu Can. Final value problem for nonlinear time fractional reaction–diffusion equation with discrete data. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301746.

Tian:2024:MPP

[TCGY24] Lulu Tian, Nattaporn Chuenjarern, Hui Guo, and Yang Yang. Maximumprinciple-preserving highorder discontinuous Galerkin methods for incompressible Euler equations on overlapping meshes. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003527.

Tyrylgin:2021:MMR

[TCVC21] Aleksei Tyrylgin, Yaoyao Chen, Maria Vasilyeva, and Eric T. Chung. Multiscale model reduction for the Allen–Cahn problem in perforated domains. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303010.

Thong:2023:SPA

[TDAT23] Duong Viet Thong, Vu Tien Dung, Pham Ky Anh, and Hoang Van Thang. A single projection algorithm with double inertial extrapolation steps for solving pseudomonotone variational inequalities in Hilbert space. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000432.

Talaska:2023:NHI

[TDD23] Tomasz Talaśka, Zofia Długosz, and Rafał Długosz. A novel hardware implemented programmable controller adapted to cooperate with AI tuning algorithms in real time systems. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001061.

Thong:2022:TFC

[TDL^{+}22] Duong Viet Thong, Qiao-Li Dong, Lu-Lu Liu, Nguyen Anh Triet, and Nguyen Phuong Lan. Two fast converging inertial subgradient extragradient algorithms with variable stepsizes for solving pseudo-monotone VIPs in Hilbert spaces. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001066.

Tank:2020:BCS

[TE20] Fatih Tank and Serkan Eryilmaz. On bivariate compound sums. Journal of Computational and Applied Mathematics, 365 (??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0377042719303747.

Tank:2024:SIC

[TE24] Fatih Tank and Serkan Eryilmaz. Special issue: Computational methods in system reliability. Journal of Computational and Applied Mathematics, 435 (??):??, ???? 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003242.

Terekhov:2020:CCFb

[Ter20] Kirill M. Terekhov. Cellcentered finite-volume method for heterogeneous anisotropic poromechanics problem. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303607.

Tourani-Farani:2023:PMT

[TFK23] Fahimeh Tourani-Farani and Iraj Kazemi. Pragmatic model transformations for analyzing bounded and positive responses. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003454.

Tamborrino:2024:EDE

[TFM24] Cristiano Tamborrino, Antonella Falini, and Francesca Mazzia. Empirical density estimation based on spline quasi-interpolation with applications to copulas clustering modeling. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003807.

Titarev:2020:CSK

[TFR^{+}20] V. A. Titarev, A. A.

Frolova, V. A. Rykov, P. V. Vashchenkov, A. A. Shevyrin, and Ye. A. Bondar. Comparison of the Shakhov kinetic equation and DSMC method as applied to space vehicle aerothermodynamics. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303577.

Takaki:2022:SCU

[TFR22] Nick Takaki, G. W. Forbes, and Jannick P. Rolland. Schemes for cubature over the unit disk found via numerical optimization. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006324.

Titi:2021:BRC

[TG21] Jihad Titi and Jürgen Garloff. Bounds for the range of a complex polynomial over a rectangular region. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306683.

Tian:2022:SAE

[TGJY22] Lulu Tian, Hui Guo, Rui Jia, and Yang Yang. Stability analysis and error estimates of fullydiscrete local discontinuous Galerkin methods for simulating wormhole propagation with Darcy– Forchheimer model. Journal of Computational and Applied Mathematics, 409 (??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722000437.

Thipwiwatpotjana:2021:STT

[TGLa21] Phantipa Thipwiwatpotjana, Artur Gorka, and Worrawate Leela-apiradee. Solution types of twosided interval linear system and their application on interval linear programming problems. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305859.

Tranquilli:2021:SAR

[TGS21] Paul Tranquilli, Ross Glandon, and Adrian Sandu. Subspace adaptivity in Rosenbrock–Krylov methods for the time integration of initial value problems. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304799.

Tang:2024:FPI

[TH24a] Shi-Ping Tang and Yu-Mei Huang. A fast preconditioning iterative method for solving the discretized second-order space-fractional advection– diffusion equations. Journal of Computational and Applied Mathematics, 438 (??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723004570.

Tuyen:2024:ESA

[TH24b] Truong Minh Tuyen and Nguyen Song Ha. An efficient self-adaptive algorithm for finding common solutions to pseudomonotone variational inequalities and split fixed point problems in Hilbert spaces. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002504.

Tian:2024:ESF

[THS24] Jia Tian, Mingyan He, and Pengtao Sun. Energystable finite element method for a class of nonlinear fourth-order parabolic equations. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005204.

Tian:2022:SFI

[Tia22]  Yan Tian.  Supercon-

vergence and fast implementation of the barycentric prolate differentiation. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000644.

Tian:2024:NCR

[Tia24]  Long Tian.  Next ca-

reer recommendation in Mississippi with artificial intelligence. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004028.

Tajrishi:2024:FFT

[TK24a] Mohammad Amin Zahedi Tajrishi and Ali Akbarzadeh Kalat. Fast finite time fractional-order robust-adaptive sliding mode control of nonlinear systems with unknown dynamics. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004983.

Tiwari:2024:MFP

[TK24b] Sudarshan Tiwari and Axel Klar. A mesh-free particle method for continuum modelling of granular flow. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004739.

Tamilalagan:2021:DFO

[TKM21] P. Tamilalagan, S. Karthiga, and P. Manivannan. Dynamics of fractional order HIV infection model with antibody and cytotoxic T-lymphocyte immune responses. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303551.

Theodosis:2024:NHE

[TKT^{+}24] Dionysios Theodosis, Iasson Karafyllis, George Titakis, Ioannis Papamichail, and Markos Papageorgiou. A nonlinear heat equation arising from automatedvehicle traffic flow models. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003874.

Tuan:2021:IQR

[TKVA21] Nguyen Huy Tuan, Vo Anh Khoa, Phan Thi Khanh Van, and Vo Van Au. An improved quasi-reversibility method for a terminalboundary value multispecies model with white Gaussian noise. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304672.

Tan:2020:ENS

[TL20] Jinying Tan and Jiangguo Liu. An efficient numerical solver for anisotropic subdiffusion problems. Journal of Computational and Applied Mathematics, 364 (??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0377042719303152.

Truhar:2021:EDN

[TL21] Ninoslav Truhar and Ren-Cang Li. On an eigenvectordependent nonlinear eigenvalue problem from the perspective of relative perturbation theory. Journal of Computational and Applied Mathematics, 395 (??):??, October 15, 2021.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721002168.

Tan:2022:SAS

[TL22] Jinying Tan and Jiangguo Liu. Solving anisotropic subdiffusion problems in annuli and shells. Journal of Computational and Applied Mathematics, 401 (??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042721003861.

Thong:2023:FRI

[TLD^{+}23] Duong Viet Thong, Lu-Lu Liu, Qiao-Li Dong, Luong Van Long, and Pham Anh Tuan. Fast relaxed inertial Tseng’s method-based algorithm for solving variational inequality and fixed point problems in Hilbert spaces. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003739.

Tang:2023:EPB

[TLY23] Yaozong Tang, Gang Luo, and Qingzhi Yang. An efficient PGM-based algorithm with backtracking strategy for solving quadratic optimization problems with spherical constraint. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005131.

Tang:2024:MTE

[TM24] Yiyi Tang and Xuerong Mao. The modified truncated Euler–Maruyama method for stochastic differential equations with concave diffusion coefficients. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006040.

Tong:2024:SAL

[TN24] Hongzhi Tong and Michael Ng. Spectral algorithms for learning with dependent observations. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003813.

Tomanovic:2024:GTQ

[Tom24] Jelena Tomanović. Gausstype quadrature rules for variable-sign weight functions. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004211.

Torrado:2021:APS

[Tor21] Nuria Torrado. On allocation policies in systems with dependence structure and random selection of components. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305653.

Thiele:2022:MPS

[TR22] Christopher Thiele and Beatrice Riviere. pmultigrid with partial smoothing: an efficient preconditioner for discontinuous Galerkin discretizations with modal bases. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004374.

Tocino:2020:MSN

[TS20] A. Tocino and M. J. Senosiain. MS-stability of nonnormal stochastic differential systems. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302417.

Tezer-Sezgin:2023:SFS

[TSA23] M. Tezer-Sezgin and S. H.

Aydin. Stabilized FEM solution of MHD duct flow with conducting cracks in the insulation. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005349.

Tamimi:2022:MNG

[TSG22] H. Tamimi, S. Saiedinezhad, and M. B. Ghaemi. The measure of noncompactness in a generalized coupled fixed point theorem and its application to an integro-differential system. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001704.

Shang:2024:SSB

[tSjTsJ24] Tong tong Shang, Guo ji Tang, and Wen sheng Jia. Solution set bounds for LCPs over tensor spaces. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003278.

Tavares:2021:SIP

[TSL^{+}21] Camila A. Tavares, Taináh M. R. Santos, Nelson H. T. Lemes, José P. C. dos Santos, José C. Ferreira, and João P. Braga. Solving ill-posed problems faster using fractionalorder Hopfield neural network. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302752.

Tatsuoka:2020:ACM

[TSMZ20] Fuminori Tatsuoka, Tomohiro Sogabe, Yuto Miyatake, and Shao-Liang Zhang. Algorithms for the computation of the matrix logarithm based on the double exponential formula. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303991.

Terekhov:2020:CCFa

[TT20] Kirill M. Terekhov and Hamdi A. Tchelepi. Cellcentered finite-volume method for elastic deformation of heterogeneous media with full-tensor properties. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303280.

Toan:2023:SAP

[TTK23] N. T. Toan, L. Q. Thuy, and D. S. Kim. Sensitivity analysis in parametric multiobjective discrete-time control via Fréchet subdifferential calculus of the frontier map. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003351.

Tubikanec:2022:QPD

[TTLB22] Irene Tubikanec, Massimiliano Tamborrino, Petr Lansky, and Evelyn Buckwar. Qualitative properties of different numerical methods for the inhomogeneous geometric Brownian motion. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005616.

Thong:2022:RLC

[TV22] Duong Viet Thong and Phan Tu Vuong. R-linear convergence analysis of inertial extragradient algorithms for strongly pseudomonotone variational inequalities. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005926.

Themistoclakis:2024:GFH

[TV24] Woula Themistoclakis and Marc Van Barel. A generalization of floater-hormann interpolants. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300626X.

Tyrylgin:2023:CMM

[TVAS23] Aleksei Tyrylgin, Maria Vasilyeva, Anatoly Alikhanov, and Dongwoo Sheen. A computational macroscale model for the time fractional poroelasticity problem in fractured and heterogeneous media. Journal of Computational and Applied Mathematics, 418 (??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042722003399.

Thomas:2021:JML

[TVB21] Abin Thomas, Gajendra K.

Vishwakarma, and Atanu Bhattacharjee. Joint modeling of longitudinal and time-to-event data on multivariate protein biomarkers. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303071.

Tyrylgin:2020:GMF

[TVSC20] Aleksei Tyrylgin, Maria Vasilyeva, Denis Spiridonov, and Eric T. Chung. Generalized Multiscale Finite Element Method for the poroelasticity problem in multicontinuum media. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300741.

Tong:2023:MMB

[TWQ^{+}23] Can Tong, Jiao Wei, Shouliang Qi, Yudong Yao, Tie Zhang, and Yueyang Teng. A majorization– minimization based solution to penalized nonnegative matrix factorization with orthogonal regularization. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004757.

Tao:2021:DGM

[TXS21] Qi Tao, Yan Xu, and Chi-Wang Shu. A discontinuous Galerkin method and its error estimate for nonlinear fourth-order wave equations. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305215.

Tang:2024:PEE

[TXZ24] Ming Tang, Xiaoqing Xing, and Liuqiang Zhong. A posteriori error estimator for mixed interior penalty discontinuous Galerkin finite element method for the H (curl)-elliptic problems. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003515.

Toraman:2023:SGA

[TY23] Sitki Can Toraman and Hamdullah Yücel. A stochastic gradient algorithm with momentum terms for optimal control problems governed by a convection–diffusion equation with random diffusivity. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005179.

Tuan:2020:WPI

[TZ20] Nguyen Huy Tuan and Yong Zhou. Well-posedness of an initial value problem for fractional diffusion equation with Caputo– Fabrizio derivative. Journal of Computational and Applied Mathematics, 375 (??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720301023.

Taherifar:2021:ADQ

[TZBK21] Reza Taherifar, Seyed Alireza Zareei, Mahmood Rabani Bidgoli, and Reza Kolahchi. Application of differential quadrature and Newmark methods for dynamic response in pad concrete foundation covered by piezoelectric layer. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303666.

Tuan:2020:ASN

[TZTC20] Nguyen Huy Tuan, Yong Zhou, Tran Ngoc Thach, and Nguyen Huu Can. An approximate solution for a nonlinear biharmonic equation with discrete random data. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300029.

Tian:2021:SRI

[TZWG21] Zhaolu Tian, Yan Zhang, Junxin Wang, and Chuanqing Gu. Several relaxed iteration methods for computing PageRank. Journal of Computational and Applied Mathematics, 388 (??):??, May 1, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305860.

Ucar:2023:AHD

[Uça23] Sümeyra Uçar. Analysis of hepatitis B disease with fractal-fractional Caputo derivative using real data from Turkey. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003508.

Usevich:2020:DMP

[UDI20] Konstantin Usevich, Philippe Dreesen, and Mariya Ishteva. Decoupling multivariate polynomials: Interconnections between tensorizations. Journal of Computational and Applied Mathematics, 363(??):22–34, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930161X.

Zaheer-ud-Din:2022:MPH

[uDuIZ22] Zaheer ud Din, Siraj ul Islam, and Sakhi Zaman. Meshless procedure for highly oscillatory kernel based one-dimensional Volterra integral equations. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001601.

Ugalde:2024:CMT

[Uga24] Esteban Segura Ugalde.

Computation of matrix pth roots using moments. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003133.

Urena:2020:SSO

[UGG^{+}20] F. Ureña, L. Gavete, A. García, J. J. Benito, and A. M. Vargas. Solving second order nonlinear hyperbolic PDEs using generalized finite difference method (GFDM). Journal of Computational and Applied Mathematics, 363(??):1–21, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302638.

Utyuzhnikov:2023:DDN

[UL23] Sergey V. Utyuzhnikov and Hongru Li. Domain decomposition with nonlocal interface boundary conditions. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004459.

Ucar:2023:QAN

[UÖA23] Esmehan Uçar, Necati Özdemir, and Eren Altun. Qualitative analysis and numerical simulations of new model describing cancer. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004976.

Upadhyay:2024:QNA

[UPPZ24] B. B. Upadhyay, Rupesh K. Pandey, Jinlan Pan, and Shengda Zeng. Quasi-Newton algorithms for solving intervalvalued multiobjective optimization problems by using their certain equivalence. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004946.

Usta:2021:NAF

[Ust21]  Fuat Usta.  Numeri-

cal analysis of fractional Volterra integral equations via Bernstein approximation method. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304891.

Ustaoglu:2023:IIR

[Ust23] Zekeriya Ustaoglu. Iterative inversion of Radon transform via discretization by fuzzy basic functions. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001851.

Upston:2023:CCC

[USTG23] J. Upston, D. Sulsky, J. D.

Tucker, and Y. Guan. CIEL^{∗} Ch color map for visualization and analysis of sea ice motion. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000705.

Urena:2024:NSP

[UV24] N. Ureña and A. M. Vargas. Numerical solution to a parabolic-ODE Solow model with spatial diffusion and technologyinduced motility. Journal of Computational and Applied Mathematics, 447 (??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724001638.

Unvan:2020:DBS

[ÜY20] Yüksel Akay Ünvan and Ibrahim Nandom Yakubu. Do bank-specific factors drive bank deposits in Ghana? Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301187.

Vigo-Aguiar:2022:CMM

[VAAR22] Jesús Vigo-Aguiar, Pedro Alonso, and Higinio Ramos. Editorial: Computational and mathematical methods in science and engineering. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000868.

Vabishchevich:2021:ARS

[Vab21] P. N. Vabishchevich. An approximate representation of a solution to fractional elliptical BVP via solution of parabolic IVP. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000790.

Vabishchevich:2022:SMS

[Vab22] Petr N. Vabishchevich.

Some methods for solving equations with an operator function and applications for problems with a fractional power of an operator. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000085.

Vabishchevich:2023:ASC

[Vab23] P. N. Vabishchevich. Approximate solution of the Cauchy problem for a first-order integrodifferential equation with solution derivative memory. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200485X.

Vabishchevich:2024:CDC

[Vab24] P. N. Vabishchevich. Computational decomposition and composition technique for approximate solution of nonstationary problems. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003601.

Venturino:2024:HON

[VAMO24] Ezio Venturino, Sebastian Anita, Domenico Mezzanotte, and Donatella Occorsio. A high order numerical scheme for a nonlinear nonlocal reaction–diffusion model arising in population theory. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003315.

Vishwakarma:2024:CIT

[VBRT24] Gajendra K. Vishwakarma, Atanu Bhattacherjee, Bhrigu K. Rajbongshi, and Abhipsa Tripathy. Censored imputation of time to event outcome through survival proximity score method. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003522.

Vasconcelos:2020:CTP

[VCNR20] Josimar M. Vasconcelos, Renato J. Cintra, Abraão D. C. Nascimento, and Leandro C. Rêgo. The compound truncated Poisson Cauchy model: a descriptor for multimodal data. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301783.

vandenBos:2021:GIA

[vdBS21] L. M. M. van den Bos and B. Sanderse. A geometrical interpretation of the addition of nodes to an interpolatory quadrature rule while preserving positive weights. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000492.

Vondrejc:2020:EBC

[VdG20] Jaroslav Vondrejc and Tom W. J. de Geus. Energybased comparison between the Fourier–Galerkin method and the finite element method. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305904.

vanderMeer:2022:OWL

[vdMOB22] Remco van der Meer, Cornelis W. Oosterlee, and Anastasia Borovykh. Optimally weighted loss functions for solving PDEs with neural networks. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005100.

Veronese:2020:OPQ

[Ver20] Daniel O. Veronese. Orthogonal polynomials and quadrature rules on the unit circle associated with perturbations of symmetric measures. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300996.

Vieira:2021:HCN

[Vie21] R. S. Vieira. How to count the number of zeros that a polynomial has on the unit circle? Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030460X.

Villarino:2023:BSP

[VLR23] Joel P. Villarino, Álvaro Leitao, and J. A. García Rodríguez. Boundarysafe PINNs extension: Application to non-linear parabolic PDEs in counterparty credit risk. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722006392.

Vasilyeva:2023:MMR

[VMP23] Maria Vasilyeva, S. M. Mallikarjunaiah, and D. Palaniappan. Multiscale model reduction technique for fluid flows with heterogeneous porous inclusions. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200574X.

Vynnycky:2022:FCL

[VNA22] M. Vynnycky, A. S. Nick, and M. Assunção. Fast computation of the Lorentz force induced by longitudinal electromagnetic stirring. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002783.

Vandenplas:2023:DTA

[VNV23] Jeremie Vandenplas, Buu-Van Nguyen, and Cornelis Vuik. Deflation techniques applied on mixed model equations. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000390.

Vrablikova:2024:IPR

[VOJS24] Jana Vráblíková, Vanessa Ortler, Bert Jüttler, and Zbynek Sír. Interpolation by polygon rolling motions for approximate sweep computation. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001407.

Valentim:2020:CFC

[VORD20] Carlos A. Valentim, Naila A.

Oliveira, José A. Rabi, and Sergio A. David. Can fractional calculus help improve tumor growth models? Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302557.

Vohra:2024:RCS

[VP24] Naren Vohra and Malgorzata Peszynska. Robust conservative scheme and nonlinear solver for phase transitions in heterogeneous permafrost. Journal of Computational and Applied Mathematics, 442 (??):??, May 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006623.

Vishwakarma:2023:ARA

[VPHE23] Gajendra K. Vishwakarma, Chinmoy Paul, Ali S. Hadi, and A. M. Elsawah. An automated robust algorithm for clustering multivariate data. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001632.

Vishwakarma:2021:CEM

[VS21] Gajendra K. Vishwakarma and Neha Singh. Computing the effect of measurement errors under additive scramble response of the sensitive variable. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002132.

Valasek:2023:AFC

[VS23] Jan Valásek and Petr Svácek. On aerodynamic force computation in fluid-structure interaction problems — comparison of different approaches. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001528.

Vasilyeva:2023:MSM

[VSP23] Maria Vasilyeva, Alexey Sadovski, and D. Palaniappan. Multiscale solver for multi-component reaction– diffusion systems in heterogeneous media. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000948.

Vasilyeva:2020:MFE

[VSSV20] Maria Vasilyeva, Sergei Stepanov, Denis Spiridonov, and Vasiliy Vasil’ev. Multiscale Finite Element Method for heat transfer problem during artificial ground freezing. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306107.

Vishwakarma:2021:WFM

[VTB21] Gajendra K. Vishwakarma, Abin Thomas, and Atanu Bhattacharjee. A weight function method for selection of proteins to predict an outcome using protein expression data. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000844.

Vasilyeva:2021:PMC

[VTBM21] Maria Vasilyeva, Aleksei Tyrylgin, Donald L. Brown, and Anirban Mondal. Preconditioning Markov Chain Monte Carlo Method for Geomechanical Subsidence using multiscale method and machine learning technique. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100039X.

Wout:2024:BEM

[vtW24] Elwin van ’t Wout. The boundary element method for acoustic transmission with nonconforming grids. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000876.

Vanderstraeten:2023:CFO

[VV23] Emile Vanderstraeten and Dries Vande Ginste. A conservative fourth-order real space method for the (2+1)D Dirac equation. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000936.

Vlasov:2021:AAH

[VVB21] A. N. Vlasov and D. B. Volkov-Bogorodsky. Application of the asymptotic homogenization in a parametric space to the modeling of structurally heterogeneous materials. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304829.

Vavpetic:2021:OPG

[VZ21] Ales Vavpetic and Emil Zagar. On optimal polynomial geometric interpolation of circular arcs according to the Hausdorff distance. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001102.

Wieloch:2021:BIC

[WA21] Victoria Wieloch and Martin Arnold. BDF integrators for constrained mechanical systems on Lie groups. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305229.

Wu:2020:SBF

[WAAM20] Wei-Tao Wu, Nadine Aubry, James F. Antaki, and Mehrdad Massoudi. Simulation of blood flow in a sudden expansion channel and a coronary artery. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301473.

Waziri:2022:MPT

[WAH22] Mohammed Yusuf Waziri, Kabiru Ahmed, and Abubakar Sani Halilu. A modified PRPtype conjugate gradient projection algorithm for solving large-scale monotone nonlinear equations with convex constraint. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006087.

Wang:2020:NPD

[Wan20a] Chunmei Wang. A new primal-dual weak Galerkin finite element method for ill-posed elliptic Cauchy problems. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930634X.

Wang:2020:BRT

[Wan20b] Wen-Kai Wang. Building recombining trinomial trees for time-homogeneous diffusion processes. Journal of Computational and Applied Mathematics, 364 (??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0377042719303541.

Wang:2021:EEP

[Wan21a] Bin Wang. Exponential energy-preserving methods for charged-particle dynamics in a strong and constant magnetic field. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306223.

Wang:2021:AVV

[Wan21b] Xingchun Wang. Analytical valuation of vulnerable European and Asian options in intensity-based models. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000315.

Wang:2024:PRE

[Wan24]  Yayun Wang.  Pricing

ratchet equity index annuity with mortality risk by complex Fourier series method. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003588.

Wang:2021:NAV

[WC21] Xilu Wang and Xiaoliang Cheng. Numerical analysis of a viscoplastic contact problem with normal compliance, unilateral constraint, memory term and friction. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002685.

Wang:2024:UES

[WC24a] Hao Wang and Yaoyao Chen. Unconditionally energy stable invariant energy quadratization finite element methods for Phase-Field Crystal equation and Swift–Hohenberg equation. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002462.

Wang:2024:EIS

[WC24b] Yibo Wang and Wanrong Cao. Exponential integrator for stochastic strongly damped wave equation based on the Wong–Zakai approximation. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272300403X.

Wang:2019:GMM

[WCC^{+}19] Min Wang, Siu Wun Cheung, Eric T. Chung, Maria Vasilyeva, and Yuhe Wang. Generalized multiscale multicontinuum model for fractured vuggy carbonate reservoirs. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303735.

Wang:2021:MGL

[WCFH21] Xia Wang, Eric Chung, Shubin Fu, and Zhaoqin Huang. Mixed GMs-FEM for linear poroelasticity problems in heterogeneous porous media. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000029.

Wang:2020:TLO

[WCK20] Junxian Wang, Eric Chung, and Hyea Hyun Kim. A two-level overlapping Schwarz method with energy-minimizing multiscale coarse basis functions. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306053.

Washio:2022:UIC

[WCK^{+}22] Takumi Washio, Xiaoke Cui, Ryo Kanada, Jun ichi Okada, Seiryo Sugiura, Yasushi Okuno, Shoji Takada, and Toshiaki Hisada. Using incomplete Cholesky factorization to increase the time step in molecular dynamics simulations. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002552.

Wang:2022:CDH

[WCL22] Yuchan Wang, Qun Chen, and Jijun Liu. On the cavity detection in a heat conductive medium from time-average boundary temperature measurement. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004027.

Wang:2020:GSA

[WCLW20] Xiuli Wang, Zhimiao Cao, Chao Liu, and Mingqiu Wang. Group selection via adjusted weighted least absolute deviation regression. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302156.

Wang:2023:DFE

[WCLY23] Ziqiang Wang, Chuanjun Chen, Yanjun Li, and Xiaofeng Yang. Decoupled finite element scheme of the variable-density and viscosity phase-field model of a two-phase incompressible fluid flow system using the volume-conserved Allen– Cahn dynamics. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003934.

Wang:2023:SSN

[WCMW23] Xuezhong Wang, Maolin Che, Changxin Mo, and Yimin Wei. Solving the system of nonsingular tensor equations via randomized Kaczmarz-like method. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272200454X.

Wang:2020:ORC

[WCPW20] Shufen Wang, Wenbin Chen, Hanshuang Pan, and Cheng Wang. Optimal rate convergence analysis of a second order scheme for a thin film model with slope selection. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301461.

Wang:2020:NNA

[WCW20] Xuezhong Wang, Maolin Che, and Yimin Wei. Neural network approach for solving nonsingular multi-linear tensor systems. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305746.

Wang:2020:CIP

[WCWZ20] Chunyan Wang, Haibin Chen, Yiju Wang, and Guanglu Zhou. On copositiveness identification of partially symmetric rectangular tensors. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306831.

Wu:2020:WBE

[WD20] Liang Wu and Yiming Ding. Wavelet-based estimations of fractional Brownian sheet: least squares versus maximum likelihood. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306144.

Wang:2024:NVT

[WDPY24] Xuejie Wang, Kangkang Deng, Zheng Peng, and Chengcheng Yan. New vector transport operators extending a Riemannian CG algorithm to generalized Stiefel manifold with low-rank applications. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002747.

Wang:2022:NCR

[WDS22] Chunfeng Wang, Yaping Deng, and Peiping Shen. A novel convex relaxationstrategy-based algorithm for solving linear multiplicative problems. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721006348.

Wang:2020:RIM

[WDTW20] Liang Wang, Sanku Dey, Yogesh Mani Tripathi, and Shuo-Jye Wu. Reliability inference for a multicomponent stress-strength model based on Kumaraswamy distribution. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030114X.

Wu:2020:IIB

[WDY20] Y. H. Wu, C. Y. Dong, and H. S. Yang. Isogeometric indirect boundary element method for solving the 3D acoustic problems. Journal of Computational and Applied Mathematics, 363(??):273–299, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302869.

Wang:2020:NAS

[WEEA20] Y. Wang, S. S. Ezz-Eldien, and A. A. Aldraiweesh. A new algorithm for the solution of nonlinear two-dimensional Volterra integro-differential equations of high-order. Journal of Computational and Applied Mathematics, 364 (??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0377042719302900.

Wen:2021:MKM

[Wen21a]  Dao-Jun Wen.  Mod-

ified Krasnoselski–Mann type iterative algorithm with strong convergence for hierarchical fixed point problem and split monotone variational inclusions. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001205.

Weng:2021:SLS

[Wen21b] Peter Chang-Yi Weng. Solving large-scale nonsymmetric algebraic Riccati equations from twodimensional transport models by doubling. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000662.

Wu:2022:GDA

[WFZ22] Peng Wu, Zhaosheng Feng, and Xuebing Zhang. Global dynamics of an age-space structured HIV/ AIDS model with viral load-dependent infection and conversion rates. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001315.

Wang:2022:CDG

[WGC22] Yue Wang, Fuzheng Gao, and Jintao Cui. A conforming discontinuous Galerkin finite element method for elliptic interface problems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001236.

Wandelt:2021:GIL

[WGM21] Michèle Wandelt, Michael Günther, and Michelle Muniz. Geometric integration on Lie groups using the Cayley transform with focus on Lattice QCD. Journal of Computational and Applied Mathematics, 387 (??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304984.

Wang:2020:WGF

[WGS20] Xiuping Wang, Fuzheng Gao, and Zhengjia Sun. Weak Galerkin finite element method for viscoelastic wave equations. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301072.

Wen:2024:MRK

[WH24] Jiao Wen and Chengming Huang. Multistep Runge– Kutta methods for Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300328X.

Wu:2020:NFD

[WHC^{+}20] Cheng-Feng Wu, Shian-Chang Huang, Tsangyao Chang, Chei-Chang Chiou, and Hsin-Pei Hsueh. The nexus of financial development and economic growth across major Asian economies: evidence from bootstrap ARDL testing and machine learning approach. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930665X.

Wang:2023:RSU

[WHlZH23] Si Wang, Ting-Zhu Huang, Xi le Zhao, and Jie Huang. Reweighted sparse unmixing for hyperspectral images with noise level estimation. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004411.

Wen:2021:APG

[WHY^{+}21] Chun Wen, Qian-Ying Hu, Guo-Jian Yin, Xian-Ming Gu, and Zhao-Li Shen. An adaptive Power– Arnoldi algorithm for computing PageRank. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305008.

Wang:2024:MMT

[WHZ24] Qi Wang, Jingyan Hu, and Guanyu Zhou. The mixed method with two Lagrange multiplier formulations for the Signorini problem. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003649.

Wang:2023:PCD

[WJ23] Hongxing Wang and Tianhe Jiang. Properties and characterizations of dual sharp orders. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002650.

Wang:2024:SVD

[WJVG24] Gang Wang, Tongsong Jiang, V. I. Vasil’ev, and Zhenwei Guo. On singular value decomposition for split quaternion matrices and applications in split quaternionic mechanics. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003916.

Wang:2020:NMT

[WK20a] Hong Wang and Hongchao Kang. Numerical methods for two classes of singularly oscillatory Bessel transforms and their error analysis. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306090.

Wang:2020:ANS

[WK20b] Wanting Wang and Muhammad Altaf Khan. Analysis and numerical simulation of fractional model of bank data with fractal–fractional Atangana–Baleanu derivative. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930651X.

Watson:2020:RBF

[WKC20] Daniel W. Watson, Andreas Karageorghis, and C. S. Chen. The radial basis function-differential quadrature method for elliptic problems in annular domains. Journal of Computational and Applied Mathematics, 363(??):53– 76, January 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0377042719302626.

Weber:2022:SAH

[WKN22] Ivy Weber, Gunilla Kreiss, and Murtazo Nazarov. Stability analysis of high order methods for the wave equation. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005239.

Wang:2021:MDD

[WL21] Jin-Liang Wang and Hui-Feng Li. Memory-dependent derivative versus fractional derivative (I): Difference in temporal modeling. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302144.

Wang:2022:PBS

[WL22] Na-Na Wang and Ji-Cheng Li. On parameterized block symmetric positive definite preconditioners for a class of block three-by-three saddle point problems. Journal of Computational and Applied Mathematics, 405 (??):??, May 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005653.

Wang:2024:RMB

[WL24a] Dan Wang and Jicheng Li. Relaxation modulusbased matrix splitting iteration method for vertical linear complementarity problem. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003746.

Wang:2024:NNA

[WL24b] Yating Wang and Liu Liu.

On a neural network approach for solving potential control problem of the semiclassical Schrödinger equation. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300448X.

Wei:2021:MFB

[WLB^{+}21] Yan-Qiao Wei, Da-Yan Liu, Driss Boutat, Hao-Ran Liu, and Chunwan Lv. Modulating functions based differentiator of the pseudo-state for a class of fractional order linear systems. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304520.

Wang:2022:NKA

[WLBG22] Qifeng Wang, Weiguo Li, Wendi Bao, and Xingqi Gao. Nonlinear Kaczmarz algorithms and their convergence. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003423.

Wen:2023:AMC

[WLC23] You-Wei Wen, Kexin Li, and Hefeng Chen. Accelerated matrix completion algorithm using continuation strategy and randomized SVD. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001590.

Wang:2023:IAS

[WLCH23] Qingsong Wang, Zehui Liu, Chunfeng Cui, and Deren Han. Inertial accelerated SGD algorithms for solving large-scale lowerrank tensor CP decomposition problems. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005465.

Wei:2024:NSP

[WLDZ24] Anli Wei, Ying Li, Wenxv Ding, and Jianli Zhao. A new structure-preserving algorithm based on the semi-tensor product of matrices for split quaternion matrix LDU decomposition and its applications. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003224.

Wang:2023:MNA

[WLH23] Zhifeng Wang, Wei Liu, and Jian Huang. Modeling and numerical analysis of compressible Darcy– Brinkman fluid flow in fractured media with finite volume method on nonmatching grids. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003946.

Wang:2020:ADG

[WLHJ20] Fei Wang, Min Ling, Weimin Han, and Feifei Jing. Adaptive discontinuous Galerkin methods for solving an incompressible Stokes flow problem with slip boundary condition of frictional type. Journal of Computational and Applied Mathematics, 371(??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719307058.

Wang:2024:NDEb

[WLJ24] Wansheng Wang, Jinping Li, and Chengyu Jin. Nonsmooth data error estimates for fully discrete finite element approximations of semilinear parabolic equations in Banach space. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001894.

Wang:2021:HOA

[WLL21] Xiang Wang, Jessica Li, and Jichun Li. High order approximation of derivatives with applications to pricing of financial derivatives. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002971.

Wang:2024:UFV

[WLM24] Yue Wang, Yonghai Li, and Xiangyun Meng. An upwind finite volume element method on a Shishkin mesh for singularly perturbed convection–diffusion problems. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004375.

Wu:2024:NFV

[WLS24] Dan Wu, Junliang Lv, and Zhiqiang Sheng. A nonlinear finite volume element method preserving the discrete maximum principle for heterogeneous anisotropic diffusion equations. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002735.

Wen:2023:NSI

[WLW23] Jin Wen, Zhuan-Xia Liu, and Shan-Shan Wang. A non-stationary iterative Tikhonov regularization method for simultaneous inversion in a timefractional diffusion equation. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000389.

Wu:2022:KMB

[WLZ22] Nian-Ci Wu, Cheng-Zhi Liu, and Qian Zuo. On the Kaczmarz methods based on relaxed greedy selection for solving matrix equation AXB = C. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001674.

Wang:2023:SEC

[WLZ23] Tongke Wang, Sijing Liu, and Zhiyue Zhang. Singular expansions and collocation methods for generalized Abel integral equations. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300184X.

Wu:2021:CAM

[WLZY21] Huixian Wu, Hezhi Luo, Fangying Zheng, and Jianfang Yang. Convergence analysis of modified p-th power Lagrangian algorithms with alternative updating strategies for constrained nonconvex optimization. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002284.

Winkler:2020:CER

[WM20] Joab R. Winkler and Marilena Mitrouli. Condition estimation for regression and feature selection. Journal of Computational and Applied Mathematics, 373(??):??, August 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719301724.

Wang:2020:CDT

[WMF20] Wei Wang, Wanbiao Ma, and Zhaosheng Feng. Complex dynamics of a time periodic nonlocal and timedelayed model of reaction– diffusion equations for modeling CD4 + T cells decline. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304339.

Wang:2022:OPE

[WMH22] Wansheng Wang, Mengli Mao, and Yi Huang. Optimal a posteriori estimators for the variable step-size BDF2 method for linear parabolic equations. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001285.

Winkler:2022:ARV

[WMK22] Joab R. Winkler, Marilena Mitrouli, and Christos Koukouvinos. The application of regularisation to variable selection in statistical modelling. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005070.

Wang:2023:MWG

[WMZZ23] Xiuli Wang, Xianglong Meng, Shangyou Zhang, and Huifang Zhou. A modified weak Galerkin finite element method for the linear elasticity problem in mixed form. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003764.

Withers:2023:NCD

[WN23] Christopher S. Withers and Saralees Nadarajah. New classes of density estimates of low bias. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001760.

Wollman:2024:FDA

[Wol24]  Stephen Wollman.  Fi-

nite difference approximations of the spatially homogeneous Fokker–Planck– Landau equation. Journal of Computational and Applied Mathematics, 449 (??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400178X.

Worz:2022:EJB

[Wör22] Sascha Wörz. An efficient Jordan basis algorithm. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004040.

Wang:2022:SOS

[WOV22] Zhibo Wang, Caixia Ou, and Seakweng Vong. A second-order scheme with nonuniform time grids for Caputo–Hadamard fractional sub-diffusion equations. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002126.

Wang:2023:PBM

[WPX23] Xiaoliang Wang, Liping Pang, and Xiantao Xiao. A proximal bundle method-based algorithm with penalty strategy and inexact oracles for constrained nonsmooth nonconvex optimization. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005477.

Wan:2020:PAE

[WQ20] Xiaole Wan and Xiaoqian Qie. Poverty alleviation ecosystem evolutionary game on smart supply chain platform under the government financial platform incentive mechanism. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306004.

Wu:2022:DTI

[WQ22] Ying Wu and Feng Qi. Discussions on two integral inequalities of Hermite– Hadamard type for convex functions. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006154.

Wang:2021:ROI

[WRZW21] Peiqi Wang, Ximin Rong, Hui Zhao, and Suxin Wang. Robust optimal investment and benefit payment adjustment strategy for target benefit pension plans under default risk. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000017.

Wang:2023:HOD

[WSC23] Dawei Wang, Kirill Serkh, and Christina Christara. A high-order deferred correction method for the solution of free boundary problems using penalty iteration, with an application to American option pricing. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002169.

Wen:2022:SCF

[WSHW22] Jing Wen, Jian Su, Yinnian He, and Zhiheng Wang. A strongly conservative finite element method for the coupled Stokes and dual-porosity model. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005021.

Wu:2023:ECG

[WSL^{+}23] Xiaoyu Wu, Hu Shao, Pengjie Liu, Yan Zhang, and Yue Zhuo. An efficient conjugate gradientbased algorithm for unconstrained optimization and its projection extension to large-scale constrained nonlinear equations with applications in signal recovery and image denoising problems. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004770.

Wang:2023:EUL

[WSLJ23] Jinhua Wang, Weiping Shen, Chong Li, and Xiaoqing Jin. An extended Ulm-like method for inverse singular value problems with multiple and/or zero singular values. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002054.

Wang:2023:IPP

[WSLW23] Xiaoquan Wang, Hu Shao, Pengjie Liu, and Ting Wu. An inertial proximal partially symmetric ADMMbased algorithm for linearly constrained multiblock nonconvex optimization problems with applications. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004198.

Wu:2021:FPP

[WSX21] Tingting Wu, Lixin Shen, and Yuesheng Xu. Fixedpoint proximity algorithms solving an incomplete Fourier transform model for seismic wavefield modeling. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304994.

Wei:2020:HVD

[WSY20] Yichen Wei, Tao Sun, and Lijun Yi. An hp-version of the discontinuous Galerkin time-stepping method for nonlinear second-order delay differential equations with vanishing delays. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303516.

Wang:2022:VCF

[WSY22] Jiangfu Wang, Zhiqiang Sheng, and Guangwei Yuan. A vertex-centered finite volume scheme preserving the discrete maximum principle for anisotropic and discontinuous diffusion equations. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004076.

Wang:2024:GCA

[WSYPD24] Xin-Yu Wang, Li-Yong Shen, Chun-Ming Yuan, and Sonia Pérez-Díaz. Globally certified G^{1} approximation of planar algebraic curves. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003436.

Wei:2020:PSR

[WSZ20] Jiaqin Wei, Yang Shen, and Qian Zhao. Portfolio selection with regime-switching and state-dependent preferences. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303644.

Wang:2022:TBH

[WT22] Yang Wang and Francesco Topputo. A TFC-based homotopy continuation algorithm with application to dynamics and control problems. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100399X.

Wang:2020:IWC

[WTL20] Liang Wang, Yogesh Mani Tripathi, and Chandrakant Lodhi. Inference for Weibull competing risks model with partially observed failure causes under generalized progressive hybrid censoring. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305424.

Wang:2021:AFF

[WTL21] Zhuoran Wang, Simon Tavener, and Jiangguo Liu. Analysis of a 2-field finite element solver for poroelasticity on quadrilateral meshes. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001588.

Wright:2021:CAS

[WTP^{+}21] Joseph P. Wright, Peng F. Tang, Jin-Song Pei, François Gay-Balmaz, and Joseph P. Havlicek. On computing the analytic-signal backbone of the unforced harmonic oscillator. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720304970.

Wang:2020:ICS

[WTWZ20] Liang Wang, Yogesh Mani Tripathi, Shuo-Jye Wu, and Meng Zhang. Inference for confidence sets of the generalized inverted exponential distribution under k-record values. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302600.

Wang:2024:LDG

[WTX24] Fan Wang, Qi Tao, and Yan Xu. The local discontinuous Galerkin method for the nonlinear quantum Zakharov system. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003606.

Wu:2022:REB

[Wu22] Xianping Wu. The refined error bounds for linear complementarity problems of H_{+} -matrices. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003733.

Wu:2024:EPA

[Wu24]  Cheng-Hsun Wu.  Ex-

act perturbation approximations for the conditional moments of a multifactor CIR term structure model with a weak mean-reversion influence. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001493.

Witherden:2021:NPS

[WV21] F. D. Witherden and P. E.

Vincent. On nodal point sets for flux reconstruction. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303058.

Wang:2022:LTA

[WW22] Bin Wang and Xinyuan Wu. Long-time analysis of an extended RKN integrator for Hamiltonian systems with a solutiondependent high frequency. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002680.

Wang:2023:OPW

[WW23] Tingchun Wang and Tingfeng Wang. Optimal point-wise error estimates of two conservative finite difference schemes for the coupled Gross–Pitaevskii equations with angular momentum rotation terms. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006549.

Wang:2024:VMC

[WW24a] Wei Wang and Jingjie Wang. A variational model for cartoon-texture decomposition of a color image. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001821.

Wang:2024:DRU

[WW24b] Yibin Wang and Haifeng Wang. Distributionally robust unsupervised domain adaptation. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003138.

Wang:2023:DRR

[WWC23] Wenyuan Wang, Ning Wang, and Mi Chen. On a doubly reflected risk process with running maximum dependent reflecting barriers. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004782.

Wang:2021:VEM

[WWH21a] Fei Wang, Bangmin Wu, and Weimin Han. The virtual element method for general elliptic hemivariational inequalities. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030621X.

Wang:2021:WGF

[WWH21b] Gang Wang, Ying Wang, and Yinnian He. A weak Galerkin finite element method based on H(div) virtual element for Darcy flow on polytopal meshes. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002995.

Wang:2022:SOL

[WWJ22] Danxia Wang, Xingxing Wang, and Hongen Jia. A second order linear energy stable numerical method for the Cahn– Hilliard–Hele–Shaw system. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004106.

Wang:2021:GSD

[WWS^{+}21] Ruibing Wang, Min Wu, Yimin Shi, Hon Keung Tony Ng, and Fode Zhang. The geometric structure on a degradation model with application to optimal design under a cost constraint. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303721.

Wang:2024:FWG

[WWTL24] Ruishu Wang, Zhuoran Wang, Simon Tavener, and Jiangguo Liu. Full weak Galerkin finite element discretizations for poroelasticity problems in the primal formulation. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000037.

Wang:2024:CAT

[WWW^{+}24] Cheng Wang, Jilu Wang, Steven M. Wise, Zeyu Xia, and Liwei Xu. Convergence analysis of a temporally second-order accurate finite element scheme for the Cahn–Hilliardmagnetohydrodynamics system of equations. Journal of Computational and Applied Mathematics, 436 (??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042723003539.

Wang:2023:NCT

[WWWY23] Yu Guang Wang, Robert S. Womersley, Hau-Tieng Wu, and Wei-Hsuan Yu. Numerical computation of triangular complex spherical designs with small mesh ratio. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004058.

Wang:2021:RCW

[WWYZ21] Chunmei Wang, Junping Wang, Xiu Ye, and Shangyou Zhang. de Rham complexes for weak Galerkin finite element spaces. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002673.

Wang:2023:WGF

[WWZ23] Chunmei Wang, Junping Wang, and Shangyou Zhang. Weak Galerkin finite element methods for quad-curl problems. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001309.

Wang:2024:WGF

[WWZ24] Chunmei Wang, Junping Wang, and Shangyou Zhang. Weak Galerkin finite element methods for optimal control problems governed by second order elliptic equations. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002322.

Wu:2020:PRK

[WX20] Nianci Wu and Hua Xiang. Projected randomized Kaczmarz methods. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306776.

Wang:2021:SNA

[WX21] Dongming Wang and Juan Xu. A symbolic-numerical algorithm for isolating real roots of certain radical expressions. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000431.

Wei:2022:DTD

[WX22] Ting Wei and Jun Xian. Determining a time-dependent coefficient in a timefractional diffusion-wave equation with the Caputo derivative by an additional integral condition. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005331.

Wang:2021:ABS

[WXS21] Dongling Wang, Aiguo Xiao, and Suzhen Sun. Asymptotic behavior of solutions to time fractional neutral functional differential equations. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303770.

Wang:2021:ENT

[WXZ21] Rui Wang, Naihua Xiu, and Shenglong Zhou. An extended Newton-type algorithm for l_{2}-regularized sparse logistic regression and its efficiency for classifying large-scale datasets. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002788.

Wang:2020:ESL

[WY20] Lin Wang and Haijun Yu. An energy stable linear diffusive Crank– Nicolson scheme for the Cahn–Hilliard gradient flow. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301710.

Wei:2021:OOF

[WY21a] Leilei Wei and Yanfang Yang. Optimal order finite difference/local discontinuous Galerkin method for variable-order timefractional diffusion equation. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304209.

Wu:2021:GHI

[WY21b] Weidong Wu and Xunnian Yang. Geometric Hermite interpolation by a family of spatial algebraic-trigonometric PH curves. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305872.

Wang:2024:CIF

[WY24a] Wei Wang and Yuming Yang. A color image fusion model by saturation-value total variation. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000815.

Wang:2024:ADC

[WY24b] Xiaoliang Wang and Gonglin Yuan. An accelerated descent CG algorithm with clustering the eigenvalues for large-scale nonconvex unconstrained optimization and its application in image restoration problems. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003989.

Wang:2024:MMI

[WYC24] Zhongqian Wang, Changqing Ye, and Eric T. Chung. A multiscale method for inhomogeneous elastic problems with high contrast coefficients. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003412.

Wang:2020:PMH

[WYD20] Zeng-Qi Wang, Jun-Feng Yin, and Quan-Yu Dou. Preconditioned modified Hermitian and skew-Hermitian splitting iteration methods for fractional nonlinear Schrödinger equations. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304236.

Wang:2023:NNI

[WYG23] Guoshun Wang, Dansheng Yu, and Lingmin Guan. Neural network interpolation operators of multivariate functions. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002108.

Wang:2022:AFE

[WYH22] Hao Wang, Wei Yang, and Yunqing Huang. Adaptive finite element method for two-dimensional timeharmonic magnetic induction intensity equations. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001352.

Wang:2023:TES

[WYJ23] Jindi Wang, Yin Yang, and Bingquan Ji. Two energy stable variable-step L1 schemes for the timefractional MBE model without slope selection. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003557.

Wang:2019:CFP

[WYTL19] Guoqiu Wang, Dingxun Yi, Rong Tang, and Wei Liang. Complete factorization for poly-phase matrix with linear phase based on semi-rank orthogonal projection matrix. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303930.

Wang:2024:SAA

[WYW^{+}24] Feiran Wang, Yuan Yuan, Xiaoquan Wang, Hu Shao, and Liang Shen. A sequential ADMM algorithm to find sparse LCP solutions using a l_{2} − l_{1} regularization technique with application in bimatrix game. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004144.

Wang:2024:MWG

[WYZ24] Chunmei Wang, Xiu Ye, and Shangyou Zhang. A modified weak Galerkin finite element method for the Maxwell equations on polyhedral meshes. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001687.

Wang:2022:WGL

[WYZZ22] Xiaoshen Wang, Xiu Ye, Shangyou Zhang, and Peng Zhu. A weak Galerkin least squares finite element method of Cauchy problem for Poisson equation. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003897.

Wang:2020:OOE

[WZ20a] Hong Wang and Xiangcheng Zheng. An optimal-order error estimate of the lowest-order ELLAM–MFEM approximation to miscible displacement in three space dimensions. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301102.

Wu:2020:CTB

[WZ20b] Yan Wu and Chun-Gang Zhu. Construction of triharmonic Bézier surfaces from boundary conditions. Journal of Computational and Applied Mathematics, 377(??):??, October 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301977.

Wang:2021:LRP

[WZ21] Chunmei Wang and Ludmil Zikatanov. Low regularity primal-dual weak Galerkin finite element methods for convection– diffusion equations. Journal of Computational and Applied Mathematics, 394 (??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272100162X.

Wang:2023:WGM

[WZ23a] Chunmei Wang and Shangyou Zhang. A weak Galerkin method for elasticity interface problems. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003673.

Wang:2023:ARH

[WZ23b] Haifeng Wang and Yufeng Zhang. Application of Riemann–Hilbert method to an extended coupled nonlinear Schrödinger equations. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004137.

Wang:2023:CSV

[WZ23c] Tong Wang and Pingping Zhao. Consumer strategy, vendor strategy and equilibrium in duopoly markets with production costs. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001486.

Wang:2024:CRP

[WZ24a] Meijuan Wang and Shugong Zhang. A class of refined preconditioners with sparse error correction for BEM linear system. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004491.

Wang:2024:PVL

[WZ24b] Xinying Wang and Ke Zhou.

Pricing vulnerable lookback options using Laplace transforms. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002644.

Wang:2022:BTP

[WZC^{+}22a] Dongwu Wang, Bin Zheng, Long Chen, Guang Lin, and Jinchao Xu. Block triangular preconditioning for stochastic Galerkin method. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001248.

Wu:2022:GBA

[WZC22b] Xiang Wu, Kanjian Zhang, and Ming Cheng. A gradient-based algorithm for non-smooth constrained optimization problems governed by discrete-time nonlinear equations with application to long-term hydrothermal optimal scheduling control. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001479.

Weng:2024:SEE

[WZD^{+}24] Zhifeng Weng, Shuying Zhai, Weizhong Dai, Yanfang Yang, and Yuchang Mo. Stability and error estimates of Strang splitting method for the nonlocal ternary conservative Allen–Cahn model. Journal of Computational and Applied Mathematics, 441 (??):??, May 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300612X.

Wu:2021:SRC

[WZH^{+}21] C. Wu, H. Zeng, Y. Huang, N. Yi, and J. Yuan. Superconvergence recovery of cubic edge elements for Maxwell’s equations. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306245.

Wang:2021:RIG

[WZJQ21] Ning Wang, Nan Zhang, Zhuo Jin, and Linyi Qian. Reinsurance-investment game between two mean-variance insurers under model uncertainty. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303861.

Wen:2023:FIP

[WZL^{+}23] Meng Wen, Yongqiang Zhang, Haiyang Li, Yuchao Tang, and Jigen Peng. A fast inertial primal-dual algorithm to composite optimization models with application to image restoration problems. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006410.

Wu:2023:DMD

[WZLT23] Weiping Wu, Ke Zhou, Zhicheng Li, and Zhenpeng Tang. Dynamic meandownside risk portfolio selection with a stochastic interest rate in continuoustime. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300047X.

Wang:2020:EFR

[WZLX20] Guochang Wang, Baoxue Zhang, Wenhui Liao, and Baojian Xie. Estimation of functional regression model via functional dimension reduction. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302399.

Wang:2024:NDEa

[WZNT24] Xuhui Wang, Yanchun Zhao, Qian Ni, and Shuo Tang. Nonparametric density estimation with nonuniform B-spline bases. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005927.

Wang:2023:FDS

[WZY23] Ziqiang Wang, Jun Zhang, and Xiaofeng Yang. Fully discrete Spectral-Galerkin scheme for a ternary Allen– Cahn type mass-conserved Nakazawa–Ohta phasefield model for triblock copolymers. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003545.

Wang:2020:EIS

[WZZ20] Xiuli Wang, Yongkui Zou, and Qilong Zhai. An effective implementation for Stokes Equation by the weak Galerkin finite element method. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305916.

Wang:2021:SGA

[WZZ21a] Fangyuan Wang, Zhongqiang Zhang, and Zhaojie Zhou. A spectral Galerkin approximation of optimal control problem governed by fractional advection– diffusion–reaction equations. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305240.

Wang:2021:CDM

[WZZ21b] Jiangxing Wang, Jiwei Zhang, and Zhimin Zhang. A CG–DG method for Maxwell’s equations in Cole–Cole dispersive media. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000996.

Wu:2024:ROM

[WZZ24] Zhizhang Wu, Cheng Zhang, and Zhiwen Zhang. Reduced-order model-based variational inference with normalizing flows for Bayesian elliptic inverse problems. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723006039.

Wei:2020:SCC

[WZZT20] Jingdong Wei, Jiangbo Zhou, Zaili Zhen, and Lixin Tian. Super-critical and critical traveling waves in a three-component delayed disease system with mixed diffusion. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304546.

Xu:2021:CDD

[XA21] Qiuyan Xu and Hengbin An. A class of domain decomposition based nonlinear explicit-implicit iteration algorithms for solving diffusion equations with discontinuous coefficient. Journal of Computational and Applied Mathematics, 386(??):??, April 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305239.

Xu:2021:IEP

[XBC21] Wei-Ru Xu, Natália Bebiano, and Guo-Liang Chen. An inverse eigenvalue problem for modified pseudo-Jacobi matrices. Journal of Computational and Applied Mathematics, 389(??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030652X.

Xu:2022:IEP

[XBC22] Wei-Ru Xu, Natália Bebiano, and Guo-Liang Chen. An inverse eigenvalue problem for doubly periodic pseudo-Jacobi matrices. Journal of Computational and Applied Mathematics, 405(??):??, May 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005641.

Xu:2020:AET

[XCWY20] Zhaoyi Xu, Xinyu Cheng, Kefei Wang, and Shenggang Yang. Analysis of the environmental trend of network finance and its influence on traditional commercial banks. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301989.

Xu:2024:PLB

[XCY24] Jingtong Xu, Xu Chen, and Yuying Yang. Pricing longevity bond with affinejump-diffusion multi-cohort mortality model. Journal of Computational and Applied Mathematics, 446 (??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724000499.

Xu:2023:SSG

[XD23] Ling Xu and Feng Ding. Separable synthesis gradient estimation methods and convergence analysis for multivariable systems. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000481.

Xing:2024:AMB

[XDP24] Haoming Xing, Feng Ding, and Feng Pan. Auxiliary model-based hierarchical stochastic gradient methods for multipleinput multiple-output systems. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006301.

Xu:2024:CAS

[XDX24] Ning Xu, Feng Ding, and Ling Xu. Convergence analysis of a synchronous gradient estimation scheme for time-varying parameter systems. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006672.

Xie:2023:WGM

[XFW23] Chun-Mei Xie, Min-Fu Feng, and Hua-Yi Wei. An H^{1} weak Galerkin mixed finite element method for Sobolev equation. Journal of Computational and Applied Mathematics, 423 (??):??, May 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005775.

Xiao:2024:DLB

[XFW24] Jichang Xiao, Fengjiang Fu, and Xiaoqun Wang. Deep learning based on randomized quasi-Monte Carlo method for solving linear Kolmogorov partial differential equation. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003376.

Xin:2023:MLE

[XGYY23] Yue Xin, Jinwu Gao, Xiangfeng Yang, and Jing Yang. Maximum likelihood estimation for uncertain autoregressive moving average model with application in financial market. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003016.

Xu:2020:CAF

[XH20a] Fei Xu and Qiumei Huang. Cascadic adaptive finite element method for nonlinear eigenvalue problem based on complementary approach. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030011X.

Xue:2020:NAS

[XH20b] Dandan Xue and Yanren Hou. Numerical analysis of a second order algorithm for a nonstationary Navier–Stokes/ Darcy model. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305849.

Xia:2022:AHP

[XHG22] Yuxin Xia, Bo Han, and Ruixue Gu. An accelerated homotopy-perturbation-Kaczmarz method for solving nonlinear inverse problems. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005203.

Xiao:2021:AZB

[XHLD21] Lin Xiao, Yongjun He, Bolin Liao, and Jianhua Dai. An accelerated ZNNbased algorithm with piecewise time-varying parameters to solve time-variant linear equations. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002879.

Xu:2024:ERE

[XHW24] Zhenghang Xu, Zhijian He, and Xiaoqun Wang. Efficient risk estimation via nested multilevel quasi-Monte Carlo simulation. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006878.

Xiao:2022:NMQ

[Xia22] Qing Xiao. A new multivariate quadrature rule for calculating statistical moments of stochastic response. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005227.

Xing:2021:NOS

[Xin21] F. Xing. New optimized Schwarz algorithms for one dimensional Schrödinger equation with general potential. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303095.

Xian:2019:CKE

[XJK19] Da-Quan Xian, Ying Jiang, and Xiao-Rong Kang. Consistent KdV expansion method and its applications to the KP equation. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304261.

Xu:2020:RBI

[XL20] Yunfu Xu and Aiya Li. The relationship between innovative human capital and interprovincial economic growth based on panel data model and spatial econometrics. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930384X.

Xiao:2023:VEM

[XL23] Wenqiang Xiao and Min Ling. The virtual element method for general variational–hemivariational inequalities with applications to contact mechanics. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000961.

Xia:2023:BTF

[XLKL23] Qing Xia, Yuehan Liu, Junseok Kim, and Yibao Li. Binary thermal fluids computation over arbitrary surfaces with secondorder accuracy and unconditional energy stability based on phase-field model. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002637.

Xu:2023:ENM

[XLL23] Zhenhua Xu, Zhanmei Lv, and Guidong Liu. Efficient numerical methods for hypersingular finite-part integrals with highly oscillatory integrands. Journal of Computational and Applied Mathematics, 433(??):??, December 1, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723002601.

Xu:2021:NAM

[XLLA21] Tao Xu, Fawang Liu, Shujuan Lü, and Vo V. Anh. Numerical approximation of 2D multi-term time and space fractional Bloch– Torrey equations involving the fractional Laplacian. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001382.

Xie:2022:FOC

[XLY22] Yaning Xie, Shuwang Li, and Wenjun Ying. A fourth-order Cartesian grid method for multiple acoustic scattering on closely packed obstacles. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005082.

Xie:2020:BFI

[XLZ20a] Meihua Xie, Haiyan Li, and Yuanjun Zhao. Blockchain financial investment based on deep learning network algorithm. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300145.

Xu:2020:EAS

[XLZ20b] Weiwei Xu, Ying Lu, and Lei Zhu. An extension of analytic solutions of a class of constrained matrix minimization problems. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306764.

Xu:2022:QOS

[XSL22] Chenglong Xu, Bihao Su, and Chan Liu. A quick operator splitting method for option pricing. Journal of Computational and Applied Mathematics, 406 (??):??, May 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005604.

Xu:2023:WSV

[XTL23] Wenhua Xu, Lihui Tan, and Rongrong Lin. Weighted singular value decomposition basis of Szegő kernel and its applications to signal reconstruction and denoising. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000110.

Xu:2020:POP

[Xu20]  Xuefeng Xu.  On the

perturbation of an L_{2}- orthogonal projection. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303243.

Xie:2021:VEA

[XW21] Cong Xie and Kun Wang.

Viscosity explicit analysis for finite element methods of time-dependent Navier–Stokes equations. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100100X.

Xiang:2022:QMC

[XW22] Jiangming Xiang and Xiaoqun Wang. Quasi-Monte Carlo simulation for American option sensitivities. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001108.

Xu:2024:AMM

[XWL24] Fei Xu, Bingyi Wang, and Fusheng Luo. Adaptive multigrid method for quantum eigenvalue problems. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003898.

Xiong:2022:LRT

[XWXX22] Zikai Xiong, Yimin Wei, Renjie Xu, and Yanwei Xu. Low-rank traffic matrix completion with marginal information. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000814.

Xia:2022:HOA

[XWXZ22] Zeyu Xia, Cheng Wang, Liwei Xu, and Zhengru Zhang. High order accurate in time, fourth order finite difference schemes for the harmonic mapping flow. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003885.

Xiao:2024:MRN

[XWZ24] Yao Xiao, Qingbiao Wu, and Yuanyuan Zhang. Minimum residual NDSS iteration method for a class of complex symmetric linear systems. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001730.

Xu:2022:COA

[XXH^{+}22] Fei Xu, Manting Xie, Qiumei Huang, Meiling Yue, and Hongkun Ma. Convergence and optimality of adaptive multigrid method for multiple eigenvalue problems. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002138.

Xiong:2021:WTF

[XXL21] Xiangtuan Xiong, Xuemin Xue, and Zhenping Li. On a weighted time-fractional asymptotical regularization method. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001990.

Xiang:2022:NSL

[XY22] Sijia Xiang and Weixin Yao. Nonparametric statistical learning based on modal regression. Journal of Computational and Applied Mathematics, 409 (??):??, August 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722000280.

Xiao:2024:ABK

[XY24] A-Qin Xiao and Jun-Feng Yin. On averaging block Kaczmarz methods for solving nonlinear systems of equations. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002917.

Xie:2023:LDP

[XYAH23] Jianqiang Xie, Xiao Yan, Muhammad Aamir Ali, and Zakia Hammouch. A linear decoupled physicalproperty-preserving difference method for fractionalorder generalized Zakharov system. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006422.

Xie:2022:RAF

[XZ22] Jiayi Xie and Zhimin Zhang. Recursive approximating to the finitetime Gerber–Shiu function in Lévy risk models under periodic observation. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003253.

Xie:2020:CAT

[XZL20] Yingying Xie, Liuqiang Zhong, and Chunmei Liu. Convergence of an AE-FEM for time-harmonic Maxwell equations with variable coefficients. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300030.

Xiao:2022:DIS

[XZP22] Yue Xiao, Shuguang Zhang, and Yu Peng. Dynamic investigations in a Stackelberg model with differentiated products and bounded rationality. Journal of Computational and Applied Mathematics, 414(??):??, November 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001881.

Xie:2020:BDS

[XZZ^{+}20] Guizhong Xie, Fenglin Zhou, Yudong Zhong, Hongrui Geng, and Changjun Wu. Bi-directional sinh transformations based on the generalized duffy space for nearly singular integrals. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302727.

Yousef:2023:COC

[YAE23] A. M. Yousef, Ahmed M.

Algelany, and A. A. Elsadany. Codimension one and codimension two bifurcations in a discrete Kolmogorov type predatorprey model. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001152.

Yakubovich:2021:OCO

[Yak21] Semyon Yakubovich. On the orthogonality and convolution orthogonality via the Kontorovich–Lebedev transform. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304696.

Yamamura:2020:FAS

[Yam20] Kiyotaka Yamamura. Finding all solution sets of piecewise-linear interval equations using an integer programming solver. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306211.

Yamamura:2021:EAF

[Yam21a] Kiyotaka Yamamura. An efficient algorithm for finding all solutions of nonlinear equations using parallelogram LP test. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272030371X.

Yamazaki:2021:NAW

[Yam21b] Kazuo Yamazaki. A note on the applications of Wick products and Feynman diagrams in the study of singular partial differential equations. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306294.

Yaman:2023:SSO

[Yam23] Gülsen Yaman. A suggestion of standard and optimized steps in the LOC (Lab on a Chip), LOD (Lab on a Disc), and POC (Point of Care) development process for biomedical applications: a case study about ESR. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003132.

Yang:2022:CEQ

[Yan22]  Zhao Yang.  Comput-

ing eigenvalues of quasigeneralized Vandermonde matrices to high relative accuracy. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006117.

Yousef:2023:MMC

[YBAE23] Ali Yousef, Fatma Bozkurt, Thabet Abdeljawad, and Emad Emreizeeq. A mathematical model of COVID-19 and the multi fears of the community during the epidemiological stage. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003120.

Yildiz:2023:SBA

[YBM23] Sevval Yildiz, Seyma Bilazeroglu, and Hüseyin Merdan. Stability and bifurcation analyses of a discrete Lotka–Volterra type predator-prey system with refuge effect. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005088.

Yang:2020:UDT

[YC20a] Shiueng-Bien Yang and Tai-Liang Chen. Uncertain decision tree for bank marketing classification. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300017.

Yu:2020:RBP

[YC20b] Dan Yu and Suman Chakravorty. A randomized balanced proper orthogonal decomposition technique. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930545X.

Yuan:2021:IEP

[YC21] Yongxin Yuan and Jinghua Chen. An inverse eigenvalue problem for Hamiltonian matrices. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303228.

Yang:2023:AES

[YC23] Zhao Yang and Tao Chen.

The accurate and efficient solutions of linear systems for generalized sign regular matrices with certain signature. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002248.

Yu:2024:LCU

[YCCZ24] Tsz Fung Yu, Eric T.

Chung, Ka Chun Cheung, and Lina Zhao. Learning computational upscaling models for a class of convection–diffusion equations. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000633.

Yiqi:2020:PSC

[YCD^{+}20] Bao Yiqi, Vicente G. Cancho, Dipak K. Dey, Narayanaswamy Balakrishnan, and Adriano K. Suzuki. Power series cure rate model for spatially correlated intervalcensored data based on generalized extreme value distribution. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303656.

Yan:2023:NSI

[YCG^{+}23] Peng Yan, Camilla Cecilia Conti, Giulio Gori, Barbara Re, and Alberto Guardone. Numerical simulation of ideal and nonideal under-expanded supersonic jets with adaptive grids. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001139.

Yu:2021:IDR

[YCH^{+}21] Bo Yu, Geyong Cao, Wendong Huo, Huanlin Zhou, and Elena Atroshchenko. Isogeometric dual reciprocity boundary element method for solving transient heat conduction problems with heat sources. Journal of Computational and Applied Mathematics, 385(??):??, March 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030488X.

Yirang:2024:UMV

[YCH24] Yuan Yirang, Li Changfeng, and Song Huailing. An upwind-mixed volume element method on changing meshes for compressible miscible displacement problem. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000785.

Yang:2020:TCE

[YCX20] Peng Yang, Zhiping Chen, and Ying Xu. Timeconsistent equilibrium reinsuranceinvestment strategy for n competitive insurers under a new interaction mechanism and a general investment framework. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300601.

Yan:2024:EBG

[YCXW24] Xiaoqiang Yan, Shi Chen, Aiguo Xiao, and Huiru Wang. The extended block generalized Störmer-Cowell methods for secondorder nonlinear delaydifferential-algebraic equations with index-1. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723005940.

Ye:2020:CRM

[YDC20] Changqing Ye, Hao Dong, and Junzhi Cui. Convergence rate of multiscale finite element method for various boundary problems. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300455.

Yeganeh:2022:MPP

[YF22] S. Mousavi Yeganeh and J. Farzi. Maximum principle and positivitypreserving high order spectral volume schemes with parametrized flux limiters for solving hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005161.

You:2020:RCF

[YFW20] Zhongli You, Michal Feckan, and JinRong Wang. Relative controllability of fractional delay differential equations via delayed perturbation of Mittag-Leffler functions. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302302.

Yang:2022:CSM

[YH22] Hongfu Yang and Jianhua Huang. Convergence and stability of modified partially truncated Euler–Maruyama method for nonlinear stochastic differential equations with Hölder continuous diffusion coefficient. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005185.

Yu:2020:CNS

[YHC20] Yang Yu, Guolin Hou, and Alatancang Chen. Comments on “A note on structured pseudospectra of block matrices” [j. comput. appl. math. 322 (2017) 18– 24]. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301709.

Yazdani:2022:CAB

[YHF22] S. Yazdani, M. Hadizadeh, and V. Fakoor. Computational analysis of the behavior of stochastic volatility models with financial applications. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001054.

Yu:2021:ECF

[YHQW21] Shao-Wen Yu, Zhuo-Heng He, Tian-Cheng Qi, and Xiang-Xiang Wang. The equivalence canonical form of five quaternion matrices with applications to imaging and Sylvester-type equations. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001138.

Yilmaz:2022:DPO

[YHSK22] Bilgi Yilmaz, A. Alper Hekimoglu, and A. Sevtap Selcuk-Kestel. Default and prepayment options pricing and default probability valuation under VG model. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721003460.

Yan:2022:TTR

[YHW22] Xiaokuai Yan, Qinglong He, and Yanfei Wang. Truncated trust region method for nonlinear inverse problems and application in full-waveform inversion. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721005197.

Yin:2023:FIR

[YJJW23] Jianghua Yin, Jinbao Jian, Xianzhen Jiang, and Xiaodi Wu. A family of inertial-relaxed DFPMbased algorithms for solving large-scale monotone nonlinear equations with application to sparse signal restoration. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003417.

Yang:2020:RSG

[YJN20] Zhen Yang, Yan-Fei Jing, and Qiang Niu. Restarted simpler GMRES augmented with harmonic Ritz vectors and approximate errors. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305709.

Yu:2023:SCI

[YJZ23] Ying-Ying Yu, Ye Ji, and Chun-Gang Zhu. Sufficient condition for injectivity of NURBS volumes by tangent cones. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002479.

Yahaya:2021:SQN

[YKAA21] Mahmoud Muhammad Yahaya, Poom Kumam, Aliyu Muhammed Awwal, and Sani Aji. A structured quasi-Newton algorithm with nonmonotone search strategy for structured NLS problems and its application in robotic motion control. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002028.

Yan:2023:ESD

[YKBH23] Ge Yan, Sharanjeet Kaur, Jeffrey W. Banks, and Jason E. Hicken. Entropystable discontinuous Galerkin difference methods for hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004836.

Yang:2024:FFP

[YKY24] Hyoseon Yang, Kyungmi Kim, and Jungho Yoon. A family of C^{2} fourpoint stationary subdivision schemes with fourthorder accuracy and shapepreserving properties. Journal of Computational and Applied Mathematics, 446 (??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400092X.

Yu:2020:PWC

[YL20] Yuncai Yu and Xinsheng Liu. Pointwise wavelet change-points estimation for dependent biased sample. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302776.

Yan:2021:UIP

[YL21a] Dong Yan and Xiaoping Lu. Utility-indifference pricing of European options with proportional transaction costs. Journal of Computational and Applied Mathematics, 397 (??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721002612.

Yao:2021:MIF

[YL21b] Xudong Yao and Zhujun Li. A Morse index formula for minimax type saddle points by a Ljusternik– Schnirelman minimax algorithm and its application in computation of multiple solutions of semilinear elliptic equation. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303678.

Yu:2021:WCP

[YL21c] Yuncai Yu and Xinsheng Liu. Wavelet change-point estimation for the density based on biased sample. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000674.

Yang:2022:TFV

[YL22] Zhen-Ping Yang and Gui-Hua Lin. Two fast variance-reduced proximal gradient algorithms for SMVIPs — Stochastic Mixed Variational Inequality Problems with suitable applications to stochastic network games and traffic assignment problems. Journal of Computational and Applied Mathematics, 408(??):??, July 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000292.

Yang:2019:MPI

[YLD19] Liu Yang, Yun Liu, and Zui-Cha Deng. Multiparameters identification problem for a degenerate parabolic equation. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930425X.

Yang:2021:NFV

[YLFT21] Shuiping Yang, Fawang Liu, Libo Feng, and Ian Turner. A novel finite volume method for the nonlinear two-sided space distributed-order diffusion equation with variable coefficients. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306282.

Yuan:2023:ROR

[YLH23] Yu Yuan, Zhibin Liang, and Xia Han. Robust optimal reinsurance in minimizing the penalized expected time to reach a goal. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004150.

Yang:2024:ELU

[YLJW24] Junxiang Yang, Mengyu Luo, Wenjing Jiang, and Jian Wang. Efficiently linear and unconditionally energy-stable timemarching schemes with energy relaxation for the phase-field surfactant model. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002899.

Yang:2024:PFM

[YLK24] Junxiang Yang, Yibao Li, and Junseok Kim. Phasefield modeling and linearly energy-stable Runge– Kutta algorithm of colloidal crystals on curved surfaces. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006921.

Yang:2020:RCL

[YLL20] Jing Yang, Fang Lu, and Xuewen Lu. Robust check loss-based inference of semiparametric models and its application in environmental data. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930250X.

Yu:2023:NUL

[YLLA23] Hui Yu, Fawang Liu, Mingxia Li, and Vo V. Anh. The non-uniform L1-type scheme coupling the finite volume method for the time–space fractional diffusion equation with variable coefficients. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001231.

Yuan:2020:MFV

[YLS20] Yirang Yuan, Changfeng Li, and Huailing Song. Mixed finite volume elementupwind mixed volume element of compressible twophase displacement and its numerical analysis. Journal of Computational and Applied Mathematics, 370 (??):??, May 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306429.

Yuan:2023:UBC

[YLS23] Yirang Yuan, Changfeng Li, and Huailing Song. An upwind-block-centered finite difference method for a semiconductor device of heat conduction and its numerical analysis. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005763.

Yang:2021:PVD

[YLWW21] Yu Yang, Shican Liu, Yonghong Wu, and Benchawan Wiwatanapataphee. Pricing of volatility derivatives in a Heston–CIR model with Markov-modulated jump diffusion. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305689.

Yang:2023:PLP

[YLXY23] Fan Yang, Zhanyang Li, Yushan Xue, and Yuehan Yang. A penalized least product relative error loss function based on wavelet decomposition for nonparametric multiplicative additive models. Journal of Computational and Applied Mathematics, 432(??):??, November 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002431.

Yang:2022:FER

[YLZ^{+}22] Hanmei Yang, Jian Lu, Heng Zhang, Ye Luo, and Jianwei Lu. Field of experts regularized nonlocal low rank matrix approximation for image denoising. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000942.

Yazdani:2022:WGF

[YMC22] A. Yazdani, H. Momeni, and M. S. Cheichan. A weak Galerkin/ finite difference method for time-fractional biharmonic problems in two dimensions. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000668.

Yanbarisov:2021:PBM

[YN21] Ruslan M. Yanbarisov and Kirill D. Nikitin. Projection-based monotone embedded discrete fracture method for flow and transport in porous media. Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001035.

Yereniuk:2020:CFC

[YO20] Michael A. Yereniuk and Sarah D. Olson. Computational framework to capture the spatiotemporal density of cells with a cumulative environmental coupling. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305771.

Ye:2022:EDS

[YOCY22] Qiongwei Ye, Zhigang Ouyang, Chuanjun Chen, and Xiaofeng Yang. Efficient decoupled secondorder numerical scheme for the flow-coupled Cahn– Hilliard phase-field model of two-phase flows. Journal of Computational and Applied Mathematics, 405 (??):??, May 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004982.

Yeung:2021:GDG

[YP21] David W. K. Yeung and Leon A. Petrosyan. Generalized dynamic games with durable strategies under uncertain planning horizon. Journal of Computational and Applied Mathematics, 395(??):??, October 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002156.

Yang:2020:FTR

[YPL20] Fan Yang, Qu Pu, and Xiao-Xiao Li. The fractional Tikhonov regularization methods for identifying the initial value problem for a time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 380 (??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042720302892.

Yang:2022:NAS

[YPLM22] Huizi Yang, Yanxi Pan, Wenxiu Liu, and Zhitong Mu. Numerical analysis of split-step θ methods with truncated Wiener process for a stochastic SIS epidemic model. Journal of Computational and Applied Mathematics, 415(??):??, December 1, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002059.

Yan:2021:TRE

[YQZS21] Jingye Yan, Xu Qian, Hong Zhang, and Songhe Song. Two regularized energypreserving finite difference methods for the logarithmic Klein–Gordon equation. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721000972.

Yan:2022:NAN

[YQZS22] Xiaoqiang Yan, Xu Qian, Hong Zhang, and Songhe Song. Numerical approximation to nonlinear delaydifferential–algebraic equations with proportional delay using block boundary value methods. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004908.

Yang:2024:IPM

[YS24a] Chen Yang and Qinghe Shi. An interval perturbation method for singular value decomposition (SVD) with unknown-but-bounded (UBB) parameters. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003801.

Yu:2024:FNS

[YS24b] Yi Yu and Marcus Sarkis.

A family of Nonoverlapping Spectral Additive Schwarz methods (NOSAS) and their economic versions. Journal of Computational and Applied Mathematics, 443(??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006763.

Yang:2020:RBF

[YSBT20] Yin Yang, Fazlollah Soleymani, Mahdiar Barfeie, and Emran Tohidi. A radial basis function– Hermite finite difference approach to tackle cashor-nothing and asset-or-nothing options. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704271930528X.

Yazdaniyan:2020:NMB

[YSFdRdP20] Z. Yazdaniyan, M. Shamsi, Z. Foroozandeh, and Maria do Rosário de Pinho. A numerical method based on the complementarity and optimal control formulations for solving a family of zero-sum pursuit-evasion differential games. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305400.

Yang:2022:LAT

[YSFS22] Ying Yang, Ruigang Shen, Mingjuan Fang, and Shi Shu. Local averaging type a posteriori error estimates for the nonlinear steady-state Poisson– Nernst–Planck equations. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721004970.

Yadav:2024:EAR

[YSN24] Abhishek Yadav, Amit Setia, and M. Thamban Nair. Error analysis of a residualbased Galerkin’s method for a system of Cauchy singular integral equations with vanishing endpoint conditions. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003096.

Yadav:2024:AFE

[YSUY24] Rohini Yadav, Housila P.

Singh, Lakshmi N. Upadhyaya, and Subhash K. Yadav. Adroit family of estimators of population mean using known auxiliary parameters. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003990.

Yoshioka:2019:ACO

[YT19] Hidekazu Yoshioka and Motoh Tsujimura. Analysis and computation of an optimality equation arising in an impulse control problem with discrete and costly observations. Journal of Computational and Applied Mathematics, 366 (??):??, ???? 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304029.

Yang:2020:SBR

[YT20] Zhen-Hang Yang and Jing-Feng Tian. Sharp bounds for the ratio of two zeta functions. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303620.

Yazici:2024:GFB

[YT24] D. Yazici and S. Topuz. Generalized fractional bi-Hamiltonian structure of Plebański’s second heavenly equation in terms of conformable fractional derivatives. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003704.

Yang:2022:AAB

[YTA22] Yunan Yang, Alex Townsend, and Daniel Appelö. Anderson acceleration based on the H^{−s} Sobolev norm for contractive and noncontractive fixed-point operators. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004672.

Yang:2023:HAC

[YTD23] Yin Yang, Emran Tohidi, and Guoting Deng. A high accurate and convergent numerical framework for solving highorder nonlinear Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004502.

Yang:2019:RCA

[YTMK19] Yin Yang, Emran Tohidi, Xiaohua Ma, and Sujuan Kang. Rigorous convergence analysis of Jacobi spectral Galerkin methods for Volterra integral equations with noncompact kernels. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304066.

Yoshioka:2023:MCC

[YTT23] Hidekazu Yoshioka, Motoh Tsujimura, and Haruka Tomobe. Modeling and computation of cost-constrained adaptive environmental management with discrete observation and intervention. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005726.

Yu:2022:SDD

[Yu22]  Yu Yu.  The stability

of a dynamic duopoly Cournot–Bertrand game model. Journal of Computational and Applied Mathematics, 413(??):??, October 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001832.

Yucel:2021:GOP

[Yüc21] Hamdullah Yücel. Goaloriented a posteriori error estimation for Dirichlet boundary control problems. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303034.

Yan:2024:EDH

[YVXX24] Fengna Yan, J. J. W.

Van der Vegt, Yinhua Xia, and Yan Xu. Entropy dissipative higher order accurate positivity preserving time-implicit discretizations for nonlinear degenerate parabolic equations. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006180.

Yu:2023:ECP

[YW23] Changjun Yu and Kar Hung Wong. An enhanced control parameterization technique with variable switching times for constrained optimal control problems with control-dependent time-delayed arguments and discrete time-delayed arguments. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300050X.

Yang:2019:TEM

[YWKM19] Hao Yang, Fuke Wu, Peter E. Kloeden, and Xuerong Mao. The truncated Euler–Maruyama method for stochastic differential equations with Hölder diffusion coefficients. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719303826.

Yuan:2024:MTP

[YWL24] Gonglin Yuan, Zhan Wang, and Pengyuan Li. A modified two-parameter scaled Broyden-type algorithm for unconstrained optimization problems. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001948.

Yu:2019:OMP

[YWTL19] Xing Yu, Zhongkai Wan, Xiaowen Tu, and Yanyin Li. The optimal multiperiod hedging model of currency futures and options with exponential utility. Journal of Computational and Applied Mathematics, 366(??):??, ???? 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304157.

Yu:2024:OMW

[YWW24] Jiao Yu, Qinyi Wang, and Chunjie Wu. Online monitoring of the Weibull distributed process based on progressive type II censoring scheme. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006866.

Yao:2024:MDH

[YWZ24] Guoqing Yao, Zhongqing Wang, and Chao Zhang. A multi-domain hybrid spectral collocation method for nonlinear Volterra integral equations with weakly singular kernel. Journal of Computational and Applied Mathematics, 444 (??):??, July 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000347.

Yan:2020:SAE

[YX20] Fengna Yan and Yan Xu. Stability analysis and error estimates of local discontinuous Galerkin methods with semi-implicit spectral deferred correction timemarching for the Allen– Cahn equation. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301485.

Yu:2021:NAD

[YY21] Qiang Yu and Jiangling Yan. A novel average dwell time strategy for stability analysis of discrete-time switched systems by T-S fuzzy modeling. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305975.

Yang:2024:CAD

[YY24a] Jinting Yang and Nianyu Yi. Convergence analysis of a decoupled pressurecorrection SAV-FEM for the Cahn–Hilliard–Navier– Stokes model. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002358.

Yuan:2024:GST

[YY24b] Long Yuan and Xiaoqiang Yue. A global spacetime Trefftz DG scheme for the time-dependent isotropic elastic wave equations. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272400181X.

Ye:2024:VCN

[YYBL24] Changlun Ye, Tingfu Yao, Hai Bi, and Xianbing Luo. A variational Crank– Nicolson ensemble Monte Carlo algorithm for a heat equation under uncertainty. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003182.

Yang:2024:OEE

[YYC24] Jinting Yang, Nianyu Yi, and Yaoyao Chen. Optimal error estimates of a SAV-FEM for the Cahn– Hilliard–Navier–Stokes model. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723005216.

Yu:2024:ETR

[YYT24] Zhensheng Yu, Yue Yuan, and Panjie Tian. An efficient trust region algorithm with bounded iteration sequence for unconstrained optimization and its application in support vector machine. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002061.

Yang:2021:SCA

[YYY21] Zhanwen Yang, Huizi Yang, and Zichen Yao. Strong convergence analysis for Volterra integrodifferential equations with fractional Brownian motions. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304477.

Yu:2024:GMB

[YYZB24] Dongmei Yu, Yifei Yuan, Yiming Zhang, and Pan Bao. Generalized modulusbased matrix splitting algorithm with Anderson acceleration strategy for vertical linear complementarity problems. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000128.

Yan:2020:NAC

[YZ20a] Xiaoqiang Yan and Chengjian Zhang. Numerical approximation to a class of nonlinear hybrid system with distributed delay via block boundary value methods. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302338.

Ye:2020:SFW

[YZ20b] Xiu Ye and Shangyou Zhang. A stabilizer-free weak Galerkin finite element method on polytopal meshes. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307046.

Yang:2021:SCE

[YZ21a] Xu Yang and Weidong Zhao. Strongly convergent error analysis for a spatially semidiscrete approximation of stochastic partial differential equations with non-globally Lipschitz continuous coefficients. Journal of Computational and Applied Mathematics, 384(??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304647.

Ye:2021:SFWb

[YZ21b] Xiu Ye and Shangyou Zhang. A stabilizer free weak Galerkin finite element method on polytopal mesh: Part II. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001448.

Ye:2021:SFWa

[YZ21c] Xiu Ye and Shangyou Zhang. A stabilizer free weak Galerkin finite element method on polytopal mesh: Part III. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001576.

Yue:2021:SCS

[YZ21d] Chao Yue and Longbin Zhao. Strong convergence of the split-step backward Euler method for stochastic delay differential equations with a nonlinear diffusion coefficient. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303782.

Yang:2023:TCF

[YZ23] Hongli Yang and Xianyang Zeng. The tri-coloured free-tree theory for symplectic multi-frequency ERKN methods. Journal of Computational and Applied Mathematics, 423 (??):??, May 15, 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722005623.

Yang:2024:HSA

[YZ24a] Zhen-Ping Yang and Yong Zhao. Hybrid SGD algorithms to solve stochastic composite optimization problems with application in sparse portfolio selection problems. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723003692.

Ye:2024:FOS

[YZ24b] Xiu Ye and Shangyou Zhang. Four-order superconvergent CDG finite elements for the biharmonic equation on triangular meshes. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004600.

Yuan:2024:SSS

[YZ24c] Haiyan Yuan and Quanxin Zhu. Some stabilities of stochastic differential equations with delay in the G-framework and Euler–Maruyama method. Journal of Computational and Applied Mathematics, 446(??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001055.

Yuan:2024:LTD

[YZ24d] Wanqiu Yuan and Chengjian Zhang. Long-term dynamics of a stabilized timespace discretization scheme for 2D time-fractional Allen–Cahn equation with double well potential. Journal of Computational and Applied Mathematics, 448(??):??, October 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002024.

Yang:2024:MGB

[YZL24] Huanfeng Yang, Shan Zhao, and Guangqing Long. A MAC grid based FFT-AMIB solver for incompressible Stokes flows with interfaces and singular forces. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002693.

Yuan:2022:AAS

[YZLL22] Yongxin Yuan, Huiting Zhang, Lina Liu, and Hao Liu. An alternative approach for solving the quadratic matrix equation C^{∗}X^{∗}AXC + C^{∗}X^{∗}B + B^{∗}XC + D = 0. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100532X.

Yang:2020:RSI

[YZM^{+}20] Jing-Hua Yang, Xi-Le Zhao, Tian-Hui Ma, Yong Chen, and Meng Ding. Remote sensing images destriping using unidirectional hybrid total variation and nonconvex low-rank regularization. Journal of Computational and Applied Mathematics, 363(??):124–144, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302766.

Yi:2023:SPD

[YZM23] Lijun Yi, Mingzhu Zhang, and Xinyu Mao. Superconvergent postprocessing of the discontinuous Galerkin time stepping method for nonlinear Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000845.

Yan:2021:IFO

[YZW21] X. B. Yan, Y. X. Zhang, and T. Wei. Identify the fractional order and diffusion coefficient in a fractional diffusion wave equation. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001163.

Ye:2020:NPW

[YZZ20a] Xiu Ye, Shangyou Zhang, and Zhimin Zhang. A new P_{1} weak Galerkin method for the biharmonic equation. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303346.

Ye:2020:DGL

[YZZ20b] Xiu Ye, Shangyou Zhang, and Peng Zhu. A discontinuous Galerkin leastsquares method for div– curl systems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304777.

Ye:2022:DLM

[YZZ22] Xiu Ye, Shangyou Zhang, and Peng Zhu. Development of a LDG method on polytopal mesh with optimal order of convergence. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722000589.

Zanella:2023:TOS

[ZAC^{+}23] Alice Zanella, Luca Abergo, Francesco Caccia, Myles Morelli, and Alberto Guardone. Towards an open-source framework for fluid-structure interaction using SU2, MBDyn and preCICE. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001553.

Zagar:2023:ALP

[Zag23] Emil Zagar. Arc length preserving G^{2} Hermite interpolation of circular arcs. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006069.

Zaman:2024:NDW

[ZAH^{+}24] U. H. M. Zaman, Mohammad Asif Arefin, Md. Akram Hossain, M. Ali Akbar, and M. Hafiz Uddin. Nonlinear dynamic wave characteristics of optical soliton solutions in ion-acoustic wave. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002930.

Zheng:2023:FDD

[ZaZWZ23] Zhihui Zheng, Guang an Zou, Bo Wang, and Wenju Zhao. A fullydecoupled discontinuous Galerkin method for the nematic liquid crystal flows with SAV approach. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001516.

Zhang:2022:ESN

[ZB22] Xiaolong Zhang and John P. Boyd. Exact solutions to a nonlinear partial differential equation: the productof-curvatures Poisson (u_{xx}u_{yy} = 1). Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004891.

Zhan:2024:QSC

[ZBH24] Xuzhou Zhan, Bohui Ban, and Yongjian Hu. On the quasi-stability criteria of monic matrix polynomials. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005046.

Zeinali:2022:RHO

[ZBS22] Masoumeh Zeinali, Fariba Bahrami, and Sedaghat Shahmorad. Recursive higher order fuzzy transform method for numerical solution of Volterra integral equation with singular and nonsingular kernels. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004775.

Zhang:2020:URP

[ZC20a] Guo-Dong Zhang and Chuanjun Chen. Uniformly robust preconditioners for incompressible MHD system. Journal of Computational and Applied Mathematics, 379(??):??, December 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302053.

Zhao:2020:ANU

[ZC20b] Lina Zhao and Eric T. Chung. An analysis of the NLMC upscaling method for high contrast problems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304832.

Zhai:2021:MPA

[ZC21] Fang-Man Zhai and Li-Qun Cao. A multiscale parallel algorithm for dualphase-lagging heat conduction equation in composite materials. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303150.

Zhang:2022:ANL

[ZC22] Jingyan Zhang and Siu Wun Cheung. Analysis of nonlocal multicontinuum upscaling for dual continuum model. Journal of Computational and Applied Mathematics, 406(??):??, May 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004969.

Zhang:2024:EED

[ZC24] Xinyue Zhang and Waixiang Cao. Error estimates of the direct discontinuous Galerkin method for two-dimensional nonlinear convection–diffusion equations: Superconvergence analysis. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003650.

Zhou:2024:GPP

[ZCCK24] Xiaolin Zhou, Gang Cai, Prasit Cholamjiak, and Suparat Kesornprom. A generalized proximal point algorithm with new step size update for solving monotone variational inequalities in real Hilbert spaces. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004624.

Zhang:2022:UPP

[ZCF^{+}22] Chun-Hua Zhang, Guang-Ze Chen, Zhi-Wei Fang, Xue lei Lin, and Hai-Wei Sun. An unconditionally positivity-preserving implicit–explicit scheme for evolutionary stable distribution model. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005069.

Zhang:2020:OLE

[ZCJL20] Jiannan Zhang, Ping Chen, Zhuo Jin, and Shuanming Li. Open-loop equilibrium strategy for mean-variance asset-liability management portfolio selection problem with debt ratio. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302429.

Zhang:2021:MAA

[ZCSL21] Yongwei Zhang, Liqun Cao, Dongyang Shi, and Yanping Lin. Multiscale analysis and algorithm of transient electromagnetic scattering from heterogeneous materials. Journal of Computational and Applied Mathematics, 391(??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000467.

Zhou:2024:MNB

[ZCSS24] Chen-Can Zhou, Yang Cao, Qin-Qin Shen, and Quan Shi. A modified Newton-based matrix splitting iteration method for generalized absolute value equations. Journal of Computational and Applied Mathematics, 442(??):??, May 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006891.

Zhou:2021:CPA

[ZCT21] Mengmeng Zhou, Jianlong Chen, and Néstor Thome. Characterizations and perturbation analysis of a class of matrices related to core-EP inverses. Journal of Computational and Applied Mathematics, 393(??):??, September 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721001151.

Zhang:2020:ENI

[ZCY^{+}20] Jun Zhang, Chuanjun Chen, Xiaofeng Yang, Yuchuan Chu, and Zeyu Xia. Efficient, noniterative, and secondorder accurate numerical algorithms for the anisotropic Allen–Cahn equation with precise nonlocal mass conservation. Journal of Computational and Applied Mathematics, 363(??):444–463, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302377.

Zhang:2020:ENS

[ZCYP20] Jun Zhang, Chuanjun Chen, Xiaofeng Yang, and Kejia Pan. Efficient numerical scheme for a penalized Allen–Cahn type Ohta–Kawasaki phase-field model for diblock copolymers. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720301965.

Zhong:2023:EVG

[ZCZ23] Wei Zhong, Zhenyu Cui, and Zhimin Zhang. Efficient valuation of guaranteed minimum maturity benefits in regime switching jump diffusion models with surrender risk. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200512X.

Zheng:2024:LSM

[ZCZ24] Weishan Zheng, Yanping Chen, and Jianwei Zhou. A Legendre spectral method for multidimensional partial Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002467.

Zhan:2021:GCH

[ZD21] Xuzhou Zhan and Alexander Dyachenko. On generalization of classical Hurwitz stability criteria for matrix polynomials. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304040.

Zlatev:2020:ANM

[ZDF^{+}20] Zahari Zlatev, Pasqua D’Ambra, István Faragó, Vladimir Shaydurov, and Lubin Vulkov. Advanced numerical methods for complex scientific and engineering problems: editorial introduction. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306016.

Zhang:2022:GIA

[ZDL22] Jicheng Zhang, Quanling Deng, and Xin Li. A generalized isogeometric analysis of elliptic eigenvalue and source problems with an interface. Journal of Computational and Applied Mathematics, 407 (??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006178.

Zhang:2024:CLM

[ZDW24] Kuokuo Zhang, Weibing Deng, and Haijun Wu. A CutFE–LOD method for the multiscale elliptic problems on complex domains. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000694.

Zhao:2020:URO

[ZDY20] Xinzhu Zhao, Bo Dong, and Bo Yu. Unfoldings and the rank-one approximation of the tensor. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300856.

Zamanpour:2023:OMM

[ZE23] I. Zamanpour and R. Ezzati. Operational matrix method for solving fractional weakly singular 2D partial Volterra integral equations. Journal of Computational and Applied Mathematics, 419(??):??, February 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003569.

Zheng:2023:APG

[ZEW23] Xiangcheng Zheng, V. J. Ervin, and Hong Wang. Analysis and Petrov– Galerkin numerical approximation for variable coefficient two-sided fractional diffusion, advection, reaction equations. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006318.

Zhang:2024:NIQ

[ZF24] Yun Zhang and Xiaoli Feng. A nonstationary iterated quasi-boundary value method for reconstructing the source term in a time-space fractional diffusion equation. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005563.

Zheng:2020:MML

[ZFH20] Sanpeng Zheng, Renzhong Feng, and Aitong Huang. A modified moving leastsquares suitable for scattered data fitting with outliers. Journal of Computational and Applied Mathematics, 370(??):??, May 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306600.

Zhao:2022:ECM

[ZFM22] Longbin Zhao, Qiongqi Fan, and Wanyuan Ming. Efficient collocation methods for Volterra integral equations with highly oscillatory kernel. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004945.

Zhao:2022:HAO

[ZFW22] Longbin Zhao, Qiongqi Fan, and Sheng Wang. High asymptotic order methods for highly oscillatory integral equations with trigonometric kernels. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722002709.

Zhang:2020:IPK

[ZFwC20] Liping Zhang, Hung-Yuan Fan, and Eric King wah Chu. Inheritance properties of Krylov subspace methods for continuoustime algebraic Riccati equations. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306909.

Zhang:2020:SOA

[ZG20a] Ye Zhang and Rongfang Gong. Second order asymptotical regularization methods for inverse problems in partial differential equations. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300893.

Zhang:2020:GFT

[ZG20b] Yue Zhang and Wenhao Gui. A goodness of fit test for the Pareto distribution with progressively type II censored data based on the cumulative hazard function. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930562X.

Zhang:2022:EBP

[ZG22] Juan Zhang and Liang Ge.

Efficient block preconditioners for integral constrained elliptic optimal control problems with finite element approximations. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721006002.

Zhang:2024:LSW

[ZG24] Rui Zhang and Yu Gao.

Learning scattering waves via coupling physics-informed neural networks and their convergence analysis. Journal of Computational and Applied Mathematics, 446 (??):??, August 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724001237.

Zocca:2023:NDA

[ZGG23] Marta Zocca, Paolo Gajoni, and Alberto Guardone. NI-MOC: a design and analysis tool for supersonic nozzles under non-ideal compressible flow conditions. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001541.

Zhang:2021:CAC

[ZGJ21] Shan Zhang, Xiaofei Guan, and Lijian Jiang. Convergence analysis of constraint energy minimizing generalized multiscale finite element method for a linear stochastic parabolic partial differential equation driven by additive noises. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720306191.

Zhu:2020:TCN

[ZGL20] Jiaqi Zhu, Guohui Guan, and Shenghong Li. Timeconsistent non-zero-sum stochastic differential reinsurance and investment game under default and volatility risks. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300285.

Zhang:2023:IMB

[ZGZ23] Zhengqiang Zhang, Shimin Guo, and Yuan-Xiang Zhang. An iterative method based on Nesterov acceleration for identifying space-dependent source term in a timefractional diffusion-wave equation. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001589.

Zulawinski:2023:IVP

[ZGZW23] Wojciech Zuławiński, Aleksandra Grzesiek, Radosław Zimroz, and Agnieszka Wyłomańska. Identification and validation of periodic autoregressive model with additive noise: finitevariance case. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000754.

Zhan:2023:RBH

[ZH23] Xuzhou Zhan and Yongjian Hu. On the relation between Hurwitz stability of matrix polynomials and matrix-valued Stieltjes functions. Journal of Computational and Applied Mathematics, 417(??):??, January 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003041.

Zhao:2024:NMH

[ZH24a] Longbin Zhao and Chengming Huang. Numerical methods for highly oscillatory Volterra integral equations with general oscillators. Journal of Computational and Applied Mathematics, 449(??):??, October 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002176.

Zhong:2024:AMR

[ZH24b] Shuangyou Zhong and Guangxin Huang. An almost-maximal residual tensor block Kaczmarz method for large tensor linear systems. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003837.

Zhou:2024:ELS

[ZH24c] Yiqing Zhou and Daan Huybrechs. Efficient least squares approximation and collocation methods using radial basis functions. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001195.

Zhang:2020:ORC

[Zha20a] Huaping Zhang. Optimization of risk control in financial markets based on particle swarm optimization algorithm. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305357.

Zhang:2020:IRM

[Zha20b] Jiboning Zhang. Investment risk model based on intelligent fuzzy neural network and VaR. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719307125.

Zhao:2020:ALD

[Zha20c] Yan-Yong Zhao. Analysis of longitudinal data with semiparametric varyingcoefficient mean-covariance models. Journal of Computational and Applied Mathematics, 363(??):485–502, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719301487.

Zhang:2021:DBS

[Zha21a]  Ning Zhang.  A dual

based semismooth Newtontype algorithm for solving large-scale sparse Tikhonov regularization problems. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002636.

Zhang:2021:PMU

[Zha21b] Wenxing Zhang. A phase model using the Huber norm for estimating point spread function under frozen flow hypothesis. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272100279X.

Zhang:2021:NMS

[Zha21c] Xian-Min Zhang. A new method for searching the integral solution of system of Riemann–Liouville fractional differential equations with non-instantaneous impulses. Journal of Computational and Applied Mathematics, 388(??):??, May 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720305987.

Zhang:2024:NEA

[Zha24] Xiaodi Zhang. New error analysis and recovery technique of a class of fully discrete finite element methods for the dynamical inductionless MHD equations. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723004855.

Zhao:2022:IPB

[ZHMD22] Jing Zhao, Jiahong He, Stanisław Migórski, and Sylwia Dudek. An inverse problem for Bingham type fluids. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S037704272100529X.

Zhou:2020:MBT

[Zho20] Weijun Zhou. A modified BFGS type quasi-Newton method with line search for symmetric nonlinear equations problems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719304571.

Zhou:2022:UCR

[Zho22]  Xingcai Zhou.  Uni-

form convergence rates for wavelet curve estimation in sup-norm loss. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003745.

Zhou:2024:BSS

[Zho24]  Zhiyong Zhou.  Block

sparse signal recovery via minimizing the block qratio sparsity. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005101.

Zhu:2020:CSC

[Zhu20a]  Ling Zhu.  A class

of strongly completely monotonic functions related to gamma function. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304728.

Zhu:2020:SIW

[Zhu20b] Tiefeng Zhu. Statistical inference of Weibull distribution based on generalized progressively hybrid censored data. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307101.

Zhu:2021:CAN

[Zhu21] Jinghao Zhu. A computational approach to nonsmooth optimization by diffusion equations. Journal of Computational and Applied Mathematics, 384 (??):??, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272030457X.

Zhu:2022:RES

[Zhu22]  Tiefeng Zhu.  Reliabil-

ity estimation of s-out-of-k system in a multicomponent stress-strength dependent model based on copula function. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005434.

Zhu:2023:IRM

[Zhu23] Tiefeng Zhu. Inference of reliability in a multicomponent stress-strength model under generalized progressive hybrid censoring. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722003004.

Zhu:2024:RIM

[Zhu24]  Tiefeng Zhu.  Reliabil-

ity inference for multicomponent stress-strength model under generalized progressive hybrid censoring. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002656.

Zhu:2020:AAB

[ZHW20] Shengfeng Zhu, Xianliang Hu, and Qingbiao Wu. On accuracy of approximate boundary and distributed H^{1} shape gradient flows for eigenvalue optimization. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303772.

Zhai:2020:PTS

[ZHWD20] Shuying Zhai, Langyang Huang, Zhifeng Weng, and Weizhong Dai. Parabolic two-step model and accurate numerical scheme for nanoscale heat conduction induced by ultrashortpulsed laser heating. Journal of Computational and Applied Mathematics, 369 (??):??, May 1, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305965.

Zhang:2020:FDL

[ZHY20] Guo-Dong Zhang, Xiaoming He, and Xiaofeng Yang. Fully decoupled, linear and unconditionally energy stable time discretization scheme for solving the magnetohydrodynamic equations. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042719306417.

Zhou:2024:NMS

[ZHZ24] Hao Zhou, Yaozhong Hu, and Jingjun Zhao. Numerical method for singular drift stochastic differential equation driven by fractional Brownian motion. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724001523.

Zhao:2021:SDH

[ZIZ21] Tengjin Zhao, Kazufumi Ito, and Zhiyue Zhang. Semi-decoupling hybrid asymptotic and augmented finite volume method for nonlinear singular interface problems. Journal of Computational and Applied Mathematics, 396(??):??, November 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002260.

Zhang:2024:CVA

[ZJ24] Xuehan Zhang and Lijian Jiang. Conditional variational autoencoder with Gaussian process regression recognition for parametric models. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004764.

Zhao:2023:GGS

[ZJBY23] Zhi Zhao, Xiao-Qing Jin, Zheng-Jian Bai, and Teng-Teng Yao. A generalized geometric spectral conjugate gradient algorithm for finding zero of a monotone tangent vector field on a constant curvature Hadamard manifold. Journal of Computational and Applied Mathematics, 422 (??):??, April 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004800.

Zhang:2020:ESP

[ZJGZ20] Jun Zhang, Maosheng Jiang, Yuezheng Gong, and Jia Zhao. Energystable predictor-corrector schemes for the Cahn– Hilliard equation. Journal of Computational and Applied Mathematics, 376 (??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042720301230.

Zhao:2022:RTS

[ZJY22] Zhi Zhao, Xiao-Qing Jin, and Teng-Teng Yao. The Riemannian two-step perturbed Gauss–Newton method for least squares inverse eigenvalue problems. Journal of Computational and Applied Mathematics, 405 (??):??, May 15, 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005707.

Zhang:2020:IFB

[ZK20] Zhiguo Zhang and Mark A.

Kon. Interpolatory filter banks and interpolatory wavelet packets. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300467.

Zuniga:2021:GLM

[ZKP21] Francesco Zuniga, Tomasz J.

Kozubowski, and Anna K. Panorska. A generalized linear model for multivariate events. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721002776.

Zhang:2020:TPN

[ZKT20] Juan Zhang, Huihui Kang, and Fangyuan Tan. Twoparameters numerical methods of the non-symmetric algebraic Riccati equation. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720302247.

Zhang:2022:BBR

[ZL22] Yaling Zhang and Hongwei Liu. A Barzilai and Borwein regularization feasible direction algorithm for convex nonlinear SOC programming with linear constraints. Journal of Computational and Applied Mathematics, 401(??):1– 15, February 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721004039.

Zheng:2023:GFM

[ZL23a] Quan Zheng and Zhongli Liu. On general families of multipoint iterations by inverse interpolation and their applications. Journal of Computational and Applied Mathematics, 426 (??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000407.

Zhou:2023:IMM

[ZL23b] Xiaojian Zhou and Baoqing Liu. Iterative methods for multiple roots with memory using selfaccelerating technique. Journal of Computational and Applied Mathematics, 428 (??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723001255.

Zhang:2024:GCN

[ZL24] Yanjun Zhang and Hanyu Li. Greedy capped nonlinear Kaczmarz methods. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003170.

Zhou:2024:DGM

[ZLC24] Mingbo Zhou, Rui Li, and Zhangxin Chen. A discontinuous Galerkin method for a coupled Stokes–Biot problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003352.

Zhang:2022:NSE

[ZLCT22] Chao Zhang, Zhipeng Liu, Sheng Chen, and DongYa Tao. New spectral element method for Volterra integral equations with weakly singular kernel. Journal of Computational and Applied Mathematics, 404 (??):??, April 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005252.

Zhang:2020:TNA

[ZLG20] Wei Zhang, Hui Liang, and Jianfang Gao. Theoretical and numerical analysis of the Euler– Maruyama method for generalized stochastic Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S037704271930367X.

Zhang:2022:SMT

[ZLJT22] Hui Zhang, Fawang Liu, Xiaoyun Jiang, and Ian Turner. Spectral method for the two-dimensional time distributed-order diffusionwave equation on a semiinfinite domain. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003344.

Zhang:2020:EIS

[ZLL20a] Ling Zhang, Danping Li, and Yongzeng Lai. Equilibrium investment strategy for a defined contribution pension plan under stochastic interest rate and stochastic volatility. Journal of Computational and Applied Mathematics, 368 (??):??, April 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305412.

Zhuang:2020:SBC

[ZLL^{+}20b] Dan Zhuang, Youbo Liu, Shuangzhe Liu, Tiefeng Ma, and Seng huat Ong. A shape-based cutting and clustering algorithm for multiple change-point detection. Journal of Computational and Applied Mathematics, 369(??):??, May 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306284.

Zhang:2023:EES

[ZLS23] Fan Zhang, Dongfang Li, and Hai-Wei Sun. Efficient and energy stable numerical schemes for the two-mode phase field crystal equation. Journal of Computational and Applied Mathematics, 427(??):??, August 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723000924.

Zheng:2020:IMN

[ZLV20] Hua Zheng, Wen Li, and Seakweng Vong. An iteration method for nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719306867.

Zhuang:2024:UVR

[ZLZ^{+}24] Jun-Hao Zhuang, Yi-Si Luo, Xi-Le Zhao, Tai-Xiang Jiang, Yi Chang, and Jun Liu. Unsupervised video rain streaks removal with deep foregroundbackground modeling. Journal of Computational and Applied Mathematics, 436 (??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723003758.

Zhang:2023:FIS

[ZLZL23] N. Zhang, J. K. Liu, L. Q.

Zhang, and Z. L. Lu. A fast inertial self-adaptive projection based algorithm for solving large-scale nonlinear monotone equations. Journal of Computational and Applied Mathematics, 426(??):??, July 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000316.

Zemlyanova:2020:NBS

[ZM20] Anna Y. Zemlyanova and Alexia Machina. A new Bspline collocation method for singular integro-differential equations of higher orders. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302405.

Zheng:2021:UNI

[ZM21a] Guang-Hui Zheng and Zhi-Qiang Miao. On uniqueness and nonuniqueness for internal potential reconstruction in quantum fields from one measurement. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303204.

Zou:2021:FGM

[ZM21b] Qinmeng Zou and Frédéric Magoulès. Fast gradient methods with alignment for symmetric linear systems without using Cauchy step. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720303241.

Zhou:2020:ATL

[ZMFL20] Xinchen Zhou, Zhaoliang Meng, Xin Fan, and Zhongxuan Luo. Analysis of two low-order equalorder finite element pairs for Stokes equations over quadrilaterals. Journal of Computational and Applied Mathematics, 364(??):??, January 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303206.

Zaman:2023:CGS

[ZMOW23] Tareq Uz Zaman, Scott P.

MacLachlan, Luke N. Olson, and Matthew West. Coarse-grid selection using simulated annealing. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002078.

Zhang:2023:NNM

[ZMYL23] Ping Zhang, Pinchao Meng, Weishi Yin, and Hongyu Liu. A neural network method for timedependent inverse source problem with limitedaperture data. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200440X.

Zhou:2024:OBB

[ZMZ^{+}24] Yan Zhou, Ruoxi Mei, Yichuan Zhao, Zongliang Hu, and Mingtao Zhao. Orthogonality-based biascorrected empirical likelihood inference for partial linear varying coefficient EV models with longitudinal data. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723006933.

Zhou:2021:CRE

[ZOK21] Guanyu Zhou, Issei Oikawa, and Takahito Kashiwabara. The Crouzeix–Raviart element for the Stokes equations with the slip boundary condition on a curved boundary. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304143.

Zhang:2021:HAL

[ZOW^{+}21] Chenhui Zhang, Jie Ouyang, Xiaodong Wang, Shuke Li, and Jiaomin Mao. Highly accurate, linear, and unconditionally energy stable large time-stepping schemes for the functionalized Cahn–Hilliard gradient flow equation.

Journal of Computational and Applied Mathematics, 392(??):??, August 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721000984.

Zhu:2024:DSO

[ZPW^{+}24] Wenxing Zhu, Mingyang Pan, Qinghe Wang, Fengyu Jiao, and Dongdong He. Decoupled second-order energy stable scheme for an electrohydrodynamic model with variable electrical conductivity. Journal of Computational and Applied Mathematics, 438(??):??, March 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004740.

Zhou:2024:HRI

[ZQL24] Guoyong Zhou, Zhijian Qiu, and Sheng Li. A hybrid reinsurance-investment game with delay and asymmetric information. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004351.

Zhao:2023:WGC

[ZR23] Jijing Zhao and Hongxing Rui. Weak Galerkin coupled with conforming finite element method for hybrid-dimensional fracture model. Journal of Computational and Applied Mathematics, 418(??):??, January 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722003703.

Zhang:2024:CGM

[ZR24a] Jing Zhang and Hongxing Rui. A coupling of Galerkin and mixed finite element methods for the quasi-static thermoporoelasticity with nonlinear convective transport. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723006167.

Zhu:2024:PGI

[ZR24b] Na Zhu and Hongxing Rui. A Petrov– Galerkin immersed finite element method for steady Navier–Stokes interface problem with nonhomogeneous jump conditions. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724000645.

Zhang:2022:SAG

[ZRQ22] Mimi Zhang, Matthew Revie, and John Quigley. Saddlepoint approximation for the generalized inverse Gaussian Lévy process. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001133.

Zhang:2020:GSM

[ZS20a] Fode Zhang and Yimin Shi. Geometry on the statistical manifold induced by the degradation model with soft failure data. Journal of Computational and Applied Mathematics, 363(??):211–222, January 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719302754.

Zhu:2020:NAN

[ZS20b] Lin Zhu and Qin Sheng.

A note on the adaptive numerical solution of a Riemann–Liouville space– fractional Kawarada problem. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300054.

Zhou:2022:BBV

[ZS22] Yongtao Zhou and Martin Stynes. Block boundary value methods for solving linear neutral Volterra integro-differential equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 401(??):??, February 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003691.

Zhang:2021:EPA

[ZSCH21] Chun Zhang, Yongzhong Song, Xingju Cai, and Deren Han. An extended proximal ADMM algorithm for three-block nonconvex optimization problems. Journal of Computational and Applied Mathematics, 398(??):??, December 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003034.

Zhang:2020:NAP

[ZSH20] Yuhong Zhang, Li Shan, and Yanren Hou. New approach to prove the stability of a decoupled algorithm for a fluid-fluid interaction problem. Journal of Computational and Applied Mathematics, 371 (??):??, June 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719307009.

Zhang:2024:EBB

[ZSHD24] Yanzhen Zhang, Peiping Shen, Bingdi Huang, and Yaping Deng. An efficient branch-and-bound algorithm using an adaptive branching rule with quadratic convex relaxation for globally solving general linear multiplicative programs. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002851.

Zhang:2022:CSS

[ZSL22] Yuhang Zhang, Minghui Song, and Mingzhu Liu. Convergence and stability of stochastic theta method for nonlinear stochastic differential equations with piecewise continuous arguments. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004726.

Zhang:2023:SCT

[ZSL23] Yuhang Zhang, Minghui Song, and Mingzhu Liu. Strong convergence of the tamed Euler method for nonlinear hybrid stochastic differential equations with piecewise continuous arguments. Journal of Computational and Applied Mathematics, 429(??):??, September 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001413.

Zambrano:2021:FMC

[ZSLG21] Miguel Zambrano, Sintya Serrano, Boyan S. Lazarov, and Juan Galvis. Fast multiscale contrast independent preconditioners for linear elastic topology optimization problems. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306579.

Zhang:2020:FBM

[ZSMG20] Deyue Zhang, Fenglin Sun, Yan Ma, and Yukun Guo. A Fourier–Bessel method with a regularization strategy for the boundary value problems of the Helmholtz equation. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719305679.

Zhao:2020:NSS

[ZSVL20] Meng Zhao, Gerard Salter, Vaughan Voller, and Shuwang Li. Nonlinear simulation of shape-preserving delta growth. Journal of Computational and Applied Mathematics, 380(??):??, December 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302582.

Zaman:2024:EHR

[ZSY24] Tolga Zaman, Usman Shazad, and Vinay Kumar Yadav. An efficient Hartley-Ross type estimators of nonsensitive and sensitive variables using robust regression methods in sample surveys. Journal of Computational and Applied Mathematics, 440 (??):??, April 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005897.

Zhong:2024:VGM

[ZSZ24] Wei Zhong, Benxuan Shi, and Zhimin Zhang. Valuation of guaranteed minimum maturity benefits under mean reversion and jump models with surrender risk. Journal of Computational and Applied Mathematics, 440(??):??, April 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723005903.

Zolfaghari:2021:SAI

[ZTS21] Reza Zolfaghari, Jacob Taylor, and Raymond J. Spiteri. Structural analysis of integro-differential– algebraic equations. Journal of Computational and Applied Mathematics, 394 (??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721001886.

Zhang:2020:MZO

[ZTY^{+}20] Chi Zhang, Guo-Liang Tian, Kam Chuen Yuen, Qin Wu, and Tao Li. Multivariate zero-and-one inflated Poisson model with applications. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303590.

Zhou:2024:CCC

[ZTY^{+}24] Jinyu Zhou, Yu Tang, Jigao Yan, Tianjiao Yan, and Jun Gu. Complete convergence and complete moment convergence for maximal weighted sums of arrays of rowwise extended negatively dependent random variables with statistical applications. Journal of Computational and Applied Mathematics, 437(??):??, February 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723004302.

Zhai:2020:SWG

[ZTZZ20] Qilong Zhai, Tian Tian, Ran Zhang, and Shangyou Zhang. A symmetric weak Galerkin method for solving non-divergence form elliptic equations. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719306983.

Zaman:2020:AHO

[ZuIH20] Sakhi Zaman, Siraj ul Islam, and Iqrar Hussain. Approximation of highly oscillatory integrals containing special functions. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303759.

Zaman:2022:NAA

[ZuIKA22] Sakhi Zaman, Siraj ul Islam, Muhammad Munib Khan, and Imtiaz Ahmad. New algorithms for approximation of Bessel transforms with high frequency parameter. Journal of Computational and Applied Mathematics, 399(??):??, January 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003277.

Zhang:2020:NAF

[ZUZ20] Jinghui Zhang, Salamat Ullah, and Yang Zhong. New analytical free vibration solutions of orthotropic rectangular thin plates using generalized integral transformation. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704271930442X.

Zielinski:2021:CSM

[ZVS21] Przemysław Zieliński, Hannes Vandecasteele, and Giovanni Samaey. Convergence and stability of a micro-macro acceleration method: Linear slow-fast stochastic differential equations with additive noise. Journal of Computational and Applied Mathematics, 387(??):??, ???? 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304935.

Zegeling:2023:GMB

[ZvS23] P. A. Zegeling and M. W. F.

van Spengler. A generalized midpoint-based boundary value method for unstable partial differential equations. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006124.

Zaky:2023:LTD

[ZVT^{+}23] M. A. Zaky, K. Van Bockstal, T. R. Taha, D. Suragan, and A. S. Hendy. An L1 type difference/Galerkin spectral scheme for variableorder time-fractional nonlinear diffusion–reaction equations with fixed delay. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004307.

Zhang:2020:RCR

[ZW20] Liping Zhang and Yimin Wei. Randomized core reduction for discrete illposed problem. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300881.

Zhao:2021:RTI

[ZW21] Youyi Zhao and Weiwei Wang. On the Rayleigh– Taylor instability in compressible viscoelastic fluids under L^{1}-norm. Journal of Computational and Applied Mathematics, 383(??):??, February 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720304210.

Zhou:2022:FQF

[ZW22] Yanhui Zhou and Jiming Wu. A family of quadratic finite volume element schemes for anisotropic diffusion problems on triangular meshes. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004167.

Zhang:2023:NBA

[ZW23] Rui Zhang and Dehui Wang. A new binomial autoregressive process with explanatory variables. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722004149.

Zhang:2024:SOS

[ZW24] Xuelin Zhang and Hanquan Wang. Second-order Sobolev gradient flows for computing ground state of ultracold Fermi gases. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003455.

Zhang:2020:RBP

[ZWH^{+}20] Feng Zhang, Wendong Wang, Jianwen Huang, Jianjun Wang, and Yao Wang. RIP-based performance guarantee for lowtubal-rank tensor recovery. Journal of Computational and Applied Mathematics, 374(??):??, August 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300583.

Zhou:2022:DMP

[ZWJ22] Huifang Zhou, Xiuli Wang, and Jiwei Jia. Discrete maximum principle for the weak Galerkin method on triangular and rectangular meshes. Journal of Computational and Applied Mathematics, 402(??):??, March 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004064.

Zhu:2024:ESI

[ZWKS24] Xingyu Zhu, Xiuping Wang, Jisheng Kou, and Shuyu Sun. An energy stable incompressible SPH method with consistent solid boundary treatment. Journal of Computational and Applied Mathematics, 436(??):??, January 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723003114.

Zhou:2021:NBM

[ZWL21] Hong-Yu Zhou, Shi-Liang Wu, and Cui-Xia Li. Newton-based matrix splitting method for generalized absolute value equation. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042721001989.

Zuo:2024:QIA

[ZWX24] Qian Zuo, Yimin Wei, and Hua Xiang. Quantuminspired algorithm for truncated total least squares solution. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724002929.

Zhang:2022:RAC

[ZWY22] Yun Zhang, Ting Wei, and Xiongbin Yan. Recovery of advection coefficient and fractional order in a time-fractional reaction– advection–diffusion-wave equation. Journal of Computational and Applied Mathematics, 411(??):??, September 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722001029.

Zhai:2023:EOS

[ZWZ^{+}23a] Shuying Zhai, Zhifeng Weng, Qingqu Zhuang, Fawang Liu, and Vo Anh. An effective operator splitting method based on spectral deferred correction for the fractional Gray– Scott model. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272200557X.

Zhang:2023:MRN

[ZWZ23b] Jianhua Zhang, Yuqing Wang, and Jing Zhao. On maximum residual nonlinear Kaczmarz-type algorithms for large nonlinear systems of equations. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723000092.

Zhao:2023:HWG

[ZWZ23c] Lidan Zhao, Ruishu Wang, and Yongkui Zou. A hybridized weak Galerkin finite element scheme for linear elasticity problem. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006227.

Zheng:2023:FAB

[ZX23a] Chunxiong Zheng and Jiangming Xie. Fast artificial boundary method for the heat equation on unbounded domains with strip tails. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006306.

Zhou:2023:RWT

[ZX23b] Zhiqiang Zhou and Wei Xu. Robust willow tree method under Lévy processes. Journal of Computational and Applied Mathematics, 424(??):??, May 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005805.

Zhong:2022:TGD

[ZXC22] Liuqiang Zhong, Yue Xuan, and Jintao Cui. Two-grid discontinuous Galerkin method for convection– diffusion–reaction equations. Journal of Computational and Applied Mathematics, 404(??):??, April 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721005264.

Zhang:2023:OCR

[ZXK23] Yali Zhang, Shuhuang Xiang, and Desong Kong. On optimal convergence rates of Laguerre polynomial expansions for piecewise functions. Journal of Computational and Applied Mathematics, 425 (??):??, June 2023. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722006513.

Zhang:2024:OPE

[ZXK24] Yali Zhang, Shuhuang Xiang, and Desong Kong. Optimal pointwise error estimates for piecewise functions expanded with Laguerre polynomials. Journal of Computational and Applied Mathematics, 443 (??):??, June 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272300691X.

Zhao:2024:SSE

[ZXZ24] Fengyang Zhao, Jiangming Xie, and Chunxiong Zheng. Sparsifiable spectral equivalence of DtN mapping and its application to elliptic problems. Journal of Computational and Applied Mathematics, 452(??):??, December 15, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003637.

Zhao:2020:WSD

[ZY20] Weifeng Zhao and Wen-An Yong. Weighted L_{2}-stability of a discrete kinetic approximation for the incompressible Navier–Stokes equations on bounded domains. Journal of Computational and Applied Mathematics, 376(??):??, October 1, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720301114.

Zuo:2022:MTC

[ZY22] Baishuai Zuo and Chuancun Yin. Multivariate tail covariance risk measure for generalized skew-elliptical distributions. Journal of Computational and Applied Mathematics, 410(??):??, August 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000772.

Zhao:2024:VFD

[ZY24] Jin Zhao and Wen-An Yong. Vectorial finitedifference-based lattice Boltzmann method: Consistency, boundary schemes and stability analysis. Journal of Computational and Applied Mathematics, 441(??):??, May 15, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042723006210.

Zuo:2021:EJM

[ZYB21] Baishuai Zuo, Chuancun Yin, and Narayanaswamy Balakrishnan. Expressions for joint moments of elliptical distributions. Journal of Computational and Applied Mathematics, 391 (??):??, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042721000376.

Zheng:2020:CGP

[ZYL20] Li Zheng, Lei Yang, and Yong Liang. A conjugate gradient projection method for solving equations with convex constraints. Journal of Computational and Applied Mathematics, 375(??):??, September 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300728.

Zhang:2020:VEL

[ZYY20] Zhimin Zhang, Yaodi Yong, and Wenguang Yu. Valuing equity-linked death benefits in general exponential Lévy models. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719303802.

Zeng:2022:MBA

[ZYZ22] Min-Li Zeng, Jun-Feng Yang, and Guo-Feng Zhang. On τ matrix-based approximate inverse preconditioning technique for diagonalplus-Toeplitz linear systems from spatial fractional diffusion equations. Journal of Computational and Applied Mathematics, 407(??):??, June 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042722000048.

Zhang:2023:SCL

[ZYZ23] Jiandong Zhang, Rongfang Yan, and Yiying Zhang. Stochastic comparisons of largest claim amount from heterogeneous and dependent insurance portfolios. Journal of Computational and Applied Mathematics, 431(??):??, October 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723002091.

Zheng:2024:COF

[ZYZ24] Shixin Zheng, Haizhao Yang, and Xiangxiong Zhang. On the convergence of orthogonalization-free conjugate gradient method for extreme eigenvalues of Hermitian matrices: a Riemannian optimization interpretation. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003030.

Zhang:2020:WGF

[ZZ20a] Tie Zhang and Shangyou Zhang. The weak Galerkin finite element method for the symmetric hyperbolic systems. Journal of Computational and Applied Mathematics, 365(??):??, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/S0377042719303784.

Zhou:2020:MBL

[ZZ20b] Weijun Zhou and Li Zhang. A modified Broyden-like quasi-Newton method for nonlinear equations. Journal of Computational and Applied Mathematics, 372 (??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720300352.

Zhang:2021:MDD

[ZZ21] Kong-Sheng Zhang and Yan-Yong Zhao. Modeling dynamic dependence between crude oil and natural gas return rates: a time-varying geometric copula approach. Journal of Computational and Applied Mathematics, 386 (??):??, April 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042720305343.

Zhang:2022:HPC

[ZZ22a] Rong Zhang and Bing Zhou. Heuristic parameter choice rule for solving linear ill-posed integral equations in finite dimensional space. Journal of Computational and Applied Mathematics, 400(??):??, January 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721003630.

Zhang:2022:EWG

[ZZ22b] Tie Zhang and Shangyou Zhang. An explicit weak Galerkin method for solving the first order hyperbolic systems. Journal of Computational and Applied Mathematics, 412(??):??, October 1, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722001327.

Zhou:2023:ICM

[ZZ23] Pei Zhou and Chun-Gang Zhu. Isogeometric collocation method based on residual parameterization of planar physical domain. Journal of Computational and Applied Mathematics, 422(??):??, April 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722004873.

Zhang:2024:GSV

[ZZ24] Chuang-Liang Zhang and Shengda Zeng. Generalized set-valued Ekeland variational principles via the null set concept with applications to some setvalued optimization problems. Journal of Computational and Applied Mathematics, 450(??):??, November 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724002619.

Zhang:2020:EAF

[ZZG20] Jun Zhang, Jia Zhao, and Yuezheng Gong. Error analysis of full-discrete invariant energy quadratization schemes for the Cahn– Hilliard type equation. Journal of Computational and Applied Mathematics, 372(??):??, July 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720300108.

Zhang:2021:OEL

[ZZK21] Yan Zhang, Peibiao Zhao, and Bingyu Kou. Optimal excess-of-loss reinsurance and investment problem with thinning dependent risks under Heston model. Journal of Computational and Applied Mathematics, 382(??):??, January 15, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720303733.

Zhu:2020:SGF

[ZZL20] Pengfei Zhu, Qinghui Zhang, and Tingyun Liu. Stable generalized finite element method (SGFEM) for parabolic interface problems. Journal of Computational and Applied Mathematics, 367(??):??, March 15, 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042719304789.

Zhang:2023:LNR

[ZZLY23] Yong Zhang, Xiaolei Zhu, Aiguo Liu, and Shichao Yi. A Lorentzian-l_{p} norm regularization based algorithm for recovering sparse signals in two types of impulsive noise. Journal of Computational and Applied Mathematics, 430(??):??, October 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042723001954.

Zhao:2020:APD

[ZZM20] Jian-Qiang Zhao, Yan-Yong Zhao, and Zhang-Xiao Miao. Analysis of panel data partially linear single-index models with serially correlated errors. Journal of Computational and Applied Mathematics, 368(??):??, April 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042719305370.

Zhang:2023:EED

[ZZMK23] Daochang Zhang, Yue Zhao, Dijana Mosić, and Vasilios N. Katsikis. Exact expressions for the Drazin inverse of anti-triangular matrices. Journal of Computational and Applied Mathematics, 428(??):??, August 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042723001310.

Zhang:2024:SWB

[ZZNL24] Borong Zhang, Leonardo Zepeda-Nunez, and Qin Li. Solving the wide-band inverse scattering problem via equivariant neural networks. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724003005.

Zhao:2021:NTH

[ZZQ21] Yan-Yong Zhao, Jian-Qiang Zhao, and Su-An Qian. A new test for heteroscedasticity in singleindex models. Journal of Computational and Applied Mathematics, 381(??):??, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720302843.

Zeng:2024:LRE

[ZZW^{+}24a] Yuping Zeng, Liuqiang Zhong, Feng Wang, Mingchao Cai, and Shangyou Zhang. Low regularity error analysis for an H(div)-conforming discontinuous Galerkin approximation of Stokes problem. Journal of Computational and Applied Mathematics, 451(??):??, December 1, 2024. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042724003674.

Zeng:2024:CPR

[ZZW^{+}24b] Yuping Zeng, Liuqiang Zhong, Feng Wang, Shangyou Zhang, and Mingchao Cai. Corrigendum to “A pressure-robust numerical scheme for the Stokes equations based on the WOP-SIP DG approach” [Journal of Computational and Applied Mathematics 445 (2024) 115819]. Journal of Computational and Applied Mathematics, 447(??):??, September 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724001420.

See [ZZW^{+}24c].

Zeng:2024:PRN

[ZZW^{+}24c] Yuping Zeng, Liuqiang Zhong, Feng Wang, Shangyou Zhang, and Mingchao Cai. A pressure-robust numerical scheme for the Stokes equations based on the WOPSIP DG approach. Journal of Computational and Applied Mathematics, 445(??):??, August 1, 2024. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0377042724000682.

See corrigendum [ZZW^{+}24b].

Zhang:2020:BMC

[ZZWS20] Na Zhang, Wenting Zeng, Yuhe Wang, and Qian Sun. Beyond multiplecontinuum modeling for the simulation of complex flow mechanisms in multiscale shale porous media. Journal of Computational and Applied Mathematics, 378(??):??, November 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S037704272030145X.

Zhang:2021:TLF

[ZZwS21] Qifeng Zhang, Lu Zhang, and Hai wei Sun. A three-level finite difference method with preconditioning technique for twodimensional nonlinear fractional complex Ginzburg– Landau equations. Journal of Computational and Applied Mathematics, 389 (??):??, June 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042720306464.

Zhao:2022:DDD

[ZZWS22] Jun Zhao, Yi Zhang, Sheng Wu, and Liming Shen. Data-driven and distribution-free estimation of tailed-related risks for GARCH models using composite asymmetric least squares regression. Journal of Computational and Applied Mathematics, 403(??):??, March 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721004854.

Zhou:2023:NSR

[ZZWX23] Jing Zhou, Dongmei Zhang, Lin Wang, and Zhijun Xu.

A new SOCP relaxation of nonconvex quadratic programming problems with a few negative eigenvalues. Journal of Computational and Applied Mathematics, 423(??):??, May 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722005428.

Zhu:2022:VVP

[ZZZ^{+}22] Aiqing Zhu, Beibei Zhu, Jiawei Zhang, Yifa Tang, and Jian Liu. VPNets: Volumepreserving neural networks for learning source-free dynamics. Journal of Computational and Applied Mathematics, 416(??):??, December 15, 2022. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042722002576.

Zhang:2021:AMC

[ZZZL21] Xi Zhang, Wenling Zhao, Guanglu Zhou, and Wanquan Liu. An accelerated monotonic convergent algorithm for a class of non-Lipschitzian NCP(F) involving an M-matrix. Journal of Computational and Applied Mathematics, 397(??):??, December 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0377042721002466.